A Complete Bibliography of Publications in Journal of Computational Physics: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

30 November 2023 Version 1.10

Title word cross-reference

+ [MC07b]. 1 [BBDE05, BWLM09, DH07, Hor02, JBHK08, LC06a, LLIK01b, LW07, Mac07, RMF08, VBL07, vDZ06]. 1 + 1 [VWW04]. 1/n [Boy09]. 1/ν [SKK^{+}08]. 1/r [HB05a]. 1000 [PWS^{+}02]. 10^{7} [MT03]. 13 [GE07]. 2 [AV05, AMXL09, BMN07, BCE^{+}09, BRC^{+}09, BBvdV06, Ber06b, BMQS02, CD03, CJSS08, CGMS06, CdHST08, CBKM00a, CBKM00b, CK03, CL00b, DS06b, DH07, Eld07, ES06, EKP06, GS05a, GS06a, GH03, GHB03, GKL03, Gui03, GLLX08, HWL08, HH01, HS08c, JHSZ07, KB04, KPP09, LGP09, LDN04, LS00, LC03, MP07a, Meh04, MAL09, OJW06, PAD07, PCP08, PVPS09, Rom02, Rom07, RFVP09, RW03, SROCdPFF05, SA09, TS02, VGS04, VL07, VGZB09, WM09, WFC09, WZ09, XMT06, YM07, YJF^{+}06, ZZ01, Zha02, ZK05, ZJW06, ZJWC08]. 2 + 1 [Wan04a]. 2.5 [CTT08]. 25 [UL06]. 3 [AFGM07, Alb00, ASQR06, AMXL09, ALGM01, AQ00, BM02, BO05, BM06, BWLM09, BPO07, BPL06, BGS08, CD04, CWJ07, CPG04, CGN^{+}07, CBB01, CLL^{+}07b, CJ07, CL00b, CP04c, Dar02, DBF08, DGP00, DB04, EHST03, EKBL09, EAY01, GH03, GS07, GW02, GSB03, GS05c, HAAO00, Her09, HSL08, JHSZ07, KKCF09, LS02b, LG05, LJ09a, LH08a, LF04, LL08b, LZH^{+}06, MKM99, MKM04, MKLU05, MG07a, MSYL00, MP01b, NCW^{+}09, OTCM08, ORM06, ON08, PSC04, PVPS09, QS07, RWWS07, RW08, RKE^{+}07, RVVL09, SWG08, SMP01, SL07a, SPLM09, Stu01, SP00, TJS03, TPR05, VGBZ09, WGSL06, YBZ06, YJF^{+}06, ZYKW01, ZTPM05, dHRvdB07, dSMN^{+}04, vZdBB07]. 4 [GBB^{+}06]. 500 [PWS^{+}02]. 96 [FVE04]. = 300 [PWS^{+}02]. ^{+} [CL07a, CL08b, GIA^{+}08, Lio06]. ^{−} [Saf00]. α [APR09, AL01, HHPW08, LNXNTX09, PHW08]. B [Saf00]. β [CK03]. C^{0} [LL06a, LLZ07]. C^{∞} [GC02a]. d^{2}/dx^{2} [Lab09]. δf [CP03c, CP07, HM09, IITV07]. f [HJKO08, IITV07]. H [CD00, BRP05, KvRvdVvdV07, RFFP06, WM09]. K [Gos02, MG05a, ML06b]. M [Edw00]. M^{3} [LMS08]. R^{3} [BGN08, VZSL07]. μ [Mac03]. N [BADG00, BDS07, DHM03, FT09, MG05a, MC07b, PGB05, PRL03, TWYC06, WPM^{+}02b]. ∇ · B = 0 [Tót00]. ν [Mac01, SH07b]. O(1) [Pau07]. O(N) [Deh02, HBHS09, YBS06, PO01]. O(n^{2}) [Gon07]. ω [AQV02, BMQS02, ML06b]. P [QS01, FOLD05, LBL06b, VBL07, WM09]. P_{N} [FKLY07, LTK^{+}02, MELD08, Ols09, WK01b, MHB08]. P_{N−2} [WK01b]. P_{N}P_{M} [DZ09b]. ±2 [WJV07]. ψ [AQV02, BMQS02]. q [CL06a]. Q^{2} [KN04]. r [CHR01, CJ04, CK07, WLT08]. R^{−ν} [SH07b]. S^{1} [COV04]. S^{2} [COV04]. S^{3} [Bey09]. SD̄ [SB06a]. z [CK07, GGRS08].

-Adaptive [CHR01, CJ04, RFFP06, WLT08]. -Body [BADG00, WPM^{+}02b, BDS07, DHM03, MG05a, MC07b, PGB05, PRL03, TWYC06]. -Branch [Gos02]. -D [Alb00, ASQR06, AMXL09, ALGM01, BO05, CJSS08, CPG04, CL00b, Dar02, EKP06, GKL03, Hor02, JHSZ07, KB04, KPP09, Mac07, MP07a, MSYL00, MP01b, PAD07, RW08, Stu01, TS02]. -diagonal [UL06]. -DSMC [Mac01, Mac03]. -Interpolation [GC02a]. -Matrix [Edw00]. -means [MG05a]. -Multigrid [FOLD05, KvRvdVvdV07, LBL06b, VBL07]. -Polarization [CD00]. -sphere [BCE^{+}09]. -stretching [GGRS08]. -T [AMXL09, JHSZ07]. -topology [Bey09]. -up [CL08b, CL07a, Lio06]. -Weighted [AL01].

/Lagrange [WZ07, Boy06]. /Lagrange-distributed [WZ07, Boy06].

1121 [Aza09]. 13-moment-equations [TS08]. 138 [DKX01]. 151 [LLIK01a]. 152 [MKM04]. 172 [MPC02]. 173 [LM03a]. 177 [Lau06]. 185 [DD03a, SCC^{+}03a]. 196 [HLWW06]. 199 [MN17]. 1D [LLIK01a].

2 [Tol08]. 200 [Tol02a]. 207 [NTYT02]. 217 [Mil07]. 218 [JJGL07]. 225 [CL08b]. 227 [HMS08b, SM09b, dTWD09]. 228 [ABRR09b, HY11, WZL09b].

3-dimensional [SP06a]. 3624 [SM09b]. 3852 [Har04].

4 [Ano04-27, TT06c]. 406 [PW01].

70 [HT00b].

Abe [WLC^{+}08]. absolute [Sus06]. absorbers [OK07b]. Absorbing [ABK09, CL00b, HLL08, NB04, Rah04, Vay00, AMR06, AK06b, AC09, Bér07, BHNPR07, GT09c, HMOG08, HK04a, HZ08, Hu05, MTH08, XHW07]. Absorption [Vay02, CFGK05]. Abstracts [Ano00q, Ano00r, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano02r]. accelerate [VTW^{+}07]. Accelerated [SH07b, BMN07, CL08d, CS08a, JH08, LCM07, PVPS09, Sam09, VS07, YWC07, ZD05]. Accelerating [BCK09, Bow01, HJM06, PS02, PFB01, Saf00]. Acceleration [Boy09, CRB^{+}08, CAL00, DIV00, FH00a, Gen01, GTD^{+}02, ISNY05, VCM00, WFTS05, CV06, FWP09, HDR^{+}06, MN09b, MYW07, PPB09, STR07b]. accelerator [ASQR06]. Accelerators [QRHD00, ZW05]. Accuracy [BS00a, BL01, CP06a, GZ01, HT00a, HT00b, IA06a, JMP02, KPB08, LP00, MOG09, MP01b, RRV01, ST01, TRL01, XSL09, YC02, ZSP02, AHNS09, AG09, AT08, BRDM09, BBC^{+}06, CS08c, DSM09b, DWC^{+}09, DET08, FDD07, FK09b, GS03a, GGF03, GP05, Hel09a, HAP06, KMS03, LL09, LH05a, Lar09, LMN^{+}09, LX09, NPPN06, OS04, RB09b, RSO04, SN06, THS07, VL07, YMWM06, YS07c, ZJS08]. Accurate [AK05, AQV02, BL08, BR09b, BCM01, BHP07, Bus00, CL02, DDF01, EGHE06, EAY01, FP02, FPT05, FF03, GCGE03, GFCK02, JL04b, KC00, KKR01a, KKR01b, KK05c, KK05d, Kul01, OLA08, Oh04, OGV02, PKP01, RGK07, SCT06, SAKDJ05, hRT02, VQSZ02, YP01, YL01, AM03, AM04, BW06, BCL06, BY07, BC05, Bea08, BSLN09, BGLN05, CL07a, CL08b, CQRW05, DPRN05, De 04, DMP08, DDS09, EG08, FKV08, FB08, GF05a, GHMP07, Gri09, HHC08, HF08b, KE06, KT06, KK05b, KFV07, KB06, KQW03b, LP06a, LRZ04, MvW08, MSS08, MHI08, MY06b, MG07a, MOG09, Mig07, MG06, MG07d, MG08, MT07b, NOG08b, OK07a, PP04, Pop09, PMP08, RVDM09, RDPN07, SGFL09, SM04, SPM03, SYC09, TMS06, TM07, VQLZ04, WM07, ZVQ07]. Accurately [KG09]. ACE [VS07]. Achieving [HAP05]. ACM [SZH07]. Acoustic [DFT01, ES03b, Man02, ME09, PL01, SFW00, VCP00, WPM02a, ACR08, BHNPR07, CDR09, FNS07, GB08a, HSQ03, IQT08, KP08, KSGF09, LS09, LP04b, MDR07, SDD07, SG03a, TJ09, Tsy03]. acoustics [CFR04, Cap05, Cap06]. across [KT06, TLL^{+}08, Wu01]. Action [SZ01, MGS09]. Actions [BLW01]. Active [HGM^{+}00, XCY06]. actuated [KB08]. actuators [LTL^{+}09]. Adaptation [DIV00, HLS02a, Hua01b, VD00, VD02, Yam01, AFGM07, BFC04a, BFC04b, DCF^{+}08, Dwi08, FL03, HZ07a, HS03b, JHSZ07, Lap04, LC06a, VD03]. adaptative [BdCB09]. Adapted [PW07]. Adaptive [AS09, ABGV02, AZC05, Bal01, BMRS01, BDR^{+}04, BL01, BK07, BJP04, CHR01, CH01, CJ04, CBL01, CFJ09, DS05a, DV02, DPR00, DGH02, Dys01, ELW04, GTD00, Ham07, HH02a, JLT06, KS08a, KKP02, LLH02, LZT09, LLdlP^{+}00, LK01, MR00, MKOW04, MC00a, MK04b, NDT06, OV00, PL04, PC06b, SMO00, Sun00, WSTW09, WGCR07, WB09b, XHW07, YC02, ZT07a, ZLAC05, ZQ09, dSAK00, AHF04, AA09, AKV06, ASPB03, AEP04, Aza07, Aza09, BGR08, BLG^{+}08, BLM04, BL05, BEG03, COQ06, CR07, CL06b, CL03a, CC07, CXZ09, CMG09, CHL06b, CY05, CFP08, DAJ07, DW09, DLD^{+}06, DLT09, DS06a, DGRS08, FYH^{+}06, FD07, FK09b, FM06, Gre04, GHMP07, HT07, HPD09, HS03a, HS06, HS08a, HAD06, HG03, JX06, KI05, KAS06, KAA^{+}07, KPP07, KLP^{+}09, LG09, LFSS07, LHR^{+}07]. adaptive [LJSM08, LP04a, LL04b, LR07, LP09, MK08a, MZ09, MCG08, MU09, MK08b, Men04, MG07b, MG06, MG07d, MC07b, MHE06, MSB07b, NT07, OK06c, PSCB08, PDHP07, PCP08, Pop03, Pop09, PRL03, RGK07, RJM07, RFFP06, RSTB03, Ryc05, SS07b, SPGR06, ST03b, SRX07, TLK07, TTZ03, TSB03, TFD06, TDGP06, TMG08, TK04, VP09a, VK05a, WK05, WW04, WT07b, WTL08, WLT08, WM09, WHS08, WKL07, XLM07, YMT^{+}04, YJL^{+}06, YBZ04, YF09, YT07, YZF^{+}06, ZK06, ZSC08, vdV08]. adaptivity [CMR08, Ngu07]. addition [BO09]. Additive [ODCK07, HC05, SRM09, XYK05]. additive-correction [XYK05]. ADER [BRDM09, KI05, SDM04, Tak06, TT05a, TT05b, TT06a, VTT08]. adhesion [ZDD09]. ADI [KZ04, NN09, SCC09, Wel07, You06]. Adiabatic [AKV00, RV09, BG07, BZ04]. Adjacent [Azm02]. Adjacent-Cell [Azm02]. Adjoint [CKvT07, CSV00, DCS00, IFZ01, LP04a, NA08, PG04, PL08, Pro08, SDCC05, UH01, VD00, WM09, DL03a, FLE03, HPS06a, MAN^{+}06, PS03a, PBH04, WGS^{+}08, YA05]. Adjoint-based [CKvT07, PL08, Pro08, WM09]. Adjoints [TH01, SZ08]. adjustable [HKM07]. Adjusting [KG09]. Admitting [BH09]. advance [VTC^{+}07]. Advanced [TBT^{+}09, CSC^{+}08]. advancement [PMP08]. Advances [CP04c, GR01]. Advancing [Set01, JP03]. Advecting [RMB07]. Advection [CL00a, DPCV02, DB00, FMO00, HFO01, HF00, MBP00, MHS02, OGV02, RS09a, TH01, TS02, XY01, AMSZ03, AMS03, AMSZ07, Bal08, BD08, CMSZ09, DPRN06, ELW04, LLTA07, LHGF04, MP08, NZ05, Pud06, RBS06, SCT09, SD06, Sou09, VSH04, XP04a, YA05]. Advection-Diffusion [CL00a, HFO01, MHS02, OGV02, DPRN06, NZ05, SCT09, YA05]. Advection-diffusion-reaction [RS09a, VSH04]. Advection-Dispersion [MBP00, LLTA07]. advection-reaction [ELW04]. advective [Löh04, PC08]. advective-dispersive [PC08]. AEGIS [ZK06]. Aequorea [SM09a]. aerial [HSW07]. aero [MWM08]. aero-optical [MWM08]. Aeroacoustic [Fre00, HRV08, KLK08, PGN08, SMS04]. Aeroacoustics [CL01b, SHA08, SSD00, AHNS09, Cap06, Kim07, MRS09, SM06a, SFMP06]. Aerodynamic [HSBG05, IFZ01, CKvT07, EHD08]. aerodynamics [Liu09a]. aeroelastic [WB09b]. aerosol [Mar09]. aerosols [DSJ03]. Affordable [IR09]. Aggregation [LBD02, BPO07, Lau06, LL06b]. agreement [OB06].

Aided [VP00]. Air [FCB02, SD00, SMO00, WB01, CSC^{+}08, CN08, NJLA06, SDCC05]. Air-Quality [SMO00]. Airflow [ZK04]. airway [ZK04]. al [SM09b]. Alamos [Har04]. ALE [AK06a, CYS06, CHCOB09, ÉGP09, FGG01, GGF03, JC02, LBL04, MY06a, MSB07b, NJX09, TT06c, VGS04]. ALE-type [NJX09]. AlGaNd [GS05a]. AlGaNd/GaN [GS05a]. Algebraic [GSV06, HH02b, HMMO05, AHPT07, HJM07, LSS^{+}09, Rah04, VSG05]. algebraically [Boy09]. algebraically-converging [Boy09]. Algorithm [AGT02, AGT05, Asl01, BFG07, BM01a, BLW01, BD01, BZB00, BK01, CS01a, CRB00, CBKM00a, CBKM00b, CYKC01, CL00b, CBL01, DKX00, DKX01, Deh02, DB00, FGOV00, FV01, GTD00, GK02, HF00, HCG01, HLKS00, JPMC01, Lar09, LBD02, LTZ02, MK02a, MD01, Mu02, Noe00, PR00, PWW00, PM00, PA00, SR00a, Shy01, SMO00, TSB01, TK02, VG01, VCP00, ZD00, ZD08, Abr07, Abr09, AA06, AL06, AMS04, BHL07, BMN07, BIVC07, BB09b, BOK^{+}06, BP07, CXB08, CGP05, CRAG07, CPKW09, CW08, CK08, CZ09, CLS09a, CJK^{+}03, CK07, Den07, Dim07, DS06a, DDS09, DTSC04, DLMK04, EKP06, EKP07, FT06, FHD^{+}09, FD09b, FBHV05, FCD^{+}06, GTRB09, GG04, GH08a, Gib04, GvH06, GSB03, GKD09, GG09b, GV06, GGCC09, HS07a, HNF07, HM05, HC05, KD09, KKS05, KZWY09]. algorithm [KGJ05, KF06, KAA^{+}07, KW03, KT07, LLY05, Lau06, LPK05, LLC06, LJSM08, LZ04, LAKD08, LJ09a, LOK05, LLGL07, LP09, MZ09, MPD03, MM03, MCGV04, MC03, Moo03, MDS03, MK07, NOG08b, NVD05, NVD07, NMG09, Nic09, NCW^{+}09, NFA03, OLLL03, OMK09, PMP08, Pro03, PRL03, RVDM09, RJ06, RBL04, RSTB03, SDGX07, SLF08, SLG^{+}03, SL07b, SA06, SMP09, SS06b, Shy04, SC09a, Sti05, SK04b, SRX07, TSB03, TC09b, VPMC04, VS09, VSV03, VBL07, Wal03, WGSL06, WB09a, WC07, XDB09, XCRX08, YMW06, YBS06, YBZ04, Yin06, Yok07, ZH04, ZSC08, ZD05, ZGSD06, vdV08]. Algorithmic [DTS05a]. Algorithms [BSJ01, BT02, Boy02a, DDF01, FKV08, GTD01, KKP02, KFV07, KYK07, Lem00, MLS01, MPC01, MPC02, OV00, PM02, RS02, RRV01, SSW01, SK06, AvdB04, ASPB03, AST09, BCK09, BDGL05, Boy04, BUEG06, CHL06a, CFS09, CB03, CC03, DVHM05, DHM03, DEHL06, ELVE07, Fas03, FD03, HJFW04, KK03b, KKS07, Kuz09, LMX^{+}08, LK09, LSW06, LCCG05, MD04, MC07a, Nit05, PP04, PVR07, Ram03, RVM07, RA09, SS09b, SMB09, SCC09, Sto07, Tyg08, Wag05, FS00a]. Alias [Pop00]. aliasing [KG08, KK03b, SVB09]. Aligned [KKR01b, GYKL05, NLLE06]. Aligning [HLKS00]. alignment [BGM08]. all-electron [HBHS09]. all-scale [PS03b]. All-Speed [sLwG08, BN09]. Alleviation [Lee09]. Alloy [SZ01, GJK09, LW06, RJM07, TZ07c, ZVHP03]. alloys [JVVS07, KGJ05, TZ07a]. allspeed [GBC06]. Almost [Tan05a]. along [JJGL06, JJGL07]. alpha [LNXNTX09]. Alternative [KJ01, Boy04, HLMM07, MKKY06, Pro08]. aluminum [MV06]. Ampère [HF01]. amplification [BCG09]. Amplified [Li01]. Amplitude [VCP00, CF06a, CSMH05, KTD03, MV08, PPCW06, TMND07]. AMR [SO08, YF09, vdHK07]. analyses [SCT09, YKK08]. Analysis [AA02, BZ04, Bod06, BC02b, BE02, CFS09, CGP02, DMG00, DE06, DCV^{+}01, ELW01, FLG01, GGL^{+}01, HA02, JM00, KMHR00, LMK03, LS02b, Mac00, MG02, PD01, PM07, Pon09, Sai02, SHWW00, SGD03, SCD00, Spo00, UH01, VCP00, Wel07, WK01b, Woo06, YXU01, YMWM06, YS07c, AA09, AJT04, AL06, AZ05, Bal08, BBD04, BV05, BDGL05, BCR04, BS09a, BCM^{+}07, BDCG03, BHP07, CLB08, CLLG09, CRAG07, CDI09, CN05, CGH05, CJ04, CFJ09, DVHM05, DL04, DWLM09, DMG04, FWK08, GB08a, GGF03, GD06a, Ham07, HNGB04, HT03, IM05, IA06a, JKL05, KK05b, KRT^{+}09, KLM07, KKS07, Kry04, LNGK04, LTZ03, LP04a, sLwG08, LY04, LJ07, LRS09, LMK09, MK08a, NU09, OK07b, PKKL05, PS07a, PIN09, Pir07, PA07a, PBH04, RVVL09, RM07, SBA07, SDCC05, SDS07, SVB09, SRNV07, SL07c]. analysis [SM06b, TX06, VCT07, VGPL09, VBJ08a, VK04, WGNT06, XS07, ZGSD06, dHRvdB07, dNWvSD07, dTWD09]. Analytic [AI09, YJ06, YMF01, BB08a]. Analytical [BEE06, LH08a, NFvS^{+}06, Ren07, SZ00, Chu09, JR03, LHD05, LT09a, NDG05, SB06a]. Analyzing [LWG03]. anchored [NT07]. anelastic [BM06, CPG04, GBC06, PS03b]. aneurysms [YXLF05]. Angle [FSY00, LWDA09]. angles [AZB09]. Angular [Car01]. animal [HSW07]. Anisotropic [BFC04a, FL03, HLS02a, JW02, LMSV00, QS01, VD03, BHR03, BFJ03, BFC04b, Bur05, BHSV07, CP05, Chr04, DT03, Hua05, JC06a, KE09, KSS09, Lar07, LLY05, MP07a, MD06, MSB07b, RBH03, SKS08, SH07c, WC07, WKL07]. Anisotropy [EV03]. Annealing [FH02, PA00, Pav07]. Announcement [Bra01, Ano05s]. Anomalies [KS02a]. antenna [LVL05]. Antennas [VR02]. Anti [XS05c]. Anti-diffusive [XS05c]. Antiplane [LAS01]. apparatus [JR03]. Appear [Ano00q, Ano00r, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano02r, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano02f]. applicability [LTZ03]. Application [AFGM07, AKH06, ADK00, ADK02, BS01, BF07, CWT00, CA06, Che00a, Chr03, DLS^{+}00, FM06, GV08, GSD01, GPH^{+}01, HCG01, IYI^{+}02, JC02, KCGH07, LX00, LS05a, MPP01, ML06a, MZ07, MSP^{+}06, ML06b, NCW^{+}09, NGC^{+}07, PSC^{+}06, PWS^{+}02, QL01, RAD07, RXH02, RFVP09, Set01, SHWC07, SSC00, UBRT07, VOD08, VVS08, Vay01, VD00, VD02, Xu01a, XHW07, ZWS07, ZLAC05, ZRR00, AC09, AD04, BBD04, BZ08, BB07b, Boy03, BRB03, CLS^{+}06, CP03a, CMSZ09, DDK06, DHM03, DBB06, Dwi08, DBS06, FDD09a, GLMH09, GS03b, Hua07, KP07, KP08, LHZW05, LLL07, LMZ^{+}08, ML05, MY03, Mil04, MG07c, NS05, PPDM08, PSC04, QS04, SR09a, SSE03, SW04b, SZC09, SLC07, SNLS03, Sme06, SRX07, Tsy04, VBL07, VB08, VD03, Wea09, WJV07, WH05, XSG08, YZLH09, YXLF05, YE05, dSHHM05]. Application [VBJ08b]. Applications [BS04d, BK01, Che00b, DC01, DR09a, LTK^{+}02, LWW04, BS05, CP06b, CP06c, DIL03, DDFT09, FVE04, FD03, FWK08, GF05a, Hoh06, Jar04, KK03b, KK04, Lee03, LG05, LXM09, LSW06, MESV09, MG07b, ON08, PR04a, RBSL06, SS08, SK08b, SK06, WSYS09, WS09, YLD09, ZZ07, ZXQX08, vdHK07, DTS05b]. Applied [BS00e, DCS00, DSS00, EAY01, PG02b, RH01b, VR02, VSMW01, BB04a, BRC^{+}09, CHG^{+}07, CMR08, DB04, FS04, HS09a, Jar04, KKO04, LC03, Mac07, MWM03, Mad06, MM03, Ram06, RVVL09, SB06b, SDD07, YSS05, ZZVM08, vDZ06]. applying [AZB09]. Approach [AS02, BL01, Bor00, BCMO01, CBMO02, DSS00, DI02, FH00b, FCB02, GPH^{+}01, LS02a, LP02, PSN00, SSC00, SSD00, AA07, AA09, AK07, AG09, AHMS03, Asl04a, AGW07, Bae03, BL09b, BF08, BB09a, BL09c, BCM^{+}07, BEA09, BHSV07, BB09c, CORT09, CGL06, CM06, CCJ07, CL07a, CL08b, CSMH05, CQRW05, CHCOB09, CVE06, DSJ03, DP09, DLW04, DST07b, FHW07, FLB03, FHLK05, FD09b, GJK09, HZ07a, HAS05, HA09, IQ08, Jia07, JZ08, JG09, JCT07, Kas07, KR09b, Lap03, LTZ03, LL04b, LDW07, LL06a, MR06a, MGCR07, Meh04, MWG^{+}06, MV06, NJLA06, NM06, Nat06, NVD07, Nis07, OF06, OVG07, Pal08, PPDM08, PGB05, PSD09, PP09, Pee03, PSC^{+}06, PCS^{+}09, PA05, RR07, RSS09, RF06, Ryc05, SB06a, Spe05, SG03b, SBC04, TC07a, TZ07c, Tan08, TGB^{+}07, TdAAP08, UYK^{+}04]. approach [VC03, VZSL07, WD07, XSG08, XSL09, YAvdB^{+}08, ZG08, ZL04, Zhe07]. Approaches [KLvBvL02, KMSH08, KK04, LTD^{+}06, MLSD07, dWKL07]. approximants [CDI09]. Approximate [LP01, SWTM01, SL03, BDRT09, Gui05, GF05b, HBLD07, MK05, RBT03, YLD09]. approximating [Boy06, Her09, Tow09b, WZ03]. Approximation [BISS01, FSY00, Her00, LLTA07, SKAS01, ACR08, BGN08, BBW06, BLL03, BS06a, CCV03, CWYM08, CHCOB09, CF04, DP08, DI09, FDK06, FKLY07, GH03, GM04, Gos04, GM06, GV06, HLO08, Kas07, KYLB07, Ngu07, PAD07, PP09, ST06, SK05, Sme06, SKW05, TMS06, Tsu06, VGCN05, WZ07, WO05, WO09, YS06, dFJS09]. Approximations [BJM02, BSJ01, Dur00, LTK^{+}02, MG02, RV00, SFY01, Boy05b, CHH06, DC07, Gro06, Gro07, KCMM03, LCB04, LX07a, Lur07, MN04, MN17, RB06, RM07, SVH^{+}06, Sou09, SN06, Tol07, Tol08, TE04, Tow08, TFDK04, XDC09, ZC09]. aquatic [HSW07]. aquifers [WGNT06]. Arakawa [DF07]. arbitrarily [BCDW06, EG08, KC06, LL07, TRSK09]. Arbitrary [AMH04, DK07, Goe00, GBS00, HPZ01, HJM07, KMS02, LY04, AEP04, BAYZ08, BR09b, BTW03, CDDL09, Cec05, COQ06, GT09b, Gel06, GF05a, HLO08, KZWY09, KK05a, KZ06, LSB04, LY07a, LHZW05, LKMU05, LS05b, LBL08, Min03, NMH^{+}07, Nik06, NPPN06, PS07a, RWWS07, SM09a, SB09, Tol07, Tol08, XDC09, YCL05, YFBH07]. arbitrary-Lagrangian [LS05b]. arbitrary-order [Tol07, Tol08]. Arbitrary-pressure [AMH04]. arc [PL09a]. arc-like [PL09a]. Architectures [BLW01]. Arctic [MR01]. area [AMSZ03, AMS04, JCT07, ZL08a]. area-preserving [AMSZ03, AMS04]. areas [GPC07]. arguments [De 04]. arising [BST03, BO04, CFS09, HJ09, TDV06]. Arnol’d [SP07]. array [CWJ07]. arrays [LVL05, LJ07]. art [VTM^{+}08]. arterial [XS07]. artery [YXLF05]. article [LM03a, MCP03, TR07]. Artificial [CS01a, FT01, HDC02, Kel05, PFSL07, RTT01, SD05a, SD05b, STiST02, Tsy03, Xu01c, BCDR06, BL09c, CL06a, EZ08b, FL07, KL08, KR09b, MLM09, Owe04, RBH03, zSW06, zS06]. artificially [ST03a]. Aspect [AJG01, Car01, BJP04]. Aspect-Ratio [Car01]. Aspects [Hua01a, HMK02, SW08c, Ano04z, MR03, Sus06, VCG03, dWKL07]. assembly [JRS05, Moo03]. Assessment [Mav02, Ano08-50, Lar09, Maz06, MGS07, SM09b]. Assimilation [DCS00, LCS02, KFH^{+}04]. Assisted [BMS00, SKXK05]. Associated [SS01a, FL06, GD06a]. assumption [CGP05]. astrophysical [BvdHKG07, KSW07, RFFP06]. Asymmetric [Vay02]. asymmetrical [FG05]. Asymptotic [BD06, CLB08, CGL08, CS04, DDSV09, DL04, GK01, JKL05, MC09, BLW04, BNNP06, CDV07, KYLB07, sLwG08, ML08]. asymptotic-preserving [ML08]. asymptotically [JAK05]. asymptotics [BLM08]. asynchronous [FDL08, KDOO05, SA06, UBRT07]. atmosphere [LHR^{+}07, SW08c]. Atmospheres [DKSW01]. Atmospheric [Bon00, GR08, RWMK03, SZ08, SMT^{+}08, SK08b, TW05, TR07]. Atom [CP00, CWWZ00, VCTS02, BRB03, BBK06]. Atomic [AC01, LD09b, SG03b, TLAD04]. Atomistic [EH02, FK09a, KZ06, LCNR07, MKL06, Ren07, WL03, WWK05]. Atomistic-Continuum [EH02, KZ06, Ren07, WWK05]. atomistic-mesoscopic-continuum [FK09a]. atomization [DMP08]. Atoms [VDM^{+}02]. attached [TLAD04]. attachment [BHL^{+}04, Lap03]. attachment-detachment [BHL^{+}04]. attitude [San03]. Attraction [Saf00]. Augmented [Geo08, HB02, IQ08, ILL09]. AUSM [CL08b, CL07a, Lio06]. AUSMPW [KKR01a]. Author [Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano00z, Ano00-27, Ano00-28, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano07-33, Ano07-34]. Authors [Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano04q, Ano04s, Ano04u, Ano04v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano07z, Ano04r, Ano04t, Ano04w, Ano04x, Ano04y, Ano05t, Ano05u, Ano05v]. Auto [VG01]. Auto-ignition [VG01]. automated [KAS08]. Automatic [CSV00, GT05]. Automatically [HvHHS05]. Automaton [LMSV00, LGN05]. autonomous [BM01c]. Autophobic [HLZ02]. auxiliary [Bae03, KKM08, MK06]. Avalanches [TNGH02, FNBB^{+}08]. average [KD09, KLLJ09, Yus06]. averages [ZIP06]. Averaging [LR01a, PS07c]. avoid [ZSW07]. avoiding [CSO09, KSGF09]. aware [ML06a]. Axis [CL02]. Axisymmetric [BBG^{+}02, Car02, CS00, GP00a, LG03a, Lem00, Mie00, Nie01, RH01b, SP00, TCM^{+}00, AK09, ACLS03, BZ08, FJ09, FBHV05, GV07, GLLN09, Kel05, LN09, OCFF08, PSCQ03, SLF08, VGBZ09, ZK06].

B [CP04a, KMJ01, WdND06]. B- [WdND06]. B-spline [DD03a, KMJ01, DD03b]. B-splines [CP04a]. B. [Aza09]. Back [DL03b]. background [LGKP07]. Backscattering [FT01, GS09a, HS07b]. Backtracking [TWS02]. backward [RFVP09]. backward-facing [RFVP09]. Baer [AW04]. bag [BB09a]. Bake [Li01]. Balance [FR02, CT08a, CPKW09, DET08, EHD08, LTD^{+}06, MKOW04, VCZS04]. Balanced [CZVS04, LR01b, Xu02b, AB05b, BES07, BKLL04, CVB06, FCD^{+}06, GPC07, Her08, KCMM03, LMNK07, Mou04, NPPN06, NXS07, RF06, WSYS09, XS06]. balanced-force [FCD^{+}06]. Balancing [DPR00, HGN00, MG05a, RBT03]. ballistic [BMN07]. Ballooning [CGC^{+}09, SHWW00]. bamboo [AINR03]. bamboo-type [AINR03]. Band [CD00, DGP00, DBB06, Dur08, KGJ05, LW09, SP05c, VTW^{+}07, WHLL03]. Bands [DFT01, BZ04]. bandwidth [MTWW06]. bandwidth-optimized [MTWW06]. bare [AINR03]. baroclinic [AL08]. barotropic [LHR^{+}07, Shy04, vBK03]. barriers [JN07]. Based [AV02, BSJ01, BW02, BMRS01, BM01b, BT02, GTD^{+}02, HMS08b, HK02, IFZ01, Jan00, JL02, KMS02, LC01, LTZ01, LLdlP^{+}00, MD01, MKR00, MOS^{+}00, Noe00, OB02, OCK^{+}02, PX02, PR01b, PM00, QV01, RS02, Rom02, SS02, SC01, Sum00, TK00, Tol02a, Tol02b, WW00, WS01, ZTZ02, AvdB04, ACGV07, AHPT07, AL06, BNV08, BAMD07, BBB08, BLG^{+}08, BdCB09, BG05a, CXB08, CKvT07, CR05, CR07, CdHST08, CS05, CP06b, CP06c, CLL^{+}07b, CBH03, CRB^{+}08, CQRW05, CGM07, CLS09b, De 04, DCF^{+}08, DFV08, DLD^{+}06, DLD08, DS05b, DDS09, Eld08a, ES03b, FSS03, Fox08, Fox09, FMR09, GGMN^{+}09, GCNB07, GKD09, GW05, Gir06, GG09b, Gra06a, GS09c, HMM08, HW08, HMS08a, HV03, HKO07, HLL08, IKS^{+}09, JD09, JRS05, JC06a, JL04b, KOQ08, KZWY09, KGJ05, KFV^{+}05, KSJ03, KNH05]. based [KL06, KS08b, KP08, KT04, KT07, Lap04, Lap03, LSA06, LWG03, LJSM08, LB03a, LZ09b, LZ09a, LTZ03, LX09, LQX06, LW07, LW09, LHGF04, LBL07, LBL08, MLSD07, kM07a, Mil06, Mil07, MT07b, MDS03, MK03, MHdB07, MO06, Neo07, NJX08a, NCS03, NFA03, NDT06, OCFF08, ODCK07, PDHP07, PNMK09, Pop03, PL08, Pro08, QT08, QL04, QKS06, RWMK03, RKE^{+}07, RK07, Ros03, SBA07, SC08a, SKWN03, Ser09, SPM03, SD05a, SD05b, SO08, SAM05, Shy06, SHP07, SS05b, SHPC09, SZH07, TBT^{+}09, Ten03, Tol07, Tol08, TU04, TCO^{+}04, Tsy04, TY07, Tuc03, VTC^{+}07, WW04, WM09, WSTW09, WFC09, WS09, XCY06, XSL09, YF09, YS06, ZR08, ZSC08, ZSP08, ZHSS09, dSM05]. Bases [ABGV02, ZSC07].

Basic [BK01, FHLO08, IX09, LVW06a, Wan02, AL06, LKNG01, Wan02]. Basis [AP02, GTD01, HL06a, OMK09, PB00, SC01, BZ08, CQO04, DR09a, FW07, FP08b, GH08a, GS06b, GG09b, HBHS09, LCW04, LJW07, LBL08, MT07b, NG06a, Ngu07, Ngu08, RA09, SC08a, SVH^{+}06, TB09, TW03, UYK^{+}04, Wag05, WFTS05, WF06, Yin06]. Baths [CS01b]. bathymetry [Che04]. Bayesian [Kou09, MNR07, MN09b]. bead [LWF^{+}08]. bead-spring [LWF^{+}08]. beads [KM08b]. Beam [QRHD00, BBF^{+}08, DZ09a, DDGS09, FCJ08a, GGRS08, QFR04, SHY07, TET09]. beam-beam [QFR04]. beams [HDBW05, KM03, LQ09]. Bearing [WB01]. BECOOL [CGC^{+}09]. Bed [PCCD00, HC08, RC06, RF06, RMF08, Sar03]. beds [QP03]. behavior [AKH06, GTRB09, KMID05, SK04a]. behaviors [LJ09b]. Behaviour [WWVG00, Ain04, DVHM05, Dur08, KKO04, sLwG08]. behind [VG01]. belonging [SS09b]. Beltrami [HZ07a]. BEM [MMS04]. Bénard [CA06, TC02]. Benchmark [FS00b, DLMK04, DOW08, LW06, SL07c, VBL03]. bend [BCZ04]. bending [DLW04, DLW06]. Bessel [GST02, Nas08, Saf02]. best [Lab09]. beta [GLN06]. Bethe [Mai03, Mai04]. between [ACK02, BBHM09, DJM05, Eg07, FG07, GHV00, HDBW05, IA06b, JA08, KK09, KMSH08, KM08b, MRS09, OB06, PC02, VLW07, VZSL07, YM07, ZKDT07]. Beyond [SDS07, PKD07]. Bézier [CH08, DAJ07]. BFGS [Abr09]. BGK [CKR01, Xu01b, CKR00, CDL04, CDL05, FH00a, FH00b, GSW00, LF06, MSJ07, Mie00, PPCW06, SY08, Xu01c, XH03, XMT05]. bi [AKH06, HHM04, KH07]. bi-period [AKH06]. bi-periodic [HHM04, KH07]. bias [ME09, TG04]. biased [BBHM09, FG05, JAK05, PYC04]. bidimensional [BS06b]. bidirectional [ES03a]. Bidomain [GGMN^{+}09]. bifluid [GV08]. Bifurcation [DSS00, SML02, dNWvSD07, dTWD09]. BiGlobal [KRT^{+}09]. biharmonic [Bia03, GD06b]. bilayer [FK06]. bimodal [Wea09]. bin [WXG07]. Binary [IYI^{+}02, SZ01, FWP09, LW06, RJM07, TBT^{+}09, Wan04b, YU05a, ZVHP03]. Bingham [VBL03]. Bins [TRL01]. biological [CDDH07, JRS05, KL06, LD06, LMZ^{+}08, MWM03]. biology [MG07b, NGC^{+}07]. biomolecular [LCM07]. biomolecule [CXB08]. biorthogonal [ELW04]. Biot [BQQ09]. Bipartite [RS02]. birth [NSC09]. Bistable [SSC00]. black [FD09a, LCH03]. black-box [FD09a]. black-oil [LCH03]. blackholes [Lau04]. BLAS [CFR08]. blast [BWLM09, SL04]. blended [Ros03]. Blending [Lar09]. blends [AKH06]. Blind [CJLS09]. Blob [CM00, BB04a]. Bloch [BBR01, BS06b, Gos04, Lin01, LW09]. Block [CP06b, MC07b, PPC00, PSH^{+}08, CHB09, Cho05, EHST03, EHS^{+}08, GGMN^{+}09, NGvdWS09, PSC^{+}06, SCT06, TDGP06, TMG08, WR09, YLD09, vdHK07]. block-adaptive [TDGP06, TMG08]. block-AMR [vdHK07]. block-triangular [GGMN^{+}09]. block-tridiagonalization [WR09]. blood [CGN^{+}07, GGCC09, LL06b, XDB09, YXLF05]. blood-tissue [XDB09]. blowup [HMR08, MJ09a]. Bluff [JML^{+}01, PW00b, PW01, PWS^{+}02, KIH09]. Bluff-Body [JML^{+}01]. blunt [BP04a, BP04b, HEN09]. blunt-body [HEN09]. board [Ano03l, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano06a, VGL^{+}07, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03m, Ano03n, Ano03o, Ano03p, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08a, Ano08b, Ano08c]. Board [Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x]. Bodies [GPH^{+}01, LAS01, PW00b, PW01, Alb09, BGS08, BP08, CYS06, CC08b, EG08, FWW04, GS05c, HB05b, KM08a, KT06, KRT^{+}09, TZL05, YM07, vZS07]. Body [BADG00, Bus00, JML^{+}01, Mai01, PWS^{+}02, WPM^{+}02b, Alb08, BCDW06, BDS07, BP04a, BP04b, DHM03, Eld07, Eld08a, Fas03, HEN09, JD04, KIH09, KC06, Mai03, Mai04, MG05a, MZ07, MC07b, PGB05, PRL03, San03, SG03b, TWYC06, Vil08, YS09, Yu05b]. Boltzmann [CGMS06, AL08, ABZ^{+}08, BKS07, BYS08, BTC05, BLM08, BdLL01, CGMS03, CA06, CDL04, Del03b, DCK08, Dys01, FM04, FR03, FSM^{+}01, FH07, GS06a, GS03b, Gua00, Hag07, HDC02, HHL00, HKG08, HNGB04, HHC08, HH07c, ISNY05, IYI^{+}02, IOTK04, IF09, JP00, JKL05, KY08, KPB08, Kwo08, LL03a, LLP07, LL01a, LL03b, LL05, LLC06, LLQ^{+}02, MP01a, MP02, MRS09, MEG02, MSYL00, MHS02, Mie00, MR07c, MAL09, NCS03, PR00, PL09b, PSCQ03, PSC04, PSC^{+}06, PA07b, PPB09, RMSB09, SCT09, SB09, SPT05, SY08, SLC07, SS03b, SS05c, Sof09, Sun00, TBJ^{+}09, VLB09, WWC07, WS09, XH03, YZ07, YGL05, YF09, ZK05, ZSC07, ZXQX08, ZSC06, ZTPM05, vdSE00]. Boltzmann-like [MEG02]. Born [CXB08]. Bose [BT03, BJM03, BW06, BCL06, BS08a, CKLS05, CLS05, VCTS02]. Bose-Condensate [VCTS02]. Boson [BTFY01]. Boson-Fermion [BTFY01]. both [SHP07]. bottom [BTT08, FG07, VTT08]. Bound [Mai01, CKLS05, GG09b, KSW03, Mai03, Mai04]. Bound-State [Mai01, Mai03, Mai04]. Boundaries [CPP02, HLKS00, UMRK01, BTC05, BF08, BJ09, CBI^{+}04, GS07, GSB03, KAK03, LL03a, LKP06, MTH08, Mil08, MDB^{+}08, PC08, SS07c, TLK09, Vik03, WS04, XW06, YB06]. Boundary [AC00, ACY00, AD03, AGH02, Bra08, CM00, Cor00, DMG00, DC01, DKSW01, DKX00, DKX01, Eli02, FVOMY00, FT01, FSY00, GZ01, GM01a, GHG01, Giv01, GVT01, Gro00, HLS01, HCG01, JSCZ08, JC06a, JL02, KG09, KAIN01, KKC01, LP00, LOK01, LFK00, MCJ01, MSYL00, NFK01, NMS07, Nys02, OKL01, OB02, PPC00, Pet01, RC00, RTT01, SFY01, SW00, SS05c, TC09a, TS08, Vay00, VR02, WHV^{+}00, YFS01, ZWL02, ZJWC08, ZP02, ZRR00, ABLS05, AST07, AMR06, AWK07, AB03, ABK09, Ata04, AG08, AKP07, AMS03, BYS08, BBD04, Bér07, BL08, BBMB07, BM05, BA03, Bet08, BGS08, BO04, BT09, BP08, CD03, Car09, CGMS06, CFR09, CP04b, COER07, CGKM06, DH09, DS05b, DT03, Dim07, DND06, DCK08, EZ08b, Eg07, Eli07, ES03a, FEL^{+}05]. boundary [FM04, FE04, FF03, FGP08, GGS09, GMD07, GS05c, GN03, GP04, Gla05, GPVB07, GP05, GHMP07, GKV09, GK04, GK07, GE07, GT09c, Gui03, HS09a, HK04a, HD07, HAS05, Hel09b, HK08b, HZ07b, HS08b, HS08c, Hu05, HLL08, HSS07, HSC09, HST09, HF08b, IKL^{+}08, IK07, IG05, IDD04, JA08, JM05, KIH09, KIHM09, KY08, Kau03, Kel05, KJ09b, KC06, Kim07, KL04, KB06, LTH08, Lau04, LWP^{+}09, Lee03, LHZW05, Li08b, LN09, LP06b, LY06, Liu09b, LDV08, LZC04, LMZ^{+}08, LCCG05, Mai04, MKLU05, MvW08, MR05, MS04, MTH08, MJ07, Mil08, MDB^{+}08, NA08, NFGK07, ND04, NN04, Nic09, NCW^{+}09, NK08, NB04, OPML07, PSC^{+}06, PH08, PS08, PK05, PWM06, Pon09, PSM08, Pro05, Pro07, RMGK04, SSN09, SS03a, SAK05, SBCL06, SW04a, SLC07, SSND03, SJHM09, SM06b, Sof09]. boundary [zSW06, zS06, SK03, SN06, SCN07, SN08, TC07a, Ten03, TE08, TF03, Tsy03, Uhl05, VB09, VVS08, VGZB09, VRM07, VHI05, VP09b, VZSL07, WT07a, WK04, WL06, WFC09, WS09, XJ07, XHW07, XD07, YB06, YP06, YZLH09, YBZ06, YH07a, YZW07, YW07, YLA08, YE07, ZKDT07, ZZ07, Zhe06, ZT03, ZW06, ZZFW06, dA04, dTDI^{+}07]. Boundary-Adjusting [KG09]. Boundary-Conforming [VR02]. boundary-fitted [PS08, SS03a, YP06, ZKDT07]. Boundary-integral [AD03, JA08]. boundary-lattice [FM04, PSC^{+}06, WS09]. boundary-lattice-Boltzmann [DCK08]. boundary-layer [ZT03]. boundary-layer-resolving [NK08]. boundary-value [ABLS05, Eg07, PSM08]. boundary/level [YS09]. Bounded [Coe02, FG02, Sum00, CLB08, CP04c, FNS07, FPK08, FGP08, HPD09, IDD04, MC06a, PPDM08, SFMP06, WD07]. bounded-obstacle [FNS07]. boundedness [HR07, RH05]. bounding [PG04]. Bounds [MPP01]. Boussinesq [ES06]. Box [PS07a, FD09a]. BPM [FCJ08a]. Branch [Gos02]. Branching [KM02, LM08b]. Brazovskii [ZZ08]. breakdown [WH05]. Breaking [DF00a, SSSWD00, KDOO05, LTD07]. Breakup [CBL01, QLS09]. brick [DR06]. brick-tetrahedron [DR06]. Bridges [LS02b]. bridging [PKKL05, WL03]. briefly [BBF^{+}08]. Brinkman [LV07]. Brownian [DHM03, SP04]. Bryan [MR01]. BSOR [CKLS05]. Bubble [Han01, YSC01, BPMR08, HY09, HY11, HL07c, LF04, MGCR07, NJLA06, YFBH07, ZEA06]. bubble-stabilized [HY09, HY11]. Bubbles [Dar00a, ZYKW01, BOK^{+}06, HSL08, Sus03, WK04]. bubbling [CGL08]. Bubbly [KS02b, MTV08]. buckling [LS08]. Buffered [SSC00]. buffers [SKR06]. Building [SSW^{+}07]. Bulk [GHG01, AKH06, MLM09, VTW^{+}07]. Buoyancy [SZ01, KIH09]. buoyancy-dominated [KIH09]. Buoyant [PG02a, SWG08]. Burgers’ [BFG07, PIN09, DP00]. Burn [BSJ01]. Burnett [LR03, OX04].

C [Thu08a, TRSK09, WdND06]. C-grid [Thu08a]. C-grids [TRSK09, WdND06]. CAA [DTMS06, RBSL06]. CABARET [KG09]. cables [GPL05]. Cage [vHBB02]. Cahn [CR07, CFP08, pHL09, KW06, KKL04, WKG06, WKL07, XXS07]. Calcium [SSC00]. Calculating [BS00e, DST07a, MBM01, MN02, PSZ09, TRL01, LWW04, MS04, RMB07]. Calculation [CTS07, CSV00, Deh02, Fed02, HO03, RS02, YSO07, BCDW06, BST03, BS09b, CSMH05, ELD08b, Fou06, HBHS09, KKD08, LZ07, LLC06, LP06a, NG06a, OLLL03, SCT06, SAM05, SMAj08, VTW^{+}07, ZSTC06]. Calculations [AC01, CY00, CWWZ00, DGP00, FJ09, HSK00, HJ02, KHV01, Lou00, Mit00, TR02a, UH01, YMF01, AST09, AT05b, BCHL07, CL06a, DT04, DL03a, DBB06, FHW07, Hof04, HRV08, KKCF09, Küm04b, LCG07, LC03, LCM07, LLRP09, PSH^{+}08, SF03, SHP07, SP05c, Tol07, Tol08, VTM^{+}08, YMW06]. Calculus [BS01, OVG07, PS07a, PS07b, PCS^{+}09]. calculus-finite [OVG07]. Calibrated [CBS05]. calibration [BV05]. call [Ano05s]. Camassa [COR08]. Cancellation [Lee07b, Lee09]. Canonical [LOK01]. canopy [Dic08]. capacitance [MS04]. capillarity [TW07]. capillarity-dominant [TW07]. Capillary [Mad05, NS04, PS05]. capsule [LS08]. capsules [SCRL08]. capture [AZ05]. Captured [YC02]. Capturing [AS02, BJ02, BS00c, EFFM02, LFK00, MC02, NFK01, RMO00, SM05, STiST02, TNGH02, Tót00, BAR08, BdCB09, BW07, CB09, DLD^{+}06, Edw06, FSS03, HJJ09, KL08, KH08, kM07a, Pir06, SYC09, SAM05, TDWY08, TY07, UTBV03, Vol04b, Wen06, XD07, YJL^{+}06, dSMN^{+}04]. Carbuncle [PD01, DMG04, NK08]. carbuncle-free [NK08]. Cardiac [Ota00]. Carlo [ABRR09b, LM03a, MCP03, ABRR09a, AMH04, BBHM09, BS07, BMDS05, BSP06, BUEG06, BB09b, CLL07a, CGMS03, CGMS06, CTW^{+}08, CV06, CF06b, CS03, CS04, Dem04, DL03a, DL04, DUEB07, DDDC07, EULM03, ED07, FG04, FG05, FT09, Gen01, GL09a, GMH06, HH07c, HGM01, IH04, KB00, KMV03, KAS08, KLW09, LSL08, LM08b, LM01, LD09b, MMKP08, MU09, MBS03, NU09, OK07b, Pal08, Pet07, PK00, PVR07, PVPS09, QL01, RRV01, RS06b, SSE03, Sch08, SL04, She08, SA06, SMSS07, UH01, VK04, VK05b, Vol04a, WBM09, WGS^{+}08, WMH07, ZSB^{+}08]. carrying [CDV05]. Cartesian [AMSZ07, CL00a, Cal02, CRB00, CBGI09, Che04, CMG09, CYS06, CGKM06, DDH01, GSB03, GS05c, HLS02a, JMK01, KKCF09, KAK03, KL04, LPK05, LJK09, LKMU05, LLB05, LBL06a, MKLU05, MCJ01, MG07d, NA08, OK06a, OSK09, RCB05, RW03, SROCdPFF05, SBCL06, SH07b, SS07b, SPGR06, TU04, UMRK01, VS07, XLM07, XLP05, YU05a, YXLF05, ZT07b, dHRvdB07, vdHK07]. Cartesian/immersed [GS05c]. cascades [Ram06]. Case [FP02, HH01, PWW00, Spo00, BMN05, CD03, CC07, CY05, DBF08, Dur08, GA09, KTD03, LRS07, QS04, QLK07, SD06, VVM05, VP09b, ZQ09]. Cases [LMS02, FGS09, GR08]. CASL [MPD03]. casting [GS03b]. castings [BEA09]. category [Cap05, Cap06]. cathode [SXyWX09]. Cauchy [CFS09, SY09b]. Caustics [BS00c]. Cavitating [SS02, LKX04, SPB09, SMS08]. cavitation [SG03a]. cavitations [Hua07, WWC07]. Cavities [CL00b, AKL^{+}08, LKD04, SS07a]. Cavity [AQV02, APQ02, AK05, DR09a, GGP06, Men04, PSC04, Woo06]. CBFEM [OMK09]. CCD [SVB09]. Cell [Azm02, Bow01, CP04c, CB02, JC02, JM00, Lap02, LDL^{+}09, MD02, MC00a, Par02, PH09, QRHD00, SMP01, Sni01, SPC01, VC00, BAMD07, BMT09, BF08, BM07, CDDL09, CKPW07, CP07, CWD08, CCF^{+}05, FHD^{+}09, FD07, FG06, GS09b, GF05b, HDR^{+}06, IITV07, JH06, JD09, KW08b, LWDA09, LLL07, LL06b, Mai09b, Mai09a, MN09a, MCG08, MSB07b, NGC^{+}07, OK06b, PPCW06, QFR04, RB05, RB09b, SS09b, SK07b, SXyWX09, TF03, WGCR07, YE07, ZSW07]. Cell- [SMP01]. Cell-Centered [MC00a, BMT09, BM07, CDDL09, CCF^{+}05, GF05b, Mai09b, Mai09a, MN09a, MCG08]. Cell-Centred [JM00]. cells [CDDH07, DPRN06, LTD07, Li08a, Liu05, MV06, RCB05, XLS09a]. Cellular [LGN05, Nov04]. Center [Saf00, Saf02, HP04b]. center-difference-WENO [HP04b]. Centered [MC00a, SMP01, BMT09, BM07, CDDL09, CCF^{+}05, GF05b, Mai09b, Mai09a, MN09a, MCG08, MCGV04, PYC04]. Central [AT05a, DPRS01, KT00a, KT00b, Liu05, QS02, TA06, BTW04, BS08a, BL03, CVB06, Cap08a, CP08, CZVS04, GS03c, JR09, KK05b, KPK09, LSB04, Li08a, MGS07, SGD03, Zie04]. central-constrained [Zie04]. central-upwind [BL03]. centre [Mot08]. Centred [JM00]. centres [SPLM09]. cerebral [YXLF05]. Cerenkov [GCLB04]. CFD [AFGM07, DTMS06, KP08, LXM09, ZWL02]. CFD-based [KP08]. CFD/CAA [DTMS06]. CG [YAvdB^{+}08]. chain [GL09a]. chains [CVE06]. Chang [Del03a]. Change [JLCD01, MR00, WW00, WHV^{+}00, YSC01, BFC04b, EKP07, GCNB07]. channel [BF07, CZVS04, DS06b, HO03, PPDM08, SS05c, TS08, VTT08]. channels [CGRGV^{+}04, NFvS^{+}06, SFX03, TCM05]. Chaos [AKY01, DGF09, LX09, MN09b, PW07, RM07, WK05, XK03, HLRZ06]. Chaotic [Lin02, YZL^{+}06]. Characteristic [DCV^{+}01, LL01a, OB02, OMK09, QS02, IX07, PL04, RLZ03, Ser09, TY07]. characteristic-based [Ser09, TY07]. characteristic-wise [RLZ03]. Characteristics [ZTZ02, HMM08, Lee03, Lee05, Neo07, NDT06, SD05a, SD05b, SZH07, TOY09, ZR08]. Characteristics-Based [ZTZ02, Neo07, NDT06, SD05a, SD05b, SZH07]. Characterization [GD06a, FH03]. characterized [RC06]. Charge [CPP02, OMG02, SUW01, Ver01, ASQR06, CK07, LSA06, WR09, XDC09]. Charge-Conserving [OMG02]. charges [CDJ07, DC07]. Chebyshev [BK08, BDCG03, BRB03, Boy04, Boy05a, GH03, JW09, Lab09, LBS^{+}04, Sar03, VB08, ZP06, ZSTC06]. Chebyshev-filtered [ZSTC06]. Chebyshev/rational [Boy05a]. Chemical [JWSC00, JW02, LX00, MEG02, San01, SD00, ACGV07, AMH04, CP06b, CP06c, ELVE07, HL07a, JW03, LGP09, Liu08, MK07, OLA08, RE07, RHPN09, SZ08]. Chemically [BM01b, Li01, CGP05, CP06a, CFL^{+}03, NS05]. chemistry [AMH04, GMAj09, LLRP09, MESV09]. chemotactic [BCGR05, SL07c]. Chen [WS04]. chimney [KW08b]. Choice [TDV06]. choosing [AMXL09, FP08b]. CIP [IX07, TOY09, YMT^{+}04]. CIP/multi [IX07]. CIP/multi-moment [IX07]. Circular [HGM^{+}00, PG02a, ACR08, GGP06, KR09a, NCS03, SLC07, SSND03]. Circulation [DOWB01, Hig02, MR01, Hig05, SP06a, TVMR03, WDÖ^{+}03]. Class [GSD01, HR01, LP01, BAMD07, BG05b, DGH08, GS03d, pHL09, KPP07, LRZ04, MY09, RP08a, Ros08, Tsu06, XS06, ZSWW03, ZWS06]. Classical [BS00d, HGM01, BCCV09, CWL08, JR03, JR04, LQ09, LTD^{+}06, QCGQ03]. Cleaning [DKK^{+}02]. Clear [Bal02]. climate [Dic08, Lap08, Lyn08, MS08b, SW08c, Thu08b, dNWvSD07, dTWD09]. climate-prediction [SW08c]. cloaking [ZH09]. clocked [Mil05]. Close [POS00, CLL07a, HO08b, ZD05]. Closed [RK07]. Closure [DK02b, HHC08, PM07, RW08, SKWN03]. Cloud [MD02, SMT^{+}08]. Cloud-in-Cell [MD02]. Clouds [VCTS02]. cluster [CD04, FT09, Nit05]. Clustering [Gut00, MK02a, MG05a]. Clusterization [PA00]. Clusters [DPRS01, KKD08, Pal08]. Cnoidal [Boy02b]. co [BBF^{+}08]. co-propagating [BBF^{+}08]. coagulation [VK04]. Coalescence [CBL01, BJP04, FLM08, LMV04]. coalescing [ADS03]. Coarse [DEHL06, KMV03, RMGK04, IM07, KEB^{+}07, THL06]. Coarse-gradient [DEHL06]. Coarse-grained [KMV03, IM07]. Coarsening [Cho00]. Coastal [SR00b]. COBRA [SHWW00]. cochlea [GB03]. COCR [JHZ^{+}09]. Code [ALGM01, BM02, BADG00, CBB01, HF01, QRHD00, SHWW00, BM06, BvdHKG07, CN08, FM06, GHB03, GLN06, GBB^{+}06, HF08a, HDR^{+}06, IITV07, KB04, LGKP07, LL08b, NC04, OPML07, Roy05, SO08, SJHM09, TVMR03, TT06c, TPR05, WGCR07, ZK06]. Codes [PFB01, SMP01, Tót00, ADS03, FG06, HM09, PH09, PL04, TS07, dSHHM05]. codimension [CFF07, Min03, Min04]. codimension-2 [CFF07]. coefficient [Ber04, BK08, HL05, JBHK08, JZ08, LT05, MGC06, UL06]. Coefficients [PL01, VSMW01, Boy09, CT04, DGH08, HH07a, HyLL07, MD06, OK06a, SRNV07, TBT^{+}09, ZZFW06]. Coherence [BTSM09]. coherent [WR09]. Cold [VCTS02]. Cold-Atom [VCTS02]. collapse [BCGR05, Sus03, TU04]. Collection [TRL01, KFV07, WXG07]. colliders [QFR04]. Colliding [MKM99, CC08b, MKM04]. Collision [ADR08, Mu02, PRT00, SR00a, DWC^{+}09, DTS05a, DTS05b, KDK^{+}07, Lar03, LWDA09]. collision-driven [DTS05a, DTS05b]. Collisional [BZB00, KK00b, CF04, FPT05, GT09a]. collisionless [LCB04, VTC^{+}07]. Collisions [SSW01, AGW07, BBDE05, HS04, SK08a, She08, SMSS07, WLC^{+}08, XCRX08]. Collocated [LP02, CEH09, FL06, IA06b, MZ07, NMM^{+}07, NMH^{+}07, Ni09, SMS04]. Collocation [CSS00, KK00b, Lay02, PB00, Rei00, VB00, YKG04, AA09, Bia03, BK08, FWK08, FH03, GZ07a, GZ07b, GH03, Hei04, HK08a, KH09, KT03, LCCG05, MZ09, MK08b, ND04, VK05a, WG09, ZG08]. Colloidal [HHL00]. colocated [HM05]. Column [SUW01]. Combination [GG00]. combinatorics [RK07]. Combined [AA02, FVOMY00, SZ01, Car09, NI03, SLV09, TZ06, VS09, WZ09]. Combining [CWD08, FHJK09, SMP01]. Combustion [FH00b, GMB01, BEG03, LG03a, LP06a, LLRP09, MMPB07, YT07, vdBG09]. Comment [Aza09, CKR01, MCP03, TR07, Xu01b, LM03a]. commentary [SM09b]. Comments [PX02]. common [EF03]. Communications [KP05]. Commutative [HV03, MVM02, CBJdlC07]. Comp [DD03a, LM03a, MKM04, SM09b]. Compact [AC00, ACY00, Bla00, CT09, Cui09, DZ00, HT00a, HT00b, Hix00, KMJ01, Lai02, LS02c, LC01, MF00, NWZL08, PKP01, PS08, Pir02, PM00, Tol02a, Tol02b, TS02, Zha02, AV05, AZ03, BACFT05, Boe05, Cap08c, Cap09, CHB09, CL08c, CS09, DE06, DS06b, GG09b, IQ08, JAK05, Jor07, Kim07, LSB04, LL09, NLF03, NI03, NF09, NS05, PKD07, PYC04, PS04, PSG05, RLZ03, SGD03, SJD05, SDR07, SLV09, SYG06, SS05a, SZ05, STZ07, TD07, Tol08, WZ09, WF06, ZJS08, ZYHS07, KG09]. Compact-Difference [MF00]. Compact-WENO [Pir02, CHB09, RLZ03]. compactons [RV07, RV09]. Comparative [KKS07, GLLX08, MC06b, SB06c, TPVG06]. Comparing [WLC^{+}08, ZRS06]. Comparison [AV02, Bar02b, Boy02a, BUEG06, CMOV02, Fas03, GH03, GHV00, GC02b, HDC02, KMSH08, KN04, Mac00, MRS09, MBS03, QS02, SS01a, WPM^{+}02b, YFS01, Yua02, ZDNP00, ABRR09a, ABRR09b, BRB03, CGMS06, Eg07, EHS^{+}08, GR04, HS04, IITV07, KSW07, LTD^{+}06, Low04, NDG05, SSW^{+}07, ZKDT07]. Comparisons [LMX^{+}08, MP01b, CGMS03, GMS06, PR04a]. compatibility [RVDM09]. compatibility-constraint [RVDM09]. Compatible [CRB00, BBC^{+}06, BAFL09, LC06b, RK07]. Compensated [PSM08]. compensation [DL03b]. Complement [ACS00, ACLS03]. Complete [CL08a]. Completely [XY01]. Complex [DDH01, FVOMY00, KKC01, MF01, RRV01, UMRK01, AB05a, AMP09, BYZ04, BGS08, BGN03, BHP07, CHM08, CB03, COER07, FLE03, GS07, GSB03, GS05c, GN07, Had05, HM08, HHMK05, JL09, JHZ^{+}09, LZL03, LG04, LZ09c, LV07, MCM04, MCGV04, MDB^{+}08, MMPB07, MK06, Pop03, RJ06, RE05, SY03, SC08b, TAL09, TF03, VSV03, XLM07, YXLF05, ZJWC08, dSMF09]. complex-step [CB03]. Complex-Valued [MF01]. Complexity [Pau07, PWW00]. compliant [LTWW07]. complicated [SZS03]. Component [YL01, CKLS05, CLS05, JVVS07, Maz06, MLS^{+}05, MGNB09, SS04, TZ07a]. Component-Wise [YL01]. Composite [BM01a, Dri02, GA09, GL06, HC09, Jor07, ZC09]. composites [WP09]. composition [CP03a]. compound [Hau08a, Hau08b]. Comprehensive [VHI05, GB03]. compressed [HO08a]. Compressibility [HDC02, VLKM02, BCDR06, Ber06a, KKS05, SD05a, SD05b]. Compressible [AK01, ACK02, BCVK02, CFA01, CR02, DLS^{+}00, GS02, Han01, HH02a, LC01, PR00, Ros00, SLY02, SBGK00, Shy01, SFMP06, Sun00, SPW^{+}00, TSB01, WLE^{+}00, WZ00, Xu02a, vdVvdV02, AS03a, BSKH07, BKST09, BALW06, BLM08, BL09c, Boe05, BB08b, CPR05, CL07a, CL08b, CHB09, CJ09, CS07a, CZ09, DT04, DP07, DP08, DF04, DND06, ECL02, FK07a, FOLD05, FD07, GXW07, GFS08, GMD07, GR04, HJ09, HH08, HM04, HM05, HK04b, HKAH06, HAI09, IAT08, JC06b, KG08, KK05c, KK05d, KK05b, KvdVvdV06a, KvdVvdV06b, KvRvdVvdV07, Kok09, KSGF09, Lar09, LMX^{+}08, LFS07, LFX05, LS07, LLS09, LJW09, LDPL08, LKW05, LV07, LP07b, LCS09, LDV08, LSW06, LB03b, LJ06, LBL06a, LBL08, LZH^{+}06, Mai09b, Mai09a, MM03, MTWW06, MC06a, MB04, MSS08, MLS^{+}05, MBP07, MG05b, NOG08b, NGvdWS09, NDT06, NT07]. compressible [OF06, OPML07, PDHP07, PS05, PvdV08, PFSL07, PWM06, QA09, QLK07, Ros03, Ros07, SRM09, SFDL07, SPB09, SWK06, SM06a, SMB09, Shy06, SY03, SC09b, SK03, SCN07, SN08, TWM07, TT09, TMD^{+}08, TW05, TT06c, TR07, WAO^{+}04, WTL08, WM09, Xia04, XAI06, XLP05, ZGG03, ZSC08, dTDI^{+}07]. Compression [HHCL01, dCNHSD07]. Compressive [CLLG09]. Compton [DWLM09]. Compton-scattering [DWLM09]. Comput [ABRR09b, CL08b, HMS08b, HY11, HLWW06, JJGL07, Lau06, MN17, Mil07, SCC^{+}03a, WZL09b, dTWD09]. Computation [AIRY01, BCB03, Bae03, CS00, CGSS00, DDG02, DD05, DP00, EKK02, GG00, GM01b, GKL00, GM01c, JTB02, Khe04, LRS07, MS01, NR01, PK07, PSN00, RS06b, Shi07, SFW00, hRT02, VLKM02, Wee02, WZ00, BBK07, BJP04, CWJ07, CFM09, DBB06, FRS08, GT09b, GXW07, GMO04, GMS06, HS08a, Heu03, HLY09, JD09, JX07, Kar04, LCB09, LBL04, MC04, MGCR07, MT08, NL08, OB06, OJW06, PLS^{+}09, RJ06, RC06, Ros03, Ros07, SP05a, SKXK05, TZL05, TJ09, Tuc03, UTBV03, VCT09, WLT08, YC06a, ZSW03, ZW05, ZIP06, dSMF09]. Computational [BMRS02, BCE^{+}09, BPS03, Bor03, CL01b, Dar00a, HMK02, JY08, KM02, KMHR00, Myo01, OP02, Sau04, SHA08, SSD00, SZS01, Abr06, AK09, AHNS09, Bod06, CFR04, Cap05, Cap06, CKPW07, Dem04, FVE04, GWF^{+}07, GE07, Kim07, LB03a, MRS09, MJ06, Meh04, MGS07, MT07b, Myo04, PBH04, Roy05, WZL09a, WZL09b, YZL^{+}06]. Computationally [EHD08, LLRP09]. Computations [AK01, CAL00, DIV00, ES03a, Fre00, Gos02, HHCL01, JK00, KKR01a, KKR01b, KLvBvL02, SS02, TBE^{+}01, BB04b, BdCB09, BLM04, BCGR05, CWL08, DH07, EG08, GS03a, GKE04, HP04a, KM03, KLK08, LJW07, MKLU05, MK06, NA08, NJX09, RMV03, SMS04, TZ03, Tan08, VOD08, XLM07, XP04b, XHC08, YP06, Yan09]. compute [CXB08, CB07, VBL04]. computed [MLSD07]. Computer [Ota00, VP00, FSS03, GHB03, KKD08, LL06b, Lyn08, MC09]. Computer-Aided [VP00]. Computers [AKY01]. Computing [BNNP06, BLW01, BBK06, CF06a, CGL06, CCJ07, CEL06, DK06, DK02a, FCT07, Fre00, GST00, Han00, HL07b, JLOT05b, JLOT05a, KG03, LM08b, LAKD08, dlFMBdlFM02, Ovt08, PS01, RS00, SP07, SP00, TMND07, Wu02, BW06, BCL06, BS08a, Boy03, CORT09, Cec05, CL07a, CL08b, Chu09, CJR04, Jao07, LLS09, LW07, LW09, MR06b, SDR07, SH07b, SFVK06, Sus03, Vos06, Wen06, XMP07]. concentrated [DMHP07]. concentration [Bil05]. concentrations [Wen06]. concept [HF08b]. Concise [VQSZ02].

Condensate [VCTS02, BT03, CKLS05, CLS05, Yam05]. condensates [BW06, BCL06, BS08a]. condensation [BJM03]. Condensed [BS07]. Condition [AGP01, LFK00, NFK01, Vay00, WHV^{+}00, APQ03, BYS08, Car09, DS05b, FGP08, GV08, GP04, GK07, HAS05, Hu05, KDOO05, KL04, Li08b, LD04, Ten03, XD07]. Conditional [LLY05, MT04]. conditioned [ILL09]. Conditions [AC00, ACY00, AGH02, DMG00, DKSW01, Eli02, FT01, FSY00, Giv01, GVT01, Gro00, JL02, LOK01, MPC01, MPC02, OB02, Pet01, RC00, RTT01, SFY01, VDM^{+}02, YFS01, AST07, AMR06, AB03, ABK09, Ata04, AG08, BNV08, BBD04, Bér07, BA03, Bra08, CGMS06, CBI^{+}04, DH09, EZ08b, Eli07, FE04, GK04, GE07, GT09c, HMOG08, HK04a, HZ08, HEN09, Hel09b, HLL08, HSC09, HF08b, IK07, JM05, Kel05, KB06, LWW04, Li08b, LP06b, Liu09b, LDV08, LZC04, LCCG05, Mai04, MTH08, MJ07, ND04, NN04, NMS07, NB04, OPML07, PH08, PK05, PWM06, Pro05, Pro07, SSN09, SS05c, Sof09, zSW06, zS06, SK03, SCN07, SN08, THL06, Tan08, Tem06, TE08, TS08, Tsy03, XHW07, YE07, Zhe06, dA04]. Conducting [CPK02, Kan02, DND06, PL09a, RVVL09]. conduction [AMXL09, DQ04, FHLO08, GIA^{+}07, GIA^{+}08, GL06, JG09, MR07c, Mou04, Ols07]. conductivities [YWC07]. Cone [SS01b]. Confined [OL01, BWLM09, Chr03, PC08, VBJ08b]. Confinement [SUW01, Gos04, SKK^{+}08]. Conformal [ZSW07, CSML06, Hum05, LMS04, NCW^{+}09, OK07b, VZSL07]. Conforming [VR02, CCV03, CEH09, KT06, SB06c]. congruent [AD04]. Conjugate [PKvdB00, AMLC08, Fen06, HC09, Ovt08, Yan09, YLD09, ZW03]. Conjunction [TK00]. Connected [BMQS02, HJ02, VRM07]. Connecting [SZ00]. Connection [Lio00, Xu01c, VLW07]. Connectivity [SJ02, TB00b, SS09b]. Conservation [Asl01, BJ00, Bar02a, BIS07, CWT00, CDKP00, CRD02, FGG01, FMO00, GC01, Han01, JTB02, KH09, KT00a, LL00, Noe00, Per00, Sti02, TS01, Vas00, VS02, Wan02, WL02, YL01, ZSP02, ZYC02, AKLMP09, BAFL09, BT05, BBCT09, BCCD08, BP03, CLG07, Cap08a, Cap08b, CP08, CGKM06, CD07, CkM07, De 04, Edw06, ÉGP09, FS09, FL06, GV07, Gui05, HLMM07, HM04, Hub07, HO03, JR09, JTL09, KI05, LL03c, LVW06b, kM07a, MY06b, ML08, MESV09, Mil04, PDL09, RLZ03, RCD05, SW04b, SYG06, SAM05, SWL06, SZLW06, SR09b, Tak06, Thu08b, TT04, TT05b, TT06a, TT06b, THS07, WZL04, WG09, XS05a, YZF07, ZYL^{+}06, vDZ06]. Conservative [Abg01, AKO09, CBKM00a, CBKM00b, CL01a, CFJ06, CRD02, DLS^{+}00, FSB01, FK02, HLS02b, HEML00, IAT08, JJGL06, JJGL07, Jan00, KKL04, LM04, MF01, MC02, MGNB09, NTYT01, NTYT02, Nic00, OF02, Pir02, THD09, Tót02, VK05b, XY01, AK06a, CS07a, CS06, CS07d, DP07, DP08, DMP08, DBBP08, EB06, FS09, HHMK05, HKAH06, IITV07, IKS^{+}09, KD09, KPK09, LCS09, MS03, MM03, MC07c, MVO04, NMM^{+}07, NMH^{+}07, NGvdWS09, OK05, OKZ07, OK06c, RVDM09, RAD07, RSTB03, SZC09, SYC09, SS07b, SA09, SPGR06, TL06, TOY09, WAO^{+}04, XP04a, ZGK09, ZWS07, vBK03]. conserve [IG05, SHP07]. conserved [XMP07]. Conserving [BS00d, KKGL01, OMG02, BYS08, DOW08, DBS06, JL04a, KJ09b, MY09, VU04]. Consistency [MPC01, MPC02, BBC^{+}06, Dom08, LSW08, ZH04]. Consistent [BKR^{+}01, LOK01, MJ07, Ni09, OB02, SUW01, TE08, WHV^{+}00, XLM07, AJ09, AL08, BEA09, CLMRP08, DST07b, Gra06a, GS03d, IK07, IR09, LG03b, OL01, Pon07a, RHPN09, SCC^{+}03a, SCC^{+}03b, SC09b, WAO^{+}04, ZSTC06]. Consisting [CFA01]. consolidation [BFG08]. Constant [HS04, BMDS05, ET06, HA09, SD05a, TLL^{+}08, ZZ09]. constant-density [HA09]. constant-volume [ZZ09]. constants [Hei05, LTL^{+}09]. Constitutive [CT07, CPG04, TdAAP08]. Constrained [CBMO02, HMS08b, KM02, PGB05, YXU01, Abr06, Abr07, Abr09, AT05a, BTWGvBW07, COV04, GS05b, GS08, HMS08a, HS09b, IX09, KSS09, LPK05, Li08a, LD04, PM08, TFD06, TA06, UYK^{+}04, YMW06, Zie04]. Constraint [BFG08, Tót00, Yon01, Abr09, BLS08, RVDM09]. Constraints [LCS02, OS01, Kau03, MS08a, MC07a]. Construct [STiST02]. Constructing [LJS08, Aza07, Aza09, Che07, YC09a]. Construction [AM03, AM04, BBD04, FDD09a, GC01, MVM02, MY06b, Ohw02, QS02, VSMW01, CK08, DBTM08, GLM07, GD06a, MGS09, VGCN05]. Contact [DK02a, KJ01, PM02, RRL01, AZB09, Khe04, Spe05, VP09b, WAO^{+}04, ZGG03, ZGK09, vLAvdV06]. Contacting [VQLZ04]. container [SJ04]. containerless [AD03]. containers [FBHV05]. Containing [CL00b, FMO00, CGDT09, DP07]. Contents [Ano07d, Ano07e, Ano07a, Ano07b, Ano07c, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano08-36, Ano08-37, Ano08-38, Ano08-39, Ano08-40, Ano08-41, Ano08-42, Ano08-43, Ano08-44, Ano08-45, Ano08-46, Ano08-47, Ano08-48, Ano08-49, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36, Ano09-37, Ano09-38, Ano09-39, Ano09-40, Ano09-41, Ano09-42, Ano09-43, Ano09-44, Ano09-45, Ano09-46, Ano09-47, Ano09-48, Ano09-49, Ano09-50, Ano09-51, Ano09-52, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09-57, Ano09-58, Ano09-59, Ano09-60, Ano09-61, Ano09-62, Ano09-63, Ano09-64]. contents [Ano09-65, Ano09-66, Ano09-67, Ano09-68, Ano09-69, Ano09-70, Ano09-71, Ano09-72]. Continuation [SML02, BHL07, BHP07, CKLS05, SNGAS04, SO08]. Continued [Lin01, Ano07d, Ano07e, Ano07a, Ano07b, Ano07c, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36, Ano09-37, Ano09-38, Ano09-39, Ano09-40, Ano09-41, Ano09-42, Ano09-43, Ano09-44, Ano09-45, Ano09-46, Ano09-47, Ano09-48]. continuity [Jar04, Tok06a]. Continuous [CJ07, DPCV02, HEML00, BBvdV06, CVE06, EZ08a, FCD^{+}06, KEB^{+}07, Kim05, Ni09, NZ07, WAH09]. continuous-time [CVE06]. continuously [MM07]. Continuum [AA02, BS01, EH02, BB09c, FK09a, HW08, JG09, KZ06, KAA^{+}07, LSL08, LZ04, LCNR07, MMKP08, Ren07, SKS08, SWB^{+}06, SSE03, SB06b, SSB07, SBC04, TKH09, WL03, WWK05, ZL09, ZRS06]. continuum-atomistic [LCNR07]. continuum-field [HW08]. continuum-particle [ZL09]. continuum-transition [LSL08]. Continuum/DSMC [AA02]. continuum/particle [SBC04]. Contour [CPP02, SJ02, SLF08, SAKDJ05, SD06, VCM00, XCY06]. Contouring [Str01a]. contracting [PK07]. contraction [APP^{+}07, TCM05]. contrast [GL06]. Contrasts [VSMW01, EG08]. contravariant [LB04]. contribution [GLM07]. contributions [FSS03]. Control [AJG01, HGM^{+}00, KMA^{+}01, PGN08, RV00, Aza06, BC08, CC07, CY05, FLB03, GKD09, GL09a, HKM07, HZ07b, HN03, HS04, KLK08, MK04b]. Controllability [HMPR07, MHPR08]. controlled [CP04b, IG05, LG03a]. controls [ZJW06]. Convection [ART02, Alb00, CWT00, GZ01, KLN^{+}01, Kul01, KT00a, MPP01, SZ01, SWL00, Str01b, TC02, vdSE00, ART04, AZ05, BKS07, CA06, CEH09, CS09, Cho05, CS07d, DGH08, DR09a, EKP06, EKP07, FBHV05, GZ07b, HK06, ID04, KKS05, KZ04, KW08b, Kuz06, LCW04, LDW07, LS05a, MZ08, MC09, NPC09a, NPC09b, PS03a, PSC04, PSMW09, TD07, Tol07, VU04, VBJ08b, WD07, You06, ZGT06]. Convection-Diffusion [CWT00, KLN^{+}01, Kul01, KT00a, vdSE00, CS09, CS07d, DGH08, KZ04, LCW04, LDW07, LS05a, NPC09a, NPC09b, TD07, You06]. Convection-Diffusion-Reaction [SWL00]. convection-radiation [BKS07]. convection-reaction [HK06]. Convective [FH07, GHG01, PR01a, Ata04, Bil05, FP08b, KG08, SPLM09, Sus06]. convective-diffusion [SPLM09]. convective/absolute [Sus06]. Convergence [CLMRP08, CAL00, DCV^{+}01, GTRB09, GTD^{+}02, GMH06, GM01c, KDK^{+}07, LZ07, Lee05, PS02, PFB01, Saf00, STR07b, SPW^{+}00, BAR08, BB08a, BS09a, CMG09, DVHM05, GS06b, GP05, Hel09a, HT03, HJM06, JS07, KJ09a, KS08b, LY07b, Maz06, NOG08a, NvL03, ODAF07, PPB09, SBA07, SY03, zSW06, zS06, Tor03, TB04, Tow08]. Convergent [DDH01, deM02, Gon07, JTL09, MGC06, TCM05, VSW04, VSW06]. converging [Boy09]. Converter [KMA^{+}01]. convex [HJJ09]. convexity [De 04, XP04a]. Convolution [RM01b, WPW02, BKM09, Boy06, GvH06, WZ07]. Convolution-Finite [WPW02]. Convolution-Thresholding [RM01b]. COOL [CGC^{+}09]. Coordinate [Bon00, FK02, HK01, MC00b, Wu02, HWW07, LRS07, LB04, SS03a, WS04, ZKDT07, dHRvdB07]. Coordinates [BM02, CSS00, CL02, NC01, VR02, Ano08-50, BN04, CJR04, CK07, DB04, GYKL05, KRT^{+}09, LGHD08, LPK05, Mea04, MVO04, Nik06, NB04, OBT06, SR09a, SM09b, SHY07, VRM07, WAH09, XLP05, Yam05, YHSX07, vdHK07]. copper [ZSB^{+}08]. Core [TR02a, HSC09, SW08c]. core-spreading [HSC09]. Cores [CKS00, LLB05, Thu08b]. Coriolis [AKO09, HC08]. Corner [HO08a, Boy05a]. corners [Boy03]. corrected [BS08b, CL05, CL08d, FWW04, MB04, Str07a, dFGLS05, dFJS09]. Correcting [SHP07, SK04a]. Correction [AV03, KLN^{+}01, KT02, MD02, MOS^{+}00, MPC01, MPC02, SM09b, ASPB03, BLM03, Che03, CL07b, DL03b, FG04, HJM06, HJM07, JH08, LM04, MTV08, NVD05, PG04, RVM07, RVDM09, SLC07, Wal03, WYS09, WS09, XYK05]. correction-based [WS09]. correction-lattice [SLC07]. Corrections [BC02a, THN^{+}07, VGCN05, XS05c]. corrector [CPKW09, CMSZ09, LRS09, TWYC06]. correlated [KS08b, AGT05]. correlation [LL04a]. correlations [MPD08]. correspondence [PHKF06]. Corrigendum [LLIK01a, MKM04, MN17, SCC^{+}03a, dTWD09]. cosmic [Min07]. cosmic-ray-hydrodynamics [Min07]. cosmological [RHPN09]. cosmology [WJV07]. Cost [LC06a, BCE^{+}09, LQ06]. Cost-effectiveness [LC06a]. Couette [LR03]. Coulomb [AKV00, DWC^{+}09, GH02, GM01b, HB05a, KK00b, Lar03, LWDA09, LJK09, PC02, Saf00, Saf02, She08, SS01b, WLC^{+}08]. Counting [Bow01]. Coupled [CFM09, DE02, FLE03, GA09, KZ06, Man02, MC02, NVD07, SP00, VDM^{+}02, AK06a, Alb08, AMS04, BKS07, BBDE05, BFG08, DSM09a, DH07, Doh09, Eld08a, GT09b, GGS09, GFR09, GGCC09, HBLD07, HMMR04, JG09, KLSW09, Mou04, NVD05, NGC^{+}07, Ols07, ODCK07, PR04b, PC06b, RBSL06, Ren07, STD^{+}05, Sus03, TC09a, WLC^{+}06, YJL^{+}06, YLD09, YSS05]. Coupling [BQQ09, ČPT01, Dar02, Del03a, Fed02, GTD00, GB08a, GL09a, SSE03, UH01, WL03, WK01a, YMF01, AHMS03, AL08, BCG09, CPKW09, CELS07, CS07c, CC08b, DM03, DDM07, DTMS06, DST07b, ED07, IA06b, KYK07, LKMK09, LM03b, MMS04, MU09, NMG09, Pon06, Yam05]. couplings [VZSL07]. Courant [KDOO05]. Covariance [SL06]. Cox [MR01]. Crack [ÁDIM09]. cracks [Oh04, PL09a]. Crank [Han00, KW08a]. Creation [OMG02]. Creep [Sie00]. creeping [Kro01, Kro02, MR06b]. Criteria [SV00, CHM08, LG08]. criterion [KP08]. criterions [HX05]. Critical [AV02, GGL^{+}01, KMJ01, Maz06, HAP05]. Critique [Mac00]. crossed [HDBW05]. crosswind [BEG03]. Crystal [JK02, LS02b, NDG05, BS05, CW08, DQA08, DBB06, GJK09, HWWL09, Lap03, LL06a, LLZ07, PSCQ03, Sau04, TBT^{+}09]. Crystal/Melt [LS02b]. Crystalline [EH02, GM04, Gos04, GM06, Tan08]. crystallization [Lap03]. Crystals [CD00, DGP00, KM02, BS06b, Chr03, DD05, LR07, ON08, YLA08]. CSP [VGCN05]. cubed [CX08, Cho05, PL07]. cubed-sphere [CX08, PL07]. Cubic [CP04b, Lay02, BIS07, CLS09b, PSC04, Zhe06]. Cumulative [Ano00-28, Ano01-28, Ano02-28]. cure [LJ09b]. Cures [sKKRH03, PD01]. Curl [CL06a, TR02b, Wel07]. Curl- [CL06a]. Curl-Preserving [TR02b]. Current [Ver01, BCDW06, BO04, CBC09, CDV05, EPW08, FM06, LTD07, NMM^{+}07, NMH^{+}07, SK05, VTC^{+}07, Wea09]. current-carrying [CDV05]. Currents [JTB02, GCW07, Pee03, SK08a, VBL04]. curse [KDOO05]. curvature [Bur05, BHSV07, ML06a, Shi07]. curvature-dependent [Bur05]. curvatures [RMB07]. curve [CFF07, SK07a, WSTW09]. Curved [MSYL00, Chr04, GH08a, JJGL06, JJGL07, JY08, KY08, KAK03, KB06, NGC^{+}07, QP03, RBL04]. Curves [BCMO01, CBMO02, KKGL01, LZ09a, MR07b]. Curvilinear [BM02, BGS08, JMK01, MR01, NC01, SK05, SCD00, VR02, VG02, XCZ02, BN04, GS07, HWW07, KL08, KRT^{+}09, Kok09, LB04, Nik06, VRM07, WS04, Yam05, vdHK07]. cut [FD07, LTD07, RCB05]. cut-cell [FD07]. Cycle [GHV00, BPM06, SJC07, XYK05]. Cyclotron [OL01, GLS03]. Cylinder [HGM^{+}00, MK02a, BC08, BT07a, BT07b, DCK08, KR09a, MPD03, NCS03, SLC07, SSND03]. Cylinders [AD01, PG02a, AL06, ACR08, MAL09, TOZP03]. Cylindrical [CSS00, CL02, CPP02, FK02, GBS00, LMS02, Nit01, Sie00, BS04c, CK07, DB04, Mai09a, MVO04, OBT06, OPML07, RRC05, XSG08]. Czochralski [JK02, PSCQ03].

D [MKM04, CHL09, AV05, AFGM07, Alb00, ASQR06, AMXL09, ALGM01, AQ00, BM02, BO05, BBDE05, BM06, BCK09, BMN07, BWLM09, BRC^{+}09, BBvdV06, Ber06b, BMQS02, BPO07, BPL06, BGS08, CD03, CJSS08, CD04, CWJ07, CPG04, CGMS06, CGN^{+}07, CdHST08, CBKM00a, CBKM00b, CK03, CBB01, CTT08, CLL^{+}07b, CJ07, CL00b, CP04c, Dar02, DBF08, DS06b, DGP00, DH07, DB04, Eld07, EHST03, EKBL09, EAY01, ES06, EKP06, GS05a, GS06a, GH03, GS07, GW02, GHB03, GSB03, GS05c, GBB^{+}06, GKL03, Gui03, GLLX08, HAAO00, HWL08, HH01, Her09, Hor02, HS08c, HSL08, JBHK08, JHSZ07, KKCF09, KB04, KPP09, LGP09, LC06a, LS02b, LDN04, LLIK01b, LJ09a, LS00, LW07, LH08a, LF04, LC03, LL08b, LZH^{+}06, Mac07, MKM99, MP07a, MKLU05, MG07a, Meh04, MSYL00, MP01b, MAL09, NCW^{+}09]. D [OTCM08, ORM06, OJW06, ON08, PAD07, PSC04, PCP08, PVPS09, QS07, RWWS07, RW08, RKE^{+}07, Rom02, Rom07, RMF08, RFVP09, RVVL09, RW03, SROCdPFF05, SWG08, SMP01, SL07a, SA09, SPLM09, Stu01, SP00, TJS03, TS02, TPR05, VGS04, VL07, VBL07, VGZB09, VGBZ09, WGSL06, WM09, WFC09, WZ09, XMT06, YM07, YBZ06, YJF^{+}06, ZYKW01, ZZ01, Zha02, ZK05, ZJW06, ZJWC08, ZTPM05, dHRvdB07, dSMN^{+}04, vDZ06, vZdBB07]. D-leaping [BCK09]. d-quadratic [CHL09]. Damped [Küm04b, Pro03]. Damping [HZ02, APT09, BCG09, CL06a, LH08b]. Darcy [EZ08a, BT09, GD07a, TC09a]. Darcy-flux [EZ08a]. Darwin [SG06]. Data [DCS00, GZ07a, Gut00, KFH^{+}04, LCS02, Mac00, TK02, WPM^{+}02b, DEHL06, DS08, GZ08, GD06a, GSK06, HS07a, KFIG06, KSJ03, KE09, RR07, RA09, SS09b, TPVG06, dCNHSD07]. Data-driven [GZ07a, GZ08]. database [TZ07c]. Daubechies [NG06a]. Davidson [BPS03, CL00b, SWTM01]. dbar [KMS04]. dbar-equation [KMS04]. DC [SUW01]. De-aliasing [KK03b, SVB09]. dealloying [EE08]. deblurring [CJLS09]. decaying [KMSH08, TMD07, YGL05]. decision [SMSS07]. decker [FK09a]. decomposed [Ber04, BUEG06]. Decomposition [BC02a, IK01, PS02, QS02, Stu01, Sum00, AV05, ABLS05, ARRS09, AAC07, AL06, BIW04, BCHL07, BL09b, BT07a, BT07b, BB09b, CC03, CELS07, CRB^{+}08, CWD08, DDK06, DLP08, Edw06, Eg07, ES03b, Gom08, IQT08, KKCF09, LVL05, LDL^{+}09, LJ07, NPH09, NL09, OMK09, PA05, SZB^{+}07, SDR07, STD^{+}05, SJC07, SXyWX09, TET09, VZSL07, WL03, ZT07a, ZSP08, dCNHSD07]. Deconvolution [AS02, AHF04, HBLD07, HAD06]. decoupling [GB08a, RVM07, SMAj08]. Dedication [RP08b]. deep [SW08c]. deep-atmosphere [SW08c]. deeply [BLW04]. defeating [Boy05a, Boy05b]. defect [CLL07a, KH08, PG04]. Defects [DDG02, VDM^{+}02, HK06]. deferred [BLM03, HJM06, HJM07, JH08, LM04]. Deflagration [GP00b]. Deflagration-to-Detonation [GP00b]. Deflated [AMLC08, VSMW01]. Deformable [TC02, ZD00, CA06, LL07, ZEA06, ZD05]. Deformation [GH09, LLdlP^{+}00, DLW06, FKK08, JA08, LS08, LQX06, MDM03, PS03b, SCRL08, VQLZ04, XMP07, ZK05, vZdBB07]. deformations [CGDT09, DT03, FGS09, MV08, ZFM08]. Deformed [AD01, AKL^{+}08]. Deforming [VG02]. Degasperis [FL09]. degeneracy [GS05a]. degenerate [BAR08, WC08]. delamination [Oh04]. Delaunay [GS09b, LQX06]. delay [GKE04, KG03]. delays [BCK09]. Deleted [Boy02b]. delta [Bea08, ETT05, MG08, Sme06, Tow07, Tow08, Tow09a, Wen07, Wen09, YZLH09]. Demonstration [TWS02]. dendrites [TZ07b]. Dendritic [ART02, EKK02, GW02, PK00, ZH01, ART04, DQA08, TZ06, TZ07a, WLT08, ZGT06, ZVHP03]. Dense [Sni01, FY07, LMV04, LZL03, MEKS03, NFvS^{+}06, SH07a, WWK05]. dense-gas [SH07a]. densities [BCDW06, SK08a, Sti05]. Density [BKR^{+}01, CYKC01, Cul01, FS00a, FS00b, GBS00, GQ00, Lou00, NFvS^{+}06, OS01, Pai01, SBGK00, Ver01, AT09, CCG08, Chr03, DBBP08, DSS07, FHW07, GS09c, HJFW04, HA09, IOTK04, Küm04b, LL05, LP06a, LF04, MP05, MP07b, MJ07, MDR07, NMM^{+}07, NMH^{+}07, Ni09, PS07d, RVM07, RVDM09, Ros03, Sam09, SF03, SD05a, SD05b, SE04, SDT08, Sur05, Tok06a, YZ07, ZSC06]. Density-Functional [Lou00]. density-functionals [Küm04b]. Density-Stratified [Pai01, SE04]. Dependent [AGH02, ACS00, ELW01, Gen01, Nys02, RTT01, VR02, AZB09, AFGM07, ACLS03, Ata04, BIW04, BH05, Bur05, CT08b, CJ07, DL04, DR09a, DKS^{+}03, FD03, FKLY07, FH03, GN03, GP04, GK07, HDBW05, JBHK08, KW03, LWG03, LP04a, LB04, ML05, MU09, OPML07, RCD05, RRW05, SV07, Ten03, WRu03, WS04, YYF09]. depending [Tok06a]. deposition [AMH04, CK07, RRV06, ZK04]. Derivation [MvW08, SZ05, AI09, LT09a, OF06, SD05a]. Derivative [TT06a, CB03, Jar04, KYLB07, RC09b, ZW04]. Derivatives [ELC02, Giv01, BEE06, BHR04, Doh09, Gro06, Gro07, HKO07, MN04, MN17, ND04]. derived [MC07a]. descent [CSMH05]. describing [CLTA07]. Description [SUW01, CHBS04, HS09a, LGKP07, LL03b, LJS08]. Design [GGF03, HFO01, KHV01, LTL^{+}09, SW00, WD07, XYK05, BHS03, CBGI09, FK09b, Hab04, Kuz06]. designing [ERVE09]. detachment [BHL^{+}04]. details [DTS05a]. Detection [GKL00, AGSX09, HD07, PW07]. Determinate [Boy02b]. Determination [Dic08, GM01b, AKL^{+}08]. determining [EN06, Pee03]. Deterministic [ELC02, BCCV09, Cha07a, GS05a]. Detonation [BJ02, BSJ01, GP00b, CDS04]. detonations [HAP06, TV08]. detrended [Ham07]. develop [LS05a, Rah04]. Developing [DZ00, DF00a, FE04, KSJ03]. Development [BW02, CKR00, CKR01, CR00, EKP06, FT06, FCB02, sKKRH03, MEKS03, SYC09, SSD00, Tol07, Tol08, WLC^{+}06, Xu01b, YS07a, ZJS08, CS09, Hig05]. deviational [HH07c]. Device [DE02, CGMS06, CELS07, LSS^{+}09]. Devices [AIRY01, MP01a, MP02, ST01, And09, CGMS03, CL03a, CL05, FH07, GS06a, dFGLS05]. Dey [NCW^{+}09]. df [Chr03]. DGBGK [NJX08a]. diagnosis [HM09]. diagnostics [ACGV07]. diagonal [Boy05b, Lur07, Tol08, UL06, WC07]. Diagonalization [TR02a, CP06b, WC08]. Diagrams [DSS00]. diameter [AV03]. diatomic [Myo04]. dielectric [CDJ07, DBF08, DC07, EG08, Mar06, ZK05]. dielectrics [WC07]. Difference [AC00, ACY00, ADK00, ADK02, Azm02, BR09a, BC02a, Bla00, CS01a, CBB01, FVOMY00, FK02, GHV00, HLS02b, HGN00, JL02, JMP02, KMJ01, MP01a, MP02, MF01, MF00, Nic00, NC01, PK00, POS00, Rem00, SV00, TK00, VCP00, Vas00, VCTS02, VG02, WA02, YP01, ZZ01, AE03, BS04a, BG07, Boe05, BMDS05, BSP06, CHH06, CdHST08, CN05, CYS06, CS06, CS07d, Cui09, DMBS05, DBBP08, DS06a, FDD09a, FDD09b, FK07b, Gro06, Gro07, GH08b, GLT07, GL09b, GL08, HP04b, HWWL09, IK07, IM05, IM07, IQ08, JD09, JAK05, JM05, Jon05, Kim07, KPP07, KPP09, LG08, LJW09, LX07a, LMS04, LSS06, LVW06a, LLTA07, LS09, MN04, MN06, MN17, MST06, MSP^{+}06, MGC06, MVO04, NI03, Nik06, NF09, PAD07, PYC04, PH06, PH08, Pir07, RB06, Rom07]. difference [SROCdPFF05, SHA08, SHWC07, SYG06, SS05a, SZ05, STZ07, SC09a, SS03b, SS05c, Sou09, SB03, zSW06, zS06, SN06, SCN07, SN08, TJ09, Tan08, TD07, TdAAP08, Tow08, Tow09b, Tsu06, VPMC04, VLW07, WG09, WF06, XS05a, XS05c, YMWM06, Yus06, ZZ07, ZH09, ZYHS07, dSHHM05, dVGLM09, CBKM00a]. difference-type [WF06]. difference/spectral [LX07a]. Differences [BBHM09, DF00b, Tol02a, Tol02b, Boy06, IOTK04, Kum04a, LRS07, MLSD07, Tow09a, WZ07]. Differencing [CM02, HH07b, Jor07, Liv07, LC06b, SZC09]. Different [WK01b, NW07, QKS06, SD05a, ZQ09]. Differential [AGT02, ABGV02, BCOS01, CKL00, GTD00, HMS08a, HMS08b, Hua01a, MF01, MOvL00, SCD00, Tuc03, VB00, APR09, AKV06, AGT05, Asl04a, BV05, CP03a, Chu09, DI09, EN06, GK03, GKE04, GBS06, HR01, HJM07, IAT08, IDD04, KG03, LP04a, LCdCN^{+}03, MZ09, MP07b, MK08b, MSO04, MT04, Ngu07, Ngu08, PSD09, PCS^{+}09, RBvdV08, RM08, SS08, SRNV07, SKW05, SG03b, TE04, VSG05, WK05, YZW05]. differential-algebraic [VSG05]. Differentiation [CSV00, BBB08, CP04a, GT05]. Diffraction [Kan02, WWVG00, BHS03, JY08]. Diffuse [DSS07, Gla01, FGS09, Kim05, RRV06, Sof09]. diffuse-interface [Kim05]. Diffusion [AGT02, BKR^{+}01, BMS00, CL00a, CWT00, DE02, EES09, GZ01, Gen01, HFO01, Her00, HGM01, JR07, JM00, KLN^{+}01, Kul01, KT00a, Li01, MHS02, MR07b, MHS01, MKR00, NGC^{+}07, OGV02, PK00, SWL00, SSC00, VDM^{+}02, WDM01, vdSE00, AS03b, ACGV07, AZ06, AINR03, BAYZ08, BMN05, BBHM09, Bar04, BBDE05, BMT09, BHR03, BM07, BSW05, Bur05, BEG03, BB08b, CLTA07, CS09, CP04b, CF06b, CS07d, Chr04, CS04, Cui09, DPRN06, DGM07, DGH08, DL04, DUEB07, EULM03, FG04, FG05, FM08, GZ07a, GLM07, GT05, GL08, Her09, HG03, HMR08, HST09, IG05, JBHK08, JHSZ07, KZ04, KLM07, KSS09, LT05, LG03a, LH05a, Lar07, LWG03, LLC06, LCW04, LLOT06, LDW07, LS05a, LX07a, LMS04, LSS06, LSSV07, LSV09, LOK05, LLGL07, LGM08, Low04, MJ09a, Mad06, MM07]. diffusion [MP07a, MEKS03, MMKP08, Maz06, MP07b, MG07b, MSP^{+}06, Moo03, Moo07, MT07b, MK03, NV09, NZ05, NPC09a, NPC09b, Nis07, OS04, Ols07, PSZ09, PS07b, Pud06, RSM05, RBH03, RSO04, RS05, RS09a, SCT09, SW04a, SH07c, SO08, Sou09, SPLM09, TMS06, TM07, TD07, VSV03, VSH04, WG06, XS09, You06, YA05, YS07c, YS08, Yus06, dFGLS05, dFJS09]. diffusion-controlled [IG05, LG03a]. diffusion-reaction [MMKP08]. diffusion-type [Lar07]. diffusions [Buc05, LN09]. Diffusive [Azm02, JP00, TAL09, XS05c]. diffusivities [PSZ09]. Diffusivity [ML01b, FL07, KL08]. digital [KSJ03]. dike [LTD04]. dilatation [BS04b]. dilute [DFV08, Fox08]. Dimension [HA02, JWSC00, BFT07, Boy03, CDDL09, Cec05, COQ06, GZ08, JW03, Min03, Min04, WO05]. Dimensional [AJG01, ART02, ACS00, BMR01, BMRS01, BMRS02, BdLL01, BZW01, Cal02, CRB00, CWT00, CMOV02, CD00, DMG00, DCV^{+}01, Del01, DK02a, DOWB01, Eli02, FVOMY00, FS00a, FS00b, Goe00, HK01, JW02, KK00b, KP00, LL00, LCS02, LK01, LMS02, Lou00, LWEM00, MR00, MR02, MC02, Pai01, PKvdB00, PL01, PWS^{+}02, RV00, SHWW00, SJ02, Sni01, VD00, VD02, VS02, WK01a, WL02, Yua02, ZSP02, ZYC02, AvdB04, ARRS09, ART04, AK05, AV03, AC05, AB03, ABK09, AI09, AT09, AMS04, AMSZ07, BTW04, BFC04b, BS04b, BS04c, BDCG03, BM07, BBK07, BH05, BH04, BHP07, BL03, BCI^{+}08, Cap06, CQO04, CKPW07, Che04, CC07, CSO09, CR09, CFGK05, CY05, DCF^{+}08, DIL03, Dim07, DLP08, DI09, DS09b, ECL02, Eli03, Eli07, FNS07, FRS08, FS09]. dimensional [FHLK05, FCGK05, FKK08, GS09b, GB03, GP04, GGP06, GWF^{+}07, GM04, Gos04, GM06, GKE04, Gro06, Gro07, Gui05, HT07, HZ08, HZGB05, HP04a, HD07, HAP06, HS08a, HT03, HGB^{+}03, HLWW04, HLWW06, HWW07, IHL03, JVVS07, JX06, JN07, JW09, KKS05, KSHS08, KK05c, KK05d, KLK08, KAK03, Kro01, KLP^{+}09, LKD04, LG09, LSD07, LWP^{+}09, Lee03, LZ09c, LDW07, LDPL08, LS05a, LR07, LT09b, LL03c, LTD^{+}06, LVW06b, LTC07, LL08a, LDV08, LJ09b, LP04b, Ma05, Mai09b, Mai09a, MMS04, MRRS05, MSJ07, MST06, MP03, Men04, MR04, MG08, MT07b, MGNB09, NTYT01, NTYT02, NTB07, NFA03, ODAF06, Ols07, OLLL03, PKKL05, Pon09, PA07b, QS04, QLK07, RB05, RRC05, RS06a, RC06, SKWN03, SBGK00, SSND03, SS07b, SP06a, SK04b, SCRL08, SS04, TM07, TOZP03, TM05, TPV07, TXCD07, TT04]. dimensional [TT05a, TOY09, TC07b, TC09b, TG08, TA06, UL06, VGCN05, VVS08, VCG03, VD03, Wag05, WK04, WZL04, WW04, Wen09, Xia04, XAI06, XHW07, XG09, YAvdB^{+}08, YYT05, YXLF05, YKK08, YW07, ZWS07, ZP05, ZH09, ZLAC05, Zhe06, ZQ09, ZT07b]. Dimensionality [MN09b]. Dimensions [BCMO01, LTZ01, LTZ02, Nys02, RW00, SWL00, TNR02, CM06, CHL06b, CCG^{+}06, Che07, DLW06, DR09b, EES09, GG04, GS08, GH02, GD06b, HLO08, HB05a, HB05b, JBF07, KLM05, LLP07, kM07a, MCG08, MR05, Moo03, Moo07, NWZL08, RS06b, SBCL06, Shy06, TTZ03, TT05b, VWW04, Wan04a, WO09, YBZ04]. Diodes [deM02, BMK^{+}06, DGM07]. diphasic [Del07]. dipole [KDK^{+}07]. dipole-wall [KDK^{+}07]. Dirac [BL04, ETT05, HL06b, HJM^{+}05, MG08, WT07b]. Direct [BRL02, CSS00, FLG01, FLM08, GPH^{+}01, HdGKG08, HPZ01, HHM04, JLCD01, KB00, KH07, PG02a, PWS^{+}02, Ros09, SW08a, SP04, SL04, SCW^{+}09, SB02, Tak06, UL06, ZS01, AMH04, BHL07, BALW06, CTW^{+}08, Chu09, CP04c, Dom08, FM05, FE04, GS06a, HK08a, HM05, JD04, KK09, KHdT^{+}08, KSJ03, KS07, LDV08, LQ06, MTWW06, MC06a, MR05, MR07a, Mot08, Pet07, Pro05, SMS08, SP05a, TWM07, Uhl05, WMH07, YS07b, ZKDT07, ZZ07, ZD08, GKJW07]. Direct-expansion [Tak06]. direct-forcing [YS07b, ZZ07]. direction [CSMH05, SMAj08, ZXQX08]. Directional [NTYT01, NTYT02, SZ01, BF08, KW08b]. Directionally [BST01, BST03]. directly [BT03, CS07b]. Dirichlet [Bia03, GP04, GK04, Gui03, HW05, Hel09a, HO03, JM05, Mil08, NR01, SSN09, TB00a, YLA08]. Dirichlet-to-Neumann [GP04, GK04, Gui03, TB00a, YLA08]. Dis [BBvdV06]. disc [Hei04]. Discharge [CYKC01, KMA^{+}01, hLA01, DMR09, SHPC09, SS04]. Discharges [HK00, HMM02, SPC01, PSCB08, SMSS07, UBRT07]. Discontinuities [Asl01, NFK01, BFT07, Boy03, FH03, HN03, kM07a, Pri08, TJ09, THS07, VVS08, WAO^{+}04]. Discontinuity [AGSX09, WC01, KL08, KYLB07, ZGG03, Zho07]. Discontinuous [BSJ01, BT02, BSB01, DPCV02, Gab07, GHW02, HH02a, HA02, Hub08, LS00, LZC04, Mac07, MPFC08, PL01, RH01a, RBvdV08, YS06, vdVvdV02, Ain04, AB07, BCDR06, BDHN09, BRC^{+}09, BTT08, Ber04, BG05b, CT04, CDI09, CC07, CELS07, CS07b, CS08b, CHG^{+}07, CLS04, CQRW05, CFP06, DD09, DLP08, DF04, DBTM08, ES06, FCJ08a, FCJ08b, FK07a, FOLD05, GLM07, GLMH09, Gir06, GR08, HH07a, HH08, JH06, JW06, KCGH07, KvdVvdV06a, KvdVvdV06b, KvRvdVvdV07, KWBH09, Kri07, KWD07, KDW08, LGHD08, LSY04, LSZZ08, LJS08, LY06, LX07b, LGM08, LBL06b, LBL07, LBL08, MRC06, MR06a, MGCR07, MN06, MHI08, MZ07, MESV09, NM06, NL08, NPC09a, NPC09b, OK04, PvdV08, QS04, QKS06, QLK07, RBS06, RC09b, SFE07, SMB09, WM07, WM09, WG09, WKG06, XXS07, XS06, XS05b, XLS09a, ZZFW06]. discontinuous [ZQSD08, ZQ09, vdVX07]. discontinuous-Galerkin [KCGH07]. Discontinuous-Pressure [BT02]. Discrete [AS03a, BSJ01, Coe02, FF02, FGG01, FHLO08, GC02a, LL01a, MD02, Mar09, Mie00, PS07b, Poz01b, RTT01, SZ08, SS00, WPW02, AST09, BBC^{+}06, Bea08, Boy06, BL03, CLS^{+}06, CL07b, CT07, EULM03, Fen06, HV03, KWD07, KT00b, KSS09, LGP09, LK09, LC06b, MN09a, MY06b, MGS09, MD06, NZ07, PS07a, PA07a, PCS^{+}09, RVDM09, SFVK06, SC09b, WZ07, YZLH09, ZXQX08]. Discrete-element [Mar09]. Discrete-Velocity [Mie00]. Discretely [RC00]. discretisation [RJM07]. discretisations [Bal08]. Discretization [Bar02b, BMS00, DMR09, Edw00, ETT05, FMO00, GFCK02, JP00, LBV00, MHS01, NE05, PYC04, SC01, Tót02, Zha02, AMR06, AB07, AB03, AKO09, BAYZ08, BB07a, BP03, BMDS05, BSP06, CS08a, Dar02, GF05a, HH08, IS04, JHSZ07, KK05c, KYK07, LL05, LSS06, LCS09, ML06a, MVD04, MKKY06, MZ07, MHPR08, NOG08a, Ols07, PvdV08, RBSL06, RWWS07, RS09b, SVB09, SP06a, TAL09, VV03, VK09, VWW04, Wan04a, BT07b]. Discretizations [Boy02b, WK01b, ZDNP00, AD04, BHvdV06, CFR09, DWLM09, DF07, EV03, FDD07, FOLD05, FD07, HMPR07, KvdVvdV06a, KvRvdVvdV07, KWD07, MGS07, MG08, MAN^{+}06, NFGK07, SMB09, TW05, TR07, WM07, ZT07b]. discretized [Chu09, DLP08]. Discretizing [Tow09a, Tow07]. disease [NLT07]. Disk [BD01, WB01, BK08, LT05]. Disks [dlFMBdlFM02]. dislocation [SFVK06, WGSL06]. Disperse [PO01]. dispersed [DDK06]. Dispersion [CL01b, MBP00, PFB01, VBL07, ZF02, CS09, CLS09a, FK07b, Kok09, LS05a, LLTA07, MST06, PSG05, SLV09]. Dispersion-Relation-Preserving [CL01b, CS09, CLS09a, LS05a, PSG05]. Dispersive [Ain04, CL01c, SW08b, BN04, BBMB07, BB04b, CJSS08, GP04, KSH^{+}06, LSY04, LZC04, MY09, MGS09, PC08, ZH09]. dispersively [SYC09]. displacement [VQLZ04, ZVQ07]. Displacement-driven [ZVQ07]. disposal [KP07]. Dissipation [SVB09, Xu01c, YVD00, Dwi08, LJ09b, PDHP07, PK03, PM08, RV09, TDWY08, VBL07]. dissipationless [ZGG03]. Dissipative [HJJ09, LMH07, MF01, MPFC08, WHV^{+}00, Ain04, AWK07, BBMB07, BB04b, PK05, VHI05, VHI06, YS07a]. Dissolution [JVVS07, EE08]. Distance [MS01, RS00, hRT02, BBK07, JC06a, Tuc03]. Distorted [Her00, YS07c]. distortion [KK09, ZJW06]. Distributed [SPT05, BYZ04, BG05a, Boy06, CV06, DLMK04, LJSM08, VB08, WZ07, vdDA06]. Distribution [Abg06, CRD02, JK02, WB01, AM03, AM04, CS06, CS07d, DPRN06, GW06, Hub08, NFvS^{+}06, Nis07, RCD05, RAD07, RB09a, Ros08, SH07a, ZZ09]. Distributions [CVB00, Pop00, VS07]. Disturbance [DC01, Sur05]. Divergence [Bal01, Bal09, DKK^{+}02, MOS^{+}00, SCC09, Tol02a, Tol02b, TR02b, AT05a, AT08, BRDM09, CLS04, CEL06, LL04b, LD04, NMS07, TA06]. Divergence- [TR02b]. Divergence-Free [Bal01, Bal09, BRDM09, CLS04, LL04b, LD04]. divergence-preserving [AT08]. DLM [SL07a, Yu05b]. DLM/FD [SL07a, Yu05b]. DNA [GPL05, vHBB02]. DNS [DHM07, KIH09, Pro07, YGL05]. DNS/LES [DHM07]. Domain [ARRS09, BIW04, BC02a, BCM09, CR08, CBB01, CC03, DDF01, GHV00, GPH^{+}01, HW02, MKL06, PS02, POS00, Rem00, SZB^{+}07, Stu01, VDM^{+}02, YP01, AvdB04, ABLS05, AA09, AL06, BCHL07, BG05a, BSLN09, BP08, BUEG06, BB09b, CdHST08, CELS07, CTT08, CWD08, CFP06, CD07, DDK06, DGMN03, DLP08, FLE03, FK07b, HZ08, IQT08, JM05, KF06, Lau04, LW06, LVL05, LL04a, LT09a, LS09, LJ07, MVD04, MLSD07, MJ06, Mil08, MPFC08, NPH09, OMK09, PAD07, RAB07, RMV03, RJ04, SDR07, SHWC07, STD^{+}05, SWZ03, SPT05, SL07b, SC09a, zSW06, zS06, SXyWX09, TZ06, VPMC04, VW02, VMN07, VS07, VZSL07, Wag05, WC08, XMP07, YCL05, YSW06, YS07b, ZH09, ZSP08, ZW06, dSHHM05, dHRvdB07, PP09]. domain-decomposition [BB09b, LJ07]. domain-type [BSLN09]. domain/finite [DGMN03]. Domains [ACS00, BC01, BW01, BMQS02, CR02, GFCK02, Goe00, HJ02, LFK00, MCJ01, PR01a, AST07, AC05, ACLS03, BB08a, BP07, CGDT09, CHCOB09, DD03a, DD03b, GS07, GF05a, GLLN07, GLLN09, IDD04, ILL09, KZ06, LG09, LF05, Mad06, MM07, MG07c, NN04, PL08, SS08, SC08b, YBZ06]. dominant [Edw06, TW07]. dominated [KIH09, TB06]. dot [HLWW04, HLWW06, VTW^{+}07, Vos06]. dots [HWW07]. Double [Che00a, Che00b, CKGL02, CKG04, LS03]. double-Fourier-series [CKG04]. downwind [LWW04]. DPD [FPK08, SK06]. DPEM [LJ09a]. Drag [HGM^{+}00, MK02a, LH05b]. Drift [BMS00, BZB00, DE02, BBDE05, DGM07, ESD05, GBB^{+}06, GD07a, dFGLS05, dFJS09]. Drift-Diffusion [BMS00, DE02, BBDE05, DGM07, dFGLS05, dFJS09]. drift-kinetic [GBB^{+}06]. Drift-Wave [BZB00]. drill [CP03b]. drill-string [CP03b]. Driven [AQV02, APQ02, DGA08, EAY01, SZ01, Str01b, AK05, CBJdlC07, DTS05a, DTS05b, GZ07a, GZ08, GGP06, HKM08, MY07, MP05, ML04, OK06b, Pau07, Pop09, RWMK03, SW04a, VQLZ04, ZVQ07, ZZ09, VS09]. Driven/Time [VS09]. Drives [WB01]. Drop [CBL01, CB09, JA08, YFLS06, ZK05]. Droplet [BW02, SR00a, JS05, KH07, LKMU05, NTB07]. droplets [RGS04, SW08a, WSI08]. Drops [HLZ02, ZD00, JA08, YZF07, ZD05]. Drum [OS01]. dry [GPC07, Vol04b]. drying [SHTB09]. DSC [WZ07, Boy06, SWZ03]. DSMC [AA02, GTRB09, GMAj09, Mac01, Mac03, MY07, OC08, SL04, WLC^{+}06]. Dual [GHG01, ZTZ02, CGH05, CS09, HC08, Hua07, LJ07, MKKY06, NPH09]. dual-compact [CS09]. dual-field [LJ07]. Dual-Reciprocity [GHG01]. dual-time [Hua07]. dual-time-stepping [HC08]. duct [Ata04, DB04, HY09, HY11]. due [BBF^{+}08, Dwi08]. Duffing [LTD^{+}06]. Dust [dlFMBdlFM02]. Dusty [Sai02]. Dusty-Gas [Sai02]. Dyadic [CY00]. Dynamic [DIV00, EH02, GC02b, HF08a, MKM99, MKM04, SM06b, THN^{+}07, vdVvdV02, AZB09, BIW04, BS03b, Che04, CSKD05, DDM07, DEHL06, FDD07, FDD09a, FDD09b, Fen06, Gra06a, HBLD07, Lap03, LDN04, LKE04, LQX06, LP06b, MG05a, MY06b, PKKL05, PS03b, TLAD04, YKG04]. Dynamic/Thermodynamic [GC02b]. dynamical [AS05a, BBF^{+}08, CBJdlC07, SW08c, Thu08b]. Dynamically [CH01, Eld08a]. Dynamics [Bar02b, BSJ01, BZW01, CPP02, DPR00, DPRS01, DGA08, GK02, Hun01, LR01a, dlFMBdlFM02, Poz01a, QRHD00, SSL00, SZS01, TTSG01, TSG02, VCG03, VCTS02, WHV^{+}00, YSC01, Yon01, ZSP02, deM02, Alb08, Alb09, AWK07, ALT08, AKP07, BIW08, BLS08, BW06, BPMR08, BS04b, BBvdV06, BDCG03, BOK^{+}06, CFM09, CELS07, CJR04, CDL04, DSJ03, Dim07, DTS05a, DTS05b, DST07a, DDDC07, Eld07, ES03a, ET06, FS04, GFS08, GCCD07, GV06, GPL05, GT09c, Har04, Her05, Hew03, HS04, IAT08, JG09, KKM08, KFH^{+}04, KG09, KK05a, KLW09, KP05, LLP07, LM04, LPK05, LRS07, Ler06, Li08b, LSK06, LL06a, LLZ07, LW04, LMH07, LZH^{+}07, MGCR07, MC07a, MPD08, ML04, MK04a, NDT06, OK07a, Pal08, PGB05, PC08, Pau07, PPCW06, PK05, Pro03, RCT07, RFFP06, SKR06, Sam09, SDS07]. dynamics [SLF08, SHY07, SFVK06, SHP07, SS09c, Sto07, SC08b, TS04, TCO^{+}04, TPR05, VS09, VGZB09, VGBZ09, Vil08, VHI05, VHI06, VCM00, WGSL06, WZ03, YWC07, YHSX07, YZL^{+}06, YZF^{+}06, ZGK09, ZRS06, dWKL07, vLAvdV06, vZS07]. dynamics/continuum [JG09]. dynamo [XSG04, XSG08]. dynamos [TFD06].

Earlier [Mac00]. early [CGN^{+}07]. earthquake [BIW08]. easily [MKLU05]. Eddy [FLG01, FG02, KK00a, LLQ^{+}02, ME09, Nov04, PPC00, TSB01, AD04, BBB08, BS03b, BO04, CM03, CSKD05, DS09a, EPW08, FDD09a, FDD09b, Gra06a, Gra06b, HBLD07, HP04b, KSJ03, KDC05, LP06b, Liu09c, LDV08, MCM04, MLM09, MGS07, MBP07, MMPB07, MHdB07, NLF03, PDHP07, PYC04, PM07, RMG^{+}09, SSW^{+}07, SFMP06, TSB03, TMD07, VK09, XLP05, YB06]. eddy-current [EPW08]. Edge [MP01b, RXH02, WS01, BHvdV06, LLB05, MP08, SS05b, VTW^{+}07]. Edge-Based [WS01, SS05b]. Edge-Plasma [RXH02]. edged [YZW07]. Editorial [Ano00-29, Ano01-29, Ano02-29, Ano03l, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano06a, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03m, Ano03n, Ano03o, Ano03p, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08a].

Editorial [Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x]. EDQNM [BBB08]. Effect [LGP09, NOG08a, WB09b, de 00, LY06, PAD07]. Effective [DLD08, LM01, PSN00, CSL08, GGRS08, LM03a, LWF^{+}08, MTWW06, MCP03, PSZ09, ZC09]. effectiveness [LC06a]. Effects [LSD07, NF09, SMAj08, VLKM02, YE07, AINR03, DM03, GS05a, HC08, HS07b, ID04, LGKP07, LLZ07, MP07a, ML08, PS05, SS03a, WWC07]. Efficiency [CGMS06, RRV01, Cam03, EKP07, LDPL08, MJ06, SROCFF03, SFVK06]. Efficient [And09, AST09, BLS08, BRDM09, BCL06, BY07, BCDW06, BST01, BIVC07, Bus00, CWJ07, CH01, CFR09, Che07, CSMH05, CSV00, DH04, DDF01, DGP00, FPC^{+}00, FGOV00, GK02, HPS06a, HWL08, HBHS09, HPS^{+}06b, JD04, JW09, KS02b, KB00, KB01, KAS08, LKE04, LCB09, LMS02, NJX08b, NZZ06, Ols07, OJW06, PHW08, PA07a, PC02, RH01a, RA09, Ros07, Saf02, SHS08, Sch08, SZS01, STiST02, TK02, Tok06b, VCT09, WLT08, WZ03, XS09, Yok07, ZD00, vBRK01, AR08, BL04, BW06, BMT09, BJ09, BSLN09, BB09b, BH04, BP07, CLG07, CRAG07, CP06b, CP06c, CW08, CFGK05, CWD08, DBF08, EKBL09, Fan08, FWR07, FCGK05, GN07, GV06, GTMC08, Gri09, HNF07, HS08b, HWWL09, HDR^{+}06, IH04, JRS05, JL04b, KK05c, KK05d, KK07, Kro01, KSW03, KR09c, KS07, KLP^{+}09]. efficient [LL09, Lar07, LLRP09, LZH^{+}06, MNR07, MBP07, NG06a, NMG09, RWMK03, SPB09, SWB^{+}06, SF03, SE09, SY09a, TAL09, VPMC04, WT07b, ZL04, ZZ08]. efficiently [EKP06]. eigen [CJSS08]. eigen-oscillation [CJSS08]. eigenelements [LM08b]. eigenfunction [GKE04]. eigenfunctions [Hau08a, Hau08b]. Eigenmode [CL00b, DMG00]. eigenmodes [DD05, LL04a]. eigenpairs [GB08b, Ovt08]. Eigenproblems [Boy02b, GG00]. Eigensolution [HA02]. eigensolver [CGC^{+}09]. eigensolvers [VTM^{+}08]. Eigenvalue [AKV00, Mit00, BBD04, CC03, DL03a, NU09, SP05c]. Eigenvalues [Mit00, Hab04, Heu03, VCT07]. Eikonal [LSZZ08, QS01, CT08b, FLZ09]. Einstein [BT03, BJM03, BW06, BCL06, BS08a, CHH06, CKLS05, CLS05]. ejection [KFV07]. Elaborating [vEB05]. Elastic [Bon00, BG09, GF02, HB02, LLN00, LAS01, MC01, WP09, APT09, AK06b, BS08b, BCZ04, CLS^{+}06, DLW04, DLW06, GH09, GFS08, HMMR04, HK08c, HS08c, IQT08, LS08, LP04b, TLL^{+}08, TC07b, TC09b, XCY06, YAvdB^{+}08]. Elastic-Plastic [GF02, HB02, MC01]. Elastic-wave [BG09]. Elastically [ATV01, LN09, RRV06]. elasticity [BDRT09, Hau08a, Hau08b, MHPR08, YHCD05]. elasto [BZ04, ZVQ07]. elasto-plastic [ZVQ07]. elasto-thermo-viscoplastic [BZ04].

Elastodynamic [Gro00]. Elastodynamics [PKvdB00, CCV03]. elastoplasticity [SKS08]. Electric [CP03b, GG00, HR08, AvdB04, AINR03, GFG09, HPS06a, HF08a]. Electrical [CCT05, GKL00, HCG01, SPC01, IKL^{+}08, LWW04, SMSS07]. electrically [AL06, FH03]. electro [KK03a, Mar06, WWC07]. electro-osmotic [WWC07]. electro-static [KK03a, Mar06]. electrocardiology [GGMN^{+}09]. electrochemical [BP07]. electrodeposition [SS08, ZSB^{+}08]. electrodynamics [BS05, BS06a, MGS09]. electrohydrodynamic [TGB^{+}07, ZK05]. electrolyte [CKPW07, SXyWX09]. Electromagnetic [BEPT09, CY00, CBB01, CP07, CL00b, FM06, HAAO00, HKM07, Kan02, OMG02, PL09a, SZ01, SFW00, Vay01, VQSZ02, AKL^{+}08, AV03, Bet08, Bot06, CJSS08, CP03c, CT07, CSML06, DS09a, EG08, FCJ08a, FWR07, GH08a, Hoh06, KT06, LZL03, MRRS05, MLFG06, NCW^{+}09, OMK09, SHWC07, SC09a, VZSL07, Woo06, ZSW03, ZW05, ZH09, dSHHM05]. electromagnetics [Bér07, SCC09]. Electromigration [AIRY01, AIR03]. Electron [HK00, Saf00, Saf02, BBF^{+}08, DGM07, ED07, GS05a, GLS03, GM04, Gos04, GM06, HBHS09, JP03, KB04, PPCW06, PA05, RRC05, SMSS07, dWKL07]. electron-molecule [SMSS07]. Electronic [CWWZ00, LCG07, TR02a, BCHL07, CGL06, HBHS09, KKCF09, Küm04b, LZ07, VTM^{+}08, WBM09, YMW06]. electrons [CP03c, Hag07, Kwo08, MK04a]. Electrophoresis [vHBB02]. Electrophysiology [Ota00]. electroseismic [GH08b]. Electrostatic [BISS01, BZB00, GBS00, HF01, GS09b, KKD08, LSA06, LWW04, LCM07]. Electrostatics [GL06, AA07, GPVB07, XJ07]. electrothermomechanical [LTM09]. Element [BMR01, BW01, CHR01, CWT00, Cod01, Dur00, GM01a, GP00a, GHG01, Han00, HH02a, HCG01, LTZ02, MPP01, OMK09, PX02, SC01, Stu01, Whi00, WK01b, YMF01, ZYC02, ZH01, vdVvdV02, AH08, Ain04, AMR06, AG09, BAYZ08, BHL^{+}04, BMN05, BGN07, BGN08, BGM08, BS04b, BBvdV06, BS03b, BS04c, BT06, BHvdV06, Boy04, BEG03, CJSS08, CCG08, CLG07, CCV03, CL03a, CXZ09, CS07b, CLL^{+}07b, CS08b, CJ04, CBH03, CQRW05, CHPR09, DR06, DW09, DGMN03, Dim07, DF04, DBB06, EGHE06, EE08, ÉGP09, FHW07, FEL^{+}05, FK06, FHLO08, FD03, FWR07, FWK08, FBHV05, Fou06, GPF03, GW05, GR08, GCCD07, GR07, GV07, GLLN09, GLT07, HPS06a, HS09a, HZGB04, HZGB05, HPD09, HdGKG08, HY09, HY11, IDD04, IHL03, Jar04, JBF07, JZ08, JLL^{+}06, KCH06, KW06, KR09b, KB08]. element [Kuz06, KSS09, KS07, LFSS07, LW06, LJSM08, LCW04, LL06a, LLZ07, LSS^{+}09, LM08c, LNXNTX09, Liu09c, LY04, LJ07, MY06a, MZ08, MK08a, Mac07, Mad05, MWM03, MR06a, MGCR07, Mar09, MB04, MP03, MCN03, MDM03, NV09, NLLE06, OVG07, PP09, PR04a, PvdV08, PR03, PR04b, Pon06, Pon07a, Pon07b, PR06, QP03, RBvdV08, RRW05, RFFP06, RJ04, SB06c, STD^{+}05, SM06b, VW02, VCM00, WK05, WK04, WLT08, WZL09a, WZL09b, WGRA09, XJ07, XP04b, ZGT06, ZYL^{+}06, ZHSS09, dFJS09, vOP04, vdBG09]. element-finite [IDD04]. element-wise [CBH03]. element/finite [SS06a]. element/volume [GV07]. Elements [BT02, BS00d, CGSS00, CSP01, HL01, KT02, MP01b, PW00a, PC02, PG02b, ST01, WPH00, WB09b, AK07, BJP04, CHL09, CP04b, CGC^{+}09, CH08, GLMH09, HK08a, HMMO05, KLP^{+}09, LJW09, Mad06, NL08, Pon07b, SGG^{+}04, SG03a, VZSL07, WHS08, YZF^{+}06]. elevation [VTT08]. Elimination [JTB02, GCLB04, LY07b]. ELLAM [LDW07, LTC07, WLE^{+}00]. ellipses [DTS05b]. ellipsoidal [LKMK09]. ellipsoids [DTS05b]. Elliptic [Che00b, OKL01, SC01, Xu01a, ABLS05, APQ03, AQ07, Ber04, Bor07, BL05, CFS09, CHL09, CT04, CC03, CGDT09, CS08a, CXZ09, CS07c, DIL03, DT03, HO08a, Hel09b, Heu03, HL05, JLT03, Kau03, KS07, MY06c, MD06, Ngu07, Ngu08, NvL03, OK06a, OSK09, PSM08, PC06b, Str07a, VK05a, YBZ06, YH07a, YZW07, ZJWC08, Zho07, ZZFW06]. elliptic-hyperbolic [PC06b]. elliptic/hyperbolic [NvL03]. elliptical [MTH08]. elongated [MR07a]. elongational [DMHP07]. Embedded [MCJ01, Bet08, CD03, CGKM06, DND06, JM05, KL04, NA08, OK06a, SSN09, SBCL06, VB09, YB06]. embedded-boundary [NA08, VB09, YB06]. embedding [RM08, dHRvdB07]. Emden [PSD09]. Emitter [ST01]. Emitting [deM02, BMK^{+}06]. Empirical [CWWZ00]. emulsion [ZD08]. emulsions [KH07]. Enablers [BBI09]. enclosed [LS08]. energetic [CDS04, HS04, YFLS06]. energies [ED07]. Energy [BV00, Bar02a, BKR^{+}01, BS00d, CRB00, CBKM00a, CBKM00b, FK02, HK00, JTB02, LW01, LW04, MR03, RSSL09, SC01, SNLS03, VP00, YC09b, BT03, Bur05, CLS05, CL03a, CL05, CL08d, DSM09b, DST07a, DLW04, DLW06, DOW08, ESD05, FL06, HM09, HDBW05, HK04c, KJ09b, KPP07, KPP09, KLSW09, LRS07, LLZ07, Mai04, MY09, NG06a, Oh04, QFR04, RC09a, RSS09, Rom07, SKK^{+}08, SLG^{+}03, SHP07, SC09b, TDWY08, Vos06, YC09a, YMW06]. Energy-Conservative [CBKM00a, CBKM00b, FK02]. energy-conserving [MY09]. Energy-Preserving [LW01]. energy-stable [KPP09]. energy-transport [CL05, Rom07]. Engineering [PVR07, Lap03]. enhance [LXM09]. Enhanced [EKP07, GE07]. Enhancing [FHD^{+}09]. enlargement [ZSW07]. ENO [WH02, CS07a, EAY01, GSD01, SM04, UTBV03, VS02]. ENO-Padé [WH02]. Enriched [CWYM08, vdBG09]. Enrichment [Dys01, ZWL02]. Ensemble [LP08, LX09, ZIP06]. Enslaved [JMP02]. enstrophy [KJ09b]. Entanglement [RS02, Kar04]. entanglements [LMK03]. entries [LAKD08, ODAF06]. entropic [DGM07]. Entropy [CL01a, Gos02, RS02, Rom02, SLY02, YVD00, Abr06, Abr07, Abr09, FDK06, GG09b, HM04, IR09, Ser09, TDWY08]. Entropy-Conservative [CL01a]. entropy-consistent [IR09]. entropy-fix [Ser09]. entry [DB04, LZ09c]. envelope [HS07b]. environment [CDR09, DQA08]. environmental [RCB05]. EPI [Tok06b]. Epitaxial [CMK^{+}01, RV00, BHL^{+}04, CLS^{+}06, CF06b, HLX06, RS06b, SSE03]. Equation [ACK02, AGH00, AGH02, AQ00, BC02a, BJM02, BD01, BdLL01, BC02b, CL00a, DV02, DP00, Dys01, Eli02, FBFF00, FSB01, FP02, FSY00, GZ01, GFCK02, GF02, GR08, GC02b, Gua00, GSW00, HMS08b, IKS01, JPMC01, JP00, JM00, KK00b, Kul01, LL01a, LFK00, LW01, LAS01, MBP00, MCJ01, Mie00, OGV02, PB00, SFY01, SDD07, SKAS01, SWL00, Shy01, SKXK05, SR00b, Stu01, TRL01, Vay00, Vay01, WL02, YFS01, ZZ01, Zha02, AST07, AHPT07, AMR06, AC05, AK07, AB03, ABK09, AL08, Asl04a, Ata04, AMLC08, BAYZ08, Bal08, BJM03, BYS08, BC05, BFT07, BFT09, BFG07, BLM08, BIS07, BS03a, Boy03, Boy05a, BZ09, BL05, CHL06a, CL08a, CPG04, CR07, CLTA07, CHL06b, CCG^{+}06, CT08b, CS09, CLS09a, CJK^{+}03, CS04, COR08, CDL04, CDL05, CkM07, Cui09, Dar02]. equation [DBF08, DSJ03, Dem04, Den07, DWLM09, DNS08, DC08, Dur08, EB06, FL09, FT05, FR03, FK09b, FH07, FF03, FLZ09, GS06b, GPVB07, GKL03, GS09c, GD06b, HPS06a, HK06, HKM08, Had05, HMOG08, HZ08, HMS08a, pHL09, HHPW08, HMPR07, HKG08, HL07a, HW05, Hel09b, HHC08, HH07c, HL05, HWWL09, HLX06, IKS^{+}09, JLOT05b, JW06, JY08, JM05, JKL05, KMS03, KW06, KNH05, KMS04, KR09c, LT05, LH05a, LLY05, LP07a, LL03b, LL05, LTE07, LTZ03, LG03b, LG04, LG07, LGK06, LX07a, LT09b, LY06, LS09, LC03, Mac07, Mai03, MRC06, MCGV04, MKKY06, MST06, MP03, MGC06, ML06b, MK07, MG05b, NV09, Nas08, Nis07, NXS07, PC08, PHW08, PIN09, Pon05, PPB09, Pud06, QA09, RSSL09, RKE^{+}07, RBK09, SFDL07, SBCL06, SCT09, SCT06, SB09, SY09b, ST04]. equation [SS07c, zSW06, zS06, SB07, TMS06, TM07, Tau07, Ten03, TW05, Tok06a, TR07, Tuc03, UL06, VWW04, Wan04a, WT07b, WXG07, WZ09, WKG06, WKL07, XSG04, XSG08, YBZ06, YA05, ZXQX08, ZKL^{+}07, dSM05, dHRvdB07]. equation-based [Tuc03]. Equation-free [SKXK05]. Equation-free/Galerkin-free [SKXK05]. Equations [AGT02, ABGV02, Asl01, ACS00, BL09a, BC01, BCOS01, BT02, BBR01, BCVK02, BM01c, BZW01, BS00e, BCM01, Cal02, Car02, Che00a, Che00b, CL01c, CKL00, CL02, DMG00, DC01, DKK^{+}02, DDH01, DF00b, DK02b, Dur00, FF02, FMO00, FR02, GTD00, Gir00, GHW02, GBGM01, Han01, HH02a, HH01, HDC02, HH02b, HW02, HF01, Hu01, Hua01a, HK01, IFZ01, JMK01, KLN^{+}01, KR02, KM00, KB01, KT00a, KT00b, LBV00, LBV01, LTK^{+}02, Lay02, LLIK01a, LLIK01b, LOK01, LC01, LL01b, Li01, Lin01, LMS02, MPP01, MR02, Mai01, MCCT02, MC00a, MF01, MF00, MG02, MLS01, MOvL00, MPC01, MPC02, Myo01, NTYT01, NTYT02, Nys02, Pai01, Pet01, Rei00, RB02, SSD00, TWS02, VB00, VS02, WDM01, WPH00, WZ02, WK01b, WA02, WS01, XCZ02, XK01, Xu01c]. Equations [Xu01a, Xu02b, YP01, ZYC02, ZCMI01, ZS01, ZDNP00, AS03a, AvdB04, APR09, APT09, AKV06, AGT05, AB07, AMXL09, AEP04, And09, AI09, ACLS03, AG08, BQQ09, BR09a, BFB08, BTW04, BLW04, BY07, BGN07, BG07, BCDR06, BV05, BB07a, BACFT05, BES07, BFG08, BRC^{+}09, BTT08, Ber04, BK08, BF07, BYZ04, Boe05, BB07b, BT06, BRP05, BJ09, BT07a, BHvdV06, Bot06, BS06b, BGLN05, BEPT09, BL03, CD03, CHH06, CVB06, CP03a, CHL09, CQO04, COQ06, CS05, CCJ07, Cha07a, Cha07b, CWL08, CC07, CTT08, CS07b, CS07a, CSL08, Chu09, CJ07, CS03, CLS04, CY05, CT07, CFP06, CFJ06, CZVS04, CFP08, DR06, DJM05, DDSV09, Del03a, Del03b, DGH08, DR09a, DD09, DH07, Doh09, DLP08, Dom08, DI09, DJTT05, DD03a, DD03b, DZ09b, DOW08, EHST03, EHS^{+}08]. equations [EG08, ES06, EN06, ES03b, ELW04, FCJ08b, FYH^{+}06, FL03, FOLD05, FD07, Fox09, FKLY07, FG06, FMR09, FH03, GPC07, GB08a, GS07, GK03, Gel06, Geo08, GSV06, GFR09, GF05a, GCNB07, GPF03, GW05, Gir06, GR08, GTMC08, GKE04, GR04, GBS06, Gri09, GE07, GLLN07, GD07b, GD05, GK05, HH07a, HLS06, HK04a, HR08, HMM08, HH08, HK08a, HC08, HO08a, HLO08, HJJ09, HLMM07, HL06b, HyLL07, HJL09, HLRZ06, HS08b, HS08c, Hu05, HLL08, HJM07, HMR08, HMM07, HC05, IX07, IM07, IHL03, IQ08, ILL09, JBF07, JR03, JR04, JHSZ07, JMC03, KE06, KOQ04, KOQ08, KSH^{+}06, KDK^{+}07, KP04, KG08, KJ09b, KAK03, KL04, KvdVvdV06a, KvdVvdV06b, KvRvdVvdV07, KL06, KS09, KT03, KT05, KG03, KLM05, KN04, KQW03a, KQW03b, KLSW09, LY07a, Lar07, LHD05, LGHD08, LS03]. equations [LMN^{+}09, LM08a, Lee05, Lee09, LFS07, LQ09, LSY04, LP04a, LSZZ08, LLOT06, LDPL08, LYC09, LMS04, LSSV07, LSV09, LRZ04, LW07, LP07b, LX07b, LW09, LCS09, Liu09b, Liv07, LGM08, LZC04, LMNK07, LB04, LBL06b, LCdCN^{+}03, MZ09, MJ09a, Mai04, MVD04, MP07a, MCG08, MR05, MY09, MSS08, Maz06, MHB08, Mea04, MK08b, MSO04, MOG09, MT04, MG06, MD06, MPFC08, Moo03, Moo07, MAN^{+}06, MT07b, MSB07b, Mou04, MDR07, NW07, NZ05, NOG08a, NFGK07, Ngu07, Ngu08, NPC09a, NPC09b, Ni09, NI03, Nik06, NMS07, NGvdWS09, OS04, OK06a, OSK09, OX04, ORM06, Ols09, OK06c, ON08, PAD07, PSD09, PNMK09, PvdV08, PCS^{+}09, PR03, PR04b, Pop03, PR06, PGN08, QS05, RC09a, Rah04, RCT07, RBH03, Rem06, RS06a, RBvdV08, RAD07, RRW05, RSO04, RM08, SSN09, Sac07, SS03a]. equations [SFDL07, SZB^{+}07, SDM04, SM06a, Ser09, SS08, SRNV07, SFE07, SMB09, SWZ03, SP04, SL07b, STZ07, SHTB09, SP05b, Soc03, SKW05, SG03b, ST03a, SCN07, SN08, STR07b, TTZ03, TL06, THL06, TBT^{+}09, Thö04, TXCD07, TKH09, TOY09, TdAAP08, TE04, TG08, TS08, Tsy04, TFDK04, VPMC04, VVM05, VW02, VSW04, VSW06, VTT08, VSG05, VK05b, WK07, WK05, WRu03, WM07, WGS^{+}08, WM09, Wel07, WKB07, WZ03, XXS07, XS05a, XS09, XS05b, XHW07, XD07, You06, YZW05, YS07c, YS08, YLA08, YE05, Yus06, ZYL^{+}06, ZW04, Zhe06, Zhe07, Zho07, ZZFW06, ZT07b, dDEK09]. Equidistribution [BMRS01, Hua01b, DCF^{+}08]. equilateral [MR03]. Equilibria [BBG^{+}02, SHWW00]. Equilibrium [BKR^{+}01, Cle00, DSS00, AZ05, BP03, BSW05, CGP05, CP07, CK08, DL04, GT09a, GT05, GWF^{+}07, GLN06, JSCZ08, JG09, MKR00, MK03, ORM06, SSB07, WSYS09, XHC08]. equivalent [LM08a, RE07]. Erratum [ABRR09b, CL08b, DKX01, DD03a, HMS08b, HT00b, HY11, HLWW06, JJGL07, Lau06, Mil07, NTYT02, PW01, Tol02a, WZL09b]. Error [Bar02a, CHR01, FWK08, GKD09, KLN^{+}01, LMK09, OV00, OP02, SDS07, ST03b, VD00, Yam01, CLMRP08, CC07, CY05, DL03b, Dur08, Dwi08, HGBH03, HNGB04, HS03b, KK09, KKO04, Lap04, MGS07, MK04b, Ngu07, PG04, RS09b, SVH^{+}06, TWM07, WK06, OV00]. error-assessment [MGS07]. Errors [FLG01, BBB08, CP06a, CM03, DL03b, GD06a, KLM07, PYC04, PM07, Vil08, VK09]. Essential [APQ03]. Essentially [Abg06, BS00a, WC01, WH02, BCCD08, CL06a, HAP05, TWM07, ZSWW03, ZWS06]. estimate [WK06]. Estimates [MP01b, OV00, HS03b, PG04]. Estimating [KS02a, KFIG06, RS09b]. Estimation [BCEG07, OV00, OP02, RMV03, VD00, BS03b, CFS09, DLD^{+}06, Dwi08, HMA05, IKL^{+}08, KK09, LJSM08, Ler06, MDJS07, Ngu07, PM08, Sti05, TPVG06, Zad08, vdDA06]. Estimator [TS01, LZ07]. estimators [SVH^{+}06]. ethylene [GIA^{+}07, GIA^{+}08]. Eu [Myo01]. Euler [TR07, AEP04, AI09, Asl04b, AG08, BZW01, Car02, CS07a, CR09, CDV07, DDK06, DMG00, DDSV09, Del03a, DOW08, GB08a, GR04, Han01, HH02a, Hu01, Hu05, HLL08, HK01, IX07, IR09, JR07, JK00, KL04, KQW03a, KQW03b, Lee05, LFS07, LW07, LCS09, LBL06b, MC00a, MOG09, MG02, MSB07b, Nat06, NOG08a, PvdV08, Pop03, PGN08, RC09a, Rah04, RBH03, RS06a, SFDL07, ST03a, TL06, TW05, VSW06, WM07, WM09, WKB07, WZ03, ZYC02]. Eulerian [AEP04, ALGM01, AV02, AHMS03, BALW06, BS00c, BR09b, CW03, DVHM05, Fed02, FLM08, FKK08, GT09b, GXW07, Her05, HG03, HPZ01, HH06, JX07, KMSH08, KMS02, LMV04, LL03b, LQ09, LHZW05, LY04, LS05b, MC01, MC02, MCN03, OF02, OCK^{+}02, QL04, RB05, RWWS07, SM09a, SCW^{+}09, SFW00, UTBV03, YA05, YFBH07]. Eulerian-Grid-Based [AV02]. Eulerian/Lagrangian [GXW07]. Evaluating [GHG01, RS09b]. Evaluation [GST02, Hau08a, Hau08b, KMJ01, LWEM00, MT04, PC02, RMG^{+}09, Saf02, ABZ^{+}08, BHS09, BO04, CRAG07, DMBS05, FT06, HO08b, KR09c, Lau04, Mar06, MG07a, PC08, VOD08, VB08, VS07, WG08]. Evanescent [BV00, BPO07]. evaporating [AJ09, SW08a]. Evaporation [HW08, LMS05, SS06b]. even [CTS07, RVM07]. Event [DGA08, ML04, OK06b, NZ07, Pau07, PA07a, ZZ09, VS09]. Event-Driven [DGA08, ML04, OK06b, ZZ09, VS09]. Event-Driven/Time-Driven [VS09]. events [MS08b]. Evidence [SS05c, BBCT09]. Evolution [ATV01, AGH00, DC01, JWSC00, JW02, LLN00, LMSW02, Nie01, Set01, SR00b, AKLMP09, AINR03, BGN07, Bey09, COQ06, CS08b, CP04b, CP05, DDD05, EN06, FM06, GFG09, JW03, KT05, KLM05, KN04, LZ09a, LMNK07, Ols09, RRV06, RSS09, SR09b, VWW04, Wan04a, ZZ09]. evolutionary [DGRS08]. Evolving [CBL01, ML05, BGN08, EE08]. EVP [Hun01]. Ewald [CWJ07, OJW06]. Exact [BDRT09, BTT08, CGP02, EZ08b, Fou06, LBD02, NN04, Ten03, UH01, VS02, Zhe06, BCZ04, BDCG03, DP07, Lau06, RS09b, Wag05, Wal03, XS05a, vZS07]. Exactly [NTYT01, NTYT02, FS09]. Example [PL01]. examples [Ram06]. exceeding [KWD07]. exceptional [LC06b]. Exchange [RH01b]. exchanges [PFSL07]. excitation [FK09b]. excitation-adaptive [FK09b].

Excited [FV01, BCL06, CGL06, Fra04]. excluded [LGP09]. execution [BDS07]. exercise [Kou07]. ExGA [MLSD07]. expanding [HDBW05, PK07]. Expansion [ADK02, Goe00, SSD00, VP00, AV03, BO05, GMH06, KYLB07, NCS03, PW07, Tak06, TPV07]. expansion- [NCS03]. Expansions [CL02, BRB03, FT05, GG09a, HLRZ06, LKNG04, SH07b, Tyg08, VS07, ZL04]. experience [BPS03]. experiment [OB06]. experimental [DGF09, NDG05, ZGSD06]. experiments [FHJK09, SSW^{+}07, XSG08]. Explicit [DZ09a, GSD01, HK06, HJL09, KM06, KM07b, Kuz09, MLSD07, PH09, QM03, RB02, VG01, VCTS02, XCZ02, YP01, AHNS09, BBMB07, BB04b, BB07b, CSL08, DR06, FDD09a, FG07, GL09b, HR08, Jia07, JL04b, KCGH07, Lar07, Löh04, LGM08, MGS09, OK07a, PH06, PH08, Sha05, zS06, TDGP06, WG08, XJ07, ZSP08]. Explicit-Implicit [RB02, DR06, HR08]. explicit/implicit [TDGP06]. Exploring [Lou00]. explosion [WK04]. explosions [KS08a]. Exponential [CM02, BIS07, Kry04, Liv07, Sti05, TWYC06, TD07, Tok06b]. expressed [NG06a]. expression [Sha05]. expressions [OLA08]. Extended [BMS00, Vay01, FHLK05, GR07, HSZ04, KFH^{+}04, KFIG06, KLK08, Nas08, WLKW07, ZGT06]. Extending [CDJ07, DC07, WS04]. Extension [Boy02a, FM08, LVW06b, NBLQ09, Pop00, SBGK00, SWL06, WL06, CB03, MBS03, WZL04, WL02]. extensional [AH08]. Extensions [HMOG08, HR07]. Exterior [MM01, ABK09, Bor07, MG07a]. External [FGOV00, HAS05, HHC08]. Extracting [WC08]. extraction [Hum05]. Extrapolation [TK00, Asl04a, CHL09, GS03a, RB06, WZ09]. extrema [CS08c, RGK07]. extrema-preserving [RGK07]. extremal [GB08b, Ham07]. Extreme [VSMW01, CS05, FGS09, MS08b, SK08a]. extremely [TAL09]. Extrinsic [KYLB07].

F [LM03a]. Faber [BS06a]. Face [Jia07, SS09b]. facing [RFVP09]. factor [Kro05, NWZL08]. Factored [Boy02b, FLZ09]. factors [RMV03]. Failure [TWS02, EV03, PW07]. falling [GMD03]. Families [Tol08, Nic09]. family [BB04b, Boy03, CGSR08, EZ08a, FDD09b, MLSD07, RC09a, SK06, ZYHS07]. Far [BBW06, Fre00, HSK00, MP01b, HKS09, SCN07]. Far-Field [Fre00, HSK00, MP01b, BBW06, HKS09, SCN07]. fashion [kM07a]. Fast [AS05a, BKM09, BD01, BO04, BK01, CY00, CdHST08, Dar00b, DFT01, EN06, FLZ09, GD06b, GD08, HAAO00, HHCL01, HKG08, Hoh06, KKD08, LDN04, LJSM08, LG07, LBS^{+}04, LQX06, LWEM00, LCCG05, MTV08, Mac01, Mar06, MS01, NL08, PRT00, SHWW00, SDM04, Set01, SSC00, Str00, Str01a, TB09, Tyg08, VS07, WJV07, WA02, YC06a, ZS01, AL06, AB05a, Bar04, Bia03, BYZ04, BPO07, BOK^{+}06, CDJ07, COQ06, CR05, CHL06b, CCG^{+}06, CJK^{+}03, CWD08, CFR08, Dar02, DH04, DC07, EHD08, FD09a, GvH06, GG09a, GKD09, GV06, GH02, HSC09, HB05b, KOQ08, KP04, KP05, LT05, LGKP07, LSZZ08, LC03, LCM07, MR05, MR07a, MR06b, NI03, OLLL03, PSH^{+}08, RVVL09, SS09a, SWG08, ST06, SH07b, SP05a, SMP09, SK04b, Tau07, Thö04, TC09b, TG08, VOD08, VBJ08a, VBJ08b]. fast [VB08, WK06, YBS06, YBZ04, Yin06, ZT07a, ZKL^{+}07]. Faster [Hel09a, BPO07]. faulting [BIW04]. FCT [BHS09, Kuz09, LOK05]. FD [SL07a, Yu05b]. FDFD [CBB01]. FDTD [Bet08, CFJ06, DS05b, FK09b, POS00, Rem06, RK07, Vay01, Wel07, WC07, XCZ02, ZSW07, ZW04, ZT07b]. FDTD-compatible [RK07]. FDTD-methods [ZSW07]. FE [AT09, BFG08, MK04b]. FE-simulation [MK04b]. features [TZ06, TMD^{+}08]. Fekete [BCEG07, BCE^{+}09]. Feller [LLTA07]. FEM [BB07a, DHOT09, GQ00, KT04, Kuz09, LNXNTX09, MMS04, MR04, NZ05, YM07]. FEM-multigrid [DHOT09]. FEM-simulation [MR04]. FENE [LC03]. Fermion [BTFY01]. fermions [Bor03]. ferromagnetic [GCW07]. FETD [CL07b]. FETI [LJ09a]. Few [GHV00, HKS09]. Few-Cycle [GHV00]. Fewer [TRL01]. Fey [Noe00]. Feynman [BLL03, HvHHS05, PWW00, SS01b]. FFT [CXB08, DBF08, LG05, PKvdB00, YAvdB^{+}08]. FFT-based [CXB08]. FHNC [Hof04]. fiber [TG06]. Fibers [BV00, TS04]. Fictitious [BCM09, GPH^{+}01, WT07a, BG05a, DGMN03, RAB07, SPT05, VMN07, YSW06, YS07b, ZW06, PP09]. fictitious-domain [ZW06]. fidelity [NT07]. Field [BISS01, Fre00, FGOV00, FV01, GW02, GKL00, HSK00, MR02, MP01b, OMG02, ST01, SSW01, AINR03, BCB03, Bae03, BCDW06, BBW06, BEA09, BJP04, CFM09, Cha09, CW08, CEL06, DDSV09, DLW04, FYH^{+}06, GFG09, HW08, HvHHS05, HKS09, HF08a, HX05, HWWL09, JOS06, KKS07, LCG07, LCB09, LW07, LJ07, LH05b, MZ07, MSP^{+}06, NDG05, NLLE06, OK07b, PH09, RJM07, SY09a, Shi07, SG03a, SB07, SS04, SCN07, TLK07, TBT^{+}09, XMT06, YFLS06, YHCD05, YZF^{+}06, YZF07, ZDD09, ZSTC06, dWKL07]. field-space-based [LW07]. Fields [DPCV02, GG00, GC02a, KMHR00, LWEM00, MN02, POS00, AV03, Bal09, DGF09, DC07, DS05b, FCJ08a, GFG09, KB04, OLLL03, SR09a, ST06, TET09, TXCD07, VOD08, VS07, XDC09, ZSW03, ZW05]. Fierz [MBM01]. fifth [GR04, HAP06, SM04, Tol07]. fifth-order [GR04, SM04]. Figures [DSS00]. Filament [ZP02]. filamentary [PSCB08]. Filamentation [DGH02]. filaments [HSS07]. filling [GS03b, Vol04a]. Film [CMK^{+}01, DK02a, ZP02, GMD03, HKM08, MO06, NTB07, SA06, SRX07]. films [AIR03, ES03a, RRV06]. Filter [CKGL02, PX02, PR01b, TR02a, HO03, IKL^{+}08, KFH^{+}04, KFIG06, KSJ03, KDC05, LX09, RMSB09, WC08, YS07a]. Filter-Based [PX02]. Filter-Diagonalization [TR02a, WC08]. filtered [MP07b, ZSTC06]. Filtering [FBFF00, VCT07, BS03b, BB07b, BdCB09, CHM08, CKG04, ES03b, HM08, HV03, KCH06, Wea09]. Filters [AA02, GSD01, MVM02, ZW03]. FIND [LAKD08]. Finding [FGOV00, FV01, LY07b]. Fine [KM02]. FINESSE [BBG^{+}02]. fingering [LLL07]. Finite [AC00, ACY00, AE03, AKLMP09, BR09a, BC02a, BHL^{+}04, BS00b, BMR01, BM01b, BT02, BW01, BS00d, Bla00, BP03, CL00a, CS01a, CP00, CHR01, CBB01, CGSS00, Cod01, DPCV02, DBB06, DF00b, DET08, Dur00, FVOMY00, FHW07, FK06, FK07b, FK02, GHV00, GW01, Gro06, Gro07, GH08b, GV07, GLT07, HLS02b, Han00, HH02a, Her00, IK07, JL02, JM00, KC00, KKC01, KT02, KMJ01, LLH02, LW06, LM01, LTZ02, LM03a, LX07a, LMSW02, MPP01, MP01a, MP02, MF01, ML01a, MST06, Nic00, Nik06, NC01, OGV02, OMK09, PR01a, PL09b, PKP01, PS04, PK00, PSG05, POS00, PL07, QP03, Rem00, SBGK00, ST01, SC09a, Sou09, Stu01, TK00, TJ09, TT04, Tol02a, Tol02b, Tow09b, VCP00, Vas00, VCTS02, VG02, WPW02, Wan02, WL02, Whi00, WA02]. Finite [WB09b, YP01, ZZ01, ZH01, ZRR00, vdVvdV02, APTJ^{+}04, APP^{+}07, AH08, Ain04, AMR06, AK07, AG09, AT05a, AT08, AKO09, BAYZ08, BS04a, BMN05, BGN07, BGN08, BG07, BGM08, BAFL09, BES07, BBvdV06, Boe05, BT06, BT07a, BT07b, BHvdV06, BKLL04, Boy06, BLM04, BEG03, BJP04, CT09, CHH06, CCG08, CLG07, CHL09, CdHST08, CL03a, CN05, CX08, CXZ09, CS07b, CS08b, CEH09, CYS06, CJ04, CS06, CS07d, CSKD05, CR09, CQRW05, CSML06, CGC^{+}09, CHPR09, CZVS04, CFP08, Cui09, CH08, DSM09a, DSM09b, De 04, DR06, DBBP08, DW09, DGMN03, DS06a, DF04, DK07, DKTT07, DBTM08, Dur08, Dwi08, Edw06, EZ08a, EGHE06, EE08, ÉGP09, FP08a, FDD09a, FDD09b, FWR07, FMR09, GPC07, GLM07, GR07, GLLN09, GL09b, GL08, HPS06a, HBHJ08, HJ09, HZGB04]. finite [HZGB05, HPD09, HLO08, Her09, Hew03, HMMO05, HY09, HY11, HWWL09, IX07, IX09, IM05, IDD04, IQ08, JD09, Jar04, JBF07, JLT03, JLT06, JL09, JZ08, JLL^{+}06, JAK05, JM05, Jon05, Jor07, KW06, KTD03, Kim07, KPB08, KLK08, Kok09, KR09b, KPP07, KPP09, Kum04a, Kuz06, KSS09, LSB04, LFSS07, LZ07, LVL05, LJSM08, LZT09, LCW04, LL06a, LLZ07, LYC09, LSS^{+}09, LMS04, LSS06, LSSV07, LSV09, LVW06b, LLTA07, LM08c, LNXNTX09, LS09, Liu09c, LY04, LJ07, LMNK07, LJ06, LZH^{+}07, MY06a, MZ08, MK08a, Mac07, Mad05, MWM03, Mad06, Mai04, MLSD07, MP07a, MR06a, MGCR07, MB04, MN04, MN06, MN17, MSJ07, MZ07, MGS07, MSP^{+}06, MGC06, MCN03, MCP03, MR07c, MDM03, MVO04, MT07b, MK07, NV09, NOG08b, NBLQ09, NLLE06, NPPN06, NXS07, OK06a, OSK09, OVG07, PAD07]. finite [PP09, PH06, PH08, PvdV08, PS08, Pir07, PR03, PR04b, Pon06, PR06, RB06, RJ06, RBvdV08, RRW05, Rom07, Ros09, RJ04, SROCdPFF05, SKS08, SJD05, SHA08, SHWC07, STD^{+}05, SMS04, SYG06, SS06a, SS05a, SL07b, SZ05, STZ07, SMAj08, SS03b, SS05c, SGG^{+}04, SB03, SC09b, SWL06, SR09b, SCN07, SN08, TVMR03, Tan08, TMND07, TD07, TdAAP08, Tor03, TA06, Tow08, Tow09a, TAL09, VPMC04, VW02, VSW04, VSW06, VZSL07, WZL04, WTL08, WLT08, WZL09a, WZL09b, WZ07, WA08, WHS08, WF06, XS05a, XS06, XCRX08, XS05c, XLP05, YMWM06, YS07c, YS08, YZF^{+}06, Yus06, YH07b, ZGT06, ZZ07, ZFM08, ZH09, ZHSS09, dSHHM05, dVGLM09, dFJS09, vDZ06, vdBG09, Lab09]. finite-band [Dur08]. Finite-Difference [AC00, ACY00, CBB01, FVOMY00, GHV00, MF01, Nic00, Rem00, VCP00, VCTS02, VG02, WA02, GH08b, IK07, Nik06, BG07, CdHST08, CYS06, HWWL09, JAK05, KPP07, KPP09, LS09, MSP^{+}06, PAD07, PH06, Pir07, RB06, SROCdPFF05, SHWC07, SS05a, VPMC04, YMWM06, ZH09, dSHHM05]. Finite-Difference-Diffusion-Monte-Carlo [PK00]. finite-differences [Kum04a]. Finite-Element [BW01, CHR01, MPP01, LW06, CQRW05, DR06, HPS06a, HPD09, LJ07, SS06a]. finite-element/finite-volume [SS06a]. finite-energy [Mai04]. finite-frequency [TMND07]. Finite-Volume [BM01b, CL00a, DPCV02, KKC01, ML01a, OGV02, PS04, PL07, TT04, Edw06, EZ08a, HBHJ08, HJ09, JLT03, JLT06, JL09, Kok09, LZT09, LJ06, MSJ07, MGS07, RJ06, Ros09, SS06a, SL07b, SC09b, XCRX08, ZH04]. FINT [LJSM08]. First [Ano05s, Boy02a, CR09, FV01, HH07b, HMMR04, BEE06, BCL06, CkM07, Gro06, Gro07, Gui03, IM07, Jar04, NPH09, Nis07, YZW05]. First- [CR09, YZW05]. first-derivative [Jar04]. First-order [HH07b, HMMR04, IM07, NPH09, Nis07]. fit [AMSZ07, Mil05, Mil06, Mil07]. fitted [PS08, SS03a, YP06, ZKDT07]. Fitting [CVE06, Che04, Sur05, TWYC06]. Five [ACK02, MG05b, QA09]. Five-Equation [ACK02, QA09]. fix [Asl04b, HK04c, Ser09]. Fixed [RMO00, YSC01, AFGM07, BL09a, BBHM09, CHCOB09, DMHP07, Mad06, MS08a, MLS^{+}05, TZ06, TLK09, ZFM08, dFJS09]. fixed-domain [TZ06]. Fixed-Grid [YSC01]. fixed-mesh [CHCOB09, ZFM08]. fixed/moving [TLK09]. Flame [CAL00, NFK01, BDR^{+}04, BDGL05, LLC06, MR04]. flames [MR04]. Flapping [ZP02]. FLAPW [YMF01]. Flexible [ZP02, Alb09, EKBL09, HSS07, Hum05, LKP06, Mad05, TS04, Tsu06, XYK05]. flexible-body [Yu05b]. flexible-cycle [XYK05]. flexible-order [EKBL09]. flight [Liu09a]. flights [Pav07]. flip [ABRR09a, ABRR09b]. floating [YM07]. Floquet [DK06, TB00a]. Flow [APQ02, BBG^{+}02, BW01, BCVK02, CFA01, CS00, ČPT01, CGSS00, CR02, CL02, DIV00, EF02, FVOMY00, FGG01, GPH^{+}01, Goe00, HLS02b, Han01, HGM^{+}00, JML^{+}01, JL02, KC00, KKC01, KLvBvL02, LLH02, LS02b, LKNG01, LRN^{+}02, LK01, Mac01, MN02, MK02b, MC01, MD01, Pai01, PR00, PG02a, PS01, PW00b, PW01, Poz01a, PO01, SBGK00, SS02, SJ02, Shy01, Sie00, Sum00, TC01b, TBE^{+}01, VD00, Xu01a, Xu02a, ZTZ02, dSAK00, vBRK01, AM03, APTJ^{+}04, AH08, AK06a, ART04, Alb08, AK05, AW04, AMP09, AT05b, BSKH07, BKST09, BS08b, BF08, Bil05, BF07, BB04b, BSLN09, BLM03, BLM04, BTW03, BP08, BP04a, BIVC07, BGN03, BK07, BB09c, CLB08, CFF07, CRAG07, CR05, CFM09, CS05, CL07a, CL08b, CT08b, CMP07, CHBS04, CHPR09, CBS05, CZVS04]. flow [CGM07, DMHP07, DS05a, DHOT09, DM03, DDK06, DT04, DVHM05, DP07, DP08, DKS^{+}03, Dim07, DS06b, DF04, DND06, DCK08, EGHE06, ECL02, ELD08b, ÉGP09, ES03b, FRS08, FS04, FK09a, FT06, FK07a, FL06, FCT07, GZ09, GMD03, GH09, GGF03, GS03b, GGP06, GC06, GGCC09, GD07a, HJ09, HL04, HP04a, HS03a, HS06, Her08, HN03, HY09, HY11, Hu05, HAI09, HT03, HLX06, Hua07, HSS07, HLY09, HH06, ID04, IK07, JLT03, JD04, JLT06, JOS06, JX07, JC06b, JP03, JS05, Kel05, KC06, KDF07, KAS06, Kok09, KSGF09, KT07, LTZ03, LFX05, LK09, LJW09, LKX04, LX07b, LY04, LR03, Löh04, LM03b, LSW06, LJ06, LMZ^{+}08, MC04, MTV08, MPD03, Mac03, MT07a, MJT06, MSJ07, MP05, ME09, MT03, MCN03, MSB07a, MVO04, MDS03, Mou04, MDR07]. flow [MGNB09, MG05b, NL08, NBLQ09, NJX08b, NJX09, OK05, OKZ07, PPDM08, PP09, PS05, PWM06, PA07b, QLK07, QLS09, RC06, RFVP09, RM07, RW03, SM09a, SWG08, SGFL09, SWK06, SMS08, SLF08, SE04, SZS03, Shi07, SLC07, Shy04, Shy06, SS05c, SCRL08, SWL06, SRX07, TZ03, TOZP03, TM05, TBJ^{+}09, TT06c, TCM05, TDV06, TF03, TJLT08, VC03, VLB09, VV03, VQLZ04, WSYS09, WFC09, WGNT06, Xia04, XK03, XH03, XMT05, Xu08, XHC08, YP06, YYT05, YC06a, YC06b, YXLF05, YKK08, YF09, YE05, Zad08, ZSWW03, ZL04, ZKY05, ZWS06, ZVQ07, ZJ09, ZKS^{+}09, ZFM08, ZLAC05, ZL08b, ZD08, vOP04, vdV08]. flow-body [Alb08]. flow-induced [SCRL08]. flow-polymer [CFM09]. flow-structure [LMZ^{+}08, ZFM08]. flow/structure [AK06a]. flowfield [KK05b]. Flowing [ZP02, CGL08]. Flows [BSJ01, BM01b, BMQS02, BL01, Bon00, CKR00, CKR01, CPK02, Cle00, Cod01, Cor00, CMOV02, DCV^{+}01, DK02a, DF00a, DLS^{+}00, EAY01, FS01, FG02, GSD01, GS02, GHG01, GW01, GM01b, GQ00, Hor02, JLCD01, JPMC01, KKR01a, KKR01b, LX00, LS00, MSYL00, MPC01, MPC02, Nic00, Nie01, PWS^{+}02, PSN00, QV01, RH01b, Ros00, Sai02, SML02, SBGK00, SSD00, Sni01, Sun00, SB02, SP00, TSB01, TCM^{+}00, UMRK01, VD02, VLKM02, VC00, WPM02a, WK01a, WPW02, WLE^{+}00, WW00, WZ00, Xu01b, Yua02, vdVvdV02, APP^{+}07, AK09, Ano04z, AMP09, AB05b, AMS03, BFB08, BHS09, BM06, BDHN09, BGM08, BALW06, BH09, BS04c, BCM09, BPL06, BKLL04, BN09, BCI^{+}08, BHSV07, BB08b, CPR05, CCG08, CGL08, CET09, CGRGV^{+}04, CFL^{+}03, CHB09, Che03, CGH05, CJ09, CYS06, CZ09, COER07, CDE06]. flows [CJR04, CHCOB09, CL03b, CMR08, CP04c, DSM09a, DBBP08, DFV08, DP09, DGMN03, DS06a, DSS07, DS09a, DDS09, DGJ03, DHM07, DBS06, FP08a, FPK08, FM05, FD03, FPT05, FL07, FD09b, FGP08, Fox08, GV08, GT09b, GXW07, GBC06, GLL03, GMD07, GSB03, GS05c, GMAj09, GS03c, Gra06a, Gra06b, Gre04, GMO04, GMS06, GAC^{+}09, GKV09, GR07, GS03d, GS09c, GLLX08, HW08, HSQ03, HAS05, HPD09, Hel05, HS08a, Her08, HM05, HVAC09, HK04c, HKAH06, HA06, HT03, HO03, IOTK04, JD09, JL04a, Jao07, JX06, JLL^{+}06, KSO^{+}05, KR09a, KHdT^{+}08, KIH09, KIHM09, KM06, KM07b, KK05c, KK05d, KPB08, KAA^{+}07, Kro01, Kro02, LTH08, LL09, LG09, LSL08, LKP06, LK07, LL05, LP06a, LZ04, LS07, Li08a, LZ09c, LLS09, LL06a, LF05, LMS08, LW04, LKO05, LV07, LCNR07, LTC07]. flows [LM08c, LKMK09, LDV08, LD09b, LHGF05, LF04, LC03, LB03b, LBL04, LBL06a, LBL08, LMK09, LZH^{+}06, MM09, Mai09b, Mai09a, MLM09, MEKS03, MKOW04, MR06a, MM03, Mar09, MB04, MY07, MJ09b, MDB^{+}08, MLS^{+}05, MK06, MT08, MAL09, Myo04, NSC09, NOG08b, Neo07, NMG09, NMM^{+}07, NMH^{+}07, NJX08a, NCS03, NPPN06, NS05, NT07, Nov04, OF06, OTCM08, OVG07, OCFF08, PKD07, PDHP07, PPDM08, PSCQ03, PSC^{+}06, PN03, PH08, PFSL07, PK07, Pon09, Pon06, Pon07a, Pon07b, Pop09, Pro05, Pro07, PS03b, PS07d, QA09, QS07, QP03, QM03, RB05, RMB07, RVM07, RVDM09, RWMK03, RJ06, RBS06, RMG^{+}09, RF06, RMF08, Ros03, Ros07, RFVP09, SDGX07, SNGAS04, SROCFF03, SROCdPFF05, SC08a, SFDL07, SPB09, SAK05, SSB07, SE09, SS07a, SD05a, SD05b, SP05a, SFX03, SMS04, SY09a]. flows [SZC09, SS06a, SYC09, SSND03, SKXK05, SY03, Spe05, SK07a, SP06a, SZH07, SC09b, SFMP06, SBC04, SSH^{+}07, SK03, TZL05, TLK07, TLL^{+}08, TLK09, TJS03, TB06, TMB07, TSG^{+}06, TT09, TMD^{+}08, TC09a, TdAAP08, TW07, TS04, TS08, Uhl05, Utn08, VCT07, VCT09, VVS08, VMN07, VGBZ09, VD03, Vik03, VBL03, Vol04b, WT07a, WAO^{+}04, WWC07, WB09a, WSI08, WWK05, Xia04, XAI06, XMP07, XP04b, XLP05, XLLZ06, YZ07, YB06, YXLF05, YSW06, YS07b, ZGG03, ZR08, ZXQX08, ZZVM08, ZL09, ZSC06, ZSC08, ZHSS09, ZW03, dSMN^{+}04, dTDI^{+}07, vBK03]. fluctuating [SP04]. fluctuation [Asl04b, DPRN05, Ham07, Hub07, Hub08, KIHM09, RDPN07]. Fluctuations [DDG02, LS02a]. Fluid [AMSZ03, BNV08, Bar02b, BW01, ČPT01, CYKC01, ELW01, Fed02, Goe00, HK00, HLS06, Har04, HF00, HPZ01, IYI^{+}02, KFV^{+}05, KLvBvL02, LKNG01, LRN^{+}02, Man02, MC02, MD01, RRL01, RR02, Shy01, Str01b, SP00, TC02, WLE^{+}00, WW00, ZSP02, AS09, APT09, ADR08, AKP07, AMS03, BQQ09, BALW06, BL08, Bod06, BGS08, BG05b, CGL08, CR05, CCV03, CN05, CZ09, CC08b, CHPR09, CBS05, CDV05, CDV07, CDL04, CDL05, DMR09, DDM07, DMP08, DFV08, Eld07, Eld08a, EF03, FRS08, FDD07, FGS09, FM04, FJ09, FLM08, FKK08, Gla05, GA09, GCCD07, Gre04, GAC^{+}09, GH08b, GGCC09, Hel05, HC09, HMMR04, HLRZ06, HG03, HAI09, IAT08, JJGL06, JJGL07, JL04a, JBF07, KG09, KSHS08, KG08, KJ09b, KDF07, KYK07, Lap03, LMV04, LWP^{+}09, LS07, LLS09, LKY03]. fluid [LKX04, LKW05, LMH07, LHGF04, LHGF05, LH08a, LZH^{+}07, MC04, Mai09b, Mai09a, MMS04, MPD08, MY03, MMPB07, NMG09, Pap08, PP09, PSC^{+}06, PP04, PK07, QA09, RSW06, RFFP06, RM07, SJ04, SPT05, SL07a, SL03, Shy04, SM06b, SG03a, Sus03, TT09, TPV07, TGB^{+}07, TDV06, TG04, VGZB09, Vik03, WTL08, WWK05, XW06, Yam05, YJL^{+}06, Yu05b, YZL^{+}06, ZKS^{+}09, ZSC08, ZTPM05, dSMN^{+}04, vBK03, vLAvdV06, vZdBB07]. fluid-body [Eld08a]. fluid-dynamic [Lap03]. fluid-dynamics [MPD08]. fluid-elastic [HMMR04]. fluid-land [KJ09b]. fluid-membrane [LWP^{+}09]. Fluid-Mixture [Shy01, Shy04]. fluid-particle [DFV08]. fluid-particles [FM04]. fluid-poroelastic [BQQ09]. fluid-saturated [GH08b]. fluid-soil-structure [SM06b]. Fluid-Solid [HPZ01, Man02, CCV03, JJGL06, JJGL07, MMS04, NMG09, Vik03]. Fluid-structure [BNV08, AKP07, GA09, GGCC09, HC09, KYK07, LZH^{+}07, Pap08, SPT05, SL07a, vLAvdV06, vZdBB07]. fluid/flexible [Yu05b]. fluid/flexible-body [Yu05b]. fluidics [RE05]. fluidized [Sar03]. Fluids [ACK02, CL01a, FS00a, FS00b, HLS01, PR00, BL09b, BL08, FCT07, HHC08, ICO04, KM08b, KKL04, PvdV08, RE05, Ren07, SPB09, SF03, SCW^{+}09, VBL03, VBL04, XLM07, YZF^{+}06]. Fluorescence [FEL^{+}05, FLE03]. Flux [Bet08, Edw00, EF02, HGN00, KP00, KT02, Lio00, Ros00, Sti02, AKLMP09, BCDR06, CWYM08, CJ09, DQ04, Edw06, EZ08a, EF03, HS09a, IR09, JTL09, KK05b, KLLJ09, KT04, Kuz06, Kuz09, LSD07, MEKS03, MM03, MY07, Ols07, OK06c, QW05, QA09, RC09a, RBT03, SJD05, SMAj08, Sof09, ST03a, Tok06a, TAL09, VCZS04, XS05c, YHSX07]. flux-conservative [OK06c]. flux-limited [Ols07]. flux-split [MM03]. Flux-Splitting [EF02, Ros00, EF03, MEKS03, QW05]. Flux-Tube [KP00]. flux-vector [DQ04, QA09, SJD05]. Fluxes [DLS^{+}00, Bil05, BB07b, GLM07, QKS06, TT06b]. FMM [ON08]. foam [VCG03]. foaming [YFBH07]. Fock [FHLK05]. Fokker [DDFT09, KB04, WO05, WO09, BC02b, CBKM00a, CBKM00b, Den07, DWLM09, FP02, Lem00, LC03, PRT00, UL06, XCRX08]. Following [Abg01, Ano08-50, SM09b, WS04]. Force [Deh02, LKMK09, LM03b, Tót02, AKO09, BBF^{+}08, Car09, CB07, DM03, FCD^{+}06, Her08, HS08c, KKD08, Kim05, LC06b, LH05b, MZ07, SHP07, TLAD04, VQLZ04, ZVQ07, THD09]. Force-coupling [LKMK09, LM03b, DM03]. force-displacement [VQLZ04, ZVQ07]. force-driven [VQLZ04]. force/work [LC06b]. forced [HLRZ06]. Forces [BCE^{+}09, FPC^{+}00, LL01b, PSN00, SZ01, BYZ04, HHC08, Ni09]. forcing [AWK07, BZ09, Dom08, FM05, PPB09, Uhl05, YS07b, ZZ07]. forecasting [Ano08-50, Jan08, LP08, SM09b, SK08b]. Form [Han01, PKvdB00, IAT08, LY07a, LMN^{+}09, LB04, MESV09, PHW08, RK07, SH07b]. Formal [LP00]. Formally [Boy02b, CLMRP08, GHMP07]. Formation [DC02, HKV01, Nit01, BEA09, CLL07a, CGN^{+}07, GPL05, GH08b, KW08b, YFLS06]. formations [LZT09]. forms [MAN^{+}06, Tak06]. Formula [AGH00]. Formulae [FF02, YMF01]. Formulas [GM01c, TR02b, WF06]. Formulation [Ano08-50, BRL02, BMQS02, BS04c, CRD02, DC01, Hua01a, HMK02, KB01, LLIK01a, LKNG01, MF00, SM09b, TRL01, VQSZ02, Wan02, vdVvdV02, AAC07, AKH06, BBC^{+}06, BACFT05, Boe05, BRP05, BW07, BP04b, BMDS05, BSP06, CQRW05, CFR08, DMBS05, DB04, DBS06, Gla05, Gra06a, GAC^{+}09, Gui03, GD07b, GK05, HLMM07, HO03, HMM07, IX07, IX09, IDD04, Kim05, LSB04, LSJA05, LVW06a, LY04, LC06b, LBL04, Mad05, MP07b, NMS07, Pap08, PL09b, Pon09, PR03, PR04b, Pon06, Pon07b, PR06, SM09a, SRM09, SKS08, SS06a, SMP09, SP05b, SS05b, VMN07, VQLZ04, WG09, Xia04, XAI06, XJ07, XLP05, YB06, YAvdB^{+}08, ZVQ07, dSMF09, LLIK01b]. Formulations [IK01, WK01b, Dem04, FHLO08, GLT07, KG08, NV09, NZ05, SB06c, SD05a, VB09, WFC09, ZW06]. forth [DL03b]. Forward [UH01, AvdB04, LDW07, NSS03, RMGK04]. Forward-Adjoint [UH01]. forward-trajectory [NSS03]. forward/reverse [RMGK04]. FOSLS [HMMR04]. Found [BS00e]. Four [LCS02, Saf00, Saf02, BUEG06]. Four-Center [Saf00, Saf02]. Four-Dimensional [LCS02]. Fourier [DK06, ZGSD06, AC05, BK08, BS04c, BTSM09, Boy02a, BRB03, Boy09, BHP07, CLTA07, Che00a, Che00b, CKGL02, CKG04, Eli02, Eli03, Eli07, FBHV05, Fou06, GTD01, GLLN09, HO03, IA06a, KDK^{+}07, KS09, LS03, LDL^{+}09, MLFG06, Nas08, OLLL03, SSN09, SB06a, SS09a, SRNV07, SZLW06, TCN09, VB08]. Fourier-sine [BRB03]. Fourier-spectral [FBHV05]. Fourier/finite [GLLN09]. Fourth [BRL02, CVB06, GBS06, Lai02, MG07a, PKP01, XCZ02, YP01, Zha02, AV05, BC05, BGN07, CC03, CFJ09, GF05a, KKM08, LM08a, Nas08, SRX07, Hau08a]. Fourth-Order [BRL02, Lai02, XCZ02, YP01, Zha02, CVB06, AV05, BC05, CC03, CFJ09, KKM08, LM08a, Nas08]. Fourth-Order-Accurate [PKP01]. Fractal [WWVG00, AST07, CMP07, PC06a]. Fraction [Lin01, CMSZ09, Lap03]. Fractional [BE02, CGP02, Cod01, LOK01, VSG05, CLTA07, Cui09, Den07, Dom08, GSV06, HVAC09, LH05a, LDW07, LX07a, LP07b, LCdCN^{+}03, MST06, PC08, PCS^{+}09, Sou09, TMS06, TM07, Yus06]. Fractional-Step [BE02, LOK01, GSV06]. Fractions [SZ00]. fracture [LMH07, PKKL05]. fractures [TM05]. Fragmentation [Hew03]. frame [DDGS09, YGL05]. frames [AKH06, HHM04, KH07, PK07]. Framework [OCK^{+}02, STiST02, Abr06, AJ09, AKO09, BG09, DBTM08, FLM08, FCD^{+}06, FMD^{+}09, GZ09, HMA05, Kou09, PHKF06, PBH04, SPT05, TCO^{+}04, TJLT08, VP09a, ZG08]. Free [Bal01, CBKM00b, DF00a, FCGK05, FG02, GHG01, HB02, OF02, SDD07, SCD00, TCM^{+}00, WSI08, ZTZ02, vBRK01, AMS03, Bal09, BRDM09, BS04b, BKM09, BPL06, BGN03, CPR05, COQ06, CS05, Cha07a, Che03, CEH09, CZ09, CLS04, DS05a, DQA08, DST07a, DKTT07, GT09b, GFR09, GS06b, GS03b, GCCD07, GAC^{+}09, GS09d, HWL08, Hel05, HZ07b, Hum05, KR09a, KK04, LZ07, LMX^{+}08, LRS07, LL04b, LSJA05, LG07, LSV09, LY04, LD04, LZH^{+}07, MRC06, NJX09, NK08, NMS07, OTCM08, PN03, QP03, RB05, RMF08, SDGX07, SE04, SKXK05, TW07, VP09b, XMP07, YMT^{+}04, YP06, YH07a, ZL09, dSMN^{+}04, vZS07]. free-boundary [AMS03, VP09b]. Free-Lagrange [HB02]. free-plasma [Hum05]. free-streaming [GS06b]. Free-Surface [DF00a, GHG01, vBRK01, WSI08, Che03, DS05a, GFR09, GS03b, GCCD07, GS09d, PN03, RMF08, SE04, YP06, ZL09]. free-surfaces [Hel05]. free/Galerkin [SKXK05]. Freezing [JC02]. frequencies [KR09c, RMV03, WC08, OS01]. Frequency [CBB01, DDF01, DFT01, ERT02, GKL00, HMM02, POS00, TK02, ACR08, BL09a, BCDW06, BO09, CJSS08, CdHST08, DH04, DNS08, FLE03, JLOT05a, JY08, MBS03, MJ06, PL09a, PS07c, RKE^{+}07, TET09, TMND07, WB09b]. Frequency-Domain [CBB01, CdHST08, FLE03]. Fresnel [YFS01]. friction [BIW04, BBF^{+}08, HC08]. frictional [ZVQ07]. Friedrichs [KOQ04]. Front [BSJ01, GNNB08, JC02, LS08, SJ02, TNGH02, TB00b, TBE^{+}01, ZH01, CB09, DFG^{+}06, Fan08, HSL08, LLP07, LLGL07, kM07a, MT08, SAM05, TZ06, TT09, WKB07, ZEA06, dSMN^{+}04]. front-capturing [CB09, dSMN^{+}04]. Front-Tracking [JC02, TNGH02, TB00b, TBE^{+}01, ZH01, Fan08, LLP07, kM07a, MT08, TZ06, TT09, dSMN^{+}04]. front-tracking/front-capturing [dSMN^{+}04]. front-tracking/ghost-fluid [TT09]. Fronts [JW02, Set01, LMS05, Vol04a]. Frozen [CS01c]. FRS [AMP09]. fuel [CKPW07, SXyWX09]. Full [Edw00, HJKO08, ZH09, EZ08a, FCJ08a, FM06, IITV07, LJW07, OX04]. full- [IITV07]. full-Burnett [OX04]. Full-wave [ZH09]. Fully [BN04, Bon00, BSW05, DOWB01, HLS02b, MVO04, APR09, ALT08, AMP09, CL06b, CN05, Dim07, Fan08, HHMK05, JLT06, KT04, LC06a, LRZ04, MG07b, ODCK07, RWMK03, RSW06, RJM07, RSTB03, STD^{+}05, SMP09, SC09b, WAO^{+}04, WDÖ^{+}03, YM07]. fully-implicit [Dim07, WDÖ^{+}03]. Function [CHR01, GST02, hRT02, Bea08, BHNPR07, BKM09, Boy06, CWJ07, CMSZ09, CTS07, DSB06, DL03b, FP08b, GS06b, JSCZ08, Khe04, KAS06, LL04a, MP07b, MJ07, OJW06, Pee03, RRC05, RK07, SC08a, SH07a, SCT06, Sme06, TB09, Tow07, TW03, WZ07, Wen07, Wen09, ZJW06, ZC09, dHRvdB07]. function-vorticity [LL04a]. Functional [FS00a, FS00b, Lou00, VD00, VD02, BT03, Chr03, ET06, FHW07, GMH06, HdGKG08, PG04, SF03, VD03]. functionally [ZB07]. functionals [Küm04b, RSS09]. Functions [BS00e, CY00, GST00, Goe00, MS01, RS00, Saf00, Saf02, AKLMP09, BZ08, CT09, CQO04, CCJ07, ETT05, FB08, FW07, GG09b, HBHS09, HKS09, HS03b, IR09, JTL09, KMID05, KLW09, KR09c, LM08a, LCW04, LJS08, LJW07, MLSD07, MG08, MT07b, PLS^{+}09, RC09a, RA09, Tow08, Tow09a, Tow09b, WF06, YZLH09, Yin06]. Fundamental [BR01, BB08a, SY09b, YJF^{+}06]. Fup [GG09b]. Further [CKG04, SVB09, CM03, Hig05]. fusion [Jar04]. Future [Ano00q, Ano00r]. FV [AT09]. FV/FE [AT09].

G [LM03a, VP09a]. G-Scheme [VP09a]. GaAs [GS06a]. Gal [WS04]. Galerkin [Ain04, AB07, AKLMP09, AQ00, APQ03, BKST09, BC01, BCDR06, BDHN09, BRC^{+}09, BS04c, BSB01, BG05b, CKLS05, CC07, CELS07, CS07b, CS08b, CHG^{+}07, CJ07, CLS04, CFP06, CR00, CHPR09, CBS05, DD09, DLP08, DF04, DBTM08, Eg07, ES06, FCJ08a, FCJ08b, FK07a, FOLD05, Gab07, GLM07, GLMH09, Gel06, GFR09, Gir00, GHW02, Gir06, GR08, GLLN07, HH02a, HH08, HA02, HEML00, HO03, IK01, JH06, KCGH07, KvdVvdV06a, KvdVvdV06b, KvRvdVvdV07, KWBH09, Kri07, KLM05, KWD07, KDW08, LGHD08, LL01a, LSY04, LSJA05, LSZZ08, LS00, LY06, LX07b, LNXNTX09, LGM08, LZC04, LMSW02, LMNK07, LBL06b, LBL07, LBL08, Ma05, Mac07, MRC06, MY09, MESV09, MPFC08, NM06, NL08, NPC09a, NPC09b, PvdV08, QS04, QKS06, QLK07, RH01a, RBS06, RBvdV08, RC09b, SFE07, SMB09, SR09b, WM07]. Galerkin [WM09, WG09, WKG06, XXS07, XS06, XS09, XS05b, XLS09a, YS06, ZQSD08, ZQ09, vdVvdV02, vdVX07]. Galerkin-free [SKXK05]. Galerkin-like [LNXNTX09]. GaN [GS05a]. Gappy [GSK06]. Gas [BZW01, CKR00, CKR01, DC02, FS01, GV02, HK00, KMA^{+}01, LZ09c, LX00, OB02, Sai02, SZS01, SPC01, SB02, TX00, Xu01b, Xu01c, Xu02b, dSAK00, AK09, BPMR08, BS04b, BDCG03, CPR05, CELS07, CJR04, CDL04, DDK06, DVHM05, Fox08, GS05a, GC06, HP04a, HH06, JX07, KKM08, KD09, KK05a, KW03, LSL08, LM04, LZ04, LF06, LL07, LKW05, MEKS03, MSJ07, Myo04, NFvS^{+}06, OLA08, OK07a, RC06, Sam09, SH07a, SSB07, SE09, SFX03, SHY07, SY08, SS09c, SHPC09, SBC04, TXCD07, TT06c, UBRT07, VVS08, WTL08, WZ03, XH03, XMT05, XHC08, YHSX07, ZXQX08, ZRS06]. Gas-Kinetic [CKR00, CKR01, LX00, TX00, Xu01b, Xu01c, Xu02b, LZ09c, JX07, LF06, MSJ07, SY08, TXCD07, XH03, XMT05, XHC08]. gas-liquid [DDK06, HP04a, LL07]. gas-particle [Fox08]. gas-phase [OLA08]. gas-solid [DVHM05, HH06, MEKS03]. gas-water [LKW05, WTL08]. Gasdynamics [Myo01, QCGQ03]. Gaseous [VG01]. gases [SM05, VS09]. GASpAR [RFFP06]. Gauge [BU02, PS07d]. Gauge-Uzawa [PS07d]. Gauges [SS01b]. Gauss [ABHT03, AB05a, CLS05, KK07, VB08, WK06, WGCE01, ZHSS09]. Gaussian [ADK00, ADK02, BZ09, Cam03, Chr03, FG04, HMA05, KKS07, LQ09, TET09]. Gaussians [PC02, TB09]. Gautschi [BHvdV06]. GDG [FCJ08a, FCJ08b]. Gegenbauer [Boy05b, Lur07, MLFG06]. General [ALT08, AG09, BLW01, CL01b, CRD02, DPCV02, Edw00, ELC02, LBD02, PW00b, PW01, AK06a, ADR08, Bar04, BRC^{+}09, BP09, CDDL09, CWYM08, CEH09, DHOT09, DSJ03, ERVE09, FLB03, GCGE03, GSB03, GBS06, Her09, HR07, KA05, KPK09, KS08b, Kuz06, Lau06, LJ09a, LH08a, LGM08, Mac03, MS03, MJT06, MSO04, MY03, NN04, RAB07, RCD05, RH05, SS05a, SP06a, XDB09, Yam05, YHSX07, ZH04, vdVvdV02]. General-Purpose [DPCV02, Kuz06]. generalization [Ber06a, PS07a]. Generalized [BTFY01, IK01, Kro05, Lin01, LR01b, Mit00, Myo01, NL09, RMF08, SKAS01, SS00, VSV03, VQSZ02, Yon01, AS07, AK09, ABRR09a, ABRR09b, BSW03, BLW04, BS08a, BBvdV06, CXB08, CJSS08, CYS06, FCJ08a, FCJ08b, FRS08, GH03, JMZ04, LY07a, LS07, LX09, Myo04, PPB09, ST06, SHY07, SK04b, WFTS05, WG08, WK05, WAH09, XK03]. generalized-Laguerre [BS08a]. generated [EES09, FNBB^{+}08, MR07b, MSB07a, SM09a, WF06]. generating [FE04, GZ08, HvHHS05, Nit05]. Generation [AJG01, CFGK05, GW06, VB00, VRM07, CGDT09, CJ04, GS09d, Hua05, Kar04, Kau03, KSJ03, LB03a, RS09b, SE09, TDWY08, ZJW06, ZJWC08]. Generator [MDJS07, Aza06, KE09, MWM03]. Genetic [HCG01, MK02a, RS02, KGJ05, MC03]. Genetic/Powell [HCG01]. Genuinely [GF02]. Geodesic [Gir00, TTSG01, JC06a]. Geodesics [MS01, YC06a]. geomaterials [MDM03]. Geometric [CK08, FGG01, Gos02, IKS01, MG07c, OCK^{+}02, SK07a, BGN07, ÉGP09, JY08, KS08b, MY06b, SS09a, YW07]. Geometrical [dSM05, AA07, AMSZ03, CMSZ09, CQRW05, JW06, LH08a, Wen06]. geometrically [AK06a, BCZ04, GS05c]. Geometries [CL00a, CL01b, DDH01, KKC01, LMS02, Mie00, BWLM09, BYZ04, BS04c, CJ07, GCGE03, GN07, GBS06, GMO04, KB06, LV07, MCM04, MCGV04, MMPB07, MK06, PKD07, PC08, Pop03, RJ06, SROCdPFF05, TAL09, YXLF05, ZJWC08]. Geometry [ART02, CRB00, Lai02, LBV00, LBV01, OS01, PW00b, PW01, AMSZ07, BAFL09, BTT08, BBW06, BP04a, CGRGV^{+}04, CR05, CP07, GIA^{+}08, HSZ04, Jao07, KZWY09, LG04, ML05, ML06a, Mai09a, MS08a, ORM06, OPML07, Pro03, RB09b, SP06a, TF03, XSG08]. geometry-aware [ML06a]. geometry-based [CR05]. geometry-compatible [BAFL09]. geometry-dependent [ML05]. Geophysical [Bar02b, RFFP06, FD09b, PS03b]. Geophysical-astrophysical [RFFP06]. Geostrophic [FR02, MPD03, TRSK09]. Gerris [Pop03]. GeSEM [CJSS08]. GFD [CYS06]. GGB [WFTS05]. Ghost [Fed02, LKY03, BF08, DMP08, FRS08, LKW05, MMPB07, TF03, WTL08]. ghost-cell [TF03]. ghost-fluid [MMPB07, TT09]. Gibbs [Boy05b, JS07]. Gilbert [dSM05]. Ginzburg [DDG02, RSS09]. giving [TW05, TR07]. GKS [GLLX08]. glass [Thö04]. Glimm [Min07]. glitch [Tan05b]. Global [DSS00, LRMB08, RTT01, BG07, CGM07, Gel06, GD05, LJS08, MC03, NSS03, NLLE06, SMT^{+}08, THN^{+}07, WFTS05, WDÖ^{+}03, dDEK09, ADK02]. Globalization [BB07a]. Globe [LR01b]. Globular [DPRS01]. Glow [hLA01, SS04]. GMRES [GKL00, NOG08a, TWS02]. Goal [BTWGvBW07, OV00]. Goal-Oriented [OV00, BTWGvBW07]. Godunov [AT05b, BSKH07, CJR04, CCF^{+}05, FF02, GS05b, GS08, GR04, Gui02, HB02, LD04, MN09a, MP05, MC01, Min07, MC07c, PK03, SWK06, SZS01, TB06, TFD06, Xu01c, vBK03]. Godunov-type [AT05b, CJR04, LD04, MN09a, PK03, TB06, vBK03, Gui02]. Good [Vas00, ZYHS07]. Gordon [BY07, HJL09, HZ08, RBK09]. GPU [ELD08b, PVPS09, YWC07]. graded [HO08a, MGC06, MG06, MG07d, ZB07]. Gradient [JLCD01, JTB02, PKvdB00, Whi00, ZCMI01, AMLC08, CSL08, DEHL06, Fen06, JBHK08, KDF07, Küm04b, Ovt08, RSS09, RBT03, Tok06a, Tow09a, Yan09]. gradient-dependent [JBHK08]. Gradients [HGN00, Cho05, DL03b, LP06a, ML05, Pro08, SNLS03]. Grain [KAIN01, CP04b, ES03a, EES09, SW04a]. Grain-Boundary [KAIN01, CP04b]. grained [IM07, KMV03]. Granular [CB02, TNGH02, FPT05, KD09, SM05, Vol04b, VQLZ04, ZVQ07, ZD08]. granular-flow [VQLZ04, ZVQ07]. granular-gas [KD09]. graph [LQX06]. graphics [ALT08, GD08, KWBH09]. grating [DBB06]. gratings [BHS03, BS04d, BS05]. gravitating [BvdHKG07]. gravitational [DHM03, NB04, TXCD07]. gravitationally [GBC06]. Gravity [BTFY01, Cha09, DQA08, GB08a, Kas07, VBL04]. Green [BKM09, CY00, CWJ07, CTS07, DSB06, FDD09a, JSCZ08, KR09c, MLSD07, MRRS05, OJW06, PLS^{+}09, RK07, SCT06, dHRvdB07]. grey [DL04, ED07, MYW07]. Grid [AJG01, ALGM01, AV02, BdLL01, CL00a, Cal02, CL01b, CR02, DCV^{+}01, DDH01, FH00a, GG00, HH07b, HLS02a, HLKS00, KT06, KKR01b, LP02, LLQ^{+}02, MCJ01, Par02, RW00, RMO00, SMP01, SY03, SMO00, TTSG01, TSG02, UMRK01, VD00, VD02, WZ00, XCZ02, Yam01, YC02, YSC01, dSAK00, vdVvdV02, AC09, Aza06, BGM08, BS04b, BCM^{+}07, CL06b, CP06b, CP06c, CK08, CGKM06, CSML06, CK07, DMHP07, DCF^{+}08, DTMS06, DG09, FS04, FL03, GZ07b, Gro06, Gro07, HZGB04, HZGB05, HC09, Her08, JX06, JX07, KE06, Kau03, KZWY09, KAK03, KAS06, Lap04, LZ07, LZ09b, LZ09a, LR07, LKMU05, LQX06, LSW08, LRS09, LBL06a, MZ09, MWM03, Mad06, MKLU05, Men04, MZ07, MD06, MHE06, MV08, MK06, MO06, NMM^{+}07, OK06a, PS03b, RW03, SS03a, SROCdPFF05]. grid [SBCL06, SMS04, SZ05, SS07b, SS05b, SP06a, ST03b, TJS03, THL06, Thu08a, TU04, VD03, WW04, YMT^{+}04, YU05a, YXLF05, YSS05, ZG08, ZT03, dTDI^{+}07]. grid-alignment [BGM08]. grid-based [CP06b, CP06c, MO06]. grid-free [BS04b]. Grid-Optimized [CL01b]. gridfree [SC08a]. Gridless [OC08, KL04, LBL06a]. Grids [BE02, Car01, DV02, DPCV02, DI02, Edw00, FGG01, GZ01, GHV00, GW01, GC02a, Gir00, HLS02b, HH01, HW02, HA02, JK00, KC00, LM01, LLdlP^{+}00, SZ00, TC01b, hRT02, Wan02, WL02, ZTZ02, Aza07, Aza09, BAYZ08, BFB08, BSKH07, BHS09, BMT09, CCV03, CWYM08, CMG09, CEH09, CYS06, CSL08, CSKD05, CT07, DSM09a, GS09b, GSB03, GF05b, HWL08, HV03, HNF07, HKG08, HS03a, HS06, HS08a, Her08, HHMK05, IA06b, JH06, KK05a, KKCF09, KK03b, KPK09, Kok09, KPP09, KDW08, LSB04, LLY05, LVL05, LM03a, LCH03, LJW09, LVW06a, LVW06b, LLB05, LH08a, LS05b, LBL06b, LBL07, LBL08, MS03, MJT06, MGC06, MG06, MG07d, MCN03, MCP03, MLS^{+}05, NE05, OSK09, PS04, PS08, PSM08, PL07, Rem06, RRW05, RWWS07, RCB05, SB06c, SS05a, STZ07, SPGR06, SWL06, TZL05, TRSK09]. grids [TDGP06, TMG08, VGPL09, VSW04, VSW06, WZL04, WG09, WA08, WdND06, XLS09a, XLS09b, YJL^{+}06, YJ06, ZT07b]. Grooving [KAIN01]. Gross [BJM03, CORT09, CJK^{+}03]. Ground [BT03, FGOV00, BCL06, CORT09, Woo06]. Ground-state [BT03]. Ground-States [FGOV00]. Groundwater [MBP00, WGNT06]. group [BFJ03, BMK^{+}06, HJFW04, San03, Har04]. group-III [BMK^{+}06]. growing [Mad06, MM07]. Growth [CMK^{+}01, GW02, GGL^{+}01, JK02, RV00, BHL^{+}04, CLL07a, CF06b, DQA08, EES09, HLX06, LB03a, ML05, ML06a, MWG^{+}06, NDG05, PSCQ03, RS06b, SSE03, SA06, Sus03, TZ07b, WLT08, YFBH07, ZVHP03]. GRP [BALW06, BF07, LLS09]. Grüneisen [Shy01]. GSHMC [AR08]. guess [TDV06]. guesses [BHL07]. guides [GP04]. guiding [Mot08]. Gyro [BB09a]. Gyro-water-bag [BB09a]. Gyrokinetic [KP00, Par02, BB09a, CW03, CP03c, CP07, HKM07, HJKO08, IITV07, LGKP07, NLLE06]. gyrokinetic-Maxwell [CW03].

Haar [LKNG04]. Haidvogel [SM09b]. HALE [AHNS09]. Half [DK02b, TFDK04]. Half-Moment [DK02b]. Hall [CK03, ODCK07, TMG08]. Hamilton [BL03, CQO04, CC07, CS07b, CY05, FF02, KOQ04, KOQ08, KR02, KT00b, QS05, TTZ03]. Hamiltonian [JW06, BM01c, BRP05, CKL00, GAC^{+}09, LY07a, Rei00, Tan05a]. Hamiltonian-preserving [JW06]. Hamiltonians [ESD05, KJ01]. Hancock [Waa09]. Handling [FPC^{+}00, BOT05]. Hard [WB01, DTSC04, DTS05a, DTS05b, DST07a, NFvS^{+}06]. hard-particle [DST07a]. hard-sphere [DTSC04, NFvS^{+}06]. Harlow [Bra04]. Harmonic [CKGL02, KMHR00, LTZ01, SS00, APT09, AG09, BCR04, BHNPR07, DLP08, EHD08, Gab07, GG09a, GD07b, LTD^{+}06, MKOW04, MHPR08, Tyg08]. harmonics [KL06, WJV07]. Hartmann [HY11, HY09]. Hartree [FHLK05]. Haurwitz [SD06]. having [Wen06]. HD [Saf00]. heap [NFA03]. heap-based [NFA03]. Heaps [Mu02]. heart [NLT07, vLAvdV06]. Heat [BW01, CS01b, CPK02, IYI^{+}02, LLH02, LTK^{+}02, MCJ01, ZZ01, AMXL09, DQ04, FHLO08, FF03, FDK06, GF05a, GL06, GYKL05, GLT07, GL09b, HL04, HC09, JG09, LG07, LCNR07, LR03, MJ09a, MSP^{+}06, MR07c, Mou04, PFSL07, PS07c, SBCL06, Tau07, Thö04, TFDK04, VB08, YYF09, YLD09, YSW06]. Heat-Transfer [IYI^{+}02]. heated [FPT05]. Heating [OL01, Rid00, DMR09, FLB03]. Heaviside [MG08, Tow09b]. HEFAT2005 [Ano05s]. Height [Bon00, FB08]. Hele [FS04, KW08b, LLL07]. helically [LP07a]. helicity [LW04]. helicopter [EHD08]. Helium [GG00, LKE04]. Helmholtz [AST07, AAC07, AHPT07, AC05, AQ00, BB08a, BFT07, BFT09, BKM09, BK08, Boy05a, CWJ07, CHL06b, CCG^{+}06, Dar02, FT05, FSY00, GKL03, Had05, HMPR07, IK01, KTD03, KR09c, LT09b, MP03, ND04, Pri08, RKE^{+}07, SFY01, ST04, VW02]. Hermite [Cap08c, BS08a, Cap08b, Cap09, GS06b, LBL07, QS04, QS05, TCN09]. Hermitian [Cap05, Cap06, Ovt08, SPLM09]. Hessian [AI09]. Heterogeneous [CGSS00, DPRS01, OV00, RE05, AE03, AZ06, CCV03, CDS04, GZ07a, GZ09, GLLN07, GLLN09, JLT06, LZT09, LTZ03, LH05b, RKE^{+}07, SMGJ09, WGNT06, YE07]. heterostructures [BNNP06, WHLL03]. Heuristic [Dwi08]. hexagonal [Thu08a]. Hexahedral [MHS01, ZYC02, Aza06, Aza07, Aza09, CFP06, FM08]. Hierarchical [Deh02, XLS09a, XLS09b, KF06, MZ09, MG05a, VCM00]. hierarchical-element [VCM00]. High [AC00, ACY00, APTJ^{+}04, AHNS09, BS00a, BFT07, BBMB07, BRC^{+}09, BLM03, BTW03, BGLN05, BK01, BL03, BSB01, Car01, CL01a, CP04a, CKGL02, CYKC01, CS06, CS07d, Coe02, CR00, CSP01, Cor00, CFR08, DT04, DZ00, DBBP08, DR09b, DDFT09, DLS^{+}00, ERT02, FT01, FR03, FCB02, FG02, GZ01, GF02, GHW02, Gir06, Giv01, GN03, GP00b, GLN06, GKL00, GL08, HMOG08, HW02, HT00a, HT00b, IX09, JH06, JK00, KZ04, KMS03, KT06, KMA^{+}01, KB01, KB06, Kry04, KT00a, KT00b, KT04, LL09, LG04, Li08a, LYC09, LS00, LCS09, LR03, LF04, MN06, MP05, MC06b, MC01, MD01, NM06, NR01, Nic00, Noe00, NXS07, NC01, NT07, Nys02, OGV02, PW00b, PW01, QS02, RH05, SLY02, SYG06, SZC09, SS05a, SHY07, SL07c, TK00, TX00, TWYC06]. High [TCN09, TD07, TB04, TS02, VCP00, Vas00, WH02, WSYS09, Wen07, Wen09, WAH09, XS05a, XS06, XK01, YHSX07, ZTZ02, ZW04, ZT03, ZW03, ZZFW06, ZS01, de 00, Ain04, AMR06, ACR08, AC09, BRDM09, BSKH07, BHS09, BFT09, BdCB09, BBCT09, Boy06, BSW05, BH04, BHP07, BJP04, CJSS08, Cap08c, Cap09, CT08a, CQO04, CK03, CTS07, CC07, CS07a, CKG04, CFP06, CC04, CF04, DSM09b, DE06, DC07, DTMS06, DDH05, DI09, DK07, DET08, DZ09b, FNS07, FOLD05, FD07, GPC07, GG04, GH08a, GSV09, Gom08, GL06, HH07b, HLS06, HWL08, HS03a, HS06, HJJ09, HY09, HY11, HH06, IQ08, JD09, JBF07, JLOT05a, JY08, JS05, KE06, KCGH07, KG09, KM06, KM07b, KK05b, KLK08, KYK07, Kok09, KPP07, KPP09, Kri07, KR09c, LG09, LL05, LTZ03, LX09, LN09]. high [LDW07, LR07, LF05, LTD^{+}06, LS09, Mai09b, Mai09a, MRS09, MY06b, MOG09, Min04, MPFC08, NLF03, NOG08a, NOG08b, NPC09a, NPC09b, NWZL08, NF09, NGvdWS09, OF06, OVG07, PPDM08, PPCW06, QW05, QA09, QFR04, RRW05, RKE^{+}07, RF06, SDM04, SWK06, SFE07, SMB09, SZS03, SZ05, STZ07, SJHM09, SY03, SGG^{+}04, SCN07, SN08, TET09, TFD06, TDWY08, VWW04, Wan04a, WL06, WM07, WGRA09, WZ07, WA08, WMH07, XCRX08, XP04b, XS05c, XYK05, YC09a, YS07a, YBZ06, You06, ZL04, ZJ09, Zho07, ZYHS07]. high- [CK03]. High-accuracy [AHNS09]. high-amplitude [PPCW06]. high-dimensional [DI09, LDW07]. High-fidelity [NT07]. High-Frequency [ERT02, ACR08, RKE^{+}07]. High-Order [AC00, ACY00, BK01, BSB01, CL01a, CKGL02, Coe02, CR00, CSP01, DZ00, DLS^{+}00, FT01, GHW02, Giv01, HW02, KB01, LS00, MC01, NR01, Nic00, NC01, Nys02, QS02, SLY02, TK00, TX00, TS02, VCP00, WH02, XK01, ZS01, de 00, APTJ^{+}04, BFT07, BBMB07, BRC^{+}09, BLM03, BL03, DR09b, Gir06, GN03, GL08, HMOG08, JH06, KT06, KB06, LL09, LF04, NM06, NXS07, RH05, SS05a, TWYC06, TCN09, TD07, WSYS09, WAH09, YHSX07, ZW04, ZT03, AC09, BFT09, BdCB09, BH04, BHP07, Cap08c, Cap09, CT08a, CC07, CKG04, CFP06, DSM09b, DC07, FNS07, FOLD05, FD07, GPC07, GG04, GH08a, Gom08, HH07b, HWL08, HJJ09, JD09, JBF07, KE06, KK05b, KLK08, KYK07, Kok09, KPP07, KPP09, Kri07, LTZ03, LN09, LF05, LS09, Mai09b, Mai09a, MY06b]. high-order [MOG09, NLF03, NOG08a, NOG08b, NPC09a, NPC09b, NF09, QW05, SFE07, SMB09, SZ05, STZ07, SGG^{+}04, SCN07, SN08, VWW04, Wan04a, WL06, WM07, WA08, WMH07, XCRX08, YC09a, YBZ06, You06, ZL04, ZJ09, Zho07, ZYHS07]. High-Order-Accurate [OGV02]. high-performance [XYK05]. high-Re [KM06, KM07b]. High-Resolution [FCB02, GF02, GP00b, JK00, KT00a, KT00b, MD01, Noe00, PW00b, PW01, ZTZ02, BTW03, GLN06, Kry04, KT04, LR03, SYG06, SL07c, KG09, BSKH07, BBCT09, CC04, OF06, SWK06, SJHM09, TDWY08]. High-Reynolds-Number [FG02]. High-Speed [KMA^{+}01, BHS09, HS03a, HS06]. high-wavenumber [CC04]. Higher [DV02, Fox09, GP05, HM04, JMC03, MBM01, PM00, SH07a, SP06b, Tol07, VGCN05, VG02, YMT^{+}04, vZdBB07, APP^{+}07, FRS08, GLT07, HMPR07, Hel09a, HMMO05, HB05b, KPB08, MC07c, MHPR08, ODAF07, PKD07, WD07, ZJS08]. Higher-Order [PM00, VG02, Fox09, JMC03, SP06b, YMT^{+}04, vZdBB07, APP^{+}07, FRS08, HMPR07, HMMO05, HB05b, KPB08, MHPR08, WD07]. Higher-than-fifth-order [Tol07]. Highly [CL02, FK02, SE09, deM02, DGH08, EG08, HH07a, HDR^{+}06, Lar07, MD06, SWB^{+}06, WGNT06, vdDA06]. highway [ZWS06]. Hill [DK06]. Hilliard [CR07, CFP08, pHL09, KW06, KKL04, WKG06, WKL07, XXS07]. histogram [BS09a]. history [BS07]. HLL [Jan00, MK05]. HLLC [Cap08c, HJ07, HAI09, KLLJ09, Li05]. HLSM [Cap09]. HOC [KR09a]. Hodge [Sum00]. hollow [DJ04]. Holm [COR08]. Homentropic [KLvBvL02]. homoenthalpic [BEA09]. Homogeneous [Cle00, SPW^{+}00, BIVC07, FR03, KM07b, KW03, PH09, SBA07, TMD07, GM04]. Homogenization [PR01b, AE03, SKR06, YH07b]. homotopy [LR07]. Hood [LSA06]. Hopper [GM01b, GMO04, GMS06]. hopping [KLW09]. hot [GS05a]. hot-phonon [GS05a]. hp [BW01, ES06, NM06, PP09, PR03, PR04b, PR06, SC01]. hp-finite [PP09]. hp-multigrid [NM06]. Hubbard [CD04]. Hugoniot [JR09]. human [BCDW06, XS07, ZK04]. Humans [POS00]. Hutter [FNBB^{+}08]. Huygens [Bér07]. HWENO [BAMD07]. Hybrid [BM02, Bow01, CS03, EFFM02, EF02, EF03, GK01, HL07a, HP04b, HCG01, JPMC01, KC00, LS02a, LM01, MPC01, MPC02, Pir02, QL01, RB02, SA06, Stu01, WWK05, vdHK07, AK06a, ABRR09a, ABRR09b, AJ09, AGW07, BFB08, BAMD07, BCCV09, BB09c, CCG08, CHB09, CYS06, CD07, CDL04, CDL05, De 04, DBF08, DR06, DUEB07, GXW07, GS05c, GN07, HR08, IH04, KA05, KK05b, Kwo08, LG08, LM03a, LKO05, LBL06a, MC07b, MCP03, NTO^{+}07, NK08, NG06b, OK06b, PLS^{+}09, PGN08, RJ06, RLZ03, SRM09, SB06b, She08, SCRL08, SBC04, SK06, TKH09, VTC^{+}07, ZH04, ZL09, ZSB^{+}08]. hybrid-Vlasov [VTC^{+}07]. hybridizable [NPC09a, NPC09b]. hybridizations [RGK07]. hydro [LW04]. hydro- [LW04]. Hydrodynamic [Myo01, NJLA06, SMP01, AK09, BS04a, BM06, BTC05, BH09, BBW06, Che04, CELS07, DJM05, HGBH03, HLFB07, HS09b, LCB04, Li08a, LYC09, Myo04, SP04, SE04, SO08, SY08]. Hydrodynamical [Rom02, ZD00]. hydrodynamically [AGW07]. hydrodynamically-interacting [AGW07]. Hydrodynamics [BKR^{+}01, BZ08, CRB00, CPK02, DW00, iI02, PM02, ZF02, ASPB03, BRDM09, BZ04, BBC^{+}06, BOT05, BD06, CGSR08, CL06a, CDDL09, CFL^{+}03, CL03b, CEL06, DST07b, ESE07, HK08b, HX05, HG03, KSW07, Li05, LC06b, LSW08, LRS09, MN09a, MC06b, MESV09, Min07, MC07c, MHW05, RHPN09, TM05, TMSW07, ZB07]. Hydrogen [CP00, GG00, BRB03]. hydromedusa [SM09a]. hydrophobic [ZTPM05]. hydrostatic [BKLL04, IHL03, SE04, SP06a]. Hyman [RV07]. HyPAM [ZL09]. Hyper [MS01, DHM03]. Hyper-Surfaces [MS01]. hyper-systolic [DHM03]. Hyperbolic [AC00, ACY00, Asl01, ADK00, ADK02, BJ00, BAFL09, DKK^{+}02, Dur00, FMO00, GC01, KKP02, LL00, LMSW02, MOS^{+}00, NTYT01, NTYT02, PL01, RC00, Sti02, TS01, Xu01a, YL01, Abg06, AC09, AKLMP09, BR09a, BBCT09, BCCD08, CLG07, Cap08a, Cap08b, Cap08c, Cap09, CT08a, CP08, CGKM06, CD07, De 04, DQ04, DK07, DKTT07, DET08, Edw06, Gir06, Gui05, HH07b, Hwa03, JR09, JTL09, JLOT05a, JAK05, KPP07, LL03c, ML08, Mil04, MC07b, NvL03, NG06b, PC06b, RLZ03, RBvdV08, RDPN07, RBL04, Ros06, SDM04, SYG06, SZLW06, SR09b, TT05a, TT05b, THD09, THS07, TY07, Wen06, XS06, vDZ06]. Hyperbolic-Elliptic [Xu01a]. hyperbolic-parabolic [BR09a]. HyperCASL [FD09b]. hyperelastic [Mil04, YH07b]. hyperreduction [Ryc05]. Hypersonic [BM01b, KKR01a, KKR01b, ELD08b, SSB07, XMT05, XHC08, ZT03]. Hyperspherical [AKV00]. hypersurfaces [BGN08]. Hyperviscosity [CC05]. hysteresis [Spe05].

IAT [Moo03]. IBC [Ano07d, Ano07e, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36, Ano09-37, Ano09-38, Ano09-39, Ano09-40, Ano09-41, Ano09-42, Ano09-43, Ano09-44, Ano09-45, Ano09-46, Ano09-47, Ano09-48]. ICCG [VSMW01]. Ice [GC02b, Hun01, SSL00, MB04, Noe00]. Ice-Sheet [GC02b]. Ice-Thickness [GC02b]. ICF [DDFT09]. Icosahedral [TTSG01, TSG02, SMT^{+}08]. Ideal [SHWW00, TX00, AT05a, CCF^{+}05, FMR09, GS05b, GS08, HLS06, HT07, HJ07, LD04, MK05, NN09, Sam09, Ser09, SHY07, Tor03, TA06, Waa09, YHSX07, ZK06, Zie04]. ideal-magnetohydrodynamics [ZK06]. ideal-MHD [Sam09]. Idempotent [KCH06]. Identification [Gut00, BG09, Kou09, PS03a, VK04]. identifying [NLT07]. IDO [IKS^{+}09, IA06a, IA06b]. IECM [CRAG07]. IEM [CRAG07]. IFC [Ano03n, Ano03o, Ano03p]. ignition [VG01]. II [ACY00, AGT05, ADK02, BT07b, Cap06, Cap08c, CBKM00b, CP06c, CFGK05, CY05, DW00, DTS05b, FS00b, Gos04, GMO04, Hau08b, HS03b, IR09, JR04, JW02, KM07b, KKR01b, KK05d, LRN^{+}02, Lio06, LL03c, LKMU05, LMS02, MS08a, kM07a, NMH^{+}07, OKZ07, PGB05, PFSL07, PCS^{+}09, SD05b, TOZP03, Tyg08, VBJ08b, WL02, XAI06, ZLAC05, ZQSD08]. III [BMK^{+}06, GM06, GMS06, HT03, JW03, YU05a]. ill [vdDA06]. ill-posed [vdDA06]. Image [CPP02, KMA^{+}01, XDC09, DC07, FSS03, XCY06]. Image-Charge [CPP02]. images [ADE^{+}08, CJLS09]. Imaginary [GST02, LTE07]. Imaging [HSZ04, BHL07, BGR08, KNH05, KE09, PL09a]. imbedded [Zho07]. IMEX [HR07]. Immersed [CM00, Cor00, FVOMY00, GKV09, KKC01, KC06, LP00, LL01b, SW00, WFC09, ZP02, AKP07, Ber04, BJ09, BGS08, BP08, CFR09, CXZ09, COER07, DCK08, FM04, GS07, GMD07, GSB03, GS05c, GP05, GHMP07, HK08b, HS08b, HS08c, HSS07, HST09, HF08b, IK07, KIH09, KIHM09, LTH08, LKP06, LWP^{+}09, LJS08, LF05, LMZ^{+}08, MvW08, MDB^{+}08, NFGK07, PSC^{+}06, Pon09, RAB07, SLC07, SSND03, TZL05, TC07a, TLL^{+}08, TLK09, TF03, Uhl05, Vik03, WZL09a, WZL09b, WS09, XW06, Xu08, YS09, YZLH09, ZZ07, Zho07, dTDI^{+}07]. Immersed-Boundary [FVOMY00, KKC01, LMZ^{+}08, YS09, ZZ07]. immersed-boundary/level-set [YS09]. immiscible [ICO04, TBJ^{+}09].

Immittance [Mac00]. Impact [SDT08, ZYKW01, BZ08, CB09, GA09, KFH^{+}04, KFIG06, KFV^{+}05, MC06b, RGS04, SL04, TU04, UTBV03]. impact-produced [KFH^{+}04, KFIG06]. impacting [LKY03, WSI08]. Impedance [HCG01, CCT05, HKS09, IKL^{+}08, Lee07a]. impinging [NTB07]. Implementation [AG08, BK01, CTW^{+}08, DCS00, Dar00b, Gel06, HHPW08, IK01, JC06b, KHV01, LLS09, MM01, MN02, PM00, Set01, Sof09, CLG07, Car09, CP06b, CP06c, CELS07, DSB06, FLE03, FYH^{+}06, HWL08, KSHS08, KB06, LJ09a, Liv07, LLRP09, LQ06, MC09, Pon07a, SKS08, SHS08, SD05b, XDB09, YBS06, Yok07, vZS07]. implementations [VK09]. implemented [ALT08, CSL08, MKLU05]. implications [GPL05, KMID05]. Implicit [AC00, ACY00, AHF04, BCOS01, BCVK02, BSP06, CBKM00a, CBKM00b, CKF02, CB02, DOWB01, ED07, Gen01, GMB01, HK00, HT00a, HT00b, HLY09, HF08b, MS01, PCP08, RH01a, RLB02, RXH02, RB02, SWL00, WM07, WA02, WZ00, WS09, Yua02, ZTZ02, APR09, Alb08, Bon00, BLM03, BMDS05, BSW05, BUEG06, CFR09, CK03, CL06b, Cha07b, DPRN05, DR06, DWLM09, Dim07, DF04, FDD09a, GL09b, HR08, HAD06, HM05, HS08b, JBF07, JLT06, KCGH07, KK07, KB04, Kuz09, LH05a, LC06a, LM04, LWP^{+}09, LM08c, LH08b, Low04, LLC^{+}08, LZH^{+}06, LZH^{+}07, MvW08, MC06a, MELD08, MU09, MBS03, ML06b, Mot08, MBP07, MK03, NL08, NPC09a, NPC09b, NZZ06, NS05, OTCM08, ODCK07, RWMK03, RSW06, RJM07, RCB05, SFMP06, TBT^{+}09, TMD07, TDGP06, TCO^{+}04, Utn08, WPM02a, WDÖ^{+}03, WZ03]. implicit [WdND06]. implicit-explicit [GL09b, KCGH07]. Implicitly [Mou04]. imploding [BPMR08]. implosion [NJLA06]. implosions [FRS08, KS08a]. importance [ASPB03]. impose [PK05]. imposition [APQ03]. Impossible [Azm02]. impregnation [Mad05]. improve [HM09, LY07b, SFVK06]. Improved [BLW01, CKR00, CKR01, EFFM02, Hei05, KK00c, MP01b, Pro03, Pro05, SK03, Xu01b, Abr07, BCCD08, BOT05, CBH03, HvHHS05, HLMM07, KZWY09, KW03, LHGF05, LP09, ML06a, MSJ07, MD06, NE05, ODAF07, Pon06, Pro07, SCW^{+}09, TMD^{+}08, TS07, WYS09, WKB07, XDB09, ZZ07]. Improvement [Bil05, FDD07, SVB09]. Improvements [CSC^{+}08, HMOG08, MPFC08]. Improving [BL01, HMM08, TB06, AG09, GS03a]. Impulse [Cor00, Sum00]. impulsively [DCK08, KR09a, NCS03, Sam09]. Impurities [Gos04]. Impurity [VDM^{+}02]. inclined [QP03]. Including [BMS00, DK02a, JP03, BBDE05, BL08, GS05a, HC08, HF08a, WG09]. Inclusions [Bal02, CGDT09, KK03a]. Incomplete [LZL03, Moo03]. Incompressible [BCM01, CFA01, ČPT01, Cod01, CMOV02, Del03b, ESE07, GPH^{+}01, Goe00, GQ00, GSW00, HLS02b, HH02b, JL02, KA05, KC00, KM00, LLH02, LOK01, LS00, LW01, MPP01, MC00a, NFK01, Pai01, PKP01, PSN00, QV01, Ros00, Sni01, SP00, TC01b, VC00, WPW02, WPH00, WS01, Yua02, BCDR06, BGM08, BF08, BS04c, BJ09, BGLN05, CPR05, CCG08, CRAG07, CYS06, CSL08, COER07, CMR08, CC08b, DHOT09, DSM09a, DDH05, DS06a, DSS07, EHST03, EHS^{+}08, ÉGP09, FP08a, FL06, FGP08, GS07, Gel06, GSV06, GCNB07, GGP06, GS03c, Gri09, GR07, GS03d, GS09c, GLLX08, HAS05, HK08a, Hel05, HVAC09, HA07, HA09, HC05, IK07, IAT08, IOTK04, JLL^{+}06, KR09a, KE06, KDK^{+}07, KDF07, KT03, LL09, LKP06, LL05, LMX^{+}08, LF05, LRZ04, LKO05, Liu09b, Löh04, LHGF05, MVD04, MR06a]. incompressible [MCG08, MG06, MCN03, MDB^{+}08, MVO04, NW07, NZ05, Neo07, NMM^{+}07, NMH^{+}07, Ni09, Nik06, OVG07, PKD07, PP09, PSC04, PN03, PvdV08, PR04b, Pon06, Pon07a, Pon07b, Pop03, PR06, PS07d, QS07, RBS06, Ros03, Ros07, RFVP09, RW03, SROCFF03, SROCdPFF05, SC08a, SFE07, SD05a, SD05b, SS06a, STZ07, SKXK05, SZH07, SSH^{+}07, TZ03, TZL05, TLK07, TLL^{+}08, TJS03, Utn08, VCT07, VSW04, WRu03, WB09a, Xia04, XAI06, XP04b, Xu08, XSL09, YZ07, YYT05, ZR08, ZKY05, ZL09, ZJ09, ZFM08, ZHSS09]. Incorporating [BHR04, HKO07]. Incorporation [DM03, Cha07b, HS07b]. Increasing [AT08]. Increasingly [BS00a, ZJS08]. Incremental [DB00]. indefinite [Had05, RS05, RS09a]. Independent [BSB01, CCJ07, CN05, CJK^{+}03, CSMH05, Lap04, LTE07, SGFL09, YBZ04, Yin06]. Index [Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano00z, Ano00-27, Ano00-28, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano07-33, Ano07-34, VSG05]. index-1 [VSG05]. Indicator [CHR01, KKP02, BAMD07, CR05, Khe04]. indicators [ZQ09]. Indirect [KS02a, WK04]. Induced [POS00, ABLS05, BCDW06, DL03b, MM07, RV09, SZC09, SCRL08, VVS08, YT07]. Induction [CBB01, FLB03, GFR09, IDD04]. inelastic [SM05]. Inert [dSAK00]. Inertia [GS02, JP03]. inertial [KMSH08, PK07]. inertio [Kas07]. inertio-gravity [Kas07]. inexact [GG09b, HC05]. inextensible [VGZB09]. inference [MN09b]. inferior [Boy06, WZ07]. Infiltration [JWSC00, JW02, JW03]. infinite [BRB03, GVT01, Mil08, ST04, zSW06, zS06, VCT07]. Inflow [LP06b, FE04, GW06, KSJ03, Lar09, MJ07, SAK05]. inflow/outflow [SAK05]. Influence [BT07a, NW07, RB09b, Wal03]. information [HMA05, MY07, Ram03]. Inherently [BS00d]. inhomogeneities [AV03, TJ09]. Inhomogeneous [FS00a, FS00b, HHCL01, LMSV00, OS01, BH04, CJSS08, DBF08, GGOB04, HLFB07, Lar07, SF03, YHCD05, ZWS06]. Initial [Kas07, KJ01, AMXL09, BHL07, BS05, CL08a, FF03, RMGK04, SN06, Tem06, TDV06]. initial-boundary [FF03, SN06]. Initial-value [Kas07, BS05]. initio [GM06, SLG^{+}03]. Injection [CVB00, FKV08, dSHHM05]. ink [YSS05]. inner [Gel06]. inorganic [MWG^{+}06]. input [GZ08]. inputs [DI09]. insect [Liu09a]. insoluble [GT09b, JL04a, LTH08]. Instabilities [PD01, KP08, LL08b, MC09, MV06, NLT08, Pri08]. Instability [FBFF00, HGB^{+}03, LS02b, Lio00, MT01, Mon00, CL07b, CGM07, FS06, KTD03, sKKRH03, LSD07, LS08, Sus06, TM05]. Instructions [Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano07z]. instrument [FHJK09]. integrable [CHL06a, CL08a]. Integral [AGH00, CHL06a, HO08a, Hel09b, HLS01, Mai01, Stu01, SG03b, AvdB04, AD03, BO04, Bot06, BT09, BEPT09, DBF08, EG08, GGS09, Gla05, GPVB07, Gui03, Hof04, HLX06, JA08, JR03, JR04, LN09, MR05, SB06c, Ten03, TC09a, TdAAP08, VGZB09, WXG07, XSG04, XSG08, YBZ06, YH07a, YLA08]. Integral-differential [SG03b]. Integrals [GM01c, Saf00, Saf02, SS01b, BLL03, GvH06, MG07a, MT04, SB06a, Wen07, Wen09]. integrate [CSC^{+}08]. Integrated [Liu09a, ZRR00, Xia04, XAI06]. integrating [JBF07, Kro05, SK06]. Integration [BKR^{+}01, BCVK02, HF01, LBV01, Lou00, MCCT02, PWW00, San01, SDD07, WDM01, Bal08, BSW05, CP03a, COR08, CSML06, DEHL06, FG06, HBHS09, JMC03, KEB^{+}07, KCMM03, KLM07, Low04, MRRS05, MELD08, MG07b, MT04, MG07c, MDM03, NWZL08, OS04, OK06c, PH09, RBSL06, RMGK04, RSO04, SV07, SHPC09, Tok06b, VVM05, WG08, Yeh07, dSM05, vZdBB07]. integrations [ZHSS09]. Integrator [LR01a, KSHS08]. Integrators [IKS01, BIS07, FDL08, MGS09, QM03, SW08b]. Integro [HR01, Chu09, IDD04, SKW05, VVS08]. Integro-differential [HR01, Chu09, IDD04, SKW05]. integro-moment [VVS08]. intended [DDFT09]. interacting [AGW07, DDD05, SKW05, YB06]. Interaction [Pir02, Sur05, ZD00, BEE06, BQQ09, BGS08, CDDH07, CWL08, FM04, FG07, GHB03, GGCC09, HK04b, HAI09, KYK07, LDL^{+}09, LMZ^{+}08, LZH^{+}07, MMS04, Pap08, QFR04, SK05, SM06b, TZ07b, XCY06, XW06, YM07, YF09, ZD05, vLAvdV06, vZdBB07]. Interactions [Han01, LTD07, VR02, AK06a, AMP09, BL08, BBW06, CC05, Eld08a, FT09, GH02, KM08b, LWP^{+}09, LJK09, LKMU05, Mar06, MSP^{+}06, SPT05, SL07a, TS04, YS09, Yu05b]. Interconnects [AIRY01, AIR03]. interdisciplinary [Ler06]. Interface [AMSZ07, ČPT01, CBI^{+}04, DDS09, EFFM02, GW01, LS02b, LL01b, PL01, RM01b, SW00, SAM05, TC02, UMRK01, Wu01, YSC01, AS07, AMS04, Ber04, BR09b, BN09, BW07, CET09, CA06, CDDH07, CS08a, CXZ09, CS07c, CB09, DR06, DDM07, DIL03, DP09, DSS07, FS04, FGS09, FCT07, GMD07, GCNB07, GAC^{+}09, Her05, HK04b, HKAH06, HAI09, JJGL06, JJGL07, Kim05, KPP07, KPP09, LLP07, LKP06, LZ09b, LJS08, LF05, LKY03, LKMU05, LHGF04, MKLU05, MDB^{+}08, NT07, NLT08, OSK09, QS07, QLS09, RMB07, RRV06, SGFL09, SYC09, SS07b, SDT08, SB07, SSH^{+}07, TLL^{+}08, TLK09, TU04, XW06, Xu08, YU05a, YJL^{+}06, YS09, YZW07, YW07, ZL08a, Zho07, ZW06, ZZFW06, vEB05]. interface-sharpening [CET09]. interface-tracking [BR09b]. Interfaces [ACK02, Gla01, SZ00, Str00, Str01a, TK02, AS03b, AMS03, BS08b, BL09b, CD03, DS08, GLL03, Gre04, Hel05, HL05, HST09, Jia07, JY08, Kro02, LL07, LP04b, ML05, OK06a, PP04, Sam09, SPB09, SS08, TLK09, TT09, UTBV03, XD07, YJ06, YZW07, ZW04]. Interfacial [Dim07, Poz01a, THL06, BB04a, CL03b, Fan08, FCD^{+}06, HW08, JL04a, KYLB07, Kro01, LTH08, MT08, NLT08, Pop09, Tan08, WB09a, XLLZ06, YZF^{+}06]. Interfacing [FK09a]. intergranular [AIR03]. interior [Bor07, GLLN07, GLLN09, HH08, ML05, Mil08, Sha05]. interlocking [SK07b]. intermolecular [SB09]. Internal [HLKS00, BBC^{+}06, Hew03, HK04c]. Interparticle [PM02]. Interpolants [GW01]. interpolated [IAT08]. Interpolating [FH03, CGSR08, HL04]. Interpolation [BR01, GC02a, Gui02, LSV09, Par02, TK00, YXU01, BP06, CP04a, CL08c, CLS09b, MS03, SZ05, TB09, TW03, UYK^{+}04, WG08, WG06, ZW06]. Interpolation-free [LSV09]. Intersecting [RMO00]. intersections [LJSM08]. Interval [JMK01, BRB03, MK07]. intravascular [ZEA06]. Introducing [BBR01]. Introduction [Bra04]. inundation [Geo08]. Invariant [BU02, KE09, vBK03]. invasion [SL07c]. Inverse [BHL07, Bor00, CSV00, IFZ01, LR01b, TK02, AvdB04, BL09a, CT04, CBGI09, CDR09, GKL03, HO08a, Hoh06, JS07, KGJ05, KNH05, Lee07a, LAKD08, MNR07, MN09b, NLT07, VP09b, YYF09, ZG08]. Inversion [JZ08, Mac00, CSMH05, DD05, HPS06a, HS07a, Kry04, PSH^{+}08, RR07]. inversions [Tol08]. Investigation [APQ02, BCZ04, MP03, CD04, CET09, KY08, LC06a, NTB07, PSCQ03]. investigations [TB04]. Inviscid [CKR00, CKR01, VD02, Xu01b, vdVvdV02, BDHN09, BP04b, BB08b, CLB08, CTT08, DF04, DS09b, FCT07, GXW07, GGF03, Kro02, NOG08b, NJX08a, NS04, SDT08, Xia04, Pro05]. Involving [KJ01, OS01, LKP06, MS08a, MR07a, ND04, SC09a]. Ion [OL01, ED07, GLS03, GIA^{+}07, GBB^{+}06, Hum05, SDD07]. ionic [BS05, DC07, XDC09]. ionization [RHPN09]. ionized [Kwo08, MD04]. ionizing [AM05]. ions [BBF^{+}08, Kwo08, XJ07]. Irregular [CL00a, Cal02, GFCK02, GC02a, LFK00, MCJ01, TS02, vdSE00, BF08, BP07, CGRGV^{+}04, HWW07, ILL09, JM05, LF05, MG07c, SROCdPFF05, SS08, VTT08]. irregularities [HMM08]. irrotational [CL06a]. ischemic [NLT07]. Isentropic [LKX04, HH06]. Ising [FGOV00, FV01, PVPS09]. Island [Cho00]. isochoric [XMP07]. ISOD [RBH03]. isolated [KK03a]. isopycnal [WAH09]. isosurfacing [Min03]. isothermal [DHOT09, Mig07]. Isotope [OL01]. Isotropic [BC02b, CL08c, Kum04a, LP02, KMSH08, YSO07, YGL05]. Isotropy [Hua01b, ZSC07]. isovalue [RBH03]. isovalue-oriented [RBH03]. ISPH [XSL09]. issue [Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano08-35, Ano08-36, Ano08-37, Ano08-38, Ano08-39, Ano08-40, Ano08-41, Ano08-42, Ano08-43, Ano08-44, Ano08-45, Ano08-46, Ano08-47, Ano08-48, Ano08-49, Ano09-49, Ano09-50, Ano09-51, Ano09-52, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09-57, Ano09-58, Ano09-59, Ano09-60, Ano09-61, Ano09-62, Ano09-63, Ano09-64, Ano09-65, Ano09-66, Ano09-67, Ano09-68, Ano09-69, Ano09-70, Ano09-71, Ano09-72, HJJ09]. Issues [Ano00q, Ano00r, Hun01, FL06, Thu08b, WZL09a, WZL09b]. iterate [AMXL09]. Iterated [SS09c, Wel07]. Iteration [Boy02b, Yan08, Küm04b, LY07a, LY07b, TB09, ZSTC06]. iterations [BEPT09, MYW07, Ovt08, YLD09]. Iterative [AvdB04, BS00e, GZ01, HBHJ08, Man02, RW00, Vos06, ZZ01, BFG08, CL03a, CL08d, FWP09, GV06, Had05, KKS05, LT05, LH08b, Mil04, Mil08, OMK09, Thö04, Tok06b, TDV06, WW04, Yon01, YHCD05]. iterative-perturbation [YHCD05]. IV [WZL04].

J [ABRR09b, CL08b, DD03a, HMS08b, HY11, HLWW06, JJGL07, Lau06, LM03a, MKM04, MN17, Mil07, SM09b, SCC^{+}03a, WZL09b, dTWD09]. Jacobi [BPS03, BL03, CQO04, CP06b, CC07, CS07b, CY05, CL00b, FF02, Had05, KOQ04, KOQ08, KR02, KT00b, QS05, SR09a, TTZ03]. Jacobian [CBKM00b, CZ09, CSV00, KK04, KMS02]. Jacobian-Free [CBKM00b, CZ09, KK04]. Jacobians [Chu09]. jammed [DST07a]. jamming [DTSC04]. Jet [ZYKW01, DB04, LLC06, YFLS06, YSS05]. jets [Cha09]. jetting [YSS07]. Joint [JPMC01, AJ09, RJ06]. Joule [DMR09]. jump [MDJS07, RC06]. jumping [LHZW05]. jumps [RAB07]. junction [LMH07]. June [Tol02a].

Kac [BLL03, PWW00]. Kalman [IKL^{+}08, KFH^{+}04, KFIG06, LX09]. Kantorovich [DCF^{+}08]. Karhunen [BP04a, ST06, ZL04]. Kármán [YKG04]. KdV [Dur08]. Keller [PS07a]. Kelvin [CPG04, KTD03, Pri08]. kernel [DLD^{+}06, MRRS05, SL06, YBZ04, Yin06, YH07a]. kernel-free [YH07a]. kernel-independent [YBZ04]. kernels [CGSR08, GvH06, HX05, Lau04, TMND07, WG06]. Kerr [de 00]. Kershaw [FM08]. Kind [GST02, Gui03, JR03, JR04]. Kinds [Boy02a]. Kinematic [TFD06, BK07, LLZ07, Pon05, XSG04, XSG08]. kinematics [BST03, LTC07, Liu09a]. Kinetic [CKR00, CKR01, CL01a, CHBS04, Del02, DQ04, FL06, HK04c, KK00b, KQW03a, Lap02, LM08a, LX00, Ohw02, San01, TX00, TRL01, Xu01b, Xu01c, Xu01a, Xu02b, ACGV07, BBHM09, CLL07a, CBC09, CV06, CP03c, CELS07, CJ09, CDL04, CDL05, DJM05, DDM07, ELVE07, Fox09, GC06, GBB^{+}06, HM09, HS07b, JX07, KK05a, KQW03b, LGKP07, LZ04, LF06, LZ09c, Liu08, MMKP08, MSJ07, NJX08b, NJX09, OX04, OK04, OF06, QW05, QA09, RCT07, RSM05, RS06b, SSE03, Sch08, SHY07, SY08, SA06, SS06b, SC09b, SK07b, TDWY08, TXCD07, TKH09, VK05b, WXG07, XH03, XMT05, XHC08, YHSX07, ZSB^{+}08]. kinetic-fluid [CDL05, DDM07]. kinetic-hydrodynamic [CELS07]. kinetic/fluid [CDL04]. Kinetics [MOvL00, SD00, BHL^{+}04, LGP09, Lap03]. kinetics-based [Lap03]. Kirchhoff [GPL05]. KIVA [TT06c]. KIVA-4 [TT06c]. Klein [BY07, HZ08, HJL09, RBK09]. kMC [RMGK04]. knot [JC06a]. knots [MR03]. known [KZWY09, Lab09]. Knudsen [KPB08]. Korteweg [CkM07, LGK06, LY06]. Krig [GSK06]. Krylov [BB07a, BT02, BEPT09, CBKM00b, CS08a, CZ09, FWP09, HJM07, JH08, KM00, KK04, MYW07, MKR00, NOG08b, SNGAS04]. Krylov-accelerated [CS08a]. Krylov-Based [BT02]. Krylov-subspace [BEPT09]. Kuramoto [CFP08]. Kutta [HyLL07, ZP06, AHNS09, Bal08, BP09, BSB01, CFR04, Dri02, HL06b, KCGH07, KHV01, KWD07, KDW08, LX07b, QS04, QKS06, QLK07, Rei00, STR07b, Tan05a, ZQSD08, ZQ09].

LA-UR-03-3852 [Har04]. laboratory [BvdHKG07, Har04]. Lacunae [QT08, Tsy04]. lacunae-based [Tsy04]. Laden [WK01a, JD09]. lag [MKKY06]. Lagrange [BG05a, DDK06, DLMK04, Gir00, HB02, SPT05, VMN07, ZSP08]. Lagrange-distributed [WZ07, Boy06]. Lagrangian [NTYT02, Tol02a, AA07, AH08, AEP04, ALGM01, AHMS03, BG07, BBC^{+}06, BS08b, BS03a, BLG^{+}08, BR09b, Bon00, CRB00, CL06a, CFF07, CDDL09, CS07a, CJR04, FF02, FP08a, Fed02, GT09a, GT09b, GXW07, GHB03, GPF03, GCCD07, GBB^{+}06, GD05, HK05, HK08c, HPZ01, IX07, Jao07, KMSH08, KMS02, LS03, LHZW05, LCS09, LY04, LS05b, LC06b, LSW08, LRS09, Mai09b, Mai09a, MN09a, MGGH00, MP08, Mea04, ML01b, MDM03, NSS03, NTYT01, OF02, RB05, RBS06, RWWS07, RCB05, SM09a, Str00, Str01a, TOY09, Tol02b, TJLT08, XY01, XK01, YSO07, YA05, YFBH07, ZWS07]. Laguerre [BS08a, BRB03]. Laminar [BCVK02, VBL03, BC08, BEG03, CFL^{+}03, FGP08, GLLX08, LLC06, MR04, MAL09, RFVP09]. Laminates [Wee02]. Lanczos [CKLS05, Bor00, BS05, JHZ^{+}09, SHS08]. Lanczos-type [JHZ^{+}09]. Land [GKL00, KJ09b]. Landau [BCG09, BC02b, DDG02, DDFT09, FP02, HZ02, Lem00, PRT00, RSS09, ZZ08, dSM05]. Lane [PSD09]. Langevin [BLW04, DEHL06]. LANS [HHPW08, PHW08]. LANS- [HHPW08, PHW08]. Laplace [GF05a, HZ07a, HSQ03, HW05, Kry04, SSN09, SY09b]. Laplacian [AHPT07, PAD07, Pon05]. Large [ATV01, BADG00, Ben02, Bor00, DF00a, ELD08b, FLG01, FG02, Gui02, KS02b, KK00a, KDC05, LLQ^{+}02, ME09, PPC00, TSB01, TR02a, ZWL02, AHNS09, AL06, AD04, BPS03, BBB08, BS03b, BSW05, BTWGvBW07, CF06a, CGDT09, Cho05, CM03, CSKD05, DT03, DSS07, DS09a, FDD09a, FDD09b, FH03, FKK08, Gra06a, Gra06b, HBLD07, Heu03, HP04b, IOTK04, KSJ03, LZL03, LVL05, LP06b, Liu09c, LDV08, LJ07, MCM04, MLM09, MGS07, MDM03, MBP07, MMPB07, MV08, MHdB07, NLF03, PDHP07, PYC04, PM07, RMG^{+}09, SSW^{+}07, Soc03, SFMP06, TSB03, TMD07, Tok06b, TC09b, VK09, VTM^{+}08, XLP05, YZ07, YB06, ZSC06]. large-amplitude [CF06a]. Large-Eddy [KK00a, ME09, PPC00, TSB01, KDC05, BBB08, CM03, CSKD05, DS09a, FDD09a, FDD09b, HBLD07, LDV08, MCM04, MLM09, MGS07, MBP07, MMPB07, PDHP07, SSW^{+}07, TSB03, YB06]. Large-Scale [ATV01, BADG00, KS02b, ZWL02, BSW05, BTWGvBW07]. large-step [AHNS09]. Large-Time-Step [Gui02]. Large-Wave [DF00a]. Laser [DGH02, DNS08, DDGS09, GHB03, HDBW05, KSHS08, LDL^{+}09, Sau04]. laser-plasma [GHB03, KSHS08, LDL^{+}09]. late [CL07b]. lateral [KJ09b]. Lattice [BTC05, BLW01, BdLL01, Del02, DC02, FH00a, FH00b, FSM^{+}01, GS03b, Gua00, GSW00, HDC02, HHL00, IYI^{+}02, LL03a, LLQ^{+}02, MSYL00, MHS02, MAL09, PR00, RMSB09, SS05c, Sun00, VLB09, XH03, vdSE00, AST09, AL08, ABZ^{+}08, BKS07, BYS08, CA06, Del03b, DCK08, FG05, FM04, GM04, Gos04, GM06, HvHHS05, HNGB04, HHC08, ISNY05, IOTK04, IF09, JKL05, KY08, KMV03, KPB08, LLP07, LL03b, LL05, LLC06, LT09b, MRS09, MR07c, NCS03, PL09b, PSCQ03, PSC04, PSC^{+}06, PA07b, PPB09, RSM05, SCT09, SPT05, SLC07, SS03b, Sof09, TBJ^{+}09, VCG03, WWC07, WS09, YZ07, YGL05, YF09, ZK05, ZSC07, ZSC06, ZTPM05]. Lattice-BGK [FH00a]. Lattice-Boltzmann [HHL00, BKS07, PA07b]. Lattices [vdSE00, CLL07a]. Law [FGG01, VPA02, De 04, ÉGP09, FS06, GD07a, LLZ07, MY06b, Mil06, Mil07]. Laws [Asl01, BJ00, CDKP00, CRD02, FMO00, GC01, KT00a, LL00, Noe00, Sti02, TS01, Wan02, WL02, YL01, AKLMP09, BAFL09, BBCT09, BCCD08, BP03, CLG07, Cap08a, Cap08b, CT08a, CP08, CGKM06, CD07, CkM07, DET08, Edw06, FS09, GV07, Gui05, Hub07, JR09, JTL09, KI05, LL03c, LVW06b, ML08, Mil04, PDL09, RLZ03, RCD05, SW04b, SYG06, SWL06, SZLW06, SR09b, Tak06, TT04, TT05b, TT06a, TT06b, THS07, VCZS04, WZL04, WG09, vDZ06]. Lax [KOQ04, LCS09]. Layer [DC01, Hu01, Hu05, Str01b, Vay00, Vay02, AK06b, AC09, BHNPR07, CGRGV^{+}04, CLL^{+}07b, GKD09, HO08b, HLL08, MT07a, NK08, RJ04, ST04, Tau07, Zhe07, ZT03]. Layered [Hig02, AC05, And09, BFT09, Hig05, IQT08, SCT06]. Layers [Bal02, ELW01, GZ01, PPC00, TC01a, BFJ03, DH07, Doh09, FE04, GGOB04, LP06b, Nat06, Rah04, SJHM09, SP05b, ZGG03]. LBB [AGP01, CHPR09]. LBE [GLLX08]. LBM [SL07a]. LDAF [WZ07, Boy06]. leaf [Dic08]. leaky [ZK05]. leaping [BCK09, RE07]. learning [Kou07]. Least [Cap09, PG02b, AMSZ07, BT05, BT06, BP04a, CSO09, DI09, GS03a, GNNB08, HV03, HK08a, HMMR04, HLMM07, HdGKG08, HY09, HY11, KH09, NCS03, PR03, PR04b, Pon06, Pon07a, Pon07b, PR06, SL07b, VB09, ZKY05]. Least-Square [Cap09]. Least-Squares [PG02b, AMSZ07, BT05, BT06, BP04a, DI09, HV03, HK08a, HLMM07, HdGKG08, HY09, HY11, KH09, PR03, PR04b, Pon06, Pon07a, Pon07b, PR06, ZKY05]. Lebesgue [Hei05]. Legendre [AQ00, APQ03, Boy03, Boy04, CDI09, KOQ08, KT03, PSD09, SS01a]. Legendre-pseudospectral [Boy03]. Legendre-transform-based [KOQ08]. length [AKP07, JG09]. Level [Asl01, BCMO01, CT04, CBGI09, CMK^{+}01, CBMO02, Cho00, EFFM02, HMS08b, Hig02, KAIN01, KLvBvL02, LLdlP^{+}00, MS08a, OF01, OS01, OCK^{+}02, PS01, SW00, Set01, SJ02, SP00, TMB07, AS03b, AS05b, AJT04, ÁDIM09, AA06, AHMS03, BHR04, BHSV07, COQ06, CM06, Che07, CSL08, CCT05, CQRW05, CC08b, DMBS05, DMP08, DL03b, EHST03, EHS^{+}08, ETT05, FSS03, GGS09, GCNB07, Hab04, HMS08a, HKO07, Her05, Her08, HK05, Hig05, JVVS07, JCT07, KH07, LW07, LW09, Liu09c, LTWW07, LLC^{+}08, LTL^{+}09, LTM09, ML06a, MRC06, MR06a, MGCR07, Min04, MG07c, MG07d, MV06, NJLA06, NLT07, NT07, OK05, OKZ07, PHKF06, QL04, RR07, SS06a, SYC09, SAKDJ05, Sme06, Spe05, Sus03, TZ06, TZ07a, TZ07b, TBJ^{+}09, Tow07, TU04, WLKW07, WSTW09, WYS09, Wen09, XLLZ06, YJL^{+}06, YSS05, ZGK09, ZLAC05, ZL08b]. level [vdDA06]. Level-Set [Asl01, CBMO02, Cho00, KLvBvL02, PS01, CBGI09, AJT04, ÁDIM09, AA06, Hab04, KH07, MG07c, RR07, Spe05, XLLZ06, YJL^{+}06, YS09, ZLAC05]. Level-Set-Based [LLdlP^{+}00]. level-set/volume-of-fluid [YJL^{+}06]. levitation [IM05]. Lévy [LLTA07, PC06a, Pav07]. Li [GIA^{+}07, MCP03, GIA^{+}08]. Li-ion [GIA^{+}07]. library [SWB^{+}06]. lid [AK05]. lid-driven [AK05]. Lie [San03]. Lifshitz [dSM05]. lifting [KRT^{+}09, WG09]. Light [KL06, SS01b, deM02, BMK^{+}06, BBK06, GGRS08, JD04]. Light-Cone [SS01b]. Light-Emitting [deM02]. Lighthill [ZSWW03]. LIGKA [LGKP07]. like [DLS^{+}00, HO06, LNXNTX09, MEG02, Mil06, Mil07, PL09a, SB06a]. likelihood [Sti05]. Limit [BKR^{+}01, DW00, Asl04b, BPM06, Boy05b, CWL08, CS04, CDV07, DP08, FPK08, JLOT05b, JLOT05a, LW09, Lur07, PSZ09, SD05a]. limit-cycle [BPM06]. limitations [CP06a]. limited [BDS07, CTT08, GD06a, LXM09, LLGL07, Ols07]. Limiter [BSB01, RM01a, CT09, CS08c, KT04, LBL07, MOG09]. Limiters [BL01, Kri07, Kuz06, NJX08a, QS04, Sof09, ZQSD08]. limiting [Bet08, KK05d, ML08, YKK08]. Limits [MHS02, Del03b]. Line [Gui02, POS00, RRL01, Khe04, MBS03, VP09b, ZGK09]. Linear [AL01, Ben02, BBCT09, CP00, KKGL01, Mav02, MYW07, NC01, QRHD00, RC00, SZ00, TS02, WC01, AMR06, AC09, BAYZ08, Bal08, BAR08, BDRT09, BM05, BB07b, BdCB09, BDCG03, BSP06, BCI^{+}08, CFS09, Cap08c, Cap09, CWJ07, Cha07b, CN05, CP06b, CP06c, CFJ09, DK06, Dem04, DTSC04, DC08, DK07, GZ08, GR04, Hau08a, Hau08b, HK08c, HR07, JHZ^{+}09, KT05, Lab09, LGKP07, LZL03, LM08b, MPD08, MGS09, MJ06, NPH09, Ngu07, NPC09a, RH05, Sam09, SDM04, SLG^{+}03, Thö04, TT05a, TT05b, UL06, VCT07, WT07b, XLS09a, YJ06, YKG04, YH07b, AGT02]. linear-scaling [SLG^{+}03]. linearity [KSW03]. linearity-and-bound-preserving [KSW03]. Linearization [GV02, Hun01, Kuz09]. Linearized [Hu01, MT03, MDR07, SM06a, BKST09, Hu05, Nat06, PGN08, Rah04, SB09]. linearly [BAR08, CJR04, Jao07, Tok06a]. linearly-perturbed [CJR04, Jao07]. Lines [DK02a, KKGL01, AINR03, Car09, JH08, Spe05]. Lineshape [KHV01]. Link [NTO^{+}07]. linked [KM08a]. Linux [CD04]. Liouville [JW06, JY08]. lipid [FK06, MK08a]. Liquid [DSS00, EKK02, JLCD01, LS02b, CPR05, Cha09, Chr03, CB09, DDK06, GGS09, GKV09, HP04a, HL07c, IM05, LMV04, LS08, LL06a, LR07, LLZ07, LL07, LL08a, RGS04, VGL^{+}07]. liquid-liquid [CB09]. Liquid-Vapor [JLCD01]. Liquids [KS02b, HSL08]. list [DTS05a, DTS05b]. lithography [BBK06]. Load [DPR00, JJGL06, JJGL07, MG05a]. Loading [CVB00, GVT01, KFV07, Li08b]. LOBPCG [HL06a]. Local [Alb00, BC02a, BS09b, DI02, GTD01, JL02, LSY04, MTH08, Min04, MHS01, MV08, OV00, OMG02, QS02, RC09b, SC08a, VDM^{+}02, VC00, XXS07, XS05b, YZW05, AMR06, BC05, BBD04, BF08, BG05b, CBH03, CFR08, CLS09b, DSM09b, DGRS08, HMOG08, HZ08, HAD06, ISNY05, JW06, KB04, KKO04, Lap04, LSA06, LSJA05, LY06, LGM08, Ma05, MCGV04, MPFC08, Pav07, PVR07, SLG^{+}03, SPLM09, SRX07, SR09b, Tsu06, UBRT07, UPKN09, YE07, ZHSS09, dTDI^{+}07]. local-orbital [SLG^{+}03]. Localization [ZGG03, Gra06a, HNF07, KZWY09, Lar09, ST03b, TB09]. Localized [KL08, DDH05, DLD08, YA05]. Locally [BS08b, CLS04, FWW04, HEML00, Str07a, AT08, FHW07, OK07b, TZHT04]. locally-conformal [OK07b]. Locally-corrected [Str07a]. locally-refined [FHW07]. Locating [TK02, SS09b]. location [HSZ04]. locking [LSJA05, MP07a]. locking-free [LSJA05]. locomotion [HSW07]. Loève [BP04a, ST06, ZL04]. Logarithmic [Mai01]. logging [GH08b]. Long [FPC^{+}00, Wee02, ZSW03, CWL08, FT09, HPS^{+}06b, LLL07, SK04a]. Long-Range [FPC^{+}00, FT09]. long-term [SK04a]. Long-time [ZSW03, LLL07]. long-wave [CWL08]. Loop [SS01b, GPL05]. loosely [GGCC09]. loosely-coupled-type [GGCC09]. Lorentz [Tót02]. Lorenz [FVE04]. losses [HR08]. lossless [LKD04, Rem06]. lossy [LZC04]. Low [BKR^{+}01, BISS01, FS01, FH00b, Nic00, POS00, RV00, SC01, WPM02a, Ano04z, AG09, BCDW06, BDHN09, BDR^{+}04, BBMB07, BO09, BB04b, BEG03, BB08b, CLB08, Cha09, DH04, Del07, DBBP08, DKS^{+}03, DST07b, EG08, HH07c, HK04c, Kok09, LG03a, Lee07b, sLwG08, LM08c, LQ06, MEKS03, MDR07, NMM^{+}07, NMH^{+}07, OTCM08, OVG07, PDHP07, RVDM09, RB09b, SDGX07, SM06a, SMS04, Soc03, SFMP06, TSG^{+}06, TMD^{+}08, VGCN05, XH03, XLP05, YS07a]. low-cost [LQ06]. low-diffusion [MEKS03]. Low-Dimensional [RV00, VGCN05]. low-dispersion [Kok09]. Low-Energy [SC01]. Low-Energy-Density [BKR^{+}01]. low-frequency [BCDW06, DH04]. Low-Mach [Nic00, LG03a, RVDM09, XH03]. Low-Order [BISS01, AG09]. Low-Speed [FS01, SMS04]. low-variance [HH07c]. lower [MM09]. Lowest [Mit00, Lab09]. lowest-order [Lab09]. LU [LZL03]. lubrication [DM03]. Luo [TK04]. LWS [DF00a].

M [LM03a]. MAC [IQ08, LW01]. MacCormack [HT00b, HT00a, HF01]. MacCormack-Type [HT00b, HT00a]. Mach [NVD05, VSW06, Ano04z, BDHN09, BDR^{+}04, BTW03, BEG03, CLB08, Del07, DBBP08, DST07b, FH00b, KKM08, LG03a, Lee07b, sLwG08, LM08c, MDR07, NVD07, Nic00, RVDM09, RB09b, SBGK00, SM06a, SSD00, SFMP06, TSG^{+}06, TMD^{+}08, WPM02a, XH03, XLP05]. Mach-uniform [VSW06, NVD07]. Mach-uniformity [NVD05]. machine [Kou07]. macro [BEA09, FT06]. macromolecular [Lap03]. macromolecular-crystal [Lap03]. macromolecules [IH04]. macroscopic [GMAj09, HA06, ZRS06]. Magic [MT01]. Magnetic [Del01, GG00, SHWW00, AvdB04, Bal09, BCDW06, BCM^{+}07, DDSV09, EPW08, HR08, IM05, IDD04, JOS06, KM08b, KB04, LKD04, LCG07, MSP^{+}06, NMM^{+}07, NMH^{+}07, PH09, SDGX07, SHS08, SS04]. Magnetically [OL01]. magnetised [GYKL05, GLT07, GL09b, SG06]. magnetization [dSMF09]. magnetized [GGOB04, Mot08, PPCW06, UPKN09, VVM05, VTC^{+}07, XCRX08]. magneto [Li05, Li08a, VOD08]. magneto-hydrodynamic [Li08a]. magneto-hydrodynamics [Li05]. magneto-static [VOD08]. magnetogasdynamics [Gom08]. magnetohydro [LW04]. magnetohydro-dynamics [LW04]. Magnetohydrodynamic [DCV^{+}01, BT07a, FJ09, GLN06, JBF07, KLM05, LTC07, Liv07, MV06, OPML07, ODCK07, Ser09, SK07b, dCNHSD07]. Magnetohydrodynamics [Bal01, Del02, GTD^{+}02, Jan00, TX00, Tót00, Asl04b, Bal09, BRDM09, GLLN09, HT07, LFSS07, LD09a, LD04, MK05, NN09, PCP08, QW05, RSW06, SGG^{+}04, Tor03, TMG08, Waa09, ZK06, Zie04, vDZ06]. magnetorheological [KM08b]. Magnetotelluric [HS07a]. Maintaining [PHKF06]. maintenance [GS09d]. Management [OK04, TS07, WGCR07]. manifold [GKE04, HE07]. manifold-mapping [HE07]. manifolds [BRC^{+}09, BBK07, KG03, MSO04, RBL04, SK07a, VGCN05]. mantle [FKK08, KKS05]. manufacturing [SS08]. Many [ZD00, CLMRP08, FHLK05, LM08a]. many-particle [FHLK05]. map [Gui03, HW05, Hel09a, dFJS09]. Mapped [HAP05, BRB03]. mapping [HE07, LQX06, MSO04, ZL08a]. Maps [LTZ01, TB00a, BBK07, YLA08]. Marangoni [LS02b, TC02]. Marching [Set01, LG03b, LG04, YBS06]. marker [AMS04, CB09, RB05]. markers [AMS03]. Markov [CVE06, GL09a, MDJS07]. MAS [KHV01]. masks [BBK06]. Mass [Lio00, OF02, BYS08, BT05, DBS06, HLMM07, KH09, KJ09b, LLGL07, RC09a, YZF07, ZH04]. mass-conserving [DBS06]. Massive [DPRS01]. Massively [KP00, SLG^{+}03, CHB09, HVAC09, KRT^{+}09]. MAST [AT09]. master [HL07a, IM07, MK07]. Matched [Hu01, Hu05, Vay02, YZW07, BFJ03, BHNPR07, CLL^{+}07b, DH07, Doh09, GKD09, GGOB04, HLL08, Nat06, OK07b, Rah04, RJ04, ST04, SP05b, YW07, Zhe07, ZW06, ZZFW06, dHRvdB07]. matching [Bor07, JJGL06, JJGL07, LVL05, NDT06, SB06c, ZW04]. Material [Bar02a, DDG02, ZZVM08, AS03b, AS07, APT09, BSKH07, BFT07, BG09, CD03, DS08, GA09, Khe04, LKY03, LBL04, MU09, Ols07, PP04, SGFL09, SC09a, WG08, XD07, ZC09, ZW04, ZD08]. material-order-independent [SGFL09]. Materials [CL00b, CB02, EH02, GM01b, HLS01, OV00, RV00, BZ04, CDS04, CP04b, CP05, EPW08, GFS08, GL06, GMO04, YU05a, Zad08, ZB07]. Mathematical [Ano04z, BTFY01, CHM08, GS02, HM08, HSW07, RBT03, LD06, SMP09]. matrices [BPS03, BT07a, DBB06, WR09]. Matrix [Bor00, BS00e, Edw00, Lin01, PCS^{+}09, PC02, SWTM01, Yon01, YMF01, Chu09, DMG04, HJFW04, HL04, Hau08a, Hau08b, HSZ04, LAKD08, LBS^{+}04, LZH^{+}07, PSH^{+}08, SBA07, UL06, Wal03]. matrix-free [LZH^{+}07]. Maximum [GG09b, Rom02, Abr06, Abr07, Abr09, KSS09, Sti05]. Maxwell [ACS00, ACLS03, BL04, BL09a, BLG^{+}08, BHvdV06, BS06b, Bra08, CD03, CW03, CXZ09, CJ07, CLS04, CFP06, CFJ06, DR06, DD05, DDH01, DLP08, DF00b, DDFT09, Eli03, Eli07, FH03, GD07b, HH07a, Hag07, HK04a, HR08, HLO08, HW02, HJM^{+}05, HGB^{+}03, HMM07, LZC04, MCCT02, MPFC08, MOS^{+}00, Nys02, ON08, PAD07, Rem06, RLB02, RRW05, RB02, SZB^{+}07, SWZ03, SL07b, SA09, SP05b, Tsy04, VPMC04, WZ02, Wel07, XCZ02, XD07, YP01, ZW04, ZT07b, de 00]. Maxwellian [CVB00, GW06]. MCC [CN08]. ME [FWK08]. mean [CB07, DSM09b, Hu05]. Means [BM01a, BHR03, BHR06, Heu03, MG05a]. Measure [RS02]. Measurement [TG04, CP06a, MSB07a]. measurements [CSC^{+}08, DGF09, HKS09]. Measures [OB06]. mechanical [DDD05, DP08, NTO^{+}07]. mechanical/molecular [NTO^{+}07]. Mechanics [Bar02a, BS01, BS00d, OP02, Bod06, BG05b, DF07, FDD07, GPL05, HLRZ06, Kou07, LNXNTX09, MK08a, Mil04]. mechanism [PK03]. Mechanisms [KLN^{+}01, RRV01, LTWW07]. Media [CS01c, CS00, CGSS00, LMSV00, LLN00, WLE^{+}00, ZF02, AT09, AZ06, BQQ09, Bar04, BFT09, BS06a, BH04, CD03, CJSS08, CPG04, CDE06, FWP09, GZ07a, GZ09, HJ09, IQT08, JLT06, KSH^{+}06, KT06, KT07, LTZ03, LMS08, LH05b, LJ06, MZ08, Mar06, MJT06, MN06, MHI08, MP05, MGS09, NL08, PC06a, Rem06, TJLT08, XD07, YE05, YE07, YH07b, ZL04]. mediated [MSP^{+}06]. Medium [CY00, BL09a, BW07, FG04, FG05, GS09a, Hoh06, KK03a, QLK07, RM07]. MEL [Wan05]. Melt [LS02b, ZGT06]. melting [Men04]. melts [HLFB07]. membrane [CKPW07, LWP^{+}09, LS08, MK08a, MSP^{+}06, TLL^{+}08]. membrane-mediated [MSP^{+}06]. membranes [DLW04, DLW06, FK06]. memoir [Bra04]. memory [HJFW04, LH05b, TS07]. MEMS [AA09, MK04b]. MEP [Rom07, TPR05]. merging [Hew03, QLS09]. meridional [TVMR03]. MESFET [GS06a, Rom02]. Mesh [Alb00, Bal01, BV05, BMR01, BMRS02, CH01, CBL01, DGH02, Dys01, FR02, Hua01a, Hua01b, ID04, LTZ01, LTZ02, LK09, MR00, MR02, MP07a, MGGH00, Mav02, OGV02, Per00, ZSP02, Zha02, AZB09, AFGM07, AEP04, BFC04a, BFC04b, BS03a, BL05, BCGR05, CR07, CGDT09, CJ04, CBH03, CBI^{+}04, CHCOB09, CFJ06, DW09, Dwi08, FL06, FYH^{+}06, FM06, HT07, HZ07a, HS06, HS08a, HG03, HS03b, Hua05, HMR08, Hum05, ISNY05, JS05, KK09, KAA^{+}07, KPP07, KE09, LC06a, LK07, LMX^{+}08, LD09a, LB03a, LP04a, LL04b, LLOT06, LKO05, MJ09a, MC07b, MSB07b, NA08, NMH^{+}07, Ni09, NLLE06, PSCB08, PDHP07, PL09b, PN03, PCP08, PL04, PC06b, QS07, QLS09, RA09, SWB^{+}06, SY09a, SHP07, SRX07, TZHT04, TLK07, TTZ03, TFD06, TK04, Wal03, WT07a, WT07b, WLC^{+}06, YMT^{+}04, YMWM06]. mesh [YF09, YT07, ZJW06, ZJWC08, ZFM08, ZSC08, vDZ06, vZdBB07]. mesh-based [SHP07]. mesh-dependent [AZB09]. mesh-free [YMT^{+}04]. Mesh-Size [Zha02]. Meshes [DLS^{+}00, Han00, Her00, MVM02, ML01a, MG02, MP01b, MHS01, TS02, Vas00, VG02, WS01, WB01, ZYC02, AK06a, AS07, AB05b, AT05b, BES07, Ber06b, BM07, Cap08a, CKvT07, CDDL09, CBGI09, CSO09, CS06, CS07d, DHOT09, DMR09, DK07, DKTT07, DBTM08, DZ09b, FM08, FHW07, GL08, HO08a, HLO08, Her09, JJGL06, JJGL07, JMC03, KA05, KOQ08, KI05, KL08, KL04, LMS04, LSS06, LSSV07, LSV09, LNXNTX09, Mai09b, MB04, MY06b, MP05, NJX08b, PS07a, RAB07, RAD07, SPM03, SP06b, THD09, TAL09, VGS04, XLM07, YA05, YS07c, YS08, ZQSD08, dVGLM09]. Meshfree [ZWL02, CYS06, KYLB07, ZKY05]. meshfree-Cartesian [CYS06]. meshing [BGR08, DS05a, YZF^{+}06]. Meshkov [LSD07]. Meshless [Ma05, BZ08, BSLN09, KJ09a, LSJA05, SB03, YYF09, YCL05]. mesoscale [GR08]. Mesoscopic [HKV01, Hor06, BL08, FK09a, HA06, ICO04, KMSH08]. metabolism [XDB09]. metal [AIR03, AINR03, MC06b, MLFG06]. metallic [ES03a, MC06b]. metals [GKV09, IM05]. Method [AKV00, Alb00, ACS00, AQ00, AP02, BR01, BC02a, BJ00, BJ02, Bar02a, BC01, BMR01, BS00c, BCE^{+}09, BM01b, BE02, CFA01, CL00a, Cal02, CHR01, Car02, CWT00, CGP02, CMK^{+}01, Cho00, Coe02, CPP02, CM00, Cor00, CB02, Dar00b, DV02, DPR00, DFT01, DGP00, Dri02, EH02, EKK02, ERT02, EFFM02, EAY01, Fed02, FBFF00, FP02, FR02, Fre00, FK02, GM01a, Gen01, GW02, GMB01, GHG01, GK01, GP00b, GBGM01, Gui02, Gut00, HHCL01, HLS02a, HMM02, Han00, HSK00, HDC02, HHL00, Her00, Hig02, HK02, HF01, HB02, HCG01, HA02, HEML00, HGM01, IYI^{+}02, IFZ01, JC02, JLCD01, JTB02, Jan00, JK02, JM00, Kan02, KS02b, KB00, KK00a, KK00b, KAIN01, KC00, KKC01, KJ01, LP00, LLH02, LL00, Lay02, LKNG01, LRN^{+}02, LL01a, LL01b, LK01, Lin01]. Method [LFK00, LS00, LMS02, LLQ^{+}02, MR00, MR02, Mac01, MD02, Man02, MKM99, MEG02, MC00a, MCJ01, MC09, MSYL00, MHS02, MC01, MC02, MKR00, MPC01, MPC02, Myo01, NFK01, Noe00, OMK09, PR01a, PKvdB00, PKP01, PS01, PL01, PB00, PK00, Pop00, PM00, PO01, QV01, QL01, RH01a, Rem00, RW00, RRL01, RR02, RH01b, RMO00, SS02, SWTM01, SY09b, SJ02, SCD00, SFW00, Sti02, Str00, Str01a, Stu01, SB02, SP00, SPW^{+}00, TK00, TX00, TMB07, TR02a, TB00b, TC01b, TBE^{+}01, TRL01, UMRK01, VB00, VCTS02, VR02, VC00, VSMW01, WPM02a, WPW02, WGCE01, Wan02, WL02, WDM01, WW00, Whi00, WA02, WS01, XCZ02, XK01, Xu01c, Xu01a, YXU01, YSC01, Yon01, Zha02, ZWL02, ZYC02, ZH01, ZTZ02, ZCMI01, ZP02, ZKK01, ZRR00, vdVvdV02, AE03, AV05]. method [AH08, APR09, AS09, ADE^{+}08, AR08, Alb08, AS05b, AJT04, AA06, AL06, AEP04, AZC05, AC05, AMP09, AKLMP09, AZ06, ADS03, ACLS03, AKP07, AMSZ03, AMS03, AMS04, Aza07, Aza09, BIW04, BIW08, BFB08, BO05, BSKH07, BHS09, BHL^{+}04, BMN05, BL04, BW06, BHL07, BS08a, BYS08, BC05, Bar04, BB08a, BGN07, BFT07, BFT09, BDHN09, BZ04, BHR03, BWLM09, BL08, BS04b, BM05, Ber06a, Ber04, BLG^{+}08, BS04c, BCM09, BG05a, BT06, BGS08, BS04d, BS05, Bor07, BDCG03, BSLN09, Boy03, BRB03, Boy05a, Boy06, BP08, BEG03, BB08b, BG05b, CD03, CDJ07, CJSS08, CCG08, CLG07, CLL07a, CP03a, Cap08c, Cap09, CWJ07, CTW^{+}08, Cec05, CR05, CR07, CFR09, CKLS05, CA06, Cha07a, Cha07b, Che03, CL03a, CP03c, Che04, CC07, CELS07, CLTA07, CL08d, CX08, CXZ09]. method [CCG^{+}06, CS07b, CS07c, CHG^{+}07, COER07, Chu09, CJR04, CS03, CS04, CY05, CFP06, CGKM06, CSML06, CC08b, CK07, CFR08, CB09, CCF^{+}05, CFP08, CkM07, Cui09, CB07, DMHP07, DM03, Dar02, DH04, DK06, DDM07, DCF^{+}08, DC07, DUEB07, DR09a, DMP08, DFV08, DTMS06, DW09, DLT09, DDH05, DLD08, DGMN03, DS09a, DG09, DLP08, DF04, Dom08, DGJ03, DLMK04, DHM07, DND06, DSB06, DD03a, DD03b, DCK08, EE08, ECL02, Eld07, EB06, EULM03, Fan08, FCJ08a, FCJ08b, FNS07, FP08a, FRS08, FS09, FS04, FEL^{+}05, FM04, FM05, Fen06, FYH^{+}06, FL09, FE04, FD07, FL07, FWR07, FHLK05, FWW04, FLZ09, FD09a, FWK08, FLM08, Fox08, FG07, FH03, FY07, FT09, FKK08, GZ07a, GT09b, GMD03, GFG09, GS03a, GS05b, GS08, GGS09, GBC06, GS07, Gel06]. method [GMD07, GFR09, GCNB07, GS05c, GG09a, GPF03, GW05, GMAj09, GWF^{+}07, Gom08, GGRS08, Gre04, GH02, GP05, GHMP07, Gri09, GKV09, GB08b, GLLN07, GLLN09, GS09c, GD06b, GL08, HPS06a, Hab04, HH07a, Had05, HS09a, HBHJ08, HJ09, HLFB07, HT07, HL04, HZGB04, HZGB05, HP04a, HSQ03, HWL08, HKM07, HPD09, HJKO08, HMPR07, HKG08, HL07a, HW05, HLO08, Her09, Her05, Her08, HAD06, HK05, HK08b, HK08c, Hig05, HP04b, HL05, HS08b, HS08c, HY09, HY11, HK04b, HKAH06, HA06, HA07, HSL08, HT03, HLX06, HSS07, HSC09, HST09, HS07b, HF08b, IX07, IX09, IK07, IKS^{+}09, ISNY05, IOTK04, IDD04, JH06, JD09, JRS05, JL04a, Jao07, JBF07, JVVS07, JLT03, JD04, JLT06, JL09, JH08, JX06, JC06a, JX07, JHZ^{+}09, JLL^{+}06, JM05, JS07, KKM08, KA05, KE06]. method [KK03a, KHdT^{+}08, KW08a, KIH09, KIHM09, KY08, KZ04, Khe04, KGJ05, KC06, KH07, KK07, KDF07, KL04, KvdVvdV06b, KFV^{+}05, KAS06, KMS04, Kok09, KS09, KT03, Kro01, KSW03, KLSW09, KB08, KSS09, KSGF09, KLP^{+}09, LKD04, LTH08, LL09, LG03a, LG09, LY07a, LL03a, LLP07, LSL08, LH05a, LFSS07, Lap03, LP07a, LGHD08, LS03, LSA06, LKP06, LWP^{+}09, Lee03, LVL05, Lee07b, LMX^{+}08, LZT09, LKE04, LTE07, LM08b, LZ09b, LZ09a, LCH03, LCW04, LHZW05, LS07, Li08a, LSZZ08, LN09, LLOT06, LJS08, LJW09, LR07, LT09a, LF05, LMS08, LSV09, LKY03, LKW05, LKO05, LKMU05, LY06, LVW06a, LVW06b, LLTA07, LV07, LW07, LP07b, LX07b, LW09, LKMK09, LNXNTX09, LS09, Liu09c, LM03b, LD04, LHGF04, LHGF05, LJ07, LS05b, LRS09, LCM07, LJ06, LBL06a, LBL06b, LBL07, LTWW07, LBL08, LMZ^{+}08, LLC^{+}08, LTL^{+}09, LMK09, LZH^{+}06, LZH^{+}07]. method [Ma05, MY06a, MZ08, Mac07, Mac03, MWM03, MCM04, MN09a, MKM04, MKOW04, MRC06, MR06a, MGCR07, MKLU05, MRS09, MvW08, MC06a, MCG08, MB04, MS04, MY07, MSJ07, MU09, MP08, MKL06, MP03, MK08b, MZ07, MESV09, Mil08, Min04, MG06, MG07d, Min07, MR07c, MDB^{+}08, MDM03, MT07b, MR06b, MMPB07, MHdB07, MK06, MT08, NLF03, NTO^{+}07, NPH09, Nas08, NA08, NBLQ09, ND04, Ngu08, NPC09a, NPC09b, NJX09, Nik06, NCS03, NG06b, NGvdWS09, NLT08, OK06a, OSK09, ODAF07, ORM06, OK05, OKZ07, OJW06, OCFF08, PDHP07, Pap08, PS03a, PPCW06, PSCQ03, PSC^{+}06, PN03, PS07a, PLS^{+}09, PFSL07, PK05, Pon09, Pon05, QW05, QA09, QL04, QS04, QKS06, QS07, QLS09, RB05, RMB07, Rah04, RAB07, RSM05, RE07, RE05, RBS06, RSW06, RMSB09, RRW05]. method [RJM07, Ros06, Ros03, RBK09, RW03, RM08, RC09b, Ryc05, RJ04, Sac07, SB06a, Sam09, SROCFF03, SROCdPFF05, Sar03, SFDL07, SZB^{+}07, SAK05, SWG08, SHS08, SBCL06, SB06b, SSB07, SWK06, SF03, SM04, SCT09, Sha05, SFE07, SH07b, SL04, SDD07, SLG^{+}03, SMS04, SY09a, She08, SPT05, SL07b, SL07a, SAKDJ05, SLC07, Shy06, SSND03, SS07b, SK04a, SHTB09, SCW^{+}09, SM06b, SPLM09, SDT08, SK04b, SCRL08, SWL06, SZLW06, SXyWX09, SR09b, Sus03, SSH^{+}07, TM07, TZ03, TZL05, TC07a, TOZP03, TLK07, TZ07b, TLL^{+}08, TLK09, TTZ03, TJS03, THL06, Tau07, TBT^{+}09, TT09, TPV07, TBJ^{+}09, TMD^{+}08, TKH09, TOY09, TW07, TC07b, TG06, TG08, TW03, TU04, TF03, UTBV03, Uhl05, UPKN09, Utn08, UYK^{+}04, VTC^{+}07, VGCN05, VW02, VOD08, VL07, VLW07, VBL07, VGPL09]. method [VVS08, VBJ08a, VBJ08b, VK05a, VGZB09, VGBZ09, VSW04, VSW06, Vik03, VK05b, VHI05, Vol04b, VCM00, WK07, WFTS05, WG08, WK05, WK04, WZL04, WW04, WL06, WT07b, WLKW07, WXG07, WTL08, WGS^{+}08, WSTW09, WFC09, WYS09, WZ09, WGRA09, WZ07, WKG06, Wen06, WWK05, WA08, WKL07, WKB07, WZ03, WMH07, WS09, XMP07, XH03, XSG04, XW06, XLLZ06, Xu08, XHC08, XSL09, XLS09b, XD07, YMT^{+}04, Yam05, YZ07, YM07, YYF09, YU05a, YJL^{+}06, YFLS06, YP06, YS09, YZLH09, YLD09, Yeh07, YC06a, YC06b, YH07a, YSO07, You06, YA05, YCL05, YJF^{+}06, YSS05, YGL05, Yu05b, YZW05, YSW06, YS07b, YZW07, YW07, YF09, YS06, YT07, YFBH07, YH07b, ZGT06, ZGK09, ZWS07, ZKY05, ZP05, ZEA06, ZYL^{+}06, ZT07a, ZZ07, ZSC07, ZB07, ZZ08, ZZVM08, ZL08a, ZKL^{+}07, ZFM08, ZZ09, ZRS06, ZP06, ZSB^{+}08, ZSP08, ZHSS09, Zho07]. method [ZW06, ZZFW06, ZL08b, ZTPM05, ZQSD08, dVGLM09, dSMN^{+}04, dTDI^{+}07, vBK03, vDZ06, vLAvdV06, vdBG09, vdVX07]. method-based [DLD08]. methodology [BdCB09, FK09b, GZ08, GS09b, KDOO05, YC09a]. Methods [AL01, AGP01, Azm02, BKR^{+}01, BMRS01, BMRS02, BM01c, Boy02b, BS00e, BCM01, BSB01, CL01c, Cod01, CKS00, CMOV02, DCS00, DDH01, ELC02, ED07, FVOMY00, FF02, FPC^{+}00, GP00a, Gir00, GHW02, HH02a, HMS08b, HW02, HKV01, HLS01, Jan00, KLN^{+}01, KR02, KMA^{+}01, KKR01a, KKR01b, KM00, KMS02, KHV01, KMJ01, LOK01, LM01, LTZ01, LLdlP^{+}00, hLA01, LMSW02, Mac00, Mav02, Mit00, ML01b, NR01, NC01, OKL01, OF01, OS01, PD01, PRT00, PX02, PW00a, PW00b, PW01, PWS^{+}02, Rei00, RXH02, RM01b, Saf00, San01, SW00, Set01, SMP01, TNGH02, TWS02, WK01a, YC02, APTJ^{+}04, ABLS05, AS03b, Ain04, ABRR09a, ABRR09b, AT05a, BB04a, BSW03, BCL06, BY07, BBHM09, BZ08, BS08b, BHR06, BB07a, BC08, BS07, BT05, Bor03, BKLL04, BS06b, BLM03]. methods [BDS07, BRB03, BCGR05, BHR04, CT09, CLS^{+}06, CL08a, CSC^{+}08, CGMS03, CGMS06, CQO04, COV04, COQ06, CM06, CLS05, CL08c, Che07, CLL^{+}07b, CJ07, CR09, CLS04, CWD08, CFJ06, CC04, CD07, CP04c, CF04, CFJ09, DSM09b, De 04, DGH08, DL04, DD09, DLP08, DL03b, ERVE09, EGHE06, EHD08, Eg07, ETT05, ES06, EN06, ÉGP09, FSS03, FWP09, FD03, FR03, FPT05, Fou06, Fox09, Gab07, GT09a, GCGE03, GLMH09, GK03, GSV06, Gir06, GR08, GR04, GKL03, GD08, GLLX08, GF05b, HD07, HMS08a, HKO07, pHL09, Hel09b, Heu03, HHMK05, HNGB04, HL06b, HyLL07, HJL09, HL07b, HS04, HJM06, HJM07, HMR08, HRV08, HR07, IF09, JHSZ07, JSCZ08, JW09, KCH06, KCGH07, KOQ08, KTD03, KKL04, KK05c, KK05d, KPB08, KYK07, KvdVvdV06a, KWBH09, KK04, KAS08, KS08b, KKO04]. methods [Kri07, Kro05, Kro02, KWD07, KDW08, KH08, KP05, LY07b, LG08, Lau04, LM04, LSY04, LM03a, LBS^{+}04, LMS04, LRZ04, LH08a, LS05b, Low05, Low04, LZC04, LB03b, LMNK07, LTM09, LCdCN^{+}03, MJ09a, ML06a, MS08a, MEKS03, MNR07, MN06, MP05, ML08, MP07b, MJ06, MST06, MJ07, MSP^{+}06, MG07c, MY06c, MCP03, MHPR08, MLS^{+}05, MK03, MO06, NW07, NM06, NU09, Ni09, NWZL08, NLT07, NB04, NZ07, OS04, ODCK07, PR04a, PS07b, PS07d, QLK07, RCT07, RRC05, Ren07, RBvdV08, RGK07, RSO04, RS05, RS09a, RH05, San03, SPB09, ST06, SK08a, SM04, SB06c, SRNV07, SS05a, SFVK06, SAM05, SY03, Sme06, Str07a, SP06b, ST03b, TZHT04, TZ06, TWM07, TCN09, TD07, Tok06b, TT06a, Tow07, Tow09b, Tsy04, TPVG06, VSG05, VK04, Vos06, WT07a, WHLL03, WWC07]. methods [WLT08, WLC^{+}08, WM09, WZL09a, WZL09b, WG09, WG06, Wen07, Wen09, WH05, XXS07, XS06, XS09, XS05b, XLS09a, Yan08, Yan09, YYT05, YKG04, YE05, Yus06, ZKDT07, ZSW07, ZH04, ZKS^{+}09, ZW04, ZQ09, vEB05, vOP04]. Metric [Hua05, Aza06, HZ07a]. metrics [OB06]. Metropolis [QL01]. MFEM [WLE^{+}00]. MHD [HY11, ALGM01, AT05a, AT08, BTW04, BBG^{+}02, BvdHKG07, BSLN09, CKF02, CK03, CCF^{+}05, CH08, DKSW01, DKK^{+}02, Del01, DZ09b, FMR09, GS05b, GS08, GLL03, GFR09, GTMC08, GKV09, GLLN07, HMM08, HJ07, HY09, Jar04, LGKP07, LW01, LL08b, Mig07, NMM^{+}07, NMH^{+}07, ORM06, RWWS07, Sam09, SDGX07, TB04, TA06, YS07a, ZYL^{+}06, vdHK07]. MIB [YZW07, YW07, ZW06]. Micro [GS02, BBD04, BEA09, CRAG07, CHBS04, FT06, LCNR07, LR03, NFvS^{+}06, RB05, RE05, SFX03, SS05c, TS08, ZXQX08]. micro- [LCNR07]. micro-channel [SS05c]. micro-channel-flows [TS08]. micro-channels [SFX03]. micro-fluidics [RE05]. Micro-Inertia [GS02]. micro-local [BBD04]. micro-macro [BEA09, FT06]. micro-mixing [CRAG07]. micro-plate [CHBS04]. micro/nano [NFvS^{+}06]. micro/nano-channels [NFvS^{+}06]. microactuators [LTM09]. microbial [PC08]. microchannel [VLB09, ZTPM05]. microchannels [WWC07]. Microelectronic [AIRY01, AIR03]. Microfluidic [AA02]. microfluidics [GV08]. microlocal [BCR04, Dar02]. micromagnetic [VOD08, dSMF09]. Micromagnetics [WGCE01, GCGE03, MO06]. Micromixing [MJ09b]. microphysics [BDR^{+}04]. Microscale [Myo01, SB02, ZZ01, KPB08, Myo04, YE07]. microscope [TLAD04]. microscopic [AKP07]. Microstructural [ATV01, LLN00, CP04b, CP05]. Microstructure [EKK02, RV00, BEA09]. microstructures [WP09]. mid [CP03a, dSM05]. mid-point [CP03a, dSM05]. Mie [Shy01]. migration [CP04b, FEL^{+}05, HS07a]. Mikhlin [HW05]. MILC [BLS08]. Million [CWWZ00]. MIMD [DPRS01]. Mimetic [CS01a, LMS04, dVGLM09, GL08, LSS06, LMS08, SP06b]. Mine [GKL00]. minimal [Cec05, KLSW09]. minimisation [CORT09]. Minimization [HdGKG08, Yam01, COV04, JCT07, Lap04, RSSL09, RSS09, SNLS03, YMW06]. minimizing [BT03]. minimum [CFR04]. mirror [DDK06]. Miscible [IYI^{+}02, TM05]. Mises [GMO04]. Mittra [NCW^{+}09]. Mixed [AP02, BFG08, CGSS00, VCTS02, dA04, AMS03, BWLM09, BG05b, CHPR09, DDK06, Doh09, GL09b, HPS06a, HBLD07, Hel09b, LJW09, MP05, MESV09, NV09, ND04, RRW05, VBJ08a, VBJ08b, WG09]. Mixed-Basis [AP02]. mixed/discontinuous [MESV09]. mixing [CRAG07]. Mixture [IYI^{+}02, Shy01, VLKM02, BW07, CET09, Shy04, TLK07, ZKS^{+}09]. Mixtures [OB02, VG01, dSAK00, AS03a, AL08, SPB09]. MLFMA [DBF08]. mobile [RF06, RMF08]. mobile-bed [RF06]. Modal [LD06, MHdB07]. modal-based [MHdB07]. mode [CTS07, DH07, HW05, LY07b, Oh04, PGN08, SRNV07]. mode-separated [Oh04]. Model [ACK02, BISS01, Bon00, BMS00, Cle00, CR02, DDG02, DE02, DOWB01, EF02, FCB02, FGOV00, FV01, GP00b, GMS06, GC02b, GSW00, Hun01, JC02, hLA01, Mil06, Mil07, ML01b, MOS^{+}00, MR01, OF02, PS07c, PCCD00, Rom02, Sni01, Sun00, Tol02a, Tol02b, TTSG01, VCG03, vHBB02, AZB09, APP^{+}07, AK09, AW04, BS04a, BBDE05, BG07, BN04, BCCV09, BL08, BBvdV06, BCZ04, BNNP06, BTWGvBW07, BJP04, CLS^{+}06, CD04, CL07a, CL08b, CL03a, Che04, CL05, CL08d, CX08, CW08, CK08, CFGK05, CDV07, CDL04, CDL05, DDK06, DSM09b, DJM05, DGM07, DDH05, DSS07, DDGS09, DDFT09, EPW08, EKBL09, EF03, FVE04, FGS09, FNBB^{+}08, FCGK05, GGMN^{+}09, GS03b, GB03, GC06, Gra06a, GD07a, GD05, HBLD07, HW08, HHPW08, ICO04, IHL03, JA08, JN07, KD09]. model [KFIG06, KB04, Kou09, Kwo08, Lar03, LHR^{+}07, LDN04, LWDA09, LB03a, LGN05, LWF^{+}08, LD09b, LF04, LC03, MWM03, MM09, MGS07, MG05b, Myo04, PM08, PSC04, PS05, PHW08, PVPS09, PS03b, QA09, QFR04, RRV06, Rom07, RFVP09, Sar03, SMT^{+}08, SWK06, SW04a, SE04, SY09a, SL03, SK08b, SS05c, Sof09, SW08c, SK07b, SXyWX09, SS04, TLK07, TLAD04, TM05, TK04, VTC^{+}07, VP09a, VK09, VP09b, VQLZ04, Wea09, WDÖ^{+}03, XCY06, XDB09, XS07, XMT06, XHC08, YH07b, ZSWW03, ZK04, ZWS06, ZVQ07, ZXQX08, ZZ08, ZL09, ZDD09, ZSC06, KN09]. Model-based [Mil06, Mil07, GGMN^{+}09]. model-constrained [BTWGvBW07]. Modeled [GW01, HR08]. Modeling [Ano08-50, BV00, BTFY01, CFM09, CS05, CDDH07, CMP07, DDG02, DC02, FSM^{+}01, GZ07a, GS02, GM01a, GIA^{+}07, GVT01, Hum05, JL09, KM08b, MT07a, MK02b, OV00, OP02, OVG07, Ota00, POS00, QLS09, SJ02, SD00, SMO00, TZ07b, VPA02, Whi00, XK03, YLA08, ZVHP03, AHF04, AMH04, ASQR06, AJ09, BOK^{+}06, Cha09, FK06, FWR07, GZ09, GR08, GH08b, HSW07, HDR^{+}06, KDOO05, KT06, LMV04, LVL05, LSS^{+}09, Liu09a, Lyn08, MZ08, Mar09, ME09, MC03, MDM03, NCW^{+}09, OK06b, RMB07, RE05, RM07, RW03, SR09a, SZ08, Sau04, SKWN03, SM09b, SCC^{+}03a, SCC^{+}03b, SMGJ09, SP06a, SBC04, TZ07c, TJLT08, VSV03, VLB09, YE07, ZKL^{+}07, ZH09]. Modelled [CGSS00]. Modelling [CDS04, CP05, Cho05, GFS08, GYKL05, Hor02, JG09, KM02, LMS05, Pri08, VHI06, WHV^{+}00, ZGT06, de 00, BQQ09, CRAG07, CP04b, CBS05, Eli03, ES06, FHLO08, GLT07, KMSH08, KAS06, LKD04, Lap08, LDL^{+}09, LKX04, MT07b, QP03, RCB05, SS06a, SS06b, SJC07, Wan05, WGRA09]. Models [BSJ01, Bla00, ČPT01, GR01, HK00, Hig02, HKV01, KK00c, MEG02, Mie00, RV00, SSC00, ACGV07, AG09, BCB03, BKST09, BTC05, BC08, BBI09, BK07, CSC^{+}08, CRAG07, CKPW07, CRB^{+}08, CDV05, DVHM05, DMR09, DP07, DP08, Dic08, FCD^{+}06, GT09a, GZ07a, GZ08, GD06a, GM06, Hag07, HK08c, Hig05, KMID05, Kim05, LM08a, LD06, LCB09, LB03b, MM03, MJ09b, ML06b, PA07b, Rah04, RW08, SDCC05, SSE03, She08, SK04a, SS03b, SK06, TSG^{+}06, TW05, Thu08b, TR07, WAH09, WdND06, YHCD05, ZRS06, dFGLS05, dFJS09, dNWvSD07, dTWD09]. Moderate [VCP00, Vik03]. Modes [GBS00, RVVL09, PPCW06, TW05, TRSK09, TR07, WC08, dSMF09]. Modest [MCP03]. Modification [SWTM01, Vik03]. Modifications [RM01a]. Modified [BZ08, BADG00, CJ09, FH02, GST02, Jon05, LSL08, TTSG01, TSG02, WG06, BZ04, BL09c, CHL06b, Eg07, MU09, MC07c, Pal08, WRu03, ZB07]. Modified-truncation [Jon05]. Modular [Str00, SSB07]. Molecular [AC01, DPR00, DGA08, Yon01, AR08, ALT08, BPMR08, DTS05a, DTS05b, DST07a, GT09c, HS04, JG09, KNH05, LPK05, Li08b, MC07a, ML04, NTO^{+}07, PGB05, Pau07, PSZ09, Pro03, SE09, SHP07, TG04, VS09, YWC07]. molecule [LR07, NTO^{+}07, SMSS07, TLAD04]. molecules [HO06, LD06]. Moment [DK02b, MHS02, Abr06, Abr07, Abr09, AS09, CX08, DFV08, DS08, Fox08, Fox09, FDK06, GE07, IX07, IX09, LTZ03, RCT07, RW08, TS08, TFDK04, TRL01, VVS08]. moment-constrained [Abr06, Abr07, Abr09]. moment-of-fluid [AS09]. Moments [BW02, DC08, FLM08, LKD04, SH07a, Xia04, XAI06]. momentum [ABRR09a, ABRR09b, KH09, SAM05, SHP07]. Monge [DCF^{+}08]. Monitor [CHR01, HS03b]. Monitoring [ESD05]. monolithic [DHOT09, GA09]. monomolecular [RE07]. Monopoles [Del01]. Monotone [Cap09, LSSV07, YS08, AM03, AM04, CL08d, DPRN05, LSV09]. monotonic [KK05c, KK05d, Yeh07]. Monotonicity [BS00a, RM01a, TS02, BD08, DT04, HR07, MD06, NE05, RGK07, RH05, SH07c]. monotonicity- [RGK07]. Monotonicity-Preserving [RM01a, TS02, DT04]. Monte [ABRR09b, LM01, LM03a, MCP03, ABRR09a, AMH04, BBHM09, BS07, BMDS05, BSP06, BUEG06, BB09b, CLL07a, CGMS03, CGMS06, CTW^{+}08, CV06, CF06b, CS03, CS04, Dem04, DL03a, DL04, DUEB07, DDDC07, EULM03, ED07, FG04, FG05, FT09, Gen01, GL09a, GMH06, HH07c, HGM01, IH04, KB00, KMV03, KAS08, KLW09, LSL08, LM08b, LD09b, MMKP08, MU09, MBS03, NU09, OK07b, Pal08, Pet07, PK00, PVR07, PVPS09, QL01, RRV01, RS06b, SSE03, Sch08, SL04, She08, SA06, SMSS07, UH01, VK04, VK05b, Vol04a, WBM09, WGS^{+}08, WMH07, ZSB^{+}08]. Moore [FS06, VPA02]. morphological [GFG09]. morphology [Liu09a]. Morse [WWVG00]. MOSFETs [BMN07, BCCV09, JSCZ08]. MoT [Noe00]. Motion [BCMO01, CBMO02, Cor00, LK01, PG02a, RMO00, RM01b, vdVvdV02, BB04a, CJLS09, DHM03, Eld07, EES09, FPT05, FG06, Kro01, LMH07, MR07b, RA09, SP04, Xu08]. motions [Fra04, LDN04, PC06a]. movement [SC09a]. Moving [BMR01, BMRS02, BW01, FGG01, GPH^{+}01, Han00, HS06, Hua01a, LTZ01, LTZ02, MJ09a, MR00, MR02, RRL01, Str00, Str01a, TZHT04, TC01b, UMRK01, Wu01, ZRR00, AT05b, Aza07, Aza09, BS08b, BL09b, BCGR05, CSO09, CYS06, CHCOB09, DDM07, DT03, FS04, FYH^{+}06, FG07, GLL03, GS05c, GNNB08, Gre04, HT07, HMR08, HF08b, IG05, JD04, Jia07, JX07, JS05, KY08, KC06, KS09, LL03a, LC06a, LZ09b, LHZW05, LLOT06, LZH^{+}06, MWM03, Mad06, MKLU05, MLS^{+}05, NXS07, OTCM08, PN03, PH06, PL08, QS07, QLS09, RW03, SS08, SY09a, Spe05, TZL05, TLL^{+}08, TLK09, VB09, WT07a, XW06, YB06, YZLH09, ZKS^{+}09, vDZ06]. moving-boundary [LHZW05]. moving-least-squares [GNNB08, VB09]. MPDATA [SS05b]. MPI [OMK09]. MRA [BLG^{+}08]. MRT [PA07b]. MSPH [BZ04, BZ08, ZB07]. Multi [AS07, COR08, CD07, HL06b, HyLL07, JLT03, LNGK04, LL07, LJ09b, MN09a, NTYT01, NTYT02, PPC00, Rei00, SBGK00, SK07b, TOY09, TRL01, YKK08, AE03, AK09, ADR08, BSKH07, BJ09, BOT05, BLM03, BL03, BK07, CLG07, CLLG09, CET09, CR05, CKLS05, CLS05, CHB09, CWYM08, CX08, DR09a, DW09, Dic08, DS08, EHST03, EHS^{+}08, FWK08, FLM08, GAC^{+}09, HJL09, HG03, HA06, HA07, HA09, HAI09, IX09, JVVS07, JLT06, JL09, JLOT05a, JW09, KSO^{+}05, KK05c, KK05d, KLK08, Kou09, LMV04, LM04, LR07, LL03c, LW07, LBL04, MVD04, Mar06, Maz06, MG08, MK05, MDS03, MLS^{+}05, MK04b, MGNB09, NGvdWS09, Ols07, OK06b, OK07a, PSC^{+}06, PA07b, RSM05, RS06a, SGFL09, SD05a, TZ07a, TW03, TJLT08, UBRT07, VP09a, VHI06, WK05, WB09b, Xia04, XAI06, XHW07, ZWS07, ZSWW03, ZWS06]. multi [ZSC08, dSMN^{+}04]. Multi-Block [PPC00, CHB09, NGvdWS09, PSC^{+}06]. multi-class [ZSWW03, ZWS06]. multi-component [CKLS05, CLS05, JVVS07, Maz06, MLS^{+}05, MGNB09, TZ07a]. multi-corrector [LRS09]. Multi-dimensional [LJ09b, NTYT01, NTYT02, SBGK00, TOY09, XAI06, YKK08, BL03, JW09, KK05c, KK05d, KLK08, LL03c, MG08, Ols07, RS06a, XHW07, ZWS07]. Multi-domain [CD07, MVD04]. multi-element [FWK08, WK05]. multi-fluid [FLM08, GAC^{+}09, HG03, HAI09, LMV04, ZSC08, dSMN^{+}04]. multi-frequency [WB09b]. multi-grid [LR07]. multi-implicit [BLM03, LM04]. multi-integrated [Xia04, XAI06]. multi-level [EHST03, EHS^{+}08]. Multi-material [AS07, BSKH07, DS08, LBL04, SGFL09]. multi-mesh [DW09]. Multi-moment [TRL01, CX08, IX07, IX09]. multi-parameter-dependent [DR09a]. multi-particle [ADR08]. multi-phase [CET09, CR05, HA06, HA07, HA09, JLT06, KSO^{+}05, Mar06, MDS03, TJLT08]. multi-physical [MK04b]. Multi-physics [LL07]. multi-point [CWYM08]. multi-quadric [TW03]. multi-relaxation [PA07b, RSM05]. Multi-resolution [LNGK04, BOT05, CLG07, Kou09]. Multi-scale [JLT03, MN09a, SK07b, AE03, BJ09, JLT06, JL09, OK06b, OK07a, UBRT07, VP09a]. multi-scattered [Dic08]. multi-species [AK09, BK07, SD05a]. multi-state [MK05]. multi-static [CLLG09]. Multi-Symplectic [Rei00, COR08, HL06b, HyLL07, HJL09]. multi-valued [JLOT05a, LW07]. multi-viscosity [VHI06]. multiband [RW08]. multiblock [RJ06]. Multicenter [GM01c, SB06a]. Multicloud [KJ09a]. Multicomponent [HLS01, LBD02, OB02, Shy01, WDM01, BGM08, BS09b, JC06b, Lau06, MC04, MM03, Shy06, WAO^{+}04]. multiconfiguration [CGL06]. Multidimensional [CRD02, GF02, Lap02, Noe00, NC01, SHA08, TX00, WB01, ZR08, Abr06, Abr07, Abr09, Asl04a, BFT09, BGN03, KT04, LD09a, LF06, LHGF04, SL06, TXCD07, XMT05]. Multidomain [GBGM01, LP07a, DDH05, DLD08, DGJ03]. multifluid [MCN03, NDT06]. Multifluids [AK01, HK04b]. multifractal [TPVG06]. Multifragmentation [PA00]. multifrequency [MYW07]. multifrequency-grey [MYW07]. Multigrid [Alb00, BZB00, BL05, DIV00, EAY01, FOLD05, GMB01, KKS05, KJ09a, KvRvdVvdV07, KM00, Mav02, MLS01, Pai01, SMB09, Soc03, VC00, WK07, Yua02, Zha02, ABHT03, AHPT07, CLS^{+}06, CS08a, DHOT09, EKP07, GT05, HH07a, Heu03, HMMO05, HWWL09, JHSZ07, KW06, KKL04, KL04, Küm04b, LLY05, LDPL08, LBL06b, LZH^{+}06, LZH^{+}07, NM06, NvL03, RKE^{+}07, SRNV07, TZ03, VBL07, WZ09, WKL07, XYK05, vdV08, SD05b]. multigrid-based [RKE^{+}07]. Multigroup [TFDK04, GS06a, SO08]. multigroup-WENO [GS06a]. Multilayered [CY00, Gut00]. multilayers [GCW07]. Multilevel [BCHL07, DJTT05, TSB01, TC09b, GKD09, Hab04, LSS^{+}09, LMS08, RAB07, TSB03, VK05a]. Multimaterial [LX00, UTBV03]. multioperators [Tol07, Tol08]. multioperators-based [Tol07, Tol08]. Multiphase [GM04, Gos04, GM06, Han01, HLS01, SJ02, Sni01, TBE^{+}01, VLKM02, VC00, YXU01, AS03a, BCB03, CL07a, CL08b, DP09, GCNB07, HJ09, ICO04, KT07, LZT09, LK09, LKO05, LJ06, LTL^{+}09, MC04, MK06, NL08, PFSL07, PA07b, QLS09, SPB09, Shi07, TBJ^{+}09, WP09, ZZVM08, ZLAC05, ZSC06]. multiphysics [FMD^{+}09]. Multiple [GB08b, LR01a, LTZ01, Mu02, RW00, TNR02, XHC08, ACR08, BS09a, BBMB07, BOK^{+}06, CJLS09, CGDT09, CK07, CB09, Del03b, ELVE07, GK04, GK07, HS09a, JG09, KCMM03, Kro02, MK07, Ngu08, RW03, SHPC09, Spe05, TZ07b, YXLF05]. multiple-grid [CK07]. Multiple-Heaps [Mu02]. multiple-scale [BBMB07]. multiplication [NU09]. Multiplicative [Ram06]. multiplier [BG05a, DLMK04, SPT05, VMN07, ZSP08]. multiplier/fictitious [SPT05]. Multipliers [TB00a]. Multiply [BMQS02, HJ02, Mil05, VRM07]. Multipoint [QL01]. Multipole [Dar00b, CDJ07, CCG^{+}06, CWD08, CFR08, Dar02, DH04, DC07, FD09a, GKD09, GH02, GD06b, GD08, KP05, LCM07, ST06, SK04b, TC09b, TG08, VOD08, YBZ04, Yin06, ZT07a, ZKL^{+}07, ZD05]. multipole-accelerated [ZD05]. multipole-to-local [CFR08]. multipoles [OLLL03]. Multiresolution [CDKP00, BK07, DGRS08, RSTB03]. Multiscale [AA02, BP06, CGP05, CF06b, FSM^{+}01, FY07, HJ09, HO06, JP00, LLIK01a, LLIK01b, LJ06, PK00, RV00, TZ07c, Vay01, AZ06, BCM^{+}07, BLM04, CELS07, ERVE09, EGHE06, FVE04, FMD^{+}09, GZ07a, GZ09, Gra06b, HBHJ08, KP07, KKS07, LZT09, LMS08, MM09, MY06c, MHdB07, NZ05, Ngu08, PBH04, RE05, SDD07, SY03, THL06, Tan08, WGRA09, YS07a, YE05, YE07, YH07b, ZSB^{+}08, ZHSS09]. multiscale/multiphysics [FMD^{+}09]. multislab [dA04]. Multispecies [BJ02, Del03a, SD05b]. multistack [DS05b]. Multistage [Löh04]. Multistate [BU02]. multistep [HR07, RH05]. multisymplectic [IS04, SW08b]. multitime [vdV08]. Multivalued [Gos02, JLOT05b, QL04]. Multivariate [WGNT06, AGSX09]. Multiwavelet [ABGV02]. Multiwavelets [JMK01, TNR02]. Mumford [ET06, RR07]. MUSCL [Ber06b, BL01, Waa09]. MUSIC [PL09a]. MUSIC-type [PL09a]. MUSTA [TT06b].

N [Aza09]. Nagumo [EV03]. Nanbu [DWC^{+}09, WLC^{+}08]. Nano [GK02, BCCV09, JSCZ08, LCNR07, VTM^{+}08, ZK04]. nano-channels [NFvS^{+}06]. nano-flows [LCNR07]. nano-MOSFETs [BCCV09, JSCZ08]. Nano-particle [GK02, ZK04]. nano-systems [VTM^{+}08]. nanodevices [CLL^{+}07b]. nanoparticle [MLFG06]. nanoparticles [FY07, MWG^{+}06]. Nanoscale [BGR08, BMN07, CL05, FH07]. nanostructures [PA05, RRC05]. National [Har04]. Natural [MPP01, SVH^{+}06, AZ05, DR09a, GZ07b, GCCD07, MZ08, PS03a, PSC04]. nature [Fen06]. Navier [DD03a, AQ09, BQQ09, BCDR06, BHR06, BB07a, BACFT05, BLM08, BCVK02, Boe05, BT06, BJ09, BCM01, BGLN05, CSL08, DC01, DR09a, DD09, DB04, Dom08, DD03b, EHST03, EHS^{+}08, FL03, FOLD05, FD07, GS07, Gel06, GSV06, GCNB07, GR08, GS03c, Gri09, GSW00, GK05, HH08, HH01, HDC02, HK08a, HH02b, HLMM07, HS08b, HLL08, HC05, ILL09, JK00, JL04b, JMC03, KA05, KE06, KDK^{+}07, KG08, KAK03, KvdVvdV06a, KvdVvdV06b, KvRvdVvdV07, KM00, KB01, KS09, KT03, LMN^{+}09, LOK01, Lee09, LC01, LL01b, LFX05, LDPL08, LRZ04, LP07b, Liu09b, LMS02, LB04, MPP01, MVD04, MRS09, MCG08, MSS08, MF00, MG06, MLS01, NW07, NZ05, Ni09, Nik06, NMS07, NGvdWS09, Pai01, PNMK09, PKP01, Pet01, PR03, PR04b, Pon07a, Pon07b, PR06, RBH03, RS06a, SML02, SNGAS04]. Navier [SFE07, SMB09, STZ07, Soc03, SCN07, SN08, STR07b, SPW^{+}00, TOZP03, TXCD07, TWS02, VSW04, WRu03, WPH00, WK01b, WS01, XK01, Xu01c, XYK05, YS07a, ZL08b, ZDNP00, vBRK01]. Near [FR02, KMID05, MK02b, OK07b, SKWN03, GLLX08, HAP05, Khe04]. Near-field [OK07b]. Near-wall [KMID05, SKWN03]. Nearest [Par02]. Nearest-Grid-Point [Par02]. Nearly [AJG01, BKLL04, DST07a, DS09b, HL07b, ZJW06]. necessary [CLMRP08, LM08a]. Negative [SHS02, MHE06, NV09]. negativity [SCT09]. NEGF [JSCZ08]. Neighbor [DTS05a, DTS05b]. neighboring [XLS09a]. Neoclassical [HKKS^{+}01]. Nested [DAJ07, ELVE07, FCB02, LJSM08, MM09, RAB07]. Network [MK02b, RCT07, XS07]. networks [KF06, LMK03]. Neumann [APQ03, GP04, GK04, Gui03, HW05, Hel09a, NR01, Poz01b, SDS07, TB00a, YLA08]. Neural [MK02b]. neuronal [RCT07]. neutral [BBK06, CDV05, GWF^{+}07, GBS00, LCB04, TPR05]. Neutrally [PG02a]. neutron [BH05, FHJK09, Mac07, NU09, RW08]. New-version-fast-multipole-method [LCM07]. Newmark [CL07b]. Newton [Yan09, BB07a, Boy02b, CBKM00b, Cha07b, CZ09, HC05, KM00, KK04, KT07, MKR00, NOG08a, NOG08b, ORM06, SNGAS04, TWS02, YLD09]. Newton-conjugate-gradient [Yan09]. Newtonian [FS04, VBL04]. NICAM [SMT^{+}08]. Nicolson [Han00, KW08a]. Niño [CC08a]. nitride [BMK^{+}06]. NMR [KHV01]. No [SN08, HSC09, PK05]. No-slip [SN08, HSC09, PK05]. Nodal [GHW02, HW02, KWBH09, WPH00, GLMH09, GW05, JH06, Pon07b, WRu03]. node [KLP^{+}09, MCGV04, WF06]. node-centered [MCGV04]. noise [BB04b, CBJdlC07, MSB07a, SMS08]. Non [BTW04, BS00a, BM01c, Dem04, GBS00, Hub07, KM07b, MKR00, NV09, SCT09, Tor03, Vas00, Yon01, Abg06, AB03, BFB08, BDRT09, Ber04, BdCB09, BCCD08, BSW05, BCI^{+}08, CFS09, Cap08a, Cap08c, Cap09, CCV03, CN05, CEH09, CSL08, CS06, CS07d, CP08, DHOT09, DSM09b, DP07, DK07, DKTT07, FS04, FR03, GS06a, GT09a, GZ08, GGS09, GHB03, GN03, GT05, GWF^{+}07, GYKL05, Hau08a, Hau08b, HMA05, HKG08, HAP05, HS06, HJJ09, Hu05, Hwa03, ISNY05, JJGL06, JJGL07, JA08, JSCZ08, JG09, KD09, KB04, KK03b, Kou09, KLSW09, LSA06, LVL05, LMS04, LCCG05, MGS09, MJ06, MGC06, MG06, MG07d, MK03, NN04, OMK09, Pav07, PWM06, PK07, SBA07, SS03a, SAK05, SSB07, SLV09, SB06c, SE04, SZ05, STZ07, TWM07, TT05a, Tok06a]. non [TT05b, TB04, TPR05, VMN07, VBL04, VZSL07, WT07b, WSYS09, WC07, XHC08, YKG04, YA05, YS06, YH07b, ZSWW03, ZIP06, ZWS06, ZT03]. non-aligned [GYKL05]. Non-autonomous [BM01c]. non-conformal [LMS04, VZSL07]. non-conforming [CCV03, CEH09, SB06c]. non-conservative [DP07, KD09]. non-convex [HJJ09]. non-diagonal [WC07]. non-dispersive [MGS09]. Non-equilibrium [MKR00, BSW05, GT09a, GT05, JSCZ08, JG09, MK03, SSB07, WSYS09, XHC08]. non-Gaussian [HMA05]. non-graded [MGC06, MG06, MG07d]. Non-homogeneous [KM07b, FR03, SBA07]. non-hybrid [BFB08]. non-hydrostatic [SE04]. non-hyperbolic [Hwa03]. non-inertial [PK07]. non-isothermal [DHOT09]. Non-iterative [Yon01, OMK09]. non-Lagrange [VMN07]. non-linear [BDRT09, BdCB09, BCI^{+}08, CFS09, Cap08c, Cap09, CN05, Dem04, GZ08, Hau08a, Hau08b, MJ06, TT05a, TT05b, WT07b, YKG04, YH07b]. non-linearly [Tok06a]. non-local [KB04, LSA06, Pav07]. non-matching [JJGL06, JJGL07, LVL05]. Non-Monte [Dem04]. Non-negative [NV09]. Non-negativity [SCT09]. Non-neutral [GBS00, GWF^{+}07, TPR05]. non-Newtonian [FS04, VBL04]. non-orthogonal [LMS04]. non-orthogonality [SS03a]. Non-oscillatory [BTW04, BS00a, Hub07, Abg06, BCCD08, CP08, DK07, DKTT07, HAP05, TWM07, ZSWW03, ZWS06]. non-overlapping [LVL05]. non-parametric [Kou09]. non-periodic [GHB03, LCCG05, SAK05, SLV09]. non-polynomial [YS06]. non-radially [KLSW09]. non-reactive [HS06]. non-reflecting [AB03, GN03, NN04, PWM06]. non-smooth [Ber04, CS06, CS07d]. non-staggered [CSL08]. non-stationary [GS06a]. non-thermal [DSM09b]. Non-uniform [Tor03, Vas00, Cap08a, HKG08, Hu05, ISNY05, KK03b, SZ05, STZ07, TB04, ZIP06, ZT03]. non-unit [JA08]. non-viscous [GGS09]. nonaffine [Ngu07]. nonaffine-parametrized [Ngu07]. Noncompact [GBGM01]. nonconforming [Fou06]. Nonconservative [Wu01, CR09, RBvdV08]. nonconvex [Ser09]. Nondissipative [KPP09]. Nonequilibrium [VDM^{+}02, KLM07, LRS07, Ols07]. Nonhomogeneous [FP02, GC01, FG04, FG05]. Nonhydrostatic [Bon00, SMT^{+}08, GR08, MM09, SK08b, SW08c]. Nonlinear [AL01, BR01, BC01, Boy02b, CKF02, CSP01, CRD02, DZ00, FGG01, FT01, Gla01, GPL05, GLLN09, HZ02, IKS01, KLN^{+}01, KK00b, KJ01, Kul01, KT00a, LMSW02, MF01, Mav02, Nie01, PR01a, Saf02, SKAS01, SGG^{+}04, Sti02, TS01, WK01b, YL01, de 00, ARRS09, AMH04, AKV06, AMXL09, AG08, BHS03, BFT07, BFT09, BN04, BB09a, BS06b, BG05b, CHL06a, CL08a, CHL09, CBJdlC07, CK03, CL06b, CC03, CCJ07, CLS09a, Chu09, CFJ09, DH07, Doh09, DKTT07, EKBL09, Fan08, FT05, FL07, FWR07, FG07, HZ08, HK08c, HL06b, HyLL07, HLL08, HWWL09, HC05, JTL09, KSH^{+}06, KLSW09, KT07, LZT09, LSY04, Low04, LL08b, Ma05, MY06a, MKOW04, MY09, MESV09, MT07b, Mou04, Nas08, NPC09b, NF09, NL09, OCFF08, PSD09, RSSL09, RGK07, Sac07, Sau04, SS07c, SP05c]. nonlinear [TTZ03, TWM07, WFTS05, Wan05, WKL07, XS05b, XHW07, XG09, YM07, YLD09, ZJS08, Zhe06, Zhe07, vdVX07]. nonlinear-multigrid [HWWL09]. nonlinearity [LY07a]. Nonlinearly [LAS01]. Nonlocal [BZB00, FS00a, FS00b]. Nonmonotone [SL07c]. Nonorthogonal [LP02, FT05]. Nonoscillatory [JMP02, TH01, WC01, WH02]. Nonparabolic [Rom02]. nonparabolicity [WHLL03]. Nonparametric [Mac00]. nonreacting [DBS06]. Nonreflecting [AGH02, Ata04, Giv01, Gro00, GK07, RC00, AG08, Zhe06]. Nonseparable [TNR02]. nonsmooth [FCJ08a, FCJ08b]. nonspherical [DTS05a, DTS05b]. Nonstationary [IKL^{+}08]. nonstiff [CR07]. Nonsymmetric [DF00b, JHZ^{+}09]. Nonuniform [GZ01, HLS02b, HA02, LLQ^{+}02, MN02, LG05, Rem06, SS09a, VB08, WA08]. norm [SVH^{+}06]. normal [ND04, RMB07, TW05, TR07]. normalization [Tow09a]. normalize [Hag07]. normally [NTB07]. normals [FB08, RMB07]. Note [Ano03y, Ano03z, Ano08-51, Del01, DF00b, Poz01b, Wu02, Ano07-32, GXW07, GJK09, GS03c, TL06, UYK^{+}04]. Notes [Ber06a]. Novel [BU02, DC01, DSS00, EG08, FH00b, YCL05, BAMD07, FGP08, LSA06, LL04b, LNXNTX09, MvW08, Pap08, SLC07, dSMF09]. nozzle [CGH05]. NS [WLC^{+}06]. Nuclear [Saf00, BDR^{+}04, KP07, PGB05]. null [CEL06]. Number [AKY01, Cor00, DKX01, FH00b, FG02, HT00b, LLIK01a, MP02, MHS02, MPC02, NTYT02, Nic00, PW01, SBGK00, SSD00, Tol02a, WPM02a, ZRR00, Ano04z, BDHN09, BDR^{+}04, BTW03, CLB08, Del07, DBBP08, DDH05, DST07b, Heu03, JS05, KKM08, LG03a, sLwG08, LM08c, MT03, MDR07, NMM^{+}07, NMH^{+}07, OTCM08, PPDM08, RE07, RB09b, SM06a, SFMP06, TSG^{+}06, TMD^{+}08, VK05b, XH03, XP04b, XLP05]. Numbers [AC01, BEPT09, CTS07, DKS^{+}03, HY09, HY11, KPB08, Lee07b, OVG07, SDGX07, Vik03]. numeric [HBHS09]. Numerical [ART02, ART04, ACS00, ACLS03, APQ02, BS04a, BSW03, BJM03, BLW04, BL09a, BST01, BMRS01, BCG09, BCR04, BA03, BS01, BRL02, BPL06, BS06b, Boy02a, BC02b, Bur05, CPR05, CFA01, CHH06, CD04, CP03a, CBJdlC07, CGRGV^{+}04, CQO04, COV04, CWL08, Cle00, CL03b, CB09, CF04, CkM07, DW00, Dar00b, Del07, Den07, DJ04, EPW08, EE08, Eld07, Eli03, FLG01, FT05, FSB01, FP02, FLM08, FCB02, GMD03, GGS09, GS02, GK01, GLS03, GPH^{+}01, GP00b, GC02b, HLFB07, HMM02, HK02, HF01, HPZ01, HL07c, HSL08, HLWW04, HLWW06, HWW07, IG05, IM05, JLCD01, JWSC00, JW02, JMZ04, JHB^{+}09, JS05, KP07, Kan02, KSH^{+}06, KAIN01, KSW07, KLM07, KMS04, KJ01, Kro02, LCB04, LP00, Lem00, Lin02, LGK06, LOK05, LP04b, LP02, LCdCN^{+}03, MR00, MCCT02, MRRS05, MC09, Men04]. Numerical [Mie00, MY06c, MC00b, MLS^{+}05, NS04, NLT08, Nys02, OKL01, OL01, PSCB08, PD01, PSCQ03, PR01b, PIN09, PWS^{+}02, PCCD00, PO01, Pud06, Ram03, RCT07, RRC05, RGS04, RRL01, RXH02, RFVP09, Saf02, Sai02, SZ01, SLY02, San01, SJ04, SK08a, SB09, SL04, SFVK06, SSND03, SSC00, SKW05, Sus06, TS01, Thu08a, TRSK09, Tok06a, TC01a, TCM^{+}00, TdAAP08, TE04, TV08, TPVG06, VC03, VR02, VPA02, VQSZ02, WHLL03, Wee02, Whi00, WO05, WO09, WB01, XMT06, XG09, YM07, YFLS06, YVD00, YE05, dWKL07, vBRK01, vZS07, vdBG09, AS03a, APP^{+}07, AK07, Ano04z, ACR08, AMP09, AM05, BL04, BW06, BCL06, BY07, Bar04, BFT07, BFT09, BV05, BDGL05, BLM08, BWLM09, BCM09, BBCT09, BW07, BP07, CLB08, CRAG07, Cec05, CMG09, CMP07, CHG^{+}07, CP05, CM03]. numerical [CHCOB09, CC04, CP04c, DMHP07, DSJ03, DLD08, DWC^{+}09, DF04, DHM07, DLW04, Eld08a, FDD07, FR03, FPT05, FF03, FD09b, GCGE03, GS07, GR04, GCLB04, GKL03, HK06, HKM08, HP04a, HKG08, HS03a, Hoh06, HM04, HL05, HM05, HLRZ06, HPS^{+}06b, HMR08, IKS^{+}09, JD04, JW03, JS07, KK03a, KW08a, KK05c, KK05d, KH07, KDF07, KSJ03, Kro01, Kry04, KN04, Lar07, Lau04, LMV04, LDN04, LSA06, LKE04, LG07, LZ09c, Liu08, LL08a, Low05, LQ06, MM07, MCM04, MLSD07, MTWW06, MC06a, ML08, MT04, MWG^{+}06, MSB07b, MGNB09, NG06a, NTB07, OPML07, OCFF08, PDHP07, PC08, PM07, PSMW09, Pro05, QKS06, Ram06, Ren07, RSW06, RMG^{+}09, Rom07, Ros09, RS09b, RV07, SB06a, SM09a, SDGX07, SROCFF03, SKS08, SFDL07, SW08a, SSB07, SMS08, SW04a, SD05b, SP04, SP05a]. numerical [Shi07, SS06b, Sme06, SP06a, SFMP06, TMS06, TM07, TWM07, TW07, TG06, Tsy03, UL06, VVM05, VSV03, VGBZ09, VGL^{+}07, VBL03, VBL04, WT07a, WB09a, Wen07, Wen09, WMH07, YXLF05, ZSW03, ZSWW03, ZWS06, ZZ08, ZT03, ZD08, dSMF09, VBJ08a]. Numerically [WDM01, RV09, Wag05]. numerically-induced [RV09]. Nunziato [AW04]. NWP [Thu08b]. Nyquist [KP08]. Nyström [FWW04, HyLL07, TC07b].

OBC [Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano08-34, Ano08-35, Ano08-36, Ano08-37, Ano08-38, Ano08-39, Ano08-40, Ano08-41, Ano08-42, Ano08-43, Ano08-44, Ano08-45, Ano08-46, Ano08-47, Ano08-48, Ano08-49, Ano09-49, Ano09-50, Ano09-51, Ano09-52, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09-57, Ano09-58, Ano09-59, Ano09-60, Ano09-61, Ano09-62, Ano09-63, Ano09-64, Ano09-65, Ano09-66, Ano09-67, Ano09-68, Ano09-69, Ano09-70, Ano09-71, Ano09-72]. Object [DPRS01, QRHD00, RFFP06, TZL05, ZSC08]. Object-Oriented [QRHD00, RFFP06, ZSC08]. objective [MC03]. objects [AvdB04, ADR08, IQT08, LZH^{+}06, RW03, TC09b, Xu08]. observables [JLOT05b, JLOT05a]. Observations [KS02a, CHM08, HM08, KYK07]. observer [BCI^{+}08, VS07]. obstacle [FNS07, GG04, Lee07a]. Obstacles [Pai01, BG09, KS09, TC07b]. obtaining [GWF^{+}07]. Ocean [Ano08-50, Bla00, DOWB01, Hig02, KN09, MR01, SM09b, HHPW08, Hig05, Jan08, Ler06, PHW08, SP06a, Wea09, WDÖ^{+}03, WAH09, WdND06, dNWvSD07, dTWD09]. ocean-climate [dNWvSD07, dTWD09]. Octant [MY07]. octree [HH07a]. odd [RVM07]. odd-even [RVM07]. ODE [HR01].

ODEs [CPKW09, Tok06b]. off [GGS09, MT07a, SZB^{+}07, Vil08, YFLS06]. offsetting [Jia07]. oil [LCH03]. One [BMRS01, CWT00, DMG00, Del01, FSY00, GKE04, LT09b, MR00, MR02, PL01, QLK07, RB06, SFY01, VD00, VC00, VS02, ZQ09, AB03, ABK09, AI09, BTW04, BFT07, BS04b, Bil05, BDCG03, Boy03, CGSR08, CC07, Cho05, CR09, DT04, DBTM08, FS09, GM04, Gos04, GM06, HH07b, HZ08, HAP06, HGB^{+}03, LKX04, NFA03, QS04, SR09a, SKWN03, SK04b, WO05, Xia04, Zhe06, HA02]. One- [LT09b, BTW04]. One-Cell [VC00]. One-Dimensional [BMRS01, CWT00, DMG00, Del01, MR00, MR02, PL01, VS02, GKE04, QLK07, ZQ09, AB03, ABK09, BDCG03, CC07, CR09, GM04, Gos04, GM06, HZ08, HAP06, HGB^{+}03, NFA03, QS04, SKWN03, SK04b, Xia04, Zhe06]. one-fluid [LKX04]. one-parameter [CGSR08]. One-sided [RB06, HH07b, SR09a]. one-step [DT04, DBTM08]. One-Way [FSY00, SFY01]. onset [CGM07]. onto [NTB07]. Open [Liu09b, SS07c, BP06, BTC05, BF07, CZVS04, JR03, JR04, LZ09a]. open-channel [CZVS04]. Operator [KLN^{+}01, KK00b, PRT00, Spo00, TK04, BG05a, CWJ07, CFR08, DD05, DWC^{+}09, FL09, IAT08, KJ09a, KK07, Lab09, PAD07, PC06a, RS05, RS09a, RBK09, SRM09, TBT^{+}09]. operator-stable [PC06a]. Operators [Edw00, Her00, Lem00, MHS01, NR01, BO04, DK06, Gra06b, Her09, Heu03, LM08b, MN04, MN17, RS05, RS09a]. Optical [BV00, GHV00, Lin01, Whi00, BBD04, FH03, KM03, KNH05, MWM08]. Optics [Gos02, OCK^{+}02, CQRW05, JW06]. Optimal [BHS03, BC08, CHG^{+}07, CKL00, FH02, HZ07b, HN03, IF09, KFIG06, KMA^{+}01, MO06, NvL03, WR09, BHNPR07, CBGI09, DCF^{+}08, FLB03, HH08, HAP05, Liu08, MG05a, NL08, Pee03, PSM08, PRL03, SY09b, TW05, TR07]. optimality [PVR07]. optimisation [Pro03]. Optimised [Kim07]. Optimization [CD00, HGM^{+}00, IFZ01, JK02, KMS02, MK02a, OS01, RBSL06, SSSWD00, SKK^{+}08, TWM07, TSG02, TS07, AS05b, AJT04, AA06, BP09, BP04a, BTWGvBW07, CKvT07, DCF^{+}08, DAJ07, FLB03, GKJW07, GJK09, Ham07, HSBG05, HE07, HS09b, LTWW07, LLC^{+}08, LTM09, MS08a, MC03, MD06, Pir07, PL08, Pro08, SHA08, TW03, WLKW07, YMW06, ZL08b, MS08a]. Optimization-Based [KMS02]. Optimized [AZ03, BS09a, BM05, CL01b, HB05a, SHWW00, WC01, ZT07b, MTWW06, PSG05]. optimizing [Hab04]. optimum [LT09a]. Orbit [YMF01, PH09]. orbital [Küm04b, LZ07, SLG^{+}03]. Orbitals [GM01c]. orbits [KG03, SNGAS04]. Order [AC00, ACY00, BS00a, BISS01, BRL02, BK01, BSB01, CL01a, CKGL02, Coe02, CR00, CSP01, DV02, DZ00, DLS^{+}00, FT01, FT09, GC01, GFCK02, GHW02, Giv01, HLS02b, HW02, KC00, KB01, LP00, Lai02, LS00, MC01, NR01, Nic00, NC01, Nys02, OGV02, PKP01, PM00, QS02, SLY02, TK00, TX00, TS01, TS02, VCP00, Vas00, VG02, WH02, XCZ02, XK01, YP01, YL01, Zha02, ZS01, de 00, AM03, AM04, AV05, APTJ^{+}04, APP^{+}07, Ain04, AMR06, AV03, AG09, AC09, AB05b, AT05b, BS04a, BC05, BKST09, BGN07, BFT07, BFT09, Bea08, BC08, BBMB07, BRC^{+}09, BdCB09, BLM03, Boy06, BSW05, BGLN05, BH04, BHP07, BL03, CT09, CVB06, CP03a, Cap08c, Cap09, CT08a, CC03, CP04a, CC07, CS07a, CKG04, CS06, CS07d, CRB^{+}08, CR09, CFP06]. order [CBS05, CF04, CFJ09, DT04, DSM09b, DPRN05, DC07, DBBP08, DTMS06, DR09b, DDFT09, DK07, DET08, DZ09b, EKBL09, FNS07, FRS08, FB08, FOLD05, FD07, FR03, FK09b, Fox08, Fox09, GPC07, GG04, GH08a, GSV09, GGF03, GF05a, Gir06, GN03, Gom08, GR04, GBS06, GP05, GHMP07, GLT07, GL08, HH07b, HMOG08, HWL08, HH08, Hau08a, Hau08b, HMPR07, HAP05, HAP06, HJJ09, HMMR04, HMMO05, HJM07, Hub07, HB05b, IX09, IM07, IQ08, JH06, JD09, JBF07, JMC03, KSO^{+}05, KKM08, KE06, KCGH07, KZ04, KT06, KK05a, KK05b, KPB08, KLK08, KYK07, Kok09, KT05, KPP07, KPP09, KB06, Kri07, KWD07, KQW03b, Lab09, LL09, LSD07, LM08a, LTZ03, LG04, Li08a, LSZZ08, LN09, LF05, LRZ04, LCS09, LS09, LF04, LB03b, LCdCN^{+}03, MZ08, Mai09b, Mai09a]. order [MS03, MRS09, MvW08, MN06, MY06b, MG07a, MP05, MOG09, MG06, MG07d, MG08, MC07c, MHPR08, MPFC08, NLF03, NPH09, Nas08, NM06, NOG08a, NOG08b, NPC09a, NPC09b, Nis07, NPPN06, NXS07, NF09, NGvdWS09, ODAF07, Ols07, Ols09, PKD07, PP04, PRL03, QW05, QA09, RP08a, RRW05, RDPN07, RH05, SZ08, SGFL09, SDM04, SM04, SPM03, SFE07, SMB09, SZC09, SS05a, SZS03, SZ05, STZ07, SY03, SP05b, SCW^{+}09, SGG^{+}04, SL06, SP06b, SRX07, Sus03, SN06, SCN07, SN08, TMS06, TM07, TLAD04, TWYC06, TFD06, TCN09, TD07, Tol07, Tol08, TB04, VGCN05, VWW04, Wan04a, Wan04b, WL06, WM07, WSYS09, WZ09, WGRA09, WD07, WZ07, Wen07, Wen09, WA08, WAH09, WMH07, XS05a, XS06, XCRX08, XS05c, YMT^{+}04, YC09a, YC09b, YP06, YHSX07, YS07a, Yeh07, YBZ06, You06]. order [YZW05, ZKDT07, ZL04, ZP05, ZJS08, ZJ09, ZW04, ZT03, Zho07, ZZFW06, ZYHS07, vZdBB07]. Order- [FT09]. ordering [NL08, SNLS03]. Orders [GST02, MBM01]. Ordinary [MOvL00]. Ordinates [Coe02]. Organic [deM02, Lap03]. organism [LB03a]. organization [FY07]. orientations [LR07]. Oriented [OV00, QRHD00, BTWGvBW07, RBH03, RFFP06, ZSC08]. origin [CL07b]. origins [Lyn08]. Ornstein [KP04]. Orr [GFR09, Meh04]. Orthogonal [AJG01, Tót02, XCZ02, Bia03, CRB^{+}08, LMS04, Nik06, ZJW06]. orthogonality [SS03a]. Orthotropic [LLN00]. Oscillating [OKL01, DGH08]. Oscillation [OF02, CJSS08, KLK08]. Oscillation-Free [OF02]. Oscillationless [XY01]. oscillations [BPM06, CBC09, Gos04]. oscillator [LTD^{+}06]. oscillatory [Abg06, BTW04, BS00a, BCCD08, CP08, DK07, DKTT07, HAP05, Hub07, TWM07, ZSWW03, ZWS06, HGM^{+}00]. Osculatory [BR01]. Osher [FSS03, LBV00]. osmotic [WWC07]. Ostwald [Hor06]. Other [Boy02b, MR03]. Out-of-Core [TR02a]. Outflow [Eli02, Eli07, FGP08, HAS05, HEN09, MJ07, SAK05]. Output [MPP01]. Outputs [VD00, VD02, VD03]. Outstanding [SS01b]. Overdetermined [Boy02b, Str07a]. overfilled [Woo06]. overlap [SB06a]. overlap-like [SB06a]. Overlapping [PW00a, WZ00, BSKH07, BHS09, DTMS06, HS03a, HS06, HS08a, KZ06, KP05, LVL05, Li08a, Liu05, TZL05]. Overlying [Str01b]. overmoded [LKD04]. Overset [BE02, FS04, SS05a, TJS03]. overset-grid [TJS03]. overtopping [LTD04]. Overview [OF01]. Oxidation [GR01]. oxide [GIA^{+}07, GIA^{+}08].

P [Gon07]. package [DFG^{+}06]. packed [CLL07a]. packet [BS04d, BS06a]. packings [DTSC04, DST07a]. Padé [WH02, CDI09, Lur07, SFY01, WH02, You06]. Padé-Gegenbauer [Lur07]. Padé-Type [WH02]. pages [DKX01, HT00b, LLIK01a, MP02, MPC02, NTYT02, PW01, SBGK00, Tol02a, ZRR00]. pair [CHPR09]. paper [Aza09]. Papers [Ano00q, Ano00r, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano02r, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano02f, Ano05s]. Parabolic [RR02, SPW^{+}00, ARRS09, AKV06, BR09a, GK03, HJ09, JH08, MP08, RSTB03, SK05, UPKN09, ZWS07, ZSP08]. paradigm [CBGI09, GK05, LSW06, OK06b, SW04b]. Parallel [ABHT03, BADG00, BLW01, BD01, CWWZ00, CP06b, CP06c, CGSS00, DPR00, DSB06, GTD00, GTD01, GW02, HS08a, KP00, LJ09a, MC04, dlFMBdlFM02, Mu02, PM00, QRHD00, RXH02, TZ03, TZL05, WZ00, ABZ^{+}08, BPS03, BBW06, BCM09, BDS07, BUEG06, CTW^{+}08, CCV03, CHB09, CWD08, CLS09b, EHST03, EHS^{+}08, FHD^{+}09, FMD^{+}09, GHB03, Gib04, HVAC09, HC05, JA08, KSHS08, KKD08, KS07, LHR^{+}07, LSS^{+}09, LLRP09, LQ06, LZH^{+}06, MC06a, MMKP08, MG05a, MC09, MC03, ML04, PSCB08, PPCW06, PLS^{+}09, QFR04, RKE^{+}07, SWB^{+}06, SDR07, SLG^{+}03, SA06, Tol07, Tol08, TDGP06, WGSL06, WZ03, XDB09, YYT05, ZH09]. parallel-wall [BBW06]. parallel/recursive [Gib04]. parallel/unstructured [LZH^{+}06]. parallel/unstructured-multigrid [LZH^{+}06]. Parallelization [DDK06, FS00a, HJFW04]. Parallelized [OMK09, ZWL02]. parallelizing [WE05]. paramagnetic [KHdT^{+}08, KM08b]. Parameter [Zad08, BV05, CGSR08, DR09a, FP08b, LJSM08, Sha05, TW03, vdDA06]. parameterization [CHG^{+}07, LTWW07]. parameterizations [DAJ07]. Parameters [GST00, HFO01, AB05a, BLW04, EN06, KFIG06, Kou09]. Parametric [Mac00, XS07, BMN05, BGN07, BGN08, BBK07, Kou09, LRMB08, SK07a]. Parametrized [Pop00, Ngu07, Ngu08]. Parasitic [JTB02]. Paraxial [QS01, DDGS09, GS09a, GGRS08, QL04]. parity [VBJ08a, VBJ08b]. parity-mixed [VBJ08a, VBJ08b]. Part [CC07, CFGK05, CY05, FCGK05, HdGKG08, HT03, IX09, KM07b, KK05c, KK05d, LG09, Lio06, NMM^{+}07, NMH^{+}07, PFSL07, SFDL07, SD05a, SD05b, TOZP03, Tol08, VBJ08a, VBJ08b]. Partial [AGT02, ABGV02, BCOS01, CKL00, FDK06, Hua01a, MF01, PCS^{+}09, VB00, AGT05, AL06, Asl04a, BV05, CGP05, DI09, GBS06, KFIG06, LP04a, LCdCN^{+}03, MK08b, MSO04, Ngu07, Ngu08, RBvdV08, RM08, SS08, SRNV07, XLS09a, YZ07, YZW05]. Partially [PWM06, Kwo08, MD04, Zad08]. Particle [AL01, Azm02, Bal02, BZ08, Bow01, CVB00, CPK02, CL01c, CSP01, CB02, ELC02, EFFM02, FPC^{+}00, FR02, GVT01, HKKS^{+}01, iI02, KB00, Lap02, LM01, LDL^{+}09, LK01, Par02, PM02, PH09, PPCW06, PFB01, QRHD00, RH01b, SSW01, Sni01, SPC01, Ver01, VCTS02, WK01a, WLC^{+}08, Wea09, WHV^{+}00, ZF02, ZKK01, ASPB03, AWK07, AMP09, ADR08, BST03, BZ04, BHR06, BOT05, BB09b, BB08b, BB09c, CGSR08, CL08a, CPKW09, CFL^{+}03, CL08c, CP03c, CP07, CL03b, CEL06, DMHP07, DL03a, DFV08, DZ09a, DST07a, ECL02, Eld07, ESE07, FS09, FKV08, FHLO08, FHD^{+}09, FHLK05, Fox08, FG06, FM06, GS09b, GK02, HGBH03, HD07, HNF07, HJKO08, HK05, HK08b, HK08c, HX05, HS04, HT03, HDR^{+}06, HRV08, HS07b, HHM04, IITV07, JH06, JD09, KZWY09, KFV07]. particle [KP05, LJM^{+}06, LGKP07, LK07, LMX^{+}08, LWDA09, LZ09b, LZ09a, LM03a, LDW07, LKO05, LMH07, MP07b, MP08, MC06b, MCP03, MHW05, NBLQ09, NLLE06, NFA03, OK06b, PK05, PMP08, QFR04, SS09b, SWB^{+}06, SWG08, SB06b, SSB07, SK08a, SK07b, SBC04, TOZP03, TM05, TMSW07, TKH09, TU04, VHI05, VHI06, Wal03, WYS09, WGCR07, ZK04, ZH04, ZP05, ZIP06, ZB07, ZL09, ZKS^{+}09, ZZ09]. particle-continuum [SB06b, SSB07]. Particle-Field [SSW01]. particle-flow [AMP09]. Particle-In-Cell [Lap02, Bow01, CB02, LDL^{+}09, Par02, QRHD00, Sni01, SPC01, PPCW06, CP07, FHD^{+}09, FG06, GS09b, HDR^{+}06, IITV07, JH06, LWDA09, OK06b, QFR04, SK07b, WGCR07]. Particle-Laden [WK01a, JD09]. particle-level [TU04]. particle-localization [HNF07]. Particle-Mesh [FR02, LK07, LKO05, SWB^{+}06, Wal03]. Particle-Particle [FPC^{+}00, TKH09, Wal03]. particle-solid [HS04]. particle-source-in-cell [JD09]. particle/finite [ZH04]. particle/finite-volume [ZH04]. Particles [AKV00, Gut00, HHL00, ADS03, AGW07, CGL08, DTS05a, DTS05b, FM04, GH09, Hew03, JVVS07, KHdT^{+}08, KMSH08, LK09, LKMK09, PH06, SK08a, SP04, SKW05, VMN07, VK05b]. Particulate [GPH^{+}01, BCM09, DM03, DGMN03, FM05, LM03b, LMK09, Mar09, PH08, RMG^{+}09, SP05a, Uhl05, WT07a, YSW06, YS07b]. partitioned [BNV08]. partitioning [MG05a]. parts [MN04, MN17]. passivation [GKJW07]. passive [BS06a, RSM05]. Past [HGM^{+}00, PW00b, PW01, CHBS04, DCK08, GPH^{+}01, JD04, KR09a, MAL09, Pai01, SLC07]. Patch [SKR06, KPP09, PDHP07, SO08]. patch-based [PDHP07, SO08]. patch-refined [KPP09]. patches [CLS09b]. Path [PWW00, Sto07, BLL03, CLMRP08, CM06, Wea09]. path-consistent [CLMRP08]. pathological [LJ09b]. paths [Liu08]. Pattern [HKV01, WWVG00, MWM03]. patterns [LT09a, SM09a]. PCICE [Ber06a].

PCM [FWK08]. PDE [CFS09, EV03, FSS03, FP08b, GS03a, HMR08, PBH04, Pro08, SBA07, SPLM09, TCO^{+}04]. PDEs [AC00, ACY00, BMRS01, BMRS02, DGRS08, Dri02, JW09, Kro05, PL08, Ram06, RSTB03, RMO00, SV07, VBJ08a, VBJ08b, YBZ06]. PDF [LM03a, MCP03, AJ09, BFB08, CP03a, CRAG07, JPMC01, JML^{+}01, LM01, LLRP09, MPC01, MPC02, RJ06, ZH04]. Péclet [MHS02]. penalization [CBGI09, KDK^{+}07, KS09, LV07]. penalty [DDH05, DLD08, GLLN07, GLLN09, HH08, JLL^{+}06, PR06, Sha05, WG09]. penalty-projection [JLL^{+}06]. penetrable [BG09, HB05b]. penetration [TU04]. Perfectly [DH07, Doh09, GGOB04, Hu01, Hu05, Kan02, RJ04, SP05b, Vay02, BFJ03, BHNPR07, CLL^{+}07b, GKD09, HLL08, Nat06, OK07b, PL09a, Rah04, ST04, Zhe07, dHRvdB07]. perforated [ME09]. Performance [ABZ^{+}08, LSS^{+}09, Pir07, STD^{+}05, SS01a, WPM^{+}02b, CGMS03, CFR08, DSM09b, NZ07, OK07b, PAD07, QKS06, WdND06, XYK05, dHRvdB07]. period [AKH06]. Periodic [BC01, TB00a, ZF02, CdHST08, Dur08, EZ08b, FCT07, GHB03, HB05a, HHM04, JCT07, KMS03, KH07, KG03, LCG07, LCCG05, MKOW04, OJW06, ON08, SNGAS04, SROCFF03, SZB^{+}07, SAK05, SLV09, vdV08]. periodized [KR09c]. PERM [MP08]. permeable [HST09]. permittivity [ZC09]. Perspective [ÉGP09]. Perturbation [LTZ03, UH01, ES03b, MDR07, Nic09, YHCD05, ZL04]. Perturbation-based [LTZ03]. perturbational [GB08a]. perturbations [CP03c, RV09]. Perturbative [NR01]. perturbed [AV03, CJR04, Jao07, LCW04, Moo07, SM06a]. Petrov [LSJA05, Ma05]. Petviashvili [LY07a]. Phase [AG09, BEA09, CFA01, ČPT01, Cha09, Cle00, EKK02, EF02, GS02, GW02, GW01, GP00b, JLCD01, MR00, MR02, MP01b, NDG05, OCK^{+}02, PS01, SP00, WW00, WHV^{+}00, Xu01a, YSC01, YZF^{+}06, AW04, AMS03, BCB03, BFC04b, BIS07, BS03a, BJP04, CET09, CR05, CA06, CW08, CQRW05, CMR08, DM03, DDK06, DP07, DP08, DSS07, DS09a, DDS09, DLW04, EGHE06, EKP06, EKP07, FRS08, FYH^{+}06, GFG09, GCNB07, GR07, GD07a, Her05, Her08, HZ07b, HA06, HA07, HA09, HWWL09, HH06, IKL^{+}08, IOTK04, IS04, IH04, JLT06, KSO^{+}05, KW03, LL05, LCB09, LMS08, LM03b, LHGF05, Low05, MR06a, Mar06, MKKY06, MR06b, MDS03, Mou04, MG05b, NMG09, OLA08, OK05, OKZ07, QS07, RMB07, RMG^{+}09, RJM07, RMF08, RM07, SWK06, SY09a, SS06a, SYC09, SNLS03, SL07c, SB07]. phase [SXyWX09, SSH^{+}07, TLK07, TMB07, TBT^{+}09, TGB^{+}07, TJLT08, XMT06, YZ07, YFLS06, YC06a, YC06b, YHCD05, YSS07, YF09, YE05, YZF07, ZDD09]. Phase-Accuracy [MP01b]. Phase-Field [GW02, MR02, BEA09, NDG05, YZF^{+}06, BJP04, FYH^{+}06, LCB09, RJM07, SY09a, SB07, TLK07, TBT^{+}09, XMT06, YFLS06, YHCD05, YZF07]. phase-lag [MKKY06]. phase-screen [DS09a]. Phase-Space-Based [OCK^{+}02, CQRW05]. phase/vapour [BW07]. Phenomena [BMS00, BP09, HGBH03]. Phenomenon [PD01, Boy05b, DMG04, JS07, Lur07, SBA07]. PHOENIX [BvdHKG07]. phonation [LMZ^{+}08]. Phonon [GT09c, VP00, GS05a]. Photon [Cul01, Bar04, BS07, Cha07a, DMBS05, FEL^{+}05, WMH07]. Photonic [CD00, DGP00, SP05c, DD05, DBB06, ON08, YLA08]. photonics [DH07]. photothermal [CR08]. Phys [ABRR09b, CL08b, DD03a, HMS08b, HY11, HLWW06, JJGL07, Lau06, LM03a, MKM04, MN17, Mil07, SM09b, SCC^{+}03a, WZL09b, dTWD09]. PHYSALIS [HT03, TOZP03, PO01]. Physical [BBR01, FLG01, Hu01, AS03a, JLOT05b, JLOT05a, KDOO05, MR03, MK04b]. Physics [GTD^{+}02, MKR00, CP06b, CP06c, LL07, MK03, PNMK09, RWMK03, SHPC09, YKK08]. Physics-Based [GTD^{+}02, MKR00, MK03, PNMK09, RWMK03]. PIC [ADS03, CN08, CK07, GCLB04, HKM07, HM09, Hew03, Kwo08, Lar03, SHPC09, TPR05, TS07]. Piecewise [BSP06, CS08a, KKGL01, SPW^{+}00, UPKN09, BAYZ08, BMDS05, ET06, GR04, Lab09, LTL^{+}09, TLL^{+}08, WA08, YJ06]. piecewise-linear [GR04, Lab09]. Piecewise-polynomial [CS08a]. piezoelectric [LTL^{+}09]. pinch [GGS09]. pinch-off [GGS09]. pinching [YFLS06]. pinching-off [YFLS06]. pinchoff [NS04]. pinned [VCG03]. pipe [MT03, XLP05]. Pipelines [DC02]. pipelining [FHD^{+}09]. Pitaevskii [BJM03, CORT09, CJK^{+}03]. planar [APP^{+}07, AH08, CJR04, JOS06]. planar-symmetric [CJR04]. Planck [BC02b, DDFT09, FP02, Lem00, PRT00, CBKM00a, CBKM00b, Den07, DWLM09, KB04, LC03, UL06, WO05, WO09, XCRX08]. Plane [Dar00a, MM01, Mie00, PG02a, PC02, BO09, DVHM05, Gab07, KMS04, SL07c, Woo06]. planning [CM06]. Plasma [CYKC01, DGH02, GBS00, GVT01, HHCL01, HMM02, hLA01, OL01, RXH02, DMR09, DNS08, GHB03, GLS03, Hag07, HLS06, HDBW05, HF08a, Hum05, KSHS08, Kwo08, Lar03, LDL^{+}09, MT07a, Mot08, NC04, OK06b, SDD07, She08, SL03, SK07b, SS04, TPV07, TPR05, VTC^{+}07, WH05, ZK06]. plasma-vacuum [KSHS08]. Plasmas [Cul01, HKKS^{+}01, OMG02, SUW01, BvdHKG07, CDV05, CF04, DSM09b, GWF^{+}07, GGOB04, GLN06, GYKL05, GLT07, GL09b, HDR^{+}06, LCB04, LL08b, MD04, MSP^{+}06, PPCW06, SG06, XCRX08]. Plastic [GF02, HB02, Hun01, MC01, VQLZ04, ZVQ07]. Plate [BISS01, CHBS04, FKK08, LWW04, YKG04]. plate-mantle [FKK08]. Plates [SCD00, GA09, JA08, LSJA05, MC06b, ME09]. Platform [DPRS01]. Plentiful [CHM08]. PLIC [LHGF05]. PML [LZC04, Rah04]. PMLs [QT08]. POD [BC08, BBI09, CBS05, SK04a, SKXK05]. POD-assisted [SKXK05]. Point [Bar02a, Par02, SMP01, TB00b, VDM^{+}02, WLE^{+}00, AFGM07, BM05, CP03a, CWJ07, CWYM08, DVHM05, Eg07, FP08a, KK07, MDM03, PHKF06, Tan05b, WG08, ZZVM08, dSM05, dFJS09]. Point-Centered [SMP01]. Point-Set [TB00b]. points [BCEG07, HAP05, LC06b, SY09b]. pointwise [SB06c]. Poiseuille [DKS^{+}03, PG02a, Sie00]. Poisson [WO05, WO09, And09, AMLC08, BR01, BC02a, BC05, BJ09, BD01, BZ09, CGMS03, CSS00, CDV07, Dys01, GS06a, GFCK02, GM06, GBGM01, GS09c, HPS06a, HZ02, IKS^{+}09, JM05, Lai02, LLY05, LFK00, LW07, ML05, MP01a, MP02, MCGV04, Mil08, MGC06, NLLE06, PB00, Poz01b, SSN09, SBCL06, VVM05, WWC07, WZ09, Zha02, ZS01]. Polar [CSS00, CL02, Lai02, MC00b, SR09a]. Polarization [CD00]. polarized [GCW07]. poles [BM05]. Pollutant [ZKK01, BES07]. Pollution [FCB02, SD00]. Poloidal [BT07a, BT07b]. Poloidal-toroidal [BT07a, BT07b]. poly [GIA^{+}07, GIA^{+}08]. polycrystalline [CP05]. polydisperse [LMV04]. Polygonal [WS01, GL08, LSSV07, LSV09, LS05b, YS08, ZL08a, dVGLM09]. polyhedral [AS07, BAYZ08, CT07, LSS06]. Polymer [GIA^{+}08, AKH06, CFM09, CKPW07, FT06, HLFB07, LMK03, LWF^{+}08, SXyWX09, VC03, YFBH07]. Polymers [FS00b, SK06]. Polymorphic [GLMH09]. Polynomial [DGF09, ABHT03, BS06a, Boy05b, CS08a, JS07, LX09, MN09b, PW07, QM03, RM07, SZ05, WK05, XK03, YS06, ZL04]. polynomials [Boy04, LBS^{+}04, SR09a]. polysaccharides [TLAD04]. pom [APP^{+}07]. pom-pom [APP^{+}07]. poor [FRS08]. population [CPKW09]. Pore [ZF02]. Pore-Scale [ZF02]. poroelastic [BQQ09]. Porosity [JWSC00, JW02, JW03, RC06]. Porous [CS01c, CS00, Str01b, WLE^{+}00, ZF02, AT09, BW07, CP05, CDE06, GZ09, GH08b, HJ09, JLT06, KT07, LTZ03, LMS08, LH05b, LJ06, MZ08, MJT06, MP05, NL08, PC06a, RGS04, RC06, RM07, TJLT08, YE05, Zad08, ZL04]. posed [Meh04, Rah04, vdDA06]. position [Che04]. Positive [Jan00, LL03c, San01, BD06, CKLS05, EZ08a, ML06b, Waa09]. positivity [AB05b]. post [RC09b]. post-processing [RC09b]. Posteriori [MPP01, Dwi08, Ngu07, SVH^{+}06]. Postexposure [Li01]. Potential [HSK00, KT07, PC02, PO01, ABK09, BEE06, CS05, Del07, GD07b, HB05a, JR03, JR04, KW08a, KJ09b, MG07a, Mil08, NG06a, NLT07, OLLL03, OCFF08, Sac07, SCT06, SB09, YCL05, YF09]. Potential-based [KT07]. Potential-Theoretic [HSK00, HSQ03]. Potentially [CH01]. Potentials [HAAO00, CB07, HO08b, SH07b, Tau07, VB08]. Potts [VCG03]. Powell [HCG01]. Power [SM04, GB08b, Mil06, Mil07]. power-law-like [Mil06, Mil07]. powerful [Sau04]. powers [Boy09]. pp [Aza09, SM09b]. PPM [CS08c, SWB^{+}06]. PQM [WA08]. Practical [Hua01a, Abr06, LC06a, Yok07]. precipitation [TMSW07]. precipitator [LWW04]. Precipitators [BISS01]. Precise [BCGR05, KW08a]. Preconditioned [DMG00, Hua07, MVD04, MG02, AMLC08, BPS03, BH04, CP06b, CP06c, HLY09, HC05, Lee05, Lee09, LZH^{+}06, MYW07]. preconditioner [AHPT07, APT09, EHST03, EKP07, GGMN^{+}09, GT05, Gri09, KSO^{+}05, Lab09, LSS^{+}09, RWMK03, RKE^{+}07]. Preconditioners [BFG08, BT09, CdHST08, EHS^{+}08, GH03, STD^{+}05]. Preconditioning [Azm02, Ben02, DD09, Gla01, Hel05, HC08, MKR00, SC01, Yam05, HO08a, IF09, LZL03, Lee07b, MY03, MK03, NOG08a, PNMK09, PPB09]. Precursor [DW09]. predictability [CC08a, HMA05]. Predicting [CSC^{+}08, CGM07, LSK06, MS08b]. Prediction [CKL00, KIHM09, SMS08, APP^{+}07, CC08a, HL04, Ler06, Löh04, Lyn08, NLT08, PGN08, SW08c]. predictions [EHD08, HPD09]. predictor [CPKW09, CMSZ09, LRS09, TWYC06]. predictor-corrector [CPKW09, CMSZ09, TWYC06]. predictor/multi [LRS09]. predictor/multi-corrector [LRS09]. Preface [Kou08, OKV07, OT01]. Prefactored [Hix00, AZ03, PSG05]. Preliminary [DP00]. premixed [MMPB07, vdBG09]. preprocessing [SK04b]. prescribed [Xu08]. Presence [CS00, AV03, HP04b, LSK06, MLFG06, PIN09]. Preservation [Car01, BD08, IS04, MY07, SCC09]. preserves [CS08c]. Preserving [BS00a, CRB00, CL01b, DDSV09, LW01, RM01a, SH07c, TR02b, TS02, AS05b, AT08, AB05b, AMSZ03, AMS04, BLM08, BD06, CGL08, CS09, CLS09a, CDV07, DT04, JW06, Kok09, KWD07, KSW03, LFS07, LS05a, LLZ07, LW04, MS03, ML08, MOG09, PSG05, QM03, RGK07, SLV09, VV03, Wen06, XP04a]. Pressure [AMLC08, BT02, Cod01, JL02, LLIK01a, LLIK01b, MD01, Pet01, SS02, AMH04, CSL08, EZ08a, GR07, GS09c, JL04b, KIHM09, MTV08, MB04, MDS03, NVD05, NMS07, Pap08, Pon06, RVM07, RVDM09, Ros03, SAM05, Utn08, vBK03]. Pressure-Based [MD01, SS02, MDS03]. pressure-corrected [MB04]. pressure-correction [MTV08, RVM07, RVDM09]. pressure-invariant [vBK03]. Pressure-Poisson [AMLC08]. pressure-velocity [Pap08]. pressure/density [Ros03]. primal [AAC07]. primitive [BG07, CTT08, HHPW08, IHL03, KSO^{+}05, PHW08, SHTB09]. primitive-equation [PHW08]. primitive-variable [SHTB09]. Princeton [KN09]. principal [LM08b]. Principle [Rom02, SSSWD00, Abr06, KSS09, MGS09, SPGR06]. principles [WD07]. prior [RBT03]. priori [KK09, Ryc05]. priority [Pau07]. probabilistic [FWK08]. Probabilistically [ABLS05]. Probability [Pop00, MJ07]. Problem [AKV00, AQV02, BS00b, BSB01, CSV00, Del01, FS00b, IYI^{+}02, IFZ01, KLvBvL02, LR01b, MPP01, MM01, MN02, Mit00, Poz01b, PG02b, Stu01, TK02, WB01, Abr07, Abr09, AW04, BL09a, BBD04, BFC04a, BCE^{+}09, BTT08, Bia03, BO04, BEA09, BP04b, CFS09, CT08a, CDR09, CMG09, DB04, DHM03, GH03, GF05a, GKL03, HEN09, HZ07b, Hoh06, KH09, KNH05, Lee07b, Lee07a, Lee09, LS07, LSS06, MRRS05, Meh04, Mil08, NLT07, QCGQ03, RVM07, SWK06, SL04, SY09b, Sou09, SRX07, TPV07, VP09b, XSG04, YYF09, YE07, dVGLM09]. Problem-Independent [BSB01]. Problems [Alb00, AL01, ADK00, ADK02, BR01, Bar02b, BMR01, BCOS01, BS01, BK01, CWT00, CM00, DCS00, DKX00, DKX01, DFT01, FGG01, GP00a, GK01, HAAO00, HFO01, Kan02, KJ01, LLH02, LTZ02, MR00, Man02, NC01, OKL01, OS01, PR01a, PL01, RW00, RRL01, ST01, VG01, VSMW01, ZRR00, AE03, AM03, AM04, Abg06, APTJ^{+}04, ABLS05, ARRS09, AFGM07, ACR08, AG09, AQ07, BS04a, BB08a, BDRT09, BBD04, BFC04b, BBMB07, BHNPR07, BM05, BF08, Bey09, BS07, BG05a, BS05, Bor07, BSLN09, BLM03, BG05b, COV04, CT04, CC03, CBGI09, CELS07, CS08a, CXZ09, CS07c, CHG^{+}07, CS06, CS07d, CT07, CFJ09, DPRN05, DPRN06, DIL03, DT03, DG09, DR09b, EZ08b, Eg07, FM04, FF03, FCGK05, Gab07, GZ07b, GT05, GN07, GM04, GP05, GK04]. problems [GL08, Hab04, HJ09, HO08a, Hel09b, HMMR04, HY09, HY11, HF08b, IG05, JBHK08, JVVS07, JLT03, JH08, JC06a, JC06b, KP07, KK03a, KW08a, KZ04, Kau03, KEB^{+}07, KFV^{+}05, KS08b, KKO04, KPP07, KSS09, KS07, Lap04, LSA06, LZL03, LZ09b, LCW04, LHZW05, LDW07, LNXNTX09, Low05, Ma05, MTV08, ML05, MS08a, MMS04, MR07a, MNR07, MN09b, MSO04, MY06c, MR07c, MG05b, NPH09, NU09, ND04, NG06b, NL09, ON08, ODCK07, Ovt08, OMK09, Pap08, PS03a, Pir07, PSM08, RMGK04, RCB05, RDPN07, SKR06, SLV09, Shy04, SC09a, SL07c, SM06b, SHPC09, SPLM09, SN06, TWYC06, TD07, Tor03, TB04, TY07, UBRT07, VGCN05, VK05a, VSH04, VZSL07, WFTS05, XMT06, XHW07, Yam05, YAvdB^{+}08, YH07a, YCL05, YJF^{+}06, YZW07, ZG08, ZSP08, vOP04, vdDA06]. Procedure [DIV00, JK00, FDD07, LP06b, MKKY06, Mil05, Mil06, Mil07, ML06b, SHPC09, UL06, WYS09, XMT06]. procedures [BNV08, Cam03, CTW^{+}08, Hua07, Roy05]. Procesi [FL09]. Process [JK02, LRN^{+}02, Li01, Cam03, GS03b, KK05d, LL03b, LLTA07, YKK08]. Processes [LBD02, AT09, Chr04, FLB03, KMV03, Lau06, MDJS07, SS06b, SL06, VSV03, VK04]. processing [ALT08, FSS03, RC09b]. processors [GD08, KWBH09]. produced [KFH^{+}04, KFIG06]. product [Gel06]. production [IR09]. Products [CSV00, DP07]. Profile [YXU01, LMK09, UYK^{+}04]. profiles [CP07]. Program [BSJ01, WBM09]. programming [CVE06, DTSC04]. Progress [Jan08]. progressive [CF06a]. Projected [SWTM01]. Projection [AGP01, AP02, BJ00, BJ02, BCM01, CM00, ERT02, GQ00, LKNG01, LRN^{+}02, LMS02, LB03b, MC00a, TC01b, VSMW01, WGCE01, AV05, GBC06, Gri09, GF05b, HO03, JLL^{+}06, KKM08, LRZ04, Löh04, MZ08, MCG08, MG06, MK07, Ni09, PFSL07, SFDL07, SB06c, TC07a, Utn08, Vos06, XSL09, YP06, YSS05, ZP06]. projection-type [Löh04]. Projections [SS01a]. Projective [KEB^{+}07, SDD07, GK03, RMGK04]. Prolate [Boy04]. Prolongation [TR02b]. proof [Bea08, Boy06, KS08b, WZ07]. Propagating [SFW00, AS03b, BBF^{+}08]. Propagation [BM01a, Dur00, ERT02, FT01, GHV00, GKL00, HHCL01, HK02, LL00, LMSV00, MN02, MHS02, Noe00, RTT01, Wee02, BP09, BG05a, BS06a, CN08, CHG^{+}07, CBI^{+}04, DNS08, DS09a, DDGS09, EV03, FCJ08a, GD06a, GGOB04, GGRS08, HLS06, HSQ03, HPS^{+}06b, KT06, Lau04, LNGK04, LKNG04, LK07, LTE07, MN06, MHI08, MR04, Pir07, PSG05, RBL04, Ros06, Shy06, Thu08a, Tok06b, Vol04b, XS07, ZB07]. Propagational [CSV00]. Propagator [WH05, IH04]. propellants [SMGJ09]. propelled [HK08b]. proper [CRB^{+}08]. Properties [JMP02, KMJ01, Per00, Saf00, Saf02, Vas00, ZSP02, BIS07, Ber06a, GMH06, HR07, Jor07, LKE04, NE05, Pir06, RH05, SW08b, SVB09, VBL07, XLM07].

Property [VS02, WP09, XS05a]. propulsion [SMP09]. PROST [RR02]. protein [GPVB07, MSP^{+}06, XJ07]. Proteus [FM05]. Protoplanetary [dlFMBdlFM02]. Prototypical [VR02]. provable [GGF03]. Pseudo [Gom08, KvdVvdV06a, WPH00, HSBG05, HL07b, KKS05, SO08, THL06, YYT05]. pseudo-compressibility [KKS05]. Pseudo-spectral [WPH00, HL07b, THL06, YYT05]. Pseudo-time [KvdVvdV06a]. pseudo-timestepping [HSBG05]. Pseudo-wave [Gom08]. Pseudopotential [CWWZ00]. Pseudospectral [BRB03, KT05, BS08a, BM05, BS05, Boy03, Boy04, BP04b, CB03, LT09a, PSD09, SZLW06]. PSM [ZWS07]. Publisher [Ano03y, Ano03z, Ano07-32, Ano08-51]. pulsating [HAP06]. Pulse [HHCL01, XS07]. Pulses [SFW00, Sau04]. pure [BACFT05, De 04, NDG05, YU05a, YSO07]. pure-compact [BACFT05]. purely [Jao07]. Purpose [DPCV02, ALT08, Kuz06]. PVM [dlFMBdlFM02]. pyramid [HLWW04, HLWW06].

QALE [YM07]. QMR [CP06c]. QR [Boy02b]. QR-Factored [Boy02b]. Quadratic [DDS09, Dur00, CHL09, CVE06]. Quadratic-Finite-Element [Dur00]. quadratically [Gon07]. Quadrature [DKTT07, SCD00, DFV08, FLM08, Fox08, Fox09, HWL08, HO08a, MRC06, NG06a]. quadrature-based [DFV08, Fox08, Fox09]. Quadrature-free [DKTT07, HWL08, MRC06]. quadratures [Chr03, DR09b]. quadric [TW03]. Quadrilateral [HLKS00, SCD00, TC01b, ZYC02, BMT09, DPRN06, KT03, KT05, MJT06, NE05, YSS05, ZSC08]. quadrilateral-mesh [ZSC08]. quadrilaterals [PR04a]. quadtree [Gre04]. Quality [CBH03, SMO00, CSC^{+}08, KK09, RMV03, SDCC05]. Quality-improved [CBH03]. quantification [BPM06, CDE06, DEHL06, KG06, PDL09, YZL^{+}06]. Quantifying [HMA05]. quantitative [GR04]. quantities [AS03b]. quantized [DJ04]. Quantum [DE02, GM01a, JMK01, Lin02, MESV09, MK04a, RS02, dFGLS05, dFJS09, BP06, BMN07, BCCV09, BNNP06, CL05, CL08d, CLL^{+}07b, DDD05, DGM07, DDDC07, HLWW04, HLWW06, HWW07, JSCZ08, JN07, Kar04, KLW09, LYC09, NTO^{+}07, PA05, PVR07, Ram03, SB06a, SHY07, TW03, VTW^{+}07, Vos06, WBM09, YHSX07]. quantum-classical [BCCV09]. Quantum-corrected [dFGLS05, dFJS09, CL08d]. quantum-mechanical [DDD05]. Quartic [SKAS01, WA08]. Quasi [CK08, CDV05, MY06a, MOvL00, QS01, VD00, AI09, BS04b, DT03, EZ08a, FHD^{+}09, FHLK05, Gla05, HH07a, LCB04, LL09, MPD03, DDS09, Yeh07]. Quasi- [QS01]. quasi-elliptic [DT03]. Quasi-equilibrium [CK08]. quasi-geostrophic [MPD03]. Quasi-monotonic [Yeh07]. Quasi-neutral [CDV05, LCB04]. Quasi-One-Dimensional [VD00, AI09, BS04b]. quasi-positive [EZ08a]. quasi-spectral [LL09]. quasi-static [FHD^{+}09, HH07a]. quasi-steady [Gla05]. Quasi-Steady-State [MOvL00]. quasicontinuum [KLP^{+}09]. Quasihyperbolic [PS02]. Quasilinear [LAS01]. Quasineutral [LJM^{+}06, CDV07]. Quasirandom [AC01]. quenching [LLOT06]. Quest [Abg01]. Queue [JRS05, Pau07]. Queue-based [JRS05]. QUICKPIC [HDR^{+}06]. Quiet [Pet07, SCW^{+}09]. QWalk [WBM09].

r [MK08a]. r-adaptive [MK08a]. R3M [YH07b]. Radial [SUW01, TW03, CQO04, FW07, FP08b, HF08a, LJW07, MT07b, RA09, SC08a, TB09, WF06, Yin06]. radially [KLSW09, LLS09]. Radiation [BKR^{+}01, DW00, DV02, HSK00, HG03, MKR00, SMP01, TSG^{+}06, UH01, BKS07, BMDS05, BSP06, BSW05, BD06, Cha07a, Cha07b, DS05b, Dic08, DST07b, ED07, GT05, GCLB04, KLM07, KAS08, Lau04, LWG03, MR07c, MK03, OS04, Ols07, Ols09, RHPN09, RW08, RV07, SO08, dA04]. radiation-diffusion [OS04]. Radiation-Hydrodynamic [SMP01, SO08]. radiation-hydrodynamics [RHPN09]. Radiative [BS00b, DK02b, Gen01, KM03, LTK^{+}02, CS03, DL04, DUEB07, EULM03, FDK06, FKLY07, KNH05, MHB08, MELD08, MU09, MAN^{+}06, PS07c, Thö04, TFDK04, WHS08, WMH07]. radiative-transfer [DUEB07]. radii [CXB08]. Radio [HMM02, GGOB04]. Radio-Frequency [HMM02]. railway [LGN05]. Raman [BCG09, HS07b]. ramified [AST07]. Random [BJ00, BJ02, DDG02, FGOV00, FV01, LRN^{+}02, ARRS09, AZ06, DGF09, DI09, DC08, FG05, GZ07a, GZ09, GS09a, KF06, KKS07, LLTA07, MZ08, MS04, Pet07, ST06, SRNV07, VSV03, WP09, XS09, ZL04]. Random-Field [FV01]. randomized [ZGSD06]. randomly [HLRZ06]. randomness [WB09b]. Range [FPC^{+}00, BDS07, FT09, HPS^{+}06b, LTZ03]. range-limited [BDS07]. Ranges [GST00]. Rankine [JR09]. RANS [KMID05, KAS06, LS02a, SRM09, Tuc03, WK07]. RANS/LES [LS02a, SRM09]. Rapid [Lau04, Saf02, hRT02, BGR08, GPVB07, Nic09]. rapidly [KB08]. Rarefied [FS01, Mac01, Myo01, AK09, BB09c, GC06, KAA^{+}07, LZ04, Mac03, Myo04, SFX03, SBC04, VS09, VVS08, ZRS06]. rate [CMG09, OLA08, Tow08]. Rates [GGL^{+}01, GP05, Oh04]. Ratio [AJG01, Car01, BJP04, JA08, LL05, LF04, YZ07, ZSC06]. Rational [PSD09, BM05, BRB03, Boy05a, CFJ09, ZC09]. ratios [DSS07]. ray [Min07, THN^{+}07, RR07]. Rayleigh [CA06, GGL^{+}01, TM05]. rays [MR06b]. RBF [SPLM09]. re [LZ09c, BEE06, KM06, KM07b, PWS^{+}02]. re-entry [LZ09c]. RE-squared [BEE06]. Reacting [ML01a, CGP05, CP06a, DHM07, DBS06, FL07, LM04, MLS^{+}05, NS05, SK03]. Reaction [BJ00, Li01, MOvL00, RRV01, SWL00, SSC00, DC07, ELW04, GC06, HK06, HMR08, LRS07, LLOT06, MJ09a, Mad06, MM07, MMKP08, MG07b, Moo03, Moo07, Pud06, RSO04, RS05, RS09a, STD^{+}05, VSH04, XDC09]. reaction-advection-diffusion [Pud06]. Reaction-Diffusion [Li01, SSC00, LLOT06, MJ09a, Mad06, MM07, MG07b, Moo03, Moo07, RSO04, RS05]. Reactions [LX00, MEG02, SSC00, BCK09, KW03, OLA08]. Reactive [BM01b, HLZ02, JPMC01, MPC01, MPC02, dSAK00, AT05b, Bil05, BLM03, CFL^{+}03, DGJ03, HS03a, HS06, LMS05, RP08a, TMSW07, dDEK09].

Reactors [PCCD00]. Real [Mit00, OB02, DDDC07, RBK09, SH07b, dWKL07]. real-space [dWKL07]. real-time [DDDC07]. realistic [CP07, FHJK09]. realizability [PSMW09]. Realization [ZSC07]. recast [MYW07]. receptivity [DS06b]. Reciprocity [GHG01]. recirculating [RMG^{+}09]. reconstructed [VCG03]. reconstructing [YJ06]. Reconstruction [BISS01, DS08, HKS09, LS02a, RR02, SJ02, SR00b, ZC09, AS07, ÁDIM09, AMS03, AMSZ07, BO05, Bal09, Boy05b, Cap09, CR08, DDS09, GSB03, JS07, LSD07, LL04b, LHGF04, MP08, OK04, SGFL09, SAKDJ05, SS07b, TMD^{+}08, VB09, XLM07, XLS09a, XLS09b]. reconstructions [MLFG06, TB06]. Record [SSSWD00]. Recreate [SSSWD00]. rectangles [Bia03]. Rectangular [BdLL01, SZ00, CN05, HK08a, KPK09, NMM^{+}07, Ni09, PKD07]. recurrence [CL08a]. recursive [DSB06, Gib04]. recycling [LP06b]. red [LL06b]. Redistancing [CT08b]. Redistribution [RW00, AMS04, DG09, TLK07, TTZ03, WW04, WT07b, YT07]. Reduced [CKF02, CKPW07, DR09a, KG08, LP00, TLAD04, BKST09, BC08, CRB^{+}08, CBS05, KT07, LP07a, LB03b, MG05b, Ngu07, Ngu08, PCP08, QA09, SVH^{+}06, SK08a, YH07b]. reduced-basis [Ngu07, Ngu08]. reduced-order [BC08, CRB^{+}08, CBS05]. Reducing [Vil08]. Reduction [hLA01, SD00, ACGV07, AG09, BPMR08, BTWGvBW07, CK08, DL03a, FK07b, GZ08, LD06, MN09b, MKL06, MV06, PS07c, RA09, RFVP09, VP09a, VK05b, ZSW07]. Reference [KMS02, PK07]. Refined [GW01, DP09, FHW07, Her08, KPP09]. Refinement [Alb00, AGT02, Bal01, DGH02, DI02, FH00a, AGT05, AEP04, BC05, BV05, BFG07, BL05, CR07, CBH03, CBI^{+}04, CFJ06, FM06, HS06, HS08a, HG03, KAA^{+}07, KPP07, LP04a, LL04b, LK09, MCGV04, MC07b, MHE06, PSCB08, PDHP07, PCP08, PL04, PC06b, RFFP06, SRX07, TFD06, TK04, YF09, dTDI^{+}07]. refinement-based [CR07]. Refining [BH09]. reflecting [AB03, GN03, NN04, PWM06]. reflection [BS04d, Sof09]. reflectometry [dSHHM05]. Reformulation [iI02]. Regime [BJM02, BKR^{+}01, BC08, GC06, LQ09, LD09b, RB09b, SKK^{+}08, SE04, VLB09]. regimes [CGL08, FK09a, JD04]. region [BC08, HE07]. Regional [Ano08-50, Lap08, SDCC05, SM09b]. Regions [Bal02, CFA01, Cal02, MG07a, VRM07]. regressing [SMGJ09]. regridding [WAH09]. regridding-remapping [WAH09]. Regular [GC02a, LSSV07, YXLF05, dA04]. Regularization [Poz01b, BT07b, BHSV07, CT04, CCT05, MK08a, PBH04, SY09b, vdDA06]. Regularized [BOT05, ADE^{+}08, CFS09, FP08a, GE07, Kry04, SDT08, TS08, WKL07]. Regularly [HM08, Moo07]. regulation [JR07]. Reinitialization [HMS08b, HMS08a]. Related [RMO00, ON08, RSSL09, RSS09, Tol08]. Relation [CL01b, LP01, CS09, CLS09a, KK09, LS05a, PSG05, Ros08, SLV09].

Relations [CL01a, SZ00, YJ06]. Relative [Cam03, RS02]. Relativistic [SZS01, BH09, BLG^{+}08, DZ09b, FKV08, HJ07, HGB^{+}03, KQW03a, KQW03b, MK04a, QW05, SK08a, SA09, ZSW03]. Relaxation [BR09a, BBR01, Car01, JP00, LP01, Pal08, BKS07, BN09, DP08, Del03b, Küm04b, LSL08, Low04, PFSL07, PA07b, RSM05, Ros08, SFDL07, SPB09]. relaxation-projection [PFSL07, SFDL07]. release [Oh04]. relevant [Vos06]. reliability [LXM09]. reliable [EN06]. Remap [ALGM01]. Remapping [DB00, NJX09, KSW03, LS05b, MS03, SPM03, WA08, WAH09]. Remapping-free [NJX09]. Remark [LS07, RS00]. Remeshed [CPK02, CFL^{+}03]. remeshing [AZC05, BIVC07, MK08a, MV08, WG06, ZLAC05]. removal [LLB05]. Removing [HS08c, DL03b]. renormalization [HJFW04, Sti05]. renormalized [CMP07]. repair [LSW06, SW04b]. Reply [CKR01, LM03a]. Representation [AKV00, FF02, MR01, Sum00, BHP07, CCT05, DGF09, HW08, KAK03, SAKDJ05, TW05, TRSK09, TR07, WL06]. representations [SL06]. representing [BP08]. repulsive [ABK09]. requirements [MWM08]. requiring [BEPT09, SS09b]. rescaling [LLL07]. research [SK08b]. reservoir [LCH03, MC04]. reservoirs [SE09]. reshocked [LSD07]. Residual [Abg06, CRD02, DPRN06, LC01, RCD05, TS01, WB01, AM03, AM04, CS06, CS07d, EULM03, Nis07, RAD07, RB09a, Ros08]. Residual-Based [LC01]. residual-distribution [Nis07]. Residuals [Boy02b, CB07]. Resistive [CKF02, RVVL09, DZ09b, GTMC08, LTC07, ODCK07, PCP08, RSW06]. resistivity [GTMC08, HPS06a]. resistor [KF06]. Resists [Li01]. REsolution [KG09, Coe02, FCB02, GF02, GP00b, JK00, KB00, KT00a, KT00b, KMJ01, MWM08, MD01, Noe00, PW00b, PW01, SZS03, WPM02a, ZTZ02, ASPB03, BSKH07, BBCT09, BOT05, BTW03, CLG07, CC04, DE06, Del07, DDFT09, FF03, GLN06, GD05, HLS06, HH06, Jor07, JS07, Kou09, Kry04, KT04, LG09, LSD07, LNGK04, LYC09, LR03, MM09, OF06, SWK06, SYG06, SHY07, SJHM09, SL07c, TDWY08, ZW03, ZYHS07]. Resolved [DSS00, AMP09, Bar04, SMP09, TV08, WB09b]. Resolving [YT07, LTC07, NK08, SMT^{+}08, SSW^{+}07]. Resonance [OL01, GLS03]. Resonances [BP06, Lin02]. resonant [BNNP06, BS04d, DGM07, RMV03]. response [CDI09, HSZ04]. responses [WB09b]. resting [NLT07]. restricted [HST09]. Restriction [TR02b, KSGF09]. restrictions [KDW08]. restructuring [SS07b]. Results [OF01, SSSWD00, MPD08, NFvS^{+}06, NDG05, VBL03]. resumming [LBS^{+}04]. resurrected [HSC09]. Retrieving [LR01b]. Revaluation [IM07]. reverse [HS07a, RMGK04]. reversibility [DOW08]. Reversible [LR01a, RE07, PH09]. Reversible-equivalent-monomolecular [RE07]. Review [Roy05, FSS03]. revised [Wan04b]. Revision [Neo07]. revisited [LOK05, MC07a]. Revisiting [FLG01, Rid00, SD06, WE05]. revolution [FWW04]. Reynolds [Cor00, DKS^{+}03, DDH05, FG02, MT03, NMM^{+}07, NMH^{+}07, OTCM08, OVG07, PPDM08, SDGX07, Vik03, XP04b]. Rezone [KMS02]. Rezoning [Lap02]. RF [hLA01]. Rheology [LL06b]. Rhonegletscher [JHB^{+}09]. Richards [ZSWW03]. Richardson [RB06]. Richtmyer [LSD07]. Riemann [AW04, BDRT09, BTT08, BZW01, BH05, CT08a, DP07, Del01, Geo08, Gui05, HAI09, iI02, KLLJ09, KSW07, KLvBvL02, LMS05, LP01, Li05, LS07, MN02, Mig07, Mil04, MK05, NK08, QCGQ03, RBT03, SWK06, SL03, SHTB09, TT06a, Tor03, TB04]. Riemann-Problem [KLvBvL02]. Rigid [Bus00, GPH^{+}01, AMP09, BCM09, BGS08, CC08b, DMHP07, Eld07, Fas03, LKP06, San03, SP05a, TZL05, TLK09, TG06, VMN07, Vil08, WT07a, Xu08, vLAvdV06, vZS07]. Rigid-Body [Bus00]. rigorous [CY05]. Rigorously [OMG02]. ring [QFR04, SDT08]. Rings [MKM99, MKM04]. ripening [Hor06]. Rising [Dar00a, HL07c, HSL08]. River [SSL00]. RK [AHNS09]. RKC [VSH04]. RKDG [BAMD07]. Robin [BNV08, JZ08, LSA06]. Robust [Azm02, JWSC00, JW02, JW03, KLLJ09, KF06, MJT06, MG08, MLS01, BBI09, BB09b, CFR09, CL07a, CL08b, DCF^{+}08, FK07a, FE04, HNF07, HM05, KK07, LH08b, MTV08, NLF03, SP06a, TAL09, VPMC04, vDZ06]. Robustness [Ber06b]. rod [BCZ04, HO06]. rod-like [HO06]. rods [GPL05, LN09]. roe [KD09, GV02, Jan00, sKKRH03, sLwG08, RBT03, RMF08]. roe-average [KD09]. Roe-type [sLwG08]. Role [AC00, ACY00, AGP01, Mai04]. Root [Bor00]. ropelength [MR03]. Rosenau [RV07]. Rosenbrock [DCS00]. Rossby [CF06a, SD06]. rotated [NK08]. rotated-hybrid [NK08]. rotating [AB07, BW06, BvdHKG07, FBHV05, GG09a, Gir06, PK07, SJ04, SS07a, VBJ08b]. Rotation [HGM^{+}00, HF08a, LMN^{+}09, YGL05, ZSC07]. Rotations [Bus00, BPO07]. rotors [EHD08]. rough [Nic09, TX06, WWC07]. Roughness [WWC07]. round [Vil08]. round-off [Vil08]. Routh [MD02]. Rudy [TK04]. Ruin [SSSWD00]. rule [dSM05]. Rules [ADK00, ADK02, HvHHS05]. Runge [AHNS09, Bal08, BP09, Boy05b, BSB01, CFR04, Dri02, HL06b, HyLL07, KCGH07, KHV01, KWD07, KDW08, LX07b, Lur07, QS04, QKS06, QLK07, Rei00, STR07b, Tan05a, ZP06, ZQSD08, ZQ09].

S [JTL09, LNXNTX09]. S-shaped [JTL09]. Saffman [FS06]. Salpeter [Mai03, Mai04]. sample [FHJK09, HMA05]. sampled [Mil05, TPVG06]. Sampling [Pop00, Cam03, CTW^{+}08, KS08b, KLW09, Sto07, Wea09, vEB05]. satisfying [CkM07, KSS09]. saturated [AT09, GH08b, Vol04b, WGNT06, Zad08]. Savage [FNBB^{+}08]. scalable [BP07, ZG08]. Scalar [BTFY01, CDKP00, GBGM01, TS02, WL02, BD08, BP03, CRAG07, FS09, GD07b, Hub07, IA06b, KI05, LY07a, LRMB08, LFX05, RSM05, RBS06, Ros09, TT05b, Asl01, Kuz06]. Scalar-Tensor [BTFY01]. Scale [AS02, ATV01, BADG00, CR02, EKK02, Gra06b, KS02b, LR01a, VG01, ZWL02, ZF02, AE03, AHF04, AC09, BBMB07, BJ09, BSW05, BTWGvBW07, HBLD07, JLT03, JLT06, JL09, KM06, KM07b, KCMM03, KE09, LZL03, MN09a, OK06b, OK07a, PKKL05, PM08, PS03b, RWMK03, SHPC09, SK07b, UBRT07, VP09a, VK09, VTM^{+}08, WL03]. scale-invariant [KE09]. Scale-separating [Gra06b]. scales [AKP07, ELVE07, JG09, Ngu08]. Scaling [PC08, Abr09, SLG^{+}03]. Scattered [WF06, Dic08]. scatterers [DBF08]. Scattering [Bal02, BK01, CY00, DDF01, DFT01, GK01, Gro00, Gut00, Lin02, Stu01, AL06, ACR08, AG09, BHL07, BL09a, BHNPR07, BA03, Bot06, BH04, CJSS08, CLLG09, CDR09, DH04, DWLM09, DBB06, DR09b, FNS07, FHJK09, FWW04, GG04, GH08a, GK04, GK07, Hoh06, HB05b, IQT08, LZL03, Lee07a, MR07a, Nic09, OMK09, SZB^{+}07, TJ09, TC07b, TC09b, Woo06, YAvdB^{+}08]. scenarios [SP07]. Schedules [FH02]. Scheme [Abg01, BR09a, BW01, Bon00, BMS00, CBKM00a, CKR00, CKR01, DPCV02, GF02, HLS02b, HF00, KKR01a, Kul01, LBV00, LS02c, LC01, LM01, LX00, LW01, MP01a, MP02, MCCT02, MF00, MT01, MHS01, NTYT01, NTYT02, Nys02, OGV02, Ros00, SWL00, SZS01, Vay01, WH02, WPH00, Wee02, Xu01b, Xu01c, Xu02a, Xu02b, YP01, YL01, ZZ01, ZSP02, Zha02, de 00, vdSE00, AK06a, AT09, Asl04b, AM05, AB05b, AT05b, BS04a, BKS07, BMT09, BACFT05, BALW06, BAFL09, BF07, BG05a, BCCD08, BHvdV06, BN09, BM07, CFR04, CP03a, Cap08a, Cap08b, Cap08c, CFF07, CDDL09, CMSZ09, CL07a, CL08b, CHB09, CMG09, CS07a, CEH09, CSO09, CYS06, CL07b, CSL08, CS09, CSKD05, CDV07, DE06, DPRN05, DDSV09, DBBP08, DLD^{+}06, DS06b, DZ09a, DGRS08, Edw06, FM08, FL06, FH07, GPC07]. scheme [GN03, GN07, GL09b, HLS06, HS03a, HHC08, HJM^{+}05, HLY09, HH06, HGB^{+}03, Hwa03, IM05, IKS^{+}09, IA06a, IA06b, IAT08, IM07, IQ08, JAK05, KR09a, KN09, KOQ04, KH09, KL08, KK05a, KJ09b, sKKRH03, KK05b, KLLJ09, KPK09, KLM05, Lab09, Lar07, LK07, LD09a, LM03a, LF06, LLL07, sLwG08, Li08a, LLS09, LS05a, LLZ07, LYC09, Liu08, Liv07, LJ09b, LGM08, LSW08, LH08b, Mai09b, Mai09a, MTWW06, MY09, MY03, MG07b, MGC06, MC07c, MCP03, MPFC08, MVO04, MSB07b, Mot08, NSS03, Neo07, NMM^{+}07, NMH^{+}07, NJX08a, NI03, NN09, NF09, NS05, OX04, PKD07, PS07a, PH06, PH08, Pon07a, RLZ03, RS06a, Rom07, SBA07, SC08a, SJD05, SDR07, SVB09, SLV09, Ser09, SHWC07, SYC09, SY08, SA09, SCW^{+}09, SB03, SC09b, SFMP06, ST03a, zSW06, zS06, SCN07, SN08, Thö04, TXCD07, TE08]. scheme [TDGP06, TAL09, TCM05, TY07, UBRT07, VVM05, VU04, Vol04a, Waa09, WRu03, WZ09, WA08, WLC^{+}06, XP04a, XH03, XMT05, YMT^{+}04, YC09b, Yok07, YHCD05, YS07c, ZW05, ZWS07, ZSWW03, ZWS06, Zie04, VP09a]. Schemes [AC00, ACY00, BS00a, Bar02b, BCVK02, Bla00, CL01a, CL01b, Coe02, CDKP00, CR00, CRD02, Del02, DZ00, DLS^{+}00, EF02, FF02, FGG01, FSB01, FH00a, FSM^{+}01, GC01, GC02b, Gui02, HL01, HT00a, HT00b, Hix00, JP00, JL02, JMP02, KT00a, KT00b, LP01, Lio00, MF01, ML01a, Mie00, Nic00, Ohw02, Per00, Pir02, QS02, RB02, SV00, SHS02, STiST02, TK00, TS01, TH01, TS02, Vas00, VG02, VS02, WC01, WB01, XY01, Yua02, AM03, AM04, Abg06, APP^{+}07, AHNS09, AB03, AT08, AZ03, BTW04, BAMD07, Bal09, BRDM09, BAR08, BES07, BLM08, BBMB07, BRC^{+}09, BP09, Ber06b, BS03a, BR09b, BB04b, BBCT09, BP03, BL03, BD06, BK07, CVB06, Cap05, Cap06, CGMS06, CGL08, CET09, CLMRP08, CL08c, CJ09, CS06, CS07d, CP08, CZVS04]. schemes [DT04, DPRN06, DQ04, DJTT05, DK07, DKTT07, DET08, DBTM08, DZ09b, DOW08, Dwi08, EZ08a, EF03, ELW04, FDD09a, FDD09b, FK07b, FK09b, FW07, FMR09, GZ07b, GLM07, GSV09, GSV06, GGF03, GS03c, GS03d, HK06, Hei04, HAP05, HJJ09, HWWL09, Hub07, HWW07, JW06, JC06b, Jon05, JMC03, KCGH07, KI05, Kim07, KLK08, KPP07, KPP09, KQW03a, KQW03b, KT04, Kuz06, LSB04, LL09, LNGK04, LWW04, LFS07, LG03b, LG04, LJS08, LSSV07, LL03c, LW04, Liu05, LCS09, LD04, Mad06, MRS09, MSS08, NL08, NJX08b, NZZ06, Nis07, NPPN06, NXS07, NF09, OK04, OF06, PAD07, PK03, PYC04, PS04, PS08, Pir06, Pir07, PSG05, QS04, QS05, RBSL06, RP08a, RAD07, RB09b, RMF08, Ros09, RDPN07, Ros08, RS09b, SDM04, SGD03, SHA08, SD05a, SD05b, SYG06]. schemes [SZS03, SHY07, SZ05, STZ07, SS09c, SPGR06, SJC07, STR07b, SK06, Tak06, Tan05a, TL06, TFD06, TDWY08, TT04, TT05a, Tol07, Tol08, TT05b, THD09, Tor03, TB04, TA06, Tsu06, VTT08, VCZS04, WG08, WSYS09, WD07, Wel07, WAH09, WZ03, XS05a, XS06, XS05c, YMT^{+}04, YC09a, YHSX07, YS07a, Yeh07, YMWM06, YS08, ZJS08, ZT03, ZYHS07]. Schmidt [JS05]. Schrödinger [BJM02, XS05b, AMR06, AB03, ABK09, BY07, BBDE05, BIS07, CCJ07, DE02, Dem04, Doh09, FCJ08b, GM06, HyLL07, HJL09, IKS01, JLOT05b, KLSW09, LTE07, LQ09, LW09, Nas08, RSSL09, Sac07, SCT06, SKAS01, SS07c, zSW06, zS06, XHW07, ZKL^{+}07, Zhe06, Zhe07]. Schrödinger-type [XHW07]. Schroedinger [And09]. Schur [NPH09]. Schwarz [BIW08, CJSS08, HC05, ODCK07, PW00a]. Schwarz-based [ODCK07]. science [KG06]. scientific [Bra04]. Scour [DC02]. scrape [MT07a]. scrape-off [MT07a]. screen [DS09a]. screened [GH02, LJK09]. Screens [Kan02]. Sea [Hun01, LTD04]. seamless [ERVE09]. search [Pav07]. searching [Sus06]. Searchlight [BS00b]. Second [AT05b, BRL02, Boy02a, FB08, GC01, GFCK02, HLS02b, JLCD01, JTB02, JR03, JR04, KC00, KQW03b, LP00, MS03, Nis07, Ols09, PP04, RP08a, SPM03, TS01, VB00, YL01, AM03, AM04, AB05b, BS04a, Bea08, CP03a, CR09, DPRN05, GGF03, GHMP07, KSO^{+}05, KK05a, KDF07, KT05, LSZZ08, LRZ04, MZ08, MvW08, MN04, MN17, MG06, MG07d, MG08, Ols07, SZ08, SGFL09, SP05b, SCW^{+}09, SL06, Sus03, TMS06, TM07, YP06, Yeh07, YZW05, ZP05]. Second-Generation [VB00]. Second-Gradient [JTB02, KDF07]. Second-Order [BRL02, GC01, HLS02b, KC00, LP00, YL01, AT05b, FB08, KQW03b, MS03, Nis07, Ols09, PP04, RP08a, AM04, AB05b, BS04a, Bea08, CP03a, CR09, GGF03, MZ08, MvW08, MG08, Ols07, SGFL09, SCW^{+}09, SL06, TMS06, TM07, YP06, Yeh07, YZW05, ZP05]. Second-Order-Accurate [GFCK02, DPRN05]. sector [Boy05a]. sediment [RF06]. sediment-transport [RF06]. seeded [LD09b]. Segment [ERT02]. segmentation [RR07, XCY06]. segregated [NVD07, Utn08]. Seidel [ABHT03, CLS05, KK07, WGCE01]. Seidel-type [CLS05]. seismic [CFS09, CSMH05, HS07a, THN^{+}07]. selection [HL06a]. Selective [BD08, LLB05, OL01, RMSB09]. Self [CBC09, OK06c, OL01, RHPN09, RV07, RMO00, SUW01, SCC^{+}03a, SCC^{+}03b, VP00, ZSTC06, BEA09, DAJ07, DS05b, FY07, HK08b, JRS05, MAN^{+}06, SMP09, TSB03]. Self-adaptive [OK06c, DAJ07, TSB03]. self-adjoint [MAN^{+}06]. self-assembly [JRS05]. Self-Consistent [SUW01, OL01, RHPN09, SCC^{+}03a, SCC^{+}03b, BEA09]. Self-consistent-field [ZSTC06]. Self-Energy [VP00]. Self-Intersecting [RMO00]. self-organization [FY07]. self-propelled [HK08b]. self-propulsion [SMP09]. Self-similar [RV07]. Self-sustained [CBC09]. self-teleportation [DS05b]. selfadjoint [Heu03]. Semi [BS03a, Bon00, FF02, GVT01, GBB^{+}06, KWD07, KT00b, MELD08, NTYT01, NTYT02, RCB05, Str00, Str01a, Tol02a, Tol02b, WPM02a, WA02, XY01, XK01, BG07, BS08b, BLG^{+}08, BRB03, BL03, CFF07, CFR09, Cha07b, CWL08, DF04, GHB03, GPF03, GGP06, GD05, HS08b, IX07, LS03, LQ09, LH08b, LLC^{+}08, MBP07, NSS03, NZZ06, RBS06, ST04, SFMP06, TBT^{+}09, TOY09, ZWS07]. semi-circular [GGP06]. semi-classical [CWL08, LQ09]. Semi-discrete [KT00b, BL03]. Semi-Implicit [WA02, Bon00, MELD08, RCB05, WPM02a, CFR09, Cha07b, DF04, HS08b, LH08b, LLC^{+}08, MBP07, NZZ06, SFMP06, TBT^{+}09]. Semi-infinite [GVT01, BRB03, ST04]. Semi-Lagrangian [NTYT02, Tol02a, BS03a, Bon00, FF02, GBB^{+}06, NTYT01, Str00, Str01a, Tol02b, XY01, XK01, BG07, BS08b, BLG^{+}08, CFF07, GHB03, GPF03, GD05, IX07, LS03, NSS03, RBS06, RCB05, TOY09, ZWS07]. Semiclassical [BJM02, DDD05, GM04, Gos04, GM06, JLOT05b, JN07, LW09, SY08]. semicoarsening [LLY05]. Semiconductor [DE02, JP00, KMA^{+}01, MP01a, MP02, And09, BS04a, CGMS03, CGMS06, CBC09, CL03a, CL05, CLL^{+}07b, FH07, KGJ05, LSS^{+}09, SS08, VTW^{+}07, WHLL03, dFGLS05]. Semilinear [Dri02]. Semirelativistic [GTD^{+}02]. Semtner [MR01]. sensing [CLLG09]. sensitivities [FLE03, HPD09]. sensitivity [AJT04, BV05, BG09, LWG03, LP04a, NA08, PA07a, SDCC05, TMND07, VK04, WGNT06]. sensitivity-based [LWG03]. Separable [TNR02, KS07]. Separate [CFA01]. separated [KRT^{+}09, Oh04]. separating [Gra06b, LG09, SPB09]. Separation [AD01, OL01, OMG02, NU09, SNLS03, SG03a]. sequel [Lio06]. Sequence [WWVG00, HK04a]. sequential [BPS03, KS08b, LCH03]. Series [Che00a, Che00b, CKGL02, CL02, VP00, BO05, BRB03, Boy09, CKG04, Fou06, KTD03, NU09, NCS03, TDV06, VCG03]. SESL [GPF03]. Set [Asl01, BCMO01, CMK^{+}01, CBMO02, Cho00, EFFM02, HMS08b, KAIN01, KLvBvL02, LLdlP^{+}00, OF01, OS01, OCK^{+}02, PS01, SW00, Set01, SP00, TMB07, TB00b, AS03b, AS05b, AJT04, ÁDIM09, AA06, AHMS03, BHR04, BHSV07, COQ06, CM06, CT04, CBGI09, Che07, CCT05, CQRW05, CC08b, DMP08, DL03b, ETT05, GGS09, GCNB07, Hab04, HMS08a, HKO07, Her05, Her08, HK05, JVVS07, JCT07, KH07, LW07, LW09, LTWW07, LLC^{+}08, LTL^{+}09, LTM09, ML06a, MS08a, MRC06, MR06a, MGCR07, Min04, MG07c, MG07d, MV06, NJLA06, NLT07, NT07, OK05, OKZ07, PHKF06, QL04, RR07, SS09b, SS06a, SYC09, Sme06, Spe05, Sus03, TZ06, TZ07a, TZ07b, TBJ^{+}09, Tow07, TU04, UYK^{+}04, WLKW07, WSTW09, WYS09, Wen09, XLLZ06, YS09, YSS05, ZGK09, ZLAC05, ZL08b, vdDA06]. set-based [TU04]. set-boundary [GGS09]. set/ghost [DMP08]. set/volume [YJL^{+}06]. set/vortex [Her05]. sets [FSS03, GR08]. Several [ZDNP00, Ovt08]. SGS [NN09]. Shadow [IH04, ESD05]. Shah [ET06, RR07]. SHAKE [BLS08, Gon07, WE05]. Shallow [BC01, CX08, Che00a, FR02, Gir00, GHW02, Hor02, LBV00, LBV01, Lay02, LLIK01a, LLIK01b, Tol02a, Tol02b, TTSG01, VS02, Xu02b, ZCMI01, AB07, AB05b, BES07, BRC^{+}09, BTT08, CVB06, CHL06a, CL08a, CGRGV^{+}04, Che03, CLS09a, CZVS04, DJTT05, GPC07, Geo08, GPF03, GW05, GD05, HC08, KJ09b, KLM05, LHD05, LGHD08, LS03, LMNK07, Mea04, MGNB09, NI03, NPPN06, NXS07, RAD07, RB09a, SS03a, SHTB09, TOY09, VTT08, XS05a]. Shallow-Water [BC01, Che00a, LLIK01a, LLIK01b, Tol02a, Tol02b, Xu02b, ZCMI01, CHL06a, CL08a, CLS09a]. Shape [AKL^{+}08, HS09b, LS02b, LTWW07, LTM09, PS03a, AS05b, ADR08, BG09, CKvT07, DAJ07, FP08b, Hab04, HPD09, HSBG05, HKO07, HWW07, LSSV07, LKMU05, LLC^{+}08, TW03, WLKW07]. shape-material [BG09]. shape-regular [LSSV07]. shaped [BCDW06, JTL09, MTH08]. shapes [HKS09]. shared [HJFW04]. shared-memory [HJFW04]. Sharp [LKMU05, MKLU05, NLT08, SB07, UMRK01, YU05a, YS09, YSC01, FCD^{+}06, GMD07, GCNB07, MDB^{+}08, OSK09, SSH^{+}07, TU04, UTBV03, WK06, YZW07]. sharp-edged [YZW07]. Sharp-Interface [YSC01]. sharpening [CET09]. Shaw [FS04, KW08b, LLL07]. Shear [ELW01, TC01a, BZ04, BCZ04, BIVC07, GH09, HO03, JOS06]. Shearing [LAS01, BM06]. Sheet [GC02b, LK01, Nit01, FM06, Her05, SDT08]. Sheets [Nie01, Alb09]. shell [CJ04, Liv07]. shields [BCM^{+}07]. shift [HHMK05]. shifted [AHPT07]. shifted-Laplacian [AHPT07]. shifts [BM05]. ship [Wan05]. Shock [AS02, BSJ01, Boy02b, FSS03, Han01, Lio00, MC02, Pir02, STiST02, TNGH02, Tót00, VG01, Wu01, Wu02, AM05, BdCB09, CLMRP08, CC05, DLD^{+}06, GA09, HMM08, HJJ09, KFH^{+}04, KFIG06, sKKRH03, KLLJ09, KH08, LM08a, LSK06, LKY03, LRS09, Low05, Pir06, SB06b, SM05, Sur05, TDWY08, TY07, UTBV03, VS09, Vol04b, YT07, KKR01b]. Shock-Aligned [KKR01b]. Shock-Bubble [Han01]. Shock-Capturing [STiST02, TNGH02, Tót00, BdCB09, DLD^{+}06, KH08, Pir06, TDWY08, TY07, UTBV03, Vol04b]. shock-induced [YT07]. shock-stable [sKKRH03]. shock-tube [Low05]. Shock-Turbulence [Pir02, CC05]. shock-wave [KFH^{+}04, KFIG06]. Shocks [DCV^{+}01, Sun00, YC02, DLT09, FL07, HP04b, IR09, JD09, MLM09, PFSL07, SPB09, SH07a]. shooting [ZK06]. shoreline [Che04]. Short [SFW00, CWL08]. short-wave [CWL08]. shortening [CFF07]. shrinkage [YZF07]. sided [HH07b, RB06, SR09a]. sign [MS03, SBA07]. sign-preserving [MS03]. signal [dSHHM05]. signals [Mil05]. Silicon [GR01, Rom02]. SIM [NLT08]. SIMD [DPRS01]. similar [RV07]. similarity [SB06a]. Simple [Fre00, Kul01, Lai02, OF06, RM01a, SPB09, SZS01, STiST02, BP08, DFG^{+}06, LL09, Mig07, NK08, RM08, Yok07, EKP07]. Simplex [COQ06]. Simplicial [Min03]. Simplification [Ber06a]. Simplified [FMO00, LTK^{+}02, RLB02, FKLY07, KL06, VGCN05]. Simulate [DPRS01, HMM02, Chr04, EKP06, MV06, PSC04, Sam09, SLC07]. Simulated [PA00, Pav07, WGNT06]. Simulating [Alb09, BBF^{+}08, Cho00, CR02, DLW06, GK02, HHL00, HDBW05, PR00, PK00, TS04, UMRK01, AGW07, BGS08, BIVC07, BB08b, Che03, DMHP07, DMP08, DP09, Dur08, GFG09, GSB03, GS05c, Gre04, Hua07, KS08a, Kwo08, LKE04, LLZ07, LF05, LKMU05, LZH^{+}06, LZH^{+}07, NZ07, VC03, VGZB09, VGBZ09, XW06, Xu08, YFBH07]. Simulation [ART02, ACK02, BM02, BST01, BHR03, BADG00, BM01a, BS01, BRL02, Bow01, Buc05, BMK^{+}06, Bus00, CS01c, CGL08, ČPT01, CPK02, CYKC01, Cle00, DNS08, DF00a, DQA08, DGH02, DDGS09, EH02, FS01, FG02, GPH^{+}01, GMAj09, HAAO00, Han01, HKKS^{+}01, HK02, HF01, HB02, HSS07, HGM01, JLCD01, JWSC00, JW02, KB00, Kar04, KW08b, KAIN01, KP00, LBD02, Li01, hLA01, LP02, Mac01, MEG02, MSB07a, Mu02, NCS03, OL01, PG02a, Par02, PR01b, PFB01, PWS^{+}02, PO01, QRHD00, RRL01, Rom02, SLY02, ST01, SSL00, SCW^{+}09, SPC01, SB02, TSB01, TC01a, TCM^{+}00, Vay01, VDM^{+}02, Ver01, WPM^{+}02b, WB01, Xu02a, ZKS^{+}09, ZTZ02, ZP02, ZF02, ZTPM05, ZKK01, AH08, AR08, AMH04, ART04, AMP09, AT09, BS04a, BPMR08, BCK09, BP06, BWLM09, BA03, BS03b, BPL06]. simulation [BEA09, BGN03, BP07, Bur05, BB09c, CPR05, CGP05, CP06a, CBJdlC07, CPG04, CGRGV^{+}04, CTW^{+}08, CFL^{+}03, CN08, CP07, CLL^{+}07b, CMP07, CF06b, CSKD05, CL03b, CSML06, CB09, DSJ03, DL03a, DDH05, DS09a, DTS05a, DTS05b, DS09b, DCK08, ELVE07, EE08, Eld07, FP08a, FG04, FG05, FDD09a, FDD09b, FT06, FD03, FD09b, FLM08, FY07, FKK08, GMD03, GS09a, GGS09, GLS03, GIA^{+}07, GBB^{+}06, Gra06a, Gra06b, HBLD07, HKM07, HJKO08, Hew03, HHMK05, HK08c, HH07c, HF08a, Hor06, HM05, HSW07, HL07c, HSL08, HT03, HPS^{+}06b, HMR08, HHM04, HLWW04, HLWW06, ICO04, JRS05, JLT03, JD04, JOS06, JW03, JMZ04, JHB^{+}09, JS05, KHdT^{+}08, KFIG06, KG06, KM06, KM07b, KFV^{+}05, KDC05, KKS07, LJM^{+}06, Lar03, LMV04, Lau06, LW06, LL05, LCH03, LLL07, LS08, LK09, LKW05, LP06b, LL06b]. simulation [LCNR07, LMH07, LWF^{+}08, LDV08, LH05b, LQ06, MC04, MCM04, MLM09, MTWW06, MC06a, MJT06, Men04, MGS07, MR04, MHE06, MWG^{+}06, MK04b, MHdB07, MGNB09, NLF03, NJLA06, NFvS^{+}06, NC04, NB04, OK06b, PSCB08, PDHP07, PYC04, PM07, PL09b, PN03, PH06, Pet07, PWM06, PA05, PVPS09, Pro05, RB05, RRC05, RGS04, RMG^{+}09, Rom07, RJM07, RFFP06, Ros09, Roy05, SM09a, SWB^{+}06, SWG08, SW08a, Sch08, SMS08, SHWC07, SP04, SL04, SP05a, SFX03, Shi07, SMP09, SSND03, SGG^{+}04, SMSS07, SK07b, TOZP03, TZ06, TZ07a, TB06, TSB03, TdAAP08, TPR05, Tsy03, Uhl05, UPKN09, VTC^{+}07, VS09, VGL^{+}07, VK05b, WT07a, WLC^{+}08, WFC09, WWK05, WMH07, XLP05, XG09, YM07, YB06, YWC07, YXLF05, YSS05, YF09, ZP05, ZSB^{+}08, ZT03, ZW03, ZD08, dSMN^{+}04, dCNHSD07, vdBG09]. Simulation-Tabulation [HGM01]. Simulations [ATV01, ALGM01, CS01b, CVB00, CTT08, CBL01, DW00, DKSW01, DE02, DPR00, FVOMY00, FPC^{+}00, FLG01, Gen01, GLL03, HAP06, HPZ01, JML^{+}01, KS02b, KK00a, KKC01, KK00c, Lap02, LS02c, LL06a, LLQ^{+}02, ME09, PPC00, PW00b, PW01, RXH02, SSW01, Sun00, TMSW07, WGCE01, Yon01, dSAK00, deM02, AS05a, AZB09, AD03, AFGM07, ALT08, AD04, AGSX09, BLW04, BDR^{+}04, BDGL05, BMN07, BBB08, BL09c, BDS07, BS09b, BTW03, CGMS06, CGN^{+}07, CV06, CP03c, CELS07, CM03, CFGK05, CWD08, CK07, CHPR09, CP04c, CH08, DUEB07, DW09, DKS^{+}03, DLD08, DZ09a, DJ04, EGHE06, Eld08a, ÉGP09, FHJK09, FPK08, FM05, FHD^{+}09, FE04, FCGK05, GCGE03, GGF03, GGRS08, GCLB04, GS09d, HGBH03, HC08, HO06, HP04b, HM04, HS04, HLX06, HS07b, ID04, IK07, ISNY05, KM08a]. simulations [KMV03, KFV07, KH07, KSJ03, KZ06, LDN04, LMN^{+}09, LWDA09, Li08b, LJ09a, Liu09c, LR03, LMK09, LL08b, MWM08, MG05a, MY06b, MP05, MKL06, MLFG06, MVO04, Mot08, MBP07, MMPB07, MO06, MAL09, NTO^{+}07, NDG05, NJX08b, NLLE06, NS04, OLA08, ODAF06, OK07b, Pau07, QLK07, QLS09, RVDM09, RB09a, SMT^{+}08, SG06, SGFL09, SE09, SK08a, STD^{+}05, SA06, SFVK06, SL07c, SHP07, SSW^{+}07, Spe05, SFMP06, TWM07, TBJ^{+}09, TMD07, TGB^{+}07, TG06, TDV06, TG04, VPMC04, VGS04, VK09, VCM00, VQLZ04, WL03, WTL08, WC08, WH05, XK03, XH03, YFLS06, YZLH09, YYT05, YZF^{+}06, YZF07, ZGG03, ZVQ07, ZLAC05, dSHHM05, vZdBB07, vdV08]. Simulator [GW02, GC06]. Simultaneous [AKV06, DVHM05, HSBG05]. Simultaneously [DSS00]. Sinc [Eg07]. Sine [Mil05, Saf02, BRB03]. Sine-fit [Mil05]. Single [JK02, JD04, LW06, NMG09, PL09a, RSW06, RM07]. Single-Crystal [JK02]. single-domain [LW06]. single-fluid [RSW06]. single-fluid-phase [RM07]. single-phase [NMG09]. Singular [ACS00, AQV02, APQ02, CH01, LL01b, LTZ02, RW00, WPW02, ZS01, ACLS03, Boy06, DG09, HL07b, LH05b, Sac07, SY09b, TE04, WZ07, ZZFW06, dA04, dCNHSD07]. singular-regular [dA04]. Singularities [Mai01, MC00b, OKL01, Boy05a, Gro06, Gro07, HO08a, VRM07, YW07]. Singularity [Nit01, CSO09, LL06a, TPVG06]. singularity-avoiding [CSO09]. singularly [LCW04, Moo07]. Sinks [WLE^{+}00]. sintering [CP05]. situ [LP09]. Sivashinsky [CFP08, KMS03]. Sixth [WZ09, Hau08b]. Size [BW02, Zha02, Hew03, YE07, ZZ09]. Skew [Coe02, DLS^{+}00]. Skew-Symmetric-like [DLS^{+}00]. skewed [TAL09, YMWM06]. skill [Ano08-50, SM09b]. Slab [BS00b]. Slater [GM01c]. Slender [KK00c, BP08]. SLICE [ZWS07]. slide [Car09]. Slider [WB01]. sliding [AKH06, HHM04, KH07]. slightly [ZD05]. slip [BIW04, HSC09, PK05, SS05c, SN08, VLB09, ZTPM05]. slip-dependent [BIW04]. slit [Mad05]. Slope [Xu02a, Boy03, ML08]. Slope-Update [Xu02a]. Sloshing [Fra04, CN05, LL08a, VGL^{+}07]. slow [GV06, Kel05]. Slowly [Wu01]. Smagorinsky [MGS07]. Small [Hix00, LWDA09, AV03, BEPT09, HMA05, KM06, KM07b, Pal08, RE07, RWMK03, TJ09]. Small-angle [LWDA09]. small-scale [KM06, KM07b, RWMK03]. Small-Stencil [Hix00]. Smearing [BU02]. Smooth [ASPB03, Ber04, CS06, CS07d, CS08c, DJM05, GP05, MC06b, YBZ06, vZdBB07]. Smoothed [BZ08, CPK02, iI02, LMK09, PM02, ZF02, BZ04, BOT05, CGSR08, CFL^{+}03, CL03b, CEL06, ESE07, HK08b, HX05, MHW05, TM05, TMSW07, Yok07, ZB07]. smoothed-particle [BZ04]. smoother [EKP07]. Smoothing [DIV00, KKGL01, ABHT03, HZGB04, HZGB05, WSTW09, YZLH09]. Smoothness [KKP02, LCS02]. Soap [ZP02]. Sobolev [RSS09, SNLS03]. soft [HK08c]. soil [SM06b]. Solar [SJC07, Dic08]. Solid [Bar02a, EKK02, FGS09, Fed02, HHL00, HPZ01, Man02, MC02, BL08, CCV03, CYS06, CP04b, DVHM05, HS04, HH06, JJGL06, JJGL07, KS09, Kou07, KB06, MEKS03, MMS04, Mil04, NMG09, SMGJ09, Vik03, VHI05, Yam05, ZKS^{+}09, ZFM08, vLAvdV06]. Solid-Fluid [MC02, FGS09, BL08, ZKS^{+}09]. Solid-Liquid [EKK02]. solid-rigid [vLAvdV06]. solid-state [CP04b]. Solidification [ART02, LLH02, LW06, MHW05, PK00, SZ01, YU05a, ZH01, AD03, ART04, KW08b, RJM07, TZ06, TZ07a, TZ07b, TZ07c, ZGT06]. Solidifying [LS02b, CDDH07]. Solids [ATV01, HB02, MC01, JG09, Tan08]. Solitary [Boy02b, LY07b, Yan08, Yan09]. Soliton [GHV00]. soluble [MT08, ZEA06]. solute [IG05]. Solution [ABGV02, ACS00, ADK00, ADK02, AQV02, BR01, BC02a, BMR01, BMRS01, BMRS02, BT02, BK01, CWT00, CL00b, Dri02, FGG01, FP02, GZ01, GC02b, HW02, Hua01a, JMK01, Kan02, LTE07, LXM09, Lin01, MR00, MR02, MN02, Pai01, PR01a, QCGQ03, Stu01, VB00, WPW02, WLE^{+}00, ZZ01, ZYC02, dFJS09, vBRK01, ARRS09, AKV06, AL06, AEP04, And09, AG09, ACLS03, BT03, BJM03, BL09a, BCDR06, BFC04a, BFG08, BTT08, Bia03, BSLN09, BSW05, BGLN05, CJK^{+}03, CFJ06, DSM09a, DVHM05, DS06a, DF04, Dom08, DR09b, EB06, FK07a, FT05, FOLD05, FH07, GS03a, GS06b, GV07, HKM08, HEN09, Hoh06, IHL03, JW09, KS08a, KK03a, KSH^{+}06, KAK03, KL04, KMS04, KT03, KT05, LH05a, LG07, LM08c, Low05, LCdCN^{+}03, MZ09, MP07a, MNR07, MKKY06, MK08b, MOG09]. solution [MSB07b, Mou04, MK07, Ols07, Pap08, Pud06, RVM07, RWMK03, RHPN09, RAD07, Roy05, SSN09, SS03a, SWZ03, SB09, SHTB09, Soc03, ST03b, Thö04, Tok06a, TV08, TDV06, VC03, WK07, WM07, YAvdB^{+}08, YKG04, ZG08, ZYL^{+}06, ZZ09, ZSC08]. solution-adaptive [ST03b]. Solution-limited [LXM09]. Solutions [CH01, Gos02, PKP01, PS01, SPW^{+}00, VQSZ02, BB08a, BDRT09, Ber04, BT09, BDCG03, CFR09, CL03a, CS08b, DD09, FCJ08b, HK06, HLRZ06, HL07b, IG05, KW08a, KMS03, KN04, KLSW09, LHD05, LW06, LRMB08, LW07, MM07, MHB08, MD06, RS09b, SY09b, TC09a, UL06, YU05a, YJF^{+}06]. Solvated [FS00b]. solvation [CXB08]. solve [BZ09, Cha07a, Cha07b, CEH09, CFP06, IM07, ND04, TBT^{+}09]. solvent [DC07]. solvents [XDC09]. Solver [BZW01, CBKM00b, CKF02, CSS00, HR01, iI02, Lai02, MOvL00, RLB02, ZS01, AQ09, ABZ^{+}08, BCCV09, BH09, Bey09, Bia03, Bil05, BYZ04, BJ09, BLM04, BL05, BH04, CW03, CGMS03, CK03, CP06b, CP06c, CHL06b, CLL^{+}07b, CLS09b, DHOT09, DSM09a, DVHM05, DBF08, DP07, GS05a, GS06a, GA09, GTMC08, Gui05, HK08a, HC09, Her08, HdGKG08, HJ07, HVAC09, HAI09, HB05b, IQT08, JR09, JTL09, KW06, KP04, KLLJ09, KAA^{+}07, KS07, LT05, LFX05, Li05, LK09, LDPL08, LC03, MR05, MR07a, MPD08, Mig07, Mil04, MK05, MBP07, NOG08a, NI03, NLLE06, NS05, Pop03, Pop09, RBT03, SO08, SL03, XYK05, YBZ06, ZJ09, dNWvSD07, dTWD09]. Solvers [AV02, CT08a, LP01, Mav02, MOS^{+}00, OMG02, QS01, SBGK00, SC01, AMLC08, APQ03, AQ07, BPO07, Bra08, BH05, DIL03, Geo08, JR07, JL04b, KDK^{+}07, KSW07, LMS05, LG07, Löh04, MTV08, NK08, SMAj08, SB03, TT06a, YLD09]. Solving [AKV00, Bot06, Cal02, CCV03, DV02, DKX00, DKX01, DFT01, HK01, HMM07, Mai03, MP01a, MP02, MSO04, MR07c, Moo07, ORM06, SS08, SKAS01, SWL00, TK02, UL06, VC00, WKL07, YA05, ZYC02, ZYL^{+}06, AMXL09, AC05, CHL09, Cap08a, Cap08b, COQ06, CS07b, CW08, CJ07, CDL04, CDL05, DLP08, Eg07, FM04, FP08b, GS07, GK05, HL05, IKS^{+}09, ILL09, KSO^{+}05, KZ04, LL03c, LRZ04, LP07b, Mai04, MR06a, OTCM08, PSD09, RCT07, RLZ03, RS06a, RRW05, RM08, SROCFF03, SL07b, SR09b, STR07b, TLL^{+}08, Tau07, WXG07, WGS^{+}08, XDB09, XMT06, YYF09, YCL05, Zho07, dDEK09]. Some [JHSZ07, OF01, Sto07, Thu08b, AST07, LM08b, VBL03]. Somerville [WS04]. Sommerfeld [GFR09, Meh04]. sonic [Asl04b, Tan05b]. Soroban [YMT^{+}04]. Sort [Bow01]. Sound [Fre00, HSK00, MN02, BA03]. Sounding [TK02]. Source [HK00, HGN00, SR00b, Xu02b, ZCMI01, BIW08, BP03, CVB06, CSMH05, ES03b, GLS03, HW08, JD09, KNH05, LT09a, MJ09a, MC07c, RBT03, TE04, TT05b, VS07, Wen06, XS06, YYF09, ZSW03]. source-independent [CSMH05]. source/observer [VS07]. Sources [GBGM01, POS00, WLE^{+}00, BCDW06, CWJ07, HO08b, OK06c, ZZFW06]. Space [AB07, CP00, CWT00, Han00, HA02, JWSC00, KvdVvdV06b, LTZ02, OCK^{+}02, PR04b, PM00, SUW01, ZYC02, ZRR00, vdVvdV02, vdVX07, AKV06, ASQR06, AK07, AGSX09, BFT07, BIS07, BS03a, BKM09, Boy03, CQRW05, CFJ06, FR03, FCGK05, GvH06, GR07, HLO08, IS04, IH04, JW03, JX07, KvdVvdV06a, KvRvdVvdV07, KT05, KLM05, LG07, LW07, LS09, kM07a, Moo03, Moo07, MR06b, RJM07, RBK09, Shy06, TT05b, TFDK04, ZYL^{+}06, dWKL07]. Space-Charge [SUW01]. spacecraft [LZ09c, VGL^{+}07]. spaced [HM08]. Spaces [FLG01, YS06]. spacewise [YYF09]. spacewise-dependent [YYF09]. SPAM [SWTM01]. Sparse [Bor00, GZ07b, GG00, WHS08, ABZ^{+}08, BPS03, DBF08, HM08, LAKD08, MZ09, ZG08, ZGSD06]. Sparsity [Lou00]. Spatial [BRL02, BCMO01, KK05c, LBV00, MAN^{+}06, OMG02, ZDNP00, Bey09, BB07b, BdCB09, Che07, CFP06, Jor07, LSD07, MPFC08, NWZL08, RBSL06, ZT07a]. Spatially [CKS00, AKLMP09, CV06, FG04, FG05, FE04, GTMC08, KSJ03, Kou09, LG09, VCZS04, ZIP06]. spatially-developing [FE04]. spatially-varying [GTMC08, Kou09]. spatio [DGF09]. spatio-temporal [DGF09]. spatiotemporal [SJC07]. Species [WDM01, AK09, BK07, LCB04, SD05a]. spectra [DK06, Mil06, Mil07, TPVG06, Yan08]. Spectral [AQ00, AGP01, AQV02, AP02, AQ07, BJM02, BK08, BS03b, Bor07, CSS00, CKGL02, CMOV02, DLMK04, Dri02, ES06, FLG01, FYH^{+}06, GT09a, GP00a, GS06b, GBGM01, HL01, Hei04, HKV01, KK00a, KK00b, KB08, LS02c, LJW09, LVW06a, LVW06b, LJW07, LMS02, PRT00, PX02, PR04a, PW00a, PR03, PR06, PG02b, RH01a, Sac07, SB06c, SC01, SG03a, SWL06, TRL01, Wan02, WL02, WZL04, WK01b, vOP04, AQ09, BM06, Bey09, BS04c, BvdHKG07, BDCG03, BLM03, Boy04, Boy05a, CJSS08, CCV03, CLL^{+}07b, CKG04, CQRW05, CD07, CFJ09, DDH05, DLD08, Dim07, DGJ03, DD03a, DD03b, FD03, FK09b, FBHV05, Fou06, GSV06, GFR09, GPF03, GW05, GR08, GD05, HWL08, pHL09, HK08a, HEN09, HdGKG08, HL07b, HJM^{+}05, HJM06, IM05, IHL03, JZ08, JW09]. spectral [KCH06, KH09, KDK^{+}07, KS09, KT03, KS07, Lab09, LL09, LP07a, LM04, LJS08, LX07a, LCCG05, MVD04, MNR07, MP03, MLFG06, ND04, NL09, NB04, PH08, Pir06, PR04b, Pon07a, Pon07b, RFFP06, SSN09, SR09a, Sar03, SS07a, SY09a, SK04a, Str07a, THL06, TBT^{+}09, TCN09, THS07, VL07, VLW07, VBL07, VGPL09, WL06, WG09, WPH00, WGRA09, XP04b, XLS09b, YYT05, YZW05, ZC09, ZYHS07]. spectral-element [Fou06, KS07, RFFP06]. Spectral-Lagrangian [GT09a]. Spectral-Projection [LMS02]. spectral-WENO [CD07]. spectral/B [DD03a, DD03b]. spectral/B-spline [DD03a, DD03b]. spectral/discontinuous [CQRW05]. Spectral/hp [ES06, PR03, PR06, SC01, PR04b]. spectrally [BW06, BCL06, DP09, HF08b]. spectrally-accurate [HF08b]. Spectrum [GBS00, VCT09]. Speed [FS01, KMA^{+}01, sLwG08, BHS09, BN09, HS03a, HS06, KSO^{+}05, SMS04]. Speeding [HK00]. Speeds [MD01, JW06, Lio06, MDS03, Soc03]. speedup [EV03]. SPH [BRP05, DKS^{+}03, DLD^{+}06, DLT09, GAC^{+}09, HA06, HA07, HA09, JOS06, KM08a, LMX^{+}08, Mon00, ODAF06, ODAF07, OBT06, Owe04, Pri08, SSL00, XSL09]. Sphere [Che00a, Che00b, CKGL02, KMHR00, Lay02, PWS^{+}02, Tol02a, Tol02b, AQ09, BAFL09, BCE^{+}09, Boy05a, BZ09, CDJ07, CF06a, CCV03, CX08, CKG04, Cho05, DC07, DTSC04, DJ04, FW07, FP08b, GW05, Gir06, LS03, LJW07, MC06b, MK08b, NFvS^{+}06, NI03, Pud06, PL07, Ros06]. Spheres [PO01, BP08, VQLZ04]. Spherical [Gir00, GHW02, LBV00, LBV01, Nit01, SS00, AQ07, GG09a, GPF03, Jao07, JD04, KL06, LHD05, LGHD08, Liv07, Mac07, NB04, OBT06, SP06a, SC08b, Tyg08, WJV07]. Spherically [HZ02]. Spheroidal [BS00e, Boy04, SJ04]. Spilling [DF00a]. Spin [GCW07, YMF01, FT09, WJV07]. Spin-Orbit [YMF01]. Spin-polarized [GCW07]. Spline [DDS09, GW01, KMJ01, Lay02, PB00, SKAS01, Ver01, Bia03, CP04b, CLS09b, DD03a, DD03b, ELW04, LHGF04, ZWS07]. splines [CP04a]. Split [HZ08, SFY01, Sti02, AMSZ07, MM03, Nas08, RBK09, SA09, SK08b]. split-conservative [SA09]. Split-Step [SFY01, Nas08]. Splitting [BJM02, BM01c, Edw00, EF02, FMR09, HH01, HH02b, HGN00, KLN^{+}01, LLIK01a, LLIK01b, MBP00, NTYT01, NTYT02, Ros00, SLY02, Spo00, VG01, YVD00, Asl04b, BG05a, CHL09, CJ09, DPRN05, DQ04, EF03, Fas03, FL09, GS03d, GS09c, HJM^{+}05, Hub07, HGB^{+}03, KN09, KKO04, LLOT06, MEKS03, MY07, NvL03, Pon07a, QW05, QA09, RC09a, RP08a, RS05, RS09a, RDPN07, Sac07, SJD05, ST03a, TBT^{+}09, TCN09, TK04, YHSX07, YZW05]. splitting-based [TBT^{+}09]. sponge [Bod06]. Spontaneous [YZF07]. Spray [BW02, FLM08]. sprays [AJ09, LMV04, TT06c]. Spread [BST01, BST03]. Spreading [HLZ02, DW09, HSC09, ZGG03]. Spring [TTSG01, TSG02, LWF^{+}08]. Spurious [DS01]. Square [Bor00, Cap09, GS03a, LL04a, SL07b]. squared [BEE06]. Squares [PG02b, AMSZ07, BT05, BT06, BP04a, CSO09, DI09, GNNB08, HV03, HK08a, HMMR04, HLMM07, HdGKG08, HY09, HY11, KH09, NCS03, PR03, PR04b, Pon06, Pon07a, Pon07b, PR06, VB09, ZKY05]. squares-based [NCS03]. stabilisation [NW07]. Stability [AC00, ACY00, APQ02, Bal08, BB08a, BFJ03, Cod01, DVHM05, DWLM09, FGG01, FDL08, GF05b, HFO01, LG08, Lee03, LRS09, MG02, NR01, PR01a, Pet01, Rem00, Rem06, RS05, RS09a, RB02, SHWW00, SV00, WK01b, BBC^{+}06, BDCG03, CHH06, CHPR09, CFJ09, DMG04, FD03, GV08, HS09b, HM04, KRT^{+}09, KWD07, LH05a, LGKP07, LZ07, MPD08, Maz06, NZ07, OCFF08, Sam09, SCT09, zSW06, zS06, VCT07, VL07, VGPL09, XSL09, Yan08, YMWM06, ZK06, ZT03]. Stabilization [PX02, San03, HH07b, QT08]. Stabilized [JML^{+}01, RB09a, XP04b, BB07a, HY09, HY11, LSS^{+}09, MZ08, MR06a, MGCR07, NZ05, SV07, ZSP08]. Stable [Azm02, BKST09, CYKC01, De 04, GGCC09, Hu01, HWWL09, IA06b, KR02, KR09c, MSS08, MHI08, Nys02, VWW04, Wan04a, WDM01, Wee02, YC09b, ZZ01, AM03, AM04, AB03, BL04, BLM08, BO09, BSLN09, DR06, FNS07, GG09a, GN07, HR08, pHL09, HX05, JAK05, JL04b, sKKRH03, KYK07, KPP07, KPP09, LL05, NFGK07, NG06b, NGvdWS09, PC06a, SLV09, SCN07, SN08, TCM05, WC07, YC09a]. stacks [CKPW07]. stage [KWD07]. stage-exceeding-order [KWD07]. Staggered [ALGM01, GHV00, GW01, HH01, Per00, XCZ02, YP01, ZSP02, Boe05, CSL08, KE06, KAK03, LD09a, LS05b, LSW08, LRS09, PN03, PS04, RCB05, SK06, VSW04, VSW06]. Staggered-Grid [XCZ02, LRS09]. stagnation [SBA07]. Standardized [BP04b]. Standing [VCP00]. Stars [BTFY01, TVMR03]. started [DCK08, KR09a, NCS03]. State [CYKC01, FV01, Mai01, MOvL00, Shy01, VTM^{+}08, BT03, BLM04, CORT09, CGH05, CC07, CP04b, CS06, CS07d, CY05, HPS06a, HJJ09, ISNY05, Kwo08, Mai03, Mai04, MK05, MK07, PPB09, SFDL07, VVS08, VTW^{+}07, Wen06, ZL08b]. State-of-the-art [VTM^{+}08]. Statement [KB01]. States [FGOV00, RS02, BCL06, BS08a, BNNP06, CL08a, CGL06, CKLS05, CLS05, FJ09, GT09a, Geo08, VTW^{+}07, Vos06]. static [CLLG09, FHD^{+}09, HH07a, KK03a, KOQ04, KOQ08, LKE04, Mar06, PH09, VOD08]. Stationary [DCV^{+}01, GS06a]. Statistical [DC08, DF07, FS01, HGBH03, SFX03, CFM09, FWP09, GT09a, PM07]. Statistics [FH02]. Steady [CYKC01, CAL00, GM01b, MOvL00, PPB09, VG01, VP09b, vBRK01, AM04, BNNP06, BLM04, BEG03, CGH05, CC07, CS06, CS07d, CY05, DR09a, FJ09, Geo08, Gla05, GMO04, GMS06, HY09, HY11, HLY09, Hub07, ISNY05, NJX08b, PR06, RFVP09, VVS08, WK07, Wen06, XMP07, XSG04, ZKY05, ZL08b]. Steady-State [CYKC01, CGH05, CY05, ISNY05, VVS08, ZL08b]. steady-states [FJ09]. steep [YM07]. steepest [CSMH05]. Stefan [BMR01, BFC04a, CMG09, GF05a, HZ07b, JVVS07]. Stellar [MC03]. stellarators [SKK^{+}08]. Stencil [Hix00, DS06a, UPKN09]. Step [BE02, CGP02, Cod01, Gui02, LOK01, MT01, SFY01, AHNS09, BBHM09, Bil05, CB03, DT04, Dom08, DBTM08, GSV06, HVAC09, HLX06, ISNY05, KDW08, KSGF09, LDW07, LP07b, Nas08, RFVP09, VSG05, ZSW07]. step-flow [HLX06]. steplength [FG04]. Stepping [Hig02, RB02, ZTZ02, BHvdV06, DR06, DGRS08, HR08, HC08, Hig05, KvdVvdV06a, LXM09, LGM08, Mad06, MPFC08, TDGP06, UBRT07, VSH04, YA05]. steps [TZHT04]. sticky [PC08]. Stiff [BJ00, BJ02, CM02, MOvL00, Spo00, VG01, APR09, CPKW09, DP08, DET08, GV06, KKO04, Kro05, MC07c, NZZ06, RE07, Tok06b]. stiffness [HS08c]. Stirred [LLQ^{+}02]. STM [BGR08]. Stochastic [AGT02, AJ09, BDGL05, DGA08, GK02, Hor02, Kou07, LKNG01, LRN^{+}02, MNR07, PFB01, TX06, WB09b, AS05a, AA07, AA09, APR09, AGT05, AWK07, AGSX09, AZ05, AZ06, AKP07, BCK09, BFG07, Cam03, CP03a, CGP05, CP06a, DSJ03, Dem04, ELVE07, EPW08, GZ07a, GZ07b, GZ08, GZ09, GD06a, JZ08, KMV03, KW03, LGP09, LK07, LRS07, LGK06, Liu08, MZ08, MZ09, MT04, NZ05, NL09, PA07a, RE07, Sto07, SL06, TJLT08, WK05, XS09, ZG08]. stochastic- [APR09]. stoichiometric [JVVS07]. Stokes [DD03a, AQ09, BQQ09, BCDR06, BS08b, BHR06, BB07a, BACFT05, BLM08, BT02, BCVK02, BYZ04, Boe05, BT06, BJ09, BT09, BP08, BCM01, BGLN05, CSL08, DC01, DR09a, Dim07, DD09, DB04, Dom08, DD03b, EHST03, EHS^{+}08, FL03, FOLD05, FD07, GH03, GS07, Gel06, GP00a, GSV06, GCNB07, GR08, GS03c, Gri09, GSW00, GK05, HH08, HH01, HDC02, HK08a, HH02b, HLMM07, HS08b, HS08c, HLL08, HC05, IQ08, ILL09, JK00, JL04b, JMC03, KA05, KE06, KH09, KDK^{+}07, KG08, KAK03, KvdVvdV06a, KvdVvdV06b, KvRvdVvdV07, KM00, KB01, KS09, KT03, LMN^{+}09, LOK01, Lee09, LC01, LL04a, LL01b, LFX05, LDPL08, LRZ04, LP07b, Liu09b, LM03b, LMS02, LB04, MPP01, MVD04, MRS09, MM01, MCG08, MSS08, MF00, MG06, MLS01, NW07, NZ05, Ni09, Nik06, NMS07]. Stokes [NGvdWS09, Pai01, PNMK09, PKP01, Pet01, PR03, PR04b, Pon07a, Pon07b, Poz01a, PR06, PG02b, RBH03, RS06a, SML02, SNGAS04, SFE07, SMB09, STZ07, Soc03, SCN07, SN08, STR07b, SPW^{+}00, TOZP03, TLK09, TXCD07, TC09a, TS04, TG08, TWS02, VSW04, WRu03, WPH00, WK01b, WS01, XK01, Xu01c, XYK05, YS07a, YJF^{+}06, ZL08b, ZDNP00, dVGLM09, vBRK01]. Stokes/MHD [YS07a]. Stokeslets [ADE^{+}08]. stopped [Buc05]. storage [CFR04]. strained [RS06b]. strains [SKS08, YH07b]. Strategies [KB00, KLN^{+}01, KMS02, BB07a, HJFW04, HM08, RMG^{+}09]. strategy [ÁDIM09, BP09, CL06b, ERVE09, FVE04, GCCD07, GE07, HE07, MMS04, MHE06, VBL03, VBL04, YYT05, dDEK09]. Stratified [CL07a, CL08b, DKSW01, Pai01, Bar04, BM06, DDH05, GBC06, KSH^{+}06, KSHS08, SE04]. stratosphere [MM09]. Stream [HF00, Pee03, LL04a]. streamer [CN08]. streamers [MHE06]. Streamfunction [Cal02, AKH06, BACFT05, GK05]. streamfunction-velocity [GK05]. Streamfunction-Vorticity [Cal02, AKH06]. streaming [GS06b, LL03b]. Streamline [HFO01, MJT06, Yeh07]. streamwise [JOS06]. Strength [RH01b, XDC09]. Stress [HJ02, BS09b, SW04a, YAvdB^{+}08, GM01b]. stress-velocity [YAvdB^{+}08]. Stressed [ATV01, LN09, RRV06]. stresses [HO03]. stretch [ID04]. stretched [dHRvdB07]. Stretching [ACGV07, GGRS08]. Stretching-based [ACGV07]. Strict [AC00, ACY00]. strides [SROCFF03]. string [CP03b]. strip [ST04]. Strong [GG00, Sun00, DDSV09, DLT09, HP04b, KLLJ09, KWD07, LP06a, LKY03, dWKL07]. strong-field [dWKL07]. strong-stability-preserving [KWD07]. Strongly [AK06a, Alb00, LG03b, KKS05, SE04, WKL07, YE07]. Structural [AJT04, SW00, LLC^{+}08, LZH^{+}07]. Structure [BADG00, CWWZ00, CD00, Cul01, DGP00, TR02a, AK06a, AKP07, BNV08, BCHL07, BGS08, GA09, GGCC09, HBHS09, HC09, IS04, KKCF09, KGJ05, KYK07, KF06, Küm04b, LZ07, LCG07, LMZ^{+}08, LZH^{+}07, MC03, MK04b, Pap08, SPT05, SL07a, SM06b, SP05c, SP06a, VTM^{+}08, WBM09, YMW06, ZFM08, vLAvdV06, vZdBB07]. Structured [DLS^{+}00, DI02, Edw00, LM01, BHS09, CSML06, LSB04, LM03a, MC07b, MCP03, RAB07, TRSK09]. structured-grid [CSML06]. Structured/Unstructured [LM01, LM03a, MCP03]. Structures [DS01, GM01a, KM02, WPM^{+}02b, WZ02, CLL07a, CGL06, CdHST08, EZ08b, LHGF05, MR07a, RVVL09, SZB^{+}07, SZS03, ZK04, ZH09]. Studied [vHBB02]. Studies [OS04, RSO04, CGSR08, LZ09c, LGK06]. Study [Dar00a, DCV^{+}01, DP00, LZ04, Lin02, PPCW06, SZ01, BBB08, CLL07a, CKG04, CP05, CM03, DS06b, DLW04, GR08, GLLX08, HMR08, IA06a, Kas07, KTD03, LCB04, LL08a, MC09, MC06b, NFA03, OCFF08, QKS06, Ren07, SS03a, SJ04, SB06c, SKW05, SCRL08, TPVG06, VL07, ZK05]. Studying [PA00, Kro01]. sub [BAMD07, BAR08, CLTA07, VS07]. sub-cell [BAMD07]. sub-diffusion [CLTA07]. sub-linear [BAR08]. sub-wavelength [VS07]. Subband [PA05]. subcell [LS05b]. Subdomain [WPW02, KT05]. subdomains [KT03]. subfilter [LDN04]. Subgrid [AS02, ML01b, AHF04, PM08, VK09, Yeh07]. Subgrid-Scale [AS02, AHF04]. subgridding [VPMC04]. subject [SG03a, VVS08]. subjected [JOS06]. Submarine [DC02, FNBB^{+}08]. Submodels [BW02]. Subsonic [SSD00, SB02, HSQ03, Pro05, Pro07, SBC04]. Subspace [SWTM01, BEPT09, ZSTC06]. substances [NDG05]. substrate [ZDD09]. Substructured [SC01]. Substructuring [Man02]. subsurface [JLT03]. Suction [CS00]. Sufficient [SV00]. sufficiently [GP05]. Suitability [MLM09, PYC04]. Suitable [Tem06, KPP07, KE09, RB06]. Summation [MN04, MN17, KTD03]. sums [LT09b]. Super [CR02, AC09, CLL07a, Sar03]. Super-Grid-Scale [CR02, AC09]. super-lattices [CLL07a]. Supercell [FHLK05]. supercells [LCG07]. superconducting [DJ04]. Superconductivity [DDG02]. superconformal [SS08]. Superconvergence [CS08b]. superconvergent [LNXNTX09]. superfluid [LKE04]. superlattices [CBC09]. superlinearly [VSW04, VSW06]. supersonic [BP04a, DGJ03, FL07, MT07a, UPKN09]. Support [MHS01, EZ08a, GG09b]. Support-Operators [MHS01]. supported [Tow07]. suppressed [Bor03]. suppression [Lur07]. supra [MGC06]. supra-convergent [MGC06]. supralinear [CMG09]. supsersonic [BP04b]. Surface [AINR03, BST01, BK01, CS00, DF00a, GHG01, GKJW07, JTB02, JK02, Nie01, RRV06, RRV01, RR02, SZ01, Str01b, TCM^{+}00, ZTZ02, ZCMI01, vBRK01, AMH04, AMS04, BMN05, Boy05a, BN09, BGN03, Bur05, CPR05, CS05, Che03, CFGK05, DS05a, EE08, EG08, FCD^{+}06, FCGK05, GV08, GT09b, GFR09, GS03b, GCCD07, GAC^{+}09, GR07, GS09d, HS09a, HZ07a, Hum05, JCT07, Kim05, KLW09, LLP07, LB03a, LN09, LY04, LF04, MS08a, Ni09, NGC^{+}07, OTCM08, Pee03, PN03, Pop09, QP03, RB05, RMF08, SDGX07, SE04, SK05, SAKDJ05, SC08b, TJ09, TW07, WSI08, XMP07, YZ07, YP06, ZKDT07, ZL09]. Surface-Tension-Driven [Str01b, Pop09]. Surfaces [BCOS01, CBMO02, CBL01, KKGL01, ML01a, MS01, BPL06, BHP07, Cec05, CDI09, Chr04, CH08, GH08a, GNNB08, Hel05, JR03, JR04, JCT07, LZ09a, LKMU05, MR07b, Nic09, RGS04, RM08, dSMN^{+}04]. Surfactant [GHG01, JL04a, LTH08, XLLZ06, ZEA06]. surfactant-conserving [JL04a]. surfactants [GT09b, MT08]. surrounded [CPR05]. Survey [Ben02, KK04]. suspended [KHdT^{+}08, KMSH08, VGZB09, VGBZ09]. suspension [AH08, FY07]. Suspensions [JC02, DMHP07, HO06, HHM04, TG06]. sustained [CBC09]. sweeping [COQ06, FLZ09, KOQ04, KOQ08, LSZZ08]. swimmers [HK08b]. swimming [KM08a]. Switch [KMA^{+}01]. Switch-On [KMA^{+}01]. switching [CBH03]. Symbolic [BMDS05, BSP06, CS03, DP00, CS04, MBS03]. Symmetric [DLS^{+}00, GFCK02, HZ02, Mit00, Ver01, BS08a, BPS03, CJR04, JLOT05a, KLSW09, LLS09]. Symmetries [WZ02, KEB^{+}07]. Symmetrized [DS06b]. Symmetry [BPMR08, CRB00, Car01, DH09, VV03, Kok09, LW04]. Symmetry-preserving [VV03, Kok09]. Symplectic [Rei00, COR08, HL06b, HyLL07, HJL09, MGS09, QM03, SHWC07, Tan05a]. Synchronous [MMKP08]. Synergia [ASQR06]. synthetic [FWP09]. System [AKV00, Ano08-50, FK02, HK01, LP02, MP01a, MP02, MCCT02, VQSZ02, Wu02, de 00, BSW03, BL04, BP06, BLG^{+}08, CGMS03, CR09, CBS05, DMBS05, Del03a, Del07, Eli03, Eli07, GS06a, GS09a, GGMN^{+}09, HK08a, HMMR04, HJM^{+}05, HGB^{+}03, ILL09, JMZ04, Nat06, NMM^{+}07, Nis07, SM09b, SA09, VVM05, WO05, WO09]. systematic [YC09a]. Systems [Ben02, CWWZ00, CPP02, CM02, FMO00, GTD00, HZ02, HR01, HPZ01, KLN^{+}01, KKP02, LMSW02, Mu02, Noe00, PS02, PO01, RC00, San01, SSC00, AS05a, AGT05, AC09, BCB03, BKS07, BS09b, BTWGvBW07, BP07, CLS^{+}06, CORT09, CGP05, CP06a, Cap08c, Cap09, CBJdlC07, CHM08, CP08, DDD05, De 04, DK07, DKTT07, ELVE07, Edw06, FVE04, FHLO08, FHLK05, FT09, GV06, Gui05, HJFW04, HH07b, HM08, Hau08a, Hau08b, HC09, HS09b, Hwa03, HWW07, JRS05, JLOT05a, JHZ^{+}09, JAK05, KDOO05, KSHS08, KMV03, Kau03, KCMM03, KB08, LZL03, LL03c, LVW06b, Liu08, LCM07, LB04, LMZ^{+}08, Mad06, MM07, MMKP08, Mil04, MG07b, MC07b, Moo03, NZZ06, NFA03, NZ07, PGB05, PA07a, PDL09, PC06b, PBH04, Pro08, RE07, RP08a, RS05, RS09a, RBL04, Ros06, Ros08, RM07]. systems [SS03a, SWB^{+}06, SHS08, Str07a, SG03b, Tan05a, TT05a, Tok06b, TT06a, TT06b, THD09, TDV06, VHI06, VTM^{+}08, WZL04, Wen06, XS06, YS07a, YKG04, ZIP06, ZFM08, dSMF09]. Systolic [DHM03].

T [Har04, AMXL09, DZ09a, JHSZ07]. T-3 [Har04]. T/TM [DZ09a]. Tabulation [HGM01, LP09]. tailored [dNWvSD07, dTWD09]. Takizuka [WLC^{+}08]. tallies [GMH06]. tandem [MAL09]. tangential [GH08a, ND04, VQLZ04, ZVQ07]. Tank [LLQ^{+}02, CN05]. tanks [Fra04, LL08a]. Tapered [Car01]. target [HZ07a, MSO04]. targets [HSZ04]. tau [RE07]. taxonomy [EHS^{+}08]. Taylor [BZ08, CR00, Dar00a, FS06, FDD09a, GGL^{+}01, KB01, LBL08, NCS03, TM05]. TDGL [WA02]. TE [ZW05]. TE/TM [ZW05]. technical [BEA09]. Technique [AA02, BU02, CL02, GG00, HPZ01, KKR01b, MBP00, NTYT01, NTYT02, SML02, SHS02, WLE^{+}00, BGM08, Bet08, BP04a, BGN03, CB03, DDK06, FK07a, FM06, GPVB07, GLLN09, HLL08, LY07b, LJM^{+}06, Lar09, LKMU05, MKLU05, MCN03, OTCM08, OK07a, RVVL09, SGFL09, SP05a, WZ09, XYK05, YZLH09, KG09]. Techniques [Ben02, HH01, MM01, MOS^{+}00, Spo00, BS03b, CR08, CRB^{+}08, Dem04, DDDC07, HV03, KS08a, LWG03, PS07c, SY09b, SP05c, SMSS07, VBJ08a]. telemetry [CP03b]. teleportation [DS05b]. Telescopic [GK03]. Temperature [ELW01, JK02, HS04, KW03, LZ07, LP06a, MDR07, NVD05, Soc03, XHC08]. Temperature-Dependent [ELW01, KW03]. temperatures [SK08a]. Temporal [CV06, GHV00, Wee02, DGF09]. Temporally [Nys02]. Tensile [Mon00]. Tension [CS00, JTB02, Nie01, RR02, SZ01, Str01b, BN09, FCD^{+}06, GV08, GR07, Kim05, LLP07, LF04, Pop09, SAKDJ05]. Tensor [BTFY01, CS01a, Edw00, KKCF09, KMHR00, ML01b, KR09b, Lar07, Owe04]. Tensor-Diffusivity [ML01b]. Tensorial [PB00, NV09]. tensors [Hua05]. Term [HK00, WK01b, CVB06, DMR09, GS06b, JL04b, PPB09, SK04a].

Terms [BJ00, BBR01, HGN00, Xu02b, ZCMI01, ASPB03, BP03, HW08, KD09, KG08, LTZ03, MPFC08, RBT03, SZC09, Tol07, TE04, TT05b, Wen06, XS06]. terrain [Ano08-50, SM09b, WS04]. terrain-following [Ano08-50, SM09b, WS04]. Test [HS07b, BZ08, CHM08, DTSC04, GR08, KZWY09, SD06, ZRS06]. testing [Hig05]. Tests [BK01, LLIK01a, LLIK01b, MEG02, SPW^{+}00, DLMK04, SD05b]. Tetrahedral [BT02, MGGH00, MP01b, CBH03, DS05a, MP05, VGPL09, YJ06]. tetrahedron [DR06]. tetrahedrons [LJSM08]. Textbook [LDPL08]. Textbook-efficiency [LDPL08]. Their [LP01, Saf00, BZ08, GLMH09, HO08b, QS04, Ros08, Tak06, VB08]. theorem [BO09, Tow08]. Theoretic [HSK00, HSQ03]. Theoretical [Wag05, ZGSD06]. Theories [BTFY01, CLMRP08, Chr03, HvHHS05, LM08a]. Theory [AKY01, FS00a, FS00b, HWL08, JTB02, YLD09, AK09, BBB08, CXB08, CBC09, CFM09, FHW07, FLE03, GD07b, HMA05, JR03, JR04, JY08, KDF07, Lau04, LS05a, MG07a, PM07, RCT07, SF03, XSG08, ZK05]. Thermal [DDG02, GR01, PR00, Sie00, Cho05, DSM09b, EULM03, FHLO08, FBHV05, LCB09, LM08c, MC04, MHB08, MELD08, MC09, PSC04, PSMW09, Sof09, TSG^{+}06, YWC07]. Thermal-Creep [Sie00]. thermally [MY07, RWMK03]. thermo [BZ04, KK03a, KP08]. thermo- [KK03a]. thermo-acoustic [KP08]. thermochemical [KW08b]. Thermodynamic [GC02b]. thermodynamically [WAO^{+}04]. thermodynamics [MY03]. Thermoelastic [BM01a]. Thermohaline [DOWB01, AT09]. Thermomechanical [SMGJ09]. thick [LSJA05]. Thickness [GC02b]. Thin [CMK^{+}01, DDF01, DK02a, KK00c, TC01a, DJ04, ES03a, HKM08, JN07, LSJA05, LHGF05, MC06b, MK04b, MO06, SA06, SRX07]. thin-structure [MK04b]. Thin-Tube [KK00c]. THINC [Yok07]. Third [Boy02a, GST02, RDPN07, YC09b, CT09, Fox08, Hub07, ZKDT07]. Third-order [YC09b, CT09, Fox08, ZKDT07]. Third-order-accurate [RDPN07]. Thomas [PM00]. Three [AKV00, ART04, BFC04b, BZ09, BCMO01, CRB00, CMOV02, DIL03, DK02a, DOWB01, FVOMY00, FS00a, FS00b, FKK08, HD07, HK01, KP00, LL00, LTZ02, LK01, LDV08, Lou00, MC02, NTB07, Pai01, PKKL05, PWS^{+}02, PA07b, Saf00, SHWW00, SJ02, SS07b, Sni01, SS01b, WK01a, YXLF05, YW07, Yua02, ZSP02, ZYC02, AvdB04, AK05, AV03, AC05, AMS04, AMSZ07, BS04c, BBK07, BHP07, BCI^{+}08, CM06, Che04, CCG^{+}06, Che07, CFGK05, Dim07, DLP08, DLW06, Eli07, EES09, FNS07, FRS08, FCGK05, GG04, GS08, GS09b, GB03, GP04, GWF^{+}07, GH02, Gro07, GD06b, HZGB05, HP04a, HS08a, HLWW04, HLWW06, HWW07, HB05b, IHL03, JVVS07, JW03, KKS05, KAK03, KLP^{+}09, LWP^{+}09, Lee03, LZ09c, LDPL08, LT09b, LVW06b, LL08a, MCG08, MRRS05, MSJ07, Moo03, Moo07, MT07b, OLLL03, Pon09]. three [RB05, RS06b, SBCL06, SCRL08, SG03b, TTZ03, TM05, TXCD07, TT04, TT05a, TC07b, TC09b, TG08, TT05b, TA06, UL06, Wag05, WK04, WW04, XG09, YAvdB^{+}08, YBZ04, YKK08, ZP05, ZH09, ZLAC05, ZT07b, LMS02]. three-body [SG03b]. Three-Center [Saf00]. Three-Dimensional [CRB00, CMOV02, DK02a, DOWB01, FVOMY00, FS00a, FS00b, HK01, KP00, LL00, LK01, Lou00, MC02, Pai01, PWS^{+}02, SHWW00, SJ02, Sni01, WK01a, Yua02, ZSP02, ZYC02, ART04, BFC04b, DIL03, FKK08, HD07, LDV08, NTB07, PKKL05, PA07b, SS07b, YXLF05, YW07, AvdB04, AK05, AC05, AMS04, AMSZ07, BS04c, BBK07, BHP07, BCI^{+}08, Che04, CFGK05, Dim07, DLP08, Eli07, FNS07, FRS08, FCGK05, GS09b, GB03, GP04, GWF^{+}07, Gro07, HZGB05, HP04a, HS08a, HWW07, IHL03, KKS05, KAK03, KLP^{+}09, LWP^{+}09, Lee03, LZ09c, LDPL08, LT09b, LVW06b, LL08a, MRRS05, MSJ07, MT07b, OLLL03, Pon09, RB05, SCRL08, TM05, TXCD07, TT04, TT05a, TC07b, TC09b, TG08, TA06, UL06, Wag05, WK04, WW04, XG09, YAvdB^{+}08, YKK08, ZP05, ZH09, ZLAC05, ZT07b, LMS02]. three-dimensions [TTZ03]. three-space [TT05b]. Threshold [ET06]. Thresholding [RM01b, Moo03]. throat [CGH05]. Thuburn [TR07]. Thue [WWVG00]. tilted [DDGS09]. Time [AGH02, ACS00, BJM02, Bar02b, BKR^{+}01, BS00d, BCVK02, CN05, CM02, DOW08, Dur08, FD03, FKLY07, FMD^{+}09, Gen01, GHV00, Gui02, Hig02, KC00, KDW08, LBV01, LR01a, LP06a, Mad06, MBP00, MT01, MHPR08, NU09, Nys02, POS00, PM00, Rem00, RTT01, RB02, VG01, VR02, YP01, ZTZ02, APT09, AKV06, AFGM07, AB07, AK07, AG09, ACLS03, Ata04, Bal08, BBHM09, Bar04, BHNPR07, BHvdV06, BSW05, BH05, CWT00, CCJ07, CT08b, CS08b, CL07b, CJK^{+}03, CJ07, CFP06, CFJ06, CVE06, CFP08, DDD05, DR06, Den07, DL04, DWLM09, DKS^{+}03, DLP08, DGRS08, DDDC07, ELVE07, FK07b, FH03, Gab07, GGF03, GvH06, GN03, GP04, GK07, GD07b, HS07a, HR08, Han00, HC08, Hig05, HDBW05, HJM^{+}05, Hua07, HGB^{+}03, ISNY05, JG09, JMC03, KN09, KvdVvdV06a]. time [KvdVvdV06b, KvRvdVvdV07, KCMM03, KLM07, KT05, KWD07, KB08, KSGF09, LSL08, Lap04, Lau04, LWG03, LTE07, LP04a, LLL07, LXM09, LX07a, LT09a, LCS09, LS09, Liv07, LGM08, LJ07, Low04, LB04, MLSD07, MKOW04, MY06b, MELD08, MU09, MC07b, MPFC08, MK07, OS04, Ols07, Ols09, OK06c, OK07a, OPML07, PAD07, PH09, PR04b, PMP08, PA07b, RBSL06, RVDM09, RSS09, RCD05, RRW05, RSO04, RJM07, RMV03, RJ04, SROCFF03, SHWC07, SWZ03, SL07b, SC09a, SK08b, SV07, SHPC09, SK06, TZHT04, Ten03, TSB03, TCN09, TDGP06, UBRT07, VPMC04, VW02, VCG03, VSH04, VS07, Wag05, WG08, WRu03, WS04, WC08, YA05, YZW05, ZSW03, ZSW07, ZYC02, ZYL^{+}06, ZH09, ZZ09, ZRR00, dSHHM05, dHRvdB07, vZdBB07, vdVvdV02, vdVX07, vdV08, HW02]. time-accuracy [GGF03]. Time-Accurate [KC00, LP06a, MY06b, OK07a, RVDM09]. time-adaptive [CFP08]. Time-Dependent [AGH02, ACS00, Gen01, RTT01, VR02, FD03, FKLY07, AFGM07, ACLS03, Ata04, CJ07, DL04, DKS^{+}03, GN03, GP04, GK07, HDBW05, LWG03, LP04a, LB04, MU09, OPML07, RCD05, SV07, Ten03, WRu03, WS04]. Time-Domain [GHV00, Rem00, YP01, Lau04, LT09a, LJ07, MLSD07, MPFC08, PAD07, SHWC07, SWZ03, SL07b, VPMC04, VW02, Wag05, WC08, ZH09, dSHHM05, dHRvdB07, HW02]. Time-Driven [VS09]. time-evolution [DDD05]. time-fractional [LX07a]. Time-harmonic [MHPR08, APT09, AG09, BHNPR07, DLP08, Gab07, GD07b]. Time-independent [CN05, CCJ07, CJK^{+}03, Lap04, LTE07]. Time-Integration [BKR^{+}01, OS04]. Time-Line [Gui02]. Time-parallel [FMD^{+}09]. time-periodic [MKOW04, vdV08]. time-resolved [Bar04]. Time-reversibility [DOW08]. time-reversible [PH09]. Time-Scale [LR01a, VG01]. time-space [LS09]. time-split [SK08b]. Time-Splitting [BJM02, MBP00, HJM^{+}05, KN09, TCN09]. Time-Stable [Nys02]. time-staggered [SK06]. Time-Stepping [Hig02, RB02, ZTZ02, Mad06, DR06, HR08, Hig05, LGM08, MPFC08, VSH04]. Times [QS01, Del03b]. Timesaving [SMSS07]. Timescale [Bar02b]. timeseries [CVE06]. Timestep [Car01]. timestepping [HSBG05]. timesteps [Pet07]. tissue [HK08c, KL06, XDB09]. Title [Ano00-28, Ano01-28, Ano02-28]. TM [DZ09a, ZW05]. tokamak [HJKO08, LL08b]. tokamaks [CTS07, LGKP07]. Tomography [CBB01, HCG01, BO05, CCT05, FLE03, IKL^{+}08, RR07, TMND07, THN^{+}07]. tongue [SP07]. tool [ASQR06, FK09b]. Tools [KT02, Küm04b, LH08a]. Topography [Hor02, BGN03, FG07, GPC07, Geo08]. topological [BHR04, HKO07, VCG03]. topologies [KT05]. topology [AS05b, AA06, Bey09, LTWW07, LLC^{+}08, LTM09, WLKW07, ZL08b]. topology-preserving [AS05b]. Toroidal [GST00, KP00, ZYKW01, BT07a, BT07b, CTS07, ORM06]. total [CT04, CCT05, SLG^{+}03, YMW06]. total-energy [SLG^{+}03]. Tracing [LM01, LM03a, MJT06, MCP03, THN^{+}07]. Tracking [Asl01, ČPT01, CSP01, JC02, NSC09, SJ02, TNGH02, TB00b, THS07, TBE^{+}01, ZH01, AMS04, BR09b, Che04, DDS09, DFG^{+}06, Fan08, FCD^{+}06, GNNB08, HSL08, KLSW09, LLP07, LS08, LDW07, LLGL07, LHGF05, kM07a, MT08, NT07, PP04, PMP08, QS07, QLS09, SPM03, Shy06, SB07, TZ06, Vol04a, WKB07, ZKDT07, ZEA06, ZL08a]. tracking/front [dSMN^{+}04]. tracking/ghost [TT09]. traction [Liu09b]. traditional [Kas07]. traffic [LGN05, ZSWW03, ZWS06]. trajectories [DDD05, MESV09]. trajectory [NSS03, TW03]. Transfer [BS00b, BW01, DK02b, Gen01, IYI^{+}02, LTK^{+}02, Cha07a, Cha07b, CS03, DL04, DUEB07, FDK06, FKLY07, HL04, HC09, HDBW05, JJGL06, JJGL07, KM03, KNH05, LSA06, LCNR07, LR03, MHB08, MELD08, MU09, MR07c, MAN^{+}06, PS07c, RW08, Thö04, TFDK04, WHS08, WMH07, YLD09, YSW06]. Transform [BTSM09, AB05a, CdHST08, HSQ03, KOQ08, OLLL03, SB06a, SS09a, WK06, ZGSD06]. transform/potential [HSQ03]. transform/potential-theoretic [HSQ03]. Transformation [MBM01, DT03, HHMK05, KR09a, SK05, WS04, ZKDT07]. transformation-free [KR09a]. Transformations [Saf02]. Transformed [Eli02, Eli03, Eli07]. Transforms [SS00, Kry04, VBJ08a, VBJ08b, VB08, WJV07]. Transient [CMR08, HLS02a, LWEM00, AFGM07, Hag07, JG09, Kwo08, MR07c, NPH09, PKD07, SO08, vOP04]. transients [CGMS03, FF03]. Transition [BRL02, GP00b, DJM05, EKP06, GC06, JOS06, LSL08, LZ04, Liu08, LD09b, Sus06, ZT03, vEB05]. transitional [DS06b, JD04]. Transitions [EKK02]. Translation [GM01c, GD07b]. Transmission [Wu01, BNV08, BS04d, PSH^{+}08]. Transonic [EAY01, MSJ07]. Transparent [AST07, DKSW01, FSY00, SFY01, YFS01, dSHHM05]. Transport [AS03b, AL01, Azm02, Bal02, Cul01, DV02, DB00, FW07, GHG01, MD04, MGGH00, Noe00, OF02, UH01, ZZ01, ZKK01, deM02, AT05a, BP06, BMN07, BCCV09, BES07, BS07, BNNP06, BMDS05, BSP06, BH05, CL03a, CL05, CL08d, CLL^{+}07b, CS04, DMBS05, DGM07, DL03a, DUEB07, DC08, FWP09, FH07, GS05a, GS05b, GS08, GC06, GYKL05, GLT07, GL09b, HLFB07, HJKO08, HF08a, JLT06, JSCZ08, JN07, KB04, KL06, KAS08, KS08b, LZT09, LRMB08, LFX05, Li08a, LD04, Mac07, MBS03, MGNB09, NSS03, NZ05, Ols09, PA05, PL07, RRC05, Rom07, RF06, Ros09, SZ08, STD^{+}05, SCC^{+}03a, SCC^{+}03b, SY08, SXyWX09, TX06, TMSW07, TFD06, TA06, UBRT07, WR09, XP04a, XDB09, Yeh07, YE05, ZWS07, ZEA06, Zie04, dA04, dDEK09, dFGLS05, DW00]. transport-diffusion [DUEB07]. Transport/Advection [DB00]. transport/reaction [STD^{+}05]. Transportation [XY01]. transported [MJ07]. transpose [JH08]. transverse [LKD04]. Trap [BMS00]. Trap-Assisted [BMS00]. trapped [LMK03]. Traps [Whi00]. Travel [QS01]. traveling [EV03, MJ09a]. travelling [Boy03]. traveltime [TMND07, THN^{+}07]. traveltimes [QL04]. Treating [SHS02, MP07b, WG06, YHCD05, YW07]. Treatment [CL02, ELC02, HK00, Li08b, ML01a, MC00b, ZCMI01, AT05a, CVB06, JL04b, LL07, LP04b, MY03, PSG05, SB06a, SAK05, TA06, WAO^{+}04, ZJWC08]. treatments [JSCZ08, KY08]. Tree [BADG00, WPM^{+}02b, COQ06, Pop03]. tree-based [Pop03]. Treecode [LK01, LJK09, Wan04b]. trees [ARRS09, CMP07]. Triangle [BM01b, GW05, Gir06, Hei05]. Triangle-Based [BM01b, GW05, Gir06]. Triangles [CDKP00, PR04a]. Triangular [HL01, WB01, FD07, GGMN^{+}09, Jar04, KI05, KDW08, LGHD08, LSSV07, LNXNTX09, MJT06, Pon07b, SPM03, YJL^{+}06, YJ06]. Triangulated [Car02, KOQ08]. triangulations [CP08]. triaxial [San03]. Trickle [PCCD00]. tridiagonal [PSH^{+}08]. tridiagonalization [WR09]. trigonometric [QM03]. Trim [BTSM09]. Trim-to-Coherence [BTSM09]. Triple [FK09a, KKGL01, Liu09c]. Triple-decker [FK09a]. triply [JCT07]. triply-periodic [JCT07]. troposphere [MM09]. Trotter [MC07a]. Trouble [Boy05b]. troubled [BAMD07]. true [HAP06]. truly [GS03d, LMX^{+}08]. Truncation [HNGB04, Yam01, Jon05, KK09, Lap04]. trust [BC08, HE07]. trust-region [BC08, HE07]. Tryggvason [Khe04]. Tsallis [FH02]. TSFP [Ano04-27]. TSFP-4 [Ano04-27]. tsunami [FNBB^{+}08]. Tube [KP00, KK00c, Sie00, Low05, ZEA06]. tubes [TX06]. tumor [ML05, ML06a]. tuned [HP04b]. tunnel [SSW^{+}07]. tunneling [DGM07]. turbid [Bar04]. Turbulence [BRL02, BZB00, FLG01, FSM^{+}01, KP00, LS02c, LP02, Pir02, SLY02, SPW^{+}00, BB09a, BL09c, CP07, CC05, DDH05, DLD08, DS09b, GBB^{+}06, GS09d, HHPW08, HHMK05, HM04, JOS06, KMID05, KMSH08, KAS06, Lar09, LDN04, LQ06, MTWW06, MC06a, ML06b, PHW08, SKWN03, SCC^{+}03a, SCC^{+}03b, TWM07, TMD07, UPKN09, WGRA09, YSO07, YGL05]. Turbulent [EAY01, GMB01, JPMC01, LS02a, MK02b, MPC01, MPC02, PPC00, SS02, TSB01, AGW07, BFB08, BIVC07, CRAG07, CMP07, CZ09, DMP08, DBBP08, DS09a, DHM07, FE04, Gra06a, Gra06b, HPD09, HM05, HO03, IK07, KIH09, KIHM09, KM06, KM07b, LP06b, MLM09, MJ09b, PPDM08, Pro05, Pro07, RJ06, SS07a, SJHM09, SFMP06, VC03, VV03, XLP05, YB06]. TVB [BBCT09]. TVD [GC01, HL04, KT04, PL09b, SPGR06, YL01]. TVD-interpolating [HL04]. Two [AJG01, ART02, ACS00, Bar02b, BMR01, BMRS02, BdLL01, BZW01, BH05, CFA01, Cal02, ČPT01, Cle00, CD00, DCV^{+}01, EKK02, Eli02, EF02, FT01, FS00a, FS00b, GS02, GW01, Goe00, GP00b, GKL03, HLS06, Hig02, JWSC00, KK00b, KLvBvL02, KMHR00, LKD04, LG09, LTZ02, Low05, LWEM00, Mai01, MR04, Nys02, ODAF06, OS01, PKvdB00, PS01, RC06, Saf00, Saf02, SWL00, SS04, SP00, TC02, TGB^{+}07, Tow07, VD02, WK01a, WL02, WB01, Xu01a, YSS07, ZYC02, ARRS09, AV03, AW04, AT09, AMS03, BTW04, BM07, BW07, BH04, CGRGV^{+}04, CA06, CHL06b, CSO09, CSL08, CY05, CMR08, CC08b, CDV07, CkM07, DMBS05, DM03, DDK06, DP07, DP08, DCF^{+}08, DSS07, DDS09, DS09b, DR09b, EGHE06, Eg07, ECL02, Eli03, EES09, EF03, FRS08, FJ09, FHLK05, FCT07]. two [GS09a, GGP06, Gro06, GR07, GD07a, Gui05, HT07, Hel05, HLO08, Her05, Her08, Hig05, HZ07b, HB05a, HT03, HH06, IOTK04, JA08, JBF07, JX06, JN07, KSHS08, KLM05, Kro01, LCB04, LLP07, LSD07, LL05, LS05a, LT09b, LMS08, LTC07, LP04b, LM03b, LHGF05, Ma05, Mai03, Mai04, Mai09b, Mai09a, kM07a, MR06a, MMS04, MR05, MST06, MP03, Men04, ML06b, Mou04, MGNB09, MG05b, MAL09, Nit05, OK05, OKZ07, QA09, QLK07, QS07, RMB07, Ram06, RRC05, RMG^{+}09, RMF08, SWK06, SY09a, SS06a, SYC09, SL03, Shy04, Shy06, SSND03, SXyWX09, SSH^{+}07, TM07, TOZP03, TTZ03, TMB07, TM05, TPV07, Tol08, VVS08, VCG03, VD03, WZL04, Wen09, WO09, YZ07, YYT05, YBZ04, YF09, YE05, ZLAC05, ZHSS09, vBK03, Cap06, JW02]. Two- [FS00a, FS00b, ZYC02, TTZ03, ZLAC05]. Two-Body [Mai01, Mai03, Mai04]. Two-component [SS04]. Two-Density [OS01]. two-diagonal [Tol08]. Two-Dimensional [AJG01, ART02, ACS00, BMR01, BMRS02, BdLL01, BZW01, Cal02, CD00, DCV^{+}01, Eli02, Goe00, KK00b, LWEM00, PKvdB00, VD02, WL02, BH05, LKD04, LG09, MR04, ODAF06, RC06, ARRS09, AT09, BTW04, BM07, BH04, CSO09, CY05, DCF^{+}08, DS09b, ECL02, Eli03, FHLK05, GGP06, Gro06, Gui05, HT07, HT03, JX06, JN07, KSHS08, Kro01, LSD07, LS05a, LT09b, LTC07, LP04b, Ma05, Mai09b, Mai09a, MMS04, MST06, MP03, Men04, MGNB09, RRC05, SSND03, SS04, TM07, TOZP03, TM05, TPV07, VVS08, VCG03, VD03, WZL04, YYT05, JW02, Cap06]. Two-Electron [Saf00, Saf02]. two-equation [ML06b]. Two-Fluid [ČPT01, HLS06, KLvBvL02, TC02, CDV07, EF03, FJ09, Hel05, JBF07, QA09, SL03, Shy04, vBK03]. two-layer [CGRGV^{+}04]. Two-Level [Hig02, CSL08, Hig05]. two-medium [QLK07]. Two-Phase [CFA01, ČPT01, Cle00, EF02, GS02, GW01, GP00b, PS01, SP00, Xu01a, Low05, TGB^{+}07, YSS07, AW04, AMS03, BW07, CA06, CMR08, DM03, DDK06, DP07, DP08, DSS07, DDS09, EGHE06, FRS08, GR07, GD07a, Her05, Her08, HZ07b, HH06, IOTK04, LL05, LMS08, LM03b, LHGF05, Mou04, QS07, RMB07, RMG^{+}09, RMF08, SWK06, SY09a, SS06a, SYC09, SXyWX09, SSH^{+}07, TMB07, YZ07, YF09, YE05]. two-phase/vapour [BW07]. two-point [Eg07]. Two-Scale [EKK02]. two-species [LCB04]. Two-Sphere [KMHR00]. Two-Timescale [Bar02b]. Two-Way [FT01, WK01a, CC08b, GS09a]. Type [Gui02, Han00, HT00a, HT00b, Shy01, WH02, AINR03, BP03, BSLN09, CLS05, CS07a, FNBB^{+}08, GGCC09, Hel09b, Lar07, LNGK04, LG05, sLwG08, LLOT06, LCS09, Löh04, NJX09, NF09, PSD09, PL09a, Shy04, TD07, Wen07, WF06, XXS07, XLS09a, AT05b, CJR04, JHZ^{+}09, LD04, MN09a, PK03, TB06, XHW07, vBK03].

Uhlenbeck [Del03a]. Ultimate [Abg01, VU04]. Ultra [HMK02, BH09, BMK^{+}06, HMM07, KQW03a, KQW03b]. ultra-relativistic [BH09, KQW03a, KQW03b]. ultra-violet [BMK^{+}06]. Ultra-Weak [HMK02]. ultrashort [Sau04]. Unbounded [CR02, BHNPR07, BP08, DD03a, DD03b, HZ08, VZSL07]. Uncertain [Hor02, EN06]. uncertainties [AA07, AA09, LSK06]. Uncertainty [BPM06, CGH05, CDE06, KG06, LKNG04, Ler06, PDL09, YZL^{+}06, CDI09, DEHL06, LNGK04, LK07, PIN09, XK03, XS07]. Unconditionally [AB03, Azm02, CYKC01, JTL09, KR02, ML06b, NFGK07, ZZ01]. under-resolved [TV08]. undergoing [CGDT09]. Underresolved [CS01b]. Understanding [DWC^{+}09]. underwater [FRS08, KS08a]. Unequal [Zha02]. uneven [DL03b]. unevenly [Mil05]. Unidirectional [dSHHM05]. Unified [HK01, KAA^{+}07, Wu02, Xia04, XAI06, DBTM08, FK09b, Jia07, JX07, LZ04, Meh04, MY03, SW08c, WD07]. Uniform [SV00, Cap08a, FCT07, HKG08, Hu05, HSS07, ISNY05, KK03b, LCG07, NVD07, SZ05, STZ07, Tor03, TB04, Vas00, VSW06, YA05, ZIP06, ZT03]. uniformity [NVD05]. Uniformly [BLM08]. unifying [WG09]. Unit [VQSZ02, Hei04, JA08]. units [ALT08]. unity [GLN06]. Universe [BADG00]. Unlimited [NT07]. unmagnetized [MD04]. unsaturated [LMH07]. unscented [IKL^{+}08]. Unsplit [Hu01, CCF^{+}05, EB06, GS05b, GS08, LD09a]. unstable [AZ05, FCT07, GKE04, KG03]. Unstaggered [GHV00]. unsteadiness [CGM07]. Unsteady [BMRS02, BMQS02, BCVK02, BL01, BGN03, GSD01, KC00, LHD05, QV01, VC00, WB09b, ZYC02, AM03, BLM04, BCI^{+}08, CTW^{+}08, DT04, DPRN05, EHD08, GS07, GMAj09, JMC03, KZ04, LDPL08, LF05, LKX04, LGM08, LZH^{+}06, MLS^{+}05, MGNB09, NJX08b, Pon06, RDPN07, SC08a, SFE07, SY03, TZ03, TZL05, TJS03, Tsy03, VBL03, Wan05, WM07, WGS^{+}08, XYK05, You06]. Unstructured [BM01b, BW01, DV02, DPCV02, Edw00, HZ07a, HW02, JK00, KC00, LM01, MVM02, Mav02, ML01a, MD06, MG02, OGV02, PW00a, Per00, SC01, SMP01, Wan02, WL02, Wan05, WPH00, WB01, ZSP02, ZTZ02, ZQSD08, dSAK00, AZC05, AB05b, BFB08, BES07, Ber06b, BS03a, BM07, CKvT07, CDDL09, CSO09, CP08, CSKD05, DSM09a, DMR09, DK07, DKTT07, DBTM08, DZ09b, GS09b, HZGB04, HZGB05, HWL08, HV03, HNF07, Her08, HHMK05, JH06, JMC03, JS05, KT03, KT05, KE09, LM03a, LCH03, LK09, LJW09, LSS06, LSSV07, LVW06a, LVW06b, LLB05, LBL06b, LBL07, LZH^{+}07, Mai09b, MB04, MCP03, NOG08a, NOG08b, NJX08b, PL09b, PN03, RAD07, RRW05, RWWS07, Ros09, SS05b, SP06a, SP06b, SWL06, TZ03, THD09, TT06c, VSW04, VSW06, WZL04, WLC^{+}06, XLS09a, XLS09b, YJL^{+}06, YA05]. unstructured [ZLAC05]. Unstructured-Grid [SMP01, SS05b]. unstructured-multigrid [LZH^{+}06]. Untangling [VGS04]. Update [Xu02a]. updated [GCCD07]. upper [GG09b, MM09, ZK04]. Upscaling [DGH08, EPW08, Kou07, Nov04, PC06a]. upstream [ST03a]. Upwind [CRD02, Hwa03, PD01, STiST02, WB01, AD04, BGN03, BL03, Cap05, Cap06, Cap08b, CS09, DE06, IM07, JAK05, KK07, LWW04, LJ09b, LD04, PYC04, RS06a, RB09b, SGD03, Ser09, SS09c, SS05c, SB03, WZ03, ZYHS07]. upwind-biased [JAK05, PYC04]. upwinding [CD03, XD07, ZKDT07, ZR08]. UR [Har04]. Use [DPRS01, MD02, PS02, TK00, VG02, DTMS06, Dic08, GS03c, KFIG06, NLT07, RB06, RBSL06, Ram03, SPLM09, VTW^{+}07, WG08]. used [KN09, Kau03]. Useful [Saf02]. Using [AC01, AZ05, BM02, BC02a, BMRS01, BMRS02, BT02, BRL02, Bon00, Bow01, BCMO01, CS01a, CBMO02, CSV00, CL02, CL00b, CB07, DDD05, DGH02, GW02, Gos02, HAAO00, HHL00, HR01, HF01, HPZ01, KMA^{+}01, LS02a, LBV00, LP02, LLQ^{+}02, MR00, MKM99, PM02, PR01b, RS02, RRL01, Saf02, SSSWD00, ST01, SJ02, SSD00, TK02, TR02a, TTSG01, hRT02, WPH00, Whi00, WHV^{+}00, ZYC02, ZF02, ZKK01, APP^{+}07, AMH04, AJT04, ÁDIM09, AA06, BLS08, BS04a, BBD04, BIS07, BPO07, BG09, BT09, BCGR05, BJP04, CJLS09, CWJ07, CPG04, CR08, CQO04, CM06, CP04a, CCT05, CEL06, DMHP07, DDK06, DK06, Del03b, DL03a, DW09, DS09a, DST07a, ELD08b, EKP07, FT05, FW07, GMD03, GGS09, GWF^{+}07, GKJW07, Gri09, GLLN09, GYKL05, HPS06a, HZ07a, HBHS09, HSBG05, HKS09]. using [HS08a, HSZ04, HL07b, HSL08, HF08b, HMM07, IKL^{+}08, IM05, ISNY05, JW09, JS05, KKS05, KW08a, KIH09, KK05b, KH07, KYLB07, KZ06, KF06, KR09b, Küm04b, KLP^{+}09, LDN04, LKNG04, LWG03, LCG07, LM08b, LAKD08, LK09, LDL^{+}09, LNXNTX09, LY04, LBL04, LTWW07, LTL^{+}09, LTM09, MKM04, MR06a, MGCR07, MPD08, Men04, MC03, MGS07, MCN03, MR07c, MHW05, NJLA06, NM06, NI03, OJW06, PS03a, PM07, Pon07b, Pro03, RRC05, RA09, RBT03, Ros09, SRM09, SDR07, SP04, SL04, SO08, SL07b, SAKDJ05, SZ05, SNLS03, SSND03, SCW^{+}09, SGG^{+}04, SP05c, SJC07, SDT08, SFMP06, zSW06, zS06, SB07, TZ03, TZL05, TLK07, TZ07b, TZ07c, TFD06, TBJ^{+}09, TGB^{+}07, TdAAP08, VCG03, VS07, WL03, WWC07, WLC^{+}06, Xia04, XAI06, XYK05, YM07, YFLS06, YYT05, YSO07, YGL05, YZF^{+}06, ZGT06, ZG08, ZC09, ZKS^{+}09, ZSTC06]. using [ZQSD08, ZQ09, dTDI^{+}07, dCNHSD07, vdBG09]. UV [Bor03]. UV-suppressed [Bor03]. Uzawa [BT02, PS07d].

V [LVW06b]. vacuum [CTS07, KSHS08]. valid [CTS07]. Validation [BP08, MHS02, OB06, BT07b, BCM^{+}07, MvW08]. validity [WZ07]. VALIS [SA09]. Value [DKX00, DKX01, KJ01, OKL01, ABLS05, BM05, BS05, Eg07, FF03, Kas07, PSM08, RMGK04, SN06, YH07a, dCNHSD07]. Valued [MF01, JLOT05a, LW07]. valve [vLAvdV06]. Vanishing [KK00a, PSZ09, SS07a]. Vapor [JLCD01, JWSC00, JW02, AMH04, JW03, Sus03]. vaporizing [TMB07]. vapour [BW07]. Variable [Alb00, BR09a, GQ00, SBGK00, Wan04b, AT09, Ber04, BK08, BRP05, CCG08, DBBP08, FG07, Geo08, GS09c, GD05, HyLL07, HL05, IQ08, KKM08, KKS05, KLP^{+}09, LT05, LP06a, MGC06, MDR07, Ni09, OK06a, PS03a, PS07d, RVM07, RVDM09, SD05a, SD05b, SHTB09, TBT^{+}09]. variable- [BRP05]. variable-density [AT09, SD05b]. variable-node [KLP^{+}09]. Variables [AD01, Hu01, BB07b, Hau08a, Hau08b, IA06b]. variance [DL03a, HH07c, VU04]. variance-conserving [VU04]. variant [GvH06]. variants [JHZ^{+}09]. variates [GL09a, HKM07]. variation [CT04, Kar04]. Variational [BCOS01, DCS00, DL03a, Hua01b, HS03b, HMK02, Lap04, Li08b, MN02, NZ05, WGRA09, AZ06, Aza06, CM06, CCT05, FDL08, GZ07a, Gra06a, Gra06b, HMM07, JCT07, MSO04, WSTW09, YFLS06, ZHSS09, ZL08b]. variations [Soc03]. various [GMO04, PL07]. Varying [CKS00, AKLMP09, GTMC08, Kou09, TZHT04, VCZS04]. Vector [BS01, CSV00, Whi00, BO05, CJ09, DQ04, FWR07, IA06b, JVVS07, LY07a, MBS03, OCFF08, QA09, RRW05, SR09a, SJD05, ST03a, YHSX07]. Vectorial [GBGM01, FCJ08a]. vectorized [FLE03]. Vectors [VSMW01, AL06, RMB07]. vehicle [ELD08b]. velocities [BFJ03]. Velocity [BRL02, Cul01, DC01, FPK08, MM07, MC07a, MF00, Mie00, BL09b, BHR03, CFS09, Car09, CEL06, DBS06, GD07a, GK05, KM06, KM07b, LY04, MC06b, NMS07, Pap08, PM08, Pon06, SH07a, SLC07, SS05c, Tan08, TG04, WFC09, WS09, YAvdB^{+}08, ZSC07, ZXQX08]. velocity-estimation [PM08]. Velocity-induced [MM07]. velocity-pressure [NMS07, Pon06]. Velocity-Vorticity [DC01, LY04]. Verification [MPD08, Roy05, Tak06, WLC^{+}06]. Verified [HPD09]. Verlet [MC07a]. versatile [HHC08, MDB^{+}08, NC04]. Version [MR01, GH02, GHMP07, LCM07, VMN07, XAI06]. Versus [Mav02, ABHT03, NVD07]. Vertical [BRL02, TW05, TR07, FCT07]. Vertically [MM09]. Very [DZ09b, GSV09, NK08, STZ07, TR02a, DET08, Heu03]. Very-high-order [GSV09]. vesicle [DLW04, DLW06, ZDD09]. vesicle-substrate [ZDD09]. vesicles [GFG09, VGZB09, VGBZ09]. vessel [CGN^{+}07]. VI [SWL06]. via [AS03b, BHP07, CFM09, Dim07, EE08, ES03b, GS05b, GS08, HS07a, JY08, JMK01, KK09, Kry04, LJS08, ML05, OVG07, OK07b, SW00, SKAS01, Sur05, TB00a, Tow09a, XK03, ZL04, ZSC07, ZW04]. Vibration [SCD00, SZC09]. vicinity [KZWY09, LL07, ZSW03]. victoria [SM09a]. violet [BMK^{+}06]. Virtual [FHJK09, GJK09, Lee03]. Viscoelastic [PS01, APTJ^{+}04, APP^{+}07, BPL06, FD03, FKK08, LC03, MDM03, TdAAP08, TCM05, VC03, VCT09, YSS07, YZF^{+}06, vOP04]. viscoplastic [BZ04]. Viscosity [Alb00, CS01a, ELW01, KK00a, LP00, SS03b, BL09c, CLG07, CL06a, Cho05, CC04, DLD08, JA08, KKS05, KR09b, Mac03, MLM09, Nov04, Owe04, RBH03, RMSB09, Sar03, SS07a, SK04a, TLL^{+}08, VHI06]. Viscous [CKR00, CKR01, CPK02, GPH^{+}01, Hun01, MK08a, PW00b, PW01, PSN00, QV01, RH01b, Sum00, TC01b, WPW02, Xu01b, ADR08, BL09b, BF08, BTW03, CN05, DS06a, DND06, FP08a, GXW07, GH09, GGS09, GGF03, GMD07, GGP06, GN07, HEN09, HL07c, HSL08, HLY09, JX06, JX07, KR09a, Kel05, LKP06, LLL07, LKO05, LX07b, LDV08, NBLQ09, NJX08a, PKD07, PSC^{+}06, PWM06, RW03, SROCdPFF05, SC08a, SZC09, SY03, SWL06, SK03, TZ03, TZL05, TLL^{+}08, VGZB09, VGBZ09, VD03, WFC09, WB09a, XH03, XMT05, Xu08, ZKY05]. Viscous-Plastic [Hun01]. Visibility [TCO^{+}04]. visible [BMK^{+}06]. vision [FSS03]. visual [Asl04b]. Vlasov [AV02, BS03a, BLG^{+}08, CDL05, CLS09b, Eli02, Eli03, Eli07, EB06, FBFF00, FSB01, GHB03, GS06b, HZ02, HF01, HGB^{+}03, IITV07, IKS^{+}09, KB04, MCCT02, SG06, SA09, VVM05, VTC^{+}07, WO05, WO09]. VOF [AZB09, GMD03, GW01, LHGF05, LF04, MZ07, Yok07, ZTZ02]. VOF-model [LF04]. void [TU04]. voids [AIR03]. Voigt [CPG04]. Volume [Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano00z, Ano00-27, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02x, Ano02y, Ano02z, Ano02-27, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano07-33, Ano07-34, AMSZ03, BM01b, CL00a, DKX01, DPCV02, HF00, Her00, HT00b, JM00, KC00, KKC01, KFV^{+}05, KKGL01, LLH02, LLIK01a, LM01, LM03a, LSW08, LMSW02, MP02, ML01a, MPC02]. Volume [NTYT02, OGV02, PKP01, PW01, RRL01, RR02, SZ00, SBGK00, SP00, Tol02a, Wan02, WL02, WW00, ZRR00, APTJ^{+}04, APP^{+}07, AZC05, AT05a, AT08, AKLMP09, AKO09, AMS03, BAFL09, BES07, BP03, Bot06, BKLL04, BLM04, CT09, CCG08, CMSZ09, CX08, CEH09, CSKD05, CR09, CZVS04, CFP08, DSM09a, DSM09b, De 04, DBF08, DK07, DKTT07, DET08, DBTM08, Dwi08, Edw06, EZ08a, FCD^{+}06, FMR09, GPC07, GLM07, GV07, HBHJ08, HJ09, HWL08, HLO08, Her09, IX07, IX09, IDD04, JL04a, JLT03, JLT06, JL09, KDK^{+}07, KLK08, Kok09, KS09, LSB04, LGP09, Lap03, LZT09, LYC09, LSSV07, LSV09, LVW06b, LHGF04, LH08a, LMNK07, LJ06, LZH^{+}07, MP07a, MSJ07, MZ07, MGS07, MCP03, MR07c, MT07b, NOG08b, NBLQ09, NPPN06, NXS07, NGC^{+}07, OK06a, OSK09]. volume [PL09b, PS04, PS08, PP04, PSG05, PL07, QM03, RJ06, Ros09, SE09, SJD05, SPM03, SMS04, SS06a, SL07b, Shy06, SMAj08, SC09b, SWL06, SR09b, Sus03, TVMR03, TPV07, TT04, TGB^{+}07, Tor03, TA06, TAL09, VL07, VLW07, VBL07, VGPL09, VSW04, VSW06, WZL04, WL06, WTL08, WG09, WA08, XS06, XCRX08, XLP05, XLS09b, YS07c, YS08, ZKDT07, ZH04, ZZ09, ZLAC05, vDZ06, Lab09]. volume-finite [CCG08]. Volume-of-Fluid [AMSZ03, KFV^{+}05, RRL01, RR02, SP00, AMS03, JL04a, PP04, Sus03, TGB^{+}07, YJL^{+}06]. volume-penalization [KDK^{+}07]. volume-preserving [QM03]. volume/difference [WG09]. Volume/PDF [LM03a, LM01, MCP03]. Volumes [GW01, XMP07]. Volumetric [ZKL^{+}07, KE09]. von-Kármán [YKG04]. Voronoi [GS09b]. VORPAL [NC04]. Vortex [BB04a, Car02, Cor00, CKS00, CMOV02, CP04c, HRV08, KK00c, LK01, MD02, MKM99, Nie01, Nit01, PW00b, PW01, PWS^{+}02, SC08b, WK01a, Alb09, BS08a, CLB08, CWD08, CC08b, ECL02, Eld07, FDD09a, Her05, HSC09, LG03a, LG09, MKM04, NSC09, SZC09, SDT08, TB06, WG06, YSO07]. vortex-dominated [TB06]. Vortex-in-Cell [CP04c, CWD08]. vortex-induced [SZC09]. Vortex/Impulse [Cor00]. Vortical [DS01, LK07]. vortices [DJ04]. Vorticity [BRL02, Cal02, Che00b, DC01, IK01, LFS07, MGGH00, MF00, QV01, Tol02a, Tol02b, AKH06, BBvdV06, CL06a, DBS06, Eld08a, KJ09b, LL04a, LY04, Pon09, PGN08, WFC09]. Vorticity-Based [QV01, Eld08a]. Vorticity-Divergence [Tol02a, Tol02b]. Vorticity-preserving [LFS07]. Vorticity-Velocity [MF00, DBS06, WFC09]. Vries [CkM07, LGK06, LY06].

Waals [CL01a, GV02]. Wachspress [AC05]. wake [BC08]. wakefield [HDR^{+}06]. walk [FG05, LLTA07, MS04, VSV03]. Wall [FG02, GGP06, BBW06, CP04c, FPK08, FGP08, GE07, HPD09, IK07, KMID05, KIHM09, KDK^{+}07, KAS06, KB06, LQ06, MC06a, NTB07, PPDM08, SKWN03, SFMP06, SN08, Tuc03, MK02b, Rid00, RVVL09, Sum00]. wall-boundary [GE07]. Wall-Bounded [FG02, CP04c, FPK08, FGP08, HPD09, MC06a, PPDM08, SFMP06]. Wall-driven [GGP06]. wall-function [KAS06]. wall-pressure [KIHM09]. walls [TX06, VHI05, ZTPM05]. Wang [Del03a]. WARP [GWF^{+}07]. waste [KP07]. Water [BC01, BST01, Che00a, FR02, Gir00, GHW02, Hor02, LBV00, LBV01, Lay02, LLIK01a, LLIK01b, Tol02a, Tol02b, TTSG01, VS02, Xu02b, ZCMI01, AB07, AB05b, BST03, BN04, BES07, BRC^{+}09, BTT08, BB09a, CVB06, CHL06a, CL08a, CGRGV^{+}04, Che03, CX08, CLS09a, CZVS04, DJTT05, EKBL09, FG07, GPC07, Geo08, GPF03, GW05, GD05, HS09a, HC08, KJ09b, KLM05, LHD05, LGHD08, LS03, LHZW05, LKW05, LMNK07, Ma05, MY06a, Mea04, MGNB09, NI03, NPPN06, NXS07, ODAF06, RAD07, RB09a, SS03a, SHTB09, TOY09, VTT08, WTL08, XS05a, XG09, vdVX07]. Wave [AGH00, AGH02, BM01a, BS06a, BZB00, CS01c, CSV00, DF00a, Dur00, ERT02, FT01, GF02, HHCL01, HK02, Kan02, LL00, LMSV00, LTD04, LAS01, LWEM00, LH05b, Noe00, Rei00, RTT01, Vay00, Vay01, VR02, Wee02, ZB07, APT09, AK07, Ata04, BN04, BP09, BPO07, BO09, BG05a, BG09, BS04d, CHL06a, CL08a, CPG04, CCJ07, CWL08, CHG^{+}07, CLS09a, CBI^{+}04, CFGK05, DNS08, DS09a, Edw06, EV03, FK09b, FCGK05, GS09a, GFS08, Gom08, GGOB04, GA09, HMOG08, HLS06, HPS^{+}06b, Jan08, JW06, KFH^{+}04, KFIG06, KSH^{+}06, KT06, KFV^{+}05, Lau04, LP07a, LZL03, LG03b, LG04, LS09, Ma05, MN06, MHI08, Pir07, PSG05, RBL04, Ros06, SDD07, SK05, Shy06, SD06, TET09, Ten03, Thu08a, TC07b, TC09b, VWW04, Vol04b, Wan04a, XS07, Yan09, YS09, ZH09, Zhe07, dHRvdB07]. wave-body [YS09]. wave-capturing [Edw06]. wave-current [SK05]. Wave-Propagation [BM01a, Noe00, MHI08, Shy06, Vol04b]. wavefield [BST03]. wavefields [BCR04]. waveform [CSMH05]. Wavefronts [RMO00, Che07]. wavefunctions [Boy04, NG06a]. Waveguide [PR01b, BBD04, FH03]. Waveguides [CdHST08, TB00a, FCJ08a]. wavelength [VS07]. Wavelet [FHLK05, HK02, PR01b, VB00, AKV06, BLG^{+}08, FH03, KDK^{+}07, MK08b, NG06a, VK05a]. Wavelet-Based [HK02, PR01b]. wavelet-MRA-based [BLG^{+}08]. Wavenumber [KLK08, TK00, CC04]. Wavenumber-extended [KLK08]. Waves [BST01, Bla00, Boy02b, DF00a, Gua00, MN02, OB02, PC02, SSC00, VCP00, Vay02, WPM02a, WC01, Wu02, AK06b, AM05, BAR08, BFJ03, BWLM09, BCZ04, Boy03, CF06a, CLMRP08, CS05, CDS04, Dur08, EV03, EKBL09, Fan08, FCT07, FG07, Gab07, GB08a, GN03, GP04, GS09d, HS09a, JY08, Kas07, LY07b, LM08a, LTD07, LP04b, MY06a, MLFG06, NB04, SB06b, SM05, Tsy03, VS09, Wan05, XG09, YM07, Yan08, vdVX07]. wavy [GMD03]. Way [FT01, FSY00, SFY01, WK01a, CC08b, GS09a]. ways [BZ09]. Weak [AGP01, BMQS02, DF00a, HMK02, KB01, PKvdB00, CP03a, HMM07, KT03]. weakform [LNXNTX09]. weakly [LMX^{+}08, SE04]. weather [Lyn08, MS08b, SK08b, SW08c]. wedge [ODAF06]. weight [MBS03]. Weighted [AL01, Azm02, BS00a, BBK07, DZ00, MS01, SK08a, SM04, WC01, Yus06, BCCD08, CB07, HAP05, KLLJ09, LCW04, NF09, TWM07, ZSWW03, ZWS06, ZJS08]. Weighted-Difference [Azm02]. Weighting [Ver01]. Weights [SHS02]. Well [BES07, BKLL04, LMNK07, NPPN06, Xu02b, AB05b, GPC07, ILL09, Meh04, NXS07, Rah04, RF06, WSYS09, XS06]. Well-Balanced [Xu02b, BES07, LMNK07, NPPN06, AB05b, GPC07, NXS07, RF06, WSYS09, XS06]. well-conditioned [ILL09]. well-posed [Meh04, Rah04]. wells [JL09]. Wendroff [LCS09]. WENO [Bal09, BRDM09, BK07, CVB06, Cap08a, Cap08b, CGMS03, CGMS06, CHB09, CS06, CS07d, CD07, CZVS04, GS06a, GR04, HP04b, HLY09, JD09, JC06b, KK05b, LSD07, LBL07, MTWW06, NJX08a, NXS07, Pir02, QS02, QS04, QS05, RLZ03, SHS02, SZS03, TT04, TB04, VS02, VCZS04, XS05a, XS06, XS05c, XLS09a, YC09a, YC09b, ZJ09, ZQSD08, GSV09]. WENO-based [LBL07]. WENO-Boltzmann [CGMS06]. WENO-solver [CGMS03]. WENO-type [XLS09a]. wetting [Gla05, SHTB09, YZ07]. which [IG05]. whistlers [LJM^{+}06]. Whitham [ZSWW03, Boy03]. Wide [FSY00, GST00]. Wide-Angle [FSY00]. wideband [CCG^{+}06]. Wiener [HLRZ06, LNGK04, LKNG04, MT04]. Wiener-type [LNGK04]. WIGGLE [LPK05]. Wigner [KLW09, RRC05]. Wind [STiST02, SSW^{+}07]. windowed [SZLW06]. windowing [SAK05]. Winds [LR01b]. Wire [BISS01, DDF01, LWW04]. Wire-Plate [BISS01, LWW04]. Wise [YL01, CBH03, RLZ03]. within [AKV00, AJ09, Bae03, FCD^{+}06, KG08, SS07a]. without [ABRR09a, ABRR09b, BIVC07, Edw06, Giv01, JP03, Kas07, KDC05, Li08a, Mon00, SJ02, TB00b, YGL05, ZSW07]. WKB [BP06, GM06]. Woollings [TR07]. Work [Mac00, LC06b]. Worst [PWW00].

X [RR07]. X-ray [RR07]. XTOR [LL08b].

Yee [LW01, MT01, TE08].

Z [JD09]. Zakharov [BSW03, JMZ04]. Zernike [KP04]. Zero [SBGK00, KKM08]. zero-Mach [KKM08]. Zonal [BDS07]. zones [BAMD07, Bod06]. zoom [KP07].

References

Aktas:2002:CCD

[AA02] Ozgur Aktas and N. R. Aluru. A combined continuum/ DSMC technique for multiscale analysis of microfluidic filters. Journal of Computational Physics, 178(2):342–372, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

               2716 (electronic). URL http://www.sciencedirect.com/
               science/article/pii/S0021999102970300.

Amstutz:2006:NAT

[AA06] Samuel Amstutz and Heiko Andrä. A new algorithm for topology optimization using a level-set method. Journal of Computational Physics, 216(2):573–588, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005656.

Agarwal:2007:SLA

[AA07] Nitin Agarwal and N. R. Aluru. A stochastic Lagrangian approach for geometrical uncertainties in electrostatics. Journal of Computational Physics, 226(1):156–179, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001489.

Agarwal:2009:DAS

[AA09] Nitin Agarwal and N. R. Aluru. A domain adaptive stochastic collocation approach for analysis of MEMS under uncertainties. Journal of Computational Physics, 228(20):7662–7688, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003866.

Ahusborde:2007:PFH

[AAC07] Etienne Ahusborde, Mejdi Azaiez, and Jean-Paul Caltagirone. A primal formulation for the Helmholtz decomposition. Journal of Computational Physics, 225(1):13–19, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001532.

Antoine:2003:USD

[AB03] X. Antoine and C. Besse. Unconditionally stable discretization schemes of non-reflecting boundary conditions for the one-dimensional Schrödinger equation. Journal of Computational Physics, 188(1):157–175, June 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001591.

Andersson:2005:FGT

[AB05a] Fredrik Andersson and Gregory Beylkin. The fast Gauss transform with complex parameters. Journal of Computational Physics, 203(1):274–286, February 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003274.

Audusse:2005:WBP

[AB05b] Emmanuel Audusse and Marie-Odile Bristeau. A wellbalanced positivity preserving “second-order” scheme for shallow water flows on unstructured meshes. Journal of Computational Physics, 206(1):311–333, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005157.

Ambati:2007:STD

[AB07] V. R. Ambati and O. Bokhove. Space–time discontinuous Galerkin discretization of rotating shallow water equations. Journal of Computational Physics, 225(2):1233–1261, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700037X.

Abgrall:2001:TUC

[Abg01] R. Abgrall. Toward the ultimate conservative scheme: Following the quest. Journal of Computational Physics, 167 (2):277–315, March 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100966725.

Abgrall:2006:ENO

[Abg06] R. Abgrall. Essentially non-oscillatory residual distribution schemes for hyperbolic problems. Journal of Computational Physics, 214(2):773–808, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004730.

Alpert:2002:ASP

[ABGV02] B. Alpert, G. Beylkin, D. Gines, and L. Vozovoi. Adaptive solution of partial differential equations in multiwavelet bases. Journal of Computational Physics, 182(1):149–190, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971603.

Adams:2003:PMS

[ABHT03] Mark Adams, Marian Brezina, Jonathan Hu, and Ray Tuminaro. Parallel multigrid smoothing: polynomial versus Gauss– Seidel. Journal of Computational Physics, 188(2):593–610, July 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103001943.

Antoine:2009:ABC

[ABK09] Xavier Antoine, Christophe Besse, and Pauline Klein. Absorbing boundary conditions for the one-dimensional Schrödinger equation with an exterior repulsive potential. Journal of Computational Physics, 228(2):312–335, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004804.

Acebron:2005:PID

[ABLS05] Juan A. Acebrón, Maria Pia Busico, Piero Lanucara, and Renato Spigler. Probabilistically induced domain decomposition methods for elliptic boundary-value problems. Journal of Computational Physics, 210(2):421–438, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002263.

Abramov:2006:PCF

[Abr06] Rafail Abramov. A practical computational framework for the multidimensional moment-constrained maximum entropy principle. Journal of Computational Physics, 211(1):198–209, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002688.

Abramov:2007:IAM

[Abr07] Rafail V. Abramov. An improved algorithm for the multidimensional moment-constrained maximum entropy problem. Journal of Computational Physics, 226(1):621–644, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107001994.

Abramov:2009:MMC

[Abr09]  Rafail V. Abramov.  The multidimensional moment-

constrained maximum entropy problem: a BFGS algorithm with constraint scaling. Journal of Computational Physics, 228(1):96–108, January 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004555.

Akhmatskaya:2009:CGH

[ABRR09a] Elena Akhmatskaya, Nawaf Bou-Rabee, and Sebastian Reich. A comparison of generalized hybrid Monte Carlo methods with and without momentum flip. Journal of Computational Physics, 228(6):2256–2265, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006323. See erratum [ABRR09b].

Akhmatskaya:2009:ECG

[ABRR09b] Elena Akhmatskaya, Nawaf Bou-Rabee, and Sebastian Reich. Erratum to “A comparison of generalized hybrid Monte Carlo methods with and without momentum flip” [J. Comput. Phys. 228 (2009) 2256–2265]. Journal of Computational Physics, 228(19):7492–7496, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003477. See [ABRR09a].

Axner:2008:PEP

[ABZ^{+}08] L. Axner, J. Bernsdorf, T. Zeiser, P. Lammers, J. Linxweiler, and A. G. Hoekstra. Performance evaluation of a parallel sparse lattice Boltzmann solver. Journal of Computational Physics, 227(10):4895–4911, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800051X.

Abarbanel:2000:SSHa

[AC00] Saul S. Abarbanel and Alina E. Chertock. Strict stability of high-order compact implicit finite-difference schemes: The role of boundary conditions for hyperbolic PDEs, I. Journal of Computational Physics, 160(1):42–66, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964209.

Alexander:2001:AMC

[AC01] S. A. Alexander and R. L. Coldwell. Atomic and molecular calculations using quasirandom numbers. Journal of Computational Physics, 172(2):908–916, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968652.

Anderson:2005:FWM

[AC05] Christopher R. Anderson and Thomas C. Cecil. A Fourier– Wachspress method for solving Helmholtz’s equation in three-dimensional layered domains. Journal of Computational Physics, 205(2):706–718, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004875.

Appelo:2009:HOS

[AC09] Daniel Appelö and Tim Colonius. A high-order super-grid-scale absorbing layer and its application to linear hyperbolic systems. Journal of Computational Physics, 228(11): 4200–4217, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001089.

Adrover:2007:SBD

[ACGV07] A. Adrover, F. Creta, M. Giona, and M. Valorani. Stretching-based diagnostics and reduction of chemical kinetic models with diffusion. Journal of Computational Physics, 225(2):1442–1471, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000629.

Allaire:2002:FEM

[ACK02] Grégoire Allaire, Sébastien Clerc, and Samuel Kokh. A five-equation model for the simulation of interfaces between compressible fluids. Journal of Computational Physics, 181(2):577–616, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971433.

Assous:2003:NST

[ACLS03] F. Assous, P. Ciarlet, Jr., S. Labrunie, and J. Segré. Numerical solution to the time-dependent Maxwell equations in axisymmetric singular domains: the singular complement method. Journal of Computational Physics, 191(1):147–176, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003097.

Antoine:2008:NAH

[ACR08] Xavier Antoine, Chokri Chniti, and Karim Ramdani. On the numerical approximation of high-frequency acoustic multiple scattering problems by circular cylinders. Journal of Computational Physics, 227(3):1754–1771, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004305.

Assous:2000:NST

[ACS00] F. Assous, P. Ciarlet, Jr., and J. Segré. Numerical solution to the time-dependent Maxwell equations in two-dimensional singular domains: The singular complement method. Journal of Computational Physics, 161(1):218–249, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964994.

Abarbanel:2000:SSHb

[ACY00] Saul S. Abarbanel, Alina E. Chertock, and Amir Yefet. Strict stability of high-order compact implicit finite-difference schemes: The role of boundary conditions for hyperbolic PDEs, II. Journal of Computational Physics, 160(1):67–87, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964210.

Aslanyan:2001:SVD

[AD01] A. Aslanyan and E. B. Davies. Separation of variables in deformed cylinders. Journal of Computational Physics, 174(1):327–344, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196913X.

Ajaev:2003:BIS

[AD03] Vladimir S. Ajaev and Stephen H. Davis. Boundary-integral simulations of containerless solidification. Journal of Computational Physics, 187(2):492–503, May 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001207.

Aprovitola:2004:ACU

[AD04] Andrea Aprovitola and Filippo Maria Denaro. On the application of congruent upwind discretizations for large eddy simulations. Journal of Computational Physics, 194(1):329– 343, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103004856.

Ainley:2008:MIR

[ADE^{+}08] Josephine Ainley, Sandra Durkin, Rafael Embid, Priya Boindala, and Ricardo Cortez. The method of images for regularized stokeslets. Journal of Computational Physics, 227(9):4600–4616, April 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000417.

Alvarez:2009:CRU

[ÁDIM09] Diego Álvarez, Oliver Dorn, Natalia Irishina, and Miguel Moscoso. Crack reconstruction using a level-set strategy. Journal of Computational Physics, 228(16):5710–5721, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002150.

Asvadurov:2000:ADG

[ADK00] Sergey Asvadurov, Vladimir Druskin, and Leonid Knizhnerman. Application of the difference Gaussian rules to solution of hyperbolic problems. Journal of Computational Physics, 158(1):116–135, February 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964100.

Asvadurov:2002:ADG

[ADK02] Sergey Asvadurov, Vladimir Druskin, and Leonid Knizhnerman. Application of the difference Gaussian rules to solution of hyperbolic problems: II. Global expansion. Journal of Computational Physics, 175(1):24–49, January 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968780.

Ardekani:2008:CMP

[ADR08] A. M. Ardekani, S. Dabiri, and R. H. Rangel. Collision of multi-particle and general shape objects in a viscous fluid. Journal of Computational Physics, 227(24):10094–10107, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004312.

Assous:2003:NMC

[ADS03] F. Assous, T. Pougeard Dulimbert, and J. Segré. A new method for coalescing particles in PIC codes. Journal of Computational Physics, 187(2):550–571, May 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001244.

Abdulle:2003:FDH

[AE03] Assyr Abdulle and Weinan E. Finite difference heterogeneous multi-scale method for homogenization problems. Journal of Computational Physics, 191(1):18–39, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003036.

Anderson:2004:ALE

[AEP04] R. W. Anderson, N. S. Elliott, and R. B. Pember. An arbitrary Lagrangian–Eulerian method with adaptive mesh refinement for the solution of the Euler equations. Journal of Computational Physics, 199(2):598–617, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400097X.

Alauzet:2007:TFP

[AFGM07] F. Alauzet, P. J. Frey, P. L. George, and B. Mohammadi. 3D transient fixed point mesh adaptation for time-dependent problems: Application to CFD simulations. Journal of Computational Physics, 222(2):592–623, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600369X.

Atassi:2008:INB

[AG08] Oliver V. Atassi and José M. Galán. Implementation of nonreflecting boundary conditions for the nonlinear Euler equations. Journal of Computational Physics, 227(3):1643–1662, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004202.

Antoine:2009:PRM

[AG09] Xavier Antoine and Christophe Geuzaine. Phase reduction models for improving the accuracy of the finite element solution of time-harmonic scattering problems I: General approach and low-order models. Journal of Computational Physics, 228(8):3114–3136, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000291.

Alpert:2000:IEF

[AGH00] Bradley Alpert, Leslie Greengard, and Thomas Hagstrom. An integral evolution formula for the wave equation. Journal of Computational Physics, 162(2):536–543, August 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965471.

Alpert:2002:NBC

[AGH02] Bradley Alpert, Leslie Greengard, and Thomas Hagstrom. Nonreflecting boundary conditions for the time-dependent wave equation. Journal of Computational Physics, 180(1): 270–296, July 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102970932.

Auteri:2001:RLC

[AGP01] F. Auteri, J.-L. Guermond, and N. Parolini. Role of the LBB condition in weak spectral projection methods. Journal of Computational Physics, 174(1):405–420, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969220.

Archibald:2009:DDM

[AGSX09] Rick Archibald, Anne Gelb, Rishu Saxena, and Dongbin Xiu. Discontinuity detection in multivariate space for stochastic simulations. Journal of Computational Physics, 228(7):2676–2689, April 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000060.

Alexander:2002:ARS

[AGT02] Francis J. Alexander, Alejandro L. Garcia, and Daniel M. Tartakovsky. Algorithm refinement for stochastic partial differential equations: I. Linear diffusion. Journal of Computational Physics, 182(1):47–66, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971494.

Alexander:2005:ARS

[AGT05] Francis J. Alexander, Alejandro L. Garcia, and Daniel M. Tartakovsky. Algorithm refinement for stochastic partial differential equations: II. Correlated systems. Journal of Computational Physics, 207(2):769–787, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000501.

Ayala:2007:HAS

[AGW07] Orlando Ayala, Wojciech W. Grabowski, and Lian-Ping Wang. A hybrid approach for simulating turbulent collisions of hydrodynamically-interacting particles. Journal of Computational Physics, 225(1):51–73, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005778.

Ahamadi:2008:LFE

[AH08] M. Ahamadi and O. G. Harlen. A Lagrangian finite element method for simulation of a suspension under planar extensional flow. Journal of Computational Physics, 227(16):7543–7560, August 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002477.

Adams:2004:ISS

[AHF04] N. A. Adams, S. Hickel, and S. Franz. Implicit subgridscale modeling by adaptive deconvolution. Journal of Computational Physics, 200(2):412–431, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001743.

Arienti:2003:LSA

[AHMS03] Marco Arienti, Patrick Hung, Eric Morano, and Joseph E. Shepherd. A level set approach to Eulerian–Lagrangian coupling. Journal of Computational Physics, 185(1):213– 251, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102000554.

Allampalli:2009:HAL

[AHNS09] Vasanth Allampalli, Ray Hixon, M. Nallasamy, and Scott D. Sawyer. High-accuracy large-step explicit Runge–Kutta (HALE–RK) schemes for computational aeroacoustics. Journal of Computational Physics, 228(10):3837–3850, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000849.

Airaksinen:2007:AMB

[AHPT07] Tuomas Airaksinen, Erkki Heikkola, Anssi Pennanen, and Jari Toivanen. An algebraic multigrid based shifted-Laplacian preconditioner for the Helmholtz equation. Journal of Computational Physics, 226(1):1196–1210, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002288.

Arian:2009:AHD

[AI09] Eyal Arian and Angelo Iollo. Analytic Hessian derivation for the quasi-one-dimensional Euler equations. Journal of Computational Physics, 228(2):476–490, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005068.

Ainsworth:2004:DDB

[Ain04] Mark Ainsworth. Dispersive and dissipative behaviour of high order discontinuous Galerkin finite element methods. Journal of Computational Physics, 198(1):106–130, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400021X.

Averbuch:2003:SEB

[AINR03] Amir Averbuch, Moshe Israeli, Menachem Nathan, and Igor Ravve. Surface evolution in bare bamboo-type metal lines under diffusion and electric field effects. Journal of Computational Physics, 188(2):640–677, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001992.

Averbuch:2003:EIV

[AIR03] Amir Averbuch, Moshe Israeli, and Igor Ravve. Electromigration of intergranular voids in metal films for microelectronic interconnects. Journal of Computational Physics, 186(2):481–502, April 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000706.

Averbuch:2001:CEI

[AIRY01] Amir Averbuch, Moshe Israeli, Igor Ravve, and Irad Yavneh. Computation for electromigration in interconnects of microelectronic devices. Journal of Computational Physics, 167(2):316–371, March 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966774.

Anand:2009:SME

[AJ09] Gaurav Anand and Patrick Jenny. Stochastic modeling of evaporating sprays within a consistent hybrid joint PDF framework. Journal of Computational Physics, 228(6):2063– 2081, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006153.

Akcelik:2001:NOT

[AJG01] Volkan Akcelik, Branislav Jaramaz, and Omar Ghattas. Nearly orthogonal two-dimensional grid generation with aspect ratio control. Journal of Computational Physics, 171(2):805–821, August 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968111.

Allaire:2004:SOU

[AJT04] Grégoire Allaire, François Jouve, and Anca-Maria Toader. Structural optimization using sensitivity analysis and a levelset method. Journal of Computational Physics, 194(1):363– 393, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910300487X.

Abgrall:2001:CCM

[AK01] Rémi Abgrall and Smadar Karni. Computations of compressible multifluids. Journal of Computational Physics, 169(2):594–623, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100966853.

Albensoeder:2005:ATD

[AK05] S. Albensoeder and H. C. Kuhlmann. Accurate threedimensional lid-driven cavity flow. Journal of Computational Physics, 206(2):536–558, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000033.

Ahn:2006:SCF

[AK06a] Hyung Taek Ahn and Yannis Kallinderis. Strongly coupled flow/structure interactions with a geometrically conservative ALE scheme on general hybrid meshes. Journal of Computational Physics, 219(2):671–696, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001987.

Appelo:2006:NAL

[AK06b] Daniel Appelö and Gunilla Kreiss. A new absorbing layer for elastic waves. Journal of Computational Physics, 215(2):642– 660, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005097.

Anderson:2007:NAS

[AK07] Matthew Anderson and Jung-Han Kimn. A numerical approach to space–time finite elements for the wave equation. Journal of Computational Physics, 226(1):466–476, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107001672.

Ahn:2009:ACM

[AK09] Jae Wan Ahn and Chongam Kim. An axisymmetric computational model of generalized hydrodynamic theory for rarefied multi-species gas flows. Journal of Computational Physics, 228(11):4088–4117, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900093X.

Anderson:2006:SVF

[AKH06] Patrick D. Anderson, Bert J. Keestra, and Martien A. Hulsen. On the streamfunction-vorticity formulation in sliding bi-period frames: Application to bulk behavior for polymer blends. Journal of Computational Physics, 212(1):268– 287, February 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105003207.

Akcelik:2008:SDD

[AKL^{+}08] Volkan Akçelik, Kwok Ko, Lie-Quan Lee, Zenghai Li, Cho-Kuen Ng, and Liling Xiao. Shape determination for deformed electromagnetic cavities. Journal of Computational Physics, 227(3):1722–1738, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004275.

Arun:2009:FVE

[AKLMP09] K. R. Arun, M. Kraft, M. Lukácová-Medvid’ová, and Phoolan Prasad. Finite volume evolution Galerkin method for hyperbolic conservation laws with spatially varying flux functions. Journal of Computational Physics, 228(2):565–590, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005317.

Audusse:2009:CDC

[AKO09] E. Audusse, R. Klein, and A. Owinoh. Conservative discretization of Coriolis force in a finite volume framework. Journal of Computational Physics, 228(8):2934–2950, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000102.

Atzberger:2007:SIB

[AKP07] Paul J. Atzberger, Peter R. Kramer, and Charles S. Peskin. A stochastic immersed boundary method for fluid-structure dynamics at microscopic length scales. Journal of Computational Physics, 224(2):1255–1292, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005766.

Abrashkevich:2000:NMS

[AKV00] A. G. Abrashkevich, M. S. Kaschiev, and S. I. Vinitsky. A new method for solving an eigenvalue problem for a system of three Coulomb particles within the hyperspherical adiabatic representation. Journal of Computational Physics, 163(2):328–348, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965628.

Alam:2006:SST

[AKV06] Jahrul M. Alam, Nicholas K.-R. Kevlahan, and Oleg V. Vasilyev. Simultaneous space–time adaptive wavelet solution of nonlinear parabolic differential equations. Journal of Computational Physics, 214(2):829–857, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004754.

Adler:2001:CNT

[AKY01] Carl Adler, Ronald Kneusel, and William Younger. Chaos, number theory, and computers. Journal of Computational Physics, 166(1):165–172, January 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101966458.

Anistratov:2001:NLW

[AL01] Dmitriy Y. Anistratov and Edward W. Larsen. Nonlinear and linear α-weighted methods for particle transport problems. Journal of Computational Physics, 173(2):664–684, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969050.

An:2006:FAB

[AL06] Xiang An and Zhi-Qing Lü. A fast algorithm based on partial basic solution vectors domain decomposition method for scattering analysis of electrically large cylinders. Journal of Computational Physics, 219(2):930–942, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003378.

Asinari:2008:CLB

[AL08] Pietro Asinari and Li-Shi Luo. A consistent lattice Boltzmann equation with baroclinic coupling for mixtures. Journal of Computational Physics, 227(8):3878–3895, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005396.

Albers:2000:LMR

[Alb00] Michael Albers. A local mesh refinement multigrid method for 3-D convection problems with strongly variable viscosity. Journal of Computational Physics, 160(1):126–150, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964386.

Alben:2008:IMC

[Alb08] Silas Alben. An implicit method for coupled flow-body dynamics. Journal of Computational Physics, 227(10):4912– 4933, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108000521.

Alben:2009:SDF

[Alb09] Silas Alben. Simulating the dynamics of flexible bodies and vortex sheets. Journal of Computational Physics, 228 (7):2587–2603, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006591.

Arber:2001:SGL

[ALGM01] T. D. Arber, A. W. Longbottom, C. L. Gerrard, and A. M. Milne. A staggered grid, Lagrangian–Eulerian remap code for 3-D MHD simulations. Journal of Computational Physics, 171(1):151–181, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967804.

Anderson:2008:GPM

[ALT08] Joshua A. Anderson, Chris D. Lorenz, and A. Travesset. General purpose molecular dynamics simulations fully implemented on graphics processing units. Journal of Computational Physics, 227(10):5342–5359, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000818.

Abgrall:2003:CSO

[AM03] Rémi Abgrall and Mohamed Mezine. Construction of second order accurate monotone and stable residual distribution schemes for unsteady flow problems. Journal of Computational Physics, 188(1):16–55, June 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000846.

Abgrall:2004:CSO

[AM04] Rémi Abgrall and Mohamed Mezine. Construction of second-order accurate monotone and stable residual distribution schemes for steady problems. Journal of Computational Physics, 195(2):474–507, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005400.

Aslan:2005:NSI

[AM05] Necdet Aslan and Michael Mond. A numerical scheme for ionizing shock waves. Journal of Computational Physics, 210(2):401–420, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002251.

Al-Mohssen:2004:APC

[AMH04] Husain A. Al-Mohssen and Nicolas G. Hadjiconstantinou. Arbitrary-pressure chemical vapor deposition modeling using direct simulation Monte Carlo with nonlinear surface chemistry. Journal of Computational Physics, 198(2):617–627, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104000543.

Aubry:2008:DPC

[AMLC08] Romain Aubry, Fernando Mut, Rainald Löhner, and Juan R. Cebral. Deflated preconditioned conjugate gradient solvers for the pressure-Poisson equation. Journal of Computational Physics, 227(24):10196–10208, December 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004518.

Apte:2009:NMF

[AMP09] Sourabh V. Apte, Mathieu Martin, and Neelesh A. Patankar. A numerical method for fully resolved simulation (FRS) of rigid particle-flow interactions in complex flows. Journal of Computational Physics, 228(8):2712–2738, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006268.

Alonso-Mallo:2006:HOF

[AMR06] Isaías Alonso-Mallo and Nuria Reguera. A high order finite element discretization with local absorbing boundary conditions of the linear Schrödinger equation. Journal of Computational Physics, 220(1):409–421, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002488.

Aulisa:2003:MMV

[AMS03] Eugenio Aulisa, Sandro Manservisi, and Ruben Scardovelli. A mixed markers and volume-of-fluid method for the reconstruction and advection of interfaces in twophase and free-boundary flows. Journal of Computational Physics, 188(2):611–639, July 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001967.

Aulisa:2004:SMA

[AMS04] Eugenio Aulisa, Sandro Manservisi, and Ruben Scardovelli. A surface marker algorithm coupled to an area-preserving marker redistribution method for three-dimensional interface tracking. Journal of Computational Physics, 197(2):555–584, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006521.

Aulisa:2003:GAP

[AMSZ03] Eugenio Aulisa, Sandro Manservisi, Ruben Scardovelli, and Stephane Zaleski. A geometrical area-preserving volumeof-fluid advection method. Journal of Computational Physics, 192(1):355–364, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003759.

Aulisa:2007:IRL

[AMSZ07] E. Aulisa, S. Manservisi, R. Scardovelli, and S. Zaleski. Interface reconstruction with least-squares fit and split advection in three-dimensional Cartesian geometry. Journal of Computational Physics, 225(2):2301–2319, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001325.

An:2009:CNI

[AMXL09] Heng-Bin An, Ze-Yao Mo, Xiao-Wen Xu, and Xu Liu. On choosing a nonlinear initial iterate for solving the 2-D 3-T heat conduction equations. Journal of Computational Physics, 228(9):3268–3287, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000345.

Anderson:2009:ESS

[And09] Christopher R. Anderson. Efficient solution of the Schroedinger– Poisson equations in layered semiconductor devices. Journal of Computational Physics, 228(13):4745–4756, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001557.

Anonymous:2000:APAc

[Ano00a] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 158(1):136–137, February 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964416.

Anonymous:2000:APAd

[Ano00b] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 158(2):263–264, March 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096457X.

Anonymous:2000:APAe

[Ano00c] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 159(1):123–124, March 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964672.

Anonymous:2000:APAf

[Ano00d] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 159(2):456–457, April 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964933.

Anonymous:2000:APAg

[Ano00e] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 160(1):401–402, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965161.

Anonymous:2000:APAh

[Ano00f] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 160(2):783–784, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965252.

Anonymous:2000:APAi

[Ano00g] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 161(1):376–377, June 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965355.

Anonymous:2000:APAj

[Ano00h] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 161(2):728–729, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965392.

Anonymous:2000:APAk

[Ano00i] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 162(1):299–300, July 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965586.

Anonymous:2000:APAl

[Ano00j] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 162(2):544–545, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965781.

Anonymous:2000:APAm

[Ano00k] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 163(1):268–269, September 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965896.

Anonymous:2000:APAn

[Ano00l] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 163(2):548–549, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966075.

Anonymous:2000:APAo

[Ano00m] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 164(1):238–239, October 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966191.

Anonymous:2000:APAp

[Ano00n] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 164(2):429–430, November 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966282.

Anonymous:2000:APAq

[Ano00o] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 165(1):307–308, November 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966518.

Anonymous:2000:APAr

[Ano00p] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 165(2):771–772, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966609.

Anonymous:2000:APAa

[Ano00q] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 157(1):417–418, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199964124.

Anonymous:2000:APAb

[Ano00r] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 157(2):801–802, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919996429X.

Anonymous:2000:AIVa

[Ano00s] Anonymous. Author index for volume 157. Journal of Computational Physics, 157(2):803, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964306.

Anonymous:2000:AIVb

[Ano00t] Anonymous. Author index for volume 158. Journal of Computational Physics, 158(2):265, March 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964581.

Anonymous:2000:AIVc

[Ano00u] Anonymous. Author index for volume 159. Journal of Computational Physics, 159(2):458, April 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964945.

Anonymous:2000:AIVd

[Ano00v] Anonymous. Author index for volume 160. Journal of Computational Physics, 160(2):785, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965264.

Anonymous:2000:AIVe

[Ano00w] Anonymous. Author index for volume 161. Journal of Computational Physics, 161(2):730, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965409.

Anonymous:2000:AIVf

[Ano00x] Anonymous. Author index for volume 162. Journal of Computational Physics, 162(2):546, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965793.

Anonymous:2000:AIVg

[Ano00y] Anonymous. Author index for volume 163. Journal of Computational Physics, 163(2):550, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966087.

Anonymous:2000:AIVh

[Ano00z] Anonymous. Author index for volume 164. Journal of Computational Physics, 164(2):431, November 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966294.

Anonymous:2000:AIVi

[Ano00-27] Anonymous. Author index for volume 165. Journal of Computational Physics, 165(2):773, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966646.

Anonymous:2000:CAT

[Ano00-28] Anonymous. Cumulative author–title index 1995–2000. Journal of Computational Physics, 165(2):774–809, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966749.

Anonymous:2000:E

[Ano00-29] Anonymous. Editorial. Journal of Computational Physics, 157(1):1, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199964112.

Anonymous:2001:APAa

[Ano01a] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 166(1):186–187, January 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101966938.

Anonymous:2001:APAb

[Ano01b] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 166(2):425–426, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967026.

Anonymous:2001:APAc

[Ano01c] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 167(1):244–245, February 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967166.

Anonymous:2001:APAd

[Ano01d] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 167(2):502–503, March 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967294.

Anonymous:2001:APAe

[Ano01e] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 168(1):265, March 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967427.

Anonymous:2001:APAf

[Ano01f] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 168(2):527–528, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967671.

Anonymous:2001:APAg

[Ano01g] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 169(1):246–247, May 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967828.

Anonymous:2001:APAh

[Ano01h] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 169(2):760–761, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968068.

Anonymous:2001:APAi

[Ano01i] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 170(1):457–458, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968330.

Anonymous:2001:APAj

[Ano01j] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 170(2):917–918, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968391.

Anonymous:2001:APAk

[Ano01k] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 171(1):423–424, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968524.

Anonymous:2001:APAl

[Ano01l] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 171(2):851–852, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968688.

Anonymous:2001:APAm

[Ano01m] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 172(1):399–400, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196886X.

Anonymous:2001:APAn

[Ano01n] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 172(2):924–925, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969207.

Anonymous:2001:APAo

[Ano01o] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 173(1):391–392, October 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969281.

Anonymous:2001:APAp

[Ano01p] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 173(2):797–798, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969499.

Anonymous:2001:APAq

[Ano01q] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 174(1):489–490, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969669.

Anonymous:2001:APAr

[Ano01r] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 174(2):947–948, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969803.

Anonymous:2001:AIVa

[Ano01s] Anonymous. Author index for volume 166. Journal of Computational Physics, 166(2):427, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967038.

Anonymous:2001:AIVb

[Ano01t] Anonymous. Author index for volume 167. Journal of Computational Physics, 167(2):504, March 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967300.

Anonymous:2001:AIVc

[Ano01u] Anonymous. Author index for volume 168. Journal of Computational Physics, 168(2):529, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967683.

Anonymous:2001:AIVd

[Ano01v] Anonymous. Author index for volume 169. Journal of Computational Physics, 169(2):762, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196807X.

Anonymous:2001:AIVe

[Ano01w] Anonymous. Author index for volume 170. Journal of Computational Physics, 170(2):921, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196841X.

Anonymous:2001:AIVf

[Ano01x] Anonymous. Author index for volume 171. Journal of Computational Physics, 171(2):853, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196869X.

Anonymous:2001:AIVg

[Ano01y] Anonymous. Author index for volume 172. Journal of Computational Physics, 172(2):926, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969219.

Anonymous:2001:AIVh

[Ano01z] Anonymous. Author index for volume 173. Journal of Computational Physics, 173(2):799, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969505.

Anonymous:2001:AIVi

[Ano01-27] Anonymous. Author index for volume 174. Journal of Computational Physics, 174(2):949, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969815.

Anonymous:2001:CAT

[Ano01-28] Anonymous. Cumulative author–title index 1996–2001. Journal of Computational Physics, 174(2):950–986, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969827.

Anonymous:2001:E

[Ano01-29] Anonymous. Editorial. Journal of Computational Physics, 174(2):491, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101969797.

Anonymous:2002:APAa

[Ano02a] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 175(1):361–362, January 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969839.

Anonymous:2002:APAb

[Ano02b] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 175(2):793–794, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969888.

Anonymous:2002:APAc

[Ano02c] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 176(1):228–229, February 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101970007.

Anonymous:2002:APAd

[Ano02d] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 176(2):507–508, March 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101970196.

Anonymous:2002:APAe

[Ano02e] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 177(1):206–207, March 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101970329.

Anonymous:2002:APAf

[Ano02f] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 177(2):451–452, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970452.

Anonymous:2002:APAg

[Ano02g] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 178(1):261–262, May 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970555.

Anonymous:2002:APAh

[Ano02h] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 178(2):563–564, May 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970695.

Anonymous:2002:APAi

[Ano02i] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 179(1):346–347, June 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971007.

Anonymous:2002:APAj

[Ano02j] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 179(2):736–737, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971147.

Anonymous:2002:APAk

[Ano02k] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 180(1):404–405, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971342.

Anonymous:2002:APAl

[Ano02l] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 180(2):775–776, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971561.

Anonymous:2002:APAm

[Ano02m] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 181(1):395–396, September 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971731.

Anonymous:2002:APAn

[Ano02n] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 181(2):760–761, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971810.

Anonymous:2002:APAo

[Ano02o] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 182(1):355–356, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972086.

Anonymous:2002:APAp

[Ano02p] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 182(2):651–652, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972438.

Anonymous:2002:APAq

[Ano02q] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 183(1):332–333, November 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972116.

Anonymous:2002:APAr

[Ano02r] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 183(2):676, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972128.

Anonymous:2002:AIVa

[Ano02s] Anonymous. Author index for volume 175. Journal of Computational Physics, 175(2):795, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196989X.

Anonymous:2002:AIVb

[Ano02t] Anonymous. Author index for volume 176. Journal of Computational Physics, 176(2):509, March 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101970202.

Anonymous:2002:AIVc

[Ano02u] Anonymous. Author index for volume 177. Journal of Computational Physics, 177(2):453, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970464.

Anonymous:2002:AIVd

[Ano02v] Anonymous. Author index for volume 178. Journal of Computational Physics, 178(2):565, May 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970701.

Anonymous:2002:AIVe

[Ano02w] Anonymous. Author index for volume 179. Journal of Computational Physics, 179(2):738, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971159.

Anonymous:2002:AIVf

[Ano02x] Anonymous. Author index for volume 180. Journal of Computational Physics, 180(2):777, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971573.

Anonymous:2002:AIVg

[Ano02y] Anonymous. Author index for volume 181. Journal of Computational Physics, 181(2):762, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971822.

Anonymous:2002:AIVh

[Ano02z] Anonymous. Author index for volume 182. Journal of Computational Physics, 182(2):653, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297244X.

Anonymous:2002:AIVi

[Ano02-27] Anonymous. Author index for volume 183. Journal of Computational Physics, 183(2):677, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297213X.

Anonymous:2002:CAT

[Ano02-28] Anonymous. Cumulative author–title index 1997–2002. Journal of Computational Physics, 183(2):678–715, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102972141.

Anonymous:2002:E

[Ano02-29] Anonymous. Editorial. Journal of Computational Physics, 181(2):397, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971883.

Anonymous:2003:EBa

[Ano03a] Anonymous. Editorial Board. Journal of Computational Physics, 184(1):??, January 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000712.

Anonymous:2003:EBb

[Ano03b] Anonymous. Editorial Board. Journal of Computational Physics, 185(1):??, February 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000354.

Anonymous:2003:EBc

[Ano03c] Anonymous. Editorial Board. Journal of Computational Physics, 185(2):??, March 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000536.

Anonymous:2003:EBd

[Ano03d] Anonymous. Editorial Board. Journal of Computational Physics, 186(1):??, March 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001062.

Anonymous:2003:EBe

[Ano03e] Anonymous. Editorial Board. Journal of Computational Physics, 187(1):??, May 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001724.

Anonymous:2003:EBf

[Ano03f] Anonymous. Editorial Board. Journal of Computational Physics, 188(1):??, June 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002390.

Anonymous:2003:EBg

[Ano03g] Anonymous. Editorial Board. Journal of Computational Physics, 189(1):??, July 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002869.

Anonymous:2003:EBh

[Ano03h] Anonymous. Editorial Board. Journal of Computational Physics, 190(1):??, September 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003577.

Anonymous:2003:EBi

[Ano03i] Anonymous. Editorial Board. Journal of Computational Physics, 190(2):??, September 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003814.

Anonymous:2003:EBj

[Ano03j] Anonymous. Editorial Board. Journal of Computational Physics, 191(1):??, October 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300442X.

Anonymous:2003:EBk

[Ano03k] Anonymous. Editorial Board. Journal of Computational Physics, 191(2):??, November 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004716.

Anonymous:2003:EBl

[Ano03l] Anonymous. Editorial board. Journal of Computational Physics, 192(1):??, November 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005205.

Anonymous:2003:EBm

[Ano03m] Anonymous. Editorial Board. Journal of Computational Physics, 192(2):??, December 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005424.

Anonymous:2003:IEBa

[Ano03n] Anonymous. IFC (Editorial Board). Journal of Computational Physics, 184(2):??, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000159.

Anonymous:2003:IEBb

[Ano03o] Anonymous. IFC (Editorial Board). Journal of Computational Physics, 187(2):??, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300216X.

Anonymous:2003:IEBc

[Ano03p] Anonymous. IFC(Editorial Board). Journal of Computational Physics, 189(2):??, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003334.

Anonymous:2003:IAa

[Ano03q] Anonymous. Instructions to authors. Journal of Computational Physics, 185(2):??, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000573.

Anonymous:2003:IAb

[Ano03r] Anonymous. Instructions to authors. Journal of Computational Physics, 186(2):??, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001438.

Anonymous:2003:IAc

[Ano03s] Anonymous. Instructions to authors. Journal of Computational Physics, 187(1):??, May 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300175X.

Anonymous:2003:IAd

[Ano03t] Anonymous. Instructions to authors. Journal of Computational Physics, 188(1):??, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002420.

Anonymous:2003:IAe

[Ano03u] Anonymous. Instructions to authors. Journal of Computational Physics, 189(1):??, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002894.

Anonymous:2003:IAf

[Ano03v] Anonymous. Instructions to authors. Journal of Computational Physics, 190(1):??, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003607.

Anonymous:2003:IAg

[Ano03w] Anonymous. Instructions to authors. Journal of Computational Physics, 191(1):??, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004455.

Anonymous:2003:IAh

[Ano03x] Anonymous. Instructions to authors. Journal of Computational Physics, 192(1):??, November 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005230.

Anonymous:2003:PNa

[Ano03y] Anonymous. Publisher’s note. Journal of Computational Physics, 184(1):v, January 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000785.

Anonymous:2003:PNb

[Ano03z] Anonymous. Publisher’s note. Journal of Computational Physics, 191(1):v, October 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004492.

Anonymous:2003:VAIa

[Ano03-27] Anonymous. Volume author index. Journal of Computational Physics, 184(2):679, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000184.

Anonymous:2003:VAIb

[Ano03-28] Anonymous. Volume author index. Journal of Computational Physics, 185(2):628, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000561.

Anonymous:2003:VAIc

[Ano03-29] Anonymous. Volume author index. Journal of Computational Physics, 186(2):704–705, April 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001426.

Anonymous:2003:VAId

[Ano03-30] Anonymous. Volume author index. Journal of Computational Physics, 187(2):716–717, May 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002195.

Anonymous:2003:VAIe

[Ano03-31] Anonymous. Volume author index. Journal of Computational Physics, 188(2):678–679, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002602.

Anonymous:2003:VAIf

[Ano03-32] Anonymous. Volume author index. Journal of Computational Physics, 189(2):676, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300336X.

Anonymous:2003:VAIg

[Ano03-33] Anonymous. Volume author index. Journal of Computational Physics, 190(2):682, September 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300384X.

Anonymous:2003:VAIh

[Ano03-34] Anonymous. Volume author index. Journal of Computational Physics, 191(2):670, November 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004741.

Anonymous:2003:VAIi

[Ano03-35] Anonymous. Volume author index. Journal of Computational Physics, 192(2):732–733, December 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300545X.

Anonymous:2004:EBa

[Ano04a] Anonymous. Editorial board. Journal of Computational Physics, 193(1):??, January 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005552.

Anonymous:2004:EBb

[Ano04b] Anonymous. Editorial board. Journal of Computational Physics, 193(2):??, January 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300634X.

Anonymous:2004:EBc

[Ano04c] Anonymous. Editorial board. Journal of Computational Physics, 194(1):??, February 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000051.

Anonymous:2004:EBd

[Ano04d] Anonymous. Editorial board. Journal of Computational Physics, 194(2):??, March 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000403.

Anonymous:2004:EBe

[Ano04e] Anonymous. Editorial board. Journal of Computational Physics, 195(1):??, March 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000695.

Anonymous:2004:EBf

[Ano04f] Anonymous. Editorial board. Journal of Computational Physics, 196(1):??, May 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001172.

Anonymous:2004:EBg

[Ano04g] Anonymous. Editorial board. Journal of Computational Physics, 196(2):??, May 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400138X.

Anonymous:2004:EBh

[Ano04h] Anonymous. Editorial board. Journal of Computational Physics, 197(1):??, June 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400155X.

Anonymous:2004:EBi

[Ano04i] Anonymous. Editorial board. Journal of Computational Physics, 197(2):??, July 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001949.

Anonymous:2004:EBj

[Ano04j] Anonymous. Editorial board. Journal of Computational Physics, 198(1):??, July 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002189.

Anonymous:2004:EBk

[Ano04k] Anonymous. Editorial board. Journal of Computational Physics, 198(2):??, August 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002463.

Anonymous:2004:EBl

[Ano04l] Anonymous. Editorial board. Journal of Computational Physics, 199(1):??, September 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002803.

Anonymous:2004:EBm

[Ano04m] Anonymous. Editorial board. Journal of Computational Physics, 200(1):??, October 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003353.

Anonymous:2004:EBn

[Ano04n] Anonymous. Editorial board. Journal of Computational Physics, 200(2):??, November 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003444.

Anonymous:2004:EBo

[Ano04o] Anonymous. Editorial board. Journal of Computational Physics, 201(1):??, November 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003535.

Anonymous:2004:EBp

[Ano04p] Anonymous. Editorial board. Journal of Computational Physics, 201(2):??, December 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003651.

Anonymous:2004:IAa

[Ano04q] Anonymous. Instructions to authors. Journal of Computational Physics, 193(1):??, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005588.

Anonymous:2004:IAb

[Ano04r] Anonymous. Instructions to authors. Journal of Computational Physics, 194(1):??, February 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000087.

Anonymous:2004:IAc

[Ano04s] Anonymous. Instructions to authors. Journal of Computational Physics, 195(1):??, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000725.

Anonymous:2004:IAd

[Ano04t] Anonymous. Instructions to authors. Journal of Computational Physics, 196(1):??, May 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001202.

Anonymous:2004:IAe

[Ano04u] Anonymous. Instructions to authors. Journal of Computational Physics, 197(1):??, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001585.

Anonymous:2004:IAf

[Ano04v] Anonymous. Instructions to authors. Journal of Computational Physics, 198(1):??, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002219.

Anonymous:2004:IAg

[Ano04w] Anonymous. Instructions to authors. Journal of Computational Physics, 199(1):??, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002839.

Anonymous:2004:IAh

[Ano04x] Anonymous. Instructions to authors. Journal of Computational Physics, 200(1):??, October 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003389.

Anonymous:2004:IAi

[Ano04y] Anonymous. Instructions to authors. Journal of Computational Physics, 201(1):??, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003560.

Anonymous:2004:MNA

[Ano04z] Anonymous. Mathematical and numerical aspects of low Mach number flows. Journal of Computational Physics, 195(2):??, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910400107X.

Anonymous:2004:T

[Ano04-27] Anonymous. TSFP-4. Journal of Computational Physics, 198(1):??, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104002268.

Anonymous:2004:VAIa

[Ano04-28] Anonymous. Volume author index. Journal of Computational Physics, 193(2):739, January 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006375.

Anonymous:2004:VAIb

[Ano04-29] Anonymous. Volume author index. Journal of Computational Physics, 194(2):809–810, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000439.

Anonymous:2004:VAIc

[Ano04-30] Anonymous. Volume author index. Journal of Computational Physics, 195(2):804–805, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001032.

Anonymous:2004:VAId

[Ano04-31] Anonymous. Volume author index. Journal of Computational Physics, 196(2):773–774, May 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400141X.

Anonymous:2004:VAIe

[Ano04-32] Anonymous. Volume author index. Journal of Computational Physics, 197(2):781–782, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001974.

Anonymous:2004:VAIf

[Ano04-33] Anonymous. Volume author index. Journal of Computational Physics, 198(2):771–772, August 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002499.

Anonymous:2004:VAIg

[Ano04-34] Anonymous. Volume author index. Journal of Computational Physics, 199(2):809–810, September 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003067.

Anonymous:2004:VAIh

[Ano04-35] Anonymous. Volume author index. Journal of Computational Physics, 200(2):795–796, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400347X.

Anonymous:2004:VAIi

[Ano04-36] Anonymous. Volume author index. Journal of Computational Physics, 201(2):798–799, December 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003687.

Anonymous:2005:EBa

[Ano05a] Anonymous. Editorial board. Journal of Computational Physics, 202(1):??, January 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003778.

Anonymous:2005:EBb

[Ano05b] Anonymous. Editorial board. Journal of Computational Physics, 202(2):??, January 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004401.

Anonymous:2005:EBc

[Ano05c] Anonymous. Editorial board. Journal of Computational Physics, 203(1):??, February 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004528.

Anonymous:2005:EBd

[Ano05d] Anonymous. Editorial board. Journal of Computational Physics, 203(2):??, March 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004978.

Anonymous:2005:EBe

[Ano05e] Anonymous. Editorial board. Journal of Computational Physics, 204(1):??, March 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000550.

Anonymous:2005:EBf

[Ano05f] Anonymous. Editorial board. Journal of Computational Physics, 204(2):??, April 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000628.

Anonymous:2005:EBg

[Ano05g] Anonymous. Editorial board. Journal of Computational Physics, 205(1):??, May 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000719.

Anonymous:2005:EBh

[Ano05h] Anonymous. Editorial board. Journal of Computational Physics, 205(2):??, May 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000975.

Anonymous:2005:EBi

[Ano05i] Anonymous. Editorial board. Journal of Computational Physics, 206(1):??, June 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001427.

Anonymous:2005:EBj

[Ano05j] Anonymous. Editorial board. Journal of Computational Physics, 206(2):??, July 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001518.

Anonymous:2005:EBk

[Ano05k] Anonymous. Editorial board. Journal of Computational Physics, 207(1):??, July 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001610.

Anonymous:2005:EBl

[Ano05l] Anonymous. Editorial board. Journal of Computational Physics, 207(2):??, August 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001683.

Anonymous:2005:EBm

[Ano05m] Anonymous. Editorial board. Journal of Computational Physics, 208(1):??, September 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002123.

Anonymous:2005:EBn

[Ano05n] Anonymous. Editorial board. Journal of Computational Physics, 208(2):??, September 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002305.

Anonymous:2005:EBo

[Ano05o] Anonymous. Editorial board. Journal of Computational Physics, 209(1):??, October 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500241X.

Anonymous:2005:EBp

[Ano05p] Anonymous. Editorial board. Journal of Computational Physics, 209(2):??, November 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002974.

Anonymous:2005:EBq

[Ano05q] Anonymous. Editorial board. Journal of Computational Physics, 210(1):??, November 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003359.

Anonymous:2005:EBr

[Ano05r] Anonymous. Editorial board. Journal of Computational Physics, 210(2):??, December 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003621.

Anonymous:2005:FAC

[Ano05s] Anonymous. First announcement and call for papers (HEFAT2005). Journal of Computational Physics, 203(2):??, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910500029X.

Anonymous:2005:IAa

[Ano05t] Anonymous. Instructions to authors. Journal of Computational Physics, 202(1):??, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003808.

Anonymous:2005:IAb

[Ano05u] Anonymous. Instructions to authors. Journal of Computational Physics, 203(1):??, February 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004553.

Anonymous:2005:IAc

[Ano05v] Anonymous. Instructions to authors. Journal of Computational Physics, 204(1):??, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000586.

Anonymous:2005:IAd

[Ano05w] Anonymous. Instructions to authors. Journal of Computational Physics, 205(1):??, May 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000744.

Anonymous:2005:IAe

[Ano05x] Anonymous. Instructions to authors. Journal of Computational Physics, 206(1):??, June 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001452.

Anonymous:2005:IAf

[Ano05y] Anonymous. Instructions to authors. Journal of Computational Physics, 207(1):??, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001646.

Anonymous:2005:IAg

[Ano05z] Anonymous. Instructions to authors. Journal of Computational Physics, 208(1):??, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002159.

Anonymous:2005:IAh

[Ano05-27] Anonymous. Instructions to authors. Journal of Computational Physics, 209(1):??, October 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002445.

Anonymous:2005:IAi

[Ano05-28] Anonymous. Instructions to authors. Journal of Computational Physics, 210(1):??, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003384.

Anonymous:2005:VAIa

[Ano05-29] Anonymous. Volume author index. Journal of Computational Physics, 202(2):794–795, January 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004437.

Anonymous:2005:VAIb

[Ano05-30] Anonymous. Volume author index. Journal of Computational Physics, 203(2):761–762, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005005.

Anonymous:2005:VAIc

[Ano05-31] Anonymous. Volume author index. Journal of Computational Physics, 204(2):805–806, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000653.

Anonymous:2005:VAId

[Ano05-32] Anonymous. Volume author index. Journal of Computational Physics, 205(2):776–777, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001002.

Anonymous:2005:VAIe

[Ano05-33] Anonymous. Volume author index. Journal of Computational Physics, 206(2):781–782, July 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001543.

Anonymous:2005:VAIf

[Ano05-34] Anonymous. Volume author index. Journal of Computational Physics, 207(2):788–789, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001713.

Anonymous:2005:VAIg

[Ano05-35] Anonymous. Volume author index. Journal of Computational Physics, 208(2):790–791, September 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002330.

Anonymous:2005:VAIh

[Ano05-36] Anonymous. Volume author index. Journal of Computational Physics, 209(2):796–797, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003001.

Anonymous:2005:VAIi

[Ano05-37] Anonymous. Volume author index. Journal of Computational Physics, 210(2):752–753, December 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003657.

Anonymous:2006:EBa

[Ano06a] Anonymous. Editorial board. Journal of Computational Physics, 211(1):??, January 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004006.

Anonymous:2006:EBb

[Ano06b] Anonymous. Editorial Board. Journal of Computational Physics, 211(2):??, January 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004389.

Anonymous:2006:EBc

[Ano06c] Anonymous. Editorial Board. Journal of Computational Physics, 212(1):??, February 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500519X.

Anonymous:2006:EBd

[Ano06d] Anonymous. Editorial Board. Journal of Computational Physics, 212(2):??, March 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005267.

Anonymous:2006:EBe

[Ano06e] Anonymous. Editorial Board. Journal of Computational Physics, 213(1):??, March 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000489.

Anonymous:2006:EBf

[Ano06f] Anonymous. Editorial Board. Journal of Computational Physics, 213(2):??, April 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000659.

Anonymous:2006:EBg

[Ano06g] Anonymous. Editorial Board. Journal of Computational Physics, 214(1):??, May 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000982.

Anonymous:2006:EBh

[Ano06h] Anonymous. Editorial Board. Journal of Computational Physics, 214(2):??, May 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001082.

Anonymous:2006:EBi

[Ano06i] Anonymous. Editorial Board. Journal of Computational Physics, 215(1):??, June 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001537.

Anonymous:2006:EBj

[Ano06j] Anonymous. Editorial Board. Journal of Computational Physics, 215(2):??, July 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001756.

Anonymous:2006:EBk

[Ano06k] Anonymous. Editorial Board. Journal of Computational Physics, 216(1):??, July 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002063.

Anonymous:2006:EBl

[Ano06l] Anonymous. Editorial Board. Journal of Computational Physics, 216(2):??, August 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600235X.

Anonymous:2006:EBm

[Ano06m] Anonymous. Editorial Board. Journal of Computational Physics, 217(1):??, September 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003251.

Anonymous:2006:EBn

[Ano06n] Anonymous. Editorial Board. Journal of Computational Physics, 217(2):??, September 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003846.

Anonymous:2006:EBo

[Ano06o] Anonymous. Editorial Board. Journal of Computational Physics, 218(1):??, October 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004013.

Anonymous:2006:EBp

[Ano06p] Anonymous. Editorial Board. Journal of Computational Physics, 218(2):??, November 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004104.

Anonymous:2006:EBq

[Ano06q] Anonymous. Editorial Board. Journal of Computational Physics, 219(1):??, November 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004578.

Anonymous:2006:EBr

[Ano06r] Anonymous. Editorial Board. Journal of Computational Physics, 219(2):??, December 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004876.

Anonymous:2006:EBs

[Ano06s] Anonymous. Editorial Board. Journal of Computational Physics, 220(1):??, December 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600502X.

Anonymous:2006:IAa

[Ano06t] Anonymous. Instructions to authors. Journal of Computational Physics, 211(1):??, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004031.

Anonymous:2006:IAb

[Ano06u] Anonymous. Instructions to authors. Journal of Computational Physics, 212(1):??, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500522X.

Anonymous:2006:IAc

[Ano06v] Anonymous. Instructions to authors. Journal of Computational Physics, 213(1):??, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000519.

Anonymous:2006:IAd

[Ano06w] Anonymous. Instructions to authors. Journal of Computational Physics, 214(1):??, May 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600101X.

Anonymous:2006:IAe

[Ano06x] Anonymous. Instructions to authors. Journal of Computational Physics, 215(1):??, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001562.

Anonymous:2006:IAf

[Ano06y] Anonymous. Instructions to authors. Journal of Computational Physics, 216(1):??, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002099.

Anonymous:2006:IAg

[Ano06z] Anonymous. Instructions to authors. Journal of Computational Physics, 217(1):??, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003287.

Anonymous:2006:IAh

[Ano06-27] Anonymous. Instructions to authors. Journal of Computational Physics, 218(1):??, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004049.

Anonymous:2006:VAIa

[Ano06-28] Anonymous. Volume author index. Journal of Computational Physics, 211(2):784–785, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004419.

Anonymous:2006:VAIb

[Ano06-29] Anonymous. Volume author index. Journal of Computational Physics, 212(2):798–799, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005310.

Anonymous:2006:VAIc

[Ano06-30] Anonymous. Volume author index. Journal of Computational Physics, 213(2):876–877, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000684.

Anonymous:2006:VAId

[Ano06-31] Anonymous. Volume author index. Journal of Computational Physics, 214(2):878–879, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001112.

Anonymous:2006:VAIe

[Ano06-32] Anonymous. Volume author index. Journal of Computational Physics, 215(2):778–779, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001781.

Anonymous:2006:VAIf

[Ano06-33] Anonymous. Volume author index. Journal of Computational Physics, 216(2):802–803, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002385.

Anonymous:2006:VAIg

[Ano06-34] Anonymous. Volume author index. Journal of Computational Physics, 217(2):868–869, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003871.

Anonymous:2006:VAIh

[Ano06-35] Anonymous. Volume author index. Journal of Computational Physics, 218(2):878–879, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600413X.

Anonymous:2006:VAIi

[Ano06-36] Anonymous. Volume author index. Journal of Computational Physics, 219(2):1013–1014, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005584.

Anonymous:2007:CCa

[Ano07a] Anonymous. Contents continued. Journal of Computational Physics, 225(1):??, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002914.

Anonymous:2007:CCb

[Ano07b] Anonymous. Contents continued. Journal of Computational Physics, 225(2):??, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002938.

Anonymous:2007:CCc

[Ano07c] Anonymous. Contents continued. Journal of Computational Physics, 226(2):??, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004111.

Anonymous:2007:CCIa

[Ano07d] Anonymous. Contents continued from IBC. Journal of Computational Physics, 227(1):??, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004573.

Anonymous:2007:CCIb

[Ano07e] Anonymous. Contents continued from IBC. Journal of Computational Physics, 227(2):??, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005074.

Anonymous:2007:EBa

[Ano07f] Anonymous. Editorial Board. Journal of Computational Physics, 220(2):??, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005122.

Anonymous:2007:EBb

[Ano07g] Anonymous. Editorial Board. Journal of Computational Physics, 221(1):??, January 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005225.

Anonymous:2007:EBc

[Ano07h] Anonymous. Editorial Board. Journal of Computational Physics, 221(2):??, February 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005304.

Anonymous:2007:EBd

[Ano07i] Anonymous. Editorial Board. Journal of Computational Physics, 222(1):??, March 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000228.

Anonymous:2007:EBe

[Ano07j] Anonymous. Editorial Board. Journal of Computational Physics, 222(2):??, March 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000514.

Anonymous:2007:EBf

[Ano07k] Anonymous. Editorial Board. Journal of Computational Physics, 223(1):??, April 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001052.

Anonymous:2007:EBg

[Ano07l] Anonymous. Editorial Board. Journal of Computational Physics, 223(2):??, May 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001131.

Anonymous:2007:EBh

[Ano07m] Anonymous. Editorial Board. Journal of Computational Physics, 224(1):??, May 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001763.

Anonymous:2007:EBi

[Ano07n] Anonymous. Editorial Board. Journal of Computational Physics, 224(2):??, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700188X.

Anonymous:2007:EBj

[Ano07o] Anonymous. Editorial Board. Journal of Computational Physics, 225(1):??, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002835.

Anonymous:2007:EBk

[Ano07p] Anonymous. Editorial Board. Journal of Computational Physics, 225(2):??, August 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002951.

Anonymous:2007:EBl

[Ano07q] Anonymous. Editorial Board. Journal of Computational Physics, 226(1):??, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003646.

Anonymous:2007:EBm

[Ano07r] Anonymous. Editorial Board. Journal of Computational Physics, 226(2):??, October 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004032.

Anonymous:2007:EBn

[Ano07s] Anonymous. Editorial Board. Journal of Computational Physics, 227(1):??, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004524.

Anonymous:2007:EBo

[Ano07t] Anonymous. Editorial Board. Journal of Computational Physics, 227(2):??, December 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004883.

Anonymous:2007:ICCa

[Ano07u] Anonymous. IBC (contents continued). Journal of Computational Physics, 225(1):??, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002860.

Anonymous:2007:ICCb

[Ano07v] Anonymous. IBC (contents continued). Journal of Computational Physics, 225(2):??, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002987.

Anonymous:2007:ICCc

[Ano07w] Anonymous. IBC (contents continued). Journal of Computational Physics, 226(2):??, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700407X.

Anonymous:2007:ICCd

[Ano07x] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(1):??, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700455X.

Anonymous:2007:ICCe

[Ano07y] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(2):??, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004913.

Anonymous:2007:IA

[Ano07z] Anonymous. Instructions to authors. Journal of Computational Physics, 221(1):??, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005250.

Anonymous:2007:OICa

[Ano07-27] Anonymous. OBC (issue contents). Journal of Computational Physics, 225(1):??, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002872.

Anonymous:2007:OICb

[Ano07-28] Anonymous. OBC (issue contents). Journal of Computational Physics, 225(2):??, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002999.

Anonymous:2007:OICc

[Ano07-29] Anonymous. OBC (issue contents). Journal of Computational Physics, 226(2):??, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004081.

Anonymous:2007:OICd

[Ano07-30] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(1):??, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004561.

Anonymous:2007:OICe

[Ano07-31] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(2):??, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004925.

Anonymous:2007:PN

[Ano07-32] Anonymous. Publisher’s note. Journal of Computational Physics, 224(2):??, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001957.

Anonymous:2007:VAIa

[Ano07-33] Anonymous. Volume author index. Journal of Computational Physics, 220(2):975–976, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005158.

Anonymous:2007:VAIb

[Ano07-34] Anonymous. Volume author index. Journal of Computational Physics, 221(2):854–855, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600533X.

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial Board. Journal of Computational Physics, 227(4):??, February 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800003X.

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial Board. Journal of Computational Physics, 227(5):??, February 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000260.

Anonymous:2008:EBc

[Ano08c] Anonymous. Editorial Board. Journal of Computational Physics, 227(6):??, March 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000624.

Anonymous:2008:EBd

[Ano08d] Anonymous. Editorial Board. Journal of Computational Physics, 227(7):??, March 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000880.

Anonymous:2008:EBe

[Ano08e] Anonymous. Editorial Board. Journal of Computational Physics, 227(8):??, April 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001289.

Anonymous:2008:EBf

[Ano08f] Anonymous. Editorial Board. Journal of Computational Physics, 227(9):??, April 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001459.

Anonymous:2008:EBg

[Ano08g] Anonymous. Editorial Board. Journal of Computational Physics, 227(10):??, May 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800185X.

Anonymous:2008:EBh

[Ano08h] Anonymous. Editorial Board. Journal of Computational Physics, 227(11):??, May 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002131.

Anonymous:2008:EBi

[Ano08i] Anonymous. Editorial Board. Journal of Computational Physics, 227(12):??, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002337.

Anonymous:2008:EBj

[Ano08j] Anonymous. Editorial Board. Journal of Computational Physics, 227(14):??, July 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002982.

Anonymous:2008:EBk

[Ano08k] Anonymous. Editorial Board. Journal of Computational Physics, 227(15):??, July 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003197.

Anonymous:2008:EBl

[Ano08l] Anonymous. Editorial Board. Journal of Computational Physics, 227(16):??, August 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003501.

Anonymous:2008:EBm

[Ano08m] Anonymous. Editorial Board. Journal of Computational Physics, 227(17):??, September 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003768.

Anonymous:2008:EBn

[Ano08n] Anonymous. Editorial Board. Journal of Computational Physics, 227(18):??, September 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003860.

Anonymous:2008:EBo

[Ano08o] Anonymous. Editorial Board. Journal of Computational Physics, 227(19):??, October 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004130.

Anonymous:2008:EBp

[Ano08p] Anonymous. Editorial Board. Journal of Computational Physics, 227(20):??, October 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800435X.

Anonymous:2008:EBq

[Ano08q] Anonymous. Editorial Board. Journal of Computational Physics, 227(21):??, November 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004610.

Anonymous:2008:EBr

[Ano08r] Anonymous. Editorial Board. Journal of Computational Physics, 227(22):??, November 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004853.

Anonymous:2008:EBs

[Ano08s] Anonymous. Editorial Board. Journal of Computational Physics, 227(23):??, December 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005147.

Anonymous:2008:EBt

[Ano08t] Anonymous. Editorial Board. Journal of Computational Physics, 227(24):??, December 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005482.

Anonymous:2008:ICCa

[Ano08u] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(10):??, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001885.

Anonymous:2008:ICCb

[Ano08v] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(11):??, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002167.

Anonymous:2008:ICCc

[Ano08w] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(12):??, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002362.

Anonymous:2008:ICCd

[Ano08x] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(13):??, June 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002763.

Anonymous:2008:ICCe

[Ano08y] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(14):??, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800301X.

Anonymous:2008:ICCf

[Ano08z] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(15):??, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003227.

Anonymous:2008:ICCg

[Ano08-27] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(16):??, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003537.

Anonymous:2008:ICCh

[Ano08-28] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(19):??, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004166.

Anonymous:2008:ICCi

[Ano08-29] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(20):??, October 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004385.

Anonymous:2008:ICCj

[Ano08-30] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(21):??, November 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004646.

Anonymous:2008:ICCk

[Ano08-31] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(22):??, November 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004889.

Anonymous:2008:ICCl

[Ano08-32] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(23):??, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005172.

Anonymous:2008:ICCm

[Ano08-33] Anonymous. IBC (contents continued). Journal of Computational Physics, 227(24):??, December 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005512.

Anonymous:2008:OC

[Ano08-34] Anonymous. OBC (contents). Journal of Computational Physics, 227(5):??, February 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000302.

Anonymous:2008:OICa

[Ano08-35] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(10):??, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001897.

Anonymous:2008:OICb

[Ano08-36] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(11):??, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002179.

Anonymous:2008:OICc

[Ano08-37] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(12):??, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002374.

Anonymous:2008:OICd

[Ano08-38] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(13):??, June 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002775.

Anonymous:2008:OICe

[Ano08-39] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(14):??, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003021.

Anonymous:2008:OICf

[Ano08-40] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(15):??, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003239.

Anonymous:2008:OICg

[Ano08-41] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(16):??, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003549.

Anonymous:2008:OICh

[Ano08-42] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(17):??, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800380X.

Anonymous:2008:OICi

[Ano08-43] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(18):??, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003902.

Anonymous:2008:OICj

[Ano08-44] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(19):??, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004178.

Anonymous:2008:OICk

[Ano08-45] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(20):??, October 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004397.

Anonymous:2008:OICl

[Ano08-46] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(21):??, November 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004658.

Anonymous:2008:OICm

[Ano08-47] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(22):??, November 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004890.

Anonymous:2008:OICn

[Ano08-48] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(23):??, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005184.

Anonymous:2008:OICo

[Ano08-49] Anonymous. OBC (issue contents). Journal of Computational Physics, 227(24):??, December 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005524.

Anonymous:2008:OFT

[Ano08-50] Anonymous. Ocean forecasting in terrain-following coordinates: Formulation and skill assessment of the regional ocean modeling system. Journal of Computational Physics, 227(7):3595–3624, March 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999107002549. See correction and commentary [SM09b].

Anonymous:2008:PN

[Ano08-51] Anonymous. Publisher’s note. Journal of Computational Physics, 227(12):5937–5938, June 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002416.

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial Board. Journal of Computational Physics, 228(1):??, January 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005767.

Anonymous:2009:EBb

[Ano09b] Anonymous. Editorial Board. Journal of Computational Physics, 228(2):??, February 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005998.

Anonymous:2009:EBc

[Ano09c] Anonymous. Editorial Board. Journal of Computational Physics, 228(3):??, February 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006414.

Anonymous:2009:EBd

[Ano09d] Anonymous. Editorial Board. Journal of Computational Physics, 228(4):??, March 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006657.

Anonymous:2009:EBe

[Ano09e] Anonymous. Editorial Board. Journal of Computational Physics, 228(5):??, March 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000175.

Anonymous:2009:EBf

[Ano09f] Anonymous. Editorial Board. Journal of Computational Physics, 228(6):??, April 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000461.

Anonymous:2009:EBg

[Ano09g] Anonymous. Editorial Board. Journal of Computational Physics, 228(7):??, April 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000722.

Anonymous:2009:EBh

[Ano09h] Anonymous. Editorial Board. Journal of Computational Physics, 228(8):??, May 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000990.

Anonymous:2009:EBi

[Ano09i] Anonymous. Editorial Board. Journal of Computational Physics, 228(9):??, May 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001193.

Anonymous:2009:EBj

[Ano09j] Anonymous. Editorial Board. Journal of Computational Physics, 228(10):??, June 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900148X.

Anonymous:2009:EBk

[Ano09k] Anonymous. Editorial Board. Journal of Computational Physics, 228(11):??, June 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001752.

Anonymous:2009:EBl

[Ano09l] Anonymous. Editorial Board. Journal of Computational Physics, 228(12):??, July 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002186.

Anonymous:2009:EBm

[Ano09m] Anonymous. Editorial Board. Journal of Computational Physics, 228(13):??, July 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002484.

Anonymous:2009:EBn

[Ano09n] Anonymous. Editorial Board. Journal of Computational Physics, 228(14):??, August 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002939.

Anonymous:2009:EBo

[Ano09o] Anonymous. Editorial Board. Journal of Computational Physics, 228(15):??, August 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003015.

Anonymous:2009:EBp

[Ano09p] Anonymous. Editorial Board. Journal of Computational Physics, 228(16):??, September 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003507.

Anonymous:2009:EBq

[Ano09q] Anonymous. Editorial Board. Journal of Computational Physics, 228(17):??, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003714.

Anonymous:2009:EBr

[Ano09r] Anonymous. Editorial Board. Journal of Computational Physics, 228(18):??, October 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003982.

Anonymous:2009:EBs

[Ano09s] Anonymous. Editorial Board. Journal of Computational Physics, 228(19):??, October 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004264.

Anonymous:2009:EBt

[Ano09t] Anonymous. Editorial Board. Journal of Computational Physics, 228(20):??, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004501.

Anonymous:2009:EBu

[Ano09u] Anonymous. Editorial Board. Journal of Computational Physics, 228(21):??, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004781.

Anonymous:2009:EBv

[Ano09v] Anonymous. Editorial Board. Journal of Computational Physics, 228(22):??, December 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005038.

Anonymous:2009:EBw

[Ano09w] Anonymous. Editorial Board. Journal of Computational Physics, 228(23):??, December 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005385.

Anonymous:2009:EBx

[Ano09x] Anonymous. Editorial Board. Journal of Computational Physics, 228(24):??, December 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005658.

Anonymous:2009:ICCa

[Ano09y] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(1):??, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005792.

Anonymous:2009:ICCb

[Ano09z] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(2):??, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006025.

Anonymous:2009:ICCc

[Ano09-27] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(3):??, February 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800644X.

Anonymous:2009:ICCd

[Ano09-28] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(4):??, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006682.

Anonymous:2009:ICCe

[Ano09-29] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(5):??, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000205.

Anonymous:2009:ICCf

[Ano09-30] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(6):??, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000497.

Anonymous:2009:ICCg

[Ano09-31] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(7):??, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000758.

Anonymous:2009:ICCh

[Ano09-32] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(8):??, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001028.

Anonymous:2009:ICCi

[Ano09-33] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(9):??, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001223.

Anonymous:2009:ICCj

[Ano09-34] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(10):??, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900151X.

Anonymous:2009:ICCk

[Ano09-35] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(11):??, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001788.

Anonymous:2009:ICCl

[Ano09-36] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(12):??, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002216.

Anonymous:2009:ICCm

[Ano09-37] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(13):??, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002514.

Anonymous:2009:ICCn

[Ano09-38] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(14):??, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002964.

Anonymous:2009:ICCo

[Ano09-39] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(15):??, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003040.

Anonymous:2009:ICCp

[Ano09-40] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(16):??, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003532.

Anonymous:2009:ICCq

[Ano09-41] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(17):??, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900374X.

Anonymous:2009:ICCr

[Ano09-42] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(18):??, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900401X.

Anonymous:2009:ICCs

[Ano09-43] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(19):??, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004318.

Anonymous:2009:ICCt

[Ano09-44] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(20):??, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004537.

Anonymous:2009:ICCu

[Ano09-45] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(21):??, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004811.

Anonymous:2009:ICCv

[Ano09-46] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(22):??, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005063.

Anonymous:2009:ICCw

[Ano09-47] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(23):??, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005415.

Anonymous:2009:ICCx

[Ano09-48] Anonymous. IBC (contents continued). Journal of Computational Physics, 228(24):??, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005683.

Anonymous:2009:OICa

[Ano09-49] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(1):??, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005809.

Anonymous:2009:OICb

[Ano09-50] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(2):??, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006037.

Anonymous:2009:OICc

[Ano09-51] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(3):??, February 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006451.

Anonymous:2009:OICd

[Ano09-52] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(4):??, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006694.

Anonymous:2009:OICe

[Ano09-53] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(5):??, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000217.

Anonymous:2009:OICf

[Ano09-54] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(6):??, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000503.

Anonymous:2009:OICg

[Ano09-55] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(7):??, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900076X.

Anonymous:2009:OICh

[Ano09-56] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(8):??, May 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900103X.

Anonymous:2009:OICi

[Ano09-57] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(9):??, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001235.

Anonymous:2009:OICj

[Ano09-58] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(10):??, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001521.

Anonymous:2009:OICk

[Ano09-59] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(11):??, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900179X.

Anonymous:2009:OICl

[Ano09-60] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(12):??, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002228.

Anonymous:2009:OICm

[Ano09-61] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(13):??, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002526.

Anonymous:2009:OICn

[Ano09-62] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(14):??, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002976.

Anonymous:2009:OICo

[Ano09-63] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(15):??, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003052.

Anonymous:2009:OICp

[Ano09-64] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(16):??, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003544.

Anonymous:2009:OICq

[Ano09-65] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(17):??, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003751.

Anonymous:2009:OICr

[Ano09-66] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(18):??, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004021.

Anonymous:2009:OICs

[Ano09-67] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(19):??, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900432X.

Anonymous:2009:OICt

[Ano09-68] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(20):??, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004549.

Anonymous:2009:OICu

[Ano09-69] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(21):??, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004823.

Anonymous:2009:OICv

[Ano09-70] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(22):??, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005075.

Anonymous:2009:OICw

[Ano09-71] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(23):??, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005427.

Anonymous:2009:OICx

[Ano09-72] Anonymous. OBC (issue contents). Journal of Computational Physics, 228(24):??, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005695.

Auteri:2002:MBS

[AP02] F. Auteri and N. Parolini. A mixed-basis spectral projection method. Journal of Computational Physics, 175(1):1–23, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910196855X.

Aguayo:2007:NPP

[APP^{+}07] J. P. Aguayo, P. M. Phillips, T. N. Phillips, H. R. Tamaddon-Jahromi, B. A. Snigerev, and M. F. Webster. The numerical prediction of planar viscoelastic contraction flows using the pom-pom model and higher-order finite volume schemes. Journal of Computational Physics, 220(2):586–611, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002518.

Auteri:2002:NIS

[APQ02] F. Auteri, N. Parolini, and L. Quartapelle. Numerical investigation on the stability of singular driven cavity flow. Journal of Computational Physics, 183(1):1–25, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971457.

Auteri:2003:EIN

[APQ03] F. Auteri, N. Parolini, and L. Quartapelle. Essential imposition of Neumann condition in Galerkin–Legendre elliptic solvers. Journal of Computational Physics, 185(2):427–444, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000645.

Ahmad:2009:FIS

[APR09] Sk. Safique Ahmad, Nigam Chandra Parida, and Soumyendu Raha. The fully implicit stochastic-α method for stiff stochastic differential equations. Journal of Computational Physics, 228(22):8263–8282, December 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004380.

Airaksinen:2009:DPT

[APT09] Tuomas Airaksinen, Anssi Pennanen, and Jari Toivanen. A damping preconditioner for time-harmonic wave equations in fluid and elastic material. Journal of Computational Physics, 228(5):1466–1479, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005688.

Aboubacar:2004:HOF

[APTJ^{+}04] M. Aboubacar, T. N. Phillips, H. R. Tamaddon-Jahromi, B. A. Snigerev, and M. F. Webster. High-order finite volume methods for viscoelastic flow problems. Journal of Computational Physics, 199(1):16–40, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000634.

Auteri:2000:GLS

[AQ00] F. Auteri and L. Quartapelle. Galerkin–Legendre spectral method for the 3D Helmholtz equation. Journal of Computational Physics, 161(2):454–483, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965045.

Auteri:2007:SSS

[AQ07] F. Auteri and L. Quartapelle. Spectral solvers for spherical elliptic problems. Journal of Computational Physics, 227 (1):36–54, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107003087.

Auteri:2009:NSS

[AQ09] F. Auteri and L. Quartapelle. Navier–Stokes spectral solver in a sphere. Journal of Computational Physics, 228(19):7197–7214, October 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003416.

Auteri:2002:ASS

[AQV02] F. Auteri, L. Quartapelle, and L. Vigevano. Accurate ω–ψ spectral solution of the singular driven cavity problem. Journal of Computational Physics, 180(2):597–615, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971081.

Akhmatskaya:2008:GEM

[AR08] Elena Akhmatskaya and Sebastian Reich. GSHMC: An efficient method for molecular simulation. Journal of Computational Physics, 227(10):4934–4954, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000533.

Acebron:2009:DDS

[ARRS09] Juan A. Acebrón, Ángel Rodríguez-Rozas, and Renato Spigler. Domain decomposition solution of nonlinear twodimensional parabolic problems by random trees. Journal of Computational Physics, 228(15):5574–5591, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002307.

Al-Rawahi:2002:NSD

[ART02] Nabeel Al-Rawahi and Gretar Tryggvason. Numerical simulation of dendritic solidification with convection: Twodimensional geometry. Journal of Computational Physics, 180(2):471–496, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102970920.

Al-Rawahi:2004:NSD

[ART04] Nabeel Al-Rawahi and Gretar Tryggvason. Numerical simulation of dendritic solidification with convection: Threedimensional flow. Journal of Computational Physics, 194 (2):677–696, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103005011.

Adams:2002:SSD

[AS02] N. A. Adams and S. Stolz. A subgrid-scale deconvolution approach for shock capturing. Journal of Computational Physics, 178(2):391–426, May 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970348.

Abgrall:2003:DEP

[AS03a] Rémi Abgrall and Richard Saurel. Discrete equations for physical and numerical compressible multiphase mixtures. Journal of Computational Physics, 186(2):361–396, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000111.

Adalsteinsson:2003:TDM

[AS03b] David Adalsteinsson and J. A. Sethian. Transport and diffusion of material quantities on propagating interfaces via level set methods. Journal of Computational Physics, 185 (1):271–288, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102000578.

Acebron:2005:FSS

[AS05a] Juan A. Acebrón and Renato Spigler. Fast simulations of stochastic dynamical systems. Journal of Computational Physics, 208(1):106–115, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000860.

Alexandrov:2005:TPL

[AS05b] Oleg Alexandrov and Fadil Santosa. A topology-preserving level set method for shape optimization. Journal of Computational Physics, 204(1):121–130, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004103.

Ahn:2007:MMI

[AS07] Hyung Taek Ahn and Mikhail Shashkov. Multi-material interface reconstruction on generalized polyhedral meshes. Journal of Computational Physics, 226(2):2096–2132, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002926.

Ahn:2009:AMF

[AS09] Hyung Taek Ahn and Mikhail Shashkov. Adaptive momentof-fluid method. Journal of Computational Physics, 228(8): 2792–2821, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006578.

Aslam:2001:LSA

[Asl01] Tariq D. Aslam. A level-set algorithm for tracking discontinuities in hyperbolic conservation laws: I. Scalar equations. Journal of Computational Physics, 167(2):413–438, March 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966865.

Aslam:2004:PDE

[Asl04a] Tariq D. Aslam. A partial differential equation approach to multidimensional extrapolation. Journal of Computational Physics, 193(1):349–355, January 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004170.

Aslan:2004:VFS

[Asl04b] Necdet Aslan. A visual fluctuation splitting scheme for magnetohydrodynamics with a new sonic fix and Euler limit. Journal of Computational Physics, 197(1):1–27, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006168.

Alimi:2003:SPH

[ASPB03] J.-M Alimi, A. Serna, C. Pastor, and G. Bernabeu. Smooth particle hydrodynamics: importance of correction terms in adaptive resolution algorithms. Journal of Computational Physics, 192(1):157–174, November 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003516.

Amundson:2006:SAM

[ASQR06] J. Amundson, P. Spentzouris, J. Qiang, and R. Ryne. Synergia: An accelerator modeling tool with 3-D space charge. Journal of Computational Physics, 211(1):229–248, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002718.

Achdou:2007:TBC

[AST07] Yves Achdou, Christophe Sabot, and Nicoletta Tchou. Transparent boundary conditions for the Helmholtz equation in some ramified domains with a fractal boundary. Journal of Computational Physics, 220(2):712–739, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002567.

Arndt:2009:EAD

[AST09] M. Arndt, V. Sorkin, and E. B. Tadmor. Efficient algorithms for discrete lattice calculations. Journal of Computational Physics, 228(13):4858–4880, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001703.

Arminjon:2005:CFV

[AT05a] P. Arminjon and R. Touma. Central finite volume methods with constrained transport divergence treatment for ideal MHD. Journal of Computational Physics, 204(2):737–759, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910400436X.

Azarenok:2005:SOG

[AT05b] Boris N. Azarenok and Tao Tang. Second-order Godunovtype scheme for reactive flow calculations on moving meshes. Journal of Computational Physics, 206(1):48–80, June 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004930.

Artebrant:2008:IAL

[AT08] Robert Artebrant and Manuel Torrilhon. Increasing the accuracy in locally divergence-preserving finite volume schemes for MHD. Journal of Computational Physics, 227(6):3405–3427, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005554.

Arico:2009:MFF

[AT09] Costanza Aricò and Tullio Tucciarelli. The MAST FV/FE scheme for the simulation of two-dimensional thermohaline processes in variable-density saturated porous media. Journal of Computational Physics, 228(4):1234–1274, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005457.

Atassi:2004:NBC

[Ata04] Oliver V. Atassi. Nonreflecting boundary conditions for the time-dependent convective wave equation in a duct. Journal of Computational Physics, 197(2):737–758, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000555.

Akaiwa:2001:LSS

[ATV01] Norio Akaiwa, K. Thornton, and P. W. Voorhees. Large-scale simulations of microstructural evolution in elastically stressed solids. Journal of Computational Physics, 173(1):61–86, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968421.

Arber:2002:CCE

[AV02] T. D. Arber and R. G. L. Vann. A critical comparison of Eulerian-grid-based Vlasov solvers. Journal of Computational Physics, 180(1):339–357, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970981.

Ammari:2003:COT

[AV03] Habib Ammari and Darko Volkov. Correction of order three for the expansion of two dimensional electromagnetic fields perturbed by the presence of inhomogeneities of small diameter. Journal of Computational Physics, 189(2):371–389, August 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002249.

Abide:2005:CFO

[AV05] Stéphane Abide and Stéphane Viazzo. A 2D compact fourthorder projection decomposition method. Journal of Computational Physics, 206(1):252–276, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005121.

Abubakar:2004:IFI

[AvdB04] Aria Abubakar and Peter M. van den Berg. Iterative forward and inverse algorithms based on domain integral equations for three-dimensional electric and magnetic objects. Journal of Computational Physics, 195(1):236–262, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910300531X.

Andrianov:2004:RPB

[AW04] Nikolai Andrianov and Gerald Warnecke. The Riemann problem for the Baer–Nunziato two-phase flow model. Journal of Computational Physics, 195(2):434–464, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005370.

Altenhoff:2007:SBF

[AWK07] Adrian M. Altenhoff, Jens H. Walther, and Petros Koumoutsakos. A stochastic boundary forcing for dissipative particle dynamics. Journal of Computational Physics, 225(1): 1125–1136, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107000320.

Ashcroft:2003:OPC

[AZ03] Graham Ashcroft and Xin Zhang. Optimized prefactored compact schemes. Journal of Computational Physics, 190(2):459–477, September 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002936.

Asokan:2005:USA

[AZ05] Badrinarayanan Velamur Asokan and Nicholas Zabaras. Using stochastic analysis to capture unstable equilibrium in natural convection. Journal of Computational Physics, 208(1):134–153, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000884.

Asokan:2006:SVM

[AZ06] Badrinarayanan Velamur Asokan and Nicholas Zabaras. A stochastic variational multiscale method for diffusion in heterogeneous random media. Journal of Computational Physics, 218(2):654–676, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001252.

Azarenok:2006:VHG

[Aza06] Boris N. Azarenok. A variational hexahedral grid generator with control metric. Journal of Computational Physics, 218(2):720–747, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001288.

Azarenok:2007:MCA

[Aza07] Boris N. Azarenok. A method of constructing adaptive hexahedral moving grids. Journal of Computational Physics, 226(1):1102–1121, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002239. See comment [Aza09].

Azarenok:2009:CPB

[Aza09] Boris Azarenok. Comment on the paper by B. N. Azarenok, “A method of constructing adaptive hexahedral moving grids” 226 (2007), pp. 1102–1121. Journal of Computational Physics, 228(13):4961, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001375. See [Aza07].

Afkhami:2009:MDM

[AZB09] S. Afkhami, S. Zaleski, and M. Bussmann. A mesh-dependent model for applying dynamic contact angles to VOF simulations. Journal of Computational Physics, 228(15):5370–5389, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002083.

Anderson:2005:AUV

[AZC05] Anthony Anderson, Xiaoming Zheng, and Vittorio Cristini. Adaptive unstructured volume remeshing — I: The method. Journal of Computational Physics, 208(2):616–625, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001221.

Azmy:2002:USR

[Azm02] Y. Y. Azmy. Unconditionally stable and robust adjacentcell diffusive preconditioning of weighted-difference particle transport methods is impossible. Journal of Computational Physics, 182(1):213–233, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971627.

Berthet:2003:NBC

[BA03] R. Berthet and D. Astruc. Numerical boundary conditions for sound scattering simulation. Journal of Computational Physics, 190(1):64–99, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002547.

Ben-Artzi:2005:PCS

[BACFT05] Matania Ben-Artzi, Jean-Pierre Croisille, Dalia Fishelov, and Shlomo Trachtenberg. A pure-compact scheme for the streamfunction formulation of Navier–Stokes equations. Journal of Computational Physics, 205(2):640–664, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004851.

Becciani:2000:MPT

[BADG00] U. Becciani, V. Antonuccio-Delogu, and M. Gambera. A modified parallel tree code for N-body simulation of the large-scale structure of the universe. Journal of Computational Physics, 163(1):118–132, September 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965574.

Baeurle:2003:CWA

[Bae03] S. A. Baeurle. Computation within the auxiliary field approach. Journal of Computational Physics, 184(2):540–558, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000360.

Ben-Artzi:2009:HCL

[BAFL09] Matania Ben-Artzi, Joseph Falcovitz, and Philippe G. LeFloch. Hyperbolic conservation laws on the sphere. A geometry-compatible finite volume scheme. Journal of Computational Physics, 228(16):5650–5668, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001880.

Balsara:2001:DFA

[Bal01] Dinshaw S. Balsara. Divergence-free adaptive mesh refinement for magnetohydrodynamics. Journal of Computational Physics, 174(2):614–648, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969177.

Bal:2002:PTT

[Bal02] Guillaume Bal. Particle transport through scattering regions with clear layers and inclusions. Journal of Computational Physics, 180(2):659–685, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971111.

Baldauf:2008:SAL

[Bal08] Michael Baldauf. Stability analysis for linear discretisations of the advection equation with Runge–Kutta time integration. Journal of Computational Physics, 227(13):6638–6659, June 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001769.

Balsara:2009:DFR

[Bal09] Dinshaw S. Balsara. Divergence-free reconstruction of magnetic fields and WENO schemes for magnetohydrodynamics.

Journal of Computational Physics, 228(14):5040–5056, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001727.

Ben-Artzi:2006:DEG

[BALW06] Matania Ben-Artzi, Jiequan Li, and Gerald Warnecke. A direct Eulerian GRP scheme for compressible fluid flows. Journal of Computational Physics, 218(1):19–43, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000581.

Balsara:2007:SCB

[BAMD07] Dinshaw S. Balsara, Christoph Altmann, Claus-Dieter Munz, and Michael Dumbser. A sub-cell based indicator for troubled zones in RKDG schemes and a novel class of hybrid RKDG+HWENO schemes. Journal of Computational Physics, 226(1):586–620, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001842.

Bardenhagen:2002:ECE

[Bar02a] S. G. Bardenhagen. Energy conservation error in the material point method for solid mechanics. Journal of Computational Physics, 180(1):383–403, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971032.

Bartello:2002:CTD

[Bar02b] P. Bartello. A comparison of time discretization schemes for two-timescale problems in geophysical fluid dynamics. Journal of Computational Physics, 179(1):268–285, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970567.

Barnett:2004:FNM

[Bar04] Alex H. Barnett. A fast numerical method for time-resolved photon diffusion in general stratified turbid media. Journal of Computational Physics, 201(2):771–797, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002645.

Banks:2008:SLC

[BAR08] J. W. Banks, T. Aslam, and W. J. Rider. On sub-linear convergence for linearly degenerate waves in capturing schemes. Journal of Computational Physics, 227(14):6985–7002, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002088.

Bailey:2008:PLF

[BAYZ08] Teresa S. Bailey, Marvin L. Adams, Brian Yang, and Michael R. Zika. A piecewise linear finite element discretization of the diffusion equation for arbitrary polyhedral grids. Journal of Computational Physics, 227(8):3738–3757, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005050.

Baker:2004:VBM

[BB04a] Gregory R. Baker and J. Thomas Beale. Vortex blob methods applied to interfacial motion. Journal of Computational Physics, 196(1):233–258, May 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005953.

Bogey:2004:FLD

[BB04b] Christophe Bogey and Christophe Bailly. A family of low dispersive and low dissipative explicit schemes for flow and noise computations. Journal of Computational Physics, 194(1):194–214, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004662.

Bellavia:2007:GSN

[BB07a] Stefania Bellavia and Stefano Berrone. Globalization strategies for Newton–Krylov methods for stabilized FEM discretization of Navier–Stokes equations. Journal of Computational Physics, 226(2):2317–2340, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003178.

Bogey:2007:AES

[BB07b] Christophe Bogey and Christophe Bailly. On the application of explicit spatial filtering to the variables or fluxes of linear equations. Journal of Computational Physics, 225(2):1211–1217, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001593.

Barnett:2008:SCM

[BB08a] A. H. Barnett and T. Betcke. Stability and convergence of the method of fundamental solutions for Helmholtz problems on analytic domains. Journal of Computational Physics, 227(14):7003–7026, July 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002106.

Burt:2008:LDP

[BB08b] Jonathan M. Burt and Iain D. Boyd. A low diffusion particle method for simulating compressible inviscid flows. Journal of Computational Physics, 227(9):4653–4670, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000454.

Besse:2009:GWB

[BB09a] Nicolas Besse and Pierre Bertrand. Gyro-water-bag approach in nonlinear gyrokinetic turbulence. Journal of Computational Physics, 228(11):3973–3995, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000862.

Brunner:2009:ERD

[BB09b] Thomas A. Brunner and Patrick S. Brantley. An efficient, robust, domain-decomposition algorithm for particle Monte Carlo. Journal of Computational Physics, 228(10):3882–3890, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000965.

Burt:2009:HPA

[BB09c] Jonathan M. Burt and Iain D. Boyd. A hybrid particle approach for continuum and rarefied flow simulation. Journal of Computational Physics, 228(2):460–475, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005044.

Berland:2008:SDE

[BBB08] J. Berland, C. Bogey, and C. Bailly. A study of differentiation errors in large-eddy simulations based on the EDQNM theory. Journal of Computational Physics, 227 (18):8314–8340, September 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002933.

Bauer:2006:ICS

[BBC^{+}06] A. L. Bauer, D. E. Burton, E. J. Caramana, R. Loubère, M. J. Shashkov, and P. P. Whalen. The internal consistency, stability, and accuracy of the discrete, compatible formulation of Lagrangian hydrodynamics. Journal of Computational Physics, 218(2):572–593, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001203.

Bona:2009:LHR

[BBCT09] C. Bona, C. Bona-Casas, and J. Terradas. Linear highresolution schemes for hyperbolic conservation laws: TVB numerical evidence. Journal of Computational Physics, 228 (6):2266–2281, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006335.

Barucq:2004:CLB

[BBD04] Hélène Barucq, Chokri Bekkey, and Rabia Djellouli. Construction of local boundary conditions for an eigenvalue problem using micro-local analysis: application to optical waveguide problems. Journal of Computational Physics, 193(2):666–696, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004558.

Baro:2005:CSD

[BBDE05] M. Baro, N. Ben Abdallah, P. Degond, and A. El Ayyadi. A 1D coupled Schrödinger drift-diffusion model including collisions. Journal of Computational Physics, 203(1):129– 153, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003213.

Bell:2008:SDF

[BBF^{+}08] George I. Bell, David L. Bruhwiler, Alexei Fedotov, Andrey Sobol, Richard S. Busby, Peter Stoltz, Dan T. Abell, Peter Messmer, Ilan Ben-Zvi, and Vladimir Litvinenko. Simulating the dynamical friction force on ions due to a briefly co-propagating electron beam. Journal of Computational Physics, 227(19):8714–8735, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003410.

Belien:2002:FAM

[BBG^{+}02] A. J. C. Beliën, M. A. Botchev, J. P. Goedbloed, B. van der Holst, and R. Keppens. FINESSE: Axisymmetric MHD equilibria with flow. Journal of Computational Physics, 182 (1):91–117, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971536.

Barlett:2009:DBF

[BBHM09] V. Ruiz Barlett, J. J. Bigeón, M. Hoyuelos, and H. O. Mártin. Differences between fixed time step and kinetic Monte Carlo methods for biased diffusion. Journal of Computational Physics, 228(16):5740–5748, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002332.

Bergmann:2009:ERP

[BBI09] M. Bergmann, C.-H. Bruneau, and A. Iollo. Enablers for robust POD models. Journal of Computational Physics, 228 (2):516–538, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800510X.

Burstedde:2006:CLM

[BBK06] Carsten Burstedde, Jürgen Braun, and Angela Kunoth. Computing light masks in neutral atom lithography. Journal of Computational Physics, 220(1):422–440, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600341X.

Bronstein:2007:WDM

[BBK07] Alexander M. Bronstein, Michael M. Bronstein, and Ron Kimmel. Weighted distance maps computation on parametric three-dimensional manifolds. Journal of Computational Physics, 225(1):771–784, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000034.

Berland:2007:HOL

[BBMB07] Julien Berland, Christophe Bogey, Olivier Marsden, and Christophe Bailly. High-order, low dispersive and low dissipative explicit schemes for multiple-scale and boundary problems. Journal of Computational Physics, 224(2):637–662, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005420.

Bidegaray:2001:IPR

[BBR01] B. Bidégaray, A. Bourgeade, and D. Reignier. Introducing physical relaxation terms in Bloch equations. Journal of Computational Physics, 170(2):603–613, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196752X.

Bernsen:2006:DCF

[BBvdV06] Erik Bernsen, Onno Bokhove, and Jaap J. W. van der Vegt. A (dis)continuous finite element model for generalized 2D vorticity dynamics. Journal of Computational Physics, 211(2):719–747, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002949.

Bhattacharya:2006:FFA

[BBW06] S. Bhattacharya, J. Bławzdziewicz, and E. Wajnryb. Far-field approximation for hydrodynamic interactions in parallel-wall geometry. Journal of Computational Physics, 212(2):718–738, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003530.

Barros:2001:NGM

[BC01] Saulo R. M. Barros and José W. Cárdenas. A nonlinear Galerkin method for the shallow-water equations on periodic domains. Journal of Computational Physics, 172(2):592–608, September 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968433.

Balls:2002:FDD

[BC02a] Gregory T. Balls and Phillip Colella. A finite difference domain decomposition method using local corrections for the solution of Poisson’s equation. Journal of Computational Physics, 180(1):25–53, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970683.

Buet:2002:NAI

[BC02b] C. Buet and S. Cordier. Numerical analysis of the isotropic Fokker–Planck–Landau equation. Journal of Computational Physics, 179(1):43–67, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970439.

Barad:2005:FOA

[BC05] Michael Barad and Phillip Colella. A fourth-order accurate local refinement method for Poisson’s equation. Journal of Computational Physics, 209(1):1–18, October 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001129.

Bergmann:2008:OCC

[BC08] M. Bergmann and L. Cordier. Optimal control of the cylinder wake in the laminar regime by trust-region methods and POD reduced-order models. Journal of Computational Physics, 227(16):7813–7840, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002659.

Badalassi:2003:CMS

[BCB03] V. E. Badalassi, H. D. Ceniceros, and S. Banerjee. Computation of multiphase systems with phase field models. Journal of Computational Physics, 190(2):371–397, September 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002808.

Borges:2008:IWE

[BCCD08] Rafael Borges, Monique Carmona, Bruno Costa, and Wai Sun Don. An improved weighted essentially non-oscillatory scheme for hyperbolic conservation laws. Journal of Computational Physics, 227(6):3191–3211, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005232.

BenAbdallah:2009:DSH

[BCCV09] N. Ben Abdallah, M. J. Cáceres, J. A. Carrillo, and F. Vecil. A deterministic solver for a hybrid quantum-classical transport model in nano-MOSFETs. Journal of Computational Physics, 228(17):6553–6571, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003180.

Bassi:2006:ACF

[BCDR06] F. Bassi, A. Crivellini, D. A. Di Pietro, and S. Rebay. An artificial compressibility flux for the discontinuous Galerkin solution of the incompressible Navier–Stokes equations. Journal of Computational Physics, 218(2):794–815, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001331.

Barchanski:2006:ECC

[BCDW06] Andreas Barchanski, Markus Clemens, Herbert De Gersem, and Thomas Weiland. Efficient calculation of current densities in the human body induced by arbitrarily shaped, low-frequency magnetic field sources. Journal of Computational Physics, 214(1):81–95, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004249.

Bendito:2009:CCF

[BCE^{+}09] E. Bendito, A. Carmona, A. M. Encinas, J. M. Gesto, A. Gómez, C. Mouriño, and M. T. Sánchez. Computational cost of the Fekete problem I: The forces method on the 2sphere. Journal of Computational Physics, 228(9):3288–3306, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000357.

Bendito:2007:EFP

[BCEG07] E. Bendito, A. Carmona, A. M. Encinas, and J. M. Gesto. Estimation of Fekete points. Journal of Computational Physics, 225(2):2354–2376, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001362.

Belaouar:2009:NCL

[BCG09] R. Belaouar, T. Colin, and G. Gallice. Numerical coupling of Landau damping and Raman amplification. Journal of Computational Physics, 228(2):387–405, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004968.

Budd:2005:PCC

[BCGR05] C. J. Budd, R. Carretero-González, and R. D. Russell. Precise computations of chemotactic collapse using moving mesh methods. Journal of Computational Physics, 202(2):463–487, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104002943.

Barrault:2007:MDD

[BCHL07] M. Barrault, E. Cancès, W. W. Hager, and C. Le Bris. Multilevel domain decomposition for electronic structure calculations. Journal of Computational Physics, 222(1):86–109, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106003184.

Buffoni:2008:NLO

[BCI^{+}08] M. Buffoni, S. Camarri, A. Iollo, E. Lombardi, and M. V. Salvetti. A non-linear observer for unsteady three-dimensional flows. Journal of Computational Physics, 227(4):2626–2643, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004858.

Bayati:2009:DLA

[BCK09] Basil Bayati, Philippe Chatelain, and Petros Koumoutsakos. D-leaping: Accelerating stochastic simulation algorithms for reactions with delays. Journal of Computational Physics, 228(16):5908–5916, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002435.

Bao:2006:ESAb

[BCL06] Weizhu Bao, I-Liang Chern, and Fong Yin Lim. Efficient and spectrally accurate numerical methods for computing ground and first excited states in Bose–Einstein condensates. Journal of Computational Physics, 219(2):836–854, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600218X.

Brown:2001:APM

[BCM01] David L. Brown, Ricardo Cortez, and Michael L. Minion. Accurate projection methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 168(2):464–499, April 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967154.

Bottauscio:2007:MAA

[BCM^{+}07] Oriano Bottauscio, Mario Chiampi, Alessandra Manzin, Paolo Emilio Roccato, and Mauro Zucca. A multiscale approach to the analysis of magnetic grid shields and its validation. Journal of Computational Physics, 227(2):1470–1482, December 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004172.

Blasco:2009:FDP

[BCM09] Jordi Blasco, M. Carmen Calzada, and Mercedes Marín. A fictitious domain, parallel numerical method for rigid particulate flows. Journal of Computational Physics, 228(20):7596–7613, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003817.

Burchard:2001:MCT

[BCMO01] Paul Burchard, Li-Tien Cheng, Barry Merriman, and Stanley Osher. Motion of curves in three spatial dimensions using a level set approach. Journal of Computational Physics, 170(2):720–741, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967580.

Bertalmiio:2001:VPP

[BCOS01] Marcelo Bertalmío, Li-Tien Cheng, Stanley Osher, and Guillermo Sapiro. Variational problems and partial differential equations on implicit surfaces. Journal of Computational Physics, 174(2):759–780, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969372.

Benamou:2004:NMA

[BCR04] Jean-David Benamou, Francis Collino, and Olof Runborg. Numerical microlocal analysis of harmonic wavefields. Journal of Computational Physics, 199(2):717–741, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001135.

Bijl:2002:ITI

[BCVK02] Hester Bijl, Mark H. Carpenter, Veer N. Vatsa, and Christopher A. Kennedy. Implicit time integration schemes for the unsteady compressible Navier–Stokes equations: Laminar flow. Journal of Computational Physics, 179(1):313–329, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970592.

Bishop:2004:IBS

[BCZ04] T. C. Bishop, R. Cortez, and O. O. Zhmudsky. Investigation of bend and shear waves in a geometrically exact elastic rod model. Journal of Computational Physics, 193(2):642–665, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004388.

Borges:2001:FPA

[BD01] Leonardo Borges and Prabir Daripa. A fast parallel algorithm for the Poisson equation on a disk. Journal of Computational Physics, 169(1):151–192, May 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967208.

Buet:2006:APP

[BD06] Christophe Buet and Bruno Despres. Asymptotic preserving and positive schemes for radiation hydrodynamics. Journal of Computational Physics, 215(2):717–740, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005127.

Blossey:2008:SMP

[BD08] Peter N. Blossey and Dale R. Durran. Selective monotonicity preservation in scalar advection. Journal of Computational Physics, 227(10):5160–5183, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000715.

Bogey:2009:SCM

[BdCB09] Christophe Bogey, Nicolas de Cacqueray, and Christophe Bailly. A shock-capturing methodology based on adaptative spatial filtering for high-order non-linear computations. Journal of Computational Physics, 228(5):1447–1465, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005676.

Boudesocque-Dubois:2003:SCM

[BDCG03] Carine Boudesocque-Dubois, Jean-Marie Clarisse, and Serge Gauthier. A spectral Chebyshev method for linear stability analysis of one-dimensional exact solutions of gas dynamics. Journal of Computational Physics, 184(2):592–618, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000396.

Bell:2005:SAA

[BDGL05] J. B. Bell, M. S. Day, J. F. Grcar, and M. J. Lijewski. Stochastic algorithms for the analysis of numerical flame simulations. Journal of Computational Physics, 202(1):262–280, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002761.

Bassi:2009:DGM

[BDHN09] F. Bassi, C. De Bartolo, R. Hartmann, and A. Nigro. A discontinuous Galerkin method for inviscid low Mach number flows. Journal of Computational Physics, 228(11):3996–4011, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000886.

Bouzidi:2001:LBE

[BdLL01] M’hamed Bouzidi, Dominique d’Humières, Pierre Lallemand, and Li-Shi Luo. Lattice Boltzmann equation on a two-dimensional rectangular grid. Journal of Computational Physics, 172(2):704–717, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968500.

Bell:2004:ALM

[BDR^{+}04] J. B. Bell, M. S. Day, C. A. Rendleman, S. E. Woosley, and M. A. Zingale. Adaptive low Mach number simulations of nuclear flame microphysics. Journal of Computational Physics, 195(2):677–694, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005783.

Barton:2009:EAS

[BDRT09] P. T. Barton, D. Drikakis, E. Romenski, and V. A. Titarev. Exact and approximate solutions of Riemann problems in non-linear elasticity. Journal of Computational Physics, 228(18):7046–7068, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003386.

Bowers:2007:ZMP

[BDS07] Kevin J. Bowers, Ron O. Dror, and David E. Shaw. Zonal methods for the parallel execution of range-limited N-body simulations. Journal of Computational Physics, 221(1):303– 329, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106002919.

Burton:2002:AFS

[BE02] Tristan M. Burton and John K. Eaton. Analysis of a fractional-step method on overset grids. Journal of Computational Physics, 177(2):336–364, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970129.

Beale:2008:PDD

[Bea08] J. Thomas Beale. A proof that a discrete delta function is second-order accurate. Journal of Computational Physics, 227(4):2195–2197, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004974.

Bottger:2009:PFS

[BEA09] B. Böttger, J. Eiken, and M. Apel. Phase-field simulation of microstructure formation in technical castings — a selfconsistent homoenthalpic approach to the micro-macro problem. Journal of Computational Physics, 228(18):6784–6795, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003210.

Babadi:2006:AFD

[BEE06] M. Babadi, M. R. Ejtehadi, and R. Everaers. Analytical first derivatives of the RE-squared interaction potential. Journal of Computational Physics, 219(2):770–779, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002130.

Burman:2003:AFE

[BEG03] Erik Burman, Alexandre Ern, and Vincent Giovangigli. An adaptive finite element method with crosswind diffusion for low Mach, steady, laminar combustion. Journal of Computational Physics, 188(2):472–492, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001888.

Benzi:2002:PTL

[Ben02] Michele Benzi. Preconditioning techniques for large linear systems: a survey. Journal of Computational Physics, 182 (2):418–477, November 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971767.

Bruno:2009:EIE

[BEPT09] Oscar Bruno, Tim Elling, Randy Paffenroth, and Catalin Turc. Electromagnetic integral equations requiring small numbers of Krylov-subspace iterations. Journal of Computational Physics, 228(17):6169–6183, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900268X.

Berthelsen:2004:DII

[Ber04] Petter Andreas Berthelsen. A decomposed immersed interface method for variable coefficient elliptic equations with non-smooth and discontinuous solutions. Journal of Computational Physics, 197(1):364–386, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006442.

Berry:2006:NPM

[Ber06a] Ray A. Berry. Notes on the PCICE method: Simplification, generalization, and compressibility properties. Journal of Computational Physics, 215(1):6–11, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005024.

Berthon:2006:RMS

[Ber06b] Christophe Berthon. Robustness of MUSCL schemes for 2D unstructured meshes. Journal of Computational Physics, 218(2):495–509, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001161.

Berenger:2007:HAB

[Bér07] Jean-Pierre Bérenger. On the Huygens absorbing boundary conditions for electromagnetics. Journal of Computational Physics, 226(1):354–378, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001611.

Benkhaldoun:2007:WBF

[BES07] Fayssal Benkhaldoun, Imad Elmahi, and Mohammed Seaid̈. Well-balanced finite volume schemes for pollutant transport by shallow water equations on unstructured meshes. Journal of Computational Physics, 226(1):180–203, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001490.

Bettencourt:2008:FLE

[Bet08] M. T. Bettencourt. Flux limiting embedded boundary technique for electromagnetic FDTD. Journal of Computational Physics, 227(6):3141–3158, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005128.

Beyer:2009:SSE

[Bey09] Florian Beyer. A spectral solver for evolution problems with spatial S^{3}-topology. Journal of Computational Physics, 228(17):6496–6513, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003131.

Birman:2007:AGS

[BF07] A. Birman and J. Falcovitz. Application of the GRP scheme to open channel flow equations. Journal of Computational Physics, 222(1):131–154, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003470.

Berthelsen:2008:LDG

[BF08] Petter A. Berthelsen and Odd M. Faltinsen. A local directional ghost cell approach for incompressible viscous flow problems with irregular boundaries. Journal of Computational Physics, 227(9):4354–4397, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000144.

Bakosi:2008:NHM

[BFB08] J. Bakosi, P. Franzese, and Z. Boybeyi. A non-hybrid method for the PDF equations of turbulent flows on unstructured grids. Journal of Computational Physics, 227(11):5896–5935, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910800123X.

Belhamadia:2004:AMA

[BFC04a] Youssef Belhamadia, André Fortin, and Éric Chamberland. Anisotropic mesh adaptation for the solution of the Stefan problem. Journal of Computational Physics, 194(1): 233–255, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103004686.

Belhamadia:2004:TDA

[BFC04b] Youssef Belhamadia, André Fortin, and Éric Chamberland. Three-dimensional anisotropic mesh adaptation for phase change problems. Journal of Computational Physics, 201(2):753–770, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002633.

Bell:2007:ARS

[BFG07] John B. Bell, Jasmine Foo, and Alejandro L. Garcia. Algorithm refinement for the stochastic Burgers’ equation. Journal of Computational Physics, 223(1):451–468, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004505.

Bergamaschi:2008:MCP

[BFG08] Luca Bergamaschi, Massimiliano Ferronato, and Giuseppe Gambolati. Mixed constraint preconditioners for the iterative solution of FE coupled consolidation equations. Journal of Computational Physics, 227(23):9885–9897, December 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800421X.

Becache:2003:SPM

[BFJ03] E. Bécache, S. Fauqueux, and P. Joly. Stability of perfectly matched layers, group velocities and anisotropic waves. Journal of Computational Physics, 188(2):399–433, July 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001840.

Baruch:2007:HON

[BFT07] G. Baruch, G. Fibich, and S. Tsynkov. High-order numerical method for the nonlinear Helmholtz equation with material discontinuities in one space dimension. Journal of Computational Physics, 227(1):820–850, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003841.

Baruch:2009:HON

[BFT09] G. Baruch, G. Fibich, and S. Tsynkov. A high-order numerical method for the nonlinear Helmholtz equation in multidimensional layered media. Journal of Computational Physics, 228(10):3789–3815, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000813.

Bokil:2005:OSS

[BG05a] Vrushali A. Bokil and Roland Glowinski. An operator splitting scheme with a distributed Lagrange multiplier based fictitious domain method for wave propagation problems. Journal of Computational Physics, 205(1):242–268, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004668.

Bustinza:2005:MLD

[BG05b] Rommel Bustinza and Gabriel N. Gatica. A mixed local discontinuous Galerkin method for a class of nonlinear problems in fluid mechanics. Journal of Computational Physics, 207(2):427–456, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000367.

Barros:2007:GFD

[BG07] Saulo R. M. Barros and Claudia I. Garcia. A global finitedifference semi-Lagrangian model for the adiabatic primitive equations. Journal of Computational Physics, 226(2): 1645–1667, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107002550.

Bonnet:2009:EWI

[BG09] Marc Bonnet and Bojan B. Guzina. Elastic-wave identification of penetrable obstacles using shape-material sensitivity framework. Journal of Computational Physics, 228(2):294–311, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004798.

Bruger:2005:HOA

[BGLN05] Arnim Brüger, Bertil Gustafsson, Per Lötstedt, and Jonas Nilsson. High order accurate solution of the incompressible Navier–Stokes equations. Journal of Computational Physics, 203(1):49–71, February 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003171.

Bejanov:2008:GAF

[BGM08] B. Bejanov, J. L. Guermond, and P. D. Minev. A grid-alignment finite element technique for incompressible multicomponent flows. Journal of Computational Physics, 227(13):6473–6489, June 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001563.

Brufau:2003:UFS

[BGN03] P. Brufau and P. García-Navarro. Unsteady free surface flow simulation over complex topography with a multidimensional upwind technique. Journal of Computational Physics, 186(2):503–526, April 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300072X.

Barrett:2007:PFE

[BGN07] John W. Barrett, Harald Garcke, and Robert Nürnberg. A parametric finite element method for fourth order geometric evolution equations. Journal of Computational Physics, 222(1):441–467, March 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003615.

Barrett:2008:PFE

[BGN08] John W. Barrett, Harald Garcke, and Robert Nürnberg. On the parametric finite element approximation of evolving hypersurfaces in R^{3}. Journal of Computational Physics, 227 (9):4281–4307, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107005141.

Bedwani:2008:NAM

[BGR08] Stéphane Bedwani, François Guibault, and Alain Rochefort. Nanoscale adaptive meshing for rapid STM imaging. Journal of Computational Physics, 227(14):6720–6726, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002040.

Borazjani:2008:CIB

[BGS08] Iman Borazjani, Liang Ge, and Fotis Sotiropoulos. Curvilinear immersed boundary method for simulating fluid structure interaction with complex 3D rigid bodies. Journal of Computational Physics, 227(16):7587–7620, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002490.

Bruno:2004:EPH

[BH04] Oscar P. Bruno and E. McKay Hyde. An efficient, preconditioned, high-order solver for scattering by twodimensional inhomogeneous media. Journal of Computational Physics, 200(2):670–694, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001846.

Brunner:2005:TDT

[BH05] Thomas A. Brunner and James Paul Holloway. Twodimensional time dependent Riemann solvers for neutron transport. Journal of Computational Physics, 210(1):386– 399, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002275.

Bernstein:2009:RRH

[BH09] J. P. Bernstein and P. A. Hughes. Refining a relativistic, hydrodynamic solver: Admitting ultra-relativistic flows. Journal of Computational Physics, 228(17):6212–6230, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910900271X.

Bansch:2004:FEM

[BHL^{+}04] Eberhard Bänsch, Frank Haußer, Omar Lakkis, Bo Li, and Axel Voigt. Finite element method for epitaxial growth with attachment-detachment kinetics. Journal of Computational Physics, 194(2):409–434, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004893.

Bao:2007:ISC

[BHL07] Gang Bao, Songming Hou, and Peijun Li. Inverse scattering by a continuation method with initial guesses from a direct imaging algorithm. Journal of Computational Physics, 227(1):755–762, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003567.

Bermudez:2007:OPM

[BHNPR07] A. Bermúdez, L. Hervella-Nieto, A. Prieto, and R. Rodríguez. An optimal perfectly matched layer with unbounded absorbing function for time-harmonic acoustic scattering problems. Journal of Computational Physics, 223(2):469–488, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106004487.

Bruno:2007:AHO

[BHP07] Oscar P. Bruno, Youngae Han, and Matthew M. Pohlman. Accurate, high-order representation of complex three-dimensional surfaces via Fourier continuation analysis. Journal of Computational Physics, 227(2):1094–1125, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003816.

Beaudoin:2003:SAD

[BHR03] A. Beaudoin, S. Huberson, and E. Rivoalen. Simulation of anisotropic diffusion by means of a diffusion velocity method. Journal of Computational Physics, 186(1):122–135, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910300024X.

Burger:2004:ITD

[BHR04] Martin Burger, Benjamin Hackl, and Wolfgang Ring. Incorporating topological derivatives into level set methods. Journal of Computational Physics, 194(1):344–362, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004868.

Beaudoin:2006:NSS

[BHR06] Anthony Beaudoin, Serge Huberson, and Elie Rivoalen. From Navier–Stokes to Stokes by means of particle methods. Journal of Computational Physics, 214(1):264–283, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004341.

Bao:2003:ODN

[BHS03] Gang Bao, Kai Huang, and Gunther Schmidt. Optimal design of nonlinear diffraction gratings. Journal of Computational Physics, 184(1):106–121, January 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000190.

Banks:2009:EFM

[BHS09] J. W. Banks, W. D. Henshaw, and J. N. Shadid. An evaluation of the FCT method for high-speed flows on structured overlapping grids. Journal of Computational Physics, 228(15):5349–5369, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002071.

Burger:2007:LSA

[BHSV07] Martin Burger, Frank Haußer, Christina Stöcker, and Axel Voigt. A level set approach to anisotropic flows with curvature regularization. Journal of Computational Physics, 225(1):183– 205, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005833.

Botchev:2006:GTS

[BHvdV06] M. A. Botchev, D. Harutyunyan, and J. J. W. van der Vegt. The Gautschi time stepping scheme for edge finite element discretizations of the Maxwell equations. Journal of Computational Physics, 216(2):654–686, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000052.

Bialecki:2003:FSO

[Bia03] Bernard Bialecki. A fast solver for the orthogonal spline collocation solution of the biharmonic Dirichlet problem on rectangles. Journal of Computational Physics, 191(2):601– 621, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003425.

Billet:2005:ICC

[Bil05] G. Billet. Improvement of convective concentration fluxes in a one step reactive flow solver. Journal of Computational Physics, 204(1):319–352, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400419X.

Berland:2007:CPS

[BIS07] Håvard Berland, Alvaro L. Islas, and Constance M. Schober. Conservation of phase space properties using exponential integrators on the cubic Schrödinger equation. Journal of Computational Physics, 225(1):284–299, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005900.

Beux:2001:ARE

[BISS01] F. Beux, A. Iollo, M. V. Salvetti, and A. Soldati. Approximation and reconstruction of the electrostatic field in wire-plate precipitators by a low-order model. Journal of Computational Physics, 170(2):893–916, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967737.

Brucker:2007:EAS

[BIVC07] Kyle A. Brucker, Juan C. Isaza, T. Vaithianathan, and Lance R. Collins. Efficient algorithm for simulating homogeneous turbulent shear flow without remeshing. Journal of Computational Physics, 225(1):20–32, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004827.

Badea:2004:DDM

[BIW04] Lori Badea, Ioan R. Ionescu, and Sylvie Wolf. Domain decomposition method for dynamic faulting under slip-dependent friction. Journal of Computational Physics, 201(2):487–510, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002396.

Badea:2008:SME

[BIW08] Lori Badea, Ioan R. Ionescu, and Sylvie Wolf. Schwarz method for earthquake source dynamics. Journal of Computational Physics, 227(8):3824–3848, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005323.

Bao:2000:RPM

[BJ00] Weizhu Bao and Shi Jin. The random projection method for hyperbolic conservation laws with stiff reaction terms. Journal of Computational Physics, 163(1):216–248, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965720.

Bao:2002:RPM

[BJ02] Weizhu Bao and Shi Jin. The random projection method for stiff multispecies detonation capturing. Journal of Computational Physics, 178(1):37–57, May 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970166.

Bonfigli:2009:EMS

[BJ09] Giuseppe Bonfigli and Patrick Jenny. An efficient multiscale Poisson solver for the incompressible Navier–Stokes equations with immersed boundaries. Journal of Computational Physics, 228(12):4568–4587, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001569.

Bao:2002:TSS

[BJM02] Weizhu Bao, Shi Jin, and Peter A. Markowich. On timesplitting spectral approximations for the schrödinger equation in the semiclassical regime. Journal of Computational Physics, 175(2):487–524, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969566.

Bao:2003:NSG

[BJM03] Weizhu Bao, Dieter Jaksch, and Peter A. Markowich. Numerical solution of the Gross–Pitaevskii equation for Bose–Einstein condensation. Journal of Computational Physics, 187(1):318–342, May 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001025.

Burman:2004:AFE

[BJP04] E. Burman, A. Jacot, and M. Picasso. Adaptive finite elements with high aspect ratio for the computation of coalescence using a phase-field model. Journal of Computational Physics, 195(1):153–174, March 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005151.

Bruno:2001:FHO

[BK01] Oscar P. Bruno and Leonid A. Kunyansky. A fast, highorder algorithm for the solution of surface scattering problems: Basic implementation, tests, and applications. Journal of Computational Physics, 169(1):80–110, May 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967142.

Burger:2007:AMW

[BK07] Raimund Bürger and Alice Kozakevicius. Adaptive multiresolution WENO schemes for multi-species kinematic flow models. Journal of Computational Physics, 224(2):1190–1222, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005729.

Bialecki:2008:SCF

[BK08] Bernard Bialecki and Andreas Karageorghis. Spectral Chebyshev–Fourier collocation for the Helmholtz and variable coefficient equations in a disk. Journal of Computational Physics, 227(19):8588–8603, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003318.

Botta:2004:WBF

[BKLL04] N. Botta, R. Klein, S. Langenberg, and S. Lützenkirchen. Well balanced finite volume methods for nearly hydrostatic flows. Journal of Computational Physics, 196(2):539–565, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006077.

Beylkin:2009:FCF

[BKM09] Gregory Beylkin, Christopher Kurcz, and Lucas Monzón. Fast convolution with the free space Helmholtz Green’s function. Journal of Computational Physics, 228(8):2770–2791, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006566.

Bates:2001:CTI

[BKR^{+}01] J. W. Bates, D. A. Knoll, W. J. Rider, R. B. Lowrie, and V. A. Mousseau. On consistent time-integration methods for radiation hydrodynamics in the equilibrium diffusion limit: Low-energy-density regime. Journal of Computational Physics, 167(1):99–130, February 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966610.

Banda:2007:LBR

[BKS07] Mapundi K. Banda, Axel Klar, and Mohammed Seaïd. A lattice-Boltzmann relaxation scheme for coupled convectionradiation systems. Journal of Computational Physics, 226 (2):1408–1431, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107002409.

Barone:2009:SGR

[BKST09] Matthew F. Barone, Irina Kalashnikova, Daniel J. Segalman, and Heidi K. Thornquist. Stable Galerkin reduced order models for linearized compressible flow. Journal of Computational Physics, 228(6):1932–1946, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006098.

Billet:2001:ALI

[BL01] G. Billet and O. Louedin. Adaptive limiters for improving the accuracy of the MUSCL approach for unsteady flows. Journal of Computational Physics, 170(1):161–183, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967312.

Bryson:2003:HOS

[BL03] Steve Bryson and Doron Levy. High-order semi-discrete central-upwind schemes for multi-dimensional Hamilton– Jacobi equations. Journal of Computational Physics, 189 (1):63–87, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103002018.

Bao:2004:ESN

[BL04] Weizhu Bao and Xiang-Gui Li. An efficient and stable numerical method for the Maxwell–Dirac system. Journal of Computational Physics, 199(2):663–687, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001111.

Brown:2005:MEE

[BL05] J. David Brown and Lisa L. Lowe. Multigrid elliptic equation solver with adaptive mesh refinement. Journal of Computational Physics, 209(2):582–598, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001890.

Berkenbos:2008:AMI

[BL08] A. Berkenbos and C. P. Lowe. Accurate method for including solid-fluid boundary interactions in mesoscopic model fluids. Journal of Computational Physics, 227(9):4589–4599, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000375.

Bao:2009:NSI

[BL09a] Gang Bao and Peijun Li. Numerical solution of an inverse medium scattering problem for Maxwell’s equations at fixed frequency. Journal of Computational Physics, 228 (12):4638–4648, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001600.

Beale:2009:VDA

[BL09b] J. Thomas Beale and Anita T. Layton. A velocity decomposition approach for moving interfaces in viscous fluids. Journal of Computational Physics, 228(9):3358–3367, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000400.

Bhagatwala:2009:MAV

[BL09c] Ankit Bhagatwala and Sanjiva K. Lele. A modified artificial viscosity approach for compressible turbulence simulations. Journal of Computational Physics, 228(14):4965–4969, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002034.

Blayo:2000:CFD

[Bla00] Eric Blayo. Compact finite difference schemes for ocean models: 1. ocean waves. Journal of Computational Physics, 164(2):241–257, November 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965653.

Besse:2008:WMB

[BLG^{+}08] Nicolas Besse, Guillaume Latu, Alain Ghizzo, Eric Sonnendrücker, and Pierre Bertrand. A wavelet-MRA-based adaptive semi-Lagrangian method for the relativistic Vlasov– Maxwell system. Journal of Computational Physics, 227 (16):7889–7916, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108002672.

Bond:2003:AFK

[BLL03] Stephen D. Bond, Brian B. Laird, and Benedict J. Leimkuhler. On the approximation of Feynman–Kac path integrals. Journal of Computational Physics, 185(2):472–483, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000669.

Bourlioux:2003:HOM

[BLM03] Anne Bourlioux, Anita T. Layton, and Michael L. Minion. High-order multi-implicit spectral deferred correction methods for problems of reactive flow. Journal of Computational Physics, 189(2):651–675, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002511.

Bramkamp:2004:AMF

[BLM04] F. Bramkamp, Ph. Lamby, and S. Müller. An adaptive multiscale finite volume solver for unsteady and steady state flow computations. Journal of Computational Physics, 197(2):460– 490, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006478.

Bennoune:2008:USN

[BLM08] Mounir Bennoune, Mohammed Lemou, and Luc Mieussens. Uniformly stable numerical schemes for the Boltzmann equation preserving the compressible Navier–Stokes asymptotics. Journal of Computational Physics, 227(8):3781–3803, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005268.

Bailey:2008:ECD

[BLS08] A. G. Bailey, C. P. Lowe, and A. P. Sutton. Efficient constraint dynamics using MILC SHAKE. Journal of Computational Physics, 227(20):8949–8959, October 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003719.

Bonnet:2001:GAI

[BLW01] Frédéric D. R. Bonnet, Derek B. Leinweber, and Anthony G. Williams. General algorithm for improved lattice actions on parallel computing architectures. Journal of Computational Physics, 170(1):1–17, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101966999.

Bao:2004:NSG

[BLW04] Jing-Dong Bao, Rong-Wu Li, and Wei Wu. Numerical simulations of generalized Langevin equations with deeply asymptotic parameters. Journal of Computational Physics, 197(1):241–252, June 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103006284.

Berezovski:2001:STW

[BM01a] A. Berezovski and G. A. Maugin. Simulation of thermoelastic wave propagation by means of a composite wave-propagation algorithm. Journal of Computational Physics, 168(1):249–264, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101966975.

Bertolazzi:2001:TBU

[BM01b] Enrico Bertolazzi and Gianmarco Manzini. A trianglebased unstructured finite-volume method for chemically reactive hypersonic flows. Journal of Computational Physics, 166(1):84–115, January 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101966446.

Blanes:2001:SMN

[BM01c] S. Blanes and P. C. Moan. Splitting methods for nonautonomous Hamiltonian equations. Journal of Computational Physics, 170(1):205–230, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967336.

Bagdonat:2002:HSC

[BM02] T. Bagdonat and U. Motschmann. 3D hybrid simulation code using curvilinear coordinates. Journal of Computational Physics, 183(2):470–485, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972037.

Berrut:2005:OPS

[BM05] Jean-Paul Berrut and Hans D. Mittelmann. Optimized point shifts and poles in the linear rational pseudospectral method for boundary value problems. Journal of Computational Physics, 204(1):292–301, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004176.

Barranco:2006:SAH

[BM06] Joseph A. Barranco and Philip S. Marcus. A 3D spectral anelastic hydrodynamic code for shearing, stratified flows. Journal of Computational Physics, 219(1):21–46, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001379.

Breil:2007:CCD

[BM07] Jérôme Breil and Pierre-Henri Maire. A cell-centered diffusion scheme on two-dimensional unstructured meshes. Journal of Computational Physics, 224(2):785–823, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005493.

Brooks:2005:SIM

[BMDS05] Eugene D. Brooks III, Michael Scott McKinley, Frank Daffin, and Abraham Szőke. Symbolic implicit Monte Carlo radiation transport in the difference formulation: a piecewise constant discretization. Journal of Computational Physics, 205(2):737–754, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004899.

Bulashevich:2006:SVU

[BMK^{+}06] K. A. Bulashevich, V. F. Mymrin, S. Yu. Karpov, I. A. Zhmakin, and A. I. Zhmakin. Simulation of visible and ultraviolet group-III nitride light emitting diodes. Journal of Computational Physics, 213(1):214–238, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003785.

Bansch:2005:FEM

[BMN05] Eberhard Bänsch, Pedro Morin, and Ricardo H. Nochetto. A finite element method for surface diffusion: the parametric case. Journal of Computational Physics, 203(1):321– 343, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003304.

BenAbdallah:2007:AAS

[BMN07] N. Ben Abdallah, M. Mouis, and C. Negulescu. An accelerated algorithm for 2D simulations of the quantum ballistic transport in nanoscale MOSFETs. Journal of Computational Physics, 225(1):74–99, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600578X.

Biava:2002:WFU

[BMQS02] M. Biava, D. Modugno, L. Quartapelle, and M. Stoppelli. Weak ψ–ω formulation for unsteady flows in 2D multiply connected domains. Journal of Computational Physics, 177(2):209–232, April 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969761.

Beckett:2001:MMF

[BMR01] G. Beckett, J. A. Mackenzie, and M. L. Robertson. A moving mesh finite element method for the solution of two-dimensional Stefan problems. Journal of Computational Physics, 168(2):500–518, April 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196721X.

Beckett:2001:NSO

[BMRS01] G. Beckett, J. A. Mackenzie, A. Ramage, and D. M. Sloan. On the numerical solution of one-dimensional PDEs using adaptive methods based on equidistribution. Journal of Computational Physics, 167(2):372–392, March 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966798.

Beckett:2002:CST

[BMRS02] G. Beckett, J. A. Mackenzie, A. Ramage, and D. M. Sloan. Computational solution of two-dimensional unsteady PDEs using moving mesh methods. Journal of Computational Physics, 182(2):478–495, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971792.

Bosisio:2000:DSE

[BMS00] F. Bosisio, S. Micheletti, and R. Sacco. A discretization scheme for an extended drift-diffusion model including trap-assisted phenomena. Journal of Computational Physics, 159(2):197–212, April 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964283.

Basko:2009:ECC

[BMT09] M. M. Basko, J. Maruhn, and An. Tauschwitz. An efficient cell-centered diffusion scheme for quadrilateral grids. Journal of Computational Physics, 228(6):2175–2193, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006232.

Beji:2004:FDN

[BN04] S. Beji and K. Nadaoka. Fully dispersive nonlinear water wave model in curvilinear coordinates. Journal of Computational Physics, 198(2):645–658, August 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000579.

Braconnier:2009:ASR

[BN09] Benjamin Braconnier and Boniface Nkonga. An all-speed relaxation scheme for interface flows with surface tension. Journal of Computational Physics, 228(16):5722–5739, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002319.

Bonnaillie-Noel:2006:CSS

[BNNP06] Virginie Bonnaillie-Noël, Francis Nier, and Yassine Patel. Computing the steady states for an asymptotic model of quantum transport in resonant heterostructures. Journal of Computational Physics, 219(2):644–670, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001975.

Badia:2008:FSP

[BNV08] Santiago Badia, Fabio Nobile, and Christian Vergara. Fluidstructure partitioned procedures based on Robin transmission conditions. Journal of Computational Physics, 227(14): 7027–7051, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108002192.

Borm:2004:FEB

[BO04] S. Börm and J. Ostrowski. Fast evaluation of boundary integral operators arising from an eddy current problem. Journal of Computational Physics, 193(1):67–85, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004194.

Balandin:2005:MSE

[BO05] A. L. Balandin and Y. Ono. The method of series expansion for 3-D vector tomography reconstruction. Journal of Computational Physics, 202(1):52–64, January 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002670.

Bogaert:2009:LFS

[BO09] Ignace Bogaert and Femke Olyslager. A low frequency stable plane wave addition theorem. Journal of Computational Physics, 228(4):1000–1016, March 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005354.

Bodony:2006:ASZ

[Bod06] Daniel J. Bodony. Analysis of sponge zones for computational fluid mechanics. Journal of Computational Physics, 212(2):681–702, March 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003517.

Boersma:2005:SCF

[Boe05] Bendiks Jan Boersma. A staggered compact finite difference formulation for the compressible Navier–Stokes equations. Journal of Computational Physics, 208(2):675–690, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001257.

Bui:2006:FAM

[BOK^{+}06] T. T. Bui, E. T. Ong, B. C. Khoo, E. Klaseboer, and K. C. Hung. A fast algorithm for modeling multiple bubbles dynamics. Journal of Computational Physics, 216(2):430–453, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005590.

Bonaventura:2000:SIS

[Bon00] Luca Bonaventura. A semi-implicit semi-Lagrangian scheme using the height coordinate for a nonhydrostatic and fully elastic model of atmospheric flows. Journal of Computational Physics, 158(2):186–213, March 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964148.

Borici:2000:LAI

[Bor00] Artan Boriçi. A Lanczos approach to the inverse square root of a large and sparse matrix. Journal of Computational Physics, 162(1):123–131, July 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096529X.

Borici:2003:CMU

[Bor03] Artan Boriçi. Computational methods for UV-suppressed fermions. Journal of Computational Physics, 189(2):454–462, August 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002274.

Boronski:2007:SMM

[Bor07] Piotr Boronski. Spectral method for matching exterior and interior elliptic problems. Journal of Computational Physics, 225(1):449–463, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006000.

Borve:2005:RSP

[BOT05] S. Børve, M. Omang, and J. Trulsen. Regularized smoothed particle hydrodynamics with improved multi-resolution handling. Journal of Computational Physics, 208(1):345–367, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105001154.

Botha:2006:SVI

[Bot06] Matthys M. Botha. Solving the volume integral equations of electromagnetic scattering. Journal of Computational Physics, 218(1):141–158, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000763.

Bowers:2001:APC

[Bow01] K. J. Bowers. Accelerating a particle-in-cell simulation using a hybrid counting sort. Journal of Computational Physics, 173(2):393–411, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968512.

Boyd:2002:CNA

[Boy02a] John P. Boyd. A comparison of numerical algorithms for Fourier extension of the first, second, and third kinds. Journal of Computational Physics, 178(1):118–160, May 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970233.

Boyd:2002:DRQ

[Boy02b] John P. Boyd. Deleted residuals, the QR-factored Newton iteration, and other methods for formally overdetermined determinate discretizations of nonlinear eigenproblems for solitary, cnoidal, and shock waves. Journal of Computational Physics, 179(1):216–237, June 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297052X.

Boyd:2003:LPM

[Boy03] John P. Boyd. A Legendre-pseudospectral method for computing travelling waves with corners (slope discontinuities) in one space dimension with application to Whitham’s equation family. Journal of Computational Physics, 189(1):98–110, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002031.

Boyd:2004:PSW

[Boy04] John P. Boyd. Prolate spheroidal wavefunctions as an alternative to Chebyshev and Legendre polynomials for spectral element and pseudospectral algorithms. Journal of Computational Physics, 199(2):688–716, September 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001123.

Boyd:2005:CRC

[Boy05a] John P. Boyd. A Chebyshev/rational Chebyshev spectral method for the Helmholtz equation in a sector on the surface of a sphere: defeating corner singularities. Journal of Computational Physics, 206(1):302–310, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005145.

Boyd:2005:TGR

[Boy05b] John P. Boyd. Trouble with Gegenbauer reconstruction for defeating Gibbs’ phenomenon: Runge phenomenon in the diagonal limit of Gegenbauer polynomial approximations. Journal of Computational Physics, 204(1):253–264, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004152.

Boyd:2006:PDS

[Boy06] John P. Boyd. A proof that the discrete singular convolution (DSC)/Lagrange-distributed approximating function (LDAF) method is inferior to high order finite differences. Journal of Computational Physics, 214(2):538–549, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004602. See comment

[WZ07].

Boyd:2009:AAC

[Boy09] John P. Boyd. Acceleration of algebraically-converging Fourier series when the coefficients have series in powers of 1/n. Journal of Computational Physics, 228(5):1404–1411, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005627.

Botchorishvili:2003:FVS

[BP03] Ramaz Botchorishvili and Olivier Pironneau. Finite volume schemes with equilibrium type discretization of source terms for scalar conservation laws. Journal of Computational Physics, 187(2):391–427, May 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300086X.

Brooks:2004:KLL

[BP04a] Gregory P. Brooks and Joseph M. Powers. A Karhunen– Loève least-squares technique for optimization of geometry of a blunt body in supersonic flow. Journal of Computational Physics, 195(1):387–412, March 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005394.

Brooks:2004:SPF

[BP04b] Gregory P. Brooks and Joseph M. Powers. Standardized pseudospectral formulation of the inviscid supsersonic blunt body problem. Journal of Computational Physics, 197(1):58–85, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006193.

BenAbdallah:2006:MST

[BP06] Naoufel Ben Abdallah and Olivier Pinaud. Multiscale simulation of transport in an open quantum system: Resonances and WKB interpolation. Journal of Computational Physics, 213(1):288–310, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003815.

Buoni:2007:ESN

[BP07] Matthew Buoni and Linda Petzold. An efficient, scalable numerical algorithm for the simulation of electrochemical systems on irregular domains. Journal of Computational Physics, 225(2):2320–2332, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001349.

Bringley:2008:VSM

[BP08] Thomas T. Bringley and Charles S. Peskin. Validation of a simple method for representing spheres and slender bodies in an immersed boundary method for Stokes flow on an unbounded domain. Journal of Computational Physics, 227(11):5397–5425, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800082X.

Bernardini:2009:GSO

[BP09] Matteo Bernardini and Sergio Pirozzoli. A general strategy for the optimization of Runge–Kutta schemes for wave propagation phenomena. Journal of Computational Physics, 228(11):4182–4199, June 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001077.

Bonito:2006:NSV

[BPL06] Andrea Bonito, Marco Picasso, and Manuel Laso. Numerical simulation of 3D viscoelastic flows with free surfaces. Journal of Computational Physics, 215(2):691–716, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005115.

Beran:2006:UQL

[BPM06] Philip S. Beran, Chris L. Pettit, and Daniel R. Millman. Uncertainty quantification of limit-cycle oscillations. Journal of Computational Physics, 217(1):217–247, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001872.

Bass:2008:SRM

[BPMR08] Alexander Bass, Seth Putterman, Barry Merriman, and Steven J. Ruuth. Symmetry reduction for molecular dynamics simulation of an imploding gas bubble. Journal of Computational Physics, 227(3):2118–2129, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004688.

Bogaert:2007:FAF

[BPO07] Ignace Bogaert, Davy Pissoort, and Femke Olyslager. A faster aggregation for 3D fast evanescent wave solvers using rotations. Journal of Computational Physics, 227(1): 557–573, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107003476.

Bergamaschi:2003:CES

[BPS03] Luca Bergamaschi, Giorgio Pini, and Flavio Sartoretto. Computational experience with sequential and parallel, preconditioned Jacobi–Davidson for large, sparse symmetric matrices. Journal of Computational Physics, 188(1):318–331, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001906.

Badia:2009:CBN

[BQQ09] Santiago Badia, Annalisa Quaini, and Alfio Quarteroni. Coupling Biot and Navier–Stokes equations for modelling fluid-poroelastic media interaction. Journal of Computational Physics, 228(21):7986–8014, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004069.

Balakrishnan:2001:OIM

[BR01] Karthik Balakrishnan and Palghat A. Ramachandran. Osculatory interpolation in the method of fundamental solution for nonlinear Poisson problems. Journal of Computational Physics, 172(1):1–18, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967968.

Bajpayi:2009:FVD

[BR09a] Mayank Bajpayi and S. V. Raghurama Rao. A finite variable difference relaxation scheme for hyperbolicparabolic equations. Journal of Computational Physics, 228(20):7513–7542, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900360X.

Bjontegaard:2009:AIT

[BR09b] Tormod Bjøntegaard and Einar M. Rønquist. Accurate interface-tracking for arbitrary Lagrangian–Eulerian schemes. Journal of Computational Physics, 228(12):4379–4399, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001326.

Brackbill:2001:A

[Bra01] J. U. Brackbill. Announcement. Journal of Computational Physics, 170(2):919–920, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968408.

Brackbill:2004:IHS

[Bra04] J. U. Brackbill. Introduction to Harlow’s scientific memoir. Journal of Computational Physics, 195(2):413, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005680.

Brackbill:2008:BCM

[Bra08] J. U. Brackbill. Boundary conditions for Maxwell solvers. Journal of Computational Physics, 227(14):6715–6719, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002003.

Boyd:2003:PMS

[BRB03] John P. Boyd, C. Rangan, and P. H. Bucksbaum. Pseudospectral methods on a semi-infinite interval with application to the hydrogen atom: a comparison of the mapped Fouriersine method with Laguerre series and rational Chebyshev expansions. Journal of Computational Physics, 188(1):56–74, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910300127X.

Bernard:2009:HOD

[BRC^{+}09] P.-E. Bernard, J.-F. Remacle, R. Comblen, V. Legat, and K. Hillewaert. High-order discontinuous Galerkin schemes on general 2D manifolds applied to the shallow water equations. Journal of Computational Physics, 228(17):6514–6535, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003143.

Balsara:2009:EHA

[BRDM09] Dinshaw S. Balsara, Tobias Rumpf, Michael Dumbser, and Claus-Dieter Munz. Efficient, high accuracy ADER–WENO schemes for hydrodynamics and divergence-free magnetohydrodynamics. Journal of Computational Physics, 228(7): 2480–2516, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006396.

Bhaganagar:2002:DNS

[BRL02] Kiran Bhaganagar, Dietmar Rempfer, and John Lumley. Direct numerical simulation of spatial transition to turbulence using fourth-order vertical velocity second-order vertical vorticity formulation. Journal of Computational Physics, 180(1):200–228, July 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970889.

Bonet:2005:HFV

[BRP05] Javier Bonet and Miguel X. Rodríguez-Paz. Hamiltonian formulation of the variable-h SPH equations. Journal of Computational Physics, 209(2):541–558, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001865.

Balsara:2000:MPW

[BS00a] Dinshaw S. Balsara and Chi-Wang Shu. Monotonicity preserving weighted essentially non-oscillatory schemes with increasingly high order of accuracy. Journal of Computational Physics, 160(2):405–452, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096443X.

Barichello:2000:SPR

[BS00b] L. B. Barichello and C. E. Siewert. The searchlight problem for radiative transfer in a finite slab. Journal of Computational Physics, 157(2):707–726, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963969.

Benamou:2000:EMC

[BS00c] Jean-David Benamou and Ian Solliec. An Eulerian method for capturing caustics. Journal of Computational Physics, 162(1):132–163, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100965306.

Betsch:2000:IEC

[BS00d] P. Betsch and P. Steinmann. Inherently energy conserving time finite elements for classical mechanics. Journal of Computational Physics, 160(1):88–116, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964271.

Brown:2000:IMA

[BS00e] D. J. Brown and R. M. Stringfield. Iterative methods applied to matrix equations found in calculating spheroidal functions. Journal of Computational Physics, 159(2):329–343, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964489.

Bessonov:2001:AVC

[BS01] Nicholas M. Bessonov and Dong Joo Song. Application of vector calculus to numerical simulation of continuum mechanics problems. Journal of Computational Physics, 167(1):22–38, February 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966531.

Besse:2003:SLS

[BS03a] N. Besse and E. Sonnendrücker. Semi-Lagrangian schemes for the Vlasov equation on an unstructured mesh of phase space. Journal of Computational Physics, 191(2):341–376, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003188.

Blackburn:2003:SEF

[BS03b] H. M. Blackburn and S. Schmidt. Spectral element filtering techniques for large eddy simulation with dynamic estimation. Journal of Computational Physics, 186(2):610–629, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000883.

Ballestra:2004:NPS

[BS04a] Luca V. Ballestra and Riccardo Sacco. Numerical problems in semiconductor simulation using the hydrodynamic model: a second-order finite difference scheme. Journal of Computational Physics, 195(1):320–340, March 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005345.

Bernard:2004:GFD

[BS04b] Peter S. Bernard and Jun Shen. A grid-free dilatation element method for quasi-one-dimensional gas dynamics. Journal of Computational Physics, 199(1):41–65, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104000646.

Blackburn:2004:FGS

[BS04c] H. M. Blackburn and S. J. Sherwin. Formulation of a Galerkin spectral element–Fourier method for three-dimensional incompressible flows in cylindrical geometries. Journal of Computational Physics, 197(2):759–778, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000907.

Borisov:2004:AWP

[BS04d] Andrei G. Borisov and Sergei V. Shabanov. Applications of the wave packet method to resonant transmission and reflection gratings. Journal of Computational Physics, 199(2):742–762, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001275.

Borisov:2005:LPM

[BS05] Andrei G. Borisov and Sergei V. Shabanov. Lanczos pseudospectral method for initial-value problems in electrodynamics and its applications to ionic crystal gratings. Journal of Computational Physics, 209(2):643–664, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001920.

Borisov:2006:WPP

[BS06a] Andrei G. Borisov and Sergei V. Shabanov. Wave packet propagation by the Faber polynomial approximation in electrodynamics of passive media. Journal of Computational Physics, 216(1):391–402, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005577.

Bourgeade:2006:NMB

[BS06b] Antoine Bourgeade and Olivier Saut. Numerical methods for the bidimensional Maxwell–Bloch equations in nonlinear crystals. Journal of Computational Physics, 213(2):823–843, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004183.

Bhan:2007:CHM

[BS07] Katherine Bhan and Jerome Spanier. Condensed history Monte Carlo methods for photon transport problems. Journal of Computational Physics, 225(2):1673–1694, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000770.

Bao:2008:GLH

[BS08a] Weizhu Bao and Jie Shen. A generalized-Laguerre–Hermite pseudospectral method for computing symmetric and central vortex states in Bose–Einstein condensates. Journal of Computational Physics, 227(23):9778–9793, December 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004063.

Beale:2008:LCS

[BS08b] J. Thomas Beale and John Strain. Locally corrected semi-Lagrangian methods for Stokes flow with moving elastic interfaces. Journal of Computational Physics, 227(8):3896–3920, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005402.

Bereau:2009:OCM

[BS09a] Tristan Bereau and Robert H. Swendsen. Optimized convergence for multiple histogram analysis. Journal of Computational Physics, 228(17):6119–6129, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002630.

Branicio:2009:LSC

[BS09b] Paulo S. Branicio and David J. Srolovitz. Local stress calculation in simulations of multicomponent systems. Journal of Computational Physics, 228(22):8467–8479, December 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004653.

Burbeau:2001:PIL

[BSB01] A. Burbeau, P. Sagaut, and Ch.-H. Bruneau. A problemindependent limiter for high-order Runge–Kutta discontinuous Galerkin methods. Journal of Computational Physics, 169(1):111–150, May 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196718X.

Bdzil:2001:PBA

[BSJ01] J. B. Bdzil, D. S. Stewart, and T. L. Jackson. Program burn algorithms based on detonation shock dynamics: Discrete approximations of detonation flows with discontinuous front models. Journal of Computational Physics, 174(2):870–902, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969426.

Banks:2007:HRG

[BSKH07] J. W. Banks, D. W. Schwendeman, A. K. Kapila, and W. D. Henshaw. A high-resolution Godunov method for compressible multi-material flow on overlapping grids. Journal of Computational Physics, 223(1):262–297, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004360.

Bourantas:2009:ASE

[BSLN09] G. C. Bourantas, E. D. Skouras, V. C. Loukopoulos, and G. C. Nikiforidis. An accurate, stable and efficient domain-type meshless method for the solution of MHD flow problems. Journal of Computational Physics, 228(21):8135–8160, November 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004227.

Brooks:2006:PLD

[BSP06] Eugene D. Brooks III, Abraham Szőke, and Jayson D. L. Peterson. Piecewise linear discretization of symbolic implicit Monte Carlo radiation transport in the difference formulation. Journal of Computational Physics, 220(1):471–497, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106003433.

Bateman:2001:ENS

[BST01] W. J. D. Bateman, C. Swan, and P. H. Taylor. On the efficient numerical simulation of directionally spread surface water waves. Journal of Computational Physics, 174(1):277–305, November 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969062.

Bateman:2003:CWP

[BST03] W. J. D. Bateman, C. Swan, and P. H. Taylor. On the calculation of the water particle kinematics arising in a directionally spread wavefield. Journal of Computational Physics, 186(1):70–92, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000123.

Bao:2003:NMG

[BSW03] Weizhu Bao, Fangfang Sun, and G. W. Wei. Numerical methods for the generalized Zakharov system. Journal of Computational Physics, 190(1):201–228, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002717.

Brown:2005:FIS

[BSW05] Peter N. Brown, Dana E. Shumaker, and Carol S. Woodward. Fully implicit solution of large-scale non-equilibrium radiation diffusion with high order time integration. Journal of Computational Physics, 204(2):760–783, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004371.

Bertrand:2002:KBU

[BT02] François Bertrand and Philippe A. Tanguy. Krylov-based Uzawa algorithms for the solution of the Stokes equations using discontinuous-pressure tetrahedral finite elements. Journal of Computational Physics, 181(2):617–638, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971445.

Bao:2003:GSS

[BT03] Weizhu Bao and Weijun Tang. Ground-state solution of Bose– Einstein condensate by directly minimizing the energy functional. Journal of Computational Physics, 187(1):230–254, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000974.

Bolton:2005:MCL

[BT05] P. Bolton and R. W. Thatcher. On mass conservation in least-squares methods. Journal of Computational Physics, 203(1):287–304, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003286.

Bolton:2006:LSF

[BT06] P. Bolton and R. W. Thatcher. A least-squares finite element method for the Navier–Stokes equations. Journal of Computational Physics, 213(1):174–183, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500375X.

Boronski:2007:PTDa

[BT07a] Piotr Boronski and Laurette S. Tuckerman. Poloidal-toroidal decomposition in a finite cylinder. I: Influence matrices for the magnetohydrodynamic equations. Journal of Computational Physics, 227(2):1523–1543, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700383X.

Boronski:2007:PTDb

[BT07b] Piotr Boronski and Laurette S. Tuckerman. Poloidaltoroidal decomposition in a finite cylinder: II. Discretization, regularization and validation. Journal of Computational Physics, 227(2):1544–1566, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003804.

Boubendir:2009:SDB

[BT09] Yassine Boubendir and Svetlana Tlupova. Stokes–Darcy boundary integral solutions using preconditioners. Journal of Computational Physics, 228(23):8627–8641, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004604.

Bennett:2005:LBO

[BTC05] A. F. Bennett, J. R. Taylor, and B. S. Chua. Lattice Boltzmann open boundaries for hydrodynamic models. Journal of Computational Physics, 203(1):89–111, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104003195.

Boyadjiev:2001:MMB

[BTFY01] T. L. Boyadjiev, M. D. Todorov, P. P. Fiziev, and S. S. Yazadjiev. Mathematical modeling of boson-fermion stars in the generalized scalar-tensor theories of gravity. Journal of Computational Physics, 166(2):253–270, January 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910096649X.

Bohm:2009:TCF

[BTSM09] M. Böhm, M. Tasche, B. Seifert, and F. Mitschke. Trimto-coherence Fourier transform. Journal of Computational Physics, 228(8):2906–2917, May 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000072.

Bernetti:2008:ESR

[BTT08] R. Bernetti, V. A. Titarev, and E. F. Toro. Exact solution of the Riemann problem for the shallow water equations with discontinuous bottom geometry. Journal of Computational Physics, 227(6):3212–3243, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005281.

Briley:2003:HRV

[BTW03] W. R. Briley, L. K. Taylor, and D. L. Whitfield. Highresolution viscous flow simulations at arbitrary Mach number. Journal of Computational Physics, 184(1):79–105, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000189.

Balbas:2004:NOC

[BTW04] Jorge Balbás, Eitan Tadmor, and Cheng-Chin Wu. Nonoscillatory central schemes for one- and two-dimensional MHD equations: I. Journal of Computational Physics, 201(1):261–285, November 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002293.

Bui-Thanh:2007:GOM

[BTWGvBW07] T. Bui-Thanh, K. Willcox, O. Ghattas, and B. van Bloemen Waanders. Goal-oriented, model-constrained optimization for reduction of large-scale systems. Journal of Computational Physics, 224(2):880–896, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005535.

Boyle:2002:NGI

[BU02] P. Boyle and UKQCD Collaboration. A novel gauge invariant multistate smearing technique. Journal of Computational Physics, 179(2):349–370, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967981.

Buchmann:2005:SSD

[Buc05] F. M. Buchmann. Simulation of stopped diffusions. Journal of Computational Physics, 202(2):446–462, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002931.

Brunner:2006:CFP

[BUEG06] Thomas A. Brunner, Todd J. Urbatsch, Thomas M. Evans, and Nicholas A. Gentile. Comparison of four parallel algorithms for domain decomposed implicit Monte Carlo. Journal of Computational Physics, 212(2):527–539, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003311.

Burger:2005:NSA

[Bur05] Martin Burger. Numerical simulation of anisotropic surface diffusion with curvature-dependent energy. Journal of Computational Physics, 203(2):602–625, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003936.

Buss:2000:AES

[Bus00] Samuel R. Buss. Accurate and efficient simulation of rigidbody rotations. Journal of Computational Physics, 164(2): 377–406, November 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100966026.

Bao:2000:MEE

[BV00] Gang Bao and Tri Van. Modeling of evanescent energy in optical fibers. Journal of Computational Physics, 161(2):700– 717, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965288.

Becker:2005:MRN

[BV05] Roland Becker and Boris Vexler. Mesh refinement and numerical sensitivity analysis for parameter calibration of partial differential equations. Journal of Computational Physics, 206(1):95–110, June 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004954.

Blokland:2007:PMS

[BvdHKG07] J. W. S. Blokland, B. van der Holst, R. Keppens, and J. P. Goedbloed. PHOENIX: MHD spectral code for rotating laboratory and gravitating astrophysical plasmas. Journal of Computational Physics, 226(1):509–533, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001702.

Beskok:2001:UHF

[BW01] Ali Beskok and Timothy C. Warburton. An unstructured hp finite-element scheme for fluid flow and heat transfer in moving domains. Journal of Computational Physics, 174(2):492–509, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968858.

Beck:2002:DSS

[BW02] J. C. Beck and A. P. Watkins. On the development of spray submodels based on droplet size moments. Journal of Computational Physics, 182(2):586–621, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297186X.

Bao:2006:ESAa

[BW06] Weizhu Bao and Hanquan Wang. An efficient and spectrally accurate numerical method for computing dynamics of rotating Bose–Einstein condensates. Journal of Computational Physics, 217(2):612–626, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000258.

Bridge:2007:MFN

[BW07] L. J. Bridge and B. R. Wetton. A mixture formulation for numerical capturing of a two-phase/vapour interface in a porous medium. Journal of Computational Physics, 225(2):2043–2068, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001192.

Benselama:2009:MMN

[BWLM09] Adel M. Benselama, Mame J.-P. William-Louis, and François Monnoyer. A 1D–3D mixed method for the numerical simulation of blast waves in confined geometries. Journal of Computational Physics, 228(18):6796–6810, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003222.

Bao:2007:EAN

[BY07] Weizhu Bao and Li Yang. Efficient and accurate numerical methods for the Klein–Gordon–Schrödinger equations. Journal of Computational Physics, 225(2):1863–1893, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000915.

Bao:2008:MCB

[BYS08] Jie Bao, Peng Yuan, and Laura Schaefer. A mass conserving boundary condition for the lattice Boltzmann equation method. Journal of Computational Physics, 227 (18):8472–8487, September 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003288.

Biros:2004:FSS

[BYZ04] George Biros, Lexing Ying, and Denis Zorin. A fast solver for the Stokes equations with distributed forces in complex geometries. Journal of Computational Physics, 193(1):317–348, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004352.

Batra:2004:AAS

[BZ04] R. C. Batra and G. M. Zhang. Analysis of adiabatic shear bands in elasto-thermo-viscoplastic materials by modified smoothed-particle hydrodynamics (MSPH) method. Journal of Computational Physics, 201(1):172–190, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002062.

Batra:2008:MSP

[BZ08] R. C. Batra and G. M. Zhang. Modified smoothed particle hydrodynamics (MSPH) basis functions for meshless methods, and their application to axisymmetric Taylor impact test. Journal of Computational Physics, 227(3):1962–1981, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004421.

Boyd:2009:TWS

[BZ09] John P. Boyd and Cheng Zhou. Three ways to solve the Poisson equation on a sphere with Gaussian forcing. Journal of Computational Physics, 228(13):4702–4713, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001338.

Bowman:2000:MAN

[BZB00] John C. Bowman, A. Zeiler, and D. Biskamp. A multigrid algorithm for nonlocal collisional electrostatic drift-wave turbulence. Journal of Computational Physics, 158(2):239–261, March 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199964173.

Brio:2001:TDR

[BZW01] M. Brio, A. R. Zakharian, and G. M. Webb. Two-dimensional Riemann solver for Euler equations of gas dynamics. Journal of Computational Physics, 167(1):177–195, February 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910096666X.

Chang:2006:ALB

[CA06] Qingming Chang and J. Iwan D. Alexander. Application of the lattice Boltzmann method to two-phase Rayleigh–Bénard convection with a deformable interface. Journal of Computational Physics, 212(2):473–489, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003293.

Cuenot:2000:CAS

[CAL00] B. Cuenot, C. Angelberger, and J.-P. Legier. Convergence acceleration for steady flame computations. Journal of Computational Physics, 161(2):718–722, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965124.

Calhoun:2002:CGM

[Cal02] Donna Calhoun. A Cartesian grid method for solving the twodimensional streamfunction-vorticity equations in irregular regions. Journal of Computational Physics, 176(2):231–275, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969700.

Cameron:2003:REG

[Cam03] Chris Cameron. Relative efficiency of Gaussian stochastic process sampling procedures. Journal of Computational Physics, 192(2):546–569, December 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300408X.

Capdeville:2005:NCH

[Cap05] G. Capdeville. A new category of Hermitian upwind schemes for computational acoustics. Journal of Computational Physics, 210(1):133–170, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002020.

Capdeville:2006:NCH

[Cap06] G. Capdeville. A new category of Hermitian upwind schemes for computational acoustics — II. Two-dimensional aeroacoustics. Journal of Computational Physics, 217(2):530– 562, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106000222.

Capdeville:2008:CWS

[Cap08a] G. Capdeville. A central WENO scheme for solving hyperbolic conservation laws on non-uniform meshes. Journal of Computational Physics, 227(5):2977–3014, February 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005256.

Capdeville:2008:HUW

[Cap08b] G. Capdeville. A Hermite upwind WENO scheme for solving hyperbolic conservation laws. Journal of Computational Physics, 227(4):2430–2454, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004731.

Capdeville:2008:TCH

[Cap08c] G. Capdeville. Towards a compact high-order method for non-linear hyperbolic systems, II. The Hermite–HLLC scheme. Journal of Computational Physics, 227(22):9428– 9462, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108003550.

Capdeville:2009:TCH

[Cap09] G. Capdeville. Towards a compact high-order method for non-linear hyperbolic systems. I: The Hermite least-square monotone (HLSM) reconstruction. Journal of Computational Physics, 228(10):3762–3788, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000783.

Caramana:2001:TRS

[Car01] E. J. Caramana. Timestep relaxation with symmetry preservation on high aspect-ratio angular or tapered grids. Journal of Computational Physics, 166(1):173–185, January 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101966483.

Carley:2002:TVM

[Car02] Michael Carley. A triangulated vortex method for the axisymmetric Euler equations. Journal of Computational Physics, 180(2):616–641, August 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297107X.

Caramana:2009:ISL

[Car09] E. J. Caramana. The implementation of slide lines as a combined force and velocity boundary condition. Journal of Computational Physics, 228(11):3911–3916, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001090.

Cummins:2002:IPC

[CB02] S. J. Cummins and J. U. Brackbill. An implicit particlein-cell method for granular materials. Journal of Computational Physics, 180(2):506–548, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971019.

Cervino:2003:ECS

[CB03] Laura I. Cerviño and Thomas R. Bewley. On the extension of the complex-step derivative technique to pseudospectral algorithms. Journal of Computational Physics, 187(2):544–549, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103001232.

Cyr:2007:UMW

[CB07] Eric C. Cyr and Stephen D. Bond. Using the method of weighted residuals to compute potentials of mean force. Journal of Computational Physics, 225(1):714–729, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600619X.

Coyajee:2009:NSD

[CB09] Emil Coyajee and Bendiks Jan Boersma. Numerical simulation of drop impact on a liquid-liquid interface with a multiple marker front-capturing method. Journal of Computational Physics, 228(12):4444–4467, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001363.

Champagne:2001:FFD

[CBB01] Nathan J. Champagne II, James G. Berryman, and H. Michael Buettner. FDFD: a 3D finite-difference frequency-domain code for electromagnetic induction tomography. Journal of Computational Physics, 170(2):830–848, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967658.

Cebrian:2009:SSC

[CBC09] E. Cebrián, L. L. Bonilla, and A. Carpio. Self-sustained current oscillations in the kinetic theory of semiconductor superlattices. Journal of Computational Physics, 228(20):7689–7705, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900388X.

Chantalat:2009:LSP

[CBGI09] Frédéric Chantalat, Charles-Henri Bruneau, Cédric Galusinski, and Angelo Iollo. Level-set, penalization and Cartesian meshes: a paradigm for inverse problems and optimal design. Journal of Computational Physics, 228 (17):6291–6315, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002757.

Choi:2003:QIL

[CBH03] Jae-Hoon Choi, Ki-Ryang Byun, and Ho-Jung Hwang. Quality-improved local refinement of tetrahedral mesh based on element-wise refinement switching. Journal of Computational Physics, 192(1):312–324, November 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003954.

Choi:2004:ICW

[CBI^{+}04] Dae-Il Choi, J. David Brown, Breno Imbiriba, Joan Centrella, and Peter MacNeice. Interface conditions for wave propagation through mesh refinement boundaries. Journal of Computational Physics, 193(2):398–425, January 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300370X.

Carbonell:2007:NSN

[CBJdlC07] F. Carbonell, R. J. Biscay, J. C. Jimenez, and H. de la Cruz. Numerical simulation of nonlinear dynamical systems driven by commutative noise. Journal of Computational Physics, 226(2):1219–1233, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002306.

Chacon:2000:IECa

[CBKM00a] L. Chacón, D. C. Barnes, D. A. Knoll, and G. H. Miley. An implicit energy-conservative 2D Fokker–Planck algorithm: I. Difference scheme. Journal of Computational Physics, 157(2):618–653, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963945.

Chacon:2000:IECb

[CBKM00b] L. Chacón, D. C. Barnes, D. A. Knoll, and G. H. Miley. An implicit energy-conservative 2D Fokker–Planck algorithm: II. Jacobian-free Newton–Krylov solver. Journal of Computational Physics, 157(2):654–682, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963957.

Cristini:2001:AMA

[CBL01] Vittorio Cristini, Jerzy Bławzdziewicz, and Michael Loewenberg. An adaptive mesh algorithm for evolving surfaces: Simulations of drop breakup and coalescence. Journal of Computational Physics, 168(2):445–463, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967130.

Cheng:2002:MCC

[CBMO02] Li-Tien Cheng, Paul Burchard, Barry Merriman, and Stanley Osher. Motion of curves constrained on surfaces using a level-set approach. Journal of Computational Physics, 175(2):604–644, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969608.

Couplet:2005:CRO

[CBS05] M. Couplet, C. Basdevant, and P. Sagaut. Calibrated reducedorder POD–Galerkin system for fluid flow modelling. Journal of Computational Physics, 207(1):192–220, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000239.

Chang:2003:DDA

[CC03] S.-L. Chang and C.-S. Chien. Domain decomposition algorithms for fourth-order nonlinear elliptic eigenvalue problems. Journal of Computational Physics, 191(2):476–501, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003279.

Cook:2004:HWV

[CC04] Andrew W. Cook and William H. Cabot. A high-wavenumber viscosity for high-resolution numerical methods. Journal of Computational Physics, 195(2):594–601, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005746.

Cook:2005:HST

[CC05] Andrew W. Cook and William H. Cabot. Hyperviscosity for shock-turbulence interactions. Journal of Computational Physics, 203(2):379–385, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004000.

Chen:2007:AHO

[CC07] Yanlai Chen and Bernardo Cockburn. An adaptive highorder discontinuous Galerkin method with error control for the Hamilton–Jacobi equations. Part I: The one-dimensional steady state case. Journal of Computational Physics, 226(1):1027–1058, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002203.

Chen:2008:NPP

[CC08a] Dake Chen and Mark A. Cane. El Niño prediction and predictability. Journal of Computational Physics, 227(7): 3625–3640, March 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910700229X.

Coquerelle:2008:VLS

[CC08b] M. Coquerelle and G.-H. Cottet. A vortex level set method for the two-way coupling of an incompressible fluid with colliding rigid bodies. Journal of Computational Physics, 227(21):9121–9137, November 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001976.

Crockett:2005:UCC

[CCF^{+}05] Robert K. Crockett, Phillip Colella, Robert T. Fisher, Richard I. Klein, and Christopher F. McKee. An unsplit, cell-centered Godunov method for ideal MHD. Journal of Computational Physics, 203(2):422–448, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003869.

Cheng:2006:WFM

[CCG^{+}06] Hongwei Cheng, William Y. Crutchfield, Zydrunas Gimbutas, Leslie F. Greengard, J. Frank Ethridge, Jingfang Huang, Vladimir Rokhlin, Norman Yarvin, and Junsheng Zhao. A wideband fast multipole method for the Helmholtz equation in three dimensions. Journal of Computational Physics, 216(1):300–325, July 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500553X.

Calgaro:2008:HFV

[CCG08] Caterina Calgaro, Emmanuel Creusé, and Thierry Goudon. An hybrid finite volume-finite element method for variable density incompressible flows. Journal of Computational Physics, 227(9):4671–4696, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000466.

Chang:2007:CWF

[CCJ07] S.-L. Chang, C.-S. Chien, and B.-W. Jeng. Computing wave functions of nonlinear Schrödinger equations: a time-independent approach. Journal of Computational Physics, 226(1):104–130, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001453.

Chung:2005:EIT

[CCT05] Eric T. Chung, Tony F. Chan, and Xue-Cheng Tai. Electrical impedance tomography using level set representation and total variational regularization. Journal of Computational Physics, 205(1):357–372, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004711.

Chaljub:2003:SEF

[CCV03] Emmanuel Chaljub, Yann Capdeville, and Jean-Pierre Vilotte. Solving elastodynamics in a fluid-solid heterogeneous sphere: a parallel spectral element approximation on non-conforming grids. Journal of Computational Physics, 187(2):457–491, May 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001190.

Cox:2000:BSO

[CD00] Steven J. Cox and David C. Dobson. Band structure optimization of two-dimensional photonic crystals in H-polarization. Journal of Computational Physics, 158(2):214–224, March 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919996415X.

Cai:2003:UEB

[CD03] Wei Cai and Shaozhong Deng. An upwinding embedded boundary method for Maxwell’s equations in media with material interfaces: 2D case. Journal of Computational Physics, 190(1):159–183, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002699.

Campos:2004:NIH

[CD04] Isabel Campos and James W. Davenport. Numerical investigation of the 3D Hubbard model on a Linux cluster. Journal of Computational Physics, 196(1):88–101, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005886.

Costa:2007:MDH

[CD07] Bruno Costa and Wai Sun Don. Multi-domain hybrid spectral-WENO methods for hyperbolic conservation laws. Journal of Computational Physics, 224(2):970–991, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005614.

Chang:2007:MIB

[CDDH07] Anthony Chang, Jonathan A. Dantzig, Brian T. Darr, and Allison Hubel. Modeling the interaction of biological cells with a solidifying interface. Journal of Computational Physics, 226(2):1808–1829, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002628.

Carre:2009:CCL

[CDDL09] G. Carré, S. Del Pino, B. Després, and E. Labourasse. A cellcentered Lagrangian hydrodynamics scheme on general unstructured meshes in arbitrary dimension. Journal of Computational Physics, 228(14):5160–5183, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900196X.

Christie:2006:UQP

[CDE06] Mike Christie, Vasily Demyanov, and Demet Erbas. Uncertainty quantification for porous media flows. Journal of Computational Physics, 217(1):143–158, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000350.

Chabory:2008:FTB

[CdHST08] A. Chabory, B. P. de Hon, W. H. A. Schilders, and A. G. Tijhuis. Fast transform based preconditioners for 2D finitedifference frequency-domain — waveguides and periodic structures. Journal of Computational Physics, 227(16):7755–7767, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108002611.

Chantrasmi:2009:PLA

[CDI09] T. Chantrasmi, A. Doostan, and G. Iaccarino. Padé– Legendre approximants for uncertainty analysis with discontinuous response surfaces. Journal of Computational Physics, 228(19):7159–7180, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900326X.

Cai:2007:EFM

[CDJ07] Wei Cai, Shaozhong Deng, and Donald Jacobs. Extending the fast multipole method to charges inside or outside a dielectric sphere. Journal of Computational Physics, 223(2):846–864, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004773.

Cohen:2000:MST

[CDKP00] Albert Cohen, Nira Dyn, Sidi Mahmoud Kaber, and Marie Postel. Multiresolution schemes on triangles for scalar conservation laws. Journal of Computational Physics, 161(1): 264–286, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100965033.

Crouseilles:2004:HKF

[CDL04] N. Crouseilles, P. Degond, and M. Lemou. A hybrid kinetic/fluid model for solving the gas dynamics Boltzmann–BGK equation. Journal of Computational Physics, 199(2):776–808, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001329.

Crouseilles:2005:HKF

[CDL05] Nicolas Crouseilles, Pierre Degond, and Mohammed Lemou. A hybrid kinetic-fluid model for solving the Vlasov–BGK equation. Journal of Computational Physics, 203(2):572–601, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104003924.

Chen:2009:ISP

[CDR09] Yu Chen, Ran Duan, and Vladimir Rokhlin. On the inverse scattering problem in the acoustic environment. Journal of Computational Physics, 228(9):3209–3231, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006724.

Chinnayya:2004:MDW

[CDS04] Ashwin Chinnayya, Eric Daniel, and Richard Saurel. Modelling detonation waves in heterogeneous energetic materials. Journal of Computational Physics, 196(2):490–538, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006065.

Crispel:2005:QNF

[CDV05] Pierre Crispel, Pierre Degond, and Marie-Hélène Vignal. Quasi-neutral fluid models for current-carrying plasmas. Journal of Computational Physics, 205(2):408–438, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004735.

Crispel:2007:APS

[CDV07] Pierre Crispel, Pierre Degond, and Marie-Hélène Vignal. An asymptotic preserving scheme for the two-fluid Euler– Poisson model in the quasineutral limit. Journal of Computational Physics, 223(1):208–234, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004335.

Cecil:2005:NMC

[Cec05] Thomas Cecil. A numerical method for computing minimal surfaces in arbitrary dimension. Journal of Computational Physics, 206(2):650–660, July 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000094.

Chenier:2009:CFV

[CEH09] Eric Chénier, Robert Eymard, and Raphaèle Herbin. A collocated finite volume scheme to solve free convection for general non-conforming grids. Journal of Computational Physics, 228(6):2296–2311, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005639.

Colin:2006:CND

[CEL06] F. Colin, R. Egli, and F. Y. Lin. Computing a null divergence velocity field using smoothed particle hydrodynamics. Journal of Computational Physics, 217(2):680–692, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000295.

Chen:2007:DGI

[CELS07] Shanqin Chen, Weinan E., Yunxian Liu, and Chi-Wang Shu. A discontinuous Galerkin implementation of a domain decomposition method for kinetic-hydrodynamic coupling multiscale problems in gas dynamics and device simulations. Journal of Computational Physics, 225(2):1314–1330, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000435.

Cassidy:2009:IIS

[CET09] Daniel A. Cassidy, Jack R. Edwards, and Ming Tian. An investigation of interface-sharpening schemes for multiphase mixture flows. Journal of Computational Physics, 228(16):5628–5649, September 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000850.

Crouseilles:2004:NAC

[CF04] Nicolas Crouseilles and Francis Filbet. Numerical approximation of collisional plasmas by high order methods. Journal of Computational Physics, 201(2):546–572, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002426.

Callaghan:2006:CLA

[CF06a] T. G. Callaghan and L. K. Forbes. Computing large-amplitude progressive Rossby waves on a sphere. Journal of Computational Physics, 217(2):845–865, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000441.

Chou:2006:MDM

[CF06b] Chuan-Chih Chou and Michael L. Falk. Multiscale diffusion Monte Carlo simulation of epitaxial growth. Journal of Computational Physics, 217(2):519–529, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000210.

Caiden:2001:NMT

[CFA01] Rachel Caiden, Ronald P. Fedkiw, and Chris Anderson. A numerical method for two-phase flow consisting of separate compressible and incompressible regions. Journal of Computational Physics, 166(1):1–27, January 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966245.

Carlini:2007:SLS

[CFF07] E. Carlini, M. Falcone, and R. Ferretti. A semi-Lagrangian scheme for the curve shortening flow in codimension-2. Journal of Computational Physics, 225(2):1388–1408, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000472.

Clamond:2005:EMT

[CFGK05] Didier Clamond, Dorian Fructus, John Grue, and Øyvind Kristiansen. An efficient model for three-dimensional surface wave simulations. Part II: Generation and absorption. Journal of Computational Physics, 205(2):686–705, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000045.

Collino:2006:CST

[CFJ06] F. Collino, T. Fouquet, and P. Joly. Conservative space–time mesh refinement methods for the FDTD solution of Maxwell’s equations. Journal of Computational Physics, 211(1):9–35, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001804.

Cueto-Felgueroso:2009:ARS

[CFJ09] Luis Cueto-Felgueroso and Ruben Juanes. Adaptive rational spectral methods for the linear stability analysis of nonlinear fourth-order problems. Journal of Computational Physics, 228(17):6536–6552, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003179.

Chaniotis:2003:RSP

[CFL^{+}03] A. K. Chaniotis, C. E. Frouzakis, J. C. Lee, A. G. Tomboulides, D. Poulikakos, and K. Boulouchos. Remeshed smoothed particle hydrodynamics for the simulation of laminar chemically reactive flows. Journal of Computational Physics, 191(1):1–17, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003024.

Ceniceros:2009:CFP

[CFM09] Hector D. Ceniceros, Glenn H. Fredrickson, and George O. Mohler. Coupled flow-polymer dynamics via statistical field theory: Modeling and computation. Journal of Computational Physics, 228(5):1624–1638, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800586X.

Cohen:2006:SHO

[CFP06] G. Cohen, X. Ferrieres, and S. Pernet. A spatial highorder hexahedral discontinuous Galerkin method to solve Maxwell’s equations in time domain. Journal of Computational Physics, 217(2):340–363, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000131.

Cueto-Felgueroso:2008:TAF

[CFP08] Luis Cueto-Felgueroso and Jaume Peraire. A time-adaptive finite volume method for the Cahn–Hilliard and Kuramoto– Sivashinsky equations. Journal of Computational Physics, 227(24):9985–10017, December 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003847.

Calvo:2004:NMS

[CFR04] M. Calvo, J. M. Franco, and L. Rández. A new minimum storage Runge–Kutta scheme for computational acoustics. Journal of Computational Physics, 201(1):1–12, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104001913.

Coulaud:2008:HPB

[CFR08] O. Coulaud, P. Fortin, and J. Roman. High performance BLAS formulation of the multipole-to-local operator in the fast multipole method. Journal of Computational Physics, 227(3):1836–1862, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004354.

Ceniceros:2009:ESR

[CFR09] Hector D. Ceniceros, Jordan E. Fisher, and Alexandre M. Roma. Efficient solutions to robust, semi-implicit discretizations of the immersed boundary method. Journal of Computational Physics, 228(19):7137–7158, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002629.

Cameron:2009:AAR

[CFS09] M. K. Cameron, S. B. Fomel, and J. A. Sethian. Analysis and algorithms for a regularized Cauchy problem arising from a non-linear elliptic PDE for seismic velocity estimation. Journal of Computational Physics, 228(19):7388–7411, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003660.

Cooper:2009:BBE

[CGC^{+}09] G. A. Cooper, J. P. Graves, W. A. Cooper, R. Gruber, and R. S. Peterson. BECOOL: Ballooning eigensolver with COOL finite elements. Journal of Computational Physics, 228(13):4911–4916, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001892.

Chatzidai:2009:EMG

[CGDT09] N. Chatzidai, A. Giannousakis, Y. Dimakopoulos, and J. Tsamopoulos. On the elliptic mesh generation in domains containing multiple inclusions and undergoing large deformations. Journal of Computational Physics, 228(6):1980–2011, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108006128.

Chen:2005:UAS

[CGH05] Qian-Yong Chen, David Gottlieb, and Jan S. Hesthaven. Uncertainty analysis for the steady-state flows in a dual throat nozzle. Journal of Computational Physics, 204(1):378–398, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004218.

Colella:2006:CGE

[CGKM06] Phillip Colella, Daniel T. Graves, Benjamin J. Keen, and David Modiano. A Cartesian grid embedded boundary method for hyperbolic conservation laws. Journal of Computational Physics, 211(1):347–366, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002780.

Cances:2006:CES

[CGL06] Éric Cancès, Hervé Galicher, and Mathieu Lewin. Computing electronic structures: a new multiconfiguration approach for excited states. Journal of Computational Physics, 212(1):73–98, February 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003104.

Carrillo:2008:SFP

[CGL08] José-Antonio Carrillo, Thierry Goudon, and Pauline Lafitte. Simulation of fluid and particles flows: Asymptotic preserving schemes for bubbling and flowing regimes. Journal of Computational Physics, 227(16):7929–7951, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002696.

Crouch:2007:POF

[CGM07] J. D. Crouch, A. Garbaruk, and D. Magidov. Predicting the onset of flow unsteadiness based on global instability. Journal of Computational Physics, 224(2):924–940, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005572.

Carrillo:2003:WST

[CGMS03] José A. Carrillo, Irene M. Gamba, Armando Majorana, and Chi-Wang Shu. A WENO-solver for the transients of Boltzmann–Poisson system for semiconductor devices: performance and comparisons with Monte Carlo methods. Journal of Computational Physics, 184(2):498–525, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000323.

Carrillo:2006:SDS

[CGMS06] José A. Carrillo, Irene M. Gamba, Armando Majorana, and Chi-Wang Shu. 2D semiconductor device simulations by WENO-boltzmann schemes: Efficiency, boundary conditions and comparison to Monte Carlo methods. Journal of Computational Physics, 214(1):55–80, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004237.

Cavalli:2007:SEB

[CGN^{+}07] F. Cavalli, A. Gamba, G. Naldi, M. Semplice, D. Valdembri, and G. Serini. 3D simulations of early blood vessel formation. Journal of Computational Physics, 225(2):2283–2300, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107001313.

Chang:2002:AEF

[CGP02] Wang Chang, Francis Giraldo, and Blair Perot. Analysis of an exact fractional step method. Journal of Computational Physics, 180(1):183–199, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970877.

Cao:2005:MSS

[CGP05] Yang Cao, Dan Gillespie, and Linda Petzold. Multiscale stochastic simulation algorithm with stochastic partial equilibrium assumption for chemically reacting systems. Journal of Computational Physics, 206(2):395–411, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005182.

Castro:2004:NST

[CGRGV^{+}04] Manuel J. Castro, José A. García-Rodríguez, José M. González-Vida, Jorge Macías, Carlos Parés, and M. Elena Vázquez-Cendón. Numerical simulation of two-layer shallow water flows through channels with irregular geometry. Journal of Computational Physics, 195(1):202–235, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005278.

Cabezon:2008:OPF

[CGSR08] Rubén M. Cabezón, Domingo García-Senz, and Antonio Relaño. A one-parameter family of interpolating kernels for smoothed particle hydrodynamics studies. Journal of Computational Physics, 227(19):8523–8540, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003148.

Cliffe:2000:PCF

[CGSS00] K. A. Cliffe, I. G. Graham, R. Scheichl, and L. Stals. Parallel computation of flow in heterogeneous media modelled by mixed finite elements. Journal of Computational Physics, 164(2):258–282, November 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965938.

Ceniceros:2001:EDA

[CH01] Hector D. Ceniceros and Thomas Y. Hou. An efficient dynamically adaptive mesh for potentially singular solutions. Journal of Computational Physics, 172(2):609–639, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968445.

Czarny:2008:BSF

[CH08] Olivier Czarny and Guido Huysmans. Bézier surfaces and finite elements for MHD simulations. Journal of Computational Physics, 227(16):7423–7445, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002118.

Chang:2007:DPF

[Cha07a] Britton Chang. A deterministic photon free method to solve radiation transfer equations. Journal of Computational Physics, 222(1):71–85, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600307X.

Chang:2007:ISI

[Cha07b] Britton Chang. The incorporation of the semi-implicit linear equations into Newton’s method to solve radiation transfer equations. Journal of Computational Physics, 226(1):852–878, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002124.

Chato:2009:PFM

[Cha09] David J. Chato. Phase field modeling of liquid jets in low gravity. Journal of Computational Physics, 228(5):1521–1540, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005718.

Chao:2009:MPM

[CHB09] J. Chao, A. Haselbacher, and S. Balachandar. A massively parallel multi-block hybrid compact-WENO scheme for compressible flows. Journal of Computational Physics, 228(19):7473–7491, October 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003805.

Christlieb:2004:KDF

[CHBS04] Andrew J. Christlieb, W. Nicholas G. Hitchon, Iain D. Boyd, and Quanhua Sun. Kinetic description of flow past a microplate. Journal of Computational Physics, 195(2):508–527, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005667.

Codina:2009:FMA

[CHCOB09] Ramon Codina, Guillaume Houzeaux, Herbert Coppola-Owen, and Joan Baiges. The fixed-mesh ALE approach for the numerical approximation of flows in moving domains. Journal of Computational Physics, 228(5):1591–1611, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005846.

Cheong:2000:ADF

[Che00a] Hyeong-Bin Cheong. Application of double Fourier series to the shallow-water equations on a sphere. Journal of Computational Physics, 165(1):261–287, November 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966154.

Cheong:2000:DFS

[Che00b] Hyeong-Bin Cheong. Double Fourier series on a sphere: Applications to elliptic and vorticity equations. Journal of Computational Physics, 157(1):327–349, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963854.

Chen:2003:FSC

[Che03] XinJian Chen. A free-surface correction method for simulating shallow water flows. Journal of Computational Physics, 189(2):557–578, August 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002341.

Chen:2004:CMF

[Che04] XinJian Chen. A Cartesian method for fitting the bathymetry and tracking the dynamic position of the shoreline in a three-dimensional, hydrodynamic model. Journal of Computational Physics, 200(2):749–768, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001871.

Cheng:2007:ELS

[Che07] Li-Tien Cheng. Efficient level set methods for constructing wavefronts in three spatial dimensions. Journal of Computational Physics, 226(2):2250–2270, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003117.

Chevaugeon:2007:ONP

[CHG^{+}07] Nicolas Chevaugeon, Koen Hillewaert, Xavier Gallez, Paul Ploumhans, and Jean-François Remacle. Optimal numerical parameterization of discontinuous Galerkin method applied to wave propagation problems. Journal of Computational Physics, 223(1):188–207, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004323.

Calabrese:2006:NSF

[CHH06] G. Calabrese, I. Hinder, and S. Husa. Numerical stability for finite difference approximations of Einstein’s equations. Journal of Computational Physics, 218(2):607–634, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106001239.

Camassa:2006:IIA

[CHL06a] Roberto Camassa, Jingfang Huang, and Long Lee. Integral and integrable algorithms for a nonlinear shallow-water wave equation. Journal of Computational Physics, 216(2):547–572, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005644.

Cheng:2006:AFS

[CHL06b] Hongwei Cheng, Jingfang Huang, and Terry Jo Leiterman. An adaptive fast solver for the modified Helmholtz equation in two dimensions. Journal of Computational Physics, 211(2):616–637, January 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002901.

Cao:2009:SES

[CHL09] Yong Cao, Xiaoming He, and Tao Lü. A splitting extrapolation for solving nonlinear elliptic equations with d-quadratic finite elements. Journal of Computational Physics, 228(1):109–122, January 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004567.

Castronovo:2008:MTC

[CHM08] E. Castronovo, J. Harlim, and A. J. Majda. Mathematical test criteria for filtering complex systems: Plentiful observations. Journal of Computational Physics, 227(7):3678–3714, March 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005372.

Chopp:2000:LSM

[Cho00] D. L. Chopp. A level-set method for simulating island coarsening. Journal of Computational Physics, 162(1):104–122, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965276.

Choblet:2005:MTC

[Cho05] Gaël Choblet. Modelling thermal convection with large viscosity gradients in one block of the ‘cubed sphere’. Journal of Computational Physics, 205(1):269–291, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400467X.

Cotter:2009:LSM

[CHPR09] Colin J. Cotter, David A. Ham, Christopher C. Pain, and Sebastian Reich. LBB stability of a mixed Galerkin finite element pair for fluid flow simulations. Journal of Computational Physics, 228(2):336–348, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004816.

Cao:2001:EIM

[CHR01] Weiming Cao, Weizhang Huang, and Robert D. Russell. An error indicator monitor function for an radaptive finite-element method. Journal of Computational Physics, 170(2):871–892, July 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967701.

Chrzanowska:2003:AGQ

[Chr03] Agnieszka Chrzanowska. Application of Gaussian quadratures to density functional (df) theories of confined liquid crystals. Journal of Computational Physics, 191(1):265–281, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003164.

Christensen:2004:HSA

[Chr04] Michael Christensen. How to simulate anisotropic diffusion processes on curved surfaces. Journal of Computational Physics, 201(2):421–438, December 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002360.

Chu:2009:DMM

[Chu09] Kevin T. Chu. A direct matrix method for computing analytical Jacobians of discretized nonlinear integro-differential equations. Journal of Computational Physics, 228(15):5526–5538, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002289.

Cho:2004:AMG

[CJ04] Maenghyo Cho and Seongki Jun. r-adaptive mesh generation for shell finite element analysis. Journal of Computational Physics, 199(1):291–316, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000853.

Ciarlet:2007:CGM

[CJ07] Patrick Ciarlet, Jr. and Erell Jamelot. Continuous Galerkin methods for solving the time-dependent Maxwell equations in 3D geometries. Journal of Computational Physics, 226(1):1122–1135, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002240.

Chen:2009:MKF

[CJ09] Yibing Chen and Song Jiang. Modified kinetic flux vector splitting schemes for compressible flows. Journal of Computational Physics, 228(10):3582–3604, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900059X.

Choi:2003:FAS

[CJK^{+}03] Yung-Sze Choi, Juha Javanainen, Israel Koltracht, Marijan Kostrun, Patrick J. McKenna, and Nataliya Savytska. A fast algorithm for the solution of the timeindependent Gross–Pitaevskii equation. Journal of Computational Physics, 190(1):1–21, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002353.

Cai:2009:BMD

[CJLS09] Jian-Feng Cai, Hui Ji, Chaoqiang Liu, and Zuowei Shen. Blind motion deblurring using multiple images. Journal of Computational Physics, 228(14):5057–5071, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001867.

Clarisse:2004:GTM

[CJR04] Jean-Marie Clarisse, Stéphane Jaouen, and Pierre-Arnaud Raviart. A Godunov-type method in Lagrangian coordinates for computing linearly-perturbed planar-symmetric flows of gas dynamics. Journal of Computational Physics, 198(1): 80–105, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104000208.

Cai:2008:SGE

[CJSS08] Wei Cai, Xia Ji, Jiguang Sun, and Sihong Shao. A Schwarz generalized eigen-oscillation spectral element method (GeSEM) for 2-D high frequency electromagnetic scattering in dispersive inhomogeneous media. Journal of Computational Physics, 227(23):9933–9954, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004324.

Chacon:2003:HHM

[CK03] L. Chacón and D. A. Knoll. A 2D high-β Hall MHD implicit nonlinear solver. Journal of Computational Physics, 188(2):573–592, July 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103001931.

Cornet:2007:NAC

[CK07] Christophe Cornet and Dixon T. K. Kwok. A new algorithm for charge deposition for multiple-grid method for PIC simulations in r–z cylindrical coordinates. Journal of Computational Physics, 225(1):808–828, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000058.

Chiavazzo:2008:QEG

[CK08] Eliodoro Chiavazzo and Iliya V. Karlin. Quasi-equilibrium grid algorithm: Geometric construction for model reduction. Journal of Computational Physics, 227(11):5535–5560, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000971.

Chacon:2002:INR

[CKF02] L. Chacón, D. A. Knoll, and J. M. Finn. An implicit, nonlinear reduced resistive MHD solver. Journal of Computational Physics, 178(1):15–36, May 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970154.

Cheong:2004:FSH

[CKG04] Hyeong-Bin Cheong, In-Hyuk Kwon, and Tae-Young Goo. Further study on the high-order double-Fourier-series spectral filtering on a sphere. Journal of Computational Physics, 193(1):180–197, January 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004261.

Cheong:2002:HOH

[CKGL02] Hyeong-Bin Cheong, In-Hyuk Kwon, Tae-Young Goo, and Myeong-Joo Lee. High-order harmonic spectral filter with the double Fourier series on a sphere. Journal of Computational Physics, 177(2):313–335, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102969974.

Chorin:2000:OPH

[CKL00] Alexandre J. Chorin, Raz Kupferman, and Doron Levy. Optimal prediction for Hamiltonian partial differential equations. Journal of Computational Physics, 162(1):267–297, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965367.

Chang:2005:CBL

[CKLS05] Shu-Ming Chang, Yuen-Cheng Kuo, Wen-Wei Lin, and Shih-Feng Shieh. A continuation BSOR–Lanczos–Galerkin method for positive bound states of a multi-component Bose–Einstein condensate. Journal of Computational Physics, 210(2):439– 458, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002287.

Cui:2007:NMS

[CkM07] Yanfen Cui and De kang Mao. Numerical method satisfying the first two conservation laws for the Korteweg– de Vries equation. Journal of Computational Physics, 227(1):376–399, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003373.

Chang:2007:RDC

[CKPW07] Paul Chang, Gwang-Soo Kim, Keith Promislow, and Brian Wetton. Reduced dimensional computational models of polymer electrolyte membrane fuel cell stacks. Journal of Computational Physics, 223(2):797–821, May 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600475X.

Chae:2000:DIG

[CKR00] Dongsuk Chae, Chongam Kim, and Oh-Hyun Rho. Development of an improved gas-kinetic BGK scheme for inviscid and viscous flows. Journal of Computational Physics, 158(1):1–27, February 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999199964008. See comment [Xu01b] and reply [CKR01].

Chae:2001:RCD

[CKR01] D. Chae, C. Kim, and O. H. Rho. Reply to comment on “Development of an Improved Gas-Kinetic BGK Scheme for Inviscid and Viscous Flows”. Journal of Computational Physics, 171(2):848–850, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968494. See [CKR00, Xu01b].

Cottet:2000:VMS

[CKS00] Georges-Henri Cottet, Petros Koumoutsakos, and Mohamed Lemine Ould Salihi. Vortex methods with spatially varying cores. Journal of Computational Physics, 162(1):164–185, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965318.

Carpentieri:2007:ABA

[CKvT07] G. Carpentieri, B. Koren, and M. J. L. van Tooren. Adjointbased aerodynamic shape optimization on unstructured meshes. Journal of Computational Physics, 224(1):267–287, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000721.

Calhoun:2000:CGF

[CL00a] Donna Calhoun and Randall J. LeVeque. A Cartesian grid finite-volume method for the advection-diffusion equation in irregular geometries. Journal of Computational Physics, 157(1):143–180, January 1, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999199963696.

Cooke:2000:ESD

[CL00b] S. J. Cooke and B. Levush. Eigenmode solution of 2-D and 3-D electromagnetic cavities containing absorbing materials using the Jacobi–Davidson algorithm. Journal of Computational Physics, 157(1):350–370, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963866.

Chalons:2001:HOE

[CL01a] C. Chalons and P. G. LeFloch. High-order entropyconservative schemes and kinetic relations for van der Waals fluids. Journal of Computational Physics, 168(1):184–206, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966907.

Cheong:2001:GOD

[CL01b] Cheolung Cheong and Soogab Lee. Grid-optimized dispersionrelation-preserving schemes on general geometries for computational aeroacoustics. Journal of Computational Physics, 174(1):248–276, November 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969049.

Chertock:2001:PMD

[CL01c] Alina Chertock and Doron Levy. Particle methods for dispersive equations. Journal of Computational Physics, 171 (2):708–730, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968032.

Constantinescu:2002:HAT

[CL02] G. S. Constantinescu and S. K. Lele. A highly accurate technique for the treatment of flow equations at the polar axis in cylindrical coordinates using series expansions. Journal of Computational Physics, 183(1):165–186, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971871.

Chen:2003:IMA

[CL03a] Ren-Chuen Chen and Jinn-Liang Liu. An iterative method for adaptive finite element solutions of an energy transport model of semiconductor devices. Journal of Computational Physics, 189(2):579–606, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300247X.

Colagrossi:2003:NSI

[CL03b] Andrea Colagrossi and Maurizio Landrini. Numerical simulation of interfacial flows by smoothed particle hydrodynamics. Journal of Computational Physics, 191(2):448–475, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003243.

Chen:2005:QCE

[CL05] Ren-Chuen Chen and Jinn-Liang Liu. A quantum corrected energy-transport model for nanoscale semiconductor devices. Journal of Computational Physics, 204(1):131–156, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004115.

Caramana:2006:CVD

[CL06a] E. J. Caramana and R. Loubère. “Curl-q”: a vorticity damping artificial viscosity for essentially irrotational Lagrangian hydrodynamics calculations. Journal of Computational Physics, 215(2):385–391, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005140.

Chacon:2006:FIN

[CL06b] L. Chacón and G. Lapenta. A fully implicit, nonlinear adaptive grid strategy. Journal of Computational Physics, 212(2):703–717, March 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003529.

Chang:2007:RAA

[CL07a] Chih-Hao Chang and Meng-Sing Liou. A robust and accurate approach to computing compressible multiphase flow: Stratified flow model and AUSM^{+}-up scheme. Journal of Computational Physics, 225(1):840–873, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000071. See erratum [CL08b].

Chilton:2007:DOF

[CL07b] Ryan A. Chilton and Robert Lee. The discrete origin of FETD–Newmark late time instability, and a correction scheme. Journal of Computational Physics, 224(2):1293–1306, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005870.

Camassa:2008:CIP

[CL08a] Roberto Camassa and Long Lee. Complete integrable particle methods and the recurrence of initial states for a nonlinear shallow-water wave equation. Journal of Computational Physics, 227(15):7206–7221, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800209X.

Chang:2008:ERA

[CL08b] Chih-Hao Chang and Meng-Sing Liou. Erratum to “A robust and accurate approach to computing compressible multiphase flow: Stratified flow model and AUSM^{+}-up scheme” [J. Comput. Phys. 225 (2007) 840–873]. Journal of Computational Physics, 227(10):5360, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800079X. See [CL07a].

Chatelain:2008:ICI

[CL08c] Philippe Chatelain and Anthony Leonard. Isotropic compact interpolation schemes for particle methods. Journal of Computational Physics, 227(6):3244–3259, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005293.

Chen:2008:AMI

[CL08d] Ren-Chuen Chen and Jinn-Liang Liu. An accelerated monotone iterative method for the quantum-corrected energy transport model. Journal of Computational Physics, 227(12): 6226–6240, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108001423.

Cadiou:2008:ANA

[CLB08] Anne Cadiou, Lionel Le Penven, and Marc Buffat. Asymptotic and numerical analysis of an inviscid bounded vortex flow at low Mach number. Journal of Computational Physics, 227(18):8268–8289, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800291X.

Clerc:2000:NSH

[Cle00] S. Clerc. Numerical simulation of the homogeneous equilibrium model for two-phase flows. Journal of Computational Physics, 161(1):354–375, June 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096515X.

Calhoun-Lopez:2007:EIF

[CLG07] Marcus Calhoun-Lopez and Max D. Gunzburger. The efficient implementation of a finite element, multi-resolution viscosity method for hyperbolic conservation laws. Journal of Computational Physics, 225(2):1288–1313, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000423.

Camarda:2007:KMC

[CLL07a] Massimo Camarda, Antonino La Magna, and Francesco La Via. A kinetic Monte Carlo method on super-lattices for the study of the defect formation in the growth of close packed structures. Journal of Computational Physics, 227(2):1075–1093, December 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003786.

Cheng:2007:QTS

[CLL^{+}07b] Candong Cheng, Joon-Ho Lee, Kim Hwa Lim, Hisham Z. Massoud, and Qing Huo Liu. 3D quantum transport solver based on the perfectly matched layer and spectral element methods for the simulation of semiconductor nanodevices. Journal of Computational Physics, 227(1):455–471, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003427.

Carin:2009:CSM

[CLLG09] Lawrence Carin, Dehong Liu, Wenbin Lin, and Bin Guo. Compressive sensing for multi-static scattering analysis. Journal of Computational Physics, 228(9):3464–3477, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000564.

Castro:2008:WMT

[CLMRP08] Manuel J. Castro, Philippe G. LeFloch, María Luz Muñoz-Ruiz, and Carlos Parés. Why many theories of shock waves are necessary: Convergence error in formally pathconsistent schemes. Journal of Computational Physics, 227(17):8107–8129, September 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002842.

Cockburn:2004:LDF

[CLS04] Bernardo Cockburn, Fengyan Li, and Chi-Wang Shu. Locally divergence-free discontinuous Galerkin methods for the Maxwell equations. Journal of Computational Physics, 194(2):588–610, March 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004960.

Chang:2005:GST

[CLS05] Shu-Ming Chang, Wen-Wei Lin, and Shih-Feng Shieh. Gauss–Seidel-type methods for energy states of a multicomponent Bose–Einstein condensate. Journal of Computational Physics, 202(1):367–390, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002906.

Caflisch:2006:AMM

[CLS^{+}06] R. E. Caflisch, Y.-J. Lee, S. Shu, Y.-X. Xiao, and J. Xu. An application of multigrid methods for a discrete elastic model for epitaxial systems. Journal of Computational Physics, 219(2):697–714, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001999.

Chiu:2009:DRP

[CLS09a] P. H. Chiu, Long Lee, and Tony W. H. Sheu. A dispersionrelation-preserving algorithm for a nonlinear shallow-water wave equation. Journal of Computational Physics, 228 (21):8034–8052, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004100.

Crouseilles:2009:PVS

[CLS09b] Nicolas Crouseilles, Guillaume Latu, and Eric Sonnendrücker. A parallel Vlasov solver based on local cubic spline interpolation on patches. Journal of Computational Physics, 228 (5):1429–1446, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005652.

Chen:2007:FMF

[CLTA07] Chang-Ming Chen, F. Liu, I. Turner, and V. Anh. A Fourier method for the fractional diffusion equation describing sub-diffusion. Journal of Computational Physics, 227 (2):886–897, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107002161.

Cortez:2000:BPM

[CM00] R. Cortez and M. Minion. The blob projection method for immersed boundary problems. Journal of Computational Physics, 161(2):428–453, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965021.

Cox:2002:ETD

[CM02] S. M. Cox and P. C. Matthews. Exponential time differencing for stiff systems. Journal of Computational Physics, 176(2):430–455, March 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102969950.

Chow:2003:FSN

[CM03] Fotini Katopodes Chow and Parviz Moin. A further study of numerical errors in large-eddy simulations. Journal of Computational Physics, 184(2):366–380, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000207.

Cecil:2006:VAP

[CM06] Thomas Cecil and Daniel E. Marthaler. A variational approach to path planning in three dimensions using level set methods. Journal of Computational Physics, 211(1):179–197, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002676.

Chen:2009:NSS

[CMG09] Han Chen, Chohong Min, and Frédéric Gibou. A numerical scheme for the Stefan problem on adaptive Cartesian grids with supralinear convergence rate. Journal of Computational Physics, 228(16):5803–5818, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002381.

Chen:2001:LSM

[CMK^{+}01] Susan Chen, Barry Merriman, Myungjoo Kang, Russel E. Caflisch, Christian Ratsch, Li-Tien Cheng, Mark Gyure, Ronald P. Fedkiw, Christopher Anderson, and Stanley Osher. A level set method for thin film epitaxial growth. Journal of Computational Physics, 167(2):475–500, March 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966890.

Cottet:2002:CSV

[CMOV02] Georges-Henri Cottet, Bertrand Michaux, Sepand Ossia, and Geoffroy VanderLinden. A comparison of spectral and vortex methods in three-dimensional incompressible flows. Journal of Computational Physics, 175(2):702–712, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969633.

Chester:2007:MTF

[CMP07] Stuart Chester, Charles Meneveau, and Marc B. Parlange. Modeling turbulent flow over fractal trees with renormalized numerical simulation. Journal of Computational Physics, 225(1):427–448, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005997.

Compere:2008:TAA

[CMR08] Gaëtan Compère, Emilie Marchandise, and Jean-François Remacle. Transient adaptivity applied to two-phase incompressible flows. Journal of Computational Physics, 227(3):1923–1942, January 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004408.

Cervone:2009:GPC

[CMSZ09] A. Cervone, S. Manservisi, R. Scardovelli, and S. Zaleski. A geometrical predictor-corrector advection scheme and its application to the volume fraction function. Journal of Computational Physics, 228(2):406–419, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800497X.

Chen:2005:TIF

[CN05] Bang-Fuh Chen and Roger Nokes. Time-independent finite difference analysis of fully non-linear and viscous fluid sloshing in a rectangular tank. Journal of Computational Physics, 209(1):47–81, October 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001348.

Chanrion:2008:PMC

[CN08] O. Chanrion and T. Neubert. A PIC–MCC code for simulation of streamer propagation in air. Journal of Computational Physics, 227(15):7222–7245, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002210.

Codina:2001:PSF

[Cod01] Ramon Codina. Pressure stability in fractional step finite element methods for incompressible flows. Journal of Computational Physics, 170(1):112–140, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967257.

Coelho:2002:BSH

[Coe02] P. J. Coelho. Bounded skew high-order resolution schemes for the discrete ordinates method. Journal of Computational Physics, 175(2):412–437, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968998.

Choi:2007:IBM

[COER07] Jung-Il Choi, Roshan C. Oberoi, Jack R. Edwards, and Jacky A. Rosati. An immersed boundary method for complex incompressible flows. Journal of Computational Physics, 224(2):757–784, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005481.

Cecil:2006:SFA

[COQ06] Thomas C. Cecil, Stanley J. Osher, and Jianliang Qian. Simplex free adaptive tree fast sweeping and evolution methods for solving level set equations in arbitrary dimension. Journal of Computational Physics, 213(2):458–473, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003888.

Cortez:2000:VIM

[Cor00] Ricardo Cortez. A vortex/impulse method for immersed boundary motion in high Reynolds number flows. Journal of Computational Physics, 160(1):385–400, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096474X.

Cohen:2008:MSI

[COR08] David Cohen, Brynjulf Owren, and Xavier Raynaud. Multisymplectic integration of the Camassa–Holm equation. Journal of Computational Physics, 227(11):5492–5512, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000855.

Caliari:2009:MAC

[CORT09] Marco Caliari, Alexander Ostermann, Stefan Rainer, and Mechthild Thalhammer. A minimisation approach for computing the ground state of Gross–Pitaevskii systems. Journal of Computational Physics, 228(2):349–360, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004828.

Cecil:2004:NMM

[COV04] Thomas Cecil, Stanley Osher, and Luminita Vese. Numerical methods for minimization problems constrained to S^{1} and S^{2}. Journal of Computational Physics, 198(2):567–579, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910400052X.

Campos:2000:HAF

[CP00] Rafael G. Campos and L. O. Pimentel. Hydrogen atom in a finite linear space. Journal of Computational Physics, 160(1):179–194, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100964465.

Cao:2003:NIS

[CP03a] Renfeng Cao and Stephen B. Pope. Numerical integration of stochastic differential equations: weak second-order midpoint scheme for application in the composition PDF method. Journal of Computational Physics, 185(1):194–212, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000542.

Carcione:2003:EDS

[CP03b] José M. Carcione and Flavio Poletto. Electric drill-string telemetry. Journal of Computational Physics, 186(2):596–609, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000858.

Chen:2003:PMG

[CP03c] Yang Chen and Scott E. Parker. A δf particle method for gyrokinetic simulations with kinetic electrons and electromagnetic perturbations. Journal of Computational Physics, 189(2):463–475, August 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002286.

Chaniotis:2004:HOI

[CP04a] A. K. Chaniotis and D. Poulikakos. High order interpolation and differentiation using B-splines. Journal of Computational Physics, 197(1):253–274, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006296.

Chng:2004:CSE

[CP04b] H. N. Ch’ng and J. Pan. Cubic spline elements for modelling microstructural evolution of materials controlled by solid-state diffusion and grain-boundary migration. Journal of Computational Physics, 196(2):724–750, May 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006144.

Cottet:2004:ADN

[CP04c] G.-H. Cottet and P. Poncet. Advances in direct numerical simulations of 3D wall-bounded flows by vortex-in-cell methods. Journal of Computational Physics, 193(1):136–158, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004248.

Chng:2005:MME

[CP05] H. N. Ch’ng and J. Pan. Modelling microstructural evolution of porous polycrystalline materials and a numerical study of anisotropic sintering. Journal of Computational Physics, 204(2):430–461, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004231.

Cao:2006:ALM

[CP06a] Yang Cao and Linda Petzold. Accuracy limitations and the measurement of errors in the stochastic simulation of chemically reacting systems. Journal of Computational Physics, 212(1):6–24, February 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003074.

Chen:2006:PIEa

[CP06b] Wenwu Chen and Bill Poirier. Parallel implementation of efficient preconditioned linear solver for grid-based applications in chemical physics. I: Block Jacobi diagonalization. Journal of Computational Physics, 219(1):185–197, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002154.

Chen:2006:PIEb

[CP06c] Wenwu Chen and Bill Poirier. Parallel implementation of efficient preconditioned linear solver for grid-based applications in chemical physics. II: QMR linear solver. Journal of Computational Physics, 219(1):198–209, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001471.

Chen:2007:EGP

[CP07] Yang Chen and Scott E. Parker. Electromagnetic gyrokinetic δf particle-in-cell turbulence simulation with realistic equilibrium profiles and geometry. Journal of Computational Physics, 220(2):839–855, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002634.

Christov:2008:NNO

[CP08] Ivan Christov and Bojan Popov. New non-oscillatory central schemes on unstructured triangulations for hyperbolic systems of conservation laws. Journal of Computational Physics, 227(11):5736–5757, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001071.

Carcione:2004:DWS

[CPG04] José M. Carcione, Flavio Poletto, and Davide Gei. 3-D wave simulation in anelastic media using the Kelvin–Voigt constitutive equation. Journal of Computational Physics, 196(1):282– 297, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005977.

Chaniotis:2002:RSP

[CPK02] A. K. Chaniotis, D. Poulikakos, and P. Koumoutsakos. Remeshed smoothed particle hydrodynamics for the simulation of viscous and heat conducting flows. Journal of Computational Physics, 182(1):67–90, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971524.

Celnik:2009:PCA

[CPKW09] Matthew Celnik, Robert Patterson, Markus Kraft, and Wolfgang Wagner. A predictor-corrector algorithm for the coupling of stiff ODEs to a particle population balance. Journal of Computational Physics, 228(8):2758–2769, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006542.

Coppa:2002:ICM

[CPP02] Gianni G. M. Coppa, Fabio Peano, and Federico Peinetti. Image-charge method for contour dynamics in systems with cylindrical boundaries. Journal of Computational Physics, 182(2):392–417, November 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971743.

Caboussat:2005:NSF

[CPR05] A. Caboussat, M. Picasso, and J. Rappaz. Numerical simulation of free surface incompressible liquid flows surrounded by compressible gas. Journal of Computational Physics, 203(2):626–649, March 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003961.

Cerne:2001:CIT

[ČPT01] Gregor Černe, Stojan Petelin, and Iztok Tiselj. Coupling of the interface tracking and the two-fluid models for the simulation of incompressible two-phase flow. Journal of Computational Physics, 171(2):776–804, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196810X.

Cecil:2004:NMH

[CQO04] Tom Cecil, Jianliang Qian, and Stanley Osher. Numerical methods for high dimensional Hamilton–Jacobi equations using radial basis functions. Journal of Computational Physics, 196(1):327–347, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005990.

Cockburn:2005:ASD

[CQRW05] Bernardo Cockburn, Jianliang Qian, Fernando Reitich, and Jing Wang. An accurate spectral/discontinuous finiteelement formulation of a phase-space-based level set approach to geometrical optics. Journal of Computational Physics, 208(1):175–195, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000914.

Colin:2000:DHO

[CR00] Olivier Colin and Michael Rudgyard. Development of highorder Taylor–Galerkin schemes for LES. Journal of Computational Physics, 162(2):338–371, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965380.

Colonius:2002:SGS

[CR02] Tim Colonius and Hongyu Ran. A super-grid-scale model for simulating compressible flow on unbounded domains. Journal of Computational Physics, 182(1):191–212, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971615.

Ceniceros:2005:MPF

[CR05] Hector D. Ceniceros and Alexandre M. Roma. A multiphase flow method with a fast, geometry-based fluid indicator. Journal of Computational Physics, 205(2):391–400, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004759.

Ceniceros:2007:NAM

[CR07] Hector D. Ceniceros and Alexandre M. Roma. A nonstiff, adaptive mesh refinement-based method for the Cahn– Hilliard equation. Journal of Computational Physics, 225 (2):1849–1862, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107000903.

Carpio:2008:DRU

[CR08] Ana Carpio and María-Luisa Rapún. Domain reconstruction using photothermal techniques. Journal of Computational Physics, 227(17):8083–8106, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002830.

Clain:2009:FSO

[CR09] Stéphane Clain and David Rochette. First- and secondorder finite volume methods for the one-dimensional nonconservative Euler system. Journal of Computational Physics, 228(22):8214–8248, December 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004185.

Cassiani:2007:EAS

[CRAG07] M. Cassiani, A. Radicchi, J. D. Albertson, and U. Giostra. An efficient algorithm for scalar PDF modelling in incompressible turbulent flow; numerical analysis with evaluation of IEM and IECM micro-mixing models. Journal of Computational Physics, 223(2):519–550, May 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004517.

Caramana:2000:CES

[CRB00] E. J. Caramana, C. L. Rousculp, and D. E. Burton. A compatible, energy and symmetry preserving Lagrangian hydrodynamics algorithm in three-dimensional Cartesian geometry. Journal of Computational Physics, 157(1):89–119, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199963684.

Cizmas:2008:ATR

[CRB^{+}08] Paul G. A. Cizmas, Brian R. Richardson, Thomas A. Brenner, Thomas J. O’Brien, and Ronald W. Breault. Acceleration techniques for reduced-order models based on proper orthogonal decomposition. Journal of Computational Physics, 227(16):7791–7812, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002647.

Csiik:2002:CFM

[CRD02] Árpád Csík, Mario Ricchiuto, and Herman Deconinck. A conservative formulation of the multidimensional upwind residual distribution schemes for general nonlinear conservation laws. Journal of Computational Physics, 179(1):286–312, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970579.

Ceniceros:2000:CAS

[CS00] Hector D. Ceniceros and Helen Si. Computation of axisymmetric suction flow through porous media in the presence of surface tension. Journal of Computational Physics, 165(1):237–260, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966130.

Campbell:2001:TAV

[CS01a] J. C. Campbell and M. J. Shashkov. A tensor artificial viscosity using a mimetic finite difference algorithm. Journal of Computational Physics, 172(2):739–765, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968561.

Cano:2001:USH

[CS01b] B. Cano and A. M. Stuart. Underresolved simulations of heat baths. Journal of Computational Physics, 169(1):193–214, May 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967221.

Carcione:2001:WSF

[CS01c] José M. Carcione and Géza Seriani. Wave simulation in frozen porous media. Journal of Computational Physics, 170(2):676– 695, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967567.

Clouet:2003:HSM

[CS03] J.-F. Clouët and G. Samba. A hybrid symbolic Monte– Carlo method for radiative transfer equations. Journal of Computational Physics, 188(1):139–156, June 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300158X.

Clouet:2004:ADL

[CS04] J.-F. Clouët and G. Samba. Asymptotic diffusion limit of the symbolic Monte–Carlo method for the transport equation.

Journal of Computational Physics, 195(1):293–319, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005333.

Chalikov:2005:MEW

[CS05] Dmitry Chalikov and Dmitry Sheinin. Modeling extreme waves based on equations of potential flow with a free surface. Journal of Computational Physics, 210(1):247–273, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002196.

Chou:2006:HOR

[CS06] Ching-Shan Chou and Chi-Wang Shu. High order residual distribution conservative finite difference WENO schemes for steady state problems on non-smooth meshes. Journal of Computational Physics, 214(2):698–724, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004699.

Cheng:2007:HOE

[CS07a] Juan Cheng and Chi-Wang Shu. A high order ENO conservative Lagrangian type scheme for the compressible Euler equations. Journal of Computational Physics, 227(2):1567–1596, December 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700424X.

Cheng:2007:DGF

[CS07b] Yingda Cheng and Chi-Wang Shu. A discontinuous Galerkin finite element method for directly solving the Hamilton–Jacobi equations. Journal of Computational Physics, 223(1):398–415, April 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004451.

Chern:2007:CIM

[CS07c] I-Liang Chern and Yu-Chen Shu. A coupling interface method for elliptic interface problems. Journal of Computational Physics, 225(2):2138–2174, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001246.

Chou:2007:HOR

[CS07d] Ching-Shan Chou and Chi-Wang Shu. High order residual distribution conservative finite difference WENO schemes for convection-diffusion steady state problems on non-smooth meshes. Journal of Computational Physics, 224(2):992–1020, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005638.

Chen:2008:PPD

[CS08a] Tianbing Chen and John Strain. Piecewise-polynomial discretization and Krylov-accelerated multigrid for elliptic interface problems. Journal of Computational Physics, 227(16):7503–7542, August 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002453.

Cheng:2008:STE

[CS08b] Yingda Cheng and Chi-Wang Shu. Superconvergence and time evolution of discontinuous Galerkin finite element solutions. Journal of Computational Physics, 227 (22):9612–9627, November 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003835.

Colella:2008:LPP

[CS08c] Phillip Colella and Michael D. Sekora. A limiter for PPM that preserves accuracy at smooth extrema. Journal of Computational Physics, 227(15):7069–7076, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001435.

Chiu:2009:DDR

[CS09] P. H. Chiu and Tony W. H. Sheu. On the development of a dispersion-relation-preserving dual-compact upwind scheme for convection-diffusion equation. Journal of Computational Physics, 228(10):3640–3655, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900062X.

Carmichael:2008:PAQ

[CSC^{+}08] Gregory R. Carmichael, Adrian Sandu, Tianfeng Chai, Dacian N. Daescu, Emil M. Constantinescu, and Youhua Tang. Predicting air quality: Improvements through advanced methods to integrate models and measurements. Journal of Computational Physics, 227(7):3540–3571, March 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000836.

Ciardi:2005:DFV

[CSKD05] M. Ciardi, P. Sagaut, M. Klein, and W. N. Dawes. A dynamic finite volume scheme for large-eddy simulation on unstructured grids. Journal of Computational Physics, 210(2):632–655, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500255X.

Chiu:2008:EEP

[CSL08] P. H. Chiu, Tony W. H. Sheu, and R. K. Lin. An effective explicit pressure gradient scheme implemented in the two-level non-staggered grids for incompressible Navier–Stokes equations. Journal of Computational Physics, 227(8):4018–4037, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005591.

Choi:2005:ECS

[CSMH05] Yunseok Choi, Changsoo Shin, Dong-Joo Min, and Taeyoung Ha. Efficient calculation of the steepest descent direction for source-independent seismic waveform inversion: An amplitude approach. Journal of Computational Physics, 208(2):455–468, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001051.

Cooke:2006:FIM

[CSML06] S. J. Cooke, R. Shtokhamer, A. A. Mondelli, and B. Levush. A finite integration method for conformal, structuredgrid, electromagnetic simulation. Journal of Computational Physics, 215(1):321–347, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004900.

Chenoweth:2009:SAM

[CSO09] Samuel K. M. Chenoweth, Julio Soria, and Andrew Ooi. A singularity-avoiding moving least squares scheme for twodimensional unstructured meshes. Journal of Computational Physics, 228(15):5592–5619, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002320.

Coppola:2001:NPT

[CSP01] G. Coppola, S. J. Sherwin, and J. Peiró. Nonlinear particle tracking for high-order elements. Journal of Computational Physics, 172(1):356–386, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968299.

Chen:2000:DSC

[CSS00] Heli Chen, Yuhong Su, and Bernie D. Shizgal. A direct spectral collocation Poisson solver in polar and cylindrical coordinates. Journal of Computational Physics, 160(2):453–469, May 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964611.

Coleman:2000:ECJ

[CSV00] Thomas F. Coleman, Fadil Santosa, and Arun Verma. Efficient calculation of Jacobian and adjoint vector products in the wave propagational inverse problem using automatic differentiation. Journal of Computational Physics, 157(1): 234–255, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199963738.

Chan:2004:LST

[CT04] Tony F. Chan and Xue-Cheng Tai. Level set and total variation regularization for elliptic inverse problems with discontinuous coefficients. Journal of Computational Physics, 193(1):40–66, January 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004169.

Codecasa:2007:CED

[CT07] Lorenzo Codecasa and Francesco Trevisan. Constitutive equations for discrete electromagnetic problems over polyhedral grids. Journal of Computational Physics, 225(2):1894–1918, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000927.

Castro:2008:SHO

[CT08a] C. E. Castro and E. F. Toro. Solvers for the high-order Riemann problem for hyperbolic balance laws. Journal of Computational Physics, 227(4):2481–2513, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004780.

Cheng:2008:RFT

[CT08b] Li-Tien Cheng and Yen-Hsi Tsai. Redistancing by flow of time dependent eikonal equation. Journal of Computational Physics, 227(8):4002–4017, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700558X.

Cada:2009:CTO

[CT09] Miroslav Cada and Manuel Torrilhon. Compact third-order limiter functions for finite volume methods. Journal of Computational Physics, 228(11):4118–4145, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000953.

Chance:2007:CVG

[CTS07] M. S. Chance, A. D. Turnbull, and P. B. Snyder. Calculation of the vacuum Green’s function valid even for high toroidal mode numbers in tokamaks. Journal of Computational Physics, 221(1):330–348, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002920.

Chen:2008:SIP

[CTT08] Qingshan Chen, Roger Temam, and Joseph J. Tribbia. Simulations of the 2.5D inviscid primitive equations in a limited domain. Journal of Computational Physics, 227(23):9865–9884, December 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004208.

Cave:2008:IUS

[CTW^{+}08] H. M. Cave, K.-C. Tseng, J.-S. Wu, M. C. Jermy, J.- C. Huang, and S. P. Krumdieck. Implementation of unsteady sampling procedures for the parallel direct simulation Monte Carlo method. Journal of Computational Physics, 227(12):6249–6271, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800154X.

Cui:2009:CFD

[Cui09] Mingrong Cui. Compact finite difference method for the fractional diffusion equation. Journal of Computational Physics, 228(20):7792–7804, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004094.

Cullen:2001:PTT

[Cul01] Jason Cullen. Photon transport through plasmas with density and velocity structure. Journal of Computational Physics, 173(1):175–186, October 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968706.

Chatterjee:2006:TAS

[CV06] Abhijit Chatterjee and Dionisios G. Vlachos. Temporal acceleration of spatially distributed kinetic Monte Carlo simulations. Journal of Computational Physics, 211(2):596–615, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002895.

Cartwright:2000:LIM

[CVB00] K. L. Cartwright, J. P. Verboncoeur, and C. K. Birdsall. Loading and injection of Maxwellian distributions in particle simulations. Journal of Computational Physics, 162(2):483–513, August 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965495.

Caleffi:2006:FOB

[CVB06] V. Caleffi, A. Valiani, and A. Bernini. Fourth-order balanced source term treatment in central WENO schemes for shallow water equations. Journal of Computational Physics, 218(1):228–245, October 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000799.

Crommelin:2006:FTC

[CVE06] D. T. Crommelin and E. Vanden-Eijnden. Fitting timeseries by continuous-time Markov chains: a quadratic programming approach. Journal of Computational Physics, 217(2):782–805, September 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000404.

Candy:2003:EGM

[CW03] J. Candy and R. E. Waltz. An Eulerian gyrokinetic-Maxwell solver. Journal of Computational Physics, 186(2):545–581, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000792.

Cheng:2008:EAS

[CW08] Mowei Cheng and James A. Warren. An efficient algorithm for solving the phase field crystal model. Journal of Computational Physics, 227(12):6241–6248, June 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001538.

Cocle:2008:CVC

[CWD08] Roger Cocle, Grégoire Winckelmans, and Goéric Daeninck. Combining the vortex-in-cell and parallel fast multipole methods for efficient domain decomposition simulations. Journal of Computational Physics, 227(21):9091–9120, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004482.

Capolino:2007:ECG

[CWJ07] F. Capolino, D. R. Wilton, and W. A. Johnson. Efficient computation of the 3D Green’s function for the Helmholtz operator for a linear array of point sources using the Ewald method. Journal of Computational Physics, 223(1):250–261, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004359.

Chang:2008:NCL

[CWL08] Qianshun Chang, Yau-Shu Wong, and Chi-Kun Lin. Numerical computations for long-wave short-wave interaction equations in semi-classical limit. Journal of Computational Physics, 227(19):8489–8507, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002854.

Chang:2000:AST

[CWT00] Sin-Chung Chang, Xiao-Yen Wang, and Wai-Ming To. Application of the space–time conservation element and solution element method to one-dimensional convectiondiffusion problems. Journal of Computational Physics, 165(1):189–215, November 20, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966105.

Canning:2000:PEP

[CWWZ00] A. Canning, L. W. Wang, A. Williamson, and A. Zunger. Parallel empirical pseudopotential electronic structure calculations for million atom systems. Journal of Computational Physics, 160(1):29–41, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964404.

Chen:2008:EMP

[CWYM08] Qian-Yong Chen, Jing Wan, Yahan Yang, and Rick T. Mifflin. Enriched multi-point flux approximation for general grids. Journal of Computational Physics, 227(3):1701–1721, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004263.

Chen:2008:SWM

[CX08] Chungang Chen and Feng Xiao. Shallow water model on cubed-sphere by multi-moment finite volume method. Journal of Computational Physics, 227(10):5019–5044, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000570.

Cai:2008:NFB

[CXB08] Wei Cai, Zhenli Xu, and Andrij Baumketner. A new FFTbased algorithm to compute Born radii in the generalized Born theory of biomolecule solvation. Journal of Computational Physics, 227(24):10162–10177, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004488.

Chen:2009:AII

[CXZ09] Zhiming Chen, Yuanming Xiao, and Linbo Zhang. The adaptive immersed interface finite element method for elliptic and Maxwell interface problems. Journal of Computational Physics, 228(14):5000–5019, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001612.

Cai:2000:FCD

[CY00] Wei Cai and Tiejun Yu. Fast calculations of dyadic Green’s functions for electromagnetic scattering in a multilayered medium. Journal of Computational Physics, 165(1):1–21, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100965835.

Cockburn:2005:AMR

[CY05] Bernardo Cockburn and Bayram Yenikaya. An adaptive method with rigorous error control for the Hamilton–Jacobi equations. Part II: The two-dimensional steady-state case. Journal of Computational Physics, 209(2):391–405, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001063.

Choe:2001:NUS

[CYKC01] Hee-Hwan Choe, N. S. Yoon, S. S. Kim, and Duk-In Choi. A new unconditionally stable algorithm for steady-state fluid simulation of high density plasma discharge. Journal of Computational Physics, 170(2):550–561, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967488.

Chew:2006:GFD

[CYS06] C. S. Chew, K. S. Yeo, and C. Shu. A generalized finitedifference (GFD) ALE scheme for incompressible flows around moving solid bodies on hybrid meshfree-Cartesian grids. Journal of Computational Physics, 218(2):510–548, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001173.

Chisholm:2009:JFN

[CZ09] Todd T. Chisholm and David W. Zingg. A Jacobian-free Newton–Krylov algorithm for compressible turbulent fluid flows. Journal of Computational Physics, 228(9):3490–3507, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000643.

Crnjaric-Zic:2004:BFV

[CZVS04] Nelida Crnjaríc-Zic, Senka Vukovíc, and Luka Sopta. Balanced finite volume WENO and central WENO schemes for the shallow water and the open-channel flow equations. Journal of Computational Physics, 200(2):512–548, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001780.

deAbreu:2004:MSR

[dA04] Marcos Pimenta de Abreu. Mixed singular-regular boundary conditions in multislab radiation transport. Journal of Computational Physics, 197(1):167–185, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006247.

Desideri:2007:NSA

[DAJ07] Jean-Antoine Désidéri, Badr Abou El Majd, and Ales Janka. Nested and self-adaptive Bézier parameterizations for shape optimization. Journal of Computational Physics, 224(1):117–131, May 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006188.

Daripa:2000:CSR

[Dar00a] Prabir Daripa. A computational study of rising plane Taylor bubbles. Journal of Computational Physics, 157(1):120–142, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199963660.

Darve:2000:FMM

[Dar00b] Eric Darve. The fast multipole method: Numerical implementation. Journal of Computational Physics, 160(1):195–240, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964519.

Darrigrand:2002:CFM

[Dar02] Eric Darrigrand. Coupling of fast multipole method and microlocal discretization for the 3-D Helmholtz equation. Journal of Computational Physics, 181(1):126–154, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970919.

Dukowicz:2000:IRT

[DB00] John K. Dukowicz and John R. Baumgardner. Incremental remapping as a transport/advection algorithm. Journal of Computational Physics, 160(1):318–335, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964659.

Domenichini:2004:FNS

[DB04] F. Domenichini and B. Baccani. A formulation of Navier– Stokes problem in cylindrical coordinates applied to the 3D entry jet in a duct. Journal of Computational Physics, 200(1):177–191, October 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001494.

Dossou:2006:FEC

[DBB06] Kokou Dossou, Michael A. Byrne, and Lindsay C. Botten. Finite element computation of grating scattering matrices and application to photonic crystal band calculations. Journal of Computational Physics, 219(1):120–143, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001434.

Desjardins:2008:HOC

[DBBP08] Olivier Desjardins, Guillaume Blanquart, Guillaume Balarac, and Heinz Pitsch. High order conservative finite difference scheme for variable density low Mach number turbulent flows. Journal of Computational Physics, 227(15):7125–7159, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001666.

DeZaeytijd:2008:EHM

[DBF08] J. De Zaeytijd, I. Bogaert, and A. Franchois. An efficient hybrid MLFMA–FFT solver for the volume integral equation in case of sparse 3D inhomogeneous dielectric scatterers. Journal of Computational Physics, 227(14):7052–7068, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910800226X.

Dworkin:2006:MCV

[DBS06] S. B. Dworkin, B. A. V. Bennett, and M. D. Smooke. A mass-conserving vorticity-velocity formulation with application to nonreacting and reacting flows. Journal of Computational Physics, 215(2):430–447, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004961.

Dumbser:2008:UFC

[DBTM08] Michael Dumbser, Dinshaw S. Balsara, Eleuterio F. Toro, and Claus-Dieter Munz. A unified framework for the construction of one-step finite volume and discontinuous Galerkin schemes on unstructured meshes. Journal of Computational Physics, 227(18):8209–8253, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002829.

Davies:2001:NVV

[DC01] Christopher Davies and Peter W. Carpenter. A novel velocityvorticity formulation of the Navier–Stokes equations with applications to boundary layer disturbance evolution. Journal of Computational Physics, 172(1):119–165, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968172.

Dupuis:2002:LGM

[DC02] Alexandre Dupuis and Bastien Chopard. Lattice gas modeling of scour formation under submarine pipelines. Journal of Computational Physics, 178(1):161–174, May 1, 2002.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970257.

Deng:2007:EFM

[DC07] Shaozhong Deng and Wei Cai. Extending the fast multipole method for charges inside a dielectric sphere in an ionic solvent: high-order image approximations for reaction fields. Journal of Computational Physics, 227(2):1246–1266, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003920.

Dorini:2008:SMR

[DC08] Fábio A. Dorini and M. Cristina C. Cunha. Statistical moments of the random linear transport equation. Journal of Computational Physics, 227(19):8541–8550, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003173.

Delzanno:2008:ORE

[DCF^{+}08] G. L. Delzanno, L. Chacón, J. M. Finn, Y. Chung, and G. Lapenta. An optimal robust equidistribution method for two-dimensional grid adaptation based on Monge– Kantorovich optimization. Journal of Computational Physics, 227(23):9841–9864, December 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004105.

Dupuis:2008:IBL

[DCK08] Alexandre Dupuis, Philippe Chatelain, and Petros Koumoutsakos. An immersed boundary-lattice-Boltzmann method for the simulation of the flow past an impulsively started cylinder. Journal of Computational Physics, 227(9):4486–4498, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108000247.

del-Castillo-Negrete:2007:CMS

[dCNHSD07] D. del Castillo-Negrete, S. P. Hirshman, D. A. Spong, and E. F. D’Azevedo. Compression of magnetohydrodynamic simulation data using singular value decomposition. Journal of Computational Physics, 222(1):265–286, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600355X.

Daescu:2000:AIR

[DCS00] Dacian Daescu, Gregory R. Carmichael, and Adrian Sandu. Adjoint implementation of Rosenbrock methods applied to variational data assimilation problems. Journal of Computational Physics, 165(2):496–510, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966221.

DeSterck:2001:STD

[DCV^{+}01] H. De Sterck, A. Csík, D. Vanden Abeele, S. Poedts, and H. Deconinck. Stationary two-dimensional magnetohydrodynamic flows with shocks: Characteristic analysis and grid convergence study. Journal of Computational Physics, 166(1):28–62, January 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966403.

Dufresne:2003:ESB

[DD03a] L. Dufresne and G. Dumas. Erratum to “A spectral/Bspline method for the Navier–Stokes equations in unbounded domains” [J. Comp. Phys. 185 (2003) 532–548]. Journal of Computational Physics, 187(1):369, May 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000809. See [DD03b].

Dufresne:2003:SBS

[DD03b] L. Dufresne and G. Dumas. A spectral/B-spline method for the Navier–Stokes equations in unbounded domains. Journal of Computational Physics, 185(2):532–548, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000020. See erratum [DD03a].

Delyon:2005:CEP

[DD05] François Delyon and Michel Duneau. Computation of eigenmodes of photonic crystals by inversion of the Maxwell operator. Journal of Computational Physics, 207(1):375–387, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000331.

Diosady:2009:PMD

[DD09] Laslo T. Diosady and David L. Darmofal. Preconditioning methods for discontinuous Galerkin solutions of the Navier–Stokes equations. Journal of Computational Physics, 228(11):3917–3935, June 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000680.

Dauger:2005:UST

[DDD05] D. E. Dauger, V. K. Decyk, and J. M. Dawson. Using semiclassical trajectories for the time-evolution of interacting quantum-mechanical systems. Journal of Computational Physics, 209(2):559–581, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001877.

Dowling:2007:MCT

[DDDC07] Mark R. Dowling, Matthew J. Davis, Peter D. Drummond, and Joel F. Corney. Monte Carlo techniques for real-time quantum dynamics. Journal of Computational Physics, 220(2):549–567, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600249X.

deDieuleveult:2009:GSS

[dDEK09] C. de Dieuleveult, J. Erhel, and M. Kern. A global strategy for solving reactive transport equations. Journal of Computational Physics, 228(17):6395–6410, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002824.

Davies:2001:AEA

[DDF01] Penny J. Davies, Dugald B. Duncan, and Stefan A. Funken. Accurate and efficient algorithms for frequency domain scattering from a thin wire. Journal of Computational Physics, 168(1):155–183, March 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966889.

Duclous:2009:HOR

[DDFT09] Roland Duclous, Bruno Dubroca, Francis Filbet, and Vladimir Tikhonchuk. High order resolution of the Maxwell– Fokker–Planck–Landau model intended for ICF applications. Journal of Computational Physics, 228(14):5072–5100, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001879.

Deang:2002:MCR

[DDG02] Jennifer Deang, Qiang Du, and Max D. Gunzburger. Modeling and computation of random thermal fluctuations and material defects in the Ginzburg–Landau model for superconductivity. Journal of Computational Physics, 181(1):45–67, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971287.

Doumic:2009:SLB

[DDGS09] Marie Doumic, Frédéric Duboc, François Golse, and Rémi Sentis. Simulation of laser beam propagation with a paraxial model in a tilted frame. Journal of Computational Physics, 228(3):861–880, February 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005287.

Ditkowski:2001:CCG

[DDH01] A. Ditkowski, K. Dridi, and J. S. Hesthaven. Convergent Cartesian grid methods for Maxwell’s equations in complex geometries. Journal of Computational Physics, 170(1):39–80, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967191.

Diamessis:2005:SMP

[DDH05] P. J. Diamessis, J. A. Domaradzki, and J. S. Hesthaven. A spectral multidomain penalty method model for the simulation of high Reynolds number localized incompressible stratified turbulence. Journal of Computational Physics, 202(1):298–322, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002876.

Darmana:2006:PEL

[DDK06] D. Darmana, N. G. Deen, and J. A. M. Kuipers. Parallelization of an Euler–Lagrange model using mixed domain decomposition and a mirror domain technique: application to dispersed gas-liquid two-phase flow. Journal of Computational Physics, 220(1):216–248, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002294.

Degond:2007:MIM

[DDM07] Pierre Degond, Giacomo Dimarco, and Luc Mieussens. A moving interface method for dynamic kinetic-fluid coupling. Journal of Computational Physics, 227(2):1176–1208, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003890.

Diwakar:2009:QSB

[DDS09] S. V. Diwakar, Sarit K. Das, and T. Sundararajan. A quadratic spline based interface (QUASI) reconstruction algorithm for accurate tracking of two-phase flows. Journal of Computational Physics, 228(24):9107–9130, December 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004999.

Degond:2009:APS

[DDSV09] P. Degond, F. Deluzet, A. Sangam, and M.-H. Vignal. An asymptotic preserving scheme for the Euler equations in a strong magnetic field. Journal of Computational Physics, 228(10):3540–3558, June 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000035.

deLaBourdonnaye:2000:HOS

[de 00] Armel de La Bourdonnaye. High-order scheme for a nonlinear Maxwell system modelling Kerr effect. Journal of Computational Physics, 160(2):500–521, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964684.

DeVuyst:2004:SAH

[De 04] Florian De Vuyst. Stable and accurate hybrid finite volume methods based on pure convexity arguments for hyperbolic systems of conservation law. Journal of Computational Physics, 193(2):426–468, January 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003747.

Degond:2002:CSD

[DE02] P. Degond and A. El Ayyadi. A coupled Schrödinger driftdiffusion model for quantum semiconductor device simulations. Journal of Computational Physics, 181(1):222– 259, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971226.

De:2006:ANH

[DE06] A. K. De and V. Eswaran. Analysis of a new high resolution upwind compact scheme. Journal of Computational Physics, 218(1):398–416, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000908.

Dehnen:2002:HFC

[Deh02] Walter Dehnen. A hierarchical O(N) force calculation algorithm. Journal of Computational Physics, 179(1):27–42, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970269.

Dostert:2006:CGL

[DEHL06] P. Dostert, Y. Efendiev, T. Y. Hou, and W. Luo. Coarsegradient Langevin algorithms for dynamic data integration and uncertainty quantification. Journal of Computational Physics, 217(1):123–142, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001380.

Dellar:2001:NMM

[Del01] Paul J. Dellar. A note on magnetic monopoles and the onedimensional MHD Riemann problem. Journal of Computational Physics, 172(1):392–398, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968159.

Dellar:2002:LKS

[Del02] Paul J. Dellar. Lattice kinetic schemes for magnetohydrodynamics. Journal of Computational Physics, 179(1):95–126, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970440.

Dellacherie:2003:CWC

[Del03a] Stéphane Dellacherie. Coupling of the Wang Chang– Uhlenbeck equations with the multispecies Euler system. Journal of Computational Physics, 189(1):239–276, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002109.

Dellar:2003:ILL

[Del03b] Paul J. Dellar. Incompressible limits of lattice Boltzmann equations using multiple relaxation times. Journal of Computational Physics, 190(2):351–370, September 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002791.

Dellacherie:2007:NRP

[Del07] Stéphane Dellacherie. Numerical resolution of a potential diphasic low Mach number system. Journal of Computational Physics, 223(1):151–187, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004311.

deMello:2002:HCS

[deM02] J. C. deMello. Highly convergent simulations of transport dynamics in organic light-emitting diodes. Journal of Computational Physics, 181(2):564–576, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971421.

Demir:2004:NMC

[Dem04] Alper Demir. Non-Monte Carlo formulations and computational techniques for the stochastic non-linear Schrödinger equation. Journal of Computational Physics, 201(1):148– 171, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104002050.

Deng:2007:NAT

[Den07] Weihua Deng. Numerical algorithm for the time fractional Fokker–Planck equation. Journal of Computational Physics, 227(2):1510–1522, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004226.

Dumbser:2008:FVS

[DET08] Michael Dumbser, Cedric Enaux, and Eleuterio F. Toro. Finite volume schemes of very high order of accuracy for stiff hyperbolic balance laws. Journal of Computational Physics, 227(8):3971–4001, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005578.

Dimas:2000:LWS

[DF00a] Athanassios A. Dimas and Laurie T. Fialkowski. Largewave simulation (LWS) of free-surface flows developing weak spilling breaking waves. Journal of Computational Physics, 159(2):172–196, April 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910096426X.

Driscoll:2000:NNF

[DF00b] Tobin A. Driscoll and Bengt Fornberg. Note on nonsymmetric finite differences for Maxwell’s equations. Journal of Computational Physics, 161(2):723–727, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965240.

Dolejsi:2004:SID

[DF04] V. Dolejší and M. Feistauer. A semi-implicit discontinuous Galerkin finite element method for the numerical solution of inviscid compressible flow. Journal of Computational Physics, 198(2):727–746, August 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000609.

Dubinkina:2007:SMA

[DF07] Svetlana Dubinkina and Jason Frank. Statistical mechanics of Arakawa’s discretizations. Journal of Computational Physics, 227(2):1286–1305, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003944.

Du:2006:SPF

[DFG^{+}06] Jian Du, Brian Fix, James Glimm, Xicheng Jia, Xiaolin Li, Yuanhua Li, and Lingling Wu. A simple package for front tracking. Journal of Computational Physics, 213(2):613–628, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003955.

deFalco:2005:QCD

[dFGLS05] Carlo de Falco, Emilio Gatti, Andrea L. Lacaita, and Riccardo Sacco. Quantum-corrected drift-diffusion models for transport in semiconductor devices. Journal of Computational Physics, 204(2):533–561, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004279.

deFalco:2009:QCD

[dFJS09] Carlo de Falco, Joseph W. Jerome, and Riccardo Sacco. Quantum-corrected drift-diffusion models: Solution fixed point map and finite element approximation. Journal of Computational Physics, 228(5):1770–1789, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005937.

Djellouli:2001:FMS

[DFT01] Rabia Djellouli, Charbel Farhat, and Radek Tezaur. A fast method for solving acoustic scattering problems in frequency bands. Journal of Computational Physics, 168(2):412–432, April 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967075.

Desjardins:2008:QBM

[DFV08] O. Desjardins, R. O. Fox, and P. Villedieu. A quadraturebased moment method for dilute fluid-particle flows. Journal of Computational Physics, 227(4):2514–2539, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004792.

Ditkowski:2009:GRM

[DG09] Adi Ditkowski and Nir Gavish. A grid redistribution method for singular problems. Journal of Computational Physics, 228(7):2354–2365, April 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800627X.

Donev:2008:SED

[DGA08] Aleksandar Donev, Alejandro L. Garcia, and Berni J. Alder. Stochastic event-driven molecular dynamics. Journal of Computational Physics, 227(4):2644–2665, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004962.

Das:2009:PCR

[DGF09] Sonjoy Das, Roger Ghanem, and Steven Finette. Polynomial chaos representation of spatio-temporal random fields from experimental measurements. Journal of Computational Physics, 228(23):8726–8751, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004677.

Dorr:2002:SLP

[DGH02] Milo R. Dorr, F. Xabier Garaizar, and Jeffrey A. F. Hittinger. Simulation of laser plasma filamentation using adaptive mesh refinement. Journal of Computational Physics, 177(2):233–263, April 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969852.

Deng:2008:UMC

[DGH08] Weibing Deng, Ji Gu, and Jianmin Huang. Upscaling methods for a class of convection-diffusion equations with highly oscillating coefficients. Journal of Computational Physics, 227(16):7621–7642, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002507.

Don:2003:MSM

[DGJ03] Wai-Sun Don, David Gottlieb, and Jae-Hun Jung. A multidomain spectral method for supersonic reactive flows. Journal of Computational Physics, 192(1):325–354, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003978.

Degond:2007:EQD

[DGM07] Pierre Degond, Samy Gallego, and Florian Méhats. An entropic quantum drift-diffusion model for electron transport in resonant tunneling diodes. Journal of Computational Physics, 221(1):226–249, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002865.

Diaz-Goano:2003:FDF

[DGMN03] C. Diaz-Goano, P. D. Minev, and K. Nandakumar. A fictitious domain/finite element method for particulate flows. Journal of Computational Physics, 192(1):105–123, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003498.

Dobson:2000:EMB

[DGP00] David C. Dobson, Jayadeep Gopalakrishnan, and Joseph E. Pasciak. An efficient method for band structure calculations in 3D photonic crystals. Journal of Computational Physics, 161(2):668–679, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965215.

Domingues:2008:AMS

[DGRS08] Margarete O. Domingues, Sônia M. Gomes, Olivier Roussel, and Kai Schneider. An adaptive multiresolution scheme with local time stepping for evolutionary PDEs. Journal of Computational Physics, 227(8):3758–3780, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005219.

Darve:2004:EFM

[DH04] Eric Darve and Pascal Havé. Efficient fast multipole method for low-frequency scattering. Journal of Computational Physics, 197(1):341–363, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006326.

Dohnal:2007:PML

[DH07] Tomás Dohnal and Thomas Hagstrom. Perfectly matched layers in photonics computations: 1D and 2D nonlinear coupled mode equations. Journal of Computational Physics, 223(2):690–710, May 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004694.

Denton:2009:SBC

[DH09] R. E. Denton and Y. Hu. Symmetry boundary conditions. Journal of Computational Physics, 228(13):4823–4835, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001673.

Dorband:2003:SHS

[DHM03] Ernst Nils Dorband, Marc Hemsendorf, and David Merritt. Systolic and hyper-systolic algorithms for the gravitational N-body problem, with an application to Brownian motion. Journal of Computational Physics, 185(2):484–511, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000670.

Doom:2007:NMD

[DHM07] Jeff Doom, Yucheng Hou, and Krishnan Mahesh. A numerical method for DNS/LES of turbulent reacting flows. Journal of Computational Physics, 226(1):1136–1151, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002252.

Damanik:2009:MFM

[DHOT09] H. Damanik, J. Hron, A. Ouazzi, and S. Turek. A monolithic FEM-multigrid solver for non-isothermal incompressible flow on general meshes. Journal of Computational Physics, 228(10):3869–3881, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000941.

deHoop:2007:WEC

[dHRvdB07] Adrianus T. de Hoop, Robert F. Remis, and Peter M. van den Berg. The 3D wave equation and its Cartesian coordinate stretched perfectly matched embedding — a time-domain Green’s function performance analysis. Journal of Computational Physics, 221(1):88–105, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002762.

Durbin:2002:ALR

[DI02] P. A. Durbin and G. Iaccarino. An approach to local refinement of structured grids. Journal of Computational Physics, 181(2):639–653, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971470.

Doostan:2009:LSA

[DI09] Alireza Doostan and Gianluca Iaccarino. A least-squares approximation of partial differential equations with highdimensional random inputs. Journal of Computational Physics, 228(12):4332–4345, July 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001144.

Dickinson:2008:DMS

[Dic08] Robert E. Dickinson. Determination of the multi-scattered solar radiation from a leaf canopy for use in climate models. Journal of Computational Physics, 227(7):3667–3677, March 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005608.

Deng:2003:TDE

[DIL03] Shaozhong Deng, Kazufumi Ito, and Zhilin Li. Threedimensional elliptic solvers for interface problems and applications. Journal of Computational Physics, 184(1):215–243, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000281.

Dimitrakopoulos:2007:IDS

[Dim07] P. Dimitrakopoulos. Interfacial dynamics in Stokes flow via a three-dimensional fully-implicit interfacial spectral boundary element algorithm. Journal of Computational Physics, 225(1):408–426, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005985.

Drikakis:2000:AMF

[DIV00] D. Drikakis, O. P. Iliev, and D. P. Vassileva. Acceleration of multigrid flow computations through dynamic adaptation of the smoothing procedure. Journal of Computational Physics, 165(2):566–591, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966336.

Du:2004:NSQ

[DJ04] Qiang Du and Lili Ju. Numerical simulations of the quantized vortices on a thin superconducting hollow sphere. Journal of Computational Physics, 201(2):511–530, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002402.

Degond:2005:STM

[DJM05] Pierre Degond, Shi Jin, and Luc Mieussens. A smooth transition model between kinetic and hydrodynamic equations. Journal of Computational Physics, 209(2):665–694, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001932.

Dubois:2005:MSS

[DJTT05] T. Dubois, F. Jauberteau, R. M. Temam, and J. Tribbia. Multilevel schemes for the shallow water equations. Journal of Computational Physics, 207(2):660–694, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910500046X.

Diez:2002:CTD

[DK02a] Javier A. Diez and L. Kondic. Computing three-dimensional thin film flows including contact lines. Journal of Computational Physics, 183(1):274–306, November 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971974.

Dubroca:2002:HMC

[DK02b] B. Dubroca and A. Klar. Half-moment closure for radiative transfer equations. Journal of Computational Physics, 180(2):584–596, August 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971068.

Deconinck:2006:CSL

[DK06] Bernard Deconinck and J. Nathan Kutz. Computing spectra of linear operators using the Floquet–Fourier–Hill method. Journal of Computational Physics, 219(1):296–321, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106001665.

Dumbser:2007:AHO

[DK07] Michael Dumbser and Martin Käser. Arbitrary high order non-oscillatory finite volume schemes on unstructured meshes for linear hyperbolic systems. Journal of Computational Physics, 221(2):693–723, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003123.

Dedner:2002:HDC

[DKK^{+}02] A. Dedner, F. Kemm, D. Kröner, C.-D. Munz, T. Schnitzer, and M. Wesenberg. Hyperbolic divergence cleaning for the MHD equations. Journal of Computational Physics, 175 (2):645–673, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910196961X.

DiGSigalotti:2003:SST

[DKS^{+}03] Leonardo Di G. Sigalotti, Jaime Klapp, Eloy Sira, Yasmin Meleán, and Anwar Hasmy. SPH simulations of timedependent Poiseuille flow at low Reynolds numbers. Journal of Computational Physics, 191(2):622–638, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003437.

Dedner:2001:TBC

[DKSW01] A. Dedner, D. Kröner, I. L. Sofronov, and M. Wesenberg. Transparent boundary conditions for MHD simulations in stratified atmospheres. Journal of Computational Physics, 171(2):448–478, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967798.

Dumbser:2007:QFN

[DKTT07] Michael Dumbser, Martin Käser, Vladimir A. Titarev, and Eleuterio F. Toro. Quadrature-free non-oscillatory finite volume schemes on unstructured meshes for nonlinear hyperbolic systems. Journal of Computational Physics, 226(1):204–243, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001520.

Deeba:2000:ASB

[DKX00] Elias Deeba, S. A. Khuri, and Shishen Xie. An algorithm for solving boundary value problems. Journal of Computational Physics, 159(2):125–138, April 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964520. See erratum [DKX01].

Deeba:2001:EAS

[DKX01] Elias Deeba, S. A. Khuri, and Shishen Xie. Erratum: “An Algorithm for Solving Boundary Value Problems”: Volume 159, Number 2 (2000), pages 125–138. Journal of Computational Physics, 170(1):448, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968317. See [DKX00].

Densmore:2003:VVR

[DL03a] Jeffery D. Densmore and Edward W. Larsen. Variational variance reduction for particle transport eigenvalue calculations using Monte Carlo adjoint simulation. Journal of Computational Physics, 192(2):387–405, December 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003991.

Dupont:2003:BFE

[DL03b] Todd F. Dupont and Yingjie Liu. Back and forth error compensation and correction methods for removing errors induced by uneven gradients of the level set function. Journal of Computational Physics, 190(1):311–324, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002766.

Densmore:2004:AED

[DL04] Jeffery D. Densmore and Edward W. Larsen. Asymptotic equilibrium diffusion analysis of time-dependent Monte Carlo methods for grey radiative transfer. Journal of Computational Physics, 199(1):175–204, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000816.

DiGSigalotti:2006:SCS

[DLD^{+}06] Leonardo Di G. Sigalotti, Hender López, Arnaldo Donoso, Eloy Sira, and Jaime Klapp. A shock-capturing SPH scheme based on adaptive kernel estimation. Journal of Computational Physics, 212(1):124–149, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500313X.

Diamessis:2008:ENV

[DLD08] P. J. Diamessis, Y. C. Lin, and J. A. Domaradzki. Effective numerical viscosity in spectral multidomain penalty method-based simulations of localized turbulence. Journal of Computational Physics, 227(17):8145–8164, September 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002891.

Marcos:2002:DDP

[dlFMBdlFM02] Carlos de la Fuente Marcos, Pierre Barge, and Raúl de la Fuente Marcos. Dust dynamics in protoplanetary disks: Parallel computing with PVM. Journal of Computational Physics, 176(2):276–294, March 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969785.

Dong:2004:SDL

[DLMK04] Suchuan Dong, Dong Liu, Martin R. Maxey, and George Em Karniadakis. Spectral distributed Lagrange multiplier method: algorithm and benchmark tests. Journal of Computational Physics, 195(2):695–717, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005795.

Dolean:2008:DDM

[DLP08] Victorita Dolean, Stéphane Lanteri, and Ronan Perrussel. A domain decomposition method for solving the threedimensional time-harmonic Maxwell equations discretized by discontinuous Galerkin methods. Journal of Computational Physics, 227(3):2044–2072, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004469.

Ducros:2000:HOF

[DLS^{+}00] F. Ducros, F. Laporte, T. Soulères, V. Guinot, P. Moinat, and B. Caruelle. High-order fluxes for conservative skewsymmetric-like schemes in structured meshes: Application to compressible flows. Journal of Computational Physics, 161(1):114–139, June 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964921.

DiGSigalotti:2009:ASM

[DLT09] Leonardo Di G. Sigalotti, Hender López, and Leonardo Trujillo. An adaptive SPH method for strong shocks. Journal of Computational Physics, 228(16):5888–5907, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002423.

Du:2004:PFA

[DLW04] Qiang Du, Chun Liu, and Xiaoqiang Wang. A phase field approach in the numerical study of the elastic bending energy for vesicle membranes. Journal of Computational Physics, 198(2):450–468, August 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000373.

Du:2006:SDV

[DLW06] Qiang Du, Chun Liu, and Xiaoqiang Wang. Simulating the deformation of vesicle membranes under elastic bending energy in three dimensions. Journal of Computational Physics, 212(2):757–777, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003566.

Dance:2003:ILE

[DM03] S. L. Dance and M. R. Maxey. Incorporation of lubrication effects into the force-coupling method for particulate two-phase flow. Journal of Computational Physics, 189(1):212–238, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002092.

Daffin:2005:EDF

[DMBS05] Frank Daffin, Michael Scott McKinley, Eugene D. Brooks III, and Abraham Szőke. An evaluation of the difference formulation for photon transport in a two level system. Journal of Computational Physics, 204(1):27–45, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400405X.

Darmofal:2000:EAB

[DMG00] David L. Darmofal, Pierre Moinier, and Michael B. Giles. Eigenmode analysis of boundary conditions for the onedimensional preconditioned Euler equations. Journal of Computational Physics, 160(1):369–384, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964726.

Dumbser:2004:MSA

[DMG04] Michael Dumbser, Jean-Marc Moschetta, and Jérémie Gressier. A matrix stability analysis of the carbuncle phenomenon. Journal of Computational Physics, 197(2):647–670, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006570.

DAvino:2007:NMS

[DMHP07] G. D’Avino, P. L. Maffettone, M. A. Hulsen, and G. W. M. Peters. A numerical method for simulating concentrated rigid particle suspensions in an elongational flow using a fixed grid. Journal of Computational Physics, 226(1):688–711, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002021.

Desjardins:2008:ACL

[DMP08] Olivier Desjardins, Vincent Moureau, and Heinz Pitsch. An accurate conservative level set/ghost fluid method for simulating turbulent atomization. Journal of Computational Physics, 227(18):8395–8416, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003112.

Deconinck:2009:DJH

[DMR09] T. Deconinck, S. Mahadevan, and L. L. Raja. Discretization of the Joule heating term for plasma discharge fluid models in unstructured meshes. Journal of Computational Physics, 228(12):4435–4443, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001351.

Dragojlovic:2006:EBM

[DND06] Zoran Dragojlovic, Farrokh Najmabadi, and Marcus Day. An embedded boundary method for viscous, conducting compressible flow. Journal of Computational Physics, 216(1):37– 51, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005413.

Desroziers:2008:SLP

[DNS08] S. Desroziers, F. Nataf, and R. Sentis. Simulation of laser propagation in a plasma with a frequency wave equation. Journal of Computational Physics, 227(4):2610–2625, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004846.

deNiet:2007:TSB

[dNWvSD07] Arie de Niet, Fred Wubs, Arjen Terwisscha van Scheltinga, and Henk A. Dijkstra. A tailored solver for bifurcation analysis of ocean-climate models. Journal of Computational Physics, 227(1):654–679, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999107003518. See corrigendum [dTWD09].

Dohnal:2009:PML

[Doh09] Tomás Dohnal. Perfectly matched layers for coupled nonlinear Schrödinger equations with mixed derivatives. Journal of Computational Physics, 228(23):8752–8765, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004689.

Domenichini:2008:CDF

[Dom08] Federico Domenichini. On the consistency of the direct forcing method in the fractional step solution of the Navier–Stokes equations. Journal of Computational Physics, 227(12):6372–6384, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001617.

Duponcheel:2008:TRE

[DOW08] M. Duponcheel, P. Orlandi, and G. Winckelmans. Timereversibility of the Euler equations as a benchmark for energy conserving schemes. Journal of Computational Physics, 227(19):8736–8752, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003422.

Dijkstra:2001:FIM

[DOWB01] Henk A. Dijkstra, Hakan Oksuzoğlu, Fred. W. Wubs, and Eugen F. F. Botta. A fully implicit model of the threedimensional thermohaline ocean circulation. Journal of Computational Physics, 173(2):685–715, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969086.

Derickson:2000:PSB

[DP00] Russell G. Derickson and Roger A. Pielke, Sr. A preliminary study of the Burgers equation with symbolic computation. Journal of Computational Physics, 162(1):219–244, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965331.

Deledicque:2007:ERS

[DP07] Vincent Deledicque and Miltiadis V. Papalexandris. An exact Riemann solver for compressible two-phase flow models containing non-conservative products. Journal of Computational Physics, 222(1):217–245, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003524.

Deledicque:2008:CAC

[DP08] Vincent Deledicque and Miltiadis V. Papalexandris. A conservative approximation to compressible two-phase flow models in the stiff mechanical relaxation limit. Journal of Computational Physics, 227(21):9241–9270, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700561X.

Desjardins:2009:SRI

[DP09] Olivier Desjardins and Heinz Pitsch. A spectrally refined interface approach for simulating multiphase flows. Journal of Computational Physics, 228(5):1658–1677, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005883.

Dendy:2002:GPF

[DPCV02] E. D. Dendy, N. T. Padial-Collins, and W. B. VanderHeyden. A general-purpose finite-volume advection scheme for continuous and discontinuous fields on unstructured grids. Journal of Computational Physics, 180(2):559–583, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971056.

Deng:2000:ALB

[DPR00] Yuefan Deng, Ronald F. Peierls, and Carlos Rivera. An adaptive load balancing method for parallel molecular dynamics simulations. Journal of Computational Physics, 161(1):250–263, June 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910096501X.

DePalma:2005:SOA

[DPRN05] P. De Palma, G. Pascazio, G. Rossiello, and M. Napolitano. A second-order-accurate monotone implicit fluctuation splitting scheme for unsteady problems. Journal of Computational Physics, 208(1):1–33, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400484X.

DePalma:2006:RDS

[DPRN06] P. De Palma, G. Pascazio, D. T. Rubino, and M. Napolitano. Residual distribution schemes for advection and advectiondiffusion problems on quadrilateral cells. Journal of Computational Physics, 218(1):159–199, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000775.

Dolcetta:2001:UHM

[DPRS01] R. Capuzzo Dolcetta, N. Pucello, V. Rosato, and F. Saraceni. On the use of a heterogeneous MIMD–SIMD platform to simulate the dynamics of globular clusters with a central massive object. Journal of Computational Physics, 174(1):208–225, November 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969025.

Dreyer:2004:KFV

[DQ04] Wolfgang Dreyer and Shamsul Qamar. Kinetic flux-vector splitting schemes for the hyperbolic heat conduction. Journal of Computational Physics, 198(2):403–423, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400035X.

Do-Quang:2008:SFD

[DQA08] Minh Do-Quang and Gustav Amberg. Simulation of free dendritic crystal growth in a gravity environment. Journal of Computational Physics, 227(3):1772–1789, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004317.

Degerfeldt:2006:BTF

[DR06] D. Degerfeldt and T. Rylander. A brick-tetrahedron finiteelement interface with stable hybrid explicit-implicit timestepping for Maxwell’s equations. Journal of Computational Physics, 220(1):383–393, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002464.

Deparis:2009:RBM

[DR09a] Simone Deparis and Gianluigi Rozza. Reduced basis method for multi-parameter-dependent steady Navier–Stokes equations: Applications to natural convection in a cavity. Journal of Computational Physics, 228(12):4359–4378, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001284.

Duan:2009:HOQ

[DR09b] Ran Duan and Vladimir Rokhlin. High-order quadratures for the solution of scattering problems in two dimensions. Journal of Computational Physics, 228(6):2152–2174, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006220.

Driscoll:2002:CRK

[Dri02] Tobin A. Driscoll. A composite Runge–Kutta method for the spectral solution of semilinear PDEs. Journal of Computational Physics, 182(2):357–367, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971275.

Drikakis:2001:SVS

[DS01] Dimitris Drikakis and Piotr K. Smolarkiewicz. On spurious vortical structures. Journal of Computational Physics, 172 (1):309–325, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968251.

Dai:2005:ATM

[DS05a] Meizhong Dai and David P. Schmidt. Adaptive tetrahedral meshing in free-surface flow. Journal of Computational Physics, 208(1):228–252, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001075.

Diaz:2005:NMR

[DS05b] Rodolfo E. Diaz and Igor Scherbatko. A new multistack radiation boundary condition for FDTD based on self-teleportation of fields. Journal of Computational Physics, 203(1):176–190, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003237.

Ding:2006:SAA

[DS06a] H. Ding and C. Shu. A stencil adaptive algorithm for finite difference solution of incompressible viscous flows. Journal of Computational Physics, 214(1):397–420, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004511.

Dipankar:2006:SCS

[DS06b] A. Dipankar and T. K. Sengupta. Symmetrized compact scheme for receptivity study of 2D transitional channel flow. Journal of Computational Physics, 215(1):245–273, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004870.

Dyadechko:2008:RMM

[DS08] Vadim Dyadechko and Mikhail Shashkov. Reconstruction of multi-material interfaces from moment data. Journal of Computational Physics, 227(11):5361–5384, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005748.

Dipankar:2009:NPS

[DS09a] A. Dipankar and P. Sagaut. A new phase-screen method for electromagnetic wave propagation in turbulent flows using large-eddy simulation. Journal of Computational Physics, 228(20):7729–7741, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003933.

Dritschel:2009:SNI

[DS09b] D. G. Dritschel and R. K. Scott. On the simulation of nearly inviscid two-dimensional turbulence. Journal of Computational Physics, 228(8):2707–2711, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000266.

daSilva:2000:UAG

[dSAK00] L. F. Figueira da Silva, João L. F. Azevedo, and Heidi Korzenowski. Unstructured adaptive grid flow simulations of inert and reactive gas mixtures. Journal of Computational Physics, 160(2):522–540, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964702.

Drouvelis:2006:PIR

[DSB06] P. S. Drouvelis, P. Schmelcher, and P. Bastian. Parallel implementation of the recursive Green’s function method. Journal of Computational Physics, 215(2):741–756, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005139.

daSilva:2005:UTS

[dSHHM05] F. da Silva, S. Heuraux, S. Hacquin, and M. E. Manso. Unidirectional transparent signal injection in finite-difference time-domain electromagnetic codes — application to reflectometry simulations. Journal of Computational Physics, 203(2):467–492, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003882.

Debry:2003:SAN

[DSJ03] Edouard Debry, Bruno Sportisse, and Benjamin Jourdain. A stochastic approach for the numerical simulation of the general dynamics equation for aerosols. Journal of Computational Physics, 184(2):649–669, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000414.

dAquino:2005:GIL

[dSM05] Massimiliano d’Aquino, Claudio Serpico, and Giovanni Miano. Geometrical integration of Landau–Lifshitz–Gilbert equation based on the mid-point rule. Journal of Computational Physics, 209(2):730–753, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500197X.

Darwish:2009:CFV

[DSM09a] M. Darwish, I. Sraj, and F. Moukalled. A coupled finite volume solver for the solution of incompressible flows on unstructured grids. Journal of Computational Physics, 228(1):180–201, January 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004725.

Davoudabadi:2009:APH

[DSM09b] M. Davoudabadi, J. S. Shrimpton, and F. Mashayek. On accuracy and performance of high-order finite volume methods in local mean energy model of non-thermal plasmas. Journal of Computational Physics, 228(7):2468–2479, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006384.

dAquino:2009:NFN

[dSMF09] Massimiliano d’Aquino, Claudio Serpico, Giovanni Miano, and Carlo Forestiere. A novel formulation for the numerical computation of magnetization modes in complex micromagnetic systems. Journal of Computational Physics, 228(17):6130–6149, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002642.

deSousa:2004:FTF

[dSMN^{+}04] F. S. de Sousa, N. Mangiavacchi, L. G. Nonato, A. Castelo, M. F. Tomé, V. G. Ferreira, J. A. Cuminato, and S. Mc-Kee. A front-tracking/front-capturing method for the simulation of 3D multi-fluid flows with free surfaces. Journal of Computational Physics, 198(2):469–499, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104000488.

Domokos:2000:SRB

[DSS00] Gábor Domokos, Imre Szeberényi, and Paul H. Steen. Simultaneously resolved bifurcation diagrams: a novel global approach applied to liquid figures of equilibrium. Journal of Computational Physics, 159(1):38–57, March 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964192.

Ding:2007:DIM

[DSS07] Hang Ding, Peter D. M. Spelt, and Chang Shu. Diffuse interface model for incompressible two-phase flows with large density ratios. Journal of Computational Physics, 226(2):2078–2095, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002793.

Donev:2007:CFE

[DST07a] Aleksandar Donev, Frank H. Stillinger, and Salvatore Torquato. Calculating the free energy of nearly jammed hard-particle packings using molecular dynamics. Journal of Computational Physics, 225(1):509–527, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006036.

Dubroca:2007:CAC

[DST07b] Bruno Dubroca, Mohammed Seaïd, and Ioan Teleaga. A consistent approach for the coupling of radiation and hydrodynamics at low Mach number. Journal of Computational Physics, 225(1):1039–1065, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000186.

Dimakopoulos:2003:QET

[DT03] Yannis Dimakopoulos and John Tsamopoulos. A quasielliptic transformation for moving boundary problems with large anisotropic deformations. Journal of Computational Physics, 192(2):494–522, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004066.

Daru:2004:HOO

[DT04] V. Daru and C. Tenaud. High order one-step monotonicitypreserving schemes for unsteady compressible flow calculations. Journal of Computational Physics, 193(2):563–594, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004327.

deTullio:2007:IBM

[dTDI^{+}07] M. D. de Tullio, P. De Palma, G. Iaccarino, G. Pascazio, and M. Napolitano. An immersed boundary method for compressible flows using local grid refinement. Journal of Computational Physics, 225(2):2098–2117, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001222.

Desquesnes:2006:UHO

[DTMS06] G. Desquesnes, M. Terracol, E. Manoha, and P. Sagaut. On the use of a high order overlapping grid method for coupling in CFD/CAA. Journal of Computational Physics, 220(1):355–382, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002452.

Donev:2005:NLCa

[DTS05a] Aleksandar Donev, Salvatore Torquato, and Frank H. Stillinger. Neighbor list collision-driven molecular dynamics simulation for nonspherical hard particles. I. Algorithmic details. Journal of Computational Physics, 202(2):737–764, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104003146.

Donev:2005:NLCb

[DTS05b] Aleksandar Donev, Salvatore Torquato, and Frank H. Stillinger. Neighbor list collision-driven molecular dynamics simulation for nonspherical hard particles. II. Applications to ellipses and ellipsoids. Journal of Computational Physics, 202(2):765–793, January 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003948.

Donev:2004:LPA

[DTSC04] Aleksandar Donev, Salvatore Torquato, Frank H. Stillinger, and Robert Connelly. A linear programming algorithm to test for jamming in hard-sphere packings. Journal of Computational Physics, 197(1):139–166, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006235.

denToom:2009:CTS

[dTWD09] Matthijs den Toom, Fred W. Wubs, and Henk A. Dijkstra. Corrigendum to “A tailored solver for bifurcation analysis of ocean-climate models” [J. Comput. Phys. 227 (2007) 654–679]. Journal of Computational Physics, 228(13):4962–4964, July 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001855. See [dNWvSD07].

Densmore:2007:HTD

[DUEB07] Jeffery D. Densmore, Todd J. Urbatsch, Thomas M. Evans, and Michael W. Buksas. A hybrid transport-diffusion method for Monte Carlo radiative-transfer simulations. Journal of Computational Physics, 222(2):485–503, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003639.

Durran:2000:WPQ

[Dur00] Dale R. Durran. Wave propagation in quadratic-finite-element approximations to hyperbolic equations. Journal of Computational Physics, 159(2):448–455, April 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964532.

Duran:2008:TBE

[Dur08] A. Durán. Time behaviour of the error when simulating finiteband periodic waves. The case of the KdV equation. Journal of Computational Physics, 227(3):2130–2153, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700472X.

Dedner:2002:AHO

[DV02] A. Dedner and P. Vollmöller. An adaptive higher order method for solving the radiation transport equation on unstructured grids. Journal of Computational Physics, 178(2):263–289, May 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970014.

daVeiga:2009:MFD

[dVGLM09] L. Beirão da Veiga, V. Gyrya, K. Lipnikov, and G. Manzini. Mimetic finite difference method for the Stokes problem on polygonal meshes. Journal of Computational Physics, 228(19):7215–7232, October 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900343X.

DeWilde:2005:SSA

[DVHM05] Juray De Wilde, Jan Vierendeels, Geraldine J. Heynderickx, and Guy B. Marin. Simultaneous solution algorithms for Eulerian–Eulerian gas-solid flow models: Stability analysis and convergence behaviour of a point and a plane solver. Journal of Computational Physics, 207(1):309–353, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000318.

Dai:2000:NSR

[DW00] W. Wenlong Dai and Paul R. Woodward. Numerical simulations for radiation hydrodynamics: II. Transport limit. Journal of Computational Physics, 157(1):199–233, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963714.

Di:2009:PSS

[DW09] Yana Di and Xiao-Ping Wang. Precursor simulations in spreading using a multi-mesh adaptive finite element method. Journal of Computational Physics, 228(5):1380–1390, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005603.

Dimits:2009:UAN

[DWC^{+}09] Andris M. Dimits, Chiaming Wang, Russel Caflisch, Bruce I. Cohen, and Yanghong Huang. Understanding the accuracy of Nanbu’s numerical Coulomb collision operator. Journal of Computational Physics, 228(13):4881–4892, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001715.

Dwight:2008:HPE

[Dwi08] Richard P. Dwight. Heuristic a posteriori estimation of error due to dissipation in finite volume schemes and application to mesh adaptation. Journal of Computational Physics, 227(5):2845–2863, February 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700513X.

deWijn:2007:NAR

[dWKL07] Astrid S. de Wijn, Stephan Kümmel, and Manfred Lein. Numerical aspects of real-space approaches to strong-field electron dynamics. Journal of Computational Physics, 226 (1):89–103, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107001441.

Densmore:2009:SAI

[DWLM09] Jeffery D. Densmore, James S. Warsa, Robert B. Lowrie, and Jim E. Morel. Stability analysis of implicit time discretizations for the Compton-scattering Fokker–Planck equation. Journal of Computational Physics, 228(16):5933–5960, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002459.

Dyshlovenko:2001:AME

[Dys01] Pavel Dyshlovenko. Adaptive mesh enrichment for the Poisson–Boltzmann equation. Journal of Computational Physics, 172(1):198–208, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968202.

Deng:2000:DHO

[DZ00] Xiaogang Deng and Hanxin Zhang. Developing high-order weighted compact nonlinear schemes. Journal of Computational Physics, 165(1):22–44, November 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096594X.

Dohlus:2009:ETT

[DZ09a] Martin Dohlus and Igor Zagorodnov. Explicit T/TM scheme for particle beam simulations. Journal of Computational Physics, 228(8):2822–2833, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006736.

Dumbser:2009:VHO

[DZ09b] Michael Dumbser and Olindo Zanotti. Very high order P_{N}P_{M} schemes on unstructured meshes for the resistive relativistic MHD equations. Journal of Computational Physics, 228(18):6991–7006, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003350.

Epstein:2001:AED

[EAY01] B. Epstein, A. Averbuch, and I. Yavneh. An accurate ENO driven multigrid method applied to 3D turbulent transonic flows. Journal of Computational Physics, 168(2):316–338, April 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101966987.

Elkina:2006:NCU

[EB06] N. V. Elkina and J. Büchner. A new conservative unsplit method for the solution of the Vlasov equation. Journal of Computational Physics, 213(2):862–875, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004201.

Eldredge:2002:VPM

[ECL02] Jeff D. Eldredge, Tim Colonius, and Anthony Leonard. A vortex particle method for two-dimensional compressible flow. Journal of Computational Physics, 179(2):371–399, July 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970609.

Evans:2007:MCR

[ED07] T. M. Evans and J. D. Densmore. Methods for coupling radiation, ion, and electron energies in grey implicit Monte Carlo. Journal of Computational Physics, 225(2):1695–1720, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000800.

Edwards:2000:MFS

[Edw00] Michael G. Edwards. M-matrix flux splitting for general full tensor discretization operators on structured and unstructured grids. Journal of Computational Physics, 160(1):1–28, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964180.

Edwards:2006:DWC

[Edw06] Michael G. Edwards. The dominant wave-capturing flux: a finite-volume scheme without decomposition for systems of hyperbolic conservation laws. Journal of Computational Physics, 218(1):275–294, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000817.

Eilks:2008:NSD

[EE08] C. Eilks and C. M. Elliott. Numerical simulation of dealloying by surface dissolution via the evolving surface finite element method. Journal of Computational Physics, 227(23):9727–9741, December 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004038.

Elsey:2009:DGM

[EES09] Matt Elsey, Selim Esedogl̄u, and Peter Smereka. Diffusion generated motion for grain growth in two and three dimensions. Journal of Computational Physics, 228(21):8015–8033, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004082.

Evje:2002:HFS

[EF02] Steinar Evje and Kjell K. Fjelde. Hybrid flux-splitting schemes for a two-phase flow model. Journal of Computational Physics, 175(2):674–701, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969621.

Evje:2003:HFS

[EF03] Steinar Evje and Tore Flåtten. Hybrid flux-splitting schemes for a common two-fluid model. Journal of Computational Physics, 192(1):175–210, November 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003711.

Enright:2002:HPL

[EFFM02] Douglas Enright, Ronald Fedkiw, Joel Ferziger, and Ian Mitchell. A hybrid particle level set method for improved interface capturing. Journal of Computational Physics, 183(1):83–116, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971664.

El-gamel:2007:CBS

[Eg07] Mohamed El-gamel. A comparison between the Sinc– Galerkin and the modified decomposition methods for solving two-point boundary-value problems. Journal of Computational Physics, 223(1):369–383, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004414.

Ergul:2008:NES

[EG08] Özgür Ergül and Levent Gürel. Novel electromagnetic surface integral equations for highly accurate computations of dielectric bodies with arbitrarily low contrasts. Journal of Computational Physics, 227(23):9898–9912, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004221.

Efendiev:2006:AMF

[EGHE06] Y. Efendiev, V. Ginting, T. Hou, and R. Ewing. Accurate multiscale finite element methods for two-phase flow simulations. Journal of Computational Physics, 220(1):155–174, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002269.

Etienne:2009:PGC

[ÉGP09] S. Étienne, A. Garon, and D. Pelletier. Perspective on the geometric conservation law and finite element methods for ALE simulations of incompressible flow. Journal of Computational Physics, 228(7):2313–2333, April 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006165.

E:2002:DAC

[EH02] Weinan E. and Zhongyi Huang. A dynamic atomisticcontinuum method for the simulation of crystalline materials. Journal of Computational Physics, 182(1):234–261, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971640.

Ekici:2008:CFH

[EHD08] Kivanc Ekici, Kenneth C. Hall, and Earl H. Dowell. Computationally fast harmonic balance methods for unsteady aerodynamic predictions of helicopter rotors. Journal of Computational Physics, 227(12):6206–6225, June 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800140X.

Elman:2008:TCP

[EHS^{+}08] Howard Elman, V. E. Howle, John Shadid, Robert Shuttleworth, and Ray Tuminaro. A taxonomy and comparison of parallel block multi-level preconditioners for the incompressible Navier–Stokes equations. Journal of Computational Physics, 227(3):1790–1808, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004330.

Elman:2003:PBM

[EHST03] Howard C. Elman, Victoria E. Howle, John N. Shadid, and Ray S. Tuminaro. A parallel block multi-level preconditioner for the 3D incompressible Navier–Stokes equations. Journal of Computational Physics, 187(2):504–523, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001219.

Engsig-Karup:2009:EFO

[EKBL09] A. P. Engsig-Karup, H. B. Bingham, and O. Lindberg. An efficient flexible-order model for 3D nonlinear water waves. Journal of Computational Physics, 228(6):2100–2118, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006190.

Eck:2002:TSM

[EKK02] Christof Eck, Peter Knabner, and Sergey Korotov. A twoscale method for the computation of solid-liquid phase transitions with dendritic microstructure. Journal of Computational Physics, 178(1):58–80, May 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297018X.

Evans:2006:DDA

[EKP06] Katherine J. Evans, D. A. Knoll, and Michael Pernice. Development of a 2-D algorithm to simulate convection and phase transition efficiently. Journal of Computational Physics, 219(1):404–417, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001690.

Evans:2007:EAE

[EKP07] Katherine J. Evans, D. A. Knoll, and Michael Pernice. Enhanced algorithm efficiency for phase change convection using a multigrid preconditioner with a SIMPLE smoother. Journal of Computational Physics, 223(1):121–126, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106004281.

Eldredge:2002:GDT

[ELC02] Jeff D. Eldredge, Anthony Leonard, and Tim Colonius. A general deterministic treatment of derivatives in particle methods. Journal of Computational Physics, 180(2):686–709, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971123.

Eldredge:2007:NSF

[Eld07] Jeff D. Eldredge. Numerical simulation of the fluid dynamics of 2D rigid body motion with the vortex particle method. Journal of Computational Physics, 221(2):626–648, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003093.

Eldredge:2008:DCF

[Eld08a] Jeff D. Eldredge. Dynamically coupled fluid-body interactions in vorticity-based numerical simulations. Journal of Computational Physics, 227(21):9170–9194, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001952.

Elsen:2008:LCF

[ELD08b] Erich Elsen, Patrick LeGresley, and Eric Darve. Large calculation of the flow over a hypersonic vehicle using a GPU. Journal of Computational Physics, 227(24):10148–10161, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004476.

Eliasson:2002:OBC

[Eli02] Bengt Eliasson. Outflow boundary conditions for the Fourier transformed two-dimensional Vlasov equation. Journal of Computational Physics, 181(1):98–125, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971214.

Eliasson:2003:NMT

[Eli03] Bengt Eliasson. Numerical modelling of the two-dimensional Fourier transformed Vlasov–Maxwell system. Journal of Computational Physics, 190(2):501–522, September 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300295X.

Eliasson:2007:OBC

[Eli07] B. Eliasson. Outflow boundary conditions for the Fourier transformed three-dimensional Vlasov–Maxwell system. Journal of Computational Physics, 225(2):1508–1532, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000666.

E:2007:NSS

[ELVE07] Weinan E., Di Liu, and Eric Vanden-Eijnden. Nested stochastic simulation algorithms for chemical kinetic systems with multiple time scales. Journal of Computational Physics, 221(1):158–180, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002804.

Estep:2001:ASL

[ELW01] Donald J. Estep, Sjoerd M. Verduyn Lunel, and Roy D. Williams. Analysis of shear layers in a fluid with temperaturedependent viscosity. Journal of Computational Physics, 173 (1):17–60, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968378.

Ewing:2004:ABS

[ELW04] Richard E. Ewing, Jiangguo Liu, and Hong Wang. Adaptive biorthogonal spline schemes for advection-reaction equations. Journal of Computational Physics, 193(1):21–39, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004145.

Estep:2006:FRM

[EN06] D. Estep and D. Neckels. Fast and reliable methods for determining the evolution of uncertain parameters in differential equations. Journal of Computational Physics, 213(2):530–556, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003918.

Eberhard:2008:NUE

[EPW08] Jens P. Eberhard, Dan Popovíc, and Gabriel Wittum. Numerical upscaling for the eddy-current model with stochastic magnetic materials. Journal of Computational Physics, 227(8):4244–4259, April 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005724.

Engquist:2002:HFW

[ERT02] Björn Engquist, Olof Runborg, and Anna-Karin Tornberg. High-frequency wave propagation by the segment projection method. Journal of Computational Physics, 178(2):373–390, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970336.

E:2009:GSD

[ERVE09] Weinan E., Weiqing Ren, and Eric Vanden-Eijnden. A general strategy for designing seamless multiscale methods. Journal of Computational Physics, 228(15):5437–5453, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002125.

Elliott:2003:CBG

[ES03a] C. M. Elliott and V. Styles. Computations of bidirectional grain boundary dynamics in thin metallic films. Journal of Computational Physics, 187(2):524–543, May 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001220.

Ewert:2003:APE

[ES03b] R. Ewert and W. Schröder. Acoustic perturbation equations based on flow decomposition via source filtering. Journal of Computational Physics, 188(2):365–398, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001682.

Eskilsson:2006:SHD

[ES06] Claes Eskilsson and Spencer J. Sherwin. Spectral/hp discontinuous Galerkin methods for modelling 2D Boussinesq equations. Journal of Computational Physics, 212(2):566–589, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003463.

Engle:2005:MED

[ESD05] Robert D. Engle, Robert D. Skeel, and Matthew Drees. Monitoring energy drift with shadow Hamiltonians. Journal of Computational Physics, 206(2):432–452, July 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005212.

Ellero:2007:ISP

[ESE07] Marco Ellero, Mar Serrano, and Pep Español. Incompressible smoothed particle hydrodynamics. Journal of Computational Physics, 226(2):1731–1752, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002598.

Esedoglu:2006:TDP

[ET06] Selim Esedoğlu and Yen-Hsi Richard Tsai. Threshold dynamics for the piecewise constant Mumford–Shah functional.

Journal of Computational Physics, 211(1):367–384, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002792.

Engquist:2005:DDD

[ETT05] Björn Engquist, Anna-Karin Tornberg, and Richard Tsai. Discretization of Dirac delta functions in level set methods. Journal of Computational Physics, 207(1):28–51, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910500015X.

Evans:2003:RMC

[EULM03] T. M. Evans, T. J. Urbatsch, H. Lichtenstein, and J. E. Morel. A residual Monte Carlo method for discrete thermal radiative diffusion. Journal of Computational Physics, 189(2):539–556, August 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910300233X.

Elmer:2003:APF

[EV03] Christopher E. Elmer and Erik S. Van Vleck. Anisotropy, propagation failure, and wave speedup in traveling waves of discretizations of a Nagumo PDE. Journal of Computational Physics, 185(2):562–582, March 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000044.

Edwards:2008:QPF

[EZ08a] Michael G. Edwards and Hongwen Zheng. A quasi-positive family of continuous darcy-flux finite-volume schemes with full pressure support. Journal of Computational Physics, 227(22):9333–9364, November 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002908.

Ehrhardt:2008:EAB

[EZ08b] Matthias Ehrhardt and Chunxiong Zheng. Exact artificial boundary conditions for problems with periodic structures. Journal of Computational Physics, 227(14):6877–6894, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002015.

Fan:2008:EFT

[Fan08] Ping Fan. An efficient front-tracking method for fully nonlinear interfacial waves. Journal of Computational Physics, 227(15):7346–7367, July 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002465.

Fasso:2003:CSA

[Fas03] Francesco Fassò. Comparison of splitting algorithms for the rigid body. Journal of Computational Physics, 189(2): 527–538, August 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103002328.

Ferdowsi:2008:SOA

[FB08] Poorya A. Ferdowsi and Markus Bussmann. Second-order accurate normals from height functions. Journal of Computational Physics, 227(22):9293–9302, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003987.

Figua:2000:IFM

[FBFF00] H. Figua, F. Bouchut, M. R. Feix, and E. Fijalkow. Instability of the filtering method for Vlasov’s equation. Journal of Computational Physics, 159(2):440–447, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964234.

Fournier:2005:FSE

[FBHV05] Alexandre Fournier, Hans-Peter Bunge, Rainer Hollerbach, and Jean-Pierre Vilotte. A Fourier-spectral element algorithm for thermal convection in rotating axisymmetric containers. Journal of Computational Physics, 204(2):462–489, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004243.

Frohn:2002:DHR

[FCB02] Lise M. Frohn, Jesper H. Christensen, and Jørgen Brandt. Development of a high-resolution nested air pollution model: The numerical approach. Journal of Computational Physics, 179(1):68–94, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970361.

Francois:2006:BFA

[FCD^{+}06] Marianne M. François, Sharen J. Cummins, Edward D. Dendy, Douglas B. Kothe, James M. Sicilian, and Matthew W. Williams. A balanced-force algorithm for continuous and sharp interfacial surface tension models within a volume tracking framework. Journal of Computational Physics, 213(1):141–173, March 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003748.

Fructus:2005:EMT

[FCGK05] Dorian Fructus, Didier Clamond, John Grue, and Øyvind Kristiansen. An efficient model for three-dimensional surface wave simulations: Part I: Free space problems. Journal of Computational Physics, 205(2):665–685, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004863.

Fan:2008:FVG

[FCJ08a] Kai Fan, Wei Cai, and Xia Ji. A full vectorial generalized discontinuous Galerkin beam propagation method (GDG–BPM) for nonsmooth electromagnetic fields in waveguides. Journal of Computational Physics, 227(15):7178–7191, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001812.

Fan:2008:GDG

[FCJ08b] Kai Fan, Wei Cai, and Xia Ji. A generalized discontinuous Galerkin (GDG) method for Schrödinger equations with nonsmooth solutions. Journal of Computational Physics, 227(4):2387–2410, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004706.

Forbes:2007:CUP

[FCT07] Lawrence K. Forbes, Michael J. Chen, and Claire E. Trenham. Computing unstable periodic waves at the interface of two inviscid fluids in uniform vertical flow. Journal of Computational Physics, 221(1):269–287, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002889.

Fietier:2003:TDA

[FD03] Nicolas Fiétier and Michel O. Deville. Time-dependent algorithms for the simulation of viscoelastic flows with spectral element methods: applications and stability. Journal of Computational Physics, 186(1):93–121, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000135.

Fidkowski:2007:TCC

[FD07] Krzysztof J. Fidkowski and David L. Darmofal. A triangular cut-cell adaptive method for high-order discretizations of the compressible Navier–Stokes equations. Journal of Computational Physics, 225(2):1653–1672, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000757.

Fong:2009:BBF

[FD09a] William Fong and Eric Darve. The black-box fast multipole method. Journal of Computational Physics, 228(23): 8712–8725, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109004665.

Fontane:2009:HAN

[FD09b] Jérôme Fontane and David G. Dritschel. The Hyper-CASL algorithm: a new approach to the numerical simulation of geophysical flows. Journal of Computational Physics, 228(17):6411–6425, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002848.

Fauconnier:2007:DPA

[FDD07] Dieter Fauconnier, Chris De Langhe, and Erik Dick. The dynamic procedure for accuracy improvement of numerical discretizations in fluid mechanics. Journal of Computational Physics, 224(2):1095–1123, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005675.

Fauconnier:2009:CEI

[FDD09a] Dieter Fauconnier, Chris De Langhe, and Erik Dick. Construction of explicit and implicit dynamic finite difference schemes and application to the large-eddy simulation of the Taylor–Green vortex. Journal of Computational Physics, 228(21):8053–8084, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004124.

Fauconnier:2009:FDF

[FDD09b] Dieter Fauconnier, Chris De Langhe, and Erik Dick. A family of dynamic finite difference schemes for large-eddy simulation. Journal of Computational Physics, 228(6):1830–1861, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005974.

Frank:2006:PME

[FDK06] Martin Frank, Bruno Dubroca, and Axel Klar. Partial moment entropy approximation to radiative heat transfer. Journal of Computational Physics, 218(1):1–18, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600057X.

Fong:2008:SAV

[FDL08] William Fong, Eric Darve, and Adrian Lew. Stability of asynchronous variational integrators. Journal of Computational Physics, 227(18):8367–8394, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003069.

Ferrante:2004:RMG

[FE04] A. Ferrante and S. E. Elghobashi. A robust method for generating inflow conditions for direct simulations of spatiallydeveloping turbulent boundary layers. Journal of Computational Physics, 198(1):372–387, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000348.

Fedkiw:2002:CEF

[Fed02] Ronald P. Fedkiw. Coupling an Eulerian fluid calculation to a Lagrangian solid calculation with the ghost fluid method. Journal of Computational Physics, 175(1):200–224, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969359.

Fedele:2005:FPM

[FEL^{+}05] Francesco Fedele, Margaret J. Eppstein, Jeffrey P. Laible, Anuradha Godavarty, and Eva M. Sevick-Muraca. Fluorescence photon migration by the boundary element method. Journal of Computational Physics, 210(1):109–132, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002007.

Feng:2006:DDN

[Fen06] Y. T. Feng. On the discrete dynamic nature of the conjugate gradient method. Journal of Computational Physics, 211(1):91–98, January 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002639.

Falcone:2002:SLS

[FF02] M. Falcone and R. Ferretti. Semi-Lagrangian schemes for Hamilton–Jacobi equations, discrete representation formulae and Godunov methods. Journal of Computational Physics, 175(2):559–575, January 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969542.

Flyer:2003:ANR

[FF03] Natasha Flyer and Bengt Fornberg. Accurate numerical resolution of transients in initial-boundary value problems for the heat equation. Journal of Computational Physics, 184(2):526–539, January 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000347.

Fureby:2002:LES

[FG02] Christer Fureby and Fernando F. Grinstein. Large eddy simulation of high-Reynolds-number free and wall-bounded flows. Journal of Computational Physics, 181(1):68–97, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971196.

Farnell:2004:MCS

[FG04] L. Farnell and W. G. Gibson. Monte Carlo simulation of diffusion in a spatially nonhomogeneous medium: correction to the Gaussian steplength. Journal of Computational Physics, 198(1):65–79, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000191.

Farnell:2005:MCS

[FG05] L. Farnell and W. G. Gibson. Monte Carlo simulation of diffusion in a spatially nonhomogeneous medium: a biased random walk on an asymmetrical lattice. Journal of Computational Physics, 208(1):253–265, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001087.

Fuchs:2006:IEM

[FG06] V. Fuchs and J. P. Gunn. On the integration of equations of motion for particle-in-cell codes. Journal of Computational Physics, 214(1):299–315, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004365.

Fructus:2007:EMN

[FG07] Dorian Fructus and John Grue. An explicit method for the nonlinear interaction between water waves and variable and moving bottom topography. Journal of Computational Physics, 222(2):720–739, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003925.

Farhat:2001:DGC

[FGG01] Charbel Farhat, Philippe Geuzaine, and Céline Grandmont. The discrete geometric conservation law and the nonlinear stability of ALE schemes for the solution of flow problems on moving grids. Journal of Computational Physics, 174(2):669–694, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969323.

Frontera:2000:EAF

[FGOV00] Carlos Frontera, Jürgen Goicoechea, Jordi Ortín, and Eduard Vives. Efficient algorithm for finding ground-states in the random field Ising model with an external field. Journal of Computational Physics, 160(1):117–125, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964374.

Fournier:2008:NOB

[FGP08] G. Fournier, F. Golanski, and A. Pollard. A novel outflow boundary condition for incompressible laminar wall-bounded flows. Journal of Computational Physics, 227(15):7077–7082, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001794.

Favrie:2009:SFD

[FGS09] N. Favrie, S. L. Gavrilyuk, and R. Saurel. Solid-fluid diffuse interface model in cases of extreme deformations. Journal of Computational Physics, 228(16):6037–6077, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002654.

Filippova:2000:ALB

[FH00a] Olga Filippova and Dieter Hänel. Acceleration of lattice-BGK schemes with grid refinement. Journal of Computational Physics, 165(2):407–427, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966178.

Filippova:2000:NLB

[FH00b] Olga Filippova and Dieter Hänel. A novel lattice BGK approach for low Mach number combustion. Journal of Computational Physics, 158(2):139–160, March 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964057.

Franz:2002:OAS

[FH02] Astrid Franz and Karl Heinz Hoffmann. Optimal annealing schedules for a modified Tsallis statistics. Journal of Computational Physics, 176(1):196–204, February 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196975X.

Fujii:2003:IWC

[FH03] Masafumi Fujii and Wolfgang J. R. Hoefer. Interpolating wavelet collocation method of time dependent Maxwell’s equations: characterization of electrically large optical waveguide discontinuities. Journal of Computational Physics, 186(2): 666–689, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103000913.

Fixel:2007:CSS

[FH07] D. A. Fixel and W. N. G. Hitchon. Convective scheme solution of the Boltzmann transport equation for nanoscale semiconductor devices. Journal of Computational Physics, 227(2):1387–1410, December 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003993.

Feng:2009:EPQ

[FHD^{+}09] B. Feng, C. Huang, V. Decyk, W. B. Mori, P. Muggli, and T. Katsouleas. Enhancing parallel quasi-static particle-in-cell simulations with a pipelining algorithm. Journal of Computational Physics, 228(15):5340–5348, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900206X.

Farhi:2009:VEC

[FHJK09] E. Farhi, V. Hugouvieux, M. R. Johnson, and W. Kob. Virtual experiments: Combining realistic neutron scattering instrument and sample simulations. Journal of Computational Physics, 228(14):5251–5261, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002009.

Flad:2005:WAQ

[FHLK05] Heinz-Jürgen Flad, Wolfgang Hackbusch, Hongjun Luo, and Dietmar Kolb. Wavelet approach to quasi two-dimensional extended many-particle systems. I. Supercell Hartree–Fock method. Journal of Computational Physics, 205(2):540–566, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004802.

Feng:2008:DTE

[FHLO08] Y. T. Feng, K. Han, C. F. Li, and D. R. J. Owen. Discrete thermal element modelling of heat conduction in particle systems: Basic formulations. Journal of Computational Physics, 227(10):5072–5089, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000594.

Fattebert:2007:FEA

[FHW07] J.-L. Fattebert, R. D. Hornung, and A. M. Wissink. Finite element approach for density functional theory calculations on locally-refined meshes. Journal of Computational Physics, 223(2):759–773, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004724.

Ferraro:2009:CTF

[FJ09] N. M. Ferraro and S. C. Jardin. Calculations of two-fluid magnetohydrodynamic axisymmetric steady-states. Journal of Computational Physics, 228(20):7742–7770, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003969.

Fukagata:2002:HEC

[FK02] Koji Fukagata and Nobuhide Kasagi. Highly energyconservative finite difference method for the cylindrical coordinate system. Journal of Computational Physics, 181(2):478–498, September 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297138X.

Feng:2006:FEM

[FK06] Feng Feng and William S. Klug. Finite element modeling of lipid bilayer membranes. Journal of Computational Physics, 220(1):394–408, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002476.

Feistauer:2007:RDG

[FK07a] M. Feistauer and V. Kucera. On a robust discontinuous Galerkin technique for the solution of compressible flow. Journal of Computational Physics, 224(1):208–221, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000411.

Finkelstein:2007:FDT

[FK07b] Bezalel Finkelstein and Raphael Kastner. Finite difference time domain dispersion reduction schemes. Journal of Computational Physics, 221(1):422–438, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002932.

Fedosov:2009:TDI

[FK09a] Dmitry A. Fedosov and George Em Karniadakis. Tripledecker: Interfacing atomistic-mesoscopic-continuum flow regimes. Journal of Computational Physics, 228(4):1157– 1171, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005421.

Finkelstein:2009:SOA

[FK09b] B. Finkelstein and R. Kastner. The spectral order of accuracy: a new unified tool in the design methodology of excitation-adaptive wave equation FDTD schemes. Journal of Computational Physics, 228(24):8958–8984, December 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004744.

Furuichi:2008:TDE

[FKK08] Mikito Furuichi, Masanori Kameyama, and Akira Kageyama. Three-dimensional Eulerian method for large deformation of viscoelastic fluid: Toward plate-mantle simulation. Journal of Computational Physics, 227(10):4977–4997, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000557.

Frank:2007:TDS

[FKLY07] Martin Frank, Axel Klar, Edward W. Larsen, and Shugo Yasuda. Time-dependent simplified P_{N} approximation to the equations of radiative transfer. Journal of Computational Physics, 226(2):2289–2305, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003154.

Feng:2008:AAR

[FKV08] Y. Feng, H. C. Kim, and J. P. Verboncoeur. Algorithms for accurate relativistic particle injection. Journal of Computational Physics, 227(3):1663–1675, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004238.

Ferm:2003:AGA

[FL03] Lars Ferm and Per Lötstedt. Anisotropic grid adaptation for Navier–Stokes’ equations. Journal of Computational Physics, 190(1):22–41, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300250X.

Felten:2006:KEC

[FL06] Frédéric N. Felten and Thomas S. Lund. Kinetic energy conservation issues associated with the collocated mesh scheme for incompressible flow. Journal of Computational Physics, 215(2):465–484, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004997.

Fiorina:2007:AND

[FL07] B. Fiorina and S. K. Lele. An artificial nonlinear diffusivity method for supersonic reacting flows with shocks. Journal of Computational Physics, 222(1):246–264, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003536.

Feng:2009:OSM

[FL09] B.-F. Feng and Y. Liu. An operator splitting method for the Degasperis–Procesi equation. Journal of Computational Physics, 228(20):7805–7820, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004112.

Favennec:2003:IHP

[FLB03] Y. Favennec, V. Labbé, and F. Bay. Induction heating processes optimization a general optimal control approach. Journal of Computational Physics, 187(1):68–94, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000810.

Fedele:2003:CCA

[FLE03] F. Fedele, J. P. Laible, and M. J. Eppstein. Coupled complex adjoint sensitivities for frequency-domain fluorescence tomography: theory and vectorized implementation. Journal of Computational Physics, 187(2):597–619, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001505.

Fedioun:2001:RNE

[FLG01] Ivan Fedioun, Nicolas Lardjane, and Iskender Gökalp. Revisiting numerical errors in direct and large eddy simulations of turbulence: Physical and spectral spaces analysis. Journal of Computational Physics, 174(2):816–851, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969396.

Fox:2008:NSS

[FLM08] R. O. Fox, F. Laurent, and M. Massot. Numerical simulation of spray coalescence in an Eulerian framework: Direct quadrature method of moments and multi-fluid method. Journal of Computational Physics, 227(6):3058–3088, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004664.

Fomel:2009:FSM

[FLZ09] Sergey Fomel, Songting Luo, and Hongkai Zhao. Fast sweeping method for the factored eikonal equation. Journal of Computational Physics, 228(17):6440–6455, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002873.

Feng:2004:IBL

[FM04] Zhi-Gang Feng and Efstathios E. Michaelides. The immersed boundary-lattice Boltzmann method for solving fluid-particles interaction problems. Journal of Computational Physics, 195(2):602–628, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005758.

Feng:2005:PDF

[FM05] Zhi-Gang Feng and Efstathios E. Michaelides. Proteus: a direct forcing method in the simulations of particulate flows. Journal of Computational Physics, 202(1):20–51, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002669.

Fujimoto:2006:EFP

[FM06] Keizo Fujimoto and Shinobu Machida. Electromagnetic full particle code with adaptive mesh refinement technique: Application to the current sheet evolution. Journal of Computational Physics, 214(2):550–566, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004614.

Fatenejad:2008:EKD

[FM08] Milad Fatenejad and Gregory A. Moses. Extension of Kershaw diffusion scheme to hexahedral meshes. Journal of Computational Physics, 227(4):2187–2194, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004810.

Frantziskonis:2009:TPM

[FMD^{+}09] G. Frantziskonis, K. Muralidharan, P. Deymier, S. Simunovic, P. Nukala, and S. Pannala. Time-parallel multiscale/ multiphysics framework. Journal of Computational Physics, 228(21):8085–8092, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004161.

Fedkiw:2000:SDS

[FMO00] Ronald P. Fedkiw, Barry Merriman, and Stanley Osher. Simplified discretization of systems of hyperbolic conservation laws containing advection equations. Journal of Computational Physics, 157(1):302–326, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963799.

Fuchs:2009:SBF

[FMR09] F. G. Fuchs, S. Mishra, and N. H. Risebro. Splitting based finite volume schemes for ideal MHD equations. Journal of Computational Physics, 228(3):641–660, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004993.

Fernandez-Nieto:2008:NSH

[FNBB^{+}08] E. D. Fernández-Nieto, F. Bouchut, D. Bresch, M. J. Castro Díaz, and A. Mangeney. A new Savage–Hutter type model for submarine avalanches and generated tsunami. Journal of Computational Physics, 227(16):7720–7754, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800260X.

Fang:2007:SHO

[FNS07] Qirong Fang, David P. Nicholls, and Jie Shen. A stable, high-order method for three-dimensional, boundedobstacle, acoustic scattering. Journal of Computational Physics, 224(2):1145–1169, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005699.

Fidkowski:2005:MSH

[FOLD05] Krzysztof J. Fidkowski, Todd A. Oliver, James Lu, and David L. Darmofal. p-multigrid solution of high-order discontinuous Galerkin discretizations of the compressible Navier–Stokes equations. Journal of Computational Physics, 207(1):92–113, July 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000185.

Fournier:2006:ECF

[Fou06] Aimé Fournier. Exact calculation of Fourier series in nonconforming spectral-element methods. Journal of Computational Physics, 215(1):1–5, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004973.

Fox:2008:QBT

[Fox08] R. O. Fox. A quadrature-based third-order moment method for dilute gas-particle flows. Journal of Computational Physics, 227(12):6313–6350, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001599.

Fox:2009:HOQ

[Fox09] R. O. Fox. Higher-order quadrature-based moment methods for kinetic equations. Journal of Computational Physics, 228(20):7771–7791, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004070.

Filbet:2002:NMA

[FP02] Francis Filbet and Lorenzo Pareschi. A numerical method for the accurate solution of the Fokker–Planck–Landau equation in the nonhomogeneous case. Journal of Computational Physics, 179(1):1–26, June 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970105.

Fang:2008:RLF

[FP08a] Jiannong Fang and Aurèle Parriaux. A regularized Lagrangian finite point method for the simulation of incompressible viscous flows. Journal of Computational Physics, 227(20):8894–8908, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003653.

Fornberg:2008:CRB

[FP08b] Bengt Fornberg and Cécile Piret. On choosing a radial basis function and a shape parameter when solving a convective PDE on a sphere. Journal of Computational Physics, 227(5):2758–2780, February 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005025.

Fangohr:2000:EMH

[FPC^{+}00] Hans Fangohr, Andrew R. Price, Simon J. Cox, Peter A. J. de Groot, Geoffrey J. Daniell, and Ken S. Thomas. Efficient methods for handling long-range forces in particle-particle simulations. Journal of Computational Physics, 162(2):372–384, August 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965410.

Fedosov:2008:VLD

[FPK08] Dmitry A. Fedosov, Igor V. Pivkin, and George Em Karniadakis. Velocity limit in DPD simulations of wall-bounded flows. Journal of Computational Physics, 227(4):2540–2559, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004809.

Filbet:2005:ANM

[FPT05] Francis Filbet, Lorenzo Pareschi, and Giuseppe Toscani. Accurate numerical methods for the collisional motion of (heated) granular flows. Journal of Computational Physics, 202(1):216–235, January 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002748.

Frank:2002:PMM

[FR02] Jason Frank and Sebastian Reich. A particle-mesh method for the shallow water equations near geostrophic balance. Journal of Computational Physics, 180(2):407–426, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970580.

Filbet:2003:HON

[FR03] Francis Filbet and Giovanni Russo. High order numerical methods for the space non-homogeneous Boltzmann equation. Journal of Computational Physics, 186(2):457–480, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000652.

Frandsen:2004:SME

[Fra04] Jannette B. Frandsen. Sloshing motions in excited tanks. Journal of Computational Physics, 196(1):53–87, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005862.

Freund:2000:SMC

[Fre00] J. B. Freund. A simple method for computing far-field sound in aeroacoustic computations. Journal of Computational Physics, 157(2):796–800, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963921.

Farhat:2008:HOG

[FRS08] Charbel Farhat, Arthur Rallu, and Sriram Shankaran. A higher-order generalized ghost fluid method for the poor for the three-dimensional two-phase flow computation of underwater implosions. Journal of Computational Physics, 227(16):7674–7700, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002568.

Frink:2000:TTDa

[FS00a] Laura J. Douglas Frink and Andrew G. Salinger. Two- and three-dimensional nonlocal density functional theory for inhomogeneous fluids: I. Algorithms and parallelization. Journal of Computational Physics, 159(2):407–424, April 10, 2000.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964544.

Frink:2000:TTDb

[FS00b] Laura J. Douglas Frink and Andrew G. Salinger. Two- and three-dimensional nonlocal density functional theory for inhomogeneous fluids: II. Solvated polymers as a benchmark problem. Journal of Computational Physics, 159(2):425–439, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964556.

Fan:2001:SSL

[FS01] Jing Fan and Ching Shen. Statistical simulation of low-speed rarefied gas flows. Journal of Computational Physics, 167(2):393–412, March 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966816.

Fast:2004:MOG

[FS04] Petri Fast and Michael J. Shelley. A moving overset grid method for interface dynamics applied to non-Newtonian Hele–Shaw flow. Journal of Computational Physics, 195(1):117–142, March 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005138.

Fast:2006:MLS

[FS06] Petri Fast and Michael J. Shelley. Moore’s law and the Saffman–Taylor instability. Journal of Computational Physics, 212(1):1–5, February 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003165.

Farjoun:2009:ECP

[FS09] Yossi Farjoun and Benjamin Seibold. An exactly conservative particle method for one dimensional scalar conservation laws. Journal of Computational Physics, 228(14):5298–5315, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002046.

Filbet:2001:CNS

[FSB01] Francis Filbet, Eric Sonnendrücker, and Pierre Bertrand. Conservative numerical schemes for the Vlasov equation. Journal of Computational Physics, 172(1):166–187, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968184.

Filippova:2001:MLB

[FSM^{+}01] Olga Filippova, Sauro Succi, Francesco Mazzocco, Cinzio Arrighetti, Gino Bella, and Dieter Hänel. Multiscale lattice Boltzmann schemes with turbulence modeling. Journal of Computational Physics, 170(2):812–829, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967646.

Fedkiw:2003:SCL

[FSS03] Ronald P. Fedkiw, Guillermo Sapiro, and Chi-Wang Shu. Shock capturing, level sets, and PDE based methods in computer vision and image processing: a review of Osher’s contributions. Journal of Computational Physics, 185(2):309–341, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000165.

Friese:2000:TBC

[FSY00] Tilmann Friese, Frank Schmidt, and David Yevick. Transparent boundary conditions for a wide-angle approximation of the one-way Helmholtz equation. Journal of Computational Physics, 165(2):645–659, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966373.

Fibich:2001:HOT

[FT01] Gadi Fibich and Semyon Tsynkov. High-order two-way artificial boundary conditions for nonlinear wave propagation with backscattering. Journal of Computational Physics, 171(2):632–677, August 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968007.

Fibich:2005:NSN

[FT05] G. Fibich and S. Tsynkov. Numerical solution of the nonlinear Helmholtz equation using nonorthogonal expansions. Journal of Computational Physics, 210(1):183–224, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002068.

Feigl:2006:DEM

[FT06] Kathleen Feigl and Franz X. Tanner. Development and evaluation of a micro-macro algorithm for the simulation of polymer flow. Journal of Computational Physics, 216(1):92–113, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005449.

Fukui:2009:OCM

[FT09] Kouki Fukui and Synge Todo. Order-N cluster Monte Carlo method for spin systems with long-range interactions. Journal of Computational Physics, 228(7):2629–2642, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006621.

Frontera:2001:AFF

[FV01] Carlos Frontera and Eduard Vives. An algorithm for finding the first excited state in the random-field Ising model. Journal of Computational Physics, 168(1):219–226, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966956.

Fatkullin:2004:CSM

[FVE04] Ibrahim Fatkullin and Eric Vanden-Eijnden. A computational strategy for multiscale systems with applications to Lorenz 96 model. Journal of Computational Physics, 200(2):605–638, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001810.

Fadlun:2000:CIB

[FVOMY00] E. A. Fadlun, R. Verzicco, P. Orlandi, and J. Mohd-Yusof. Combined immersed-boundary finite-difference methods for three-dimensional complex flow simulations. Journal of Computational Physics, 161(1):35–60, June 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964842.

Flyer:2007:TSS

[FW07] Natasha Flyer and Grady B. Wright. Transport schemes on a sphere using radial basis functions. Journal of Computational Physics, 226(1):1059–1084, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002215.

Foo:2008:MEP

[FWK08] Jasmine Foo, Xiaoliang Wan, and George Em Karniadakis. The multi-element probabilistic collocation method (ME– PCM): Error analysis and applications. Journal of Computational Physics, 227(22):9572–9595, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003707.

Fichtl:2009:KIM

[FWP09] Erin D. Fichtl, James S. Warsa, and Anil K. Prinja. Krylov iterative methods and synthetic acceleration for transport in binary statistical media. Journal of Computational Physics, 228(22):8413–8426, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004562.

Fisher:2007:EVF

[FWR07] A. Fisher, D. White, and G. Rodrigue. An efficient vector finite element method for nonlinear electromagnetic modeling. Journal of Computational Physics, 225(2):1331–1346, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000447.

Fleming:2004:LCN

[FWW04] John L. Fleming, Aihua W. Wood, and William D. Wood, Jr. Locally corrected Nyström method for EM scattering by bodies of revolution. Journal of Computational Physics, 196(1):41–52, May 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005850.

Fujita:2007:MSM

[FY07] M. Fujita and Y. Yamaguchi. Multiscale simulation method for self-organization of nanoparticles in dense suspension. Journal of Computational Physics, 223(1):108–120, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600427X.

Feng:2006:SIA

[FYH^{+}06] W. M. Feng, P. Yu, S. Y. Hu, Z. K. Liu, Q. Du, and L. Q. Chen. Spectral implementation of an adaptive moving mesh method for phase-field equations. Journal of Computational Physics, 220(1):498–510, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003445.

Gong:2009:CFS

[GA09] Minwei Gong and Yiannis Andreopoulos. Coupled fluidstructure solver: The case of shock wave impact on monolithic and composite material plates. Journal of Computational Physics, 228(12):4400–4434, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900134X.

Gabard:2007:DGM

[Gab07] Gwénaël Gabard. Discontinuous Galerkin methods with plane waves for time-harmonic problems. Journal of Computational Physics, 225(2):1961–1984, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000964.

Grenier:2009:HIS

[GAC^{+}09] N. Grenier, M. Antuono, A. Colagrossi, D. Le Touzé, and B. Alessandrini. An Hamiltonian interface SPH formulation for multi-fluid and free surface flows. Journal of Computational Physics, 228(22):8380–8393, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900446X.

Givelberg:2003:CTD

[GB03] Edward Givelberg and Julian Bunn. A comprehensive threedimensional model of the cochlea. Journal of Computational Physics, 191(2):377–391, November 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300319X.

Gatti-Bono:2008:CDA

[GB08a] Caroline Gatti-Bono. Coupling and decoupling of the acoustic and gravity waves through perturbational analysis of the Euler equations. Journal of Computational Physics, 227(3):1609–1612, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004378.

Gubernatis:2008:MEE

[GB08b] J. E. Gubernatis and T. E. Booth. Multiple extremal eigenpairs by the power method. Journal of Computational Physics, 227(19):8508–8522, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003136.

Grandgirard:2006:DKS

[GBB^{+}06] V. Grandgirard, M. Brunetti, P. Bertrand, N. Besse, X. Garbet, P. Ghendrih, G. Manfredi, Y. Sarazin, O. Sauter, E. Sonnendrücker, J. Vaclavik, and L. Villard. A drift-kinetic semi-Lagrangian 4D code for ion turbulence simulation. Journal of Computational Physics, 217(2):395–423, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000155.

Gatti-Bono:2006:AAP

[GBC06] Caroline Gatti-Bono and Phillip Colella. An anelastic allspeed projection method for gravitationally stratified flows. Journal of Computational Physics, 216(2):589–615, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000027.

Grandclement:2001:MSM

[GBGM01] P. Grandclément, S. Bonazzola, E. Gourgoulhon, and J.-A. Marck. A multidomain spectral method for scalar and vectorial Poisson equations with noncompact sources. Journal of Computational Physics, 170(1):231–260, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967348.

Goswami:2000:SEM

[GBS00] Priyanka Goswami, S. N. Bhattacharyya, and A. Sen. Spectrum of electrostatic modes in a cylindrical non-neutral plasma of arbitrary density. Journal of Computational Physics, 159(2):312–328, April 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964477.

Greer:2006:FOP

[GBS06] John B. Greer, Andrea L. Bertozzi, and Guillermo Sapiro. Fourth order partial differential equations on general geometries. Journal of Computational Physics, 216(1):216–246, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005498.

Gascon:2001:CSO

[GC01] Ll. Gascón and J. M. Corberán. Construction of secondorder TVD schemes for nonhomogeneous hyperbolic conservation laws. Journal of Computational Physics, 172(1): 261–297, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968238.

Giona:2002:IDF

[GC02a] Massimiliano Giona and Stefano Cerbelli. C^{∞}-interpolation of discrete fields on regular and irregular grids. Journal of Computational Physics, 176(1):145–169, February 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969724.

Greve:2002:CNS

[GC02b] Ralf Greve and Reinhard Calov. Comparison of numerical schemes for the solution of the ice-thickness equation in a dynamic/thermodynamic ice-sheet model. Journal of Computational Physics, 179(2):649–664, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970816.

Gobbert:2006:KTR

[GC06] Matthias K. Gobbert and Timothy S. Cale. A kinetic transport and reaction model and simulator for rarefied gas flow in the transition regime. Journal of Computational Physics, 213(2):591–612, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003943.

Gonzalez:2007:NEU

[GCCD07] D. González, E. Cueto, F. Chinesta, and M. Doblaré. A natural element updated Lagrangian strategy for free-surface fluid dynamics. Journal of Computational Physics, 223(1):127–150, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004293.

Garcia-Cervera:2003:ANM

[GCGE03] Carlos J. García-Cervera, Zydrunas Gimbutas, and Weinan E. Accurate numerical methods for micromagnetics simulations with general geometries. Journal of Computational Physics, 184(1):37–52, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000141.

Greenwood:2004:ENC

[GCLB04] Andrew D. Greenwood, Keith L. Cartwright, John W. Luginsland, and Ernest A. Baca. On the elimination of numerical Cerenkov radiation in PIC simulations. Journal of Computational Physics, 201(2):665–684, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002608.

Gibou:2007:LSB

[GCNB07] Frédéric Gibou, Liguo Chen, Duc Nguyen, and Sanjoy Banerjee. A level set based sharp interface method for the multiphase incompressible Navier–Stokes equations with phase change. Journal of Computational Physics, 222(2):536–555, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106003652.

Garcia-Cervera:2007:SPC

[GCW07] Carlos J. García-Cervera and Xiao-Ping Wang. Spin-polarized currents in ferromagnetic multilayers. Journal of Computational Physics, 224(2):699–711, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005456.

Guo:2005:GSL

[GD05] Daniel X. Guo and John B. Drake. A global semi-Lagrangian spectral model of the shallow water equations with variable resolution. Journal of Computational Physics, 206(2):559– 577, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000057.

Ghanem:2006:CAS

[GD06a] Roger G. Ghanem and Alireza Doostan. On the construction and analysis of stochastic models: Characterization and propagation of the errors associated with limited data. Journal of Computational Physics, 217(1):63–81, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000386.

Gumerov:2006:FMM

[GD06b] Nail A. Gumerov and Ramani Duraiswami. Fast multipole method for the biharmonic equation in three dimensions. Journal of Computational Physics, 215(1):363–383, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004924.

Guillard:2007:DLD

[GD07a] H. Guillard and F. Duval. A Darcy law for the drift velocity in a two-phase flow model. Journal of Computational Physics, 224(1):288–313, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000769.

Gumerov:2007:SPF

[GD07b] Nail A. Gumerov and Ramani Duraiswami. A scalar potential formulation and translation theory for the timeharmonic Maxwell equations. Journal of Computational Physics, 225(1):206–236, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005845.

Gumerov:2008:FMM

[GD08] Nail A. Gumerov and Ramani Duraiswami. Fast multipole methods on graphics processors. Journal of Computational Physics, 227(18):8290–8313, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002921.

Gu:2007:CSR

[GE07] X. J. Gu and D. R. Emerson. A computational strategy for the regularized 13 moment equations with enhanced wall-boundary conditions. Journal of Computational Physics, 225(1):263–283, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005882.

Gelfgat:2006:IAI

[Gel06] Alexander Yu. Gelfgat. Implementation of arbitrary inner product in the global Galerkin method for incompressible Navier–Stokes equations. Journal of Computational Physics, 211(2):513–530, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002858.

Gentile:2001:IMC

[Gen01] N. A. Gentile. Implicit Monte Carlo diffusion — an acceleration method for Monte Carlo time-dependent radiative transfer simulations. Journal of Computational Physics, 172(2):543–571, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968366.

George:2008:ARS

[Geo08] David L. George. Augmented Riemann solvers for the shallow water equations over variable topography with steady states and inundation. Journal of Computational Physics, 227(6):3089–3113, March 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004767.

Giese:2002:GMH

[GF02] Guido Giese and Michael Fey. A genuinely multidimensional high-resolution scheme for the elastic-plastic wave equation. Journal of Computational Physics, 181(1):338–353, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971299.

Gibou:2005:FOA

[GF05a] Frédéric Gibou and Ronald Fedkiw. A fourth order accurate discretization for the Laplace and heat equations on arbitrary domains, with applications to the Stefan problem. Journal of Computational Physics, 202(2):577–601, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002980.

Guy:2005:SAP

[GF05b] Robert D. Guy and Aaron L. Fogelson. Stability of approximate projection methods on cell-centered grids. Journal of Computational Physics, 203(2):517–538, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003900.

Gibou:2002:SOA

[GFCK02] Frederic Gibou, Ronald P. Fedkiw, Li-Tien Cheng, and Myungjoo Kang. A second-order-accurate symmetric discretization of the Poisson equation on irregular domains. Journal of Computational Physics, 176(1):205–227, February 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969773.

Gao:2009:PFM

[GFG09] Ling-Tian Gao, Xi-Qiao Feng, and Huajian Gao. A phase field method for simulating morphological evolution of vesicles in electric fields. Journal of Computational Physics, 228(11):4162–4181, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001065.

Giannakis:2009:SGM

[GFR09] Dimitrios Giannakis, Paul F. Fischer, and Robert Rosner. A spectral Galerkin method for the coupled Orr–Sommerfeld and induction equations for free-surface MHD. Journal of Computational Physics, 228(4):1188–1233, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005445.

Gavrilyuk:2008:MWD

[GFS08] S. L. Gavrilyuk, N. Favrie, and R. Saurel. Modelling wave dynamics of compressible elastic materials. Journal of Computational Physics, 227(5):2941–2969, February 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005220.

Garcke:2000:CEH

[GG00] Jochen Garcke and Michael Griebel. On the computation of the eigenproblems of hydrogen and helium in strong magnetic and electric fields with the sparse grid combination technique. Journal of Computational Physics, 165(2):694–716, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966270.

Ganesh:2004:HOA

[GG04] M. Ganesh and I. G. Graham. A high-order algorithm for obstacle scattering in three dimensions. Journal of Computational Physics, 198(1):211–242, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000257.

Gimbutas:2009:FSM

[GG09a] Z. Gimbutas and L. Greengard. A fast and stable method for rotating spherical harmonic expansions. Journal of Computational Physics, 228(16):5621–5627, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002691.

Gotovac:2009:MEA

[GG09b] Hrvoje Gotovac and Blaz Gotovac. Maximum entropy algorithm with inexact upper entropy bound based on fup basis functions with compact support. Journal of Computational Physics, 228(24):9079–9091, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004963.

Guidoboni:2009:SLC

[GGCC09] Giovanna Guidoboni, Roland Glowinski, Nicola Cavallini, and Suncica Canic. Stable loosely-coupled-type algorithm for fluid-structure interaction in blood flow. Journal of Computational Physics, 228(18):6916–6937, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003295.

Geuzaine:2003:DAA

[GGF03] Philippe Geuzaine, Céline Grandmont, and Charbel Farhat. Design and analysis of ALE schemes with provable secondorder time-accuracy for inviscid and viscous flow simulations. Journal of Computational Physics, 191(1):206–227, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003115.

Glimm:2001:CAR

[GGL^{+}01] J. Glimm, J. W. Grove, X. L. Li, W. Oh, and D. H. Sharp. A critical analysis of Rayleigh–Taylor growth rates. Journal of Computational Physics, 169(2):652–677, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965902.

Gerardo-Giorda:2009:MBB

[GGMN^{+}09] L. Gerardo-Giorda, L. Mirabella, F. Nobile, M. Perego, and A. Veneziani. A model-based block-triangular preconditioner for the bidomain system in electrocardiology. Journal of Computational Physics, 228(10):3625–3639, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000618.

Gondarenko:2004:PML

[GGOB04] Natalia A. Gondarenko, Parvez N. Guzdar, Sidney L. Ossakow, and Paul A. Bernhardt. Perfectly matched layers for radio wave propagation in inhomogeneous magnetized plasmas. Journal of Computational Physics, 194(2):481–504, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004923.

Glowinski:2006:WDI

[GGP06] R. Glowinski, G. Guidoboni, and T.-W. Pan. Wall-driven incompressible viscous flow in a two-dimensional semi-circular cavity. Journal of Computational Physics, 216(1):76–91, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005437.

Gonzalez:2008:ESM

[GGRS08] Leonel Gonzalez, Shekhar Guha, James W. Rogers, and Qin Sheng. An effective z-stretching method for paraxial light beam propagation simulations. Journal of Computational Physics, 227(15):7264–7278, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002258.

Garzon:2009:NSN

[GGS09] M. Garzon, L. J. Gray, and J. A. Sethian. Numerical simulation of non-viscous liquid pinch-off using a coupled level set-boundary integral method. Journal of Computational Physics, 228(17):6079–6106, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002344.

Greengard:2002:NVF

[GH02] Leslie F. Greengard and Jingfang Huang. A new version of the fast multipole method for screened Coulomb interactions in three dimensions. Journal of Computational Physics, 180(2):642–658, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297110X.

Garba:2003:CPC

[GH03] Abdou Garba and Pierre Haldenwang. Comparison of preconditioners for collocation Chebyshev approximation of 2D and 3D generalized Stokes problem. Journal of Computational Physics, 191(1):282–304, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003176.

Ganesh:2008:HOT

[GH08a] M. Ganesh and S. C. Hawkins. A high-order tangential basis algorithm for electromagnetic scattering by curved surfaces. Journal of Computational Physics, 227(9):4543–4562, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108000351.

Guan:2008:FDM

[GH08b] Wei Guan and Hengshan Hu. Finite-difference modeling of the electroseismic logging in a fluid-saturated porous formation. Journal of Computational Physics, 227(11):5633–5648, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001022.

Gao:2009:DEP

[GH09] Tong Gao and Howard H. Hu. Deformation of elastic particles in viscous shear flow. Journal of Computational Physics, 228(6):2132–2151, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006219.

Ghizzo:2003:NPS

[GHB03] A. Ghizzo, F. Huot, and P. Bertrand. A non-periodic 2D semi-Lagrangian Vlasov code for laser-plasma interaction on parallel computer. Journal of Computational Physics, 186(1):47–69, March 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300010X.

Ghadiali:2001:DRB

[GHG01] Samir N. Ghadiali, David Halpern, and Donald P. Gaver III. A dual-reciprocity boundary element method for evaluating bulk convective transport of surfactant in free-surface flows. Journal of Computational Physics, 171(2):534–559, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967920.

Griffith:2007:AFS

[GHMP07] Boyce E. Griffith, Richard D. Hornung, David M. McQueen, and Charles S. Peskin. An adaptive, formally second order accurate version of the immersed boundary method. Journal of Computational Physics, 223(1):10–49, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004207.

Gilles:2000:CBS

[GHV00] L. Gilles, S. C. Hagness, and L. Vázquez. Comparison between staggered and unstaggered finite-difference time-domain grids for few-cycle temporal optical soliton propagation. Journal of Computational Physics, 161(2):379–400, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096460X.

Giraldo:2002:NHO

[GHW02] F. X. Giraldo, J. S. Hesthaven, and T. Warburton. Nodal high-order discontinuous Galerkin methods for the spherical shallow water equations. Journal of Computational Physics, 181(2):499–525, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971391.

Gitelman:2007:MSL

[GIA^{+}07] L. Gitelman, M. Israeli, A. Averbuch, M. Nathan, Z. Schuss, and D. Golodnitsky. Modeling and simulation of li-ion conduction in poly(ethylene oxide). Journal of Computational Physics, 227(2):1162–1175, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003865.

Gitelman:2008:PGL

[GIA^{+}08] L. Gitelman, M. Israeli, A. Averbuch, M. Nathan, Z. Schuss, and D. Golodnitsky. Polymer geometry and Li ^{+} conduction in poly(ethylene oxide). Journal of Computational Physics, 227(18):8437–8447, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003264.

Gibbs:2004:PRA

[Gib04] W. R. Gibbs. A parallel/recursive algorithm. Journal of Computational Physics, 201(2):573–585, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002438.

Giraldo:2000:LGM

[Gir00] Francis X. Giraldo. Lagrange–Galerkin methods on spherical geodesic grids: The shallow water equations. Journal of Computational Physics, 160(1):336–368, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964696.

Giraldo:2006:HOT

[Gir06] Francis X. Giraldo. High-order triangle-based discontinuous Galerkin methods for hyperbolic equations on a rotating sphere. Journal of Computational Physics, 214(2):447–465, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004535.

Givoli:2001:HON

[Giv01] Dan Givoli. High-order nonreflecting boundary conditions without high-order derivatives. Journal of Computational Physics, 170(2):849–870, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196766X.

Graf:2009:NVC

[GJK09] Peter A. Graf, Wesley B. Jones, and Kwiseon Kim. A note on the virtual crystal approach to alloy optimization. Journal of Computational Physics, 228(12):4309–4311, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001454.

Giladi:2001:HNA

[GK01] Eldar Giladi and Joseph B. Keller. A hybrid numerical asymptotic method for scattering problems. Journal of Computational Physics, 174(1):226–247, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969037.

Goodson:2002:ESA

[GK02] Michael Goodson and Markus Kraft. An efficient stochastic algorithm for simulating nano-particle dynamics. Journal of Computational Physics, 183(1):210–232, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971925.

Gear:2003:TPM

[GK03] C. W. Gear and Ioannis G. Kevrekidis. Telescopic projective methods for parabolic differential equations. Journal of Computational Physics, 187(1):95–109, May 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000822.

Grote:2004:DNB

[GK04] Marcus J. Grote and Christoph Kirsch. Dirichlet-to-Neumann boundary conditions for multiple scattering problems. Journal of Computational Physics, 201(2):630–650, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400258X.

Gupta:2005:NPS

[GK05] Murli M. Gupta and Jiten C. Kalita. A new paradigm for solving Navier–Stokes equations: streamfunction-velocity formulation. Journal of Computational Physics, 207(1):52–68, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000161.

Grote:2007:NBC

[GK07] Marcus J. Grote and Christoph Kirsch. Nonreflecting boundary condition for time-dependent multiple scattering. Journal of Computational Physics, 221(1):41–62, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002749.

Ginste:2009:ECP

[GKD09] Dries Vande Ginste, Luc Knockaert, and Daniël De Zutter. Error control in the perfectly matched layer based multilevel fast multipole algorithm. Journal of Computational Physics, 228(13):4811–4822, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001661.

Green:2004:ODU

[GKE04] Kirk Green, Bernd Krauskopf, and Koen Engelborghs. Onedimensional unstable eigenfunction and manifold computations in delay differential equations. Journal of Computational Physics, 197(1):86–98, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300620X.

Graf:2007:SPO

[GKJW07] Peter A. Graf, Kwiseon Kim, Wesley B. Jones, and Lin-Wang Wang. Surface passivation optimization using DIRECT. Journal of Computational Physics, 224(2):824–835, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600550X.

Gryazin:2000:GCH

[GKL00] Yuriy A. Gryazin, Michael V. Klibanov, and Thomas R. Lucas. GMRES computation of high frequency electrical field propagation in land mine detection. Journal of Computational Physics, 158(1):98–115, February 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964094.

Gryazin:2003:TNM

[GKL03] Yuriy A. Gryazin, Michael V. Klibanov, and Thomas R. Lucas. Two numerical methods for an inverse problem for the 2-D Helmholtz equation. Journal of Computational Physics, 184(1):122–148, January 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000232.

Grigoriadis:2009:IBM

[GKV09] D. G. E. Grigoriadis, S. C. Kassinos, and E. V. Votyakov. Immersed boundary method for the MHD flows of liquid metals. Journal of Computational Physics, 228(3):903–920, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005305.

Greengard:2006:EHC

[GL06] Leslie Greengard and June-Yub Lee. Electrostatics and heat conduction in high contrast composite materials. Journal of Computational Physics, 211(1):64–76, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002615.

Gyrya:2008:HOM

[GL08] Vitaliy Gyrya and Konstantin Lipnikov. High-order mimetic finite difference method for diffusion problems on polygonal meshes. Journal of Computational Physics, 227(20): 8841–8854, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108003483.

Goodman:2009:CCV

[GL09a] Jonathan B. Goodman and Kevin K. Lin. Coupling control variates for Markov chain Monte Carlo. Journal of Computational Physics, 228(19):7127–7136, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001685.

Gunter:2009:MIE

[GL09b] S. Günter and K. Lackner. A mixed implicit-explicit finite difference scheme for heat transport in magnetised plasmas. Journal of Computational Physics, 228(2):282–293, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004786.

Glasner:2001:NPD

[Gla01] Karl Glasner. Nonlinear preconditioning for diffuse interfaces. Journal of Computational Physics, 174(2):695–711, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969335.

Glasner:2005:BIF

[Gla05] K. B. Glasner. A boundary integral formulation of quasi-steady fluid wetting. Journal of Computational Physics, 207(2):529–541, August 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000392.

Gerbeau:2003:SMF

[GLL03] J.-F. Gerbeau, T. Lelièvre, and C. Le Bris. Simulations of MHD flows with moving interfaces. Journal of Computational Physics, 184(1):163–191, January 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000256.

Guermond:2007:IPG

[GLLN07] J.-L. Guermond, R. Laguerre, J. Léorat, and C. Nore. An interior penalty Galerkin method for the MHD equations in heterogeneous domains. Journal of Computational Physics, 221(1):349–369, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002944.

Guermond:2009:NMA

[GLLN09] J.-L. Guermond, R. Laguerre, J. Léorat, and C. Nore. Nonlinear magnetohydrodynamics in axisymmetric heterogeneous domains using a Fourier/finite element technique and an interior penalty method. Journal of Computational Physics, 228(8):2739–2757, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006463.

Guo:2008:CSL

[GLLX08] Zhaoli Guo, Hongwei Liu, Li-Shi Luo, and Kun Xu. A comparative study of the LBE and GKS methods for 2D near incompressible laminar flows. Journal of Computational Physics, 227(10):4955–4976, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000545.

Gassner:2007:CCD

[GLM07] Gregor Gassner, Frieder Lörcher, and Claus-Dieter Munz. A contribution to the construction of diffusion fluxes for finite volume and discontinuous Galerkin schemes. Journal of Computational Physics, 224(2):1049–1063, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005651.

Gassner:2009:PNE

[GLMH09] Gregor J. Gassner, Frieder Lörcher, Claus-Dieter Munz, and Jan S. Hesthaven. Polymorphic nodal elements and their application in discontinuous Galerkin methods. Journal of Computational Physics, 228(5):1573–1590, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005834.

Gourdain:2006:HRM

[GLN06] P.-A. Gourdain, J.-N. Leboeuf, and R. Y. Neches. Highresolution magnetohydrodynamic equilibrium code for unity beta plasmas. Journal of Computational Physics, 216(1): 275–299, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105005528.

Girard:2003:NSP

[GLS03] Alain Girard, Christian Lécot, and Konstantin Serebrennikov. Numerical simulation of the plasma of an electron cyclotron resonance ion source. Journal of Computational Physics, 191(1):228–248, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003127.

Gunter:2007:FEH

[GLT07] S. Günter, K. Lackner, and C. Tichmann. Finite element and higher order difference formulations for modelling heat transport in magnetised plasmas. Journal of Computational Physics, 226(2):2306–2316, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003166.

Gelbard:2001:MQS

[GM01a] F. Gelbard and K. J. Malloy. Modeling quantum structures with the boundary element method. Journal of Computational Physics, 172(1):19–39, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967518.

Gremaud:2001:CSH

[GM01b] Pierre A. Gremaud and John V. Matthews. On the computation of steady hopper flows: I. Stress determination for Coulomb materials. Journal of Computational Physics, 166(1):63–83, January 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966415.

Guseinov:2001:CTF

[GM01c] I. I. Guseinov and B. A. Mamedov. Convergence of translation formulas for the computation of multicenter integrals over Slater orbitals. Journal of Computational Physics, 174(1):428–437, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969244.

Gosse:2004:MSAa

[GM04] Laurent Gosse and Peter A. Markowich. Multiphase semiclassical approximation of an electron in a one-dimensional crystalline lattice: I. Homogeneous problems. Journal of Computational Physics, 197(2):387–417, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006454.

Gosse:2006:MSA

[GM06] Laurent Gosse and Norbert J. Mauser. Multiphase semiclassical approximation of an electron in a one-dimensional crystalline lattice — III. From ab initio models to WKB for Schrödinger–Poisson. Journal of Computational Physics, 211(1):326–346, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002779.

Goldsworthy:2009:SUF

[GMAj09] Mark Goldsworthy, Michael Macrossan, and Madhat Abdeljawad. Simulation of unsteady flows by the DSMC macroscopic chemistry method. Journal of Computational Physics, 228(4):976–982, March 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004750.

Gerlinger:2001:IMM

[GMB01] Peter Gerlinger, Helge Möbus, and Dieter Brüggemann. An implicit multigrid method for turbulent combustion. Journal of Computational Physics, 167(2):247–276, March 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966713.

Gao:2003:NSW

[GMD03] D. Gao, N. B. Morley, and V. Dhir. Numerical simulation of wavy falling film flow using VOF method. Journal of Computational Physics, 192(2):624–642, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910300411X.

Ghias:2007:SII

[GMD07] R. Ghias, R. Mittal, and H. Dong. A sharp interface immersed boundary method for compressible viscous flows. Journal of Computational Physics, 225(1):528–553, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006048.

Griesheimer:2006:CPM

[GMH06] David P. Griesheimer, William R. Martin, and James Paul Holloway. Convergence properties of Monte Carlo functional expansion tallies. Journal of Computational Physics, 211 (1):129–153, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002652.

Gremaud:2004:CSH

[GMO04] Pierre A. Gremaud, John V. Matthews, and Meghan O’Malley. On the computation of steady hopper flows: II: von Mises materials in various geometries. Journal of Computational Physics, 200(2):639–653, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001822.

Gremaud:2006:CSH

[GMS06] Pierre A. Gremaud, John V. Matthews, and David G. Schaeffer. On the computation of steady hopper flows III: Model comparisons. Journal of Computational Physics, 219(1):443–454, November 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001719.

Givoli:2003:HON

[GN03] Dan Givoli and Beny Neta. High-order non-reflecting boundary scheme for time-dependent waves. Journal of Computational Physics, 186(1):24–46, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000056.

Gong:2007:SEH

[GN07] Jing Gong and Jan Nordström. A stable and efficient hybrid scheme for viscous problems in complex geometries. Journal of Computational Physics, 226(2):1291–1309, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002343.

Gois:2008:FTM

[GNNB08] João Paulo Gois, Anderson Nakano, Luis Gustavo Nonato, and Gustavo C. Buscaglia. Front tracking with movingleast-squares surfaces. Journal of Computational Physics, 227(22):9643–9669, November 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003951.

Goedbloed:2000:EFT

[Goe00] J. P. Goedbloed. Expansion functions for two-dimensional incompressible fluid flow in arbitrary domains. Journal of Computational Physics, 160(1):283–297, May 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964623.

Gomez:2008:PWD

[Gom08] T. Gomez. Pseudo-wave decomposition high-order method for magnetogasdynamics. Journal of Computational Physics, 227(20):8909–8921, October 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003689.

Gonnet:2007:PSQ

[Gon07] Pedro Gonnet. P–SHAKE: a quadratically convergent SHAKE in O(n^{2}). Journal of Computational Physics, 220 (2):740–750, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106002579.

Gosse:2002:UBE

[Gos02] Laurent Gosse. Using K-branch entropy solutions for multivalued geometric optics computations. Journal of Computational Physics, 180(1):155–182, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970853.

Gosse:2004:MSAb

[Gos04] Laurent Gosse. Multiphase semiclassical approximation of an electron in a one-dimensional crystalline lattice II. Impurities, confinement and Bloch oscillations. Journal of Computational Physics, 201(1):344–375, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002335.

Gerritsma:2000:SEM

[GP00a] M. I. Gerritsma and T. N. Phillips. Spectral element methods for axisymmetric Stokes problems. Journal of Computational Physics, 164(1):81–103, October 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965744.

Gonthier:2000:HRN

[GP00b] Keith A. Gonthier and Joseph M. Powers. A highresolution numerical method for a two-phase model of deflagration-to-detonation transition. Journal of Computational Physics, 163(2):376–433, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965690.

Givoli:2004:DNB

[GP04] Dan Givoli and Igor Patlashenko. Dirichlet-to-Neumann boundary condition for time-dependent dispersive waves in three-dimensional guides. Journal of Computational Physics, 199(1):339–354, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000877.

Griffith:2005:OAI

[GP05] Boyce E. Griffith and Charles S. Peskin. On the order of accuracy of the immersed boundary method: Higher order convergence rates for sufficiently smooth problems. Journal of Computational Physics, 208(1):75–105, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000835.

Gallardo:2007:WBH

[GPC07] José M. Gallardo, Carlos Parés, and Manuel Castro. On a well-balanced high-order finite volume scheme for shallow water equations with topography and dry areas. Journal of Computational Physics, 227(1):574–601, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003488.

Giraldo:2003:SES

[GPF03] F. X. Giraldo, J. B. Perot, and P. F. Fischer. A spectral element semi-Lagrangian (SESL) method for the spherical shallow water equations. Journal of Computational Physics, 190(2):623–650, September 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003000.

Glowinski:2001:FDA

[GPH^{+}01] R. Glowinski, T. W. Pan, T. I. Hesla, D. D. Joseph, and J. Périaux. A fictitious domain approach to the direct numerical simulation of incompressible viscous flow past moving rigid bodies: Application to particulate flow. Journal of Computational Physics, 169(2):363–426, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965422.

Goyal:2005:NDL

[GPL05] S. Goyal, N. C. Perkins, and C. L. Lee. Nonlinear dynamics and loop formation in Kirchhoff rods with implications to the mechanics of DNA and cables. Journal of Computational Physics, 209(1):371–389, October 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500183X.

Grandison:2007:RBI

[GPVB07] Scott Grandison, Robert Penfold, and Jean-Marc Vanden-Broeck. A rapid boundary integral equation technique for protein electrostatics. Journal of Computational Physics, 224(2):663–680, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106005432.

Guermond:2000:PFV

[GQ00] J.-L. Guermond and L. Quartapelle. A projection FEM for variable density incompressible flows. Journal of Computational Physics, 165(1):167–188, November 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966099.

Garikipati:2001:RAM

[GR01] Krishna Garikipati and Vinay S. Rao. Recent advances in models for thermal oxidation of silicon. Journal of Computational Physics, 174(1):138–170, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968846.

Greenough:2004:QCN

[GR04] J. A. Greenough and W. J. Rider. A quantitative comparison of numerical methods for the compressible Euler equations: fifth-order WENO and piecewise-linear Godunov. Journal of Computational Physics, 196(1):259–281, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005965.

Gross:2007:EPF

[GR07] Sven Groß and Arnold Reusken. An extended pressure finite element space for two-phase incompressible flows with surface tension. Journal of Computational Physics, 224(1):40–58, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106006097.

Giraldo:2008:SSE

[GR08] F. X. Giraldo and M. Restelli. A study of spectral element and discontinuous Galerkin methods for the Navier– Stokes equations in nonhydrostatic mesoscale atmospheric modeling: Equation sets and test cases. Journal of Computational Physics, 227(8):3849–3877, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005384.

Gravemeier:2006:CDL

[Gra06a] Volker Gravemeier. A consistent dynamic localization model for large eddy simulation of turbulent flows based on a variational formulation. Journal of Computational Physics, 218(2):677–701, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001264.

Gravemeier:2006:SSO

[Gra06b] Volker Gravemeier. Scale-separating operators for variational multiscale large eddy simulation of turbulent flows. Journal of Computational Physics, 212(2):400–435, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003268.

Greaves:2004:QAM

[Gre04] Deborah Greaves. A quadtree adaptive method for simulating fluid flows with moving interfaces. Journal of Computational Physics, 194(1):35–56, February 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004595.

Griffith:2009:AEM

[Gri09] Boyce E. Griffith. An accurate and efficient method for the incompressible Navier–Stokes equations using the projection method as a preconditioner. Journal of Computational Physics, 228(20):7565–7595, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003787.

Grote:2000:NBC

[Gro00]  Marcus J. Grote.  Nonreflecting boundary conditions

for elastodynamic scattering. Journal of Computational Physics, 161(1):331–353, June 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965094.

Grogger:2006:FDA

[Gro06] Herwig A. Grogger. Finite difference approximations of first derivatives for two-dimensional grid singularities. Journal of Computational Physics, 217(2):642–657, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000271.

Grogger:2007:FDA

[Gro07] Herwig A. Grogger. Finite difference approximations of first derivatives for three-dimensional grid singularities. Journal of Computational Physics, 225(2):2377–2397, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001386.

Gavrilyuk:2002:MNM

[GS02] Sergey Gavrilyuk and Richard Saurel. Mathematical and numerical modeling of two-phase compressible flows with microinertia. Journal of Computational Physics, 175(1):326–360, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969517.

Garbey:2003:LSE

[GS03a] M. Garbey and W. Shyy. A least square extrapolation method for improving solution accuracy of PDE computations. Journal of Computational Physics, 186(1):1–23, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000803.

Ginzburg:2003:LBM

[GS03b] Irina Ginzburg and Konrad Steiner. Lattice Boltzmann model for free-surface flow and its application to filling process in casting. Journal of Computational Physics, 185(1): 61–99, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102000487.

Grauer:2003:NUC

[GS03c] R. Grauer and F. Spanier. A note on the use of central schemes for incompressible Navier–Stokes flows. Journal of Computational Physics, 192(2):727–731, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004030.

Guermond:2003:NCT

[GS03d] J. L. Guermond and Jie Shen. A new class of truly consistent splitting schemes for incompressible flows. Journal of Computational Physics, 192(1):262–276, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003784.

Galler:2005:DST

[GS05a] M. Galler and F. Schürrer. A deterministic solver for the transport of the AlGaNd/GaN 2D electron gas including hot-phonon and degeneracy effects. Journal of Computational Physics, 210(2):519–534, December 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002494.

Gardiner:2005:UGM

[GS05b] Thomas A. Gardiner and James M. Stone. An unsplit Godunov method for ideal MHD via constrained transport. Journal of Computational Physics, 205(2):509–539, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004784.

Gilmanov:2005:HCI

[GS05c] Anvar Gilmanov and Fotis Sotiropoulos. A hybrid Cartesian/immersed boundary method for simulating flows with 3D, geometrically complex, moving bodies. Journal of Computational Physics, 207(2):457–492, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000379.

Galler:2006:DMW

[GS06a] M. Galler and F. Schürrer. A direct multigroup-WENO solver for the 2D non-stationary Boltzmann–Poisson system for GaAs devices: GaAs–MESFET. Journal of Computational Physics, 212(2):778–797, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003724.

Gibelli:2006:SCH

[GS06b] Livio Gibelli and Bernie D. Shizgal. Spectral convergence of the Hermite basis function solution of the Vlasov equation: The free-streaming term. Journal of Computational Physics, 219(2):477–488, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002890.

Ge:2007:NMS

[GS07] Liang Ge and Fotis Sotiropoulos. A numerical method for solving the 3D unsteady incompressible Navier–Stokes equations in curvilinear domains with complex immersed boundaries. Journal of Computational Physics, 225(2):1782–1809, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000873.

Gardiner:2008:UGM

[GS08] Thomas A. Gardiner and James M. Stone. An unsplit Godunov method for ideal MHD via constrained transport in three dimensions. Journal of Computational Physics, 227(8):4123–4141, April 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005669.

Garnier:2009:TWP

[GS09a] Josselin Garnier and Knut Sølna. A two-way paraxial system for simulation of wave backscattering by a random medium. Journal of Computational Physics, 228(9):3307–3325, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000370.

Gatsonis:2009:TDE

[GS09b]  Nikolaos A. Gatsonis and Anton Spirkin.  A three-

dimensional electrostatic particle-in-cell methodology on unstructured Delaunay–Voronoi grids. Journal of Computational Physics, 228(10):3742–3761, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000692.

Guermond:2009:SMI

[GS09c] J.-L. Guermond and Abner Salgado. A splitting method for incompressible flows with variable density based on a pressure Poisson equation. Journal of Computational Physics, 228(8):2834–2846, May 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006748.

Guo:2009:GMW

[GS09d] Xin Guo and Lian Shen. On the generation and maintenance of waves and turbulence in simulations of free-surface turbulence. Journal of Computational Physics, 228(19):7313–7332, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003581.

Gilmanov:2003:GRA

[GSB03] A. Gilmanov, F. Sotiropoulos, and E. Balaras. A general reconstruction algorithm for simulating flows with complex 3D immersed boundaries on Cartesian grids. Journal of Computational Physics, 191(2):660–669, November 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003218.

Garnier:2001:CEE

[GSD01] Eric Garnier, Pierre Sagaut, and Michel Deville. A class of explicit ENO filters with application to unsteady flows. Journal of Computational Physics, 170(1):184–204, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967324.

Gunes:2006:GDK

[GSK06] Hasan Gunes, Sirod Sirisup, and George Em Karniadakis. Gappy data: To krig or not to krig? Journal of Computational Physics, 212(1):358–382, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003256.

Gil:2000:CTF

[GST00] Amparo Gil, Javier Segura, and Nico M. Temme. Computing toroidal functions for wide ranges of the parameters. Journal of Computational Physics, 161(1):204–217, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964982.

Gil:2002:EMB

[GST02] Amparo Gil, Javier Segura, and Nico M. Temme. Evaluation of the modified Bessel function of the third kind of imaginary orders. Journal of Computational Physics, 175(2):398–411, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968949.

Gervasio:2006:AFS

[GSV06] Paola Gervasio, Fausto Saleri, and Alessandro Veneziani. Algebraic fractional-step schemes with spectral methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 214(1):347–365, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004493.

Gerolymos:2009:VHO

[GSV09] G. A. Gerolymos, D. Sénéchal, and I. Vallet. Very-high-order Weno schemes. Journal of Computational Physics, 228(23):8481–8524, December 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003908.

Guo:2000:LBM

[GSW00] Zhaoli Guo, Baochang Shi, and Nengchao Wang. Lattice BGK model for incompressible Navier–Stokes equation. Journal of Computational Physics, 165(1):288–306, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966166.

Glowinski:2005:MPA

[GT05] Roland Glowinski and Jari Toivanen. A multigrid preconditioner and automatic differentiation for non-equilibrium radiation diffusion problems. Journal of Computational Physics, 207(1):354–374, July 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500032X.

Gamba:2009:SLM

[GT09a] Irene M. Gamba and Sri Harsha Tharkabhushanam. Spectral-Lagrangian methods for collisional models of non-equilibrium statistical states. Journal of Computational Physics, 228 (6):2012–2036, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800613X.

Ganesan:2009:CAL

[GT09b] Sashikumaar Ganesan and Lutz Tobiska. A coupled arbitrary Lagrangian–Eulerian and Lagrangian method for computation of free surface flows with insoluble surfactants. Journal of Computational Physics, 228(8):2859–2873, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000047.

Guddati:2009:PAB

[GT09c] Murthy N. Guddati and Senganal Thirunavukkarasu. Phonon absorbing boundary conditions for molecular dynamics. Journal of Computational Physics, 228(21):8112–8134, November 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004197.

Garbey:2000:PAC

[GTD00] M. Garbey and D. Tromeur-Dervout. A parallel adaptive coupling algorithm for systems of differential equations. Journal of Computational Physics, 161(2):401–427, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965008.

Garbey:2001:PAL

[GTD01] Marc Garbey and Damien Tromeur-Dervout. Parallel algorithms with local Fourier basis. Journal of Computational Physics, 173(2):575–599, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968950.

Gombosi:2002:SMP

[GTD^{+}02] Tamas I. Gombosi, Gábor Tóth, Darren L. De Zeeuw, Kenneth C. Hansen, Konstantin Kabin, and Kenneth G. Powell. Semirelativistic magnetohydrodynamics and physicsbased convergence acceleration. Journal of Computational Physics, 177(1):176–205, March 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970099.

Graves:2008:ESE

[GTMC08] D. T. Graves, D. Trebotich, G. H. Miller, and P. Colella. An efficient solver for the equations of resistive MHD with spatially-varying resistivity. Journal of Computational Physics, 227(10):4797–4804, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000193.

Gallis:2009:CBN

[GTRB09] M. A. Gallis, J. R. Torczynski, D. J. Rader, and G. A. Bird. Convergence behavior of a new DSMC algorithm. Journal of Computational Physics, 228(12):4532–4548, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001429.

Guangwu:2000:LBE

[Gua00] Yan Guangwu. A lattice Boltzmann equation for waves. Journal of Computational Physics, 161(1):61–69, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964866.

Guinot:2002:TLI

[Gui02] Vincent Guinot. The time-line interpolation method for large-time-step Godunov-type schemes. Journal of Computational Physics, 177(2):394–417, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970130.

Guidotti:2003:NFK

[Gui03] Patrick Guidotti. A new first kind boundary integral formulation for the Dirichlet-to-Neumann map in 2D. Journal of Computational Physics, 190(1):325–345, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002778.

Guinot:2005:ATD

[Gui05] Vincent Guinot. An approximate two-dimensional Riemann solver for hyperbolic systems of conservation laws. Journal of Computational Physics, 205(1):292–314, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004681.

Gutman:2000:IMP

[Gut00] S. Gutman. Identification of multilayered particles from scattering data by a clustering method. Journal of Computational Physics, 163(2):529–546, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965847.

Guardone:2002:RLV

[GV02] A. Guardone and L. Vigevano. Roe linearization for the van der Waals gas. Journal of Computational Physics, 175(1):

50–78, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101969153.

Goussis:2006:EIA

[GV06] Dimitris A. Goussis and Mauro Valorani. An efficient iterative algorithm for the approximation of the fast and slow dynamics of stiff systems. Journal of Computational Physics, 214(1):316–346, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004481.

Guardone:2007:FEV

[GV07] A. Guardone and L. Vigevano. Finite element/volume solution to axisymmetric conservation laws. Journal of Computational Physics, 224(2):489–518, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004232.

Galusinski:2008:SCB

[GV08] Cédric Galusinski and Paul Vigneaux. On stability condition for bifluid flows with surface tension: Application to microfluidics. Journal of Computational Physics, 227(12):6140–6164, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001253.

Gilad:2006:FAC

[GvH06] Erez Gilad and Jost von Hardenberg. A fast algorithm for convolution integrals with space and time variant kernels. Journal of Computational Physics, 216(1):326–336, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005541.

Gozadinos:2001:BCP

[GVT01] G. Gozadinos, D. Vender, and M. M. Turner. Boundary conditions and particle loading for the modeling of a semi-infinite plasma. Journal of Computational Physics, 172(1):348–355, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968275.

Ginzburg:2001:TPF

[GW01] I. Ginzburg and G. Wittum. Two-phase flows on interface refined grids modeled with VOF, staggered finite volumes, and spline interpolants. Journal of Computational Physics, 166(2):302–335, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966555.

George:2002:PDG

[GW02] William L. George and James A. Warren. A parallel 3D dendritic growth simulator using the phase-field method. Journal of Computational Physics, 177(2):264–283, April 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970051.

Giraldo:2005:NTB

[GW05] F. X. Giraldo and T. Warburton. A nodal triangle-based spectral element method for the shallow water equations on the sphere. Journal of Computational Physics, 207(1):129–150, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000203.

Garcia:2006:GMI

[GW06] Alejandro L. Garcia and Wolfgang Wagner. Generation of the Maxwellian inflow distribution. Journal of Computational Physics, 217(2):693–708, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000313.

Gomberoff:2007:MOT

[GWF^{+}07] K. Gomberoff, J. Wurtele, A. Friedman, D. P. Grote, and J.-L. Vay. A method for obtaining three-dimensional computational equilibrium of non-neutral plasmas using WARP. Journal of Computational Physics, 225(2):1736–1752, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000848.

Gao:2007:NHE

[GXW07] Bo Gao, Shan-Shu Xu, and Zi-Niu Wu. A note on hybrid Eulerian/Lagrangian computation of compressible inviscid and viscous flows. Journal of Computational Physics, 226(1):1–16, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002446.

Gunter:2005:MHT

[GYKL05] S. Günter, Q. Yu, J. Krüger, and K. Lackner. Modelling of heat transport in magnetised plasmas using non-aligned coordinates. Journal of Computational Physics, 209(1):354–370, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001373.

Ge:2001:HAI

[GZ01] Lixin Ge and Jun Zhang. High accuracy iterative solution of convection diffusion equation with boundary layers on nonuniform grids. Journal of Computational Physics, 171(2):560–578, August 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967944.

Ganapathysubramanian:2007:MDR

[GZ07a] Baskar Ganapathysubramanian and Nicholas Zabaras. Modeling diffusion in random heterogeneous media: Datadriven models, stochastic collocation and the variational multiscale method. Journal of Computational Physics, 226(1):326–353, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700160X.

Ganapathysubramanian:2007:SGC

[GZ07b] Baskar Ganapathysubramanian and Nicholas Zabaras. Sparse grid collocation schemes for stochastic natural convection problems. Journal of Computational Physics, 225(1):652–685, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106006152.

Ganapathysubramanian:2008:NLD

[GZ08] Baskar Ganapathysubramanian and Nicholas Zabaras. A non-linear dimension reduction methodology for generating data-driven stochastic input models. Journal of Computational Physics, 227(13):6612–6637, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001721.

Ganapathysubramanian:2009:SMF

[GZ09] B. Ganapathysubramanian and N. Zabaras. A stochastic multiscale framework for modeling flow through random heterogeneous porous media. Journal of Computational Physics, 228(2):591–618, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005329.

Hu:2002:EAD

[HA02] Fang Q. Hu and Harold L. Atkins. Eigensolution analysis of the discontinuous Galerkin method with nonuniform grids: I. One space dimension. Journal of Computational Physics, 182(2):516–545, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971846.

Hu:2006:MPS

[HA06] X. Y. Hu and N. A. Adams. A multi-phase SPH method for macroscopic and mesoscopic flows. Journal of Computational Physics, 213(2):844–861, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004195.

Hu:2007:IMP

[HA07] X. Y. Hu and N. A. Adams. An incompressible multiphase SPH method. Journal of Computational Physics, 227(1):264–278, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003300.

Hu:2009:CDA

[HA09] X. Y. Hu and N. A. Adams. A constant-density approach for incompressible multi-phase SPH. Journal of Computational Physics, 228(6):2082–2091, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006177.

Haber:2000:FSE

[HAAO00] E. Haber, U. M. Ascher, D. A. Aruliah, and D. W. Oldenburg. Fast simulation of 3D electromagnetic problems using potentials. Journal of Computational Physics, 163(1): 150–171, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100965458.

Haber:2004:MLS

[Hab04] E. Haber. A multilevel, level-set method for optimizing eigenvalues in shape design problems. Journal of Computational Physics, 198(2):518–534, August 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000506.

Hadley:2005:CJI

[Had05] G. Ronald Hadley. A complex Jacobi iterative method for the indefinite Helmholtz equation. Journal of Computational Physics, 203(1):358–370, February 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004012.

Hickel:2006:ALD

[HAD06] Stefan Hickel, Nikolaus A. Adams, and J. Andrzej Domaradzki. An adaptive local deconvolution method for implicit LES. Journal of Computational Physics, 213(1):413– 436, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105003864.

Hagelaar:2007:HNM

[Hag07] G. J. M. Hagelaar. How to normalize Maxwell–Boltzmann electrons in transient plasma models. Journal of Computational Physics, 227(2):871–876, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004329.

Hu:2009:HRS

[HAI09] X. Y. Hu, N. A. Adams, and G. Iaccarino. On the HLLC Riemann solver for interface interaction in compressible multi-fluid flow. Journal of Computational Physics, 228(17):6572–6589, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003192.

Hamacher:2007:AEO

[Ham07] K. Hamacher. Adaptive extremal optimization by detrended fluctuation analysis. Journal of Computational Physics, 227(2):1500–1509, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004214.

Hansbo:2000:CNT

[Han00] Peter Hansbo. A Crank–Nicolson type space–time finite element method for computing on moving meshes. Journal of Computational Physics, 159(2):274–289, April 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964362.

Hankin:2001:EEM

[Han01] Robin K. S. Hankin. The Euler equations for multiphase compressible flow in conservation form: Simulation of shock-bubble interactions. Journal of Computational Physics, 172(2):808–826, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968597.

Henrick:2005:MWE

[HAP05] Andrew K. Henrick, Tariq D. Aslam, and Joseph M. Powers. Mapped weighted essentially non-oscillatory schemes: Achieving optimal order near critical points. Journal of Computational Physics, 207(2):542–567, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000409.

Henrick:2006:SPO

[HAP06] Andrew K. Henrick, Tariq D. Aslam, and Joseph M. Powers. Simulations of pulsating one-dimensional detonations with true fifth order accuracy. Journal of Computational Physics, 213(1):311–329, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003827.

Harlow:2004:FDG

[Har04] Francis H. Harlow. Fluid dynamics in Group T-3 Los Alamos National Laboratory: (LA-UR-03-3852). Journal of Computational Physics, 195(2):414–433, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005692.

Hasan:2005:OBC

[HAS05] Nadeem Hasan, Syed Fahad Anwer, and Sanjeev Sanghi. On the outflow boundary condition for external incompressible flows: a new approach. Journal of Computational Physics, 206(2):661–683, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000100.

Haughton:2008:EECa

[Hau08a] D. M. Haughton. Evaluation of eigenfunctions from compound matrix variables in non-linear elasticity — I. Fourth order systems. Journal of Computational Physics, 227(9):4478–4485, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108000211.

Haughton:2008:EECb

[Hau08b] D. M. Haughton. Evaluation of eigenfunctions from compound matrix variables in non-linear elasticity — II. Sixth order systems. Journal of Computational Physics, 227(20):8960–8967, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108003720.

Howell:2002:FLA

[HB02] B. P. Howell and G. J. Ball. A free-Lagrange augmented Godunov method for the simulation of elastic-plastic solids. Journal of Computational Physics, 175(1):128–167, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969311.

Holzmann:2005:OPC

[HB05a] Markus Holzmann and Bernard Bernu. Optimized periodic 1/r Coulomb potential in two dimensions. Journal of Computational Physics, 206(1):111–121, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005030.

Hyde:2005:FHO

[HB05b] E. McKay Hyde and Oscar P. Bruno. A fast, higher-order solver for scattering by penetrable bodies in three dimensions. Journal of Computational Physics, 202(1):236–261, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400275X.

Hajibeygi:2008:IMF

[HBHJ08] Hadi Hajibeygi, Giuseppe Bonfigli, Marc Andre Hesse, and Patrick Jenny. Iterative multiscale finite-volume method. Journal of Computational Physics, 227(19):8604–8621, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910800332X.

Havu:2009:EIA

[HBHS09] V. Havu, V. Blum, P. Havu, and M. Scheffler. Efficient O(N) integration for all-electron electronic structure calculation using numeric basis functions. Journal of Computational Physics, 228(22):8367–8379, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004458.

Habisreutinger:2007:CAD

[HBLD07] Marc A. Habisreutinger, Roland Bouffanais, Emmanuel Leriche, and Michel O. Deville. A coupled approximate deconvolution and dynamic mixed scale model for large-eddy simulation. Journal of Computational Physics, 224(1):241–266, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910700071X.

Hwang:2005:PNA

[HC05] Feng-Nan Hwang and Xiao-Chuan Cai. A parallel nonlinear additive Schwarz preconditioned inexact Newton algorithm for incompressible Navier–Stokes equations. Journal of Computational Physics, 204(2):666–691, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004322.

Helenbrook:2008:PDT

[HC08] B. T. Helenbrook and G. W. Cowles. Preconditioning for dual-time-stepping simulations of the shallow water equations including Coriolis and bed friction effects. Journal of Computational Physics, 227(9):4425–4440, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800017X.

Henshaw:2009:CGS

[HC09] William D. Henshaw and Kyle K. Chand. A composite grid solver for conjugate heat transfer in fluid-structure systems. Journal of Computational Physics, 228(10):3708–3741, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000667.

Hsiao:2001:AHG

[HCG01] Chao-Tsung Hsiao, Georges Chahine, and Nail Gumerov. Application of a hybrid genetic/Powell algorithm and a boundary element method to electrical impedance tomography. Journal of Computational Physics, 173(2):433–454, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968664.

Haque:2007:TDB

[HD07] Aamer Haque and Gary A. Dilts. Three-dimensional boundary detection for particle methods. Journal of Computational Physics, 226(2):1710–1730, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002586.

Hittinger:2005:STD

[HDBW05] J. A. F. Hittinger, M. R. Dorr, R. L. Berger, and E. A. Williams. Simulating time-dependent energy transfer between crossed laser beams in an expanding plasma. Journal of Computational Physics, 209(2):695–729, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001944.

He:2002:CLB

[HDC02] Xiaoyi He, Gary D. Doolen, and T. Clark. Comparison of the lattice Boltzmann method and the artificial compressibility method for Navier–Stokes equations. Journal of Computational Physics, 179(2):439–451, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970646.

Hoitinga:2008:DML

[HdGKG08] Wijnand Hoitinga, Roel de Groot, Marcel Kwakkel, and Marc Gerritsma. Direct minimization of the least-squares spectral element functional — Part I: Direct solver. Journal of Computational Physics, 227(4):2411–2429, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004718.

Huang:2006:QHE

[HDR^{+}06] C. Huang, V. K. Decyk, C. Ren, M. Zhou, W. Lu, W. B. Mori, J. H. Cooley, T. M. Antonsen, Jr., and T. Katsouleas. QUICKPIC: a highly efficient particle-in-cell code for modeling wakefield acceleration in plasmas. Journal of Computational Physics, 217(2):658–679, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000283.

Hemker:2007:TRS

[HE07] P. W. Hemker and D. Echeverría. A trust-region strategy for manifold-mapping optimization. Journal of Computational Physics, 224(1):464–475, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001519.

Heinrichs:2004:SCS

[Hei04] Wilhelm Heinrichs. Spectral collocation schemes on the unit disc. Journal of Computational Physics, 199(1):66– 86, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104000658.

Heinrichs:2005:ILC

[Hei05] Wilhelm Heinrichs. Improved Lebesgue constants on the triangle. Journal of Computational Physics, 207(2):625–638, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000446.

Helenbrook:2005:PIF

[Hel05] B. T. Helenbrook. Preconditioning for incompressible flows with free-surfaces and two-fluid interfaces. Journal of Computational Physics, 207(1):282–308, July 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000306.

Helsing:2009:FCH

[Hel09a] Johan Helsing. Faster convergence and higher accuracy for the Dirichlet–Neumann map. Journal of Computational Physics, 228(7):2578–2586, April 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800658X.

Helsing:2009:IEM

[Hel09b] Johan Helsing. Integral equation methods for elliptic problems with boundary conditions of mixed type. Journal of Computational Physics, 228(23):8892–8907, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004896.

Hughes:2000:CGM

[HEML00] Thomas J. R. Hughes, Gerald Engel, Luca Mazzei, and Mats G. Larson. The continuous Galerkin method is locally conservative. Journal of Computational Physics, 163(2):467–488, September 20, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910096577X.

Hejranfar:2009:OCS

[HEN09] Kazem Hejranfar, Vahid Esfahanian, and Mehdi Najafi. On the outflow conditions for spectral solution of the viscous blunt-body problem. Journal of Computational Physics, 228(11):3936–3972, June 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000795.

Hermeline:2000:FVM

[Her00] F. Hermeline. A finite volume method for the approximation of diffusion operators on distorted meshes. Journal of Computational Physics, 160(2):481–499, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964660.

Herrmann:2005:ELS

[Her05] M. Herrmann. A Eulerian level set/vortex sheet method for two-phase interface dynamics. Journal of Computational Physics, 203(2):539–571, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003912.

Herrmann:2008:BFR

[Her08] M. Herrmann. A balanced force refined level set grid method for two-phase flows on unstructured flow solver grids. Journal of Computational Physics, 227(4):2674–2706, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004998.

Hermeline:2009:FVM

[Her09] F. Hermeline. A finite volume method for approximating 3D diffusion operators on general meshes. Journal of Computational Physics, 228(16):5763–5786, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002368.

Heuveline:2003:CVL

[Heu03] Vincent Heuveline. On the computation of a very large number of eigenvalues for selfadjoint elliptic operators by means of multigrid methods. Journal of Computational Physics, 184(1):321–337, January 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000438.

Hewett:2003:FMI

[Hew03] Dennis W. Hewett. Fragmentation, merging, and internal dynamics for PIC simulation with finite size particles. Journal of Computational Physics, 189(2):390–426, August 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002250.

Harvie:2000:NVF

[HF00] Dalton J. E. Harvie and David F. Fletcher. A new volume of fluid advection algorithm: The stream scheme. Journal of Computational Physics, 162(1):1–32, July 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965100.

Horne:2001:NCE

[HF01] Richard B. Horne and Mervyn P. Freeman. A new code for electrostatic simulation by numerical integration of the Vlasov and Ampère equations using MacCormack’s method. Journal of Computational Physics, 171(1):182–200, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967816.

Honda:2008:DTS

[HF08a] M. Honda and A. Fukuyama. Dynamic transport simulation code including plasma rotation and radial electric field. Journal of Computational Physics, 227(5):2808–2844, February 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005062.

Husain:2008:ISA

[HF08b] S. Z. Husain and J. M. Floryan. Implicit spectrallyaccurate method for moving boundary problems using immersed boundary conditions concept. Journal of Computational Physics, 227(9):4459–4477, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800020X.

Harari:2001:SDS

[HFO01] Isaac Harari, Leopold P. Franca, and Saulo P. Oliveira. Streamline design of stability parameters for advectiondiffusion problems. Journal of Computational Physics, 171 (1):115–131, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101967774.

Howell:2003:RDM

[HG03] Louis H. Howell and Jeffrey A. Greenough. Radiation diffusion for multi-fluid Eulerian hydrodynamics with adaptive mesh refinement. Journal of Computational Physics, 184(1):53–78, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000153.

Huot:2003:ITS

[HGB^{+}03] F. Huot, A. Ghizzo, P. Bertrand, E. Sonnendrücker, and O. Coulaud. Instability of the time splitting scheme for the one-dimensional and relativistic Vlasov–Maxwell system. Journal of Computational Physics, 185(2):512–531, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000797.

Hadjiconstantinou:2003:SEP

[HGBH03] Nicolas G. Hadjiconstantinou, Alejandro L. Garcia, Martin Z. Bazant, and Gang He. Statistical error in particle simulations of hydrodynamic phenomena. Journal of Computational Physics, 187(1):274–297, May 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000998.

He:2000:ACD

[HGM^{+}00] J.-W. He, R. Glowinski, R. Metcalfe, A. Nordlander, and J. Periaux. Active control and drag optimization for flow past a circular cylinder: I. Oscillatory cylinder rotation. Journal of Computational Physics, 163(1):83–117, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965562.

Hwang:2001:STM

[HGM01] Chi-Ok Hwang, James A. Given, and Michael Mascagni. The simulation-tabulation method for classical diffusion Monte Carlo. Journal of Computational Physics, 174(2):925–946, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969475.

Hubbard:2000:FDS

[HGN00] M. E. Hubbard and P. Garcia-Navarro. Flux difference splitting and the balancing of source terms and flux gradients. Journal of Computational Physics, 165(1):89–125, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966038.

Haschke:2001:STS

[HH01] Heike Haschke and Wilhelm Heinrichs. Splitting techniques with staggered grids for the Navier–Stokes equations in the 2D case. Journal of Computational Physics, 168(1):131–154, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966841.

Hartmann:2002:ADG

[HH02a] Ralf Hartmann and Paul Houston. Adaptive discontinuous Galerkin finite element methods for the compressible Euler equations. Journal of Computational Physics, 183(2):508–532, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102972062.

Henriksen:2002:ASI

[HH02b] Martin Ofstad Henriksen and Jens Holmen. Algebraic splitting for incompressible Navier–Stokes equations. Journal of Computational Physics, 175(2):438–453, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969074.

Hudson:2006:HRS

[HH06] Justin Hudson and David Harris. A high resolution scheme for Eulerian gas-solid two-phase isentropic flow. Journal of Computational Physics, 216(2):494–525, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005620.

Haber:2007:OMM

[HH07a] Eldad Haber and Stefan Heldmann. An octree multigrid method for quasi-static Maxwell’s equations with highly discontinuous coefficients. Journal of Computational Physics, 223(2):783–796, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004748.

Hagstrom:2007:GSH

[HH07b] Thomas Hagstrom and George Hagstrom. Grid stabilization of high-order one-sided differencing I: First-order hyperbolic systems. Journal of Computational Physics, 223(1):316–340, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004396.

Homolle:2007:LVD

[HH07c] Thomas M. M. Homolle and Nicolas G. Hadjiconstantinou. A low-variance deviational simulation Monte Carlo for the Boltzmann equation. Journal of Computational Physics, 226(2):2341–2358, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700318X.

Hartmann:2008:OOI

[HH08] Ralf Hartmann and Paul Houston. An optimal order interior penalty discontinuous Galerkin discretization of the compressible Navier–Stokes equations. Journal of Computational Physics, 227(22):9670–9685, November 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003975.

Hollis:2008:AVL

[HHC08] A. P. Hollis, I. Halliday, and C. M. Care. An accurate and versatile lattice closure scheme for lattice Boltzmann equation fluids under external forces. Journal of Computational Physics, 227(17):8065–8082, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002817.

Hacquin:2001:FCW

[HHCL01] S. Hacquin, S. Heuraux, M. Colin, and G. Leclert. Fast computations of wave propagation in an inhomogeneous plasma by a pulse compression method. Journal of Computational Physics, 174(1):1–11, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968676.

Heemels:2000:SSC

[HHL00] M. W. Heemels, M. H. J. Hagen, and C. P. Lowe. Simulating solid colloidal particles using the lattice-Boltzmann method. Journal of Computational Physics, 164(1):48–61, October 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965641.

Hwang:2004:DSP

[HHM04] Wook Ryol Hwang, Martien A. Hulsen, and Han E. H. Meijer. Direct simulation of particle suspensions in sliding bi-periodic frames. Journal of Computational Physics, 194(2):742–772, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005060.

Hicken:2005:STF

[HHMK05] J. E. Hicken, F. E. Ham, J. Militzer, and M. Koksal. A shift transformation for fully conservative methods: turbulence simulation on complex, unstructured grids. Journal of Computational Physics, 208(2):704–734, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001270.

Hecht:2008:ILT

[HHPW08] Matthew W. Hecht, Darryl D. Holm, Mark R. Petersen, and Beth A. Wingate. Implementation of the LANS-α turbulence model in a primitive equation ocean model. Journal of Computational Physics, 227(11):5691–5716, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001125.

Higdon:2002:TLT

[Hig02] Robert L. Higdon. A two-level time-stepping method for layered ocean circulation models. Journal of Computational Physics, 177(1):59–94, March 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970038.

Higdon:2005:TLT

[Hig05] Robert L. Higdon. A two-level time-stepping method for layered ocean circulation models: further development and testing. Journal of Computational Physics, 206(2):463–504, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104005236.

Hixon:2000:PSS

[Hix00] R. Hixon. Prefactored small-stencil compact schemes. Journal of Computational Physics, 165(2):522–541, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966312.

Helsing:2002:SCM

[HJ02] Johan Helsing and Anders Jonsson. Stress calculations on multiply connected domains. Journal of Computational Physics, 176(2):456–482, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102969962.

Honkkila:2007:HSI

[HJ07] V. Honkkila and P. Janhunen. HLLC solver for ideal relativistic MHD. Journal of Computational Physics, 223(2):643–656, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004669.

Hajibeygi:2009:MFV

[HJ09] Hadi Hajibeygi and Patrick Jenny. Multiscale finite-volume method for parabolic problems arising from compressible multiphase flow in porous media. Journal of Computational Physics, 228(14):5129–5147, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001946.

Hager:2004:PSD

[HJFW04] G. Hager, E. Jeckelmann, H. Fehske, and G. Wellein. Parallelization strategies for density matrix renormalization group algorithms on shared-memory systems. Journal of Computational Physics, 194(2):795–808, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005084.

Heuze:2009:DIH

[HJJ09] Olivier Heuzé, Stéphane Jaouen, and Hervé Jourdren. Dissipative issue of high-order shock capturing schemes with non-convex equations of state. Journal of Computational Physics, 228(3):833–860, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005275.

Heikkinen:2008:FGM

[HJKO08] J. A. Heikkinen, S. J. Janhunen, T. P. Kiviniemi, and F. Ogando. Full f gyrokinetic method for particle simulation of tokamak transport. Journal of Computational Physics, 227(11):5582–5609, May 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001009.

Hong:2009:EMS

[HJL09] Jialin Hong, Shanshan Jiang, and Chun Li. Explicit multisymplectic methods for Klein–Gordon–Schrödinger equations. Journal of Computational Physics, 228(9):3517–3532, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000709.

Huang:2005:TSS

[HJM^{+}05] Zhongyi Huang, Shi Jin, Peter A. Markowich, Christof Sparber, and Chunxiong Zheng. A time-splitting spectral scheme for the Maxwell–Dirac system. Journal of Computational Physics, 208(2):761–789, September 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001300.

Huang:2006:ACS

[HJM06] Jingfang Huang, Jun Jia, and Michael Minion. Accelerating the convergence of spectral deferred correction methods. Journal of Computational Physics, 214(2):633–656, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004663.

Huang:2007:AOK

[HJM07] Jingfang Huang, Jun Jia, and Michael Minion. Arbitrary order Krylov deferred correction methods for differential algebraic equations. Journal of Computational Physics, 221(2):739–760, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003147.

Hagelaar:2000:SFM

[HK00] G. J. M. Hagelaar and G. M. W. Kroesen. Speeding up fluid models for gas discharges by implicit treatment of the electron energy source term. Journal of Computational Physics, 159(1):1–12, March 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964453.

Hui:2001:UCS

[HK01] W. H. Hui and S. Kudriakov. A unified coordinate system for solving the three-dimensional Euler equations. Journal of Computational Physics, 172(1):235–260, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968226.

Hong:2002:WBM

[HK02] Tae-Kyung Hong and B. L. N. Kennett. On a wavelet-based method for the numerical simulation of wave propagation. Journal of Computational Physics, 183(2):577–622, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102972025.

Hall:2004:SAB

[HK04a] William F. Hall and Adour V. Kabakian. A sequence of absorbing boundary conditions for Maxwell’s equations. Journal of Computational Physics, 194(1):140–155, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004637.

Hu:2004:IIM

[HK04b] X. Y. Hu and B. C. Khoo. An interface interaction method for compressible multifluids. Journal of Computational Physics, 198(1):35–64, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000178.

Hu:2004:KEF

[HK04c] X. Y. Hu and B. C. Khoo. Kinetic energy fix for low internal energy flows. Journal of Computational Physics, 193(1):243–259, January 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004297.

Hieber:2005:LPL

[HK05] Simone E. Hieber and Petros Koumoutsakos. A Lagrangian particle level set method. Journal of Computational Physics, 210(1):342–367, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002238.

Ha:2006:ESC

[HK06] Youngsoo Ha and Yong Jung Kim. Explicit solutions to a convection-reaction equation and defects of numerical schemes. Journal of Computational Physics, 220(1):511– 531, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106003457.

Heinrichs:2008:DSL

[HK08a] Wilhelm Heinrichs and Thorsten Kattelans. A direct solver for the least-squares spectral collocation system on rectangular elements for the incompressible Navier–Stokes equations. Journal of Computational Physics, 227(9):4776–4796, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000508.

Hieber:2008:IBM

[HK08b] S. E. Hieber and P. Koumoutsakos. An immersed boundary method for smoothed particle hydrodynamics of selfpropelled swimmers. Journal of Computational Physics, 227(19):8636–8654, October 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003343.

Hieber:2008:LPM

[HK08c] Simone E. Hieber and Petros Koumoutsakos. A Lagrangian particle method for the simulation of linear and nonlinear elastic models of soft tissue. Journal of Computational Physics, 227(21):9195–9215, November 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002714.

Hu:2006:CIM

[HKAH06] X. Y. Hu, B. C. Khoo, N. A. Adams, and F. L. Huang. A conservative interface method for compressible flows. Journal of Computational Physics, 219(2):553–578, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001926.

Heintz:2008:FNM

[HKG08] Alexei Heintz, Piotr Kowalczyk, and Richards Grzhibovskis. Fast numerical method for the Boltzmann equation on non-uniform grids. Journal of Computational Physics, 227(13):6681–6695, June 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001800.

Heikkinen:2001:PSN

[HKKS^{+}01] J. A. Heikkinen, T. P. Kiviniemi, T. Kurki-Suonio, A. G. Peeters, and S. K. Sipilä. Particle simulation of the neoclassical plasmas. Journal of Computational Physics, 173 (2):527–548, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968913.

Hatzky:2007:EGP

[HKM07] R. Hatzky, A. Könies, and A. Mishchenko. Electromagnetic gyrokinetic PIC simulation with an adjustable control variates method. Journal of Computational Physics, 225(1):568–590, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106006085.

Ha:2008:NSD

[HKM08] Youngsoo Ha, Yong-Jung Kim, and Tim G. Myers. On the numerical solution of a driven thin film equation. Journal of Computational Physics, 227(15):7246–7263, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002222.

He:2007:ITD

[HKO07] Lin He, Chiu-Yen Kao, and Stanley Osher. Incorporating topological derivatives into shape derivatives based level set methods. Journal of Computational Physics, 225(1):891–909, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000095.

He:2009:RSI

[HKS09] Lin He, Stefan Kindermann, and Mourad Sini. Reconstruction of shapes and impedance functions using few far-field measurements. Journal of Computational Physics, 228(3):717–730, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800507X.

Horntrop:2001:SMM

[HKV01] David J. Horntrop, Markos A. Katsoulakis, and Dionisios G. Vlachos. Spectral methods for mesoscopic models of pattern formation. Journal of Computational Physics, 173(1):364–390, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968834.

Heinrichs:2001:SST

[HL01] Wilhelm Heinrichs and Birgit I. Loch. Spectral schemes on triangular elements. Journal of Computational Physics, 173 (1):279–301, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968767.

Hanich:2004:TIM

[HL04] L. Hanich and M. Louaked. TVD-interpolating matrix method prediction of flow with heat transfer. Journal of Computational Physics, 193(1):226–242, January 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004285.

Hou:2005:NMS

[HL05] Songming Hou and Xu-Dong Liu. A numerical method for solving variable coefficient elliptic equation with interfaces. Journal of Computational Physics, 202(2):411–445, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400292X.

Hetmaniuk:2006:BSL

[HL06a] U. Hetmaniuk and R. Lehoucq. Basis selection in LOBPCG. Journal of Computational Physics, 218(1):324–332, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000866.

Hong:2006:MSR

[HL06b] Jialin Hong and Chun Li. Multi-symplectic Runge–Kutta methods for nonlinear Dirac equations. Journal of Computational Physics, 211(2):448–472, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002822.

Hellander:2007:HMC

[HL07a] Andreas Hellander and Per Lötstedt. Hybrid method for the chemical master equation. Journal of Computational Physics, 227(1):100–122, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003221.

Hou:2007:CNS

[HL07b] Thomas Y. Hou and Ruo Li. Computing nearly singular solutions using pseudo-spectral methods. Journal of Computational Physics, 226(1):379–397, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001623.

Hua:2007:NSB

[HL07c] Jinsong Hua and Jing Lou. Numerical simulation of bubble rising in viscous liquid. Journal of Computational Physics, 222(2):769–795, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003949.

Lin:2001:SMR

[hLA01] Yi hung Lin and Raymond A. Adomaitis. Simulation and model reduction methods for an RF plasma glow discharge. Journal of Computational Physics, 171(2):731–752, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968081.

Hall:2007:NMH

[HLFB07] David M. Hall, Turab Lookman, Glenn H. Fredrickson, and Sanjoy Banerjee. Numerical method for hydrodynamic transport of inhomogeneous polymer melts. Journal of Computational Physics, 224(2):681–698, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005444.

Hyman:2000:AAQ

[HLKS00] James M. Hyman, Shengtai Li, Patrick Knupp, and Mikhail Shashkov. An algorithm for aligning a quadrilateral grid with internal boundaries. Journal of Computational Physics, 163(1):133–149, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965604.

Hu:2008:ABC

[HLL08] Fang Q. Hu, X. D. Li, and D. K. Lin. Absorbing boundary conditions for nonlinear Euler and Navier–Stokes equations based on the perfectly matched layer technique. Journal of Computational Physics, 227(9):4398–4424, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000168.

Heys:2007:ALS

[HLMM07] J. J. Heys, E. Lee, T. A. Manteuffel, and S. F. McCormick. An alternative least-squares formulation of the Navier–Stokes equations with improved mass conservation. Journal of Computational Physics, 226(1):994–1006, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002185.

Hermeline:2008:FVM

[HLO08] F. Hermeline, S. Layouni, and P. Omnes. A finite volume method for the approximation of Maxwell’s equations in two space dimensions on arbitrary meshes. Journal of Computational Physics, 227(22):9365–9388, November 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002945.

Hou:2006:WCE

[HLRZ06] Thomas Y. Hou, Wuan Luo, Boris Rozovskii, and Hao-Min Zhou. Wiener Chaos expansions and numerical solutions of randomly forced equations of fluid mechanics. Journal of Computational Physics, 216(2):687–706, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000064.

Hou:2001:BIM

[HLS01] T. Y. Hou, J. S. Lowengrub, and M. J. Shelley. Boundary integral methods for multicomponent fluids and multiphase materials. Journal of Computational Physics, 169(2):302–362, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966269.

Ham:2002:CGM

[HLS02a] F. E. Ham, F. S. Lien, and A. B. Strong. A Cartesian grid method with transient anisotropic adaptation. Journal of Computational Physics, 179(2):469–494, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970671.

Ham:2002:FCS

[HLS02b] F. E. Ham, F. S. Lien, and A. B. Strong. A fully conservative second-order finite difference scheme for incompressible flow on nonuniform grids. Journal of Computational Physics, 177(1):117–133, March 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970063.

Hakim:2006:HRW

[HLS06] A. Hakim, J. Loverich, and U. Shumlak. A high resolution wave propagation scheme for ideal two-fluid plasma equations. Journal of Computational Physics, 219(1):418–442, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106001707.

Hwang:2004:NST

[HLWW04] Tsung-Min Hwang, Wen-Wei Lin, Wei-Cheng Wang, and Weichung Wang. Numerical simulation of three dimensional pyramid quantum dot. Journal of Computational Physics, 196(1):208–232, May 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005941. See erratum [HLWW06].

Hwang:2006:ENS

[HLWW06] Tsung-Min Hwang, Wen-Wei Lin, Wei-Cheng Wang, and Weichung Wang. Erratum to “Numerical simulation of three dimensional pyramid quantum dot” [J. Comput. Phys. 196 (2004) 208–232]. Journal of Computational Physics, 215(1):384, June 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500495X. See [HLWW04].

Huang:2006:IEM

[HLX06] Jingfang Huang, Ming-Chih Lai, and Yang Xiang. An integral equation method for epitaxial step-flow growth simulations. Journal of Computational Physics, 216(2):724–743, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000088.

Huang:2009:IPW

[HLY09] Juan-Chen Huang, Herng Lin, and Jaw-Yen Yang. Implicit preconditioned WENO scheme for steady viscous flow computation. Journal of Computational Physics, 228(2):420–438, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005019.

Hunter:2002:RAS

[HLZ02] John K. Hunter, Zhilin Li, and Hongkai Zhao. Reactive autophobic spreading of drops. Journal of Computational Physics, 183(2):335–366, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971688.

Honein:2004:HEC

[HM04] Albert E. Honein and Parviz Moin. Higher entropy conservation and numerical stability of compressible turbulence simulations. Journal of Computational Physics, 201(2):531– 545, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104002414.

Hou:2005:RCI

[HM05] Yucheng Hou and Krishnan Mahesh. A robust, colocated, implicit algorithm for direct numerical simulation of compressible, turbulent flows. Journal of Computational Physics, 205(1):205–221, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004644.

Harlim:2008:MSF

[HM08] J. Harlim and A. J. Majda. Mathematical strategies for filtering complex systems: Regularly spaced sparse observations. Journal of Computational Physics, 227(10):5304–5341, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000806.

Hess:2009:HID

[HM09] S. Hess and F. Mottez. How to improve the diagnosis of kinetic energy in δf PIC codes. Journal of Computational Physics, 228(18):6670–6681, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002812.

Haven:2005:QPT

[HMA05] Kyle Haven, Andrew Majda, and Rafail Abramov. Quantifying predictability through information theory: small sample estimation in a non-Gaussian framework. Journal of Computational Physics, 206(1):334–362, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005169.

Huttunen:2002:CAU

[HMK02] Tomi Huttunen, Peter Monk, and Jari P. Kaipio. Computational aspects of the ultra-weak variational formulation. Journal of Computational Physics, 182(1):27–46, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971482.

Hammond:2002:NMS

[HMM02] E. P. Hammond, K. Mahesh, and P. Moin. A numerical method to simulate radio-frequency plasma discharges. Journal of Computational Physics, 176(2):402–429, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969943.

Huttunen:2007:SME

[HMM07] T. Huttunen, M. Malinen, and P. Monk. Solving Maxwell’s equations using the ultra weak variational formulation. Journal of Computational Physics, 223(2):731–758, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004712.

Hanawa:2008:ISI

[HMM08] Tomoyuki Hanawa, Hayato Mikami, and Tomoaki Matsumoto. Improving shock irregularities based on the characteristics of the MHD equations. Journal of Computational Physics, 227(16):7952–7976, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002702.

Heys:2005:AMH

[HMMO05] J. J. Heys, T. A. Manteuffel, S. F. McCormick, and L. N. Olson. Algebraic multigrid for higher-order finite elements. Journal of Computational Physics, 204(2):520–532, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004267.

Heys:2004:FOS

[HMMR04] J. J. Heys, T. A. Manteuffel, S. F. McCormick, and J. W. Ruge. First-order system least squares (FOSLS) for coupled fluid-elastic problems. Journal of Computational Physics, 195(2):560–575, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005722.

Hagstrom:2008:HOL

[HMOG08] Thomas Hagstrom, Assaf Mar-Or, and Dan Givoli. Highorder local absorbing conditions for the wave equation: Extensions and improvements. Journal of Computational Physics, 227(6):3322–3357, March 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005360.

Heikkola:2007:CMH

[HMPR07] Erkki Heikkola, Sanna Mönkölä, Anssi Pennanen, and Tuomo Rossi. Controllability method for the Helmholtz equation with higher-order discretizations. Journal of Computational Physics, 225(2):1553–1576, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700068X.

Huang:2008:SMM

[HMR08] Weizhang Huang, Jingtang Ma, and Robert D. Russell. A study of moving mesh PDE methods for numerical simulation of blowup in reaction diffusion equations. Journal of Computational Physics, 227(13):6532–6552, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001678.

Hartmann:2008:DEB

[HMS08a] Daniel Hartmann, Matthias Meinke, and Wolfgang Schröder. Differential equation based constrained reinitialization for level set methods. Journal of Computational Physics, 227 (14):6821–6845, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108001964.

See erratum [HMS08b].

Hartmann:2008:EDE

[HMS08b] Daniel Hartmann, Matthias Meinke, and Wolfgang Schröder. Erratum to “Differential Equation Based Constrained Reinitialization for Level Set Methods” [J. Comput. Phys. 227 (2008) 6821–6845]. Journal of Computational Physics, 227 (22):9696, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800418X.

See [HMS08a].

Homescu:2003:OCF

[HN03] Chris Homescu and I. M. Navon. Optimal control of flow with discontinuities. Journal of Computational Physics, 187 (2):660–682, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103001542.

Haselbacher:2007:ERP

[HNF07] A. Haselbacher, F. M. Najjar, and J. P. Ferry. An efficient and robust particle-localization algorithm for unstructured grids. Journal of Computational Physics, 225(2):2198–2213, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700126X.

Holdych:2004:TEA

[HNGB04] David J. Holdych, David R. Noble, John G. Georgiadis, and Richard O. Buckius. Truncation error analysis of lattice Boltzmann methods. Journal of Computational Physics, 193(2):595–619, January 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004364.

Hughes:2003:CSS

[HO03] Thomas J. R. Hughes and Assad A. Oberai. Calculation of shear stresses in the Fourier–Galerkin formulation of turbulent channel flows: projection, the Dirichlet filter and conservation. Journal of Computational Physics, 188(1):281–295, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103001670.

Helzel:2006:MSS

[HO06] Christiane Helzel and Felix Otto. Multiscale simulations for suspensions of rod-like molecules. Journal of Computational Physics, 216(1):52–75, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005425.

Helsing:2008:CSE

[HO08a] Johan Helsing and Rikard Ojala. Corner singularities for elliptic problems: Integral equations, graded meshes, quadrature, and compressed inverse preconditioning. Journal of Computational Physics, 227(20):8820–8840, October 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003471.

Helsing:2008:ELP

[HO08b] Johan Helsing and Rikard Ojala. On the evaluation of layer potentials close to their sources. Journal of Computational Physics, 227(5):2899–2921, February 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005190.

Hoffman:2004:IFC

[Hof04] Gary G. Hoffman. An integral for FHNC calculations. Journal of Computational Physics, 194(2):659–676, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300500X.

Hohage:2006:FNS

[Hoh06] Thorsten Hohage. Fast numerical solution of the electromagnetic medium scattering problem and applications to the inverse problem. Journal of Computational Physics, 214(1):224– 238, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004328.

Horritt:2002:SMD

[Hor02] M. S. Horritt. Stochastic modelling of 1-D shallow water flows over uncertain topography. Journal of Computational Physics, 180(1):327–338, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297097X.

Horntrop:2006:MSO

[Hor06] David J. Horntrop. Mesoscopic simulation of Ostwald ripening. Journal of Computational Physics, 218(1):429–441, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000921.

Hao:2004:NMT

[HP04a] Yue Hao and Andrea Prosperetti. A numerical method for three-dimensional gas-liquid flow computations. Journal of Computational Physics, 196(1):126–144, May 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005904.

Hill:2004:HTC

[HP04b] D. J. Hill and D. I. Pullin. Hybrid tuned centerdifference-WENO method for large eddy simulations in the presence of strong shocks. Journal of Computational Physics, 194(2):435–450, March 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300490X.

Hay:2009:VPS

[HPD09] A. Hay, D. Pelletier, and R. Di Caro. Verified predictions of shape sensitivities in wall-bounded turbulent flows by an adaptive finite-element method. Journal of Computational Physics, 228(12):4510–4531, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001417.

Ha:2006:EER

[HPS06a] Taeyoung Ha, Sukjoon Pyun, and Changsoo Shin. Efficient electric resistivity inversion using adjoint state of mixed finite-element method for Poisson’s equation. Journal of Computational Physics, 214(1):171–186, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004286.

Huang:2006:ENS

[HPS^{+}06b] Kai Huang, George Papanicolaou, Knut Solna, Chrysoula Tsogka, and Hongkai Zhao. Efficient numerical simulation for long range wave propagation. Journal of Computational Physics, 215(2):448–464, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004985.

Hu:2001:DNS

[HPZ01] Howard H. Hu, N. A. Patankar, and M. Y. Zhu. Direct numerical simulations of fluid-solid systems using the arbitrary Lagrangian–Eulerian technique. Journal of Computational Physics, 169(2):427–462, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965926.

Hindmarsh:2001:UOS

[HR01] Alan C. Hindmarsh and Mark D. Rotter. Using an ODE solver for a class of integro-differential systems. Journal of Computational Physics, 168(2):267–285, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966804.

Hundsdorfer:2007:IEL

[HR07] Willem Hundsdorfer and Steven J. Ruuth. IMEX extensions of linear multistep methods with general monotonicity and boundedness properties. Journal of Computational Physics, 225(2):2016–2042, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001003.

Hallerod:2008:EML

[HR08] Tomas Halleröd and Thomas Rylander. Electric and magnetic losses modeled by a stable hybrid with explicit-implicit time-stepping for Maxwell’s equations. Journal of Computational Physics, 227(9):4499–4511, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000314.

Tsai:2002:RAC

[hRT02] Yen hsi Richard Tsai. Rapid and accurate computation of the distance function using grids. Journal of Computational Physics, 178(1):175–195, May 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970282.

Huberson:2008:VPM

[HRV08] Serge Huberson, Elie Rivoalen, and Spyros Voutsinas. Vortex particle methods in aeroacoustic calculations. Journal of Computational Physics, 227(21):9216–9240, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108003124.

Henshaw:2003:ANS

[HS03a] William D. Henshaw and Donald W. Schwendeman. An adaptive numerical scheme for high-speed reactive flow on overlapping grids. Journal of Computational Physics, 191 (2):420–447, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003231.

Huang:2003:VMA

[HS03b] Weizhang Huang and Weiwei Sun. Variational mesh adaptation II: error estimates and monitor functions. Journal of Computational Physics, 184(2):619–648, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000402.

Hu:2004:CTM

[HS04] Yanhong Hu and Susan B. Sinnott. Constant temperature molecular dynamics simulations of energetic particle-solid collisions: comparison of temperature control methods. Journal of Computational Physics, 200(1):251–266, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001664.

Henshaw:2006:MOG

[HS06] William D. Henshaw and Donald W. Schwendeman. Moving overlapping grids with adaptive mesh refinement for highspeed reactive and non-reactive flow. Journal of Computational Physics, 216(2):744–779, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600009X.

Ha:2007:MIR

[HS07a] Taeyoung Ha and Changsoo Shin. Magnetotelluric inversion via reverse time migration algorithm of seismic data. Journal of Computational Physics, 225(1):237–262, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005869.

Hur:2007:TPM

[HS07b] Min Sup Hur and Hyyong Suk. Test particle method for incorporation of the kinetic effects into the envelope simulations of Raman backscattering. Journal of Computational Physics, 226(2):2133–2145, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003002.

Henshaw:2008:PCT

[HS08a] William D. Henshaw and Donald W. Schwendeman. Parallel computation of three-dimensional flows using overlapping grids with adaptive mesh refinement. Journal of Computational Physics, 227(16):7469–7502, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800243X.

Hou:2008:ESI

[HS08b] Thomas Y. Hou and Zuoqiang Shi. An efficient semi-implicit immersed boundary method for the Navier–Stokes equations. Journal of Computational Physics, 227(20):8968–8991, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108003732.

Hou:2008:RSE

[HS08c] Thomas Y. Hou and Zuoqiang Shi. Removing the stiffness of elastic force from the immersed boundary method for the 2D Stokes equations. Journal of Computational Physics, 227(21):9138–9169, November 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800137X.

Hague:2009:MFB

[HS09a] C. H. Hague and C. Swan. A multiple flux boundary element method applied to the description of surface water waves. Journal of Computational Physics, 228(14):5111–5128, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001934.

Heuveline:2009:SOT

[HS09b] Vincent Heuveline and Frank Strauß. Shape optimization towards stability in constrained hydrodynamic systems. Journal of Computational Physics, 228(4):938–951, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003380.

Hazra:2005:ASO

[HSBG05] S. B. Hazra, V. Schulz, J. Brezillon, and N. R. Gauger. Aerodynamic shape optimization using simultaneous pseudotimestepping. Journal of Computational Physics, 204(1): 46–64, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104004061.

Huang:2009:FRC

[HSC09] Mei-Jiau Huang, Huan-Xun Su, and Li-Chieh Chen. A fast resurrected core-spreading vortex method with no-slip boundary conditions. Journal of Computational Physics, 228(6):1916–1931, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006086.

Hariharan:2000:PTM

[HSK00] S. I. Hariharan, J. R. Scott, and K. L. Kreider. A potentialtheoretic method for far-field sound radiation calculations. Journal of Computational Physics, 164(1):143–164, October 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965914.

Hua:2008:NSB

[HSL08] Jinsong Hua, Jan F. Stene, and Ping Lin. Numerical simulation of 3D bubbles rising in viscous liquids using a front tracking method. Journal of Computational Physics, 227(6):3358–3382, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005530.

Hariharan:2003:LTP

[HSQ03] S. I. Hariharan, Scott Sawyer, and D. Dane Quinn. A Laplace transform/potential-theoretic method for acoustic propagation in subsonic flows. Journal of Computational Physics, 185(1):252–270, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000566.

Huang:2007:SFF

[HSS07] Wei-Xi Huang, Soo Jai Shin, and Hyung Jin Sung. Simulation of flexible filaments in a uniform flow by the immersed boundary method. Journal of Computational Physics, 226(2):2206–2228, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003051.

Huang:2009:IBM

[HST09] Huaxiong Huang, Kazuyasu Sugiyama, and Shu Takagi. An immersed boundary method for restricted diffusion with permeable interfaces. Journal of Computational Physics, 228(15):5317–5322, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002265.

Hou:2007:MMS

[HSW07] T. Y. Hou, V. G. Stredie, and T. Y. Wu. Mathematical modeling and simulation of aquatic and aerial animal locomotion. Journal of Computational Physics, 225(2):1603–1631, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000733.

Hou:2004:ILG

[HSZ04] Songming Hou, Knut Solna, and Hongkai Zhao. Imaging of location and geometry for extended targets using the response matrix. Journal of Computational Physics, 199(1): 317–338, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104000865.

Hixon:2000:CIM

[HT00a] R. Hixon and E. Turkel. Compact implicit MacCormacktype schemes with high accuracy. Journal of Computational Physics, 158(1):51–70, February 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964069. See erratum [HT00b].

Hixon:2000:ECI

[HT00b] R. Hixon and E. Turkel. Erratum: “Compact Implicit MacCormack-Type Schemes with High Accuracy”: Volume 158, Number 1 (2000), pages 51–70. Journal of Computational Physics, 163(2):547, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965598. See [HT00a].

Huang:2003:PNM

[HT03] Huaxiong Huang and Shu Takagi. PHYSALIS: a new method for particle flow simulation. Part III: convergence analysis of two-dimensional flows. Journal of Computational Physics, 189(2):493–511, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002304.

Han:2007:AMM

[HT07] Jianqiang Han and Huazhong Tang. An adaptive moving mesh method for two-dimensional ideal magnetohydrodynamics. Journal of Computational Physics, 220(2):791–812, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002609.

Hu:2001:SPM

[Hu01] Fang Q. Hu. A stable, perfectly matched layer for linearized Euler equations in unsplit physical variables. Journal of Computational Physics, 173(2):455–480, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968871.

Hu:2005:PML

[Hu05] Fang Q. Hu. A perfectly matched layer absorbing boundary condition for linearized Euler equations with a nonuniform mean flow. Journal of Computational Physics, 208(2):469–492, September 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001117.

Huang:2001:PAF

[Hua01a] Weizhang Huang. Practical aspects of formulation and solution of moving mesh partial differential equations. Journal of Computational Physics, 171(2):753–775, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968093.

Huang:2001:VMA

[Hua01b] Weizhang Huang. Variational mesh adaptation: Isotropy and equidistribution. Journal of Computational Physics, 174 (2):903–924, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101969451.

Huang:2005:MTA

[Hua05] Weizhang Huang. Metric tensors for anisotropic mesh generation. Journal of Computational Physics, 204(2):633–665, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004310.

Huang:2007:PDT

[Hua07] D. G. Huang. Preconditioned dual-time procedures and its application to simulating the flow with cavitations. Journal of Computational Physics, 223(2):685–689, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004682.

Hubbard:2007:NOT

[Hub07] Matthew Hubbard. Non-oscillatory third order fluctuation splitting schemes for steady scalar conservation laws. Journal of Computational Physics, 222(2):740–768, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003937.

Hubbard:2008:DFD

[Hub08] Matthew Hubbard. Discontinuous fluctuation distribution. Journal of Computational Physics, 227(24):10125–10147, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004452.

Humphries:2005:MIE

[Hum05] Stanley Humphries, Jr. Modeling ion extraction from a freeplasma surface with a flexible conformal mesh. Journal of Computational Physics, 204(2):587–597, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004292.

Hunke:2001:VPS

[Hun01] Elizabeth C. Hunke. Viscous-plastic sea ice dynamics with the EVP model: Linearization issues. Journal of Computational Physics, 170(1):18–38, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967105.

Haselbacher:2003:CDF

[HV03] Andreas Haselbacher and Oleg V. Vasilyev. Commutative discrete filtering on unstructured grids based on least-squares techniques. Journal of Computational Physics, 187(1):197– 211, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000950.

Houzeaux:2009:MPF

[HVAC09] G. Houzeaux, M. Vázquez, R. Aubry, and J. M. Cela. A massively parallel fractional step solver for incompressible flows. Journal of Computational Physics, 228 (17):6316–6332, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002769.

Hart:2005:AGF

[HvHHS05] A. Hart, G. M. von Hippel, R. R. Horgan, and L. C. Storoni. Automatically generating Feynman rules for improved lattice field theories. Journal of Computational Physics, 209(1):340–353, October 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500135X.

Hesthaven:2002:NHO

[HW02] J. S. Hesthaven and T. Warburton. Nodal high-order methods on unstructured grids: I. Time-domain solution of Maxwell’s equations. Journal of Computational Physics, 181(1):186–221, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971184.

Helsing:2005:LED

[HW05] Johan Helsing and Eddie Wadbro. Laplace’s equation and the Dirichlet–Neumann map: a new mode for Mikhlin’s method. Journal of Computational Physics, 202(2):391–410, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002918.

Hardt:2008:EMI

[HW08] S. Hardt and F. Wondra. Evaporation model for interfacial flows based on a continuum-field representation of the source terms. Journal of Computational Physics, 227(11):5871–5895, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001228.

Hwang:2003:USN

[Hwa03] Yao-Hsin Hwang. Upwind scheme for non-hyperbolic systems. Journal of Computational Physics, 192(2):643–676, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004121.

Harris:2008:EQF

[HWL08] Rob Harris, Z. J. Wang, and Yen Liu. Efficient quadraturefree high-order spectral volume method on unstructured grids: Theory and 2D implementation. Journal of Computational Physics, 227(3):1620–1642, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004160.

Hwang:2007:NST

[HWW07] Tsung-Min Hwang, Wei-Cheng Wang, and Weichung Wang. Numerical schemes for three-dimensional irregular shape quantum dots over curvilinear coordinate systems. Journal of Computational Physics, 226(1):754–773, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002057.

Hu:2009:SEF

[HWWL09] Z. Hu, S. M. Wise, C. Wang, and J. S. Lowengrub. Stable and efficient finite-difference nonlinear-multigrid schemes for the phase field crystal equation. Journal of Computational Physics, 228(15):5323–5339, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002058.

Hongbin:2005:CSP

[HX05] Jin Hongbin and Ding Xin. On criterions for smoothed particle hydrodynamics kernels in stable field. Journal of Computational Physics, 202(2):699–709, January 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003122.

Hsieh:2009:BSL

[HY09] Po-Wen Hsieh and Suh-Yuh Yang. A bubble-stabilized leastsquares finite element method for steady MHD duct flow problems at high Hartmann numbers. Journal of Computational Physics, 228(22):8301–8320, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004410. See erratum [HY11].

Hsieh:2011:EBS

[HY11] Po-Wen Hsieh and Suh-Yuh Yang. Erratum to “A bubblestabilized least-squares finite element method for steady MHD duct flow problems at high Hartmann numbers” [J. Comput. Phys. 228 (2009) 8301–8320]. Journal of Computational Physics, 230(2):474–475, January 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005474. See [HY09].

Hong:2007:MSR

[HyLL07] Jialin Hong, Xiao yan Liu, and Chun Li. Multi-symplectic Runge–Kutta–Nyström methods for nonlinear Schrödinger equations with variable coefficients. Journal of Computational Physics, 226(2):1968–1984, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002720.

Heerlein:2002:NLD

[HZ02] Claus Heerlein and Günter Zwicknagel. Nonlinear Landau damping in spherically symmetric Vlasov Poisson systems. Journal of Computational Physics, 180(2):497–505, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970968.

Hansen:2007:USM

[HZ07a] Glen Hansen and Andrew Zardecki. Unstructured surface mesh adaptation using the Laplace–Beltrami target metric approach. Journal of Computational Physics, 225(1):165–182, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005821.

Hinze:2007:OCF

[HZ07b] Michael Hinze and Stefan Ziegenbalg. Optimal control of the free boundary in a two-phase Stefan problem. Journal of Computational Physics, 223(2):657–684, May 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004670.

Han:2008:SLA

[HZ08] Houde Han and Zhiwen Zhang. Split local absorbing conditions for one-dimensional nonlinear Klein–Gordon equation on unbounded domain. Journal of Computational Physics, 227(20):8992–9004, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003823.

Hansen:2004:FEM

[HZGB04] Glen Hansen, Andrew Zardecki, Doran Greening, and Randy Bos. A finite element method for unstructured grid smoothing. Journal of Computational Physics, 194(2):611–631, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004984.

Hansen:2005:FEM

[HZGB05] Glen Hansen, Andrew Zardecki, Doran Greening, and Randy Bos. A finite element method for three-dimensional unstructured grid smoothing. Journal of Computational Physics, 202(1):281–297, January 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002864.

Imai:2006:ASI

[IA06a] Yohsuke Imai and Takayuki Aoki. Accuracy study of the IDO scheme by Fourier analysis. Journal of Computational Physics, 217(2):453–472, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000179.

Imai:2006:SCB

[IA06b] Yohsuke Imai and Takayuki Aoki. Stable coupling between vector and scalar variables for the IDO scheme on collocated grids. Journal of Computational Physics, 215(1):81–97, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004808.

Imai:2008:CFI

[IAT08] Yohsuke Imai, Takayuki Aoki, and Kenji Takizawa. Conservative form of interpolated differential operator scheme for compressible and incompressible fluid dynamics. Journal of Computational Physics, 227(4):2263–2285, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700530X.

Inoue:2004:MSM

[ICO04] Yasuhiro Inoue, Yu Chen, and Hirotada Ohashi. A mesoscopic simulation model for immiscible multiphase fluids. Journal of Computational Physics, 201(1):191–203, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002074.

Ikeda:2004:MSE

[ID04] Tomoaki Ikeda and Paul A. Durbin. Mesh stretch effects on convection in flow simulations. Journal of Computational Physics, 199(1):110–125, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000786.

Iskakov:2004:IDF

[IDD04] A. B. Iskakov, S. Descombes, and E. Dormy. An integrodifferential formulation for magnetic induction in bounded domains: boundary element-finite volume method. Journal of Computational Physics, 197(2):540–554, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300651X.

Izquierdo:2009:OPL

[IF09] Salvador Izquierdo and Norberto Fueyo. Optimal preconditioning of lattice Boltzmann methods. Journal of Computational Physics, 228(17):6479–6495, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900312X.

Iollo:2001:AOM

[IFZ01] A. Iollo, M. Ferlauto, and L. Zannetti. An aerodynamic optimization method based on the inverse problem adjoint equations. Journal of Computational Physics, 173(1):87–115, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968457.

Illingworth:2005:NSD

[IG05] T. C. Illingworth and I. O. Golosnoy. Numerical solutions of diffusion-controlled moving boundary problems which conserve solute. Journal of Computational Physics, 209(1):207– 225, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105000859.

Izaguirre:2004:SHM

[IH04] Jesús A. Izaguirre and Scott S. Hampton. Shadow hybrid Monte Carlo: an efficient propagator in phase space of macromolecules. Journal of Computational Physics, 200 (2):581–604, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104001809.

Iskandarani:2003:TDS

[IHL03] M. Iskandarani, D. B. Haidvogel, and J. C. Levin. A three-dimensional spectral element model for the solution of the hydrostatic primitive equations. Journal of Computational Physics, 186(2):397–425, April 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000251.

Inutsuka:2002:RSP

[iI02] Shu ichiro Inutsuka. Reformulation of smoothed particle hydrodynamics with Riemann solver. Journal of Computational Physics, 179(1):238–267, June 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970531.

Idomura:2007:NCG

[IITV07] Yasuhiro Idomura, Masato Ida, Shinji Tokuda, and Laurent Villard. New conservative gyrokinetic full-f Vlasov code and its comparison to gyrokinetic δf particle-in-cell code. Journal of Computational Physics, 226(1):244–262, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001556.

Ingber:2001:GIG

[IK01] Marc S. Ingber and Steven N. Kempka. A Galerkin implementation of the generalized Helmholtz decomposition for vorticity formulations. Journal of Computational Physics, 169(1):215–237, May 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967245.

Ikeno:2007:FDI

[IK07] Tsutomu Ikeno and Takeo Kajishima. Finite-difference immersed boundary method consistent with wall conditions for incompressible turbulent flow simulations. Journal of Computational Physics, 226(2):1485–1508, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002458.

Ijaz:2008:NPB

[IKL^{+}08] Umer Zeeshan Ijaz, Anil Kumar Khambampati, Jeong Seong Lee, Sin Kim, and Kyung Youn Kim. Nonstationary phase boundary estimation in electrical impedance tomography using unscented Kalman filter. Journal of Computational Physics, 227(15):7089–7112, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000132.

Islas:2001:GIN

[IKS01] A. L. Islas, D. A. Karpeev, and C. M. Schober. Geometric integrators for the nonlinear Schrödinger equation. Journal of Computational Physics, 173(1):116–148, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968548.

Imadera:2009:NMS

[IKS^{+}09] Kenji Imadera, Yasuaki Kishimoto, Daisuke Saito, Jiquan Li, and Takayuki Utsumi. A numerical method for solving the Vlasov–Poisson equation based on the conservative IDO scheme. Journal of Computational Physics, 228(23):8919–8943, December 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004938.

Ito:2009:WCA

[ILL09] Kazufumi Ito, Ming-Chih Lai, and Zhilin Li. A wellconditioned augmented system for solving Navier–Stokes equations in irregular domains. Journal of Computational Physics, 228(7):2616–2628, April 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800661X.

Im:2005:NAM

[IM05] Kichang Im and Yoshihiro Mochimaru. Numerical analysis on magnetic levitation of liquid metals, using a spectral finite difference scheme. Journal of Computational Physics, 203(1):112–128, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003201.

Ishida:2007:RFO

[IM07] Hideshi Ishida and Kazunari Momose. Revaluation of the firstorder upwind difference scheme to solve coarse-grained master equations. Journal of Computational Physics, 221(1):106–121, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002774.

Inamuro:2004:LBM

[IOTK04] T. Inamuro, T. Ogata, S. Tajima, and N. Konishi. A lattice Boltzmann method for incompressible two-phase flows with large density differences. Journal of Computational Physics, 198(2):628–644, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000567.

Ito:2008:HOC

[IQ08] Kazufumi Ito and Zhonghua Qiao. A high order compact MAC finite difference scheme for the Stokes equations: Augmented variable approach. Journal of Computational Physics, 227(17):8177–8190, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003070.

Ito:2008:DDS

[IQT08] Kazufumi Ito, Zhonghua Qiao, and Jari Toivanen. A domain decomposition solver for acoustic scattering by elastic objects in layered media. Journal of Computational Physics, 227(19):8685–8698, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003379.

Ismail:2009:AEC

[IR09] Farzad Ismail and Philip L. Roe. Affordable, entropyconsistent Euler flux functions II: Entropy production at shocks. Journal of Computational Physics, 228(15):5410– 5436, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109002113.

Islas:2004:PPS

[IS04] A. L. Islas and C. M. Schober. On the preservation of phase space structure under multisymplectic discretization. Journal of Computational Physics, 197(2):585–609, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006533.

Imamura:2005:ASS

[ISNY05] Taro Imamura, Kojiro Suzuki, Takashi Nakamura, and Masahiro Yoshida. Acceleration of steady-state lattice Boltzmann simulations on non-uniform mesh using local time step method. Journal of Computational Physics, 202(2):645–663, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104003006.

Ii:2007:CMM

[IX07] S. Ii and F. Xiao. CIP/multi-moment finite volume method for Euler equations: a semi-Lagrangian characteristic formulation. Journal of Computational Physics, 222(2):849–871, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004189.

Ii:2009:HOM

[IX09] Satoshi Ii and Feng Xiao. High order multi-moment constrained finite volume method. Part I: Basic formulation. Journal of Computational Physics, 228(10):3669–3707, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000655.

Inamuro:2002:LBM

[IYI^{+}02] Takaji Inamuro, Masato Yoshino, Hiroshi Inoue, Riki Mizuno, and Fumimaru Ogino. A lattice Boltzmann method for a binary miscible fluid mixture and its application to a heat-transfer problem. Journal of Computational Physics, 179(1):201–215, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970518.

Janssen:2008:BIM

[JA08] P. J. A. Janssen and P. D. Anderson. A boundary-integral model for drop deformation between two parallel plates with non-unit viscosity ratio drops. Journal of Computational Physics, 227(20):8807–8819, October 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800346X.

Jocksch:2005:ASC

[JAK05] A. Jocksch, N. A. Adams, and L. Kleiser. An asymptotically stable compact upwind-biased finite-difference scheme for hyperbolic systems. Journal of Computational Physics, 208(2):435–454, September 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000938.

Janhunen:2000:PCM

[Jan00] P. Janhunen. A positive conservative method for magnetohydrodynamics based on HLL and Roe methods. Journal of Computational Physics, 160(2):649–661, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964799.

Janssen:2008:POW

[Jan08] Peter A. E. M. Janssen. Progress in ocean wave forecasting. Journal of Computational Physics, 227(7):3572–3594, March 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107001659.

Jaouen:2007:PLM

[Jao07] Stéphane Jaouen. A purely Lagrangian method for computing linearly-perturbed flows in spherical geometry. Journal of Computational Physics, 225(1):464–490, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006012.

Jardin:2004:TFE

[Jar04] S. C. Jardin. A triangular finite element with first-derivative continuity applied to fusion MHD applications. Journal of Computational Physics, 200(1):133–152, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001366.

Jardin:2007:HOI

[JBF07] S. C. Jardin, J. Breslau, and N. Ferraro. A high-order implicit finite element method for integrating the two-fluid magnetohydrodynamic equations in two dimensions. Journal of Computational Physics, 226(2):2146–2174, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003038.

Jardin:2008:DPG

[JBHK08] S. C. Jardin, G. Bateman, G. W. Hammett, and L. P. Ku. On 1D diffusion problems with a gradient-dependent diffusion coefficient. Journal of Computational Physics, 227(20):8769–8775, October 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003616.

Jaeger:2002:FTA

[JC02] M. Jaeger and M. Carin. The front-tracking ALE method: Application to a model of the freezing of cell suspensions. Journal of Computational Physics, 179(2):704–735, July 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970841.

Jin:2006:BKM

[JC06a] Bangti Jin and Wen Chen. Boundary knot method based on geodesic distance for anisotropic problems. Journal of Computational Physics, 215(2):614–629, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005073.

Johnsen:2006:IWS

[JC06b] Eric Johnsen and Tim Colonius. Implementation of WENO schemes in compressible multicomponent flow problems. Journal of Computational Physics, 219(2):715–732, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002014.

Jung:2007:VLS

[JCT07] Y. Jung, K. T. Chu, and S. Torquato. A variational level set approach for surface area minimization of triply-periodic surfaces. Journal of Computational Physics, 223(2):711–730, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004700.

Jenny:2004:ENM

[JD04] Mathieu Jenny and Jan Dusek. Efficient numerical method for the direct numerical simulation of the flow past a single light moving spherical body in transitional regimes. Journal of Computational Physics, 194(1):215–232, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004674.

Jacobs:2009:HOW

[JD09] Gustaaf B. Jacobs and Wai-Sun Don. A high-order WENO–Z finite difference based particle-source-in-cell method for computation of particle-laden flows with shocks. Journal of Computational Physics, 228(5):1365–1379, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005585.

Jolley:2009:MTH

[JG09] Kenny Jolley and Simon P. A. Gill. Modelling transient heat conduction in solids at multiple length and time scales: a coupled non-equilibrium molecular dynamics/continuum approach. Journal of Computational Physics, 228(19):7412–7425, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003672.

Jacobs:2006:HON

[JH06] G. B. Jacobs and J. S. Hesthaven. High-order nodal discontinuous Galerkin particle-in-cell method on unstructured grids. Journal of Computational Physics, 214(1):96–121, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004250.

Jia:2008:KDC

[JH08] Jun Jia and Jingfang Huang. Krylov deferred correction accelerated method of lines transpose for parabolic problems. Journal of Computational Physics, 227(3):1739–1753, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004287.

Jouvet:2009:NSR

[JHB^{+}09] Guillaume Jouvet, Matthias Huss, Heinz Blatter, Marco Picasso, and Jacques Rappaz. Numerical simulation of Rhonegletscher from 1874 to 2100. Journal of Computational Physics, 228(17):6426–6439, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900285X.

Jiang:2007:SND

[JHSZ07] Jun Jiang, Yunqing Huang, Shi Shu, and Shi Zeng. Some new discretization and adaptation and multigrid methods for 2-D 3-T diffusion equations. Journal of Computational Physics, 224(1):168–181, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000150.

Jing:2009:LTV

[JHZ^{+}09] Yan-Fei Jing, Ting-Zhu Huang, Yong Zhang, Liang Li, Guang-Hui Cheng, Zhi-Gang Ren, Yong Duan, Tomohiro Sogabe, and Bruno Carpentieri. Lanczos-type variants of the COCR method for complex nonsymmetric linear systems. Journal of Computational Physics, 228(17):6376–6394, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002800.

Jiao:2007:FOU

[Jia07] Xiangmin Jiao. Face offsetting: a unified approach for explicit moving interfaces. Journal of Computational Physics, 220(2):612–625, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600252X.

Jaiman:2006:CLT

[JJGL06] R. K. Jaiman, X. Jiao, P. H. Geubelle, and E. Loth. Conservative load transfer along curved fluid-solid interface with non-matching meshes. Journal of Computational Physics, 218(1):372–397, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000891. See erratum [JJGL07].

Jaiman:2007:ECL

[JJGL07] R. K. Jaiman, X. Jiao, P. H. Geubelle, and E. Loth. Erratum to “Conservative load transfer along curved fluidsolid interface with non-matching meshes” [J. Comput. Phys. 218(1) (2006) 372–397]. Journal of Computational Physics, 224(2):1307, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005924. See [JJGL06].

Jawahar:2000:HRP

[JK00] P. Jawahar and Hemant Kamath. A high-resolution procedure for Euler and Navier–Stokes computations on unstructured grids. Journal of Computational Physics, 164(1):165–203, October 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965963.

Jeong:2002:OCS

[JK02] Ja Hoon Jeong and In Seok Kang. Optimization of the crystal surface temperature distribution in the single-crystal growth process by the Czochralski method. Journal of Computational Physics, 177(2):284–312, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970117.

Junk:2005:AAL

[JKL05] Michael Junk, Axel Klar, and Li-Shi Luo. Asymptotic analysis of the lattice Boltzmann equation. Journal of Computational Physics, 210(2):676–704, December 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002573.

Johnston:2002:FDS

[JL02] Hans Johnston and Jian-Guo Liu. Finite difference schemes for incompressible flow based on local pressure boundary conditions. Journal of Computational Physics, 180(1):120–154, July 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970798.

James:2004:SCV

[JL04a] Ashley J. James and John Lowengrub. A surfactantconserving volume-of-fluid method for interfacial flows with insoluble surfactant. Journal of Computational Physics, 201(2):685–722, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400261X.

Johnston:2004:ASE

[JL04b] Hans Johnston and Jian-Guo Liu. Accurate, stable and efficient Navier–Stokes solvers based on explicit treatment of the pressure term. Journal of Computational Physics, 199(1):221–259, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400083X.

Jenny:2009:MCW

[JL09] Patrick Jenny and Ivan Lunati. Modeling complex wells with the multi-scale finite-volume method. Journal of Computational Physics, 228(3):687–702, February 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005020.

Jamet:2001:SGM

[JLCD01] D. Jamet, O. Lebaigue, N. Coutris, and J. M. Delhaye. The second gradient method for the direct numerical simulation of liquid-vapor flows with phase change. Journal of Computational Physics, 169(2):624–651, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966920.

Jobelin:2006:FEP

[JLL^{+}06] M. Jobelin, C. Lapuerta, J.-C. Latché, Ph. Angot, and B. Piar. A finite element penalty-projection method for incompressible flows. Journal of Computational Physics, 217(2):502–518, September 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000209.

Jin:2005:CMV

[JLOT05a] Shi Jin, Hailiang Liu, Stanley Osher, and Richard Tsai. Computing multi-valued physical observables for the high frequency limit of symmetric hyperbolic systems. Journal of Computational Physics, 210(2):497–518, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002378.

Jin:2005:CMP

[JLOT05b] Shi Jin, Hailiang Liu, Stanley Osher, and Yen-Hsi Richard Tsai. Computing multivalued physical observables for the semiclassical limit of the Schrödinger equation. Journal of Computational Physics, 205(1):222–241, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004656.

Jenny:2003:MSF

[JLT03] P. Jenny, S. H. Lee, and H. A. Tchelepi. Multi-scale finitevolume method for elliptic problems in subsurface flow simulation. Journal of Computational Physics, 187(1):47–67, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000755.

Jenny:2006:AFI

[JLT06] P. Jenny, S. H. Lee, and H. A. Tchelepi. Adaptive fully implicit multi-scale finite-volume method for multi-phase flow and transport in heterogeneous porous media. Journal of Computational Physics, 217(2):627–641, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600026X.

Jones:2000:ACC

[JM00] W. P. Jones and K. R. Menzies. Analysis of the cell-centred finite volume method for the diffusion equation. Journal of Computational Physics, 165(1):45–68, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965951.

Jomaa:2005:EFD

[JM05] Z. Jomaa and C. Macaskill. The embedded finite difference method for the Poisson equation in a domain with an irregular boundary and Dirichlet boundary conditions. Journal of Computational Physics, 202(2):488–506, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002955.

Jothiprasad:2003:HOT

[JMC03] Giridhar Jothiprasad, Dimitri J. Mavriplis, and David A. Caughey. Higher-order time integration schemes for the unsteady Navier–Stokes equations on unstructured meshes. Journal of Computational Physics, 191(2):542–566, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003309.

Johnson:2001:SCC

[JMK01] Bruce R. Johnson, Jeffrey L. Mackey, and James L. Kinsey. Solution of Cartesian and curvilinear quantum equations via multiwavelets on the interval. Journal of Computational Physics, 168(2):356–383, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967014.

Jenny:2001:PSB

[JML^{+}01] P. Jenny, M. Muradoğlu, K. Liu, S. B. Pope, and D. A. Caughey. PDF simulations of a bluff-body stabilized flow. Journal of Computational Physics, 169(1):1–23, May 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196704X.

Jones:2002:ANP

[JMP02] Don A. Jones, Len G. Margolin, and Andrew C. Poje. Accuracy and nonoscillatory properties of enslaved difference schemes. Journal of Computational Physics, 181(2):705– 728, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910297155X.

Jin:2004:NSG

[JMZ04] Shi Jin, Peter A. Markowich, and Chunxiong Zheng. Numerical simulation of a generalized Zakharov system. Journal of Computational Physics, 201(1):376–395, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002347.

Jin:2007:STM

[JN07] Shi Jin and Kyle A. Novak. A semiclassical transport model for two-dimensional thin quantum barriers. Journal of Computational Physics, 226(2):1623–1644, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002525.

Jones:2005:MTF

[Jon05] Don A. Jones. Modified-truncation finite difference schemes. Journal of Computational Physics, 209(1):322–339, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001324.

Jordan:2007:SRP

[Jor07] Stephen A. Jordan. The spatial resolution properties of composite compact finite differencing. Journal of Computational Physics, 221(2):558–576, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003032.

Jiang:2006:SST

[JOS06] Fangming Jiang, Mónica S. A. Oliveira, and Antonio C. M. Sousa. SPH simulation of transition to turbulence for planar shear flow subjected to a streamwise magnetic field. Journal of Computational Physics, 217(2):485–501, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000192.

Jin:2000:DMS

[JP00] Shi Jin and Lorenzo Pareschi. Discretization of the multiscale semiconductor Boltzmann equation by diffusive relaxation schemes. Journal of Computational Physics, 161(1):312–330, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965069.

Jones:2003:IEI

[JP03] Samuel T. Jones and Scott E. Parker. Including electron inertia without advancing electron flow. Journal of Computational Physics, 191(1):322–327, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300305X.

Jenny:2001:HAJ

[JPMC01] P. Jenny, S. B. Pope, M. Muradoğlu, and D. A. Caughey. A hybrid algorithm for the joint PDF equation of turbulent reactive flows. Journal of Computational Physics, 166(2):218–252, January 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966464.

Jiang:2003:SKI

[JR03] Shidong Jiang and Vladimir Rokhlin. Second kind integral equations for the classical potential theory on open surfaces I: analytical apparatus. Journal of Computational Physics, 191(1):40–74, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003048.

Jiang:2004:SKI

[JR04] Shidong Jiang and Vladimir Rokhlin. Second kind integral equations for the classical potential theory on open surfaces II. Journal of Computational Physics, 195(1):1–16, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005096.

Jaisankar:2007:DRE

[JR07] S. Jaisankar and S. V. Raghurama Rao. Diffusion regulation for Euler solvers. Journal of Computational Physics, 221(2):577–599, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003044.

Jaisankar:2009:CRH

[JR09] S. Jaisankar and S. V. Raghurama Rao. A central Rankine– Hugoniot solver for hyperbolic conservation laws. Journal of Computational Physics, 228(3):770–798, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005123.

Jamalyaria:2005:QBM

[JRS05] Farokh Jamalyaria, Rori Rohlfs, and Russell Schwartz. Queue-based method for efficient simulation of biological self-assembly systems. Journal of Computational Physics, 204(1):100–120, March 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004097.

Jung:2005:NSH

[JS05] Rho-Taek Jung and Toru Sato. Numerical simulation of high Schmidt number flow over a droplet by using moving unstructured mesh. Journal of Computational Physics, 203(1):221–249, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003250.

Jung:2007:NCI

[JS07] Jae-Hun Jung and Bernie D. Shizgal. On the numerical convergence with the inverse polynomial reconstruction method for the resolution of the Gibbs phenomenon. Journal of Computational Physics, 224(2):477–488, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000332.

Jiang:2008:BTN

[JSCZ08] Haiyan Jiang, Sihong Shao, Wei Cai, and Pingwen Zhang. Boundary treatments in non-equilibrium Green’s function (NEGF) methods for quantum transport in nano-MOSFETs. Journal of Computational Physics, 227(13):6553–6573, June 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800168X.

Jamet:2002:TCS

[JTB02] Didier Jamet, David Torres, and J. U. Brackbill. On the theory and computation of surface tension: The elimination of parasitic currents through energy conservation in the second-gradient method. Journal of Computational Physics, 182(1):262–276, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971652.

Jenny:2009:UCN

[JTL09] Patrick Jenny, Hamdi A. Tchelepi, and Seong H. Lee. Unconditionally convergent nonlinear solver for hyperbolic conservation laws with S-shaped flux functions. Journal of Computational Physics, 228(20):7497–7512, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003301.

Javierre:2007:LSM

[JVVS07] E. Javierre, C. Vuik, F. J. Vermolen, and A. Segal. A level set method for three dimensional vector Stefan problems: Dissolution of stoichiometric particles in multi-component alloys. Journal of Computational Physics, 224(1):222–240, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000599.

Jin:2002:RNS

[JW02] Shi Jin and Xuelei Wang. Robust numerical simulation of porosity evolution in chemical vapor infiltration: II. Two-dimensional anisotropic fronts. Journal of Computational Physics, 179(2):557–577, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970749.

Jin:2003:RNS

[JW03] Shi Jin and Xuelei Wang. Robust numerical simulation of porosity evolution in chemical vapor infiltration III: three space dimension. Journal of Computational Physics, 186(2):582–595, April 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000834.

Jin:2006:HPS

[JW06] Shi Jin and Xin Wen. Hamiltonian-preserving schemes for the Liouville equation of geometrical optics with discontinuous local wave speeds. Journal of Computational Physics, 214(2):672–697, May 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004687.

Julien:2009:EMD

[JW09] Keith Julien and Mike Watson. Efficient multi-dimensional solution of PDEs using Chebyshev spectral methods. Journal of Computational Physics, 228(5):1480–1503, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800569X.

Jin:2000:RNS

[JWSC00] Shi Jin, Xuelei Wang, Thomas L. Starr, and Xinfu Chen. Robust numerical simulation of porosity evolution in chemical vapor infiltration I: Two space dimension. Journal of Computational Physics, 162(2):467–482, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965501.

Jin:2006:AGM

[JX06] Changqiu Jin and Kun Xu. An adaptive grid method for two-dimensional viscous flows. Journal of Computational Physics, 218(1):68–81, October 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600060X.

Jin:2007:UMG

[JX07] Changqiu Jin and Kun Xu. A unified moving grid gaskinetic method in Eulerian space for viscous flow computation. Journal of Computational Physics, 222(1):155–175, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106003494.

Jin:2008:CHF

[JY08] Shi Jin and Dongsheng Yin. Computational high frequency waves through curved interfaces via the Liouville equation and geometric theory of diffraction. Journal of Computational Physics, 227(12):6106–6139, June 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001241.

Jin:2008:IRC

[JZ08] Bangti Jin and Jun Zou. Inversion of Robin coefficient by a spectral stochastic finite element approach. Journal of Computational Physics, 227(6):3282–3306, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005335.

Kallinderis:2005:INS

[KA05] Y. Kallinderis and H. T. Ahn. Incompressible Navier–Stokes method with general hybrid meshes. Journal of Computational Physics, 210(1):75–108, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001993.

Kolobov:2007:USR

[KAA^{+}07] V. I. Kolobov, R. R. Arslanbekov, V. V. Aristov, A. A. Frolova, and S. A. Zabelok. Unified solver for rarefied and continuum flows with adaptive mesh and algorithm refinement. Journal of Computational Physics, 223(2):589–608, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106004633.

Khenner:2001:NSG

[KAIN01] M. Khenner, A. Averbuch, M. Israeli, and M. Nathan. Numerical simulation of grain-boundary grooving by level set method. Journal of Computational Physics, 170(2):764–784, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967609.

Kirkpatrick:2003:RCB

[KAK03] M. P. Kirkpatrick, S. W. Armfield, and J. H. Kent. A representation of curved boundaries for the solution of the Navier– Stokes equations on a staggered three-dimensional Cartesian grid. Journal of Computational Physics, 184(1):1–36, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910200013X.

Kanaun:2002:NMS

[Kan02] S. K. Kanaun. A numerical method for the solution of electromagnetic wave diffraction problems on perfectly conducting screens. Journal of Computational Physics, 176 (1):170–195, February 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101969748.

Karafyllidis:2004:SEG

[Kar04] Ioannis G. Karafyllidis. Simulation of entanglement generation and variation in quantum computation. Journal of Computational Physics, 200(1):383–397, October 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400172X.

Knopp:2006:GFA

[KAS06] Tobias Knopp, Thomas Alrutz, and Dieter Schwamborn. A grid and flow adaptive wall-function method for RANS turbulence modelling. Journal of Computational Physics, 220 (1):19–40, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106002208.

Kasahara:2007:IVA

[Kas07] Akira Kasahara. Initial-value approach to study the inertiogravity waves without the ‘traditional approximation’. Journal of Computational Physics, 225(2):2175–2197, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001258.

Kong:2008:EAM

[KAS08] Rong Kong, Martin Ambrose, and Jerome Spanier. Efficient, automated Monte Carlo methods for radiation transport. Journal of Computational Physics, 227(22):9463– 9476, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108003628.

Kaul:2003:NBC

[Kau03] Upender K. Kaul. New boundary constraints for elliptic systems used in grid generation problems. Journal of Computational Physics, 189(2):476–492, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002298.

Kannenberg:2000:SEP

[KB00] Keith C. Kannenberg and Iain D. Boyd. Strategies for efficient particle resolution in the direct simulation Monte Carlo method. Journal of Computational Physics, 157(2):727–745, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199963970.

Kolesnikov:2001:EHO

[KB01] A. Kolesnikov and A. J. Baker. An efficient high-order Taylor weak statement formulation for the Navier–Stokes equations. Journal of Computational Physics, 173(2):549–574, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968937.

Kingham:2004:IVF

[KB04] R. J. Kingham and A. R. Bell. An implicit Vlasov– Fokker–Planck code to model non-local electron transport in 2-D with magnetic fields. Journal of Computational Physics, 194(1):1–34, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004583.

Krivodonova:2006:HOA

[KB06] Lilia Krivodonova and Marsha Berger. High-order accurate implementation of solid wall boundary conditions in curved geometries. Journal of Computational Physics, 211 (2):492–512, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002846.

Kurdi:2008:SEM

[KB08] Mohammad H. Kurdi and Philip S. Beran. Spectral element method in time for rapidly actuated systems. Journal of Computational Physics, 227(3):1809–1835, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004342.

Kim:2000:SOT

[KC00] Dongjoo Kim and Haecheon Choi. A second-order timeaccurate finite volume method for unsteady incompressible flow on hybrid unstructured grids. Journal of Computational Physics, 162(2):411–428, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096546X.

Kim:2006:IBM

[KC06] Dokyun Kim and Haecheon Choi. Immersed boundary method for flow around an arbitrarily moving body. Journal of Computational Physics, 212(2):662–680, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003505.

Kanevsky:2007:AIE

[KCGH07] Alex Kanevsky, Mark H. Carpenter, David Gottlieb, and Jan S. Hesthaven. Application of implicit-explicit high order Runge–Kutta methods to discontinuous-Galerkin schemes. Journal of Computational Physics, 225(2):1753–1781, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000861.

Kanevsky:2006:IFS

[KCH06] Alex Kanevsky, Mark H. Carpenter, and Jan S. Hesthaven. Idempotent filtering in spectral and spectral element methods. Journal of Computational Physics, 220(1):41–58, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910600221X.

Knoll:2003:BAT

[KCMM03] D. A. Knoll, L. Chacon, L. G. Margolin, and V. A. Mousseau. On balanced approximations for time integration of multiple time scale systems. Journal of Computational Physics, 185(2):583–611, March 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000081.

Kamath:2009:RAA

[KD09] Hemant Kamath and Xiaoju Du. A roe-average algorithm for a granular-gas model with non-conservative terms. Journal of Computational Physics, 228(21):8187–8202, November 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004422.

Knaepen:2005:LES

[KDC05] Bernard Knaepen, Olivier Debliquy, and Daniele Carati. Large-eddy simulation without filter. Journal of Computational Physics, 205(1):98–107, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004590.

Kim:2007:NMS

[KDF07] Tae-Yeon Kim, John Dolbow, and Eliot Fried. A numerical method for a second-gradient theory of incompressible fluid flow. Journal of Computational Physics, 223(2):551–570, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106004529.

Keetels:2007:FSW

[KDK^{+}07] G. H. Keetels, U. D’Ortona, W. Krämer, H. J. H. Clercx, K. Schneider, and G. J. F. van Heijst. Fourier spectral and wavelet solvers for the incompressible Navier–Stokes equations with volume-penalization: Convergence of a dipole-wall collision. Journal of Computational Physics, 227(2):919–945, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003336.

Karimabadi:2005:NAM

[KDOO05] H. Karimabadi, J. Driscoll, Y. A. Omelchenko, and N. Omidi. A new asynchronous methodology for modeling of physical systems: breaking the curse of Courant condition. Journal of Computational Physics, 205(2):755–775, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004917.

Kubatko:2008:TSR

[KDW08] Ethan J. Kubatko, Clint Dawson, and Joannes J. Westerink. Time step restrictions for Runge–Kutta discontinuous Galerkin methods on triangular grids. Journal of Computational Physics, 227(23):9697–9710, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003963.

Kampanis:2006:SGH

[KE06] Nikolaos A. Kampanis and John A. Ekaterinaris. A staggered grid, high-order accurate method for the incompressible Navier–Stokes equations. Journal of Computational Physics, 215(2):589–613, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005061.

Kuprat:2009:ASI

[KE09] Andrew P. Kuprat and Daniel R. Einstein. An anisotropic scale-invariant unstructured mesh generator suitable for volumetric imaging data. Journal of Computational Physics, 228 (3):619–640, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108004981.

Kavousanakis:2007:PCP

[KEB^{+}07] M. E. Kavousanakis, R. Erban, A. G. Boudouvis, C. W. Gear, and I. G. Kevrekidis. Projective and coarse projective integration for problems with continuous symmetries. Journal of Computational Physics, 225(1):382–407, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005973.

Keller:2005:ABC

[Kel05] Joseph B. Keller. Artificial boundary conditions for axisymmetric slow viscous flow. Journal of Computational Physics, 204(2):399–403, April 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004346.

Knudsen:2006:RAR

[KF06] Henning Arendt Knudsen and Sándor Fazekas. Robust algorithm for random resistor networks using hierarchical domain structure. Journal of Computational Physics, 211(2):700–718, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002937.

Kao:2004:DAE

[KFH^{+}04] Jim Kao, Dawn Flicker, Rudy Henninger, Sarah Frey, Michael Ghil, and Kayo Ide. Data assimilation with an extended Kalman filter for impact-produced shock-wave dynamics. Journal of Computational Physics, 196(2):705–723, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006132.

Kao:2006:EMP

[KFIG06] Jim Kao, Dawn Flicker, Kayo Ide, and Michael Ghil. Estimating model parameters for an impact-produced shockwave simulation: Optimal use of partial data with the extended Kalman filter. Journal of Computational Physics, 214(2):725–737, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105004705.

Kleefsman:2005:VFB

[KFV^{+}05] K. M. T. Kleefsman, G. Fekken, A. E. P. Veldman, B. Iwanowski, and B. Buchner. A volume-of-fluid based simulation method for wave impact problems. Journal of Computational Physics, 206(1):363–393, June 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104005170.

Kim:2007:AAC

[KFV07] H. C. Kim, Y. Feng, and J. P. Verboncoeur. Algorithms for accurate collection, ejection, and loading in particle simulations. Journal of Computational Physics, 223(2):629–642, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004657.

Krauskopf:2003:CUM

[KG03] Bernd Krauskopf and Kirk Green. Computing unstable manifolds of periodic orbits in delay differential equations. Journal of Computational Physics, 186(1):230–249, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000500.

Karniadakis:2006:UQS

[KG06] George Em Karniadakis and James Glimm. Uncertainty quantification in simulation science. Journal of Computational Physics, 217(1):1–4, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002841.

Kennedy:2008:RAF

[KG08] Christopher A. Kennedy and Andrea Gruber. Reduced aliasing formulations of the convective terms within the Navier– Stokes equations for a compressible fluid. Journal of Computational Physics, 227(3):1676–1700, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004251.

Karabasov:2009:CAB

[KG09] S. A. Karabasov and V. M. Goloviznin. Compact Accurately Boundary-Adjusting high-REsolution Technique [CABARET] for fluid dynamics. Journal of Computational Physics, 228(19):7426–7451, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003684.

Kim:2005:GAB

[KGJ05] Kwiseon Kim, Peter A. Graf, and Wesley B. Jones. A genetic algorithm based inverse band structure method for semiconductor alloys. Journal of Computational Physics, 208(2):735–760, September 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001282.

Kim:2007:DNS

[KH07] See Jo Kim and Wook Ryol Hwang. Direct numerical simulations of droplet emulsions in sliding bi-periodic frames using the level-set method. Journal of Computational Physics, 225(1):615–634, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006115.

Kudriakov:2008:NDS

[KH08] S. Kudriakov and W. H. Hui. On a new defect of shockcapturing methods. Journal of Computational Physics, 227(3):2105–2117, January 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700469X.

Kattelans:2009:CMM

[KH09] Thorsten Kattelans and Wilhelm Heinrichs. Conservation of mass and momentum of the least-squares spectral collocation scheme for the Stokes problem. Journal of Computational Physics, 228(13):4649–4664, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001168.

Kang:2008:DSM

[KHdT^{+}08] Tae Gon Kang, Martien A. Hulsen, Jaap M. J. den Toonder, Patrick D. Anderson, and Han E. H. Meijer. A direct simulation method for flows with suspended paramagnetic particles. Journal of Computational Physics, 227(9):4441–4458, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108000181.

Khenner:2004:CMI

[Khe04] M. Khenner. Computation of the material indicator function near the contact line (in Tryggvason’s method). Journal of Computational Physics, 200(1):1–7, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001354.

Kristensen:2001:DIR

[KHV01] J. H. Kristensen, G. L. Hoatson, and R. L. Vold. Design and implementation of Runge–Kutta methods for MAS NMR lineshape calculations. Journal of Computational Physics, 170(1):415–447, June 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967440.

Kaser:2005:ASA

[KI05] Martin Käser and Armin Iske. ADER schemes on adaptive triangular meshes for scalar conservation laws. Journal of Computational Physics, 205(2):486–508, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004772.

Kang:2009:DBD

[KIH09] Seongwon Kang, Gianluca Iaccarino, and Frank Ham. DNS of buoyancy-dominated turbulent flows on a bluff body using the immersed boundary method. Journal of Computational Physics, 228(9):3189–3208, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006517.

Kang:2009:PWP

[KIHM09] Seongwon Kang, Gianluca Iaccarino, Frank Ham, and Parviz Moin. Prediction of wall-pressure fluctuation in turbulent flows with an immersed boundary method. Journal of Computational Physics, 228(18):6753–6772, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002915.

Kim:2005:CST

[Kim05] Junseok Kim. A continuous surface tension force formulation for diffuse-interface models. Journal of Computational Physics, 204(2):784–804, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004383.

Kim:2007:OBC

[Kim07] Jae Wook Kim. Optimised boundary compact finite difference schemes for computational aeroacoustics. Journal of Computational Physics, 225(1):995–1019, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000149.

Kostrun:2001:ANM

[KJ01] Marijan Kostrun and Juha Javanainen. An alternative numerical method for initial value problems involving the contact nonlinear Hamiltonians. Journal of Computational Physics, 172(1):298–308, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196824X.

Katz:2009:MMC

[KJ09a] Aaron Katz and Antony Jameson. Multicloud: Multigrid convergence with a meshless operator. Journal of Computational Physics, 228(14):5237–5250, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001995.

Ketefian:2009:MEV

[KJ09b] G. S. Ketefian and M. Z. Jacobson. A mass, energy, vorticity, and potential enstrophy conserving lateral fluid-land boundary scheme for the shallow water equations. Journal of Computational Physics, 228(1):1–32, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004191.

Karamanos:2000:SVV

[KK00a] G-S. Karamanos and G. E. Karniadakis. A spectral vanishing viscosity method for large-eddy simulations. Journal of Computational Physics, 163(1):22–50, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965525.

Khabibrakhmanov:2000:SCM

[KK00b] Ildar. K. Khabibrakhmanov and George. V. Khazanov. The spectral collocation method for the kinetic equation with the nonlinear two-dimensional Coulomb collisional operator. Journal of Computational Physics, 161(2):558–575, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965136.

Knio:2000:ITT

[KK00c] Omar M. Knio and Rupert Klein. Improved thin-tube models for slender vortex simulations. Journal of Computational Physics, 163(1):68–82, September 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965550.

Kanaun:2003:NMS

[KK03a] S. K. Kanaun and S. Babaii Kochekseraii. A numerical method for the solution of thermo- and electro-static problems for a medium with isolated inclusions. Journal of Computational Physics, 192(2):471–493, December 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004054.

Kirby:2003:ANU

[KK03b] Robert M. Kirby and George Em Karniadakis. De-aliasing on non-uniform grids: algorithms and applications. Journal of Computational Physics, 191(1):249–264, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003140.

Knoll:2004:JFN

[KK04] D. A. Knoll and D. E. Keyes. Jacobian-free Newton–Krylov methods: a survey of approaches and applications. Journal of Computational Physics, 193(2):357–397, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004340.

Keen:2005:SOK

[KK05a] Benjamin Keen and Smadar Karni. A second order kinetic scheme for gas dynamics on arbitrary grids. Journal of Computational Physics, 205(1):108–130, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004607.

Kim:2005:HOA

[KK05b] Dehee Kim and Jang Hyuk Kwon. A high-order accurate hybrid scheme using a central flux scheme and a WENO scheme for compressible flowfield analysis. Journal of Computational Physics, 210(2):554–583, December 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002512.

Kim:2005:AEMa

[KK05c] Kyu Hong Kim and Chongam Kim. Accurate, efficient and monotonic numerical methods for multi-dimensional compressible flows: Part I: Spatial discretization. Journal of Computational Physics, 208(2):527–569, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001208.

Kim:2005:AEMb

[KK05d] Kyu Hong Kim and Chongam Kim. Accurate, efficient and monotonic numerical methods for multi-dimensional compressible flows: Part II: Multi-dimensional limiting process. Journal of Computational Physics, 208(2):570–615, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910500121X.

Kim:2007:ERI

[KK07] Joo Sung Kim and Oh Joon Kwon. An efficient and robust implicit operator for upwind point Gauss–Seidel method. Journal of Computational Physics, 224(2):1124–1144, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005687.

Kallinderis:2009:PMQ

[KK09] Y. Kallinderis and C. Kontzialis. A priori mesh quality estimation via direct relation between truncation error and mesh distortion. Journal of Computational Physics, 228 (3):881–902, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005299.

Kim:2001:IBF

[KKC01] Jungwoo Kim, Dongjoo Kim, and Haecheon Choi. An immersed-boundary finite-volume method for simulations of flow in complex geometries. Journal of Computational Physics, 171(1):132–150, July 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967786.

Khoromskij:2009:TDE

[KKCF09] B. N. Khoromskij, V. Khoromskaia, S. R. Chinnamsetty, and H.-J. Flad. Tensor decomposition in electronic structure calculations on 3D Cartesian grids. Journal of Computational Physics, 228(16):5749–5762, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002356.

Kia:2008:FEF

[KKD08] Amirali Kia, Daejoong Kim, and Eric Darve. Fast electrostatic force calculation on parallel computer clusters. Journal of Computational Physics, 227(19):8551–8567, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800329X.

Kuprat:2001:VCS

[KKGL01] Andrew Kuprat, Ahmed Khamayseh, Denise George, and Levi Larkey. Volume conserving smoothing for piecewise linear curves, surfaces, and triple lines. Journal of Computational Physics, 172(1):99–118, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968160.

Kim:2004:CMM

[KKL04] Junseok Kim, Kyungkeun Kang, and John Lowengrub. Conservative multigrid methods for Cahn–Hilliard fluids. Journal of Computational Physics, 193(2):511–543, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003966.

Kadioglu:2008:FOA

[KKM08] Samet Y. Kadioğlu, Rupert Klein, and Michael L. Minion. A fourth-order auxiliary variable projection method for zero-Mach number gas dynamics. Journal of Computational Physics, 227(3):2012–2043, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004445.

Kozlov:2004:BLE

[KKO04] Roman Kozlov, Anne Kværnø, and Brynjulf Owren. The behaviour of the local error in splitting methods applied to stiff problems. Journal of Computational Physics, 195(2):576–593, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005734.

Karni:2002:SIA

[KKP02] Smadar Karni, Alexander Kurganov, and Guergana Petrova. A smoothness indicator for adaptive algorithms for hyperbolic systems. Journal of Computational Physics, 178(2):323–341, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970245.

Kim:2001:MACa

[KKR01a] Kyu Hong Kim, Chongam Kim, and Oh-Hyun Rho. Methods for the accurate computations of hypersonic flows: I. AUSMPW+ scheme. Journal of Computational Physics, 174(1):38–80, November 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968731.

Kim:2001:MACb

[KKR01b] Kyu Hong Kim, Chongam Kim, and Oh-Hyun Rho. Methods for the accurate computations of hypersonic flows: II. Shock-aligned grid technique. Journal of Computational Physics, 174(1):81–119, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968962.

Kameyama:2005:MIA

[KKS05] Masanori Kameyama, Akira Kageyama, and Tetsuya Sato. Multigrid iterative algorithm using pseudo-compressibility for three-dimensional mantle convection with strongly variable viscosity. Journal of Computational Physics, 206(1):162–181, June 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104005066.

Kramer:2007:CAM

[KKS07] Peter R. Kramer, Orazgeldi Kurbanmuradov, and Karl Sabelfeld. Comparative analysis of multiscale Gaussian random field simulation algorithms. Journal of Computational Physics, 226(1):897–924, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002148.

Kirshman:2004:GBC

[KL04] D. J. Kirshman and F. Liu. A gridless boundary condition method for the solution of the Euler equations on embedded Cartesian meshes with multigrid. Journal of Computational Physics, 201(1):119–147, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002049.

Klose:2006:LTB

[KL06] Alexander D. Klose and Edward W. Larsen. Light transport in biological tissue based on the simplified spherical harmonics equations. Journal of Computational Physics, 220(1):441–470, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003421.

Kawai:2008:LAD

[KL08] S. Kawai and S. K. Lele. Localized artificial diffusivity scheme for discontinuity capturing on curvilinear meshes. Journal of Computational Physics, 227(22):9498–9526, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108003641.

Kim:2008:WEH

[KLK08] Sungtae Kim, Soogab Lee, and Kyu Hong Kim. Wavenumberextended high-order oscillation control finite volume schemes for multi-dimensional aeroacoustic computations. Journal of Computational Physics, 227(8):4089–4122, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005657.

Kim:2009:RHR

[KLLJ09] Sung Don Kim, Bok Jik Lee, Hyoung Jin Lee, and In-Seuck Jeung. Robust HLLC Riemann solver with weighted average flux scheme for strong shock. Journal of Computational Physics, 228(20):7634–7642, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003842.

Kroger:2005:EGS

[KLM05] Tim Kröger and Mária Lukácová-Medvid’ová. An evolution Galerkin scheme for the shallow water magnetohydrodynamic equations in two space dimensions. Journal of Computational Physics, 206(1):122–149, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005042.

Knoll:2007:NAT

[KLM07] D. A. Knoll, R. B. Lowrie, and J. E. Morel. Numerical analysis of time integration errors for nonequilibrium radiation diffusion. Journal of Computational Physics, 226(2):1332–1347, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002367.

Karlsen:2001:OSM

[KLN^{+}01] K. Hvistendahl Karlsen, K.-A Lie, J. R. Natvig, H. F. Nordhaug, and H. K. Dahle. Operator splitting methods for systems of convection-diffusion equations: Nonlinear error mechanisms and correction strategies. Journal of Computational Physics, 173(2):636–663, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969013.

Kwon:2009:ETD

[KLP^{+}09] Sungjin Kwon, Youngmin Lee, Jong Youn Park, Dongwoo Sohn, Jae Hyuk Lim, and Seyoung Im. An efficient three-dimensional adaptive quasicontinuum method using variable-node elements. Journal of Computational Physics, 228(13):4789–4810, July 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900165X.

Kuo:2009:MET

[KLSW09] Yueh-Cheng Kuo, Wen-Wei Lin, Shih-Feng Shieh, and Weichung Wang. A minimal energy tracking method for non-radially symmetric solutions of coupled nonlinear Schrödinger equations. Journal of Computational Physics, 228(21):7941–7956, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003921.

Koren:2002:RPL

[KLvBvL02] B. Koren, M. R. Lewis, E. H. van Brummelen, and B. van Leer. Riemann-problem and level-set approaches for homentropic two-fluid flow computations. Journal of Computational Physics, 181(2):654–674, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971500.

Kube:2009:MCS

[KLW09] Susanna Kube, Caroline Lasser, and Marcus Weber. Monte Carlo sampling of Wigner functions and surface hopping quantum dynamics. Journal of Computational Physics, 228 (6):1947–1962, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006104.

Knoll:2000:NKM

[KM00] D. A. Knoll and V. A. Mousseau. On Newton–Krylov multigrid methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 163(1):262–267, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965616.

Kloucek:2002:CMB

[KM02] Petr Kloucek and Luis A. Melara. The computational modelling of branching fine structures in constrained crystals. Journal of Computational Physics, 183(2):623–651, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102972050.

Kim:2003:RTC

[KM03] Arnold D. Kim and Miguel Moscoso. Radiative transfer computations for optical beams. Journal of Computational Physics, 185(1):50–60, February 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000475.

Kemenov:2006:ESS

[KM06] K. A. Kemenov and S. Menon. Explicit small-scale velocity simulation for high-Re turbulent flows. Journal of Computational Physics, 220(1):290–311, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002324.

Mao:2007:TFT

[kM07a] De kang Mao. Towards front-tracking based on conservation in two space dimensions II, tracking discontinuities in capturing fashion. Journal of Computational Physics, 226(2):1550–1588, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002495.

Kemenov:2007:ESS

[KM07b] K. A. Kemenov and S. Menon. Explicit small-scale velocity simulation for high-Re turbulent flows. Part II: Nonhomogeneous flows. Journal of Computational Physics, 222 (2):673–701, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106003731.

Kajtar:2008:SSS

[KM08a] J. Kajtar and J. J. Monaghan. SPH simulations of swimming linked bodies. Journal of Computational Physics, 227(19):8568–8587, October 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003306.

Keaveny:2008:MMI

[KM08b] Eric E. Keaveny and Martin R. Maxey. Modeling the magnetic interactions between paramagnetic beads in magnetorheological fluids. Journal of Computational Physics, 227(22):9554–9571, November 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003677.

Kim:2001:HSS

[KMA^{+}01] J.-H. R. Kim, H. Maurer, Yu. A. Astrov, M. Bode, and H.-G. Purwins. High-speed switch-on of a semiconductor gas discharge image converter using optimal control methods. Journal of Computational Physics, 170(1):395–414, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967415.

Kostelec:2000:CHA

[KMHR00] Peter J. Kostelec, David K. Maslen, Dennis M. Healy, Jr., and Daniel N. Rockmore. Computational harmonic analysis for tensor fields on the two-sphere. Journal of Computational Physics, 162(2):514–535, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965513.

Kalitzin:2005:NWB

[KMID05] Georgi Kalitzin, Gorazd Medic, Gianluca Iaccarino, and Paul Durbin. Near-wall behavior of RANS turbulence models and implications for wall functions. Journal of Computational Physics, 204(1):265–291, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004164.

Kwok:2001:CER

[KMJ01] Wai Yip Kwok, Robert D. Moser, and Javier Jiménez. A critical evaluation of the resolution properties of B-spline and compact finite difference methods. Journal of Computational Physics, 174(2):510–551, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969190.

Knupp:2002:RJO

[KMS02] Patrick Knupp, Len G. Margolin, and Mikhail Shashkov. Reference Jacobian optimization-based rezone strategies for arbitrary Lagrangian Eulerian methods. Journal of Computational Physics, 176(1):93–128, February 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969694.

Karlin:2003:HAP

[KMS03] V. Karlin, V. Maz’ya, and G. Schmidt. High accuracy periodic solutions to the Sivashinsky equation. Journal of Computational Physics, 188(1):209–231, June 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001645.

Knudsen:2004:NSM

[KMS04] K. Knudsen, J. Mueller, and S. Siltanen. Numerical solution method for the dbar-equation in the plane. Journal of Computational Physics, 198(2):500–517, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910400049X.

Kaufmann:2008:CBL

[KMSH08] A. Kaufmann, M. Moreau, O. Simonin, and J. Helie. Comparison between Lagrangian and mesoscopic Eulerian modelling approaches for inertial particles suspended in decaying isotropic turbulence. Journal of Computational Physics, 227(13):6448–6472, June 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001411.

Katsoulakis:2003:CGS

[KMV03] Markos A. Katsoulakis, Andrew J. Majda, and Dionisios G. Vlachos. Coarse-grained stochastic processes and Monte Carlo simulations in lattice systems. Journal of Computational Physics, 186(1):250–278, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000512.

Kumano:2004:CNS

[KN04] S. Kumano and T.-H. Nagai. Comparison of numerical solutions for Q^{2} evolution equations. Journal of Computational Physics, 201(2):651–664, December 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002591.

Kamenkovich:2009:TSS

[KN09] V. M. Kamenkovich and D. A. Nechaev. On the time-splitting scheme used in the Princeton Ocean Model. Journal of Computational Physics, 228(8):2874–2905, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000059.

Klose:2005:ISP

[KNH05] Alexander D. Klose, Vasilis Ntziachristos, and Andreas H. Hielscher. The inverse source problem based on the radiative transfer equation in optical molecular imaging. Journal of Computational Physics, 202(1):323–345, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002888.

Kok:2009:HOL

[Kok09] J. C. Kok. A high-order low-dispersion symmetry-preserving finite-volume method for compressible flow on curvilinear grids. Journal of Computational Physics, 228(18):6811–6832, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003234.

Kao:2004:LFS

[KOQ04] Chiu Yen Kao, Stanley Osher, and Jianliang Qian. Lax– Friedrichs sweeping scheme for static Hamilton–Jacobi equations. Journal of Computational Physics, 196(1):367–391, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006016.

Kao:2008:LTB

[KOQ08] Chiu-Yen Kao, Stanley Osher, and Jianliang Qian. Legendretransform-based fast sweeping methods for static Hamilton– Jacobi equations on triangulated meshes. Journal of Computational Physics, 227(24):10209–10225, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800452X.

Koutsourelakis:2007:SUS

[Kou07] P. S. Koutsourelakis. Stochastic upscaling in solid mechanics: An exercise in machine learning. Journal of Computational Physics, 226(1):301–325, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001581.

Koumoutsakos:2008:P

[Kou08] Petros Koumoutsakos. Preface. Journal of Computational Physics, 227(21):9005, November 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003586.

Koutsourelakis:2009:MRN

[Kou09] P. S. Koutsourelakis. A multi-resolution, non-parametric, Bayesian framework for identification of spatially-varying model parameters. Journal of Computational Physics, 228(17):6184–6211, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002708.

Kim:2000:MPT

[KP00] Charlson C. Kim and Scott E. Parker. Massively parallel three-dimensional toroidal gyrokinetic flux-tube turbulence simulation. Journal of Computational Physics, 161(2):589– 604, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965185.

Kelley:2004:FSO

[KP04] C. T. Kelley and B. Montgomery Pettitt. A fast solver for the Ornstein–Zernike equations. Journal of Computational Physics, 197(2):491–501, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300648X.

Kurzak:2005:COF

[KP05] Jakub Kurzak and B. Montgomery Pettitt. Communications overlapping in fast multipole particle dynamics methods. Journal of Computational Physics, 203(2):731–743, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004024.

Kamga:2007:NZM

[KP07] Jean-Baptiste Apoung Kamga and Olivier Pironneau. Numerical zoom for multiscale problems with an application to nuclear waste disposal. Journal of Computational Physics, 224(1):403–413, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001337.

Kopitz:2008:CBA

[KP08] J. Kopitz and W. Polifke. CFD-based application of the Nyquist criterion to thermo-acoustic instabilities. Journal of Computational Physics, 227(14):6754–6778, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001733.

Kim:2008:AHO

[KPB08] Seung Hyun Kim, Heinz Pitsch, and Iain D. Boyd. Accuracy of higher-order lattice Boltzmann methods for microscale flows with finite Knudsen numbers. Journal of Computational Physics, 227(19):8655–8671, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003355.

Kissmann:2009:CCS

[KPK09] R. Kissmann, J. Pomoell, and W. Kley. A central conservative scheme for general rectangular grids. Journal of Computational Physics, 228(6):2119–2131, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006207.

Kramer:2007:CES

[KPP07] R. M. J. Kramer, C. Pantano, and D. I. Pullin. A class of energy stable, high-order finite-difference interface schemes suitable for adaptive mesh refinement of hyperbolic problems. Journal of Computational Physics, 226(2):1458–1484, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002434.

Kramer:2009:NES

[KPP09] R. M. J. Kramer, C. Pantano, and D. I. Pullin. Nondissipative and energy-stable high-order finite-difference interface schemes for 2-D patch-refined grids. Journal of Computational Physics, 228(14):5280–5297, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002022.

Kunik:2003:KSU

[KQW03a] Matthias Kunik, Shamsul Qamar, and Gerald Warnecke. Kinetic schemes for the ultra-relativistic Euler equations. Journal of Computational Physics, 187(2):572–596, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001256.

Kunik:2003:SOA

[KQW03b] Matthias Kunik, Shamsul Qamar, and Gerald Warnecke. Second-order accurate kinetic schemes for the ultra-relativistic Euler equations. Journal of Computational Physics, 192(2): 695–726, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103004157.

Karlsen:2002:USM

[KR02] Kenneth Hvistendahl Karlsen and Nils Henrik Risebro. Unconditionally stable methods for Hamilton–Jacobi equations. Journal of Computational Physics, 180(2):710–735, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971135.

Kalita:2009:TFH

[KR09a] Jiten C. Kalita and Rajendra K. Ray. A transformationfree HOC scheme for incompressible viscous flows past an impulsively started circular cylinder. Journal of Computational Physics, 228(14):5207–5236, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001983.

Kolev:2009:TAV

[KR09b] Tz. V. Kolev and R. N. Rieben. A tensor artificial viscosity using a finite element approach. Journal of Computational Physics, 228(22):8336–8366, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004446.

Kurkcu:2009:SEE

[KR09c] Harun Kurkcu and Fernando Reitich. Stable and efficient evaluation of periodized Green’s functions for the Helmholtz equation at high frequencies. Journal of Computational Physics, 228(1):75–95, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004531.

Krivodonova:2007:LHO

[Kri07] Lilia Krivodonova. Limiters for high-order discontinuous Galerkin methods. Journal of Computational Physics, 226(1):879–896, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002136.

Kropinski:2001:ENM

[Kro01] M. C. A. Kropinski. An efficient numerical method for studying interfacial motion in two-dimensional creeping flows. Journal of Computational Physics, 171(2):479–508, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967877.

Kropinski:2002:NMM

[Kro02] M. C. A. Kropinski. Numerical methods for multiple inviscid interfaces in creeping flows. Journal of Computational Physics, 180(1):1–24, July 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969530.

Krogstad:2005:GIF

[Kro05] S. Krogstad. Generalized integrating factor methods for stiff PDEs. Journal of Computational Physics, 203(1):72– 88, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003183.

Kitsios:2009:BSA

[KRT^{+}09] Vassili Kitsios, Daniel Rodríguez, Vassilis Theofilis, Andrew Ooi, and Julio Soria. BiGlobal stability analysis in curvilinear coordinates of massively separated lifting bodies. Journal of Computational Physics, 228(19):7181–7196, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003349.

Kryzhniy:2004:HRE

[Kry04] V. V. Kryzhniy. High-resolution exponential analysis via regularized numerical inversion of Laplace transforms. Journal of Computational Physics, 199(2):618–630, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104000981.

Kaipio:2002:EAI

[KS02a] Jari P. Kaipio and Erkki Somersalo. Estimating anomalies from indirect observations. Journal of Computational Physics, 181(2):398–406, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971093.

Kang:2002:EML

[KS02b] Sang-Yoon Kang and Ashok S. Sangani. An efficient method for large-scale simulations of bubbly liquids. Journal of Computational Physics, 179(1):330–345, June 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970658.

Kwan:2007:EDP

[KS07] Yuen-Yick Kwan and Jie Shen. An efficient direct parallel spectral-element solver for separable elliptic problems. Journal of Computational Physics, 225(2):1721–1735, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000824.

Kadioglu:2008:AST

[KS08a] Samet Y. Kadioğlu and Mark Sussman. Adaptive solution techniques for simulating underwater explosions and implosions. Journal of Computational Physics, 227(3):2083–2104, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004652.

Kong:2008:NPG

[KS08b] Rong Kong and Jerome Spanier. A new proof of geometric convergence for general transport problems based on sequential correlated sampling methods. Journal of Computational Physics, 227(23):9762–9777, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004051.

Kolomenskiy:2009:FSM

[KS09] Dmitry Kolomenskiy and Kai Schneider. A Fourier spectral method for the Navier–Stokes equations with volume penalization for moving solid obstacles. Journal of Computational Physics, 228(16):5687–5709, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002149.

Kwatra:2009:MAA

[KSGF09] Nipun Kwatra, Jonathan Su, Jón T. Grétarsson, and Ronald Fedkiw. A method for avoiding the acoustic time step restriction in compressible flow. Journal of Computational Physics, 228(11):4146–4161, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000977.

Karle:2006:NSN

[KSH^{+}06] Ch. Karle, J. Schweitzer, M. Hochbruck, E. W. Laedke, and K. H. Spatschek. Numerical solution of nonlinear wave equations in stratified dispersive media. Journal of Computational Physics, 216(1):138–152, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005462.

Karle:2008:PIT

[KSHS08] Ch. Karle, J. Schweitzer, M. Hochbruck, and K. H. Spatschek. A parallel implementation of a two-dimensional fluid laserplasma integrator for stratified plasma-vacuum systems. Journal of Computational Physics, 227(16):7701–7719, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002581.

Klein:2003:DFB

[KSJ03] M. Klein, A. Sadiki, and J. Janicka. A digital filter based generation of inflow data for spatially developing direct numerical or large eddy simulations. Journal of Computational Physics, 186(2):652–665, April 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000901.

Kadioglu:2005:SOP

[KSO^{+}05] Samet Y. Kadioğlu, Mark Sussman, Stanley Osher, Joseph P. Wright, and Myungjoo Kang. A second order primitive preconditioner for solving all speed multi-phase flows. Journal of Computational Physics, 209(2):477–503, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001798.

Kuzmin:2009:CFE

[KSS09] D. Kuzmin, M. J. Shashkov, and D. Svyatskiy. A constrained finite element method satisfying the discrete maximum principle for anisotropic diffusion problems. Journal of Computational Physics, 228(9):3448–3463, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000552.

Kucharik:2003:ELB

[KSW03] Milan Kucharik, Mikhail Shashkov, and Burton Wendroff. An efficient linearity-and-bound-preserving remapping method. Journal of Computational Physics, 188(2):462–471, July 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001876.

Klingenberg:2007:NCR

[KSW07] Christian Klingenberg, Wolfram Schmidt, and Knut Waagan. Numerical comparison of Riemann solvers for astrophysical hydrodynamics. Journal of Computational Physics, 227(1):12–35, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003063.

Kurganov:2000:NHRa

[KT00a] Alexander Kurganov and Eitan Tadmor. New highresolution central schemes for nonlinear conservation laws and convection-diffusion equations. Journal of Computational Physics, 160(1):241–282, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964593.

Kurganov:2000:NHRb

[KT00b] Alexander Kurganov and Eitan Tadmor. New high-resolution semi-discrete central schemes for Hamilton–Jacobi equations. Journal of Computational Physics, 160(2):720–742, May 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964854.

Kuzmin:2002:FCT

[KT02] D. Kuzmin and S. Turek. Flux correction tools for finite elements. Journal of Computational Physics, 175(2):525–558, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969554.

Kondaxakis:2003:WLC

[KT03] Dimitris Kondaxakis and Sokrates Tsangaris. A weak Legendre collocation spectral method for the solution of the incompressible Navier–Stokes equations in unstructured quadrilateral subdomains. Journal of Computational Physics, 192(1):124–156, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003504.

Kuzmin:2004:HRF

[KT04] D. Kuzmin and S. Turek. High-resolution FEM–TVD schemes based on a fully multidimensional flux limiter. Journal of Computational Physics, 198(1):131–158, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000221.

Kondaxakis:2005:PSL

[KT05] D. Kondaxakis and S. Tsangaris. Pseudospectral solution of linear evolution equations of second order in space and time on unstructured quadrilateral subdomain topologies. Journal of Computational Physics, 202(2):533–576, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002979.

Kashdan:2006:HOA

[KT06] Eugene Kashdan and Eli Turkel. High-order accurate modeling of electromagnetic wave propagation across media — grid conforming bodies. Journal of Computational Physics, 218(2):816–835, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001343.

Kwok:2007:PBR

[KT07] Felix Kwok and Hamdi Tchelepi. Potential-based reduced Newton algorithm for nonlinear multiphase flow in porous media. Journal of Computational Physics, 227(1):706–727, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003543.

Khan:2003:CSM

[KTD03] M. A. H. Khan, Y. Tourigny, and P. G. Drazin. A case study of methods of series summation: Kelvin–Helmholtz instability of finite amplitude. Journal of Computational Physics, 187(1):212–229, May 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000962.

Kulikovsky:2001:SAS

[Kul01] A. A. Kulikovsky. Simple and accurate scheme for nonlinear convection-diffusion equation. Journal of Computational Physics, 173(2):716–729, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969098.

Kumar:2004:IFD

[Kum04a] Anand Kumar. Isotropic finite-differences. Journal of Computational Physics, 201(1):109–118, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002037.

Kummel:2004:DGI

[Küm04b] Stephan Kümmel. Damped gradient iteration and multigrid relaxation: tools for electronic structure calculations using orbital density-functionals. Journal of Computational Physics, 201(1):333–343, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002323.

Kuzmin:2006:DGP

[Kuz06] D. Kuzmin. On the design of general-purpose flux limiters for finite element schemes. I. Scalar convection. Journal of Computational Physics, 219(2):513–531, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001902.

Kuzmin:2009:EIF

[Kuz09] Dmitri Kuzmin. Explicit and implicit FEM–FCT algorithms with flux linearization. Journal of Computational Physics, 228(7):2517–2534, April 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006475.

Klaij:2006:PTS

[KvdVvdV06a] C. M. Klaij, J. J. W. van der Vegt, and H. van der Ven. Pseudo-time stepping methods for space–time discontinuous Galerkin discretizations of the compressible Navier–Stokes equations. Journal of Computational Physics, 219(2):622– 643, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106001951.

Klaij:2006:STD

[KvdVvdV06b] C. M. Klaij, J. J. W. van der Vegt, and H. van der Ven. Space–time discontinuous Galerkin method for the compressible Navier–Stokes equations. Journal of Computational Physics, 217(2):589–611, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000246.

Klaij:2007:MST

[KvRvdVvdV07] C. M. Klaij, M. H. van Raalte, H. van der Ven, and J. J. W. van der Vegt. h-multigrid for space–time discontinuous Galerkin discretizations of the compressible Navier–Stokes equations. Journal of Computational Physics, 227(2):1024–1045, December 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003762.

Kraft:2003:ISA

[KW03]  Markus Kraft and Wolfgang Wagner.  An improved

stochastic algorithm for temperature-dependent homogeneous gas phase reactions. Journal of Computational Physics, 185(1):139–157, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000517.

Kay:2006:MFE

[KW06] David Kay and Richard Welford. A multigrid finite element solver for the Cahn–Hilliard equation. Journal of Computational Physics, 212(1):288–304, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003219.

Kang:2008:PNS

[KW08a] Daekyoung Kang and E. Won. Precise numerical solutions of potential problems using the Crank–Nicolson method. Journal of Computational Physics, 227(5):2970–2976, February 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005244.

Katz:2008:SDS

[KW08b] Richard F. Katz and M. Grae Worster. Simulation of directional solidification, thermochemical convection, and chimney formation in a Hele–Shaw cell. Journal of Computational Physics, 227(23):9823–9840, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004099.

Klockner:2009:NDG

[KWBH09] A. Klöckner, T. Warburton, J. Bridge, and J. S. Hesthaven. Nodal discontinuous Galerkin methods on graphics processors. Journal of Computational Physics, 228 (21):7863–7882, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003647.

Kubatko:2007:SDD

[KWD07] Ethan J. Kubatko, Joannes J. Westerink, and Clint Dawson. Semi discrete discontinuous Galerkin methods and stage-exceeding-order, strong-stability-preserving Runge– Kutta time discretizations. Journal of Computational Physics, 222(2):832–848, March 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003974.

Kwok:2008:HBE

[Kwo08] Dixon T. K. Kwok. A hybrid Boltzmann electrons and PIC ions model for simulating transient state of partially ionized plasma. Journal of Computational Physics, 227(11): 5758–5777, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108001113.

Kao:2008:ICM

[KY08] P.-H. Kao and R.-J. Yang. An investigation into curved and moving boundary treatments in the lattice Boltzmann method. Journal of Computational Physics, 227(11):5671– 5690, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108001058.

Kirby:2007:TSC

[KYK07] Robert M. Kirby, Zohar Yosibash, and George Em Karniadakis. Towards stable coupling methods for high-order discretization of fluid-structure interaction: Algorithms and observations. Journal of Computational Physics, 223(2):489– 518, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004499.

Kim:2007:EMA

[KYLB07] Do Wan Kim, Young-Cheol Yoon, Wing Kam Liu, and Ted Belytschko. Extrinsic meshfree approximation using asymptotic expansion for interfacial discontinuity of derivative. Journal of Computational Physics, 221(1):370–394, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002956.

Karaa:2004:HOA

[KZ04] Samir Karaa and Jun Zhang. High order ADI method for solving unsteady convection-diffusion problems. Journal of Computational Physics, 198(1):1–9, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400018X.

Klein:2006:CAC

[KZ06] P. A. Klein and J. A. Zimmerman. Coupled atomisticcontinuum simulations using arbitrary overlapping domains. Journal of Computational Physics, 213(1):86–116, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003608.

Ke:2009:IKV

[KZWY09] Peng Ke, Shuguang Zhang, Jianghao Wu, and Chunxin Yang. An improved known vicinity algorithm based on geometry test for particle localization in arbitrary grid. Journal of Computational Physics, 228(24):9001–9019, December 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004884.

Labrosse:2009:PLF

[Lab09] G. Labrosse. The piecewise-linear Finite Volume scheme: The best known lowest-order preconditioner for the d^{2}/dx^{2} Chebyshev spectral operator. Journal of Computational Physics, 228(12):4491–4509, July 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001405.

Lai:2002:SCF

[Lai02] Ming-Chih Lai. A simple compact fourth-order Poisson solver on polar geometry. Journal of Computational Physics, 182(1):337–345, October 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297172X.

Li:2008:CEI

[LAKD08] S. Li, S. Ahmed, G. Klimeck, and E. Darve. Computing entries of the inverse of a sparse matrix using the FIND algorithm. Journal of Computational Physics, 227(22):9408–9427, November 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003458.

Lapenta:2002:PRM

[Lap02] Giovanni Lapenta. Particle rezoning for multidimensional kinetic particle-in-cell simulations. Journal of Computational Physics, 181(1):317–337, September 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971263.

Lappa:2003:AKB

[Lap03] Marcello Lappa. An “attachment kinetics-based” volume of fraction method for organic crystallization: a fluid-dynamic approach to macromolecular-crystal engineering. Journal of Computational Physics, 191(1):97–129, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003073.

Lapenta:2004:VGA

[Lap04] Giovanni Lapenta. Variational grid adaptation based on the minimization of local truncation error: time-independent problems. Journal of Computational Physics, 193(1):159–179, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910300425X.

Laprise:2008:RCM

[Lap08] René Laprise. Regional climate modelling. Journal of Computational Physics, 227(7):3641–3666, March 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005407.

Larson:2003:CCM

[Lar03] David J. Larson. A Coulomb collision model for PIC plasma simulation. Journal of Computational Physics, 188 (1):123–138, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103001578.

Larroche:2007:EEN

[Lar07] O. Larroche. An efficient explicit numerical scheme for diffusion-type equations with a highly inhomogeneous and highly anisotropic diffusion tensor. Journal of Computational Physics, 223(1):436–450, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004475.

Larsson:2009:BTC

[Lar09] Johan Larsson. Blending technique for compressible inflow turbulence: Algorithm localization and accuracy assessment. Journal of Computational Physics, 228(4):933–937, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005597.

Lott:2001:QWE

[LAS01] Dawn A. Lott, Stuart S. Antman, and William G. Szymczak. The quasilinear wave equation for antiplane shearing of nonlinearly elastic bodies. Journal of Computational Physics, 171(1):201–226, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196783X.

Lau:2004:RER

[Lau04] Stephen R. Lau. Rapid evaluation of radiation boundary kernels for time-domain wave propagation on blackholes: theory and numerical methods. Journal of Computational Physics, 199(1):376–422, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002086.

Laurenzi:2006:EGA

[Lau06] Ian J. Laurenzi. Erratum to “A general algorithm for exact simulation of multicomponent aggregation processes” [J. Comput. Phys. 177 (2002) 418–449]. Journal of Computational Physics, 217(2):866–867, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001306. See [LBD02].

Layton:2002:CSC

[Lay02] Anita T. Layton. Cubic spline collocation method for the shallow water equations on the sphere. Journal of Computational Physics, 179(2):578–592, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970750.

Leung:2003:CMO

[LB03a] C. H. Leung and M. Berzins. A computational model for organism growth based on surface mesh generation. Journal of Computational Physics, 188(1):75–99, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001530.

Lucia:2003:PMR

[LB03b] David J. Lucia and Philip S. Beran. Projection methods for reduced order models of compressible flows. Journal of Computational Physics, 188(1):252–280, June 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001669.

Luo:2004:CFN

[LB04] Haoxiang Luo and Thomas R. Bewley. On the contravariant form of the Navier–Stokes equations in timedependent curvilinear coordinate systems. Journal of Computational Physics, 199(1):355–375, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000889.

Laurenzi:2002:GAE

[LBD02] Ian J. Laurenzi, John D. Bartels, and Scott L. Diamond. A general algorithm for exact simulation of multicomponent aggregation processes. Journal of Computational Physics, 177(2):418–449, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970178. See erratum [Lau06].

Luo:2004:CMM

[LBL04] Hong Luo, Joseph D. Baum, and Rainald Löhner. On the computation of multi-material flows using ALE formulation. Journal of Computational Physics, 194(1):304–328, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004844.

Luo:2006:HCG

[LBL06a] Hong Luo, Joseph D. Baum, and Rainald Löhner. A hybrid Cartesian grid and gridless method for compressible flows. Journal of Computational Physics, 214(2):618–632, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004651.

Luo:2006:MDG

[LBL06b] Hong Luo, Joseph D. Baum, and Rainald Löhner. A p-multigrid discontinuous Galerkin method for the Euler equations on unstructured grids. Journal of Computational Physics, 211(2):767–783, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003128.

Luo:2007:HWB

[LBL07] Hong Luo, Joseph D. Baum, and Rainald Löhner. A Hermite WENO-based limiter for discontinuous Galerkin method on unstructured grids. Journal of Computational Physics, 225(1):686–713, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006164.

Luo:2008:DGM

[LBL08] Hong Luo, Joseph D. Baum, and Rainald Löhner. A discontinuous Galerkin method based on a Taylor basis for the compressible flows on arbitrary grids. Journal of Computational Physics, 227(20):8875–8893, October 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003604.

Liang:2004:FMR

[LBS^{+}04] WanZhen Liang, Roi Baer, Chandra Saravanan, Yihan Shao, Alexis T. Bell, and Martin Head-Gordon. Fast methods for resumming matrix polynomials and Chebyshev matrix polynomials. Journal of Computational Physics, 194(2):575–587, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004959.

Lanser:2000:SDS

[LBV00] D. Lanser, J. G. Blom, and J. G. Verwer. Spatial discretization of the shallow water equations in spherical geometry using Osher’s scheme. Journal of Computational Physics, 165(2):542–565, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966324.

Lanser:2001:TIS

[LBV01] D. Lanser, J. G. Blom, and J. G. Verwer. Time integration of the shallow water equations in spherical geometry. Journal of Computational Physics, 171(1):373–393, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968020.

Lerat:2001:RBC

[LC01] Alain Lerat and Christophe Corre. A residual-based compact scheme for the compressible Navier–Stokes equations. Journal of Computational Physics, 170(2):642–675, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967555.

Lozinski:2003:FSF

[LC03] Alexei Lozinski and Cédric Chauvière. A fast solver for Fokker–Planck equation applied to viscoelastic flows calculations: 2D FENE model. Journal of Computational Physics, 189(2):607–625, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002481.

Lapenta:2006:CEF

[LC06a] Giovanni Lapenta and Luis Chacón. Cost-effectiveness of fully implicit moving mesh adaptation: a practical investigation in 1D. Journal of Computational Physics, 219(1): 86–103, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106001410.

Loubere:2006:FWD

[LC06b] R. Loubère and E. J. Caramana. The force/work differencing of exceptional points in the discrete, compatible formulation of Lagrangian hydrodynamics. Journal of Computational Physics, 216(1):1–18, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005395.

Labrunie:2004:NSH

[LCB04] Simon Labrunie, José A. Carrillo, and Pierre Bertrand. Numerical study on hydrodynamic and quasi-neutral approximations for collisionless two-species plasmas. Journal of Computational Physics, 200(1):267–298, October 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001676.

Li:2009:ETF

[LCB09] Jing-Rebecca Li, Donna Calhoun, and Lucien Brush. Efficient thermal field computation in phase-field models. Journal of Computational Physics, 228(24):8945–8957, December 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004392.

Lyons:2005:FAS

[LCCG05] W. Lyons, H. D. Ceniceros, S. Chandrasekaran, and M. Gu. Fast algorithms for spectral collocation with non-periodic boundary conditions. Journal of Computational Physics, 207(1):173–191, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000227.

Lynch:2003:NMS

[LCdCN^{+}03] V. E. Lynch, B. A. Carreras, D. del Castillo-Negrete, K. M. Ferreira-Mejias, and H. R. Hicks. Numerical methods for the solution of partial differential equations of fractional order. Journal of Computational Physics, 192(2):406–421, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004005.

Lee:2007:ESC

[LCG07] Eunseok Lee, Wei Cai, and Giulia A. Galli. Electronic structure calculations in a uniform magnetic field using periodic supercells. Journal of Computational Physics, 226(2):1310–1331, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002355.

Li:2003:SMB

[LCH03] Baoyan Li, Zhangxin Chen, and Guanren Huan. The sequential method for the black-oil reservoir simulation on unstructured grids. Journal of Computational Physics, 192 (1):36–72, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003462.

Lu:2007:NVF

[LCM07] Benzhuo Lu, Xiaolin Cheng, and J. Andrew McCammon. “New-version-fast-multipole-method” accelerated electrostatic calculations in biomolecular systems. Journal of Computational Physics, 226(2):1348–1366, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002379.

Liu:2007:CAS

[LCNR07] Jin Liu, Shiyi Chen, Xiaobo Nie, and Mark O. Robbins. A continuum-atomistic simulation of heat transfer in micro- and nano-flows. Journal of Computational Physics, 227(1):279–291, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003324.

Lin:2002:SCF

[LCS02] Ching-Long Lin, Tianfeng Chai, and Juanzhen Sun. On the smoothness constraints for four-dimensional data assimilation. Journal of Computational Physics, 181(2):430–453, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971366.

Liu:2009:HOC

[LCS09] Wei Liu, Juan Cheng, and Chi-Wang Shu. High order conservative Lagrangian schemes with Lax–Wendroff type time discretization for the compressible Euler equations. Journal of Computational Physics, 228(23):8872–8891, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004860.

Li:2004:FEM

[LCW04] Xiang-Gui Li, C. K. Chan, and Song Wang. The finite element method with weighted basis functions for singularly perturbed convection-diffusion problems. Journal of Computational Physics, 195(2):773–789, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005825.

Londrillo:2004:DFC

[LD04] P. Londrillo and L. Del Zanna. On the divergence-free condition in Godunov-type schemes for ideal magnetohydrodynamics: the upwind constrained transport method. Journal of Computational Physics, 195(1):17–48, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005102.

Li:2006:MRM

[LD06] Aiqin Li and Earl H. Dowell. Modal reduction of mathematical models of biological molecules. Journal of Computational Physics, 211(1):262–288, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002743.

Lee:2009:USM

[LD09a] Dongwook Lee and Anil E. Deane. An unsplit staggered mesh scheme for multidimensional magnetohydrodynamics. Journal of Computational Physics, 228(4):952–975, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004506.

Longo:2009:MCM

[LD09b] S. Longo and P. Diomede. A Monte Carlo model for seeded atomic flows in the transition regime. Journal of Computational Physics, 228(10):3851–3857, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000874.

Lifschitz:2009:PCM

[LDL^{+}09] A. F. Lifschitz, X. Davoine, E. Lefebvre, J. Faure, C. Rechatin, and V. Malka. Particle-in-cell modelling of laser-plasma interaction using Fourier decomposition. Journal of Computational Physics, 228(5):1803–1814, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005950.

Laval:2004:FNS

[LDN04] J.-P. Laval, B. Dubrulle, and S. V. Nazarenko. Fast numerical simulations of 2D turbulence using a dynamic model for subfilter motions. Journal of Computational Physics, 196(1):184– 207, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910300593X.

Liao:2008:TEM

[LDPL08] Wei Liao, Boris Diskin, Yan Peng, and Li-Shi Luo. Textbookefficiency multigrid solver for three-dimensional unsteady compressible Navier–Stokes equations. Journal of Computational Physics, 227(15):7160–7177, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001757.

Lodato:2008:TDB

[LDV08] Guido Lodato, Pascale Domingo, and Luc Vervisch. Threedimensional boundary conditions for direct and large-eddy simulation of compressible viscous flows. Journal of Computational Physics, 227(10):5105–5143, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000685.

Liang:2007:FSE

[LDW07] Dong Liang, Chuanbin Du, and Hong Wang. A fractional step ELLAM approach to high-dimensional convection-diffusion problems with forward particle tracking. Journal of Computational Physics, 221(1):198–225, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600283X.

Lee:2003:SCV

[Lee03] Changhoon Lee. Stability characteristics of the virtual boundary method in three-dimensional applications. Journal of Computational Physics, 184(2):559–591, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000384.

Lee:2005:CCP

[Lee05]  Sang-Hyeon Lee.  Convergence characteristics of pre-

conditioned Euler equations. Journal of Computational Physics, 208(1):266–288, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001099.

Lee:2007:ISP

[Lee07a] Kuo-Ming Lee. An inverse scattering problem from an impedance obstacle. Journal of Computational Physics, 227(1):431–439, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003397.

Lee:2007:CPP

[Lee07b] Sang-Hyeon Lee. Cancellation problem of preconditioning method at low Mach numbers. Journal of Computational Physics, 225(2):1199–1210, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001544.

Lee:2009:ACP

[Lee09] Sang-Hyeon Lee. Alleviation of cancellation problem of preconditioned Navier–Stokes equations. Journal of Computational Physics, 228(14):4970–4975, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002162.

Lemou:2000:NAA

[Lem00] M. Lemou. Numerical algorithms for axisymmetric Fokker– Planck–Landau operators. Journal of Computational Physics, 157(2):762–786, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199919996401X.

Lermusiaux:2006:UEP

[Ler06] Pierre F. J. Lermusiaux. Uncertainty estimation and prediction for interdisciplinary ocean dynamics. Journal of Computational Physics, 217(1):176–199, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000829.

Lorstad:2004:HOS

[LF04] Daniel Lörstad and Laszlo Fuchs. High-order surface tension VOF-model for 3D bubble flows with high density ratio. Journal of Computational Physics, 200(1):153–176, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104001482.

Linnick:2005:HOI

[LF05] Mark N. Linnick and Hermann F. Fasel. A high-order immersed interface method for simulating unsteady incompressible flows on irregular domains. Journal of Computational Physics, 204(1):157–192, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004127.

Li:2006:MGK

[LF06] Qibing Li and Song Fu. On the multidimensional gas-kinetic BGK scheme. Journal of Computational Physics, 220(1): 532–548, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106003469.

Liu:2000:BCC

[LFK00] Xu-Dong Liu, Ronald P. Fedkiw, and Myungjoo Kang. A boundary condition capturing method for Poisson’s equation on irregular domains. Journal of Computational Physics, 160(1):151–178, May 1, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964441.

Lerat:2007:VPS

[LFS07] A. Lerat, F. Falissard, and J. Sidès. Vorticity-preserving schemes for the compressible Euler equations. Journal of Computational Physics, 225(1):635–651, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006140.

Lankalapalli:2007:AFE

[LFSS07] S. Lankalapalli, J. E. Flaherty, M. S. Shephard, and H. Strauss. An adaptive finite element method for magnetohydrodynamics. Journal of Computational Physics, 225(1):363– 381, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910600595X.

Li:2005:CNS

[LFX05] Qibing Li, Song Fu, and Kun Xu. A compressible Navier– Stokes flow solver with scalar transport. Journal of Computational Physics, 204(2):692–714, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004334.

Lakkis:2003:AVM

[LG03a] Issam Lakkis and Ahmed F. Ghoniem. Axisymmetric vortex method for low-Mach number, diffusion-controlled combustion. Journal of Computational Physics, 184(2):435–475, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910200030X.

Li:2003:SCM

[LG03b] Jing-Rebecca Li and Leslie Greengard. Strongly consistent marching schemes for the wave equation. Journal of Computational Physics, 188(1):194–208, June 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001633.

Li:2004:HOM

[LG04] Jing-Rebecca Li and Leslie Greengard. High order marching schemes for the wave equation in complex geometry. Journal of Computational Physics, 198(1):295–309, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104000294.

Lee:2005:TNF

[LG05] June-Yub Lee and Leslie Greengard. The type 3 nonuniform FFT and its applications. Journal of Computational Physics, 206(1):1–5, June 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400511X.

Li:2007:NSH

[LG07] Jing-Rebecca Li and Leslie Greengard. On the numerical solution of the heat equation I: Fast solvers in free space. Journal of Computational Physics, 226(2):1891–1901, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002689.

Larsson:2008:SCH

[LG08] Johan Larsson and Bertil Gustafsson. Stability criteria for hybrid difference methods. Journal of Computational Physics, 227(5):2886–2898, February 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005189.

Lakkis:2009:HRS

[LG09] Issam Lakkis and Ahmed Ghoniem. A high resolution spatially adaptive vortex method for separating flows. Part I: Two-dimensional domains. Journal of Computational Physics, 228(2):491–515, February 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005093.

Lauter:2008:DGM

[LGHD08] Matthias Läuter, Francis X. Giraldo, Dörthe Handorf, and Klaus Dethloff. A discontinuous Galerkin method for the shallow water equations in spherical triangular coordinates. Journal of Computational Physics, 227(24):10226–10242, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004543.

Lin:2006:NSS

[LGK06] Guang Lin, Leopold Grinberg, and George Em Karniadakis. Numerical studies of the stochastic Korteweg–de Vries equation. Journal of Computational Physics, 213(2):676–703, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004122.

Lauber:2007:LLG

[LGKP07] Ph. Lauber, S. Günter, A. Könies, and S. D. Pinches. LIGKA: a linear gyrokinetic code for the description of background kinetic and fast particle effects on the MHD stability in tokamaks. Journal of Computational Physics, 226(1):447–465, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001660.

Lorcher:2008:EDG

[LGM08] Frieder Lörcher, Gregor Gassner, and Claus-Dieter Munz. An explicit discontinuous Galerkin scheme with local timestepping for general unsteady diffusion equations. Journal of Computational Physics, 227(11):5649–5670, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001046.

Li:2005:CAM

[LGN05] KePing Li, ZiYou Gao, and Bin Ning. Cellular automaton model for railway traffic. Journal of Computational Physics, 209(1):179–192, October 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001889.

Lampoudi:2009:EEV

[LGP09] Sotiria Lampoudi, Dan T. Gillespie, and Linda R. Petzold. Effect of excluded volume on 2D discrete stochastic chemical kinetics. Journal of Computational Physics, 228(10):3656–3668, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000631.

Langlands:2005:ASI

[LH05a] T. A. M. Langlands and B. I. Henry. The accuracy and stability of an implicit solution method for the fractional diffusion equation. Journal of Computational Physics, 205(2):719–736, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004887.

Lu:2005:WFS

[LH05b] Jian-Fei Lu and Andrzej Hanyga. Wave field simulation for heterogeneous porous media with singular memory drag force. Journal of Computational Physics, 208(2):651–674, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001245.

Lopez:2008:AGT

[LH08a] J. López and J. Hernández. Analytical and geometrical tools for 3D volume of fluid methods in general grids. Journal of Computational Physics, 227(12):5939–5948, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001629.

Loureiro:2008:ISI

[LH08b] N. F. Loureiro and G. W. Hammett. An iterative semiimplicit scheme with robust damping. Journal of Computational Physics, 227(9):4518–4542, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800034X.

Lauter:2005:UAS

[LHD05] Matthias Läuter, Dörthe Handorf, and Klaus Dethloff. Unsteady analytical solutions of the spherical shallow water equations. Journal of Computational Physics, 210(2):535–553, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002500.

Lopez:2004:VFM

[LHGF04] J. López, J. Hernández, P. Gómez, and F. Faura. A volume of fluid method based on multidimensional advection and spline interface reconstruction. Journal of Computational Physics, 195(2):718–742, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005801.

Lopez:2005:IPV

[LHGF05] J. López, J. Hernández, P. Gómez, and F. Faura. An improved PLIC–VOF method for tracking thin fluid structures in incompressible two-phase flows. Journal of Computational Physics, 208(1):51–74, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000513.

Lauter:2007:PAB

[LHR^{+}07] Matthias Läuter, Dörthe Handorf, Natalja Rakowsky, Jörn Behrens, Stephan Frickenhaus, Meike Best, Klaus Dethloff, and Wolfgang Hiller. A parallel adaptive barotropic model of the atmosphere. Journal of Computational Physics, 223(2):609–628, May 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004645.

Li:2005:ALE

[LHZW05] Jie Li, Marc Hesse, Johanna Ziegler, and Andrew W. Woods. An arbitrary Lagrangian Eulerian method for movingboundary problems and its application to jumping over water. Journal of Computational Physics, 208(1):289–314, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001130.

Li:2001:SPB

[Li01] Tsung-Lung Li. Simulation of the postexposure bake process of chemically amplified resists by reaction-diffusion equations. Journal of Computational Physics, 173(1):348–363, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968810.

Li:2005:HRS

[Li05] Shengtai Li. An HLLC Riemann solver for magnetohydrodynamics. Journal of Computational Physics, 203(1): 344–357, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003857.

Li:2008:HOC

[Li08a] Shengtai Li. High order central scheme on overlapping cells for magneto-hydrodynamic flows with and without constrained transport method. Journal of Computational Physics, 227(15):7368–7393, July 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002519.

Li:2008:VBC

[Li08b] Xiantao Li. Variational boundary conditions for molecular dynamics simulations: Treatment of the loading condition. Journal of Computational Physics, 227(24):10078–10093, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004300.

Ling:2001:SGO

[Lin01] Hong Y. Ling. Solution of generalized optical Bloch equations by the method of matrix continued fraction. Journal of Computational Physics, 171(1):264–271, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967865.

Lin:2002:NSQ

[Lin02] Kevin K. Lin. Numerical study of quantum resonances in chaotic scattering. Journal of Computational Physics, 176 (2):295–329, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101969864.

Liou:2000:MFS

[Lio00] Meng-Sing Liou. Mass flux schemes and connection to shock instability. Journal of Computational Physics, 160(2): 623–648, May 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100964787.

Liou:2006:SAP

[Lio06] Meng-Sing Liou. A sequel to AUSM, Part II: AUSM^{+}-up for all speeds. Journal of Computational Physics, 214(1):137–170, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004274.

Liu:2005:CSO

[Liu05] Yingjie Liu. Central schemes on overlapping cells. Journal of Computational Physics, 209(1):82–104, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001786.

Liu:2008:NSO

[Liu08] Di Liu. A numerical scheme for optimal transition paths of stochastic chemical kinetic systems. Journal of Computational Physics, 227(19):8672–8684, October 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003367.

Liu:2009:IMI

[Liu09a] Hao Liu. Integrated modeling of insect flight: From morphology, kinematics to aerodynamics. Journal of Computational Physics, 228(2):439–459, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005032.

Liu:2009:OTB

[Liu09b] Jie Liu. Open and traction boundary conditions for the incompressible Navier–Stokes equations. Journal of Computational Physics, 228(19):7250–7267, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003453.

Liu:2009:TLF

[Liu09c]  Weiming Liu.  A triple level finite element method

for large eddy simulations. Journal of Computational Physics, 228(7):2690–2706, April 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006505.

Livermore:2007:IET

[Liv07] Philip W. Livermore. An implementation of the exponential time differencing scheme to the magnetohydrodynamic equations in a spherical shell. Journal of Computational Physics, 220(2):824–838, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002622.

Lunati:2006:MFV

[LJ06] Ivan Lunati and Patrick Jenny. Multiscale finite-volume method for compressible multiphase flow in porous media. Journal of Computational Physics, 216(2):616–636, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000039.

Lou:2007:DFD

[LJ07] Zheng Lou and Jian-Ming Jin. A dual-field domaindecomposition method for the time-domain finite-element analysis of large finite arrays. Journal of Computational Physics, 222(1):408–427, March 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003597.

Li:2009:PIF

[LJ09a] Yu-Jia Li and Jian-Ming Jin. Parallel implementation of the FETI–DPEM algorithm for general 3D EM simulations. Journal of Computational Physics, 228(9):3255–3267, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900031X.

Loh:2009:MDD

[LJ09b] Ching Y. Loh and Philip C. E. Jorgenson. Multi-dimensional dissipation for cure of pathological behaviors of upwind scheme. Journal of Computational Physics, 228(5):1343–1346, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005664.

Li:2009:CTS

[LJK09] Peijun Li, Hans Johnston, and Robert Krasny. A Cartesian treecode for screened Coulomb interactions. Journal of Computational Physics, 228(10):3858–3868, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000916.

Lampe:2006:QPS

[LJM^{+}06] Martin Lampe, Glenn Joyce, Wallace M. Manheimer, Anatoly Streltsov, and Gurudas Ganguli. Quasineutral particle simulation technique for whistlers. Journal of Computational Physics, 214(1):284–298, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004353.

Liang:2008:CSS

[LJS08] An Liang, Xiaodong Jing, and Xiaofeng Sun. Constructing spectral schemes of the immersed interface method via a global description of discontinuous functions. Journal of Computational Physics, 227(18):8341–8366, September 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002957.

Lee:2008:FIN

[LJSM08] Jae Hoon Lee, Amit Joshi, and Eva M. Sevick-Muraca. Fast intersections on nested tetrahedrons (FINT): An algorithm for adaptive finite element based distributed parameter estimation. Journal of Computational Physics, 227(11):5778–5798, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001137.

Livermore:2007:SRB

[LJW07] Philip W. Livermore, Chris A. Jones, and Steven J. Worland. Spectral radial basis functions for full sphere computations. Journal of Computational Physics, 227(2):1209–1224, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003907.

Liang:2009:SDM

[LJW09] Chunlei Liang, Antony Jameson, and Z. J. Wang. Spectral difference method for compressible flow on unstructured grids with mixed elements. Journal of Computational Physics, 228(8):2847–2858, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000023.

Lindsay:2001:PMA

[LK01] Keith Lindsay and Robert Krasny. A particle method and adaptive treecode for vortex sheet motion in threedimensional flow. Journal of Computational Physics, 172(2):879–907, September 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968627.

LeMaitre:2007:SPM

[LK07] Olivier P. Le Maître and Omar M. Knio. A stochastic particlemesh scheme for uncertainty propagation in vortical flows. Journal of Computational Physics, 226(1):645–671, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002008.

Li:2009:MRA

[LK09] Yuanhong Li and Song-Charng Kong. Mesh refinement algorithms in an unstructured solver for multiphase flow simulation using discrete particles. Journal of Computational Physics, 228(17):6349–6360, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002782.

Laermans:2004:TDM

[LKD04] E. Laermans, L. Knockaert, and D. De Zutter. Twodimensional method of moments modelling of lossless overmoded transverse magnetic cavities. Journal of Computational Physics, 198(1):326–348, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000312.

Lehtovaara:2004:ENM

[LKE04] L. Lehtovaara, T. Kiljunen, and J. Eloranta. Efficient numerical method for simulating static and dynamic properties of superfluid helium. Journal of Computational Physics, 194(1):78–91, February 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004613.

Liu:2009:FCM

[LKMK09] D. Liu, E. E. Keaveny, M. R. Maxey, and G. E. Karniadakis. Force-coupling method for flows with ellipsoidal particles. Journal of Computational Physics, 228(10):3559–3581, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000369.

Liu:2005:SIC

[LKMU05] H. Liu, S. Krishnan, S. Marella, and H. S. Udaykumar. Sharp interface Cartesian grid method II: a technique for simulating droplet interactions with surfaces of arbitrary shape. Journal of Computational Physics, 210(1):32–54, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001968.

LeMaitre:2001:SPM

[LKNG01] Olivier P. Le Maître, Omar M. Knio, Habib N. Najm, and Roger G. Ghanem. A stochastic projection method for fluid flow: I. Basic formulation. Journal of Computational Physics, 173(2):481–511, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968895.

LeMaitre:2004:UPU

[LKNG04] O. P. Le Maître, O. M. Knio, H. N. Najm, and R. G. Ghanem. Uncertainty propagation using Wiener–Haar expansions. Journal of Computational Physics, 197(1):28–57, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006181.

Liu:2005:HPM

[LKO05] Jie Liu, Seiichi Koshizuka, and Yoshiaki Oka. A hybrid particle-mesh method for viscous, incompressible, multiphase flows. Journal of Computational Physics, 202(1):65–93, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104002682.

Le:2006:IIM

[LKP06] D. V. Le, B. C. Khoo, and J. Peraire. An immersed interface method for viscous incompressible flows involving rigid and flexible boundaries. Journal of Computational Physics, 220(1):109–138, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002245.

Liu:2005:GFM

[LKW05] T. G. Liu, B. C. Khoo, and C. W. Wang. The ghost fluid method for compressible gas-water simulation. Journal of Computational Physics, 204(1):193–221, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004139.

Liu:2004:IOF

[LKX04] T. G. Liu, B. C. Khoo, and W. F. Xie. Isentropic one-fluid modelling of unsteady cavitating flow. Journal of Computational Physics, 201(1):80–108, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002025.

Liu:2003:GFM

[LKY03] T. G. Liu, B. C. Khoo, and K. S. Yeo. Ghost fluid method for strong shock impacting on material interface. Journal of Computational Physics, 190(2):651–681, September 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003012.

Langseth:2000:WPM

[LL00] Jan Olav Langseth and Randall J. LeVeque. A wave propagation method for three-dimensional hyperbolic conservation laws. Journal of Computational Physics, 165(1):126– 166, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100966063.

Lee:2001:CGM

[LL01a] Taehun Lee and Ching-Long Lin. A characteristic Galerkin method for discrete Boltzmann equation. Journal of Computational Physics, 171(1):336–356, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967919.

Li:2001:IIM

[LL01b] Zhilin Li and Ming-Chih Lai. The immersed interface method for the Navier–Stokes equations with singular forces. Journal of Computational Physics, 171(2):822–842, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968135.

Lallemand:2003:LBM

[LL03a] Pierre Lallemand and Li-Shi Luo. Lattice Boltzmann method for moving boundaries. Journal of Computational Physics, 184(2):406–421, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000220.

Lee:2003:EDS

[LL03b] Taehun Lee and Ching-Long Lin. An Eulerian description of the streaming process in the lattice Boltzmann equation. Journal of Computational Physics, 185(2):445–471, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000657.

Liu:2003:PSS

[LL03c] Xu-Dong Liu and Peter D. Lax. Positive schemes for solving multi-dimensional hyperbolic systems of conservation laws II. Journal of Computational Physics, 187(2):428–440, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001001.

Leriche:2004:SES

[LL04a] E. Leriche and G. Labrosse. Stokes eigenmodes in square domain and the stream function-vorticity correlation. Journal of Computational Physics, 200(2):489–511, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001779.

Li:2004:NAD

[LL04b] Shengtai Li and Hui Li. A novel approach of divergencefree reconstruction for adaptive mesh refinement. Journal of Computational Physics, 199(1):1–15, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000622.

Lee:2005:SDL

[LL05] Taehun Lee and Ching-Long Lin. A stable discretization of the lattice Boltzmann equation for simulation of incompressible two-phase flows at high density ratio. Journal of Computational Physics, 206(1):16–47, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004929.

Lin:2006:SSD

[LL06a] Ping Lin and Chun Liu. Simulations of singularity dynamics in liquid crystal flows: a C^{0} finite element approach. Journal of Computational Physics, 215(1):348–362, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004912.

Liu:2006:RRB

[LL06b] Yaling Liu and Wing Kam Liu. Rheology of red blood cell aggregation by computer simulation. Journal of Computational Physics, 220(1):139–154, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002257.

Liovic:2007:MPT

[LL07] Petar Liovic and Djamel Lakehal. Multi-physics treatment in the vicinity of arbitrarily deformable gas-liquid interfaces. Journal of Computational Physics, 222(2):504–535, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003640.

Liu:2008:NST

[LL08a] Dongming Liu and Pengzhi Lin. A numerical study of three-dimensional liquid sloshing in tanks. Journal of Computational Physics, 227(8):3921–3939, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005414.

Lutjens:2008:XCN

[LL08b] Hinrich Lütjens and Jean-François Luciani. The XTOR code for nonlinear 3D simulations of MHD instabilities in tokamak plasmas. Journal of Computational Physics, 227(14):6944–6966, July 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002064.

Laizet:2009:HOC

[LL09] Sylvain Laizet and Eric Lamballais. High-order compact schemes for incompressible flows: a simple and efficient method with quasi-spectral accuracy. Journal of Computational Physics, 228(16):5989–6015, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002587.

Lohner:2005:SER

[LLB05] Rainald Löhner, Hong Luo, and Joseph D. Baum. Selective edge removal for unstructured grids with Cartesian cores. Journal of Computational Physics, 206(1):208–226, June 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104005091.

Lee:2006:LBA

[LLC06] Taehun Lee, Ching-Long Lin, and Lea-Der Chen. A lattice Boltzmann algorithm for calculation of the laminar jet diffusion flame. Journal of Computational Physics, 215(1): 133–152, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105004821.

Luo:2008:SIL

[LLC^{+}08] Junzhao Luo, Zhen Luo, Liping Chen, Liyong Tong, and Michael Yu Wang. A semi-implicit level set method for structural shape and topology optimization. Journal of Computational Physics, 227(11):5561–5581, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000995.

Liao:2000:LSB

[LLdlP^{+}00] Guojun Liao, Feng Liu, Gary C. de la Pena, Danping Peng, and Stanley Osher. Level-set-based deformation methods for adaptive grids. Journal of Computational Physics, 159(1):103–122, March 20, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964325.

Liu:2007:FTA

[LLGL07] Xinfeng Liu, Yuanhua Li, J. Glimm, and X. L. Li. A front tracking algorithm for limited mass diffusion. Journal of Computational Physics, 222(2):644–653, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003718.

Lan:2002:AFV

[LLH02] C. W. Lan, C. C. Liu, and C. M. Hsu. An adaptive finite volume method for incompressible heat flow problems in solidification. Journal of Computational Physics, 178(2):464–497, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970373.

LeMaitre:2001:CMP

[LLIK01a] Olivier Le Maître, Julia Levin, Mohamed Iskandarani, and Omar M. Knio. Corrigendum: “A Multiscale Pressure Splitting of the Shallow-Water Equations. I. Formulation and 1D Tests”: Volume 166, Number 1 (2001), pages 116–151. Journal of Computational Physics, 167(2):501, March 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967178. See [LLIK01b].

LeMaitre:2001:MPS

[LLIK01b] Olivier Le Maître, Julia Levin, Mohamed Iskandarani, and Omar M. Knio. A multiscale pressure splitting of the shallow-water equations: I. Formulation and 1D tests. Journal of Computational Physics, 166(1):116–151, January 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101966471. See corrigen-

dum [LLIK01a].

Li:2007:RSA

[LLL07] Shuwang Li, John S. Lowengrub, and Perry H. Leo. A rescaling scheme with application to the long-time simulation of viscous fingering in a Hele–Shaw cell. Journal of Computational Physics, 225(1):554–567, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006073.

Leo:2000:MEO

[LLN00] P. H. Leo, J. S. Lowengrub, and Qing Nie. Microstructural evolution in orthotropic elastic media. Journal of Computational Physics, 157(1):44–88, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963593.

Liang:2006:SMM

[LLOT06] Kewei Liang, Ping Lin, Ming Tze Ong, and Roger C. E. Tan. A splitting moving mesh method for reaction-diffusion equations of quenching type. Journal of Computational Physics, 215(2):757–777, July 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005383.

Lallemand:2007:LBF

[LLP07] Pierre Lallemand, Li-Shi Luo, and Yan Peng. A lattice Boltzmann front-tracking method for interface dynamics with surface tension in two dimensions. Journal of Computational Physics, 226(2):1367–1384, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002380.

Lu:2002:LES

[LLQ^{+}02] Zhenyu Lu, Ying Liao, Dongying Qian, J. B. McLaughlin, J. J. Derksen, and K. Kontomaris. Large eddy simulations of a stirred tank using the lattice Boltzmann method on a nonuniform grid. Journal of Computational Physics, 181(2):675–704, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971512.

Lu:2009:CEI

[LLRP09] Liuyan Lu, Steven R. Lantz, Zhuyin Ren, and Stephen B. Pope. Computationally efficient implementation of combustion chemistry in parallel PDF calculations. Journal of Computational Physics, 228(15):5490–5525, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002277.

Li:2009:IGS

[LLS09] Jiequan Li, Tiegang Liu, and Zhongfeng Sun. Implementation of the GRP scheme for computing radially symmetric compressible fluid flows. Journal of Computational Physics, 228(16):5867–5887, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002411.

Liu:2007:ALF

[LLTA07] Q. Liu, F. Liu, I. Turner, and V. Anh. Approximation of the Lévy–Feller advection-dispersion process by random walk and finite difference method. Journal of Computational Physics, 222(1):57–70, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002828.

Larsson:2005:CSM

[LLY05] J. Larsson, F. S. Lien, and E. Yee. Conditional semicoarsening multigrid algorithm for the Poisson equation on anisotropic grids. Journal of Computational Physics, 208 (1):368–383, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105001166.

Lin:2007:ELP

[LLZ07] Ping Lin, Chun Liu, and Hui Zhang. An energy law preserving C^{0} finite element scheme for simulating the kinematic effects in liquid crystal dynamics. Journal of Computational Physics, 227(2):1411–1427, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004007.

Li:2001:EPT

[LM01] Genong Li and Michael F. Modest. An effective particle tracing scheme on structured/unstructured grids in hybrid finite volume/PDF Monte Carlo methods. Journal of Computational Physics, 173(1):187–207, October 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968718. See comment [MCP03]

and reply [LM01].

Li:2003:RCA

[LM03a] G. Li and M. F. Modest. Reply to the comment on the article “An effective particle tracing scheme on structured/unstructured grids in hybrid Finite Volume/PDF Monte Carlo methods” by G. Li and M. F. Modest [J. Comp. Phys. 173 (2001) 187–207]. Journal of Computational Physics, 186(1):359, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300007X. See [LM01].

Lomholt:2003:FCM

[LM03b] Sune Lomholt and Martin R. Maxey. Force-coupling method for particulate two-phase flow: Stokes flow. Journal of Computational Physics, 184(2):381–405, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000219.

Layton:2004:CMI

[LM04] Anita T. Layton and Michael L. Minion. Conservative multiimplicit spectral deferred correction methods for reacting gas dynamics. Journal of Computational Physics, 194(2):697–715, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005023.

LeFloch:2008:WMT

[LM08a] Philippe G. LeFloch and Majid Mohammadian. Why many theories of shock waves are necessary: Kinetic functions, equivalent equations, and fourth-order models. Journal of Computational Physics, 227(8):4162–4189, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005694.

Lejay:2008:CPE

[LM08b] Antoine Lejay and Sylvain Maire. Computing the principal eigenelements of some linear operators using a branching Monte Carlo method. Journal of Computational Physics, 227(23):9794–9806, December 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004075.

Liu:2008:IFE

[LM08c] Weiming Liu and Georgy Makhviladze. An implicit finite element solution of thermal flows at low Mach number. Journal of Computational Physics, 227(5):2743–2757, February 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004743.

Liu:2007:DPD

[LMH07] Moubin Liu, Paul Meakin, and Hai Huang. Dissipative particle dynamics simulation of fluid motion through an unsaturated fracture and fracture junction. Journal of Computational Physics, 222(1):110–130, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003202.

Lang:2003:ATE

[LMK03] Michael Lang, Wolfgang Michalke, and Stefan Kreitmeier. Analysis of trapped entanglements in polymer networks. Journal of Computational Physics, 185(2):549–561, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000032.

Luo:2009:SPM

[LMK09] Xian Luo, Martin R. Maxey, and George Em Karniadakis. Smoothed profile method for particulate flows: Error analysis and simulations. Journal of Computational Physics, 228(5):1750–1769, March 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005925.

Layton:2009:ARF

[LMN^{+}09] William Layton, Carolina C. Manica, Monika Neda, Maxim Olshanskii, and Leo G. Rebholz. On the accuracy of the rotation form in simulations of the Navier–Stokes equations. Journal of Computational Physics, 228(9):3433–3447, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000540.

Lukacova-Medvidova:2007:WBF

[LMNK07] M. Lukácová-Medvid’ová, S. Noelle, and M. Kraft. Wellbalanced finite volume evolution Galerkin methods for the shallow water equations. Journal of Computational Physics, 221(1):122–147, January 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002786.

Lopez:2002:ESP

[LMS02] J. M. Lopez, F. Marques, and Jie Shen. An efficient spectralprojection method for the Navier–Stokes equations in cylindrical geometries: II. Three-dimensional cases. Journal of Computational Physics, 176(2):384–401, March 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102969937.

Lipnikov:2004:MFD

[LMS04] Konstantin Lipnikov, Jim Morel, and Mikhail Shashkov. Mimetic finite difference methods for diffusion equations on non-orthogonal non-conformal meshes. Journal of Computational Physics, 199(2):589–597, September 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000968.

LeMetayer:2005:MEF

[LMS05] O. Le Métayer, J. Massoni, and R. Saurel. Modelling evaporation fronts with reactive Riemann solvers. Journal of Computational Physics, 205(2):567–610, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004814.

Lipnikov:2008:MMM

[LMS08] K. Lipnikov, J. D. Moulton, and D. Svyatskiy. A multilevel multiscale mimetic (M^{3}) method for two-phase flows in porous media. Journal of Computational Physics, 227(14):6727–6753, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001654.

Legrand:2000:WAW

[LMSV00] Olivier Legrand, Fabrice Mortessagne, Patrick Sebbah, and Christian Vanneste. A wave automaton for wave propagation in inhomogeneous anisotropic media. Journal of Computational Physics, 160(2):541–563, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964714.

Lukacova-Medvidova:2002:FVE

[LMSW02] M. Lukácová-Medvid’ová, J. Saibertová, and G. Warnecke. Finite volume evolution Galerkin methods for nonlinear hyperbolic systems. Journal of Computational Physics, 183(2):533–562, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102972074.

Laurent:2004:EMF

[LMV04] Frédérique Laurent, Marc Massot, and Philippe Villedieu. Eulerian multi-fluid modeling for the numerical simulation of coalescence in polydisperse dense liquid sprays. Journal of Computational Physics, 194(2):505–543, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004935.

Lee:2008:CWC

[LMX^{+}08] E.-S. Lee, C. Moulinec, R. Xu, D. Violeau, D. Laurence, and P. Stansby. Comparisons of weakly compressible and truly incompressible algorithms for the SPH mesh free particle method. Journal of Computational Physics, 227(18):8417–8436, September 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800315X.

Luo:2008:IBM

[LMZ^{+}08] Haoxiang Luo, Rajat Mittal, Xudong Zheng, Steven A. Bielamowicz, Raymond J. Walsh, and James K. Hahn. An immersed-boundary method for flow-structure interaction in biological systems with application to phonation. Journal of Computational Physics, 227(22):9303–9332, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002623.

Li:2009:HOB

[LN09] Xiaofan Li and Qing Nie. A high-order boundary integral method for surface diffusions on elastically stressed axisymmetric rods. Journal of Computational Physics, 228 (12):4625–4637, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001594.

LeMaitre:2004:MRA

[LNGK04] O. P. Le Maître, H. N. Najm, R. G. Ghanem, and O. M. Knio. Multi-resolution analysis of Wiener-type uncertainty propagation schemes. Journal of Computational Physics, 197(2):502– 531, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006491.

Liu:2009:NGL

[LNXNTX09] G. R. Liu, H. Nguyen-Xuan, T. Nguyen-Thoi, and X. Xu. A novel Galerkin-like weakform and a superconvergent alpha finite element method (S α FEM) for mechanics problems using triangular meshes. Journal of Computational Physics, 228(11):4055–4087, June 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000928.

Lohner:2004:MEA

[Löh04] Rainald Löhner. Multistage explicit advective prediction for projection-type incompressible flow solvers. Journal of Computational Physics, 195(1):143–152, March 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300514X.

Lee:2001:CFS

[LOK01] Moon J. Lee, Byung Do Oh, and Young Bae Kim. Canonical fractional-step methods and consistent boundary conditions for the incompressible Navier–Stokes equations. Journal of Computational Physics, 168(1):73–100, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966828.

Liu:2005:NDF

[LOK05] Junhui Liu, Elaine S. Oran, and Carolyn R. Kaplan. Numerical diffusion in the FCT algorithm, revisited. Journal of Computational Physics, 208(2):416–434, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000896.

Lou:2000:EST

[Lou00] L. Lou. Exploring sparsity in three-dimensional integration for density-functional calculations. Journal of Computational Physics, 157(1):404–416, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919996388X.

Lowrie:2004:CIT

[Low04] Robert B. Lowrie. A comparison of implicit time integration methods for nonlinear relaxation and diffusion. Journal of Computational Physics, 196(2):566–590, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006089.

Lowe:2005:TPS

[Low05] C. A. Lowe. Two-phase shock-tube problems and numerical methods of solution. Journal of Computational Physics, 204(2):598–632, April 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004309.

Lai:2000:IBM

[LP00] Ming-Chih Lai and Charles S. Peskin. An immersed boundary method with formal second-order accuracy and reduced numerical viscosity. Journal of Computational Physics, 160(2):705–719, May 20, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964830.

LeVeque:2001:CAR

[LP01] Randall J. LeVeque and Marica Pelanti. A class of approximate Riemann solvers and their relation to relaxation schemes. Journal of Computational Physics, 172(2):572–591, September 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196838X.

Lopes:2002:NSI

[LP02] A. Silva Lopes and J. M. L. M. Palma. Numerical simulation of isotropic turbulence using a collocated approach and a nonorthogonal grid system. Journal of Computational Physics, 175(2):713–738, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969645.

Li:2004:ASA

[LP04a] Shengtai Li and Linda Petzold. Adjoint sensitivity analysis for time-dependent partial differential equations with adaptive mesh refinement. Journal of Computational Physics, 198(1):310–325, July 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000300.

Lombard:2004:NTT

[LP04b] Bruno Lombard and Joël Piraux. Numerical treatment of twodimensional interfaces for acoustic and elastic waves. Journal of Computational Physics, 195(1):90–116, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005126.

Lessani:2006:TAC

[LP06a] Bamdad Lessani and Miltiadis V. Papalexandris. Timeaccurate calculation of variable density flows with strong temperature gradients and combustion. Journal of Computational Physics, 212(1):218–246, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003189.

Liu:2006:ICL

[LP06b] Kunlun Liu and Richard H. Pletcher. Inflow conditions for the large eddy simulation of turbulent boundary layers: a dynamic recycling procedure. Journal of Computational Physics, 219(1):1–6, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001963.

Lau:2007:MSM

[LP07a] Stephen R. Lau and Richard H. Price. Multidomain spectral method for the helically reduced wave equation. Journal of Computational Physics, 227(2):1126–1161, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003828.

Liu:2007:FSM

[LP07b] Kunlun Liu and Richard H. Pletcher. A fractional step method for solving the compressible Navier–Stokes equations. Journal of Computational Physics, 226(2):1930–1951, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002707.

Leutbecher:2008:EF

[LP08] M. Leutbecher and T. N. Palmer. Ensemble forecasting. Journal of Computational Physics, 227(7):3515–3539, March 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000812.

Lu:2009:IAS

[LP09] Liuyan Lu and Stephen B. Pope. An improved algorithm for in situ adaptive tabulation. Journal of Computational Physics, 228(2):361–386, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800483X.

Lee:2005:WNC

[LPK05] Sang-Ho Lee, Kim Palmo, and Samuel Krimm. WIG-GLE: a new constrained molecular dynamics algorithm in Cartesian coordinates. Journal of Computational Physics, 210(1):171–182, November 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002056.

Luchini:2006:LCP

[LQ06] Paolo Luchini and Maurizio Quadrio. A low-cost parallel implementation of direct numerical simulation of wall turbulence. Journal of Computational Physics, 211(2):551–571, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002871.

Leung:2009:EGB

[LQ09] Shingyu Leung and Jianliang Qian. Eulerian Gaussian beams for Schrödinger equations in the semi-classical regime. Journal of Computational Physics, 228(8):2951–2977, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000114.

Liu:2006:FDG

[LQX06] Xueqiang Liu, Ning Qin, and Hao Xia. Fast dynamic grid deformation based on Delaunay graph mapping. Journal of Computational Physics, 211(2):405–423, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002809.

Leimkuhler:2001:RAI

[LR01a] Ben Leimkuhler and Sebastian Reich. A reversible averaging integrator for multiple time-scale dynamics. Journal of Computational Physics, 171(1):95–114, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967749.

Lu:2001:RBW

[LR01b] Huei-Iin Lu and Franklin R. Robertson. Retrieving the balanced winds on the globe as a generalized inverse problem. Journal of Computational Physics, 170(1):299–319, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967385.

Lockerby:2003:HRB

[LR03] Duncan A. Lockerby and Jason M. Reese. High-resolution Burnett simulations of micro Couette flow and heat transfer. Journal of Computational Physics, 188(2):333–347, July 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001621.

Lin:2007:AHM

[LR07] Ping Lin and Thomas Richter. An adaptive homotopy multi-grid method for molecule orientations of high dimensional liquid crystals. Journal of Computational Physics, 225(2):2069–2082, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001209.

Lester:2008:GPS

[LRMB08] D. R. Lester, M. Rudman, G. Metcalfe, and H. M. Blackburn. Global parametric solutions of scalar transport. Journal of Computational Physics, 227(6):3032–3057, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004603.

LeMaitre:2002:SPM

[LRN^{+}02] Olivier P. Le Maître, Matthew T. Reagan, Habib N. Najm, Roger G. Ghanem, and Omar M. Knio. A stochastic projection method for fluid flow: II. random process. Journal of Computational Physics, 181(1):9–44, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971044.

Lelievre:2007:CFE

[LRS07] Tony Lelièvre, Mathias Rousset, and Gabriel Stoltz. Computation of free energy differences through nonequilibrium stochastic dynamics: The reaction coordinate case. Journal of Computational Physics, 222(2):624–643, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003706.

Love:2009:SAP

[LRS09] E. Love, W. J. Rider, and G. Scovazzi. Stability analysis of a predictor/multi-corrector method for staggered-grid Lagrangian shock hydrodynamics. Journal of Computational Physics, 228(20):7543–7564, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003635.

Liu:2004:CFS

[LRZ04] Miao’er Liu, Yu-Xin Ren, and Hanxin Zhang. A class of fully second order accurate projection methods for solving the incompressible Navier–Stokes equations. Journal of Computational Physics, 200(1):325–346, October 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400169X.

Liu:2000:HOD

[LS00] Jian-Guo Liu and Chi-Wang Shu. A high-order discontinuous Galerkin method for 2D incompressible flows. Journal of Computational Physics, 160(2):577–596, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964751.

Labourasse:2002:RTF

[LS02a] E. Labourasse and P. Sagaut. Reconstruction of turbulent fluctuations using a hybrid RANS/LES approach. Journal of Computational Physics, 182(1):301–336, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910297169X.

Lappa:2002:ACM

[LS02b] M. Lappa and R. Savino. 3D analysis of crystal/melt interface shape and Marangoni flow instability in solidifying liquid bridges. Journal of Computational Physics, 180(2):751–774, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971317.

Lee:2002:NCS

[LS02c] Changhoon Lee and Youngchwa Seo. A new compact spectral scheme for turbulence simulations. Journal of Computational Physics, 183(2):438–469, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972013.

Layton:2003:SLD

[LS03] Anita T. Layton and William F. Spotz. A semi-Lagrangian double Fourier method for the shallow water equations on the sphere. Journal of Computational Physics, 189(1):180–196, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002079.

Lin:2005:ADR

[LS05a] R. K. Lin and Tony W. H. Sheu. Application of dispersionrelation-preserving theory to develop a two-dimensional convection-diffusion scheme. Journal of Computational Physics, 208(2):493–526, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001191.

Loubere:2005:SRM

[LS05b] Raphaël Loubère and Mikhail J. Shashkov. A subcell remapping method on staggered polygonal grids for arbitrary-Lagrangian–Eulerian methods. Journal of Computational Physics, 209(1):105–138, October 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105001841.

Li:2007:RGR

[LS07] Jiequan Li and Zhongfeng Sun. Remark on the generalized Riemann problem method for compressible fluid flows. Journal of Computational Physics, 222(2):796–808, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003950.

Li:2008:FTS

[LS08] Xiaoyi Li and Kausik Sarkar. Front tracking simulation of deformation and buckling instability of a liquid capsule enclosed by an elastic membrane. Journal of Computational Physics, 227(10):4998–5018, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000569.

Liu:2009:NTS

[LS09] Yang Liu and Mrinal K. Sen. A new time-space domain highorder finite-difference method for the acoustic wave equation. Journal of Computational Physics, 228(23):8779–8806, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004707.

Lazic:2006:RHM

[LSA06] Predrag Lazíc, Hrvoje Stefancíc, and Hrvoje Abraham. The Robin hood method — a novel numerical method for electrostatic problems based on a non-local charge transfer. Journal of Computational Physics, 213(1):117–140, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003712.

Lacor:2004:FVF

[LSB04] Chris Lacor, Sergey Smirnov, and Martine Baelmans. A finite volume formulation of compact central schemes on arbitrary structured grids. Journal of Computational Physics, 198(2):535–566, August 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000518.

Latini:2007:EWF

[LSD07] Marco Latini, Oleg Schilling, and Wai Sun Don. Effects of WENO flux reconstruction order and spatial resolution on reshocked two-dimensional Richtmyer–Meshkov instability. Journal of Computational Physics, 221(2):805–836, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106003196.

Li:2005:LFM

[LSJA05] Q. Li, J. Soric, T. Jarak, and S. N. Atluri. A lockingfree meshless local Petrov–Galerkin formulation for thick and thin plates. Journal of Computational Physics, 208 (1):116–133, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105000872.

Lin:2006:PSD

[LSK06] G. Lin, C.-H. Su, and G. E. Karniadakis. Predicting shock dynamics in the presence of uncertainties. Journal of Computational Physics, 217(1):260–276, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001197.

Lan:2008:MRT

[LSL08] Xudong Lan, Jun Sun, and Zhixin Li. Modified relaxation time Monte Carlo method for continuum-transition gas flows. Journal of Computational Physics, 227(9):4763–4775, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000491.

Lipnikov:2006:MFD

[LSS06] Konstantin Lipnikov, Mikhail Shashkov, and Daniil Svyatskiy. The mimetic finite difference discretization of diffusion problem on unstructured polyhedral meshes. Journal of Computational Physics, 211(2):473–491, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002834.

Lin:2009:PPA

[LSS^{+}09] Paul T. Lin, John N. Shadid, Marzio Sala, Raymond S. Tuminaro, Gary L. Hennigan, and Robert J. Hoekstra. Performance of a parallel algebraic multilevel preconditioner for stabilized finite element semiconductor device modeling. Journal of Computational Physics, 228(17):6250–6267, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002733.

Lipnikov:2007:MFV

[LSSV07] K. Lipnikov, M. Shashkov, D. Svyatskiy, and Yu. Vassilevski. Monotone finite volume schemes for diffusion equations on unstructured triangular and shape-regular polygonal meshes. Journal of Computational Physics, 227(1):492–512, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003440.

Lipnikov:2009:IFM

[LSV09] K. Lipnikov, D. Svyatskiy, and Y. Vassilevski. Interpolationfree monotone finite volume method for diffusion equations on polygonal meshes. Journal of Computational Physics, 228(3):703–716, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005056.

Loubere:2006:RPN

[LSW06] Raphaël Loubère, Martin Staley, and Burton Wendroff. The repair paradigm: New algorithms and applications to compressible flow. Journal of Computational Physics, 211(2): 385–404, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002706.

Loubere:2008:VCS

[LSW08] R. Loubère, M. Shashkov, and B. Wendroff. Volume consistency in a staggered grid Lagrangian hydrodynamics scheme. Journal of Computational Physics, 227(8):3731–3737, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000223.

Levy:2004:LDG

[LSY04] Doron Levy, Chi-Wang Shu, and Jue Yan. Local discontinuous Galerkin methods for nonlinear dispersive equations. Journal of Computational Physics, 196(2):751–772, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006156.

Li:2008:SOD

[LSZZ08] Fengyan Li, Chi-Wang Shu, Yong-Tao Zhang, and Hongkai Zhao. A second order discontinuous Galerkin fast sweeping method for eikonal equations. Journal of Computational Physics, 227(17):8191–8208, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003082.

Lai:2005:FIS

[LT05] Ming-Chih Lai and Yu-Hou Tseng. A fast iterative solver for the variable coefficient diffusion equation on a disk. Journal of Computational Physics, 208(1):196–205, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000926.

Lin:2009:ADO

[LT09a] Zhili Lin and Lars Thylén. An analytical derivation of the optimum source patterns for the pseudospectral time-domain method. Journal of Computational Physics, 228(19):7375–7387, October 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003659.

Linton:2009:OTD

[LT09b] C. M. Linton and I. Thompson. One- and two-dimensional lattice sums for the three-dimensional Helmholtz equation. Journal of Computational Physics, 228(6):1815–1829, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005962.

Liu:2007:ERK

[LTC07] Jiangguo Liu, Simon Tavener, and Hongsen Chen. EL-LAM for resolving the kinematics of two-dimensional resistive magnetohydrodynamic flows. Journal of Computational Physics, 227(2):1372–1386, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003981.

Li:2004:WOS

[LTD04] Tingqiu Li, Peter Troch, and Julien De Rouck. Wave overtopping over a sea dike. Journal of Computational Physics, 198(2):686–726, August 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000592.

Liu:2006:CCH

[LTD^{+}06] L. Liu, J. P. Thomas, E. H. Dowell, P. Attar, and K. C. Hall. A comparison of classical and high dimensional harmonic balance approaches for a Duffing oscillator. Journal of Computational Physics, 215(1):298–320, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004894.

Li:2007:IBW

[LTD07] Tingqiu Li, Peter Troch, and Julien De Rouck. Interactions of breaking waves with a current over cut cells. Journal of Computational Physics, 223(2):865–897, May 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004785.

Lehtovaara:2007:STI

[LTE07] L. Lehtovaara, J. Toivanen, and J. Eloranta. Solution of time-independent Schrödinger equation by the imaginary time propagation method. Journal of Computational Physics, 221(1):148–157, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002798.

Lai:2008:IBM

[LTH08] Ming-Chih Lai, Yu-Hau Tseng, and Huaxiong Huang. An immersed boundary method for interfacial flows with insoluble surfactant. Journal of Computational Physics, 227 (15):7279–7293, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108002283.

Larsen:2002:SAE

[LTK^{+}02] Edward W. Larsen, Guido Thömmes, Axel Klar, Mohammed Seaïd, and Thomas Götz. Simplified P_{N} approximations to the equations of radiative heat transfer and applications. Journal of Computational Physics, 183(2):652–675, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102972104.

Luo:2009:DPA

[LTL^{+}09] Zhen Luo, Liyong Tong, Junzhao Luo, Peng Wei, and Michael Yu Wang. Design of piezoelectric actuators using a multiphase level set method of piecewise constants. Journal of Computational Physics, 228(7):2643–2659, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006633.

Luo:2009:STO

[LTM09] Zhen Luo, Liyong Tong, and Haitao Ma. Shape and topology optimization for electrothermomechanical microactuators using level set methods. Journal of Computational Physics, 228(9):3173–3181, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000084.

Luo:2007:STO

[LTWW07] Zhen Luo, Liyong Tong, Michael Yu Wang, and Shengyin Wang. Shape and topology optimization of compliant mechanisms using a parameterization level set method. Journal of Computational Physics, 227(1):680–705, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003531.

Li:2001:MMM

[LTZ01] Ruo Li, Tao Tang, and Pingwen Zhang. Moving mesh methods in multiple dimensions based on harmonic maps. Journal of Computational Physics, 170(2):562–588, July 1, 2001.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196749X.

Li:2002:MMF

[LTZ02] Ruo Li, Tao Tang, and Pingwen Zhang. A moving mesh finite element algorithm for singular problems in two and three space dimensions. Journal of Computational Physics, 177(2):365–393, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970026.

Li:2003:PBM

[LTZ03] Liyong Li, Hamdi A. Tchelepi, and Dongxiao Zhang. Perturbation-based moment equation approach for flow in heterogeneous porous media: applicability range and analysis of high-order terms. Journal of Computational Physics, 188(1):296–317, June 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001864.

Lurati:2007:PGS

[Lur07] Laura B. Lurati. Padé-Gegenbauer suppression of Runge phenomenon in the diagonal limit of Gegenbauer approximations. Journal of Computational Physics, 222(1):1–8, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003056.

Liu:2007:BPM

[LV07] Qianlong Liu and Oleg V. Vasilyev. A Brinkman penalization method for compressible flows in complex geometries. Journal of Computational Physics, 227(2):946–966, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003403.

Lee:2005:NOD

[LVL05] Seung-Cheol Lee, Marinos N. Vouvakis, and Jin-Fa Lee. A non-overlapping domain decomposition method with nonmatching grids for modeling large finite antenna arrays. Journal of Computational Physics, 203(1):1–21, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003158.

Liu:2006:SDM

[LVW06a] Yen Liu, Marcel Vinokur, and Z. J. Wang. Spectral difference method for unstructured grids I: Basic formulation. Journal of Computational Physics, 216(2):780–801, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000106.

Liu:2006:SFV

[LVW06b] Yen Liu, Marcel Vinokur, and Z. J. Wang. Spectral (finite) volume method for conservation laws on unstructured grids V: Extension to three-dimensional systems. Journal of Computational Physics, 212(2):454–472, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003281.

Liu:2001:EPM

[LW01] Jian-Guo Liu and Wei-Cheng Wang. An energy-preserving MAC–Yee scheme for the incompressible MHD equation. Journal of Computational Physics, 174(1):12–37, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967725.

Liu:2004:EHP

[LW04] Jian-Guo Liu and Wei-Cheng Wang. Energy and helicity preserving schemes for hydro- and magnetohydrodynamics flows with symmetry. Journal of Computational Physics, 200(1):8–33, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001287.

LeBars:2006:SBA

[LW06] Michael Le Bars and M. Grae Worster. Solidification of a binary alloy: Finite-element, single-domain simulation and new benchmark solutions. Journal of Computational Physics, 216(1):247–263, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005504.

Liu:2007:FSB

[LW07] Hailiang Liu and Zhongming Wang. A field-space-based level set method for computing multi-valued solutions to 1D Euler–Poisson equations. Journal of Computational Physics, 225(1):591–614, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006103.

Liu:2009:BBB

[LW09] Hailiang Liu and Zhongming Wang. A Bloch band based level set method for computing the semiclassical limit of Schrödinger equations. Journal of Computational Physics, 228(9):3326–3344, May 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000382.

Lemons:2009:SAC

[LWDA09] Don S. Lemons, Dan Winske, William Daughton, and Brian Albright. Small-angle Coulomb collision model for particle-in-cell simulations. Journal of Computational Physics, 228(5):1391–1403, March 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005615.

Lu:2000:FET

[LWEM00] Mingyu Lu, Jianguo Wang, A. Arif Ergin, and Eric Michielssen. Fast evaluation of two-dimensional transient wave fields. Journal of Computational Physics, 158(2): 161–185, March 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199964070.

Liu:2008:EBS

[LWF^{+}08] Bin Liu, Jizeng Wang, Xiaojun Fan, Yong Kong, and Huajian Gao. An effective bead-spring model for polymer simulation.

Journal of Computational Physics, 227(5):2794–2807, February 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005049.

Lee:2003:ARD

[LWG03] Steven L. Lee, Carol S. Woodward, and Frank Graziani. Analyzing radiation diffusion using time-dependent sensitivitybased techniques. Journal of Computational Physics, 192(1):211–230, November 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003723.

Le:2009:IIB

[LWP^{+}09] D. V. Le, J. White, J. Peraire, K. M. Lim, and B. C. Khoo. An implicit immersed boundary method for threedimensional fluid-membrane interactions. Journal of Computational Physics, 228(22):8427–8445, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004574.

Lei:2004:AUD

[LWW04] Hong Lei, Lian-Ze Wang, and Zi-Niu Wu. Applications of upwind and downwind schemes for calculating electrical conditions in a wire-plate electrostatic precipitator. Journal of Computational Physics, 193(2):697–707, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910300456X.

Lian:2000:GKS

[LX00] Y. S. Lian and K. Xu. A gas-kinetic scheme for multimaterial flows and its application in chemical reactions. Journal of Computational Physics, 163(2):349–375, September 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965719.

Lin:2007:FDS

[LX07a] Yumin Lin and Chuanju Xu. Finite difference/spectral approximations for the time-fractional diffusion equation. Journal of Computational Physics, 225(2):1533–1552, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000678.

Liu:2007:RKD

[LX07b] Hongwei Liu and Kun Xu. A Runge–Kutta discontinuous Galerkin method for viscous flow equations. Journal of Computational Physics, 224(2):1223–1242, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005730.

Li:2009:GPC

[LX09] Jia Li and Dongbin Xiu. A generalized polynomial chaos based ensemble Kalman filter with high accuracy. Journal of Computational Physics, 228(15):5454–5469, August 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002137.

Lian:2009:SLT

[LXM09] Chenzhou Lian, Guoping Xia, and Charles L. Merkle. Solution-limited time stepping to enhance reliability in CFD applications. Journal of Computational Physics, 228(13): 4836–4857, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001697.

Lo:2004:ALE

[LY04] D. C. Lo and D. L. Young. Arbitrary Lagrangian– Eulerian finite element analysis of free surface flow using a velocity-vorticity formulation. Journal of Computational Physics, 195(1):175–201, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103005163.

Liu:2006:LDG

[LY06] Hailiang Liu and Jue Yan. A local discontinuous Galerkin method for the Korteweg–de Vries equation with boundary effect. Journal of Computational Physics, 215(1):197–218, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004857.

Lakoba:2007:GPI

[LY07a] T. I. Lakoba and J. Yang. A generalized Petviashvili iteration method for scalar and vector Hamiltonian equations with arbitrary form of nonlinearity. Journal of Computational Physics, 226(2):1668–1692, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002562.

Lakoba:2007:MET

[LY07b] T. I. Lakoba and J. Yang. A mode elimination technique to improve convergence of iteration methods for finding solitary waves. Journal of Computational Physics, 226(2):1693–1709, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002641.

Lin:2009:HRF

[LYC09] Chin-Tien Lin, Jia-Yi Yeh, and Jiun-Yeu Chen. High resolution finite volume scheme for the quantum hydrodynamic equations. Journal of Computational Physics, 228(5):1713– 1732, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005901.

Lynch:2008:OCW

[Lyn08] Peter Lynch. The origins of computer weather prediction and climate modeling. Journal of Computational Physics, 227(7):3431–3444, March 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000952.

Li:2004:SGK

[LZ04] Zhi-Hui Li and Han-Xin Zhang. Study on gas kinetic unified algorithm for flows from rarefied transition to continuum. Journal of Computational Physics, 193(2):708–738, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004571.

LeRoux:2007:CSE

[LZ07] S. Le Roux and G. Zérah. Convergence stability and estimator in orbital free electronic structure calculation on a grid at finite temperature. Journal of Computational Physics, 226(2):2063–2077, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002781.

Leung:2009:GBPb

[LZ09a] Shingyu Leung and Hongkai Zhao. A grid based particle method for evolution of open curves and surfaces. Journal of Computational Physics, 228(20):7706–7728, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900391X.

Leung:2009:GBPa

[LZ09b] Shingyu Leung and Hongkai Zhao. A grid based particle method for moving interface problems. Journal of Computational Physics, 228(8):2993–3024, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000138.

Li:2009:GKN

[LZ09c] Zhi-Hui Li and Han-Xin Zhang. Gas-kinetic numerical studies of three-dimensional complex flows on spacecraft re-entry. Journal of Computational Physics, 228(4):1116–1138, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005408.

Lu:2004:DGM

[LZC04] Tiao Lu, Pingwen Zhang, and Wei Cai. Discontinuous Galerkin methods for dispersive and lossy Maxwell’s equations and PML boundary conditions. Journal of Computational Physics, 200(2):549–580, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001792.

Lv:2006:EPU

[LZH^{+}06] X. Lv, Y. Zhao, X. Y. Huang, G. H. Xia, and Z. J. Wang. An efficient parallel/unstructured-multigrid preconditioned implicit method for simulating 3D unsteady compressible flows with moving objects. Journal of Computational Physics, 215(2):661–690, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005103.

Lv:2007:MFI

[LZH^{+}07] X. Lv, Y. Zhao, X. Y. Huang, G. H. Xia, and X. H. Su. A matrix-free implicit unstructured multigrid finite volume method for simulating structural dynamics and fluid-structure interaction. Journal of Computational Physics, 225(1):120– 144, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005808.

Lee:2003:ILP

[LZL03] Jeonghwa Lee, Jun Zhang, and Cai-Cheng Lu. Incomplete LU preconditioning for large scale dense complex linear systems from electromagnetic wave scattering problems. Journal of Computational Physics, 185(1):158–175, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000529.

Lee:2009:AMF

[LZT09] S. H. Lee, H. Zhou, and H. A. Tchelepi. Adaptive multiscale finite-volume method for nonlinear multiphase transport in heterogeneous formations. Journal of Computational Physics, 228(24):9036–9058, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900494X.

Ma:2005:MLP

[Ma05] Q. W. Ma. Meshless local Petrov–Galerkin method for twodimensional nonlinear water wave problems. Journal of Computational Physics, 205(2):611–625, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004826.

Macdonald:2000:CPN

[Mac00] J. Ross Macdonald. Comparison of parametric and nonparametric methods for the analysis and inversion of immittance data: Critique of earlier work. Journal of Computational Physics, 157(1):280–301, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963787.

Macrossan:2001:DFS

[Mac01] M. N. Macrossan. ν-DSMC: a fast simulation method for rarefied flow. Journal of Computational Physics, 173(2): 600–619, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968986.

Macrossan:2003:DGV

[Mac03] M. N. Macrossan. μ-DSMC: a general viscosity method for rarefied flow. Journal of Computational Physics, 185(2): 612–627, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103000093.

Machorro:2007:DGF

[Mac07] Eric Machorro. Discontinuous Galerkin finite element method applied to the 1-D spherical neutron transport equation. Journal of Computational Physics, 223(1):67–81, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004244.

Madasu:2005:CIF

[Mad05] Srinath Madasu. Capillary impregnation in a flexible slit: a finite element formulation. Journal of Computational Physics, 206(1):277–301, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104005133.

Madzvamuse:2006:TSS

[Mad06] Anotida Madzvamuse. Time-stepping schemes for moving grid finite elements applied to reaction-diffusion systems on fixed and growing domains. Journal of Computational Physics, 214(1):239–263, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500433X.

Mainland:2001:LST

[Mai01] G. B. Mainland. Logarithmic singularities in two-body, bound-state integral equations. Journal of Computational Physics, 174(2):852–869, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969414.

Mainland:2003:STB

[Mai03] G. B. Mainland. Solving the two-body, bound-state Bethe– Salpeter equation. Journal of Computational Physics, 192 (1):21–35, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003450.

Mainland:2004:RBC

[Mai04] G. B. Mainland. The role of boundary conditions in solving finite-energy, two-body, bound-state Bethe–Salpeter equations. Journal of Computational Physics, 197(2):610–623, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006545.

Maire:2009:HOCb

[Mai09a] Pierre-Henri Maire. A high-order cell-centered Lagrangian scheme for compressible fluid flows in two-dimensional cylindrical geometry. Journal of Computational Physics, 228(18):6882–6915, October 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003283.

Maire:2009:HOCa

[Mai09b] Pierre-Henri Maire. A high-order cell-centered Lagrangian scheme for two-dimensional compressible fluid flows on unstructured meshes. Journal of Computational Physics, 228 (7):2391–2425, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006359.

Mussa:2009:LBS

[MAL09] Alberto Mussa, Pietro Asinari, and Li-Shi Luo. Lattice Boltzmann simulations of 2D laminar flows past two tandem cylinders. Journal of Computational Physics, 228(4):983–999, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005342.

Mandel:2002:ISM

[Man02] Jan Mandel. An iterative substructuring method for coupled fluid-solid acoustic problems. Journal of Computational Physics, 177(1):95–116, March 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297004X.

Morel:2006:SDS

[MAN^{+}06] Jim E. Morel, B. Todd Adams, Taewan Noh, John M. McGhee, Thomas M. Evans, and Todd J. Urbatsch. Spatial discretizations for self-adjoint forms of the radiative transfer equations. Journal of Computational Physics, 214(1):12–40, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004213.

Martinsson:2006:FEE

[Mar06] Per-Gunnar Martinsson. Fast evaluation of electro-static interactions in multi-phase dielectric media. Journal of Computational Physics, 211(1):289–299, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002755.

Marshall:2009:DEM

[Mar09] J. S. Marshall. Discrete-element modeling of particulate aerosol flows. Journal of Computational Physics, 228(5): 1541–1561, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800572X.

Mavriplis:2002:ALV

[Mav02] Dimitri J. Mavriplis. An assessment of linear versus nonlinear multigrid methods for unstructured mesh solvers. Journal of Computational Physics, 175(1):302–325, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969487.

Mazumder:2006:CAS

[Maz06] Sandip Mazumder. Critical assessment of the stability and convergence of the equations of multi-component diffusion. Journal of Computational Physics, 212(2):383–392, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003554.

Martineau:2004:PCI

[MB04] Richard C. Martineau and Ray A. Berry. The pressurecorrected ICE finite element method for compressible flows on unstructured meshes. Journal of Computational Physics, 198(2):659–685, August 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000580.

Maruhn:2001:CFT

[MBM01] J. A. Maruhn, T. Bürvenich, and D. G. Madland. Calculating the Fierz transformation for higher orders. Journal of Computational Physics, 169(1):238–245, May 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967270.

Mazzia:2000:TST

[MBP00] Annamaria Mazzia, Luca Bergamaschi, and Mario Putti. A time-splitting technique for the advection-dispersion equation in groundwater. Journal of Computational Physics, 157 (1):181–198, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199963702.

Moureau:2007:ESI

[MBP07] V. Moureau, C. Bérat, and H. Pitsch. An efficient semiimplicit compressible solver for large-eddy simulations. Journal of Computational Physics, 226(2):1256–1270, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700232X.

McKinley:2003:CIS

[MBS03] Michael Scott McKinley, Eugene D. Brooks III, and Abraham Szoke. Comparison of implicit and symbolic implicit Monte Carlo line transport with frequency weight vector extension. Journal of Computational Physics, 189(1):330–349, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002134.

Martin:2000:CCA

[MC00a] Daniel F. Martin and Phillip Colella. A cell-centered adaptive projection method for the incompressible Euler equations. Journal of Computational Physics, 163(2):271–312, September 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965756.

Mohseni:2000:NTP

[MC00b] Kamran Mohseni and Tim Colonius. Numerical treatment of polar coordinate singularities. Journal of Computational Physics, 157(2):787–795, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963829.

Miller:2001:HOE

[MC01] G. H. Miller and P. Colella. A high-order Eulerian Godunov method for elastic-plastic flow in solids. Journal of Computational Physics, 167(1):131–176, February 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966658.

Miller:2002:CTD

[MC02] G. H. Miller and P. Colella. A conservative three-dimensional Eulerian method for coupled solid-fluid shock capturing. Journal of Computational Physics, 183(1):26–82, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971585.

Metcalfe:2003:SSM

[MC03] Travis S. Metcalfe and Paul Charbonneau. Stellar structure modeling using a parallel genetic algorithm for objective global optimization. Journal of Computational Physics, 185(1):176–193, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000530.

Ma:2004:PCR

[MC04] Yuanle Ma and Zhangxin Chen. Parallel computation for reservoir thermal simulation of multicomponent and multiphase fluid flow. Journal of Computational Physics, 201(1):224–237, November 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002104.

Martin:2006:PIM

[MC06a] M. Pino Martín and Graham V. Candler. A parallel implicit method for the direct numerical simulation of wall-bounded compressible turbulence. Journal of Computational Physics, 215(1):153–171, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004833.

Mehra:2006:HVI

[MC06b] Vishal Mehra and Shashank Chaturvedi. High velocity impact of metal sphere on thin metallic plates: a comparative smooth particle hydrodynamics study. Journal of Computational Physics, 212(1):318–337, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003232.

Marry:2007:TDA

[MC07a] V. Marry and G. Ciccotti. Trotter derived algorithms for molecular dynamics with constraints: Velocity Verlet revisited. Journal of Computational Physics, 222(1):428–440, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106003603.

Miniati:2007:BSA

[MC07b] Francesco Miniati and Phillip Colella. Block structured adaptive mesh and time refinement for hybrid, hyperbolic + N-body systems. Journal of Computational Physics, 227(1):400–430, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003385.

Miniati:2007:MHO

[MC07c] Francesco Miniati and Phillip Colella. A modified higher order Godunov’s scheme for stiff source conservative hydrodynamics. Journal of Computational Physics, 224(2):519–538, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004803.

Medale:2009:PCI

[MC09] Marc Medale and Bruno Cochelin. A parallel computer implementation of the asymptotic numerical method to study thermal convection instabilities. Journal of Computational Physics, 228(22):8249–8262, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004215.

Mangeney:2002:NSI

[MCCT02] A. Mangeney, F. Califano, C. Cavazzoni, and P. Travnicek. A numerical scheme for the integration of the Vlasov– Maxwell system of equations. Journal of Computational Physics, 179(2):495–538, July 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970713.

Martin:2008:CCA

[MCG08] Daniel F. Martin, Phillip Colella, and Daniel Graves. A cell-centered adaptive projection method for the incompressible Navier–Stokes equations in three dimensions. Journal of Computational Physics, 227(3):1863–1886, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004366.

McCorquodale:2004:NCL

[MCGV04] Peter McCorquodale, Phillip Colella, David P. Grote, and Jean-Luc Vay. A node-centered local refinement algorithm for Poisson’s equation in complex geometries. Journal of Computational Physics, 201(1):34–60, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001901.

McCorquodale:2001:CGE

[MCJ01] Peter McCorquodale, Phillip Colella, and Hans Johansen. A Cartesian grid embedded boundary method for the heat equation on irregular domains. Journal of Computational Physics, 173(2):620–635, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969001.

Mahesh:2004:NML

[MCM04] K. Mahesh, G. Constantinescu, and P. Moin. A numerical method for large-eddy simulation in complex geometries. Journal of Computational Physics, 197(1):215–240, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006272.

Minev:2003:FET

[MCN03] P. D. Minev, T. Chen, and K. Nandakumar. A finite element technique for multifluid incompressible flow using Eulerian grids. Journal of Computational Physics, 187(1):255–273, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000986.

Minier:2003:CAE

[MCP03] J.-P. Minier, R. Cao, and S. B. Pope. Comment on the article “An effective particle tracing scheme on structured/unstructured grids in hybrid finite volume/PDF Monte Carlo methods” by Li and Modest. Journal of Computational Physics, 186(1):356–358, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000068. See [LM01, LM03a].

Moukalled:2001:HRP

[MD01] F. Moukalled and M. Darwish. A high-resolution pressurebased algorithm for fluid flow at all speeds. Journal of Computational Physics, 168(1):101–130, March 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096683X.

Malarkey:2002:URC

[MD02] J. Malarkey and A. G. Davies. Use of Routh’s correction in the cloud-in-cell discrete vortex method. Journal of Computational Physics, 181(2):753–759, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971639.

Magin:2004:TAP

[MD04] Thierry E. Magin and Gérard Degrez. Transport algorithms for partially ionized and unmagnetized plasmas. Journal of Computational Physics, 198(2):424–449, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000361.

Mlacnik:2006:UGO

[MD06] Martin J. Mlacnik and Louis J. Durlofsky. Unstructured grid optimization for improved monotonicity of discrete solutions of elliptic equations with highly anisotropic coefficients. Journal of Computational Physics, 216(1):337–361, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005553.

Mittal:2008:VSI

[MDB^{+}08] R. Mittal, H. Dong, M. Bozkurttas, F. M. Najjar, A. Vargas, and A. von Loebbecke. A versatile sharp interface immersed boundary method for incompressible flows with complex boundaries. Journal of Computational Physics, 227(10):4825–4852, May 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000235.

Metzner:2007:GEM

[MDJS07] P. Metzner, E. Dittmer, T. Jahnke, and Ch. Schütte. Generator estimation of Markov jump processes. Journal of Computational Physics, 227(1):353–375, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003361.

Moresi:2003:LIP

[MDM03] L. Moresi, F. Dufour, and H.-B. Mühlhaus. A Lagrangian integration point finite element method for large deformation modeling of viscoelastic geomaterials. Journal of Computational Physics, 184(2):476–497, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000311.

Munz:2007:LAP

[MDR07] Claus-Dieter Munz, Michael Dumbser, and Sabine Roller. Linearized acoustic perturbation equations for low Mach number flow with variable density and temperature. Journal of Computational Physics, 224(1):352–364, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700085X.

Moukalled:2003:PBA

[MDS03] F. Moukalled, M. Darwish, and B. Sekar. A pressure-based algorithm for multi-phase flow at all speeds. Journal of Computational Physics, 190(2):550–571, September 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002973.

Mendez:2009:AMP

[ME09] S. Mendez and J. D. Eldredge. Acoustic modeling of perforated plates with bias flow for large-eddy simulations. Journal of Computational Physics, 228(13):4757–4772, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001624.

Mead:2004:SWE

[Mea04] J. L. Mead. The shallow water equations in Lagrangian coordinates. Journal of Computational Physics, 200(2):654– 669, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104001834.

Marinescu:2002:TSM

[MEG02] Dorin Marinescu, Aude Espesset, and Cecil P. Grünfeld. Tests of a simulation method for Boltzmann-like models with chemical reactions. Journal of Computational Physics, 175(1):225–248, January 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969402.

Mehta:2004:UWP

[Meh04] Prashant G. Mehta. A unified well-posed computational approach for the 2D Orr–Sommerfeld problem. Journal of Computational Physics, 199(2):541–557, September 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000944.

Mao:2003:DLD

[MEKS03] Deming Mao, Jack R. Edwards, Andrey V. Kuznetsov, and Ravi K. Srivastava. Development of low-diffusion flux-splitting methods for dense gas-solid flows. Journal of Computational Physics, 185(1):100–119, February 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000499.

McClarren:2008:SIT

[MELD08] Ryan G. McClarren, Thomas M. Evans, Robert B. Lowrie, and Jeffery D. Densmore. Semi-implicit time integration for P_{N} thermal radiative transfer. Journal of Computational Physics, 227(16):7561–7586, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002489.

Mencinger:2004:NSM

[Men04] Jure Mencinger. Numerical simulation of melting in twodimensional cavity using adaptive grid. Journal of Computational Physics, 198(1):243–264, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000269.

Michoski:2009:QHT

[MESV09] C. Michoski, J. A. Evans, P. G. Schmitz, and A. Vasseur. Quantum hydrodynamics with trajectories: The nonlinear conservation form mixed/discontinuous Galerkin method with applications in chemistry. Journal of Computational Physics, 228(23):8589–8608, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004483.

Meitz:2000:CDS

[MF00] Hubert L. Meitz and Hermann F. Fasel. A compact-difference scheme for the Navier–Stokes equations in vorticity-velocity formulation. Journal of Computational Physics, 157(1):371– 403, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199963878.

Matsuo:2001:DCF

[MF01] Takayasu Matsuo and Daisuke Furihata. Dissipative or conservative finite-difference schemes for complex-valued nonlinear partial differential equations. Journal of Computational Physics, 171(2):425–447, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967750.

Moinier:2002:SAP

[MG02] P. Moinier and M. B. Giles. Stability analysis of preconditioned approximations of the Euler equations on unstructured meshes. Journal of Computational Physics, 178(2):498–519, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970385.

Marzouk:2005:MCO

[MG05a] Youssef M. Marzouk and Ahmed F. Ghoniem. K-means clustering for optimal partitioning and dynamic load balancing of parallel hierarchical N-body simulations. Journal of Computational Physics, 207(2):493–528, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000380.

Murrone:2005:FER

[MG05b] Angelo Murrone and Hervé Guillard. A five equation reduced model for compressible two phase flow problems. Journal of Computational Physics, 202(2):664–698, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104003018.

Min:2006:SOA

[MG06] Chohong Min and Frédéric Gibou. A second order accurate projection method for the incompressible Navier–Stokes equations on non-graded adaptive grids. Journal of Computational Physics, 219(2):912–929, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003366.

Mayo:2007:FOA

[MG07a] Anita Mayo and Anne Greenbaum. Fourth order accurate evaluation of integrals in potential theory on exterior 3D regions. Journal of Computational Physics, 220(2):900–914, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002671.

Miller:2007:FAR

[MG07b] David J. Miller and Avijit Ghosh. A fully adaptive reactiondiffusion integration scheme with applications to systems biology. Journal of Computational Physics, 226(2):1509–1531, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002471.

Min:2007:GII

[MG07c] Chohong Min and Frédéric Gibou. Geometric integration over irregular domains with application to level-set methods. Journal of Computational Physics, 226(2):1432–1443, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002410.

Min:2007:SOA

[MG07d] Chohong Min and Frédéric Gibou. A second order accurate level set method on non-graded adaptive Cartesian grids. Journal of Computational Physics, 225(1):300–321, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005912.

Min:2008:RSO

[MG08] Chohong Min and Frédéric Gibou. Robust second-order accurate discretizations of the multi-dimensional Heaviside and Dirac delta functions. Journal of Computational Physics, 227(22):9686–9695, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003999.

Min:2006:SCF

[MGC06] Chohong Min, Frédéric Gibou, and Hector D. Ceniceros. A supra-convergent finite difference scheme for the variable coefficient Poisson equation on non-graded grids. Journal of Computational Physics, 218(1):123–140, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000635.

Marchandise:2007:SFE

[MGCR07] Emilie Marchandise, Philippe Geuzaine, Nicolas Chevaugeon, and Jean-François Remacle. A stabilized finite element method using a discontinuous level set approach for the computation of bubble dynamics. Journal of Computational Physics, 225(1):949–974, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000125.

Marshall:2000:VTL

[MGGH00] J. S. Marshall, J. R. Grant, A. A. Gossler, and S. A. Huyer. Vorticity transport on a Lagrangian tetrahedral mesh. Journal of Computational Physics, 161(1):85–113, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964908.

Murillo:2009:CNS

[MGNB09] J. Murillo, P. García-Navarro, and J. Burguete. Conservative numerical simulation of multi-component transport in twodimensional unsteady shallow water flow. Journal of Computational Physics, 228(15):5539–5573, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002290.

Meyers:2007:CEA

[MGS07] J. Meyers, B. J. Geurts, and P. Sagaut. A computational error-assessment of central finite-volume discretizations in large-eddy simulation using a Smagorinsky model. Journal of Computational Physics, 227(1):156–173, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003257.

McMahon:2009:DAP

[MGS09] Jeffrey M. McMahon, Stephen K. Gray, and George C. Schatz. A discrete action principle for electrodynamics and the construction of explicit symplectic integrators for linear, non-dispersive media. Journal of Computational Physics, 228(9):3421–3432, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000448.

McClarren:2008:SET

[MHB08] Ryan G. McClarren, James Paul Holloway, and Thomas A. Brunner. On solutions to the P_{n} equations for thermal radiative transfer. Journal of Computational Physics, 227(5):2864–2885, February 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005153.

Munts:2007:MBM

[MHdB07] E. A. Munts, S. J. Hulshoff, and R. de Borst. A modalbased multiscale method for large eddy simulation. Journal of Computational Physics, 224(1):389–402, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001015.

Montijn:2006:AGR

[MHE06] C. Montijn, W. Hundsdorfer, and U. Ebert. An adaptive grid refinement strategy for the simulation of negative streamers. Journal of Computational Physics, 219(2):801–835, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002178.

Mattsson:2008:SAW

[MHI08] Ken Mattsson, Frank Ham, and Gianluca Iaccarino. Stable and accurate wave-propagation in discontinuous media.

Journal of Computational Physics, 227(19):8753–8767, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108003434.

Monkola:2008:THE

[MHPR08] Sanna Mönkölä, Erkki Heikkola, Anssi Pennanen, and Tuomo Rossi. Time-harmonic elasticity with controllability and higher-order discretization methods. Journal of Computational Physics, 227(11):5513–5534, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000867.

Morel:2001:LSO

[MHS01] J. E. Morel, Michael L. Hall, and Mikhail J. Shashkov. A local support-operators diffusion discretization scheme for hexahedral meshes. Journal of Computational Physics, 170(1): 338–372, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101967361.

Merks:2002:MPM

[MHS02] R. M. H. Merks, A. G. Hoekstra, and P. M. A. Sloot. The moment propagation method for advection-diffusion in the lattice Boltzmann method: Validation and Péclet number limits. Journal of Computational Physics, 183(2):563–576, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102972098.

Monaghan:2005:SUS

[MHW05] Joseph J. Monaghan, Herbert E. Huppert, and M. Grae Worster. Solidification using smoothed particle hydrodynamics. Journal of Computational Physics, 206(2):684–705, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000112.

Mieussens:2000:DVM

[Mie00] Luc Mieussens. Discrete-velocity models and numerical schemes for the Boltzmann–BGK equation in plane and axisymmetric geometries. Journal of Computational Physics, 162(2):429–466, August 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965483.

Mignone:2007:SAR

[Mig07] A. Mignone. A simple and accurate Riemann solver for isothermal MHD. Journal of Computational Physics, 225 (2):1427–1441, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107000617.

Miller:2004:IRS

[Mil04] G. H. Miller. An iterative Riemann solver for systems of hyperbolic conservation laws, with application to hyperelastic solid mechanics. Journal of Computational Physics, 193(1):198–225, January 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004273.

Milotti:2005:SFP

[Mil05] Edoardo Milotti. Sine-fit procedure for unevenly sampled, multiply clocked signals. Journal of Computational Physics, 202(1):134–149, January 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002700.

Milotti:2006:MBF

[Mil06] Edoardo Milotti. Model-based fit procedure for power-law-like spectra. Journal of Computational Physics, 217(2): 834–844, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910600043X.

See erratum [Mil07].

Milotti:2007:EMB

[Mil07] E. Milotti. Erratum to “Model-based fit procedure for powerlaw-like spectra” [J. Comput. Phys. 217 (2006) 834–844]. Journal of Computational Physics, 227(2):1608, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003798. See [Mil06].

Miller:2008:IBP

[Mil08] G. H. Miller. An iterative boundary potential method for the infinite domain Poisson problem with interior Dirichlet boundaries. Journal of Computational Physics, 227(16):7917–7928, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108002684.

Min:2003:SIA

[Min03] Chohong Min. Simplicial isosurfacing in arbitrary dimension and codimension. Journal of Computational Physics, 190(1):295–310, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002754.

Min:2004:LLS

[Min04]  Chohong Min.  Local level set method in high di-

mension and codimension. Journal of Computational Physics, 200(1):368–382, October 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104001718.

Miniati:2007:GGM

[Min07] Francesco Miniati. Glimm–Godunov’s method for cosmicray-hydrodynamics. Journal of Computational Physics, 227(1):776–796, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003592.

Mitin:2000:NMC

[Mit00] Alexander V. Mitin. New methods for calculations of the lowest eigenvalues of the real symmetric generalized eigenvalue problem. Journal of Computational Physics, 161(2):653–667, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965203.

McMullen:2006:CEN

[MJ06] M. S. McMullen and A. Jameson. The computational efficiency of non-linear frequency domain methods. Journal of Computational Physics, 212(2):637–661, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003499.

Meyer:2007:CIO

[MJ07] D. W. Meyer and P. Jenny. Consistent inflow and outflow boundary conditions for transported probability density function methods. Journal of Computational Physics, 226(2):1859–1873, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002665.

Ma:2009:MMM

[MJ09a] Jingtang Ma and Yingjun Jiang. Moving mesh methods for blowup in reaction-diffusion equations with traveling heat source. Journal of Computational Physics, 228(18):6977–6990, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003337.

Meyer:2009:MMT

[MJ09b] Daniel W. Meyer and Patrick Jenny. Micromixing models for turbulent flows. Journal of Computational Physics, 228 (4):1275–1293, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005469.

Matringe:2006:RST

[MJT06] Sébastien F. Matringe, Ruben Juanes, and Hamdi A. Tchelepi. Robust streamline tracing for the simulation of porous media flow on general triangular and quadrilateral grids. Journal of Computational Physics, 219(2):992–1012, December 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003408.

Milano:2002:CGA

[MK02a] Michele Milano and Petros Koumoutsakos. A clustering genetic algorithm for cylinder drag optimization. Journal of Computational Physics, 175(1):79–107, January 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968822.

Milano:2002:NNM

[MK02b] Michele Milano and Petros Koumoutsakos. Neural network modeling for near Wall turbulent flow. Journal of Computational Physics, 182(1):1–26, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971469.

Mousseau:2003:NPB

[MK03] V. A. Mousseau and D. A. Knoll. New physics-based preconditioning of implicit methods for non-equilibrium radiation diffusion. Journal of Computational Physics, 190(1): 42–51, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103002523.

Mocken:2004:QDR

[MK04a] Guido R. Mocken and Christoph H. Keitel. Quantum dynamics of relativistic electrons. Journal of Computational Physics, 199(2):558–588, September 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000956.

Muller:2004:AEC

[MK04b] Jens Müller and Jan G. Korvink. Adaptive error control in multi-physical thin-structure MEMS FE-simulation. Journal of Computational Physics, 196(1):145–172, May 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005916.

Miyoshi:2005:MSH

[MK05] Takahiro Miyoshi and Kanya Kusano. A multi-state HLL approximate Riemann solver for ideal magnetohydrodynamics.

Journal of Computational Physics, 208(1):315–344, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001142.

Muradoglu:2006:AGM

[MK06] Metin Muradoğlu and Arif Doruk Kayaalp. An auxiliary grid method for computations of multiphase flows in complex geometries. Journal of Computational Physics, 214(2):858–877, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004766.

Munsky:2007:MTI

[MK07] Brian Munsky and Mustafa Khammash. A multiple time interval finite state projection algorithm for the solution to the chemical master equation. Journal of Computational Physics, 226(1):818–835, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002100.

Ma:2008:VRR

[MK08a] Lin Ma and William S. Klug. Viscous regularization and r-adaptive remeshing for finite element analysis of lipid membrane mechanics. Journal of Computational Physics, 227(11):5816–5835, May 10, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001150.

Mehra:2008:AWC

[MK08b] Mani Mehra and Nicholas K.-R. Kevlahan. An adaptive wavelet collocation method for the solution of partial differential equations on the sphere. Journal of Computational Physics, 227(11):5610–5632, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001010.

McDonough:2006:ADS

[MKKY06] J. M. McDonough, I. Kunadian, R. R. Kumar, and T. Yang. An alternative discretization and solution procedure for the dual phase-lag equation. Journal of Computational Physics, 219(1):163–171, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001458.

Medyanik:2006:DRM

[MKL06] Sergey N. Medyanik, Eduard G. Karpov, and Wing Kam Liu. Domain reduction method for atomistic simulations. Journal of Computational Physics, 218(2):836–859, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001355.

Marella:2005:SIC

[MKLU05] S. Marella, S. Krishnan, H. Liu, and H. S. Udaykumar. Sharp interface Cartesian grid method I: An easily implemented technique for 3D moving boundary computations. Journal of Computational Physics, 210(1):1–31, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001956.

Mansfield:1999:DCV

[MKM99] John R. Mansfield, Omar M. Knio, and Charles Meneveau. Dynamic LES of colliding vortex rings using a 3D vortex method. Journal of Computational Physics, 152(1):305–345, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199962587. See erra-

tum [MKM04].

Mansfield:2004:CDC

[MKM04] John R. Mansfield, Omar M. Knio, and Charles Meneveau. Corrigendum to “Dynamic LES of colliding vortex rings using a 3D vortex method” [J. Comp. Phys. 152 (1999) 305– 345]. Journal of Computational Physics, 197(2):779–780, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006569. See [MKM99].

Maple:2004:AHB

[MKOW04] Raymond C. Maple, Paul I. King, Paul D. Orkwis, and J. Mitch Wolff. Adaptive harmonic balance method for nonlinear time-periodic flows. Journal of Computational Physics, 193(2):620–641, January 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004376.

Mousseau:2000:PBP

[MKR00] V. A. Mousseau, D. A. Knoll, and W. J. Rider. Physicsbased preconditioning and the Newton–Krylov method for non-equilibrium radiation diffusion. Journal of Computational Physics, 160(2):743–765, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096488X.

Mazumder:2001:TRS

[ML01a] Sandip Mazumder and Samuel A. Lowry. The treatment of reacting surfaces for finite-volume schemes on unstructured meshes. Journal of Computational Physics, 173(2): 512–526, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968901.

Moeleker:2001:LMT

[ML01b] Piet Moeleker and Anthony Leonard. Lagrangian methods for the tensor-diffusivity subgrid model. Journal of Computational Physics, 167(1):1–21, February 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966695.

Miller:2004:EDM

[ML04] S. Miller and S. Luding. Event-driven molecular dynamics in parallel. Journal of Computational Physics, 193(1):306–316, January 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004339.

Macklin:2005:EIG

[ML05] Paul Macklin and John Lowengrub. Evolving interfaces via gradients of geometry-dependent interior Poisson problems: application to tumor growth. Journal of Computational Physics, 203(1):191–220, February 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003249.

Macklin:2006:IGA

[ML06a] Paul Macklin and John Lowengrub. An improved geometryaware curvature discretization for level set methods: Application to tumor growth. Journal of Computational Physics, 215(2):392–401, July 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005152.

Moryossef:2006:UPI

[ML06b] Y. Moryossef and Y. Levy. Unconditionally positive implicit procedure for two-equation turbulence models: Application to k–ω turbulence models. Journal of Computational Physics, 220(1):88–108, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002233.

McClarren:2008:ESL

[ML08] Ryan G. McClarren and Robert B. Lowrie. The effects of slope limiting on asymptotic-preserving numerical methods for hyperbolic conservation laws. Journal of Computational Physics, 227(23):9711–9726, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004002.

Min:2006:FSS

[MLFG06] Mi-Sun Min, Tae-Woo Lee, Paul F. Fischer, and Stephen K. Gray. Fourier spectral simulations and Gegenbauer reconstructions for electromagnetic waves in the presence of a metal nanoparticle. Journal of Computational Physics, 213 (2):730–747, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105004146.

Mani:2009:SAB

[MLM09] Ali Mani, Johan Larsson, and Parviz Moin. Suitability of artificial bulk viscosity for large-eddy simulation of turbulent flows with shocks. Journal of Computational Physics, 228(19):7368–7374, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003623.

Montero:2001:RMA

[MLS01] Ruben S. Montero, Ignacio M. Llorente, and Manuel D. Salas. Robust multigrid algorithms for the Navier–Stokes equations. Journal of Computational Physics, 173(2):412–432, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968639.

Moureau:2005:NMU

[MLS^{+}05] V. Moureau, G. Lartigue, Y. Sommerer, C. Angelberger, O. Colin, and T. Poinsot. Numerical methods for unsteady compressible multi-component reacting flows on fixed and moving grids. Journal of Computational Physics, 202 (2):710–736, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003134.

Mansur:2007:ETD

[MLSD07] W. J. Mansur, F. S. Loureiro, D. Soares, Jr., and C. Dors. Explicit time-domain approaches based on numerical Green’s functions computed by finite differences — the ExGA family. Journal of Computational Physics, 227(1):851–870, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003877.

Marquez:2001:NIT

[MM01] Antonio Márquez and Salim Meddahi. New implementation techniques for the exterior Stokes problem in the plane. Journal of Computational Physics, 172(2):685–703, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968482.

Marquina:2003:FSA

[MM03]  Antonio Marquina and Pep Mulet.  A flux-split al-

gorithm applied to conservative models for multicomponent compressible flows. Journal of Computational Physics, 185(1):120–138, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102000505.

Madzvamuse:2007:VIN

[MM07] Anotida Madzvamuse and Philip K. Maini. Velocityinduced numerical solutions of reaction-diffusion systems on continuously growing domains. Journal of Computational Physics, 225(1):100–119, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005791.

Mahalov:2009:VNN

[MM09] Alex Mahalov and Mohamed Moustaoui. Vertically nested nonhydrostatic model for multiscale resolution of flows in the upper troposphere and lower stratosphere. Journal of Computational Physics, 228(4):1294–1311, March 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800555X.

Martinez:2008:SPK

[MMKP08] E. Martínez, J. Marian, M. H. Kalos, and J. M. Perlado. Synchronous parallel kinetic Monte Carlo for continuum diffusion-reaction systems. Journal of Computational Physics, 227(8):3804–3823, April 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700527X.

Moureau:2007:GFM

[MMPB07] V. Moureau, P. Minot, H. Pitsch, and C. Bérat. A ghost-fluid method for large-eddy simulations of premixed combustion in complex geometries. Journal of Computational Physics, 221(2):600–614, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003068.

Marquez:2004:NBF

[MMS04] A. Márquez, S. Meddahi, and V. Selgas. A new BEM– FEM coupling strategy for two-dimensional fluid-solid interaction problems. Journal of Computational Physics, 199 (1):205–220, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104000828.

Menshov:2002:IVR

[MN02] Igor Men’shov and Yoshiaki Nakamura. Implementation of the variational Riemann problem solution for calculating propagation of sound waves in nonuniform flow fields. Journal of Computational Physics, 182(1):118–148, October 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971548.

Mattsson:2004:SPO

[MN04] Ken Mattsson and Jan Nordström. Summation by parts operators for finite difference approximations of second derivatives. Journal of Computational Physics, 199(2):503–540, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104000932. See corri-

gendum [MN17].

Mattsson:2006:HOF

[MN06] Ken Mattsson and Jan Nordström. High order finite difference methods for wave propagation in discontinuous media. Journal of Computational Physics, 220(1):249–269, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002300.

Maire:2009:MSG

[MN09a] Pierre-Henri Maire and Boniface Nkonga. Multi-scale Godunov-type method for cell-centered discrete Lagrangian hydrodynamics. Journal of Computational Physics, 228(3): 799–821, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800524X.

Marzouk:2009:DRP

[MN09b] Youssef M. Marzouk and Habib N. Najm. Dimensionality reduction and polynomial chaos acceleration of Bayesian inference in inverse problems. Journal of Computational Physics, 228(6):1862–1902, April 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006062.

Mattsson:2017:CSP

[MN17] Ken Mattsson and Jan Nordström. Corrigendum to “Summation by parts operators for finite difference approximations of second derivatives” [J. Comput. Phys. 199 (2004) 503–540]. Journal of Computational Physics, 351(??):535, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307118. See [MN04].

Marzouk:2007:SSM

[MNR07] Youssef M. Marzouk, Habib N. Najm, and Larry A. Rahn. Stochastic spectral methods for efficient Bayesian solution of inverse problems. Journal of Computational Physics, 224(2):560–586, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004839.

Muratov:2006:OGB

[MO06] C. B. Muratov and V. V. Osipov. Optimal grid-based methods for thin film micromagnetics simulations. Journal of Computational Physics, 216(2):637–653, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000040.

Michalak:2009:APL

[MOG09] Christopher Michalak and Carl Ollivier-Gooch. Accuracy preserving limiter for the high-order accurate solution of the Euler equations. Journal of Computational Physics, 228(23):8693–8711, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004641.

Monaghan:2000:STI

[Mon00] J. J. Monaghan. SPH without a tensile instability. Journal of Computational Physics, 159(2):290–311, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964398.

Moore:2003:IAT

[Moo03] Peter K. Moore. An incomplete assembly with thresholding algorithm for systems of reaction-diffusion equations in three space dimensions IAT for reaction-diffusion systems. Journal of Computational Physics, 189(1):130–158, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002055.

Moore:2007:SRS

[Moo07] Peter K. Moore. Solving regularly and singularly perturbed reaction-diffusion equations in three space dimensions. Journal of Computational Physics, 224(2):601–615, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005390.

Munz:2000:DCT

[MOS^{+}00] C.-D. Munz, P. Omnes, R. Schneider, E. Sonnendrücker, and U. Voß. Divergence correction techniques for Maxwell solvers based on a hyperbolic model. Journal of Computational Physics, 161(2):484–511, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965070.

Mottez:2008:GCD

[Mot08] F. Mottez. A guiding centre direct implicit scheme for magnetized plasma simulations. Journal of Computational Physics, 227(6):3260–3281, March 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005311.

Mousseau:2004:IBS

[Mou04] V. A. Mousseau. Implicitly balanced solution of the twophase flow equations coupled to nonlinear heat conduction. Journal of Computational Physics, 200(1):104–132, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001342.

Mott:2000:QSS

[MOvL00] David R. Mott, Elaine S. Oran, and Bram van Leer. A quasi-steady-state solver for the stiff ordinary differential equations of reaction kinetics. Journal of Computational Physics, 164(2):407–428, November 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966051.

Majorana:2001:FDS

[MP01a] A. Majorana and R. M. Pidatella. A finite difference scheme solving the Boltzmann–Poisson system for semiconductor devices. Journal of Computational Physics, 174(2):649–668, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969293.

Monk:2001:PAC

[MP01b] Peter Monk and Kevin Parrott. Phase-accuracy comparisons and improved far-field estimates for 3-D edge elements on tetrahedral meshes. Journal of Computational Physics, 170(2):614–641, July 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967531.

Majorana:2002:FDS

[MP02] A. Majorana and R. M. Pidatella. A finite difference scheme solving the Boltzmann–Poisson system for semiconductor devices: Volume 174, number 2 (2001), pages 649–668. Journal of Computational Physics, 177(2):450, April 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970312.

Mehdizadeh:2003:ITD

[MP03] Omid Z. Mehdizadeh and Marius Paraschivoiu. Investigation of a two-dimensional spectral element method for Helmholtz’s equation. Journal of Computational Physics, 189(1):111–129, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002043.

Mazzia:2005:HOG

[MP05] Annamaria Mazzia and Mario Putti. High order Godunov mixed methods on tetrahedral meshes for density driven flow simulations in porous media. Journal of Computational Physics, 208(1):154–174, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000902.

Manzini:2007:MLE

[MP07a] Gianmarco Manzini and Mario Putti. Mesh locking effects in the finite volume solution of 2-D anisotropic diffusion equations. Journal of Computational Physics, 220(2):751–771, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002580.

McDermott:2007:PFT

[MP07b] R. McDermott and S. B. Pope. A particle formulation for treating differential diffusion in filtered density function methods. Journal of Computational Physics, 226(1):947–993, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002173.

McDermott:2008:PER

[MP08] R. McDermott and S. B. Pope. The parabolic edge reconstruction method (PERM) for Lagrangian particle advection. Journal of Computational Physics, 227(11):5447–5491, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000843.

Muradoglu:2001:HMP

[MPC01] Metin Muradoğlu, Stephen B. Pope, and David A. Caughey. The hybrid method for the PDF equations of turbulent reactive flows: Consistency conditions and correction algorithms. Journal of Computational Physics, 172(2):841–878, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968615. See corri-

gendum [MPC02].

Muradoglu:2002:HMP

[MPC02] Metin Muradoğlu, Stephen B. Pope, and David A. Caughey. The hybrid method for the PDF equations of turbulent reactive flows: Consistency conditions and correction algorithms: Volume 172, number 2 (2001), pages 841–878. Journal of Computational Physics, 178(1):260, May 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970427. See [MPC01].

Macaskill:2003:CAQ

[MPD03] C. Macaskill, W. E. P. Padden, and D. G. Dritschel. The CASL algorithm for quasi-geostrophic flow in a cylinder. Journal of Computational Physics, 188(1):232–251, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103001657.

Matheou:2008:VFD

[MPD08] G. Matheou, C. Pantano, and P. E. Dimotakis. Verification of a fluid-dynamics solver using correlations with linear stability results. Journal of Computational Physics, 227 (11):5385–5396, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108000703.

Montseny:2008:DTL

[MPFC08] E. Montseny, S. Pernet, X. Ferriéres, and G. Cohen. Dissipative terms and local time-stepping improvements in a spatial high order discontinuous Galerkin scheme for the time-domain Maxwell’s equations. Journal of Computational Physics, 227(14):6795–6820, July 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001836.

Machiels:2001:PFE

[MPP01] L. Machiels, J. Peraire, and A. T. Patera. A posteriori finiteelement output bounds for the incompressible Navier–Stokes equations: Application to a natural convection problem. Journal of Computational Physics, 172(2):401–425, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967695.

Mackenzie:2000:NSO

[MR00] J. A. Mackenzie and M. L. Robertson. The numerical solution of one-dimensional phase change problems using an adaptive moving mesh method. Journal of Computational Physics, 161(2):537–557, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965112.

Murray:2001:CVB

[MR01] Ross J. Murray and C. J. C. Reason. A curvilinear version of the Bryan–Cox–Semtner ocean model and its representation of the Arctic circulation. Journal of Computational Physics, 171(1):1–46, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967610.

Mackenzie:2002:MMM

[MR02] J. A. Mackenzie and M. L. Robertson. A moving mesh method for the solution of the one-dimensional phase-field equations. Journal of Computational Physics, 181(2):526–544, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971408.

Millett:2003:ERO

[MR03] Kenneth C. Millett and Eric J. Rawdon. Energy, ropelength, and other physical aspects of equilateral knots. Journal of Computational Physics, 186(2):426–456, April 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000263.

Michaelis:2004:FSL

[MR04] B. Michaelis and B. Rogg. FEM-simulation of laminar flame propagation. I: Two-dimensional flames. Journal of Computational Physics, 196(2):417–447, May 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300603X.

Martinsson:2005:FDS

[MR05] P. G. Martinsson and V. Rokhlin. A fast direct solver for boundary integral equations in two dimensions. Journal of Computational Physics, 205(1):1–23, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004462.

Marchandise:2006:SFE

[MR06a] Emilie Marchandise and Jean-François Remacle. A stabilized finite element method using a discontinuous level set approach for solving two phase incompressible flows. Journal of Computational Physics, 219(2):780–800, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002142.

Motamed:2006:FPS

[MR06b] Mohammad Motamed and Olof Runborg. A fast phase space method for computing creeping rays. Journal of Computational Physics, 219(1):276–295, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001653.

Martinsson:2007:FDS

[MR07a] P. G. Martinsson and V. Rokhlin. A fast direct solver for scattering problems involving elongated structures. Journal of Computational Physics, 221(1):288–302, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002907.

Merriman:2007:DGM

[MR07b] Barry Merriman and Steven J. Ruuth. Diffusion generated motion of curves on surfaces. Journal of Computational Physics, 225(2):2267–2282, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001301.

Mishra:2007:STC

[MR07c] Subhash C. Mishra and Hillol K. Roy. Solving transient conduction and radiation heat transfer problems using the lattice Boltzmann method and the finite volume method. Journal of Computational Physics, 223(1):89–107, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106004268.

Marchandise:2006:QFD

[MRC06] Emilie Marchandise, Jean-François Remacle, and Nicolas Chevaugeon. A quadrature-free discontinuous Galerkin method for the level set equation. Journal of Computational Physics, 212(1):338–357, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003244.

Masserey:2005:NIT

[MRRS05] Alexandre Masserey, Jacques Rappaz, Roland Rozsnyo, and Marek Swierkosz. Numerical integration of the threedimensional Green kernel for an electromagnetic problem. Journal of Computational Physics, 205(1):48–71, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004486.

Marie:2009:CBL

[MRS09] Simon Marié, Denis Ricot, and Pierre Sagaut. Comparison between lattice Boltzmann method and Navier–Stokes high order schemes for computational aeroacoustics. Journal of Computational Physics, 228(4):1056–1070, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800538X.

Memoli:2001:FCW

[MS01] Facundo Mémoli and Guillermo Sapiro. Fast computation of weighted distance functions and geodesics on implicit hyper-surfaces. Journal of Computational Physics, 173(2): 730–764, November 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101969104.

Margolin:2003:SOS

[MS03] L. G. Margolin and Mikhail Shashkov. Second-order signpreserving conservative interpolation (remapping) on general grids. Journal of Computational Physics, 184(1):266–298, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000335.

Mascagni:2004:RWB

[MS04] Michael Mascagni and Nikolai A. Simonov. The random walk on the boundary method for calculating capacitance. Journal of Computational Physics, 195(2):465–473, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005382.

Maitre:2008:LSM

[MS08a] Emmanuel Maitre and Fadil Santosa. Level set methods for optimization problems involving geometry and constraints II. Optimization over a fixed surface. Journal of Computational Physics, 227(22):9596–9611, November 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003744.

Miller:2008:PWC

[MS08b] Martin J. Miller and Piotr K. Smolarkiewicz. Predicting weather, climate and extreme events. Journal of Computational Physics, 227(7):3429–3430, March 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000156.

Moore:2007:SMF

[MSB07a] Peter Moore, Harmen Slot, and Bendiks Jan Boersma. Simulation and measurement of flow generated noise. Journal of Computational Physics, 224(1):449–463, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001507.

Morrell:2007:CCA

[MSB07b] J. M. Morrell, P. K. Sweby, and A. Barlow. A cell by cell anisotropic adaptive mesh ALE scheme for the numerical solution of the Euler equations. Journal of Computational Physics, 226(1):1152–1180, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002264.

May:2007:IGK

[MSJ07] Georg May, Balaji Srinivasan, and Antony Jameson. An improved gas-kinetic BGK finite-volume method for threedimensional transonic flow. Journal of Computational Physics, 220(2):856–878, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002646.

Memoli:2004:SVP

[MSO04] Facundo Mémoli, Guillermo Sapiro, and Stanley Osher. Solving variational problems and partial differential equations mapping into general target manifolds. Journal of Computational Physics, 195(1):263–292, March 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005321.

Miloshevsky:2006:AFD

[MSP^{+}06] Gennady V. Miloshevsky, Valeryi A. Sizyuk, Michael B. Partenskii, Ahmed Hassanein, and Peter C. Jordan. Application of finite-difference methods to membrane-mediated protein interactions and to heat and magnetic field diffusion in plasmas. Journal of Computational Physics, 212 (1):25–51, February 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105003086.

Mattsson:2008:SAS

[MSS08] K. Mattsson, M. Svärd, and M. Shoeybi. Stable and accurate schemes for the compressible Navier–Stokes equations. Journal of Computational Physics, 227(4):2293–2316, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004627.

Meerschaert:2006:FDM

[MST06] Mark M. Meerschaert, Hans-Peter Scheffler, and Charles Tadjeran. Finite difference methods for two-dimensional fractional dispersion equation. Journal of Computational Physics, 211(1):249–261, January 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002731.

Mei:2000:LBM

[MSYL00] Renwei Mei, Wei Shyy, Dazhi Yu, and Li-Shi Luo. Lattice Boltzmann method for 3-D flows with curved boundary. Journal of Computational Physics, 161(2):680–699, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965227.

Min:2001:IYS

[MT01] M. S. Min and C. H. Teng. The instability of the Yee scheme for the “Magic Time Step”. Journal of Computational Physics, 166(2):418–424, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966506.

Meseguer:2003:LPF

[MT03] Á. Meseguer and L. N. Trefethen. Linearized pipe flow to Reynolds number 10^{7}. Journal of Computational Physics, 186(1):178–197, March 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000299.

Milstein:2004:ECW

[MT04] G. N. Milstein and M. V. Tretyakov. Evaluation of conditional Wiener integrals by numerical integration of stochastic differential equations. Journal of Computational Physics, 197(1):275–298, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103006302.

Marchuk:2007:MSP

[MT07a] O. Marchuk and M. Z. Tokar. Modeling of supersonic plasma flow in the scrape-off layer. Journal of Computational Physics, 227(2):1597–1607, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004299.

Moroney:2007:TDF

[MT07b] T. J. Moroney and I. W. Turner. A three-dimensional finite volume method based on radial basis functions for the accurate computational modelling of nonlinear diffusion equations. Journal of Computational Physics, 225(2):1409–1426, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000484.

Muradoglu:2008:FTM

[MT08]  Metin Muradoğlu and Gretar Tryggvason.  A front-

tracking method for computation of interfacial flows with soluble surfactants. Journal of Computational Physics, 227(4):2238–2262, February 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700438X.

Medvinsky:2008:LAB

[MTH08] M. Medvinsky, E. Turkel, and U. Hetmaniuk. Local absorbing boundary conditions for elliptical shaped boundaries. Journal of Computational Physics, 227(18):8254–8267, September 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108002866.

MacLachlan:2008:FRS

[MTV08] S. P. MacLachlan, J. M. Tang, and C. Vuik. Fast and robust solvers for pressure-correction in bubbly flow problems. Journal of Computational Physics, 227(23):9742–9761, December 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800404X.

Martin:2006:BOW

[MTWW06] M. P. Martín, E. M. Taylor, M. Wu, and V. G. Weirs. A bandwidth-optimized WENO scheme for the effective direct numerical simulation of compressible turbulence. Journal of Computational Physics, 220(1):270–289, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002312.

Mu:2002:MHA

[Mu02] Mo Mu. A multiple-heaps algorithm for parallel simulation of collision systems. Journal of Computational Physics, 179(2):539–556, July 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970737.

McClarren:2009:MIM

[MU09] Ryan G. McClarren and Todd J. Urbatsch. A modified implicit Monte Carlo method for time-dependent radiative transfer with adaptive material coupling. Journal of Computational Physics, 228(16):5669–5686, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002095.

Munger:2006:LSA

[MV06] David Munger and Alain Vincent. A level set approach to simulate magnetohydrodynamic instabilities in aluminum reduction cells. Journal of Computational Physics, 217(2):295–311, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000118.

Moyle:2008:LRL

[MV08] Keri R. Moyle and Yiannis Ventikos. Local remeshing for large amplitude grid deformations. Journal of Computational Physics, 227(5):2781–2793, February 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005037.

Manna:2004:PSM

[MVD04] Marcello Manna, Andrea Vacca, and Michel O. Deville. Preconditioned spectral multi-domain discretization of the incompressible Navier–Stokes equations. Journal of Computational Physics, 201(1):204–223, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002098.

Marsden:2002:CCF

[MVM02] Alison L. Marsden, Oleg V. Vasilyev, and Parviz Moin. Construction of commutative filters for LES on unstructured meshes. Journal of Computational Physics, 175(2):584–603, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910196958X.

Morinishi:2004:FCF

[MVO04] Youhei Morinishi, Oleg V. Vasilyev, and Takeshi Ogi. Fully conservative finite difference scheme in cylindrical coordinates for incompressible flow simulations. Journal of Computational Physics, 197(2):686–710, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006594.

Mark:2008:DVN

[MvW08] Andreas Mark and Berend G. M. van Wachem. Derivation and validation of a novel implicit second-order accurate immersed boundary method. Journal of Computational Physics, 227(13):6660–6680, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001770.

Morgan:2006:NNA

[MWG^{+}06] N. M. Morgan, C. G. Wells, M. J. Goodson, M. Kraft, and W. Wagner. A new numerical approach for the simulation of the growth of inorganic nanoparticles. Journal of Computational Physics, 211(2):638–658, January 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002913.

Madzvamuse:2003:MGF

[MWM03] Anotida Madzvamuse, Andrew J. Wathen, and Philip K. Maini. A moving grid finite element method applied to a model biological pattern generator. Journal of Computational Physics, 190(2):478–500, September 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002948.

Mani:2008:RRA

[MWM08] Ali Mani, Meng Wang, and Parviz Moin. Resolution requirements for aero-optical simulations. Journal of Computational Physics, 227(21):9008–9020, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800096X.

Meng:2003:UTG

[MY03] Hua Meng and Vigor Yang. A unified treatment of general fluid thermodynamics and its application to a preconditioning scheme. Journal of Computational Physics, 189(1):277–304, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002110.

Ma:2006:QAF

[MY06a] Q. W. Ma and S. Yan. Quasi ALE finite element method for nonlinear water waves. Journal of Computational Physics, 212(1):52–72, February 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003098.

Mavriplis:2006:CDG

[MY06b] Dimitri J. Mavriplis and Zhi Yang. Construction of the discrete geometric conservation law for high-order timeaccurate simulations on dynamic meshes. Journal of Computational Physics, 213(2):557–573, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500392X.

Ming:2006:NMM

[MY06c] Pingbing Ming and Xingye Yue. Numerical methods for multiscale elliptic problems. Journal of Computational Physics, 214(1):421–445, May 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004523.

Masters:2007:OFS

[MY07] Nathan D. Masters and Wenjing Ye. Octant flux splitting information preservation DSMC method for thermally driven flows. Journal of Computational Physics, 226(2):2044–2062, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910700277X.

Matsuo:2009:ECG

[MY09] Takayasu Matsuo and Hisashi Yamaguchi. An energyconserving Galerkin scheme for a class of nonlinear dispersive equations. Journal of Computational Physics, 228(12): 4346–4358, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910900117X.

Myong:2001:CME

[Myo01] R. S. Myong. A computational method for Eu’s generalized hydrodynamic equations of rarefied and microscale gasdynamics. Journal of Computational Physics, 168(1):47–72, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966786.

Myong:2004:GHC

[Myo04] R. S. Myong. A generalized hydrodynamic computational model for rarefied and microscale diatomic gas flows. Journal of Computational Physics, 195(2):655–676, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005771.

Morel:2007:LMG

[MYW07] Jim E. Morel, T.-Y. Brian Yang, and James S. Warsa. Linear multifrequency-grey acceleration recast for preconditioned Krylov iterations. Journal of Computational Physics, 227(1):244–263, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003294.

Mencinger:2007:FVD

[MZ07] Jure Mencinger and Iztok Zun. On the finite volume discretization of discontinuous body force field on collocated grid: Application to VOF method. Journal of Computational Physics, 221(2):524–538, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003019.

Ma:2008:SSF

[MZ08] Xiang Ma and Nicholas Zabaras. A stabilized stochastic finite element second-order projection method for modeling natural convection in random porous media. Journal of Computational Physics, 227(18):8448–8471, September 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003276.

Ma:2009:AHS

[MZ09] Xiang Ma and Nicholas Zabaras. An adaptive hierarchical sparse grid collocation algorithm for the solution of stochastic differential equations. Journal of Computational Physics, 228(8):3084–3113, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900028X.

Nemec:2008:ASC

[NA08] Marian Nemec and Michael J. Aftosmis. Adjoint sensitivity computations for an embedded-boundary Cartesian mesh method. Journal of Computational Physics, 227(4):2724–2742, February 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005013.

Nash:2008:NFO

[Nas08] Patrick L. Nash. A new fourth-order Fourier–Bessel splitstep method for the extended nonlinear Schrödinger equation. Journal of Computational Physics, 227(3):2073–2082, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004470.

Nataf:2006:NAP

[Nat06] Frédéric Nataf. A new approach to perfectly matched layers for the linearized Euler system. Journal of Computational Physics, 214(2):757–772, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004729.

Novak:2004:ABC

[NB04] Jérôme Novak and Silvano Bonazzola. Absorbing boundary conditions for simulation of gravitational waves with spectral methods in spherical coordinates. Journal of Computational Physics, 197(1):186–196, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006259.

Nestor:2009:EFV

[NBLQ09] Ruairi M. Nestor, Mihai Basa, Martin Lastiwka, and Nathan J. Quinlan. Extension of the finite volume particle method to viscous flow. Journal of Computational Physics, 228(5):1733–1749, March 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005913.

Nordstrom:2001:HOF

[NC01] Jan Nordström and Mark H. Carpenter. High-order finite difference methods, multidimensional linear problems, and curvilinear coordinates. Journal of Computational Physics, 173(1):149–174, October 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968640.

Nieter:2004:VVP

[NC04] Chet Nieter and John R. Cary. VORPAL: a versatile plasma simulation code. Journal of Computational Physics, 196(2):448–473, May 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103006041.

Niu:2003:SFA

[NCS03] X. D. Niu, Y. T. Chew, and C. Shu. Simulation of flows around an impulsively started circular cylinder by Taylor series expansion- and least squares-based lattice Boltzmann method. Journal of Computational Physics, 188(1):176–193, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910300161X.

Nieter:2009:ADM

[NCW^{+}09] C. Nieter, John R. Cary, Gregory R. Werner, David N. Smithe, and Peter H. Stoltz. Application of Dey–Mittra conformal boundary algorithm to 3D electromagnetic modeling. Journal of Computational Physics, 228(21):7902–7916, November 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003829.

Nguyen:2004:SCM

[ND04] S. Nguyen and C. Delcarte. A spectral collocation method to solve Helmholtz problems with boundary conditions involving mixed tangential and normal derivatives. Journal of Computational Physics, 200(1):34–49, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001299.

Nestler:2005:CGP

[NDG05] B. Nestler, D. Danilov, and P. Galenko. Crystal growth of pure substances: Phase-field simulations in comparison with analytical and experimental results. Journal of Computational Physics, 207(1):221–239, July 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000240.

Nourgaliev:2006:ACB

[NDT06] R. R. Nourgaliev, T. N. Dinh, and T. G. Theofanous. Adaptive characteristics-based matching for compressible multifluid dynamics. Journal of Computational Physics, 213(2):500–529, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003906.

Nordbotten:2005:DQG

[NE05] Jan Martin Nordbotten and Geir Terje Eigestad. Discretization on quadrilateral grids with improved monotonicity properties. Journal of Computational Physics, 203(2):744–760, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004036.

Neofytou:2007:RCB

[Neo07] Panagiotis Neofytou. Revision of the characteristics-based scheme for incompressible flows. Journal of Computational Physics, 222(2):475–484, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004852.

Nonomura:2009:EDS

[NF09] Taku Nonomura and Kozo Fujii. Effects of difference scheme type in high-order weighted compact nonlinear schemes. Journal of Computational Physics, 228(10):3533–3539, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000837.

Noullez:2003:HBA

[NFA03] Alain Noullez, Duccio Fanelli, and Erik Aurell. A heap-based algorithm for the study of one-dimensional particle systems. Journal of Computational Physics, 186(2):697–703, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000482.

Newren:2007:USD

[NFGK07] Elijah P. Newren, Aaron L. Fogelson, Robert D. Guy, and Robert M. Kirby. Unconditionally stable discretizations of the immersed boundary equations. Journal of Computational Physics, 222(2):702–719, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003743.

Nguyen:2001:BCC

[NFK01] Duc Q. Nguyen, Ronald P. Fedkiw, and Myungjoo Kang. A boundary condition capturing method for incompressible flame discontinuities. Journal of Computational Physics, 172 (1):71–98, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968123.

Nedea:2006:DDD

[NFvS^{+}06] S. V. Nedea, A. J. H. Frijns, A. A. van Steenhoven, A. P. J. Jansen, A. J. Markvoort, and P. A. J. Hilbers. Density distribution for a dense hard-sphere gas in micro/nano-channels: Analytical and simulation results. Journal of Computational Physics, 219(2):532–552, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001914.

Neelov:2006:ENQ

[NG06a] A. I. Neelov and S. Goedecker. An efficient numerical quadrature for the calculation of the potential energy of wavefunctions expressed in the Daubechies wavelet basis. Journal of Computational Physics, 217(2):312–339, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600012X.

Nordstrom:2006:SHM

[NG06b] Jan Nordström and Jing Gong. A stable hybrid method for hyperbolic problems. Journal of Computational Physics, 212(2):436–453, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910500327X.

Novak:2007:DCS

[NGC^{+}07] Igor L. Novak, Fei Gao, Yung-Sze Choi, Diana Resasco, James C. Schaff, and Boris M. Slepchenko. Diffusion on a curved surface coupled to diffusion in the volume: Application to cell biology. Journal of Computational Physics, 226(2):1271–1290, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002331.

Nguyen:2007:PEE

[Ngu07] N. C. Nguyen. A posteriori error estimation and basis adaptivity for reduced-basis approximation of nonaffine-parametrized linear elliptic partial differential equations. Journal of Computational Physics, 227(2):983–1006, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003749.

Nguyen:2008:MRB

[Ngu08] N. C. Nguyen. A multiscale reduced-basis method for parametrized elliptic partial differential equations with multiple scales. Journal of Computational Physics, 227(23): 9807–9822, December 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108004087.

Nordstrom:2009:SCH

[NGvdWS09] Jan Nordström, Jing Gong, Edwin van der Weide, and Magnus Svärd. A stable and conservative high order multi-block method for the compressible Navier–Stokes equations. Journal of Computational Physics, 228(24):9020–9035, December 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004902.

Nihei:2003:FSS

[NI03] Tomonori Nihei and Katsuya Ishii. A fast solver of the shallow water equations on a sphere using a combined compact difference scheme. Journal of Computational Physics, 187(2):639–659, May 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001529.

Ni:2009:CPM

[Ni09] Ming-Jiu Ni. Consistent projection methods for variable density incompressible Navier–Stokes equations with continuous surface forces on a rectangular collocated mesh. Journal of Computational Physics, 228(18):6938–6956, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003313.

Nicoud:2000:CHO

[Nic00] F. Nicoud. Conservative high-order finite-difference schemes for low-Mach number flows. Journal of Computational Physics, 158(1):71–97, February 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999199964082.

Nicholls:2009:RBP

[Nic09] David P. Nicholls. A rapid boundary perturbation algorithm for scattering by families of rough surfaces. Journal of Computational Physics, 228(9):3405–3420, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000436.

Nie:2001:NEV

[Nie01] Qing Nie. The nonlinear evolution of vortex sheets with surface tension in axisymmetric flows. Journal of Computational Physics, 174(1):438–459, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969268.

Nikitin:2006:FDM

[Nik06] Nikolay Nikitin. Finite-difference method for incompressible Navier–Stokes equations in arbitrary orthogonal curvilinear coordinates. Journal of Computational Physics, 217(2):759–781, September 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000398.

Nishikawa:2007:FOS

[Nis07] Hiroaki Nishikawa. A first-order system approach for diffusion equation. I: Second-order residual-distribution schemes. Journal of Computational Physics, 227(1):315–352, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910700335X.

Nitsche:2001:SFC

[Nit01] Monika Nitsche. Singularity formation in a cylindrical and a spherical vortex sheet. Journal of Computational Physics, 173(1):208–230, October 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196872X.

Nita:2005:ATC

[Nit05] Gelu M. Nita. About two cluster generating algorithms. Journal of Computational Physics, 206(2):578–596, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000069.

Nagrath:2006:HSA

[NJLA06] Sunitha Nagrath, Kenneth Jansen, Richard T. Lahey, Jr., and Iskander Akhatov. Hydrodynamic simulation of air bubble implosion using a level set approach. Journal of Computational Physics, 215(1):98–132, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500481X.

Ni:2008:DSB

[NJX08a] Guoxi Ni, Song Jiang, and Kun Xu. A DGBGK scheme based on WENO limiters for viscous and inviscid flows. Journal of Computational Physics, 227(11):5799–5815, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001149.

Ni:2008:EKS

[NJX08b] Guoxi Ni, Song Jiang, and Kun Xu. Efficient kinetic schemes for steady and unsteady flow simulations on unstructured meshes. Journal of Computational Physics, 227(6):3015–3031, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002574.

Ni:2009:RFA

[NJX09] Guoxi Ni, Song Jiang, and Kun Xu. Remapping-free ALEtype kinetic method for flow computations. Journal of Computational Physics, 228(8):3154–3171, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000333.

Nishikawa:2008:VSC

[NK08] Hiroaki Nishikawa and Keiichi Kitamura. Very simple, carbuncle-free, boundary-layer-resolving, rotated-hybrid Riemann solvers. Journal of Computational Physics, 227(4):2560–2581, February 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004822.

Natvig:2008:FCM

[NL08] Jostein R. Natvig and Knut-Andreas Lie. Fast computation of multiphase flow in porous media by implicit discontinuous Galerkin schemes with optimal ordering of elements. Journal of Computational Physics, 227(24):10108–10124, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108004440.

Nouy:2009:GSD

[NL09] Anthony Nouy and Olivier P. Le Maître. Generalized spectral decomposition for stochastic nonlinear problems. Journal of Computational Physics, 228(1):202–235, January 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004737.

Nagarajan:2003:RHO

[NLF03] Santhanam Nagarajan, Sanjiva K. Lele, and Joel H. Ferziger. A robust high-order compact method for large eddy simulation. Journal of Computational Physics, 191(2):392– 419, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910300322X.

Nishimura:2006:FEP

[NLLE06] Y. Nishimura, Z. Lin, J. L. V. Lewandowski, and S. Ethier. A finite element Poisson solver for gyrokinetic particle simulations in a global field aligned mesh. Journal of Computational Physics, 214(2):657–671, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004675.

Nielsen:2007:URP

[NLT07] Bjørn Fredrik Nielsen, Marius Lysaker, and Aslak Tveito. On the use of the resting potential and level set methods for identifying ischemic heart disease: An inverse problem. Journal of Computational Physics, 220(2):772–790, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002592.

Nourgaliev:2008:NPI

[NLT08] R. R. Nourgaliev, M.-S. Liou, and T. G. Theofanous. Numerical prediction of interfacial instabilities: Sharp interface method (SIM). Journal of Computational Physics, 227 (8):3940–3970, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107005566.

Nastase:2006:HOD

[NM06] Cristian R. Nastase and Dimitri J. Mavriplis. High-order discontinuous Galerkin methods using an hp-multigrid approach. Journal of Computational Physics, 213(1):330–357, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003839.

Ng:2009:EFS

[NMG09] Yen Ting Ng, Chohong Min, and Frédéric Gibou. An efficient fluid-solid coupling algorithm for single-phase flows. Journal of Computational Physics, 228(23):8807–8829, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004719.

Ni:2007:CDCb

[NMH^{+}07] Ming-Jiu Ni, Ramakanth Munipalli, Peter Huang, Neil B. Morley, and Mohamed A. Abdou. A current density conservative scheme for incompressible MHD flows at a low magnetic Reynolds number. Part II: On an arbitrary collocated mesh. Journal of Computational Physics, 227(1):205–228, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003269.

Ni:2007:CDCa

[NMM^{+}07] Ming-Jiu Ni, Ramakanth Munipalli, Neil B. Morley, Peter Huang, and Mohamed A. Abdou. A current density conservative scheme for incompressible MHD flows at a low magnetic Reynolds number. Part I: On a rectangular collocated grid system. Journal of Computational Physics, 227(1):174–204, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003270.

Nordstrom:2007:BCD

[NMS07] Jan Nordström, Ken Mattsson, and Charles Swanson. Boundary conditions for a divergence free velocity-pressure formulation of the Navier–Stokes equations. Journal of Computational Physics, 225(1):874–890, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000083.

Nicholls:2004:ENR

[NN04] David P. Nicholls and Nilima Nigam. Exact non-reflecting boundary conditions on general domains. Journal of Computational Physics, 194(1):278–303, February 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004832.

Nishida:2009:ASS

[NN09] Hiroyuki Nishida and Taku Nonomura. ADI–SGS scheme on ideal magnetohydrodynamics. Journal of Computational Physics, 228(9):3182–3188, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000588.

Noelle:2000:MIN

[Noe00] Sebastian Noelle. The MoT–ICE: a new high-resolution wavepropagation algorithm for multidimensional systems of conservation laws based on Fey’s method of transport. Journal of Computational Physics, 164(2):283–334, November 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965987.

Nejat:2008:EDO

[NOG08a] Amir Nejat and Carl Ollivier-Gooch. Effect of discretization order on preconditioning and convergence of a high-order unstructured Newton–GMRES solver for the Euler equations. Journal of Computational Physics, 227(4):2366–2386, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004676.

Nejat:2008:HOA

[NOG08b] Amir Nejat and Carl Ollivier-Gooch. A high-order accurate unstructured finite volume Newton–Krylov algorithm for inviscid compressible flows. Journal of Computational Physics, 227(4):2582–2609, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004834.

Novikov:2004:EVC

[Nov04] Alexei Novikov. Eddy viscosity of cellular flows by upscaling. Journal of Computational Physics, 195(1):341–354, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005357.

Nguyen:2009:IHOa

[NPC09a] N. C. Nguyen, J. Peraire, and B. Cockburn. An implicit high-order hybridizable discontinuous Galerkin method for linear convection-diffusion equations. Journal of Computational Physics, 228(9):3232–3254, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000308.

Nguyen:2009:IHOb

[NPC09b] N. C. Nguyen, J. Peraire, and B. Cockburn. An implicit highorder hybridizable discontinuous Galerkin method for nonlinear convection-diffusion equations. Journal of Computational Physics, 228(23):8841–8855, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004756.

Nakshatrala:2009:DSD

[NPH09] K. B. Nakshatrala, A. Prakash, and K. D. Hjelmstad. On dual Schur domain decomposition method for linear first-order transient problems. Journal of Computational Physics, 228(21):7957–7985, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003945.

Noelle:2006:WBF

[NPPN06] Sebastian Noelle, Normann Pankratz, Gabriella Puppo, and Jostein R. Natvig. Well-balanced finite volume schemes of arbitrary order of accuracy for shallow water flows. Journal of Computational Physics, 213(2):474–499, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500389X.

Nicholls:2001:SHO

[NR01] David P. Nicholls and Fernando Reitich. Stability of high-order perturbative methods for the computation of Dirichlet–Neumann operators. Journal of Computational Physics, 170(1):276–298, June 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967373.

Nitsche:2004:NSI

[NS04] Monika Nitsche and Paul H. Steen. Numerical simulations of inviscid capillary pinchoff. Journal of Computational Physics, 200(1):299–324, October 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001688.

Noskov:2005:ICS

[NS05] Mikhail Noskov and Mitchell D. Smooke. An implicit compact scheme solver with application to chemically reacting flows. Journal of Computational Physics, 203(2):700–730, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104003997.

Nam:2009:TBV

[NSC09] Jaewook Nam, L. E. Scriven, and Marcio S. Carvalho. Tracking birth of vortex in flows. Journal of Computational Physics, 228(12):4549–4567, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001430.

Nair:2003:FTG

[NSS03] Ramachandran D. Nair, Jeffrey S. Scroggs, and Frederick H. M. Semazzi. A forward-trajectory global semi-Lagrangian transport scheme. Journal of Computational Physics, 190(1):275–294, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002742.

Nourgaliev:2007:HFI

[NT07] R. R. Nourgaliev and T. G. Theofanous. High-fidelity interface tracking in compressible flows: Unlimited anchored adaptive level set. Journal of Computational Physics, 224(2):836–866, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005511.

Nikolopoulos:2007:TDN

[NTB07] N. Nikolopoulos, A. Theodorakakos, and G. Bergeles. Threedimensional numerical investigation of a droplet impinging normally onto a wall film. Journal of Computational Physics, 225(1):322–341, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005936.

Nakamura:2007:LMM

[NTO^{+}07] Yoshimichi Nakamura, Norihiko Takahashi, Masakuni Okamoto, Tsuyoshi Uda, and Takahisa Ohno. Link molecule method for quantum mechanical/molecular mechanical hybrid simulations. Journal of Computational Physics, 225(2):1985–1993, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000988.

Nakamura:2001:ECS

[NTYT01] Takashi Nakamura, Ryotaro Tanaka, Takashi Yabe, and Kenji Takizawa. Exactly conservative semi-Lagrangian scheme for multi-dimensional hyperbolic equations with directional splitting technique. Journal of Computational Physics, 174(1):171–207, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968883. See erratum [NTYT02].

Nakamura:2002:EEC

[NTYT02] Takashi Nakamura, Ryotaro Tanaka, Takashi Yabe, and Kenji Takizawa. Erratum: “Exactly Conservative Semi-Lagrangian Scheme for Multi-dimensional Hyperbolic Equations with Directional Splitting Technique”: Volume 174, Number 1 (2001), pages 171–207. Journal of Computational Physics, 175(2):792, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969840. See [NTYT01].

Nease:2009:TSA

[NU09] Brian R. Nease and Taro Ueki. Time series analysis and Monte Carlo methods for eigenvalue separation in neutron multiplication problems. Journal of Computational Physics, 228(23):8525–8547, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003957.

Nakshatrala:2009:NNM

[NV09] K. B. Nakshatrala and A. J. Valocchi. Non-negative mixed finite element formulations for a tensorial diffusion equation. Journal of Computational Physics, 228(18):6726–6752, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002897.

Nerinckx:2005:MUT

[NVD05] Krista Nerinckx, Jan Vierendeels, and Erik Dick. Machuniformity through the coupled pressure and temperature correction algorithm. Journal of Computational Physics, 206(2):597–623, July 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000070.

Nerinckx:2007:MUA

[NVD07] Krista Nerinckx, Jan Vierendeels, and Erik Dick. A Machuniform algorithm: Coupled versus segregated approach.

Journal of Computational Physics, 224(1):314–331, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000782.

Nishikawa:2003:OMC

[NvL03] Hiroaki Nishikawa and Bram van Leer. Optimal multigrid convergence by elliptic/hyperbolic splitting. Journal of Computational Physics, 190(1):52–63, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002535.

Nagele:2007:IDS

[NW07]  Sandra Nägele and Gabriel Wittum.  On the influ-

ence of different stabilisation methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 224(1):100–116, May 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006176.

Nie:2008:CIF

[NWZL08] Qing Nie, Frederic Y. M. Wan, Yong-Tao Zhang, and Xin-Feng Liu. Compact integration factor methods in high spatial dimensions. Journal of Computational Physics, 227(10): 5238–5255, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108000752.

Noelle:2007:HOW

[NXS07] Sebastian Noelle, Yulong Xing, and Chi-Wang Shu. Highorder well-balanced finite volume WENO schemes for shallow water equation with moving water. Journal of Computational Physics, 226(1):29–58, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001416.

Nystrom:2002:HOT

[Nys02] J. F. Nystrom. High-order time-stable numerical boundary scheme for the temporally dependent Maxwell equations in two dimensions. Journal of Computational Physics, 178(2):290–306, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102970142.

Narayanan:2005:VMS

[NZ05] Velamur Asokan Badri Narayanan and Nicholas Zabaras. Variational multiscale stabilized FEM formulations for transport equations: stochastic advection-diffusion and incompressible stochastic Navier–Stokes equations. Journal of Computational Physics, 202(1):94–133, January 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002694.

Nutaro:2007:SPD

[NZ07] James Nutaro and Bernard Zeigler. On the stability and performance of discrete event methods for simulating continuous systems. Journal of Computational Physics, 227(1):797–819, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003609.

Nie:2006:ESI

[NZZ06] Qing Nie, Yong-Tao Zhang, and Rui Zhao. Efficient semi-implicit schemes for stiff systems. Journal of Computational Physics, 214(2):521–537, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004596.

Okongo:2002:CBC

[OB02] Nora Okong’o and Josette Bellan. Consistent boundary conditions for multicomponent real gas mixtures based on characteristic waves. Journal of Computational Physics, 176(2): 330–344, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102969901.

Oberkampf:2006:MAB

[OB06] William L. Oberkampf and Matthew F. Barone. Measures of agreement between computation and experiment: Validation metrics. Journal of Computational Physics, 217(1):

5–36, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106001860.

Omang:2006:SSC

[OBT06] M. Omang, S. Børve, and J. Trulsen. SPH in spherical and cylindrical coordinates. Journal of Computational Physics, 213(1):391–412, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003852.

Olson:2008:GD

[OC08] Spencer E. Olson and Andrew J. Christlieb. Gridless DSMC. Journal of Computational Physics, 227(17):8035– 8064, September 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800257X.

Ortega-Casanova:2008:NMS

[OCFF08] J. Ortega-Casanova and R. Fernandez-Feria. A numerical method for the study of nonlinear stability of axisymmetric flows based on the vector potential. Journal of Computational Physics, 227(6):3307–3321, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005347.

Osher:2002:GOP

[OCK^{+}02] Stanley Osher, Li-Tien Cheng, Myungjoo Kang, Hyeseon Shim, and Yen-Hsi Tsai. Geometric optics in a phasespace-based level set and Eulerian framework. Journal of Computational Physics, 179(2):622–648, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970804.

Oger:2006:TDS

[ODAF06] G. Oger, M. Doring, B. Alessandrini, and P. Ferrant. Twodimensional SPH simulations of wedge water entries. Journal of Computational Physics, 213(2):803–822, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004171.

Oger:2007:ISM

[ODAF07] G. Oger, M. Doring, B. Alessandrini, and P. Ferrant. An improved SPH method: Towards higher order convergence. Journal of Computational Physics, 225(2):1472–1492, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000630.

Ovtchinnikov:2007:ASB

[ODCK07] S. Ovtchinnikov, F. Dobrian, X.-C. Cai, and D. E. Keyes. Additive Schwarz-based fully coupled implicit methods for resistive Hall magnetohydrodynamic problems. Journal of Computational Physics, 225(2):1919–1936, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000939.

Osher:2001:LSMa

[OF01] Stanley Osher and Ronald P. Fedkiw. Level set methods: An overview and some recent results. Journal of Computational Physics, 169(2):463–502, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966361.

Oliveira:2002:TOF

[OF02] Anabela Oliveira and André B. Fortunato. Toward an oscillation-free, mass conservative, Eulerian–Lagrangian transport model. Journal of Computational Physics, 183(1):142–164, November 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971809.

Ohwada:2006:SDH

[OF06] Taku Ohwada and Satoshi Fukata. Simple derivation of highresolution schemes for compressible flows by kinetic approach. Journal of Computational Physics, 211(2):424–447, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002810.

Ollivier-Gooch:2002:HOA

[OGV02] Carl Ollivier-Gooch and Michael Van Altena. A highorder-accurate unstructured mesh finite-volume scheme for the advection-diffusion equation. Journal of Computational Physics, 181(2):729–752, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971597.

Oh:2004:AMS

[Oh04] Hae-Soo Oh. Accurate mode-separated energy release rates for delamination cracks. Journal of Computational Physics, 193(1):86–114, January 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004200.

Ohwada:2002:CKS

[Ohw02] Taku Ohwada. On the construction of kinetic schemes. Journal of Computational Physics, 177(1):156–175, March 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970087.

Oroskar:2006:ECP

[OJW06] Siddharth Oroskar, David R. Jackson, and Donald R. Wilton. Efficient computation of the 2D periodic Green’s function using the Ewald method. Journal of Computational Physics, 219(2):899–911, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003354.

Ohwada:2004:MDR

[OK04] Taku Ohwada and Seijiro Kobayashi. Management of discontinuous reconstruction in kinetic schemes. Journal of Computational Physics, 197(1):116–138, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006223.

Olsson:2005:CLS

[OK05] Elin Olsson and Gunilla Kreiss. A conservative level set method for two phase flow. Journal of Computational Physics, 210(1):225–246, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002184.

Oevermann:2006:CGF

[OK06a] M. Oevermann and R. Klein. A Cartesian grid finite volume method for elliptic equations with variable coefficients and embedded interfaces. Journal of Computational Physics, 219(2):749–769, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002129.

Omelchenko:2006:EDH

[OK06b] Y. A. Omelchenko and H. Karimabadi. Event-driven, hybrid particle-in-cell simulation: a new paradigm for multi-scale plasma modeling. Journal of Computational Physics, 216(1):153–178, July 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005474.

Omelchenko:2006:SAT

[OK06c] Y. A. Omelchenko and H. Karimabadi. Self-adaptive time integration of flux-conservative equations with sources. Journal of Computational Physics, 216(1):179–194, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005619.

Omelchenko:2007:TAE

[OK07a] Y. A. Omelchenko and H. Karimabadi. A time-accurate explicit multi-scale technique for gas dynamics. Journal of Computational Physics, 226(1):282–300, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700157X.

Ozgun:2007:NFP

[OK07b] Ozlem Ozgun and Mustafa Kuzuoğlu. Near-field performance analysis of locally-conformal perfectly matched absorbers via Monte Carlo simulations. Journal of Computational Physics, 227(2):1225–1245, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003919.

Oh:2001:NMO

[OKL01] Hae-Soo Oh, Hoonjoo Kim, and Sung-Jin Lee. The numerical methods for oscillating singularities in elliptic boundary value problems. Journal of Computational Physics, 170(2):742–763, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967592.

Oosterlee:2007:P

[OKV07] Kees Oosterlee, Barry Koren, and Kees Vuik. Preface. Journal of Computational Physics, 224(1):1–2, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001398.

Olsson:2007:CLS

[OKZ07] Elin Olsson, Gunilla Kreiss, and Sara Zahedi. A conservative level set method for two phase flow II. Journal of Computational Physics, 225(1):785–807, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000046.

Omnes:2001:SCN

[OL01] P. Omnes and P. Louvet. Self-consistent numerical simulation of isotope separation by selective ion cyclotron resonance heating in a magnetically confined plasma. Journal of Computational Physics, 172(1):326–347, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968263.

OConnor:2008:ARE

[OLA08] Patrick D. O’Connor, Lyle N. Long, and James B. Anderson. Accurate rate expressions for simulations of gas-phase chemical reactions. Journal of Computational Physics, 227(16):7664–7673, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002556.

Ong:2003:FAT

[OLLL03] E. T. Ong, K. M. Lim, K. H. Lee, and H. P. Lee. A fast algorithm for three-dimensional potential fields calculation: fast Fourier transform on multipoles. Journal of Computational Physics, 192(1):244–261, November 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003772.

Olson:2007:ESM

[Ols07] Gordon L. Olson. Efficient solution of multi-dimensional fluxlimited nonequilibrium radiation diffusion coupled to material conduction with second-order time discretization. Journal of Computational Physics, 226(1):1181–1195, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002276.

Olson:2009:SOT

[Ols09] Gordon L. Olson. Second-order time evolution of P_{N} equations for radiation transport. Journal of Computational Physics, 228(8):3072–3083, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000278.

Othmer:2002:CSC

[OMG02] C. Othmer, U. Motschmann, and K. H. Glassmeier. Creation of spatial charge separation in plasmas with rigorously charge-conserving local electromagnetic field solvers. Journal of Computational Physics, 180(1):99–109, July 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970774.

Ozgun:2009:PCB

[OMK09] Ozlem Ozgun, Raj Mittra, and Mustafa Kuzuoğlu. Parallelized characteristic basis finite element method (CBFEM– MPI) — a non-iterative domain decomposition algorithm for electromagnetic scattering problems. Journal of Computational Physics, 228(6):2225–2238, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006293.

Otani:2008:PFM

[ON08] Yoshihiro Otani and Naoshi Nishimura. A periodic FMM for Maxwell’s equations in 3D and its applications to problems related to photonic crystals. Journal of Computational Physics, 227(9):4630–4652, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000442.

Oden:2002:EME

[OP02] J. Tinsley Oden and Serge Prudhomme. Estimation of modeling error in computational mechanics. Journal of Computational Physics, 182(2):496–515, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971834.

Onofri:2007:CMN

[OPML07] Marco Onofri, Leonardo Primavera, Francesco Malara, and Pasquale Londrillo. A compressible magnetohydrodynamic numerical code with time-dependent boundary conditions in cylindrical geometry. Journal of Computational Physics, 226(2):1874–1890, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002677.

Oliver:2006:SME

[ORM06] H. J. Oliver, A. H. Reiman, and D. A. Monticello. Solving the 3D MHD equilibrium equations in toroidal geometry by Newton’s method. Journal of Computational Physics, 211(1):99–128, January 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002640.

Osher:2001:LSMb

[OS01] Stanley J. Osher and Fadil Santosa. Level set methods for optimization problems involving geometry and constraints: I. Frequencies of a two-density inhomogeneous drum. Journal of Computational Physics, 171(1):272–288, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967890.

Ober:2004:SAT

[OS04] Curtis C. Ober and John N. Shadid. Studies on the accuracy of time-integration methods for the radiation-diffusion equations. Journal of Computational Physics, 195(2):743–772, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005813.

Oevermann:2009:SIF

[OSK09] M. Oevermann, C. Scharfenberg, and R. Klein. A sharp interface finite volume method for elliptic equations on Cartesian grids. Journal of Computational Physics, 228(14):5184–5206, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001971.

Osher:2001:P

[OT01] S. J. Osher and G. Tryggvason. PREFACE. Journal of Computational Physics, 169(2):249, May 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968044.

Otani:2000:CMC

[Ota00] Niels F. Otani. Computer modeling in cardiac electrophysiology. Journal of Computational Physics, 161(1):21–34, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964428.

Oishi:2008:ITS

[OTCM08] Cassio M. Oishi, Murilo F. Tomé, José A. Cuminato, and Sean McKee. An implicit technique for solving 3D low Reynolds number moving free surface flows. Journal of Computational Physics, 227(16):7446–7468, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002271.

Oden:2000:ELM

[OV00] J. Tinsley Oden and Kumar S. Vemaganti. Estimation of local modeling error and goal-oriented adaptive modeling of heterogeneous materials: I. Error estimates and adaptive algorithms. Journal of Computational Physics, 164(1):22–47, October 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965859.

Onate:2007:MIF

[OVG07] Eugenio Oñate, Aleix Valls, and Julio García. Modeling incompressible flows at low and high Reynolds numbers via a finite calculus-finite element approach. Journal of Computational Physics, 224(1):332–351, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000794.

Ovtchinnikov:2008:CSE

[Ovt08] E. E. Ovtchinnikov. Computing several eigenpairs of Hermitian problems by conjugate gradient iterations. Journal of Computational Physics, 227(22):9477–9497, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800363X.

Owen:2004:TAV

[Owe04] J. Michael Owen. A tensor artificial viscosity for SPH. Journal of Computational Physics, 201(2):601–629, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002578.

Ohwada:2004:KSF

[OX04] Taku Ohwada and Kun Xu. The kinetic scheme for the full-Burnett equations. Journal of Computational Physics, 201(1):315–332, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002311.

Puri:2000:SAC

[PA00] Rajeev K. Puri and Joerg Aichelin. Simulated annealing clusterization algorithm for studying the multifragmentation. Journal of Computational Physics, 162(1):245–266, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965343.

Polizzi:2005:SDA

[PA05] E. Polizzi and N. Ben Abdallah. Subband decomposition approach for the simulation of quantum electron transport in nanostructures. Journal of Computational Physics, 202(1):150–180, January 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002712.

Plyasunov:2007:ESS

[PA07a] Sergey Plyasunov and Adam P. Arkin. Efficient stochastic sensitivity analysis of discrete event systems. Journal of Computational Physics, 221(2):724–738, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003135.

Premnath:2007:TDM

[PA07b] Kannan N. Premnath and John Abraham. Three-dimensional multi-relaxation time (MRT) lattice-Boltzmann models for multiphase flow. Journal of Computational Physics, 224(2): 539–559, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106004815.

Panaretos:2007:EDL

[PAD07] Anastasios H. Panaretos, James T. Aberle, and Rodolfo E. Díaz. The effect of the 2-D Laplacian operator approximation on the performance of finite-difference time-domain schemes for Maxwell’s equations. Journal of Computational Physics, 227(1):513–536, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003452.

Paisley:2001:MSI

[Pai01] M. F. Paisley. Multigrid solution of the incompressible Navier–Stokes equations for density-stratified flow past three-dimensional obstacles. Journal of Computational Physics, 170(2):785–811, July 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967622.

Pal:2008:RDS

[Pal08] Barnana Pal. Relaxation dynamics in small clusters: a modified Monte Carlo approach. Journal of Computational Physics, 227(4):2666–2673, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004986.

Papadakis:2008:NPV

[Pap08] George Papadakis. A novel pressure-velocity formulation and solution method for fluid-structure interaction problems. Journal of Computational Physics, 227(6):3383–3404, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005542.

Parker:2002:NGP

[Par02] Scott E. Parker. Nearest-grid-point interpolation in gyrokinetic particle-in-cell simulation. Journal of Computational Physics, 178(2):520–532, May 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970397.

Paul:2007:CPQ

[Pau07] Gerald Paul. A complexity O(1) priority queue for event driven molecular dynamics simulations. Journal of Computational Physics, 221(2):615–625, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003081.

Pavlyukevich:2007:LFN

[Pav07] Ilya Pavlyukevich. Lévy flights, non-local search and simulated annealing. Journal of Computational Physics, 226 (2):1830–1844, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910700263X.

Plagne:2000:TBS

[PB00] Laurent Plagne and Jean-Yves Berthou. Tensorial basis spline collocation method for Poisson’s equation. Journal of Computational Physics, 157(2):419–440, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963386.

Protas:2004:CFR

[PBH04] Bartosz Protas, Thomas R. Bewley, and Greg Hagen. A computational framework for the regularization of adjoint analysis in multiscale PDE systems. Journal of Computational Physics, 195(1):49–89, March 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005114.

Polasek:2002:EEM

[PC02] Martin Polásek and Petr Cársky. Efficient evaluation of the matrix elements of the Coulomb potential between plane waves and Gaussians. Journal of Computational Physics, 181(1):1–8, September 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297124X.

Park:2006:UOS

[PC06a] Moongyu Park and John H. Cushman. On upscaling operatorstable Lévy motions in fractal porous media. Journal of Computational Physics, 217(1):159–165, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000362.

Pretorius:2006:AMR

[PC06b] Frans Pretorius and Matthew W. Choptuik. Adaptive mesh refinement for coupled elliptic-hyperbolic systems. Journal of Computational Physics, 218(1):246–274, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000805.

Parashar:2008:SFA

[PC08] R. Parashar and J. H. Cushman. Scaling the fractional advective-dispersive equation for numerical evaluation of microbial dynamics in confined geometries with sticky boundaries. Journal of Computational Physics, 227(13):6598–6611, June 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910800171X.

Propp:2000:NMT

[PCCD00] Richard M. Propp, Phillip Colella, William Y. Crutchfield, and Marcus S. Day. A numerical model for trickle bed reactors. Journal of Computational Physics, 165(2):311–333, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910096604X.

Philip:2008:IAM

[PCP08] Bobby Philip, Luis Chacón, and Michael Pernice. Implicit adaptive mesh refinement for 2D reduced resistive magnetohydrodynamics. Journal of Computational Physics, 227(20):8855–8874, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003598.

Podlubny:2009:MAD

[PCS^{+}09] Igor Podlubny, Aleksei Chechkin, Tomas Skovranek, YangQuan Chen, and Blas M. Vinagre Jara. Matrix approach to discrete fractional calculus II: Partial fractional differential equations. Journal of Computational Physics, 228(8):3137–3153, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000321.

Pandolfi:2001:NIU

[PD01] Maurizio Pandolfi and Domenic D’Ambrosio. Numerical instabilities in upwind methods: Analysis and cures for the “carbuncle” phenomenon. Journal of Computational Physics, 166(2):271–301, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096652X.

Pantano:2007:LND

[PDHP07] C. Pantano, R. Deiterding, D. J. Hill, and D. I. Pullin. A low numerical dissipation patch-based adaptive mesh refinement method for large-eddy simulation of compressible flows. Journal of Computational Physics, 221(1):63–87, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002750.

Poette:2009:UQS

[PDL09] Gaël Poëtte, Bruno Després, and Didier Lucor. Uncertainty quantification for systems of conservation laws. Journal of Computational Physics, 228(7):2443–2467, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006372.

Peeren:2003:SFA

[Pee03] G. N. Peeren. Stream function approach for determining optimal surface currents. Journal of Computational Physics, 191(1):305–321, October 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003206.

Perot:2000:CPU

[Per00] Blair Perot. Conservation properties of unstructured staggered mesh schemes. Journal of Computational Physics, 159(1):58–89, March 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100964246.

Petersson:2001:SPB

[Pet01] N. Anders Petersson. Stability of pressure boundary conditions for Stokes and Navier–Stokes equations. Journal of Computational Physics, 172(1):40–70, September 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967543.

Peter:2007:QDS

[Pet07] William Peter. Quiet direct simulation Monte–Carlo with random timesteps. Journal of Computational Physics, 221(1):1–8, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002713.

Picard:2001:ACS

[PFB01] Richard R. Picard, Mark Fitzgerald, and Michael J. Brown. Accelerating convergence in stochastic particle dispersion simulation codes. Journal of Computational Physics, 173(1):231–255, October 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968743.

Petitpas:2007:RPM

[PFSL07] Fabien Petitpas, Erwin Franquet, Richard Saurel, and Olivier Le Métayer. A relaxation-projection method for compressible flows. Part II: Artificial heat exchanges for multiphase shocks. Journal of Computational Physics, 225(2):2214–2248, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107001283.

Pan:2002:DSM

[PG02a] Tsorng-Whay Pan and Roland Glowinski. Direct simulation of the motion of neutrally buoyant circular cylinders in plane Poiseuille flow. Journal of Computational Physics, 181(1):260–279, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971238.

Proot:2002:LSS

[PG02b] M. M. J. Proot and M. I. Gerrtisma. Least-squares spectral elements applied to the Stokes problem. Journal of Computational Physics, 181(2):454–477, September 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971378.

Pierce:2004:ADE

[PG04] Niles A. Pierce and Michael B. Giles. Adjoint and defect error bounding and correction for functional estimates. Journal of Computational Physics, 200(2):769–794, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001883.

Papa:2005:CMD

[PGB05] M. Papa, G. Giuliani, and A. Bonasera. Constrained molecular dynamics II: An N-body approach to nuclear systems. Journal of Computational Physics, 208(2):403–415, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000847.

Prax:2008:CVM

[PGN08] C. Prax, F. Golanski, and L. Nadal. Control of the vorticity mode in the linearized Euler equations for hybrid aeroacoustic prediction. Journal of Computational Physics, 227(12):6044–6057, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001186.

Perrin:2006:EFD

[PH06] A. Perrin and H. H. Hu. An explicit finite-difference scheme for simulation of moving particles. Journal of Computational Physics, 212(1):166–187, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003153.

Perrin:2008:EFD

[PH08] A. Perrin and H. H. Hu. An explicit finite difference scheme with spectral boundary conditions for particulate flows. Journal of Computational Physics, 227(20):8776–8791, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108003161.

Patacchini:2009:ETR

[PH09] L. Patacchini and I. H. Hutchinson. Explicit timereversible orbit integration in particle in cell codes with static homogeneous magnetic field. Journal of Computational Physics, 228(7):2604–2615, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006608.

Pons:2006:MPC

[PHKF06] J.-P. Pons, G. Hermosillo, R. Keriven, and O. Faugeras. Maintaining the point correspondence in the level set framework. Journal of Computational Physics, 220(1):339–354, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002440.

He:2009:CSS

[pHL09] Li ping He and Yunxian Liu. A class of stable spectral methods for the Cahn–Hilliard equation. Journal of Computational Physics, 228(14):5101–5110, August 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001909.

Petersen:2008:EFL

[PHW08] Mark R. Petersen, Matthew W. Hecht, and Beth A. Wingate. Efficient form of the LANS-α turbulence model in a primitive-equation ocean model. Journal of Computational Physics, 227(11):5717–5735, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910800106X.

Pettersson:2009:NAB

[PIN09] Per Pettersson, Gianluca Iaccarino, and Jan Nordström. Numerical analysis of the Burgers’ equation in the presence of uncertainty. Journal of Computational Physics, 228(22):8394–8412, December 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004471.

Pirozzoli:2002:CHC

[Pir02] Sergio Pirozzoli. Conservative hybrid compact-WENO schemes for shock-turbulence interaction. Journal of Computational Physics, 178(1):81–117, May 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297021X.

Pirozzoli:2006:SPS

[Pir06] Sergio Pirozzoli. On the spectral properties of shock-capturing schemes. Journal of Computational Physics, 219(2):489– 497, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106003482.

Pirozzoli:2007:PAO

[Pir07] Sergio Pirozzoli. Performance analysis and optimization of finite-difference schemes for wave propagation problems. Journal of Computational Physics, 222(2):809–831, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003962.

Plapp:2000:MFD

[PK00] Mathis Plapp and Alain Karma. Multiscale finite-difference-diffusion-Monte-Carlo method for simulating dendritic solidification. Journal of Computational Physics, 165(2):592– 619, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100966348.

Park:2003:DMG

[PK03] Soo Hyung Park and Jang Hyuk Kwon. On the dissipation mechanism of Godunov-type schemes. Journal of Computational Physics, 188(2):524–542, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001918.

Pivkin:2005:NMI

[PK05] Igor V. Pivkin and George Em Karniadakis. A new method to impose no-slip boundary conditions in dissipative particle dynamics. Journal of Computational Physics, 207(1):114–128, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000197.

Poludnenko:2007:CFF

[PK07] Alexei Y. Poludnenko and Alexei M. Khokhlov. Computation of fluid flows in non-inertial contracting, expanding, and rotating reference frames. Journal of Computational Physics, 220(2):678–711, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002555.

Pandit:2007:THO

[PKD07] Swapan K. Pandit, Jiten C. Kalita, and D. C. Dalal. A transient higher order compact scheme for incompressible viscous flows on geometries beyond rectangular. Journal of Computational Physics, 225(1):1100–1124, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000319.

Park:2005:TDB

[PKKL05] Harold S. Park, Eduard G. Karpov, Patrick A. Klein, and Wing Kam Liu. Three-dimensional bridging scale analysis of dynamic fracture. Journal of Computational Physics, 207(2):588–609, August 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000422.

Pereira:2001:FOA

[PKP01] J. M. C. Pereira, M. H. Kobayashi, and J. C. F. Pereira. A fourth-order-accurate finite volume compact method for the incompressible Navier–Stokes solutions. Journal of Computational Physics, 167(1):217–243, February 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966737.

Pelekanos:2000:WFC

[PKvdB00] George Pelekanos, Ralph E. Kleinman, and Peter M. van den Berg. A weak form of the conjugate gradient FFT method for two-dimensional elastodynamics. Journal of Computational Physics, 160(2):597–611, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964763.

Piraux:2001:NIM

[PL01] Joël Piraux and Bruno Lombard. A new interface method for hyperbolic problems with discontinuous coefficients: Onedimensional acoustic example. Journal of Computational Physics, 168(1):227–248, March 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101966963.

Pretorius:2004:AMR

[PL04] Frans Pretorius and Luis Lehner. Adaptive mesh refinement for characteristic codes. Journal of Computational Physics, 198(1):10–34, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000166.

Putman:2007:FVT

[PL07] William M. Putman and Shian-Jiann Lin. Finite-volume transport on various cubed-sphere grids. Journal of Computational Physics, 227(1):55–78, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003105.

Protas:2008:ABOa

[PL08] Bartosz Protas and Wenyuan Liao. Adjoint-based optimization of PDEs in moving domains. Journal of Computational Physics, 227(4):2707–2723, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005001.

Park:2009:EMT

[PL09a] Won-Kwang Park and Dominique Lesselier. Electromagnetic MUSIC-type imaging of perfectly conducting, arclike cracks at single frequency. Journal of Computational Physics, 228(21):8093–8111, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004173.

Patil:2009:FVT

[PL09b] Dhiraj V. Patil and K. N. Lakshmisha. Finite volume TVD formulation of lattice Boltzmann simulation on unstructured mesh. Journal of Computational Physics, 228(14):5262–5279, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002010.

Petersen:2009:HMP

[PLS^{+}09] Dan Erik Petersen, Song Li, Kurt Stokbro, Hans Henrik B. Sørensen, Per Christian Hansen, Stig Skelboe, and Eric Darve. A hybrid method for the parallel computation of Green’s functions. Journal of Computational Physics, 228(14):5020–5039, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001648.

Povitsky:2000:HOC

[PM00] Alex Povitsky and Philip J. Morris. A higher-order compact method in space and time based on parallel implementation of the Thomas algorithm. Journal of Computational Physics, 161(1):182–203, June 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964970.

Parshikov:2002:SPH

[PM02] Anatoly N. Parshikov and Stanislav A. Medin. Smoothed particle hydrodynamics using interparticle contact algorithms. Journal of Computational Physics, 180(1):358–382, July 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970993.

Park:2007:ANE

[PM07] Noma Park and Krishnan Mahesh. Analysis of numerical errors in large eddy simulation using statistical closure theory. Journal of Computational Physics, 222(1):194–216, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003512.

Park:2008:VES

[PM08] Noma Park and Krishnan Mahesh. A velocity-estimation subgrid model constrained by subgrid scale dissipation. Journal of Computational Physics, 227(8):4190–4206, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005700.

Popov:2008:ATA

[PMP08] Pavel P. Popov, Randall McDermott, and Stephen B. Pope. An accurate time advancement algorithm for particle tracking. Journal of Computational Physics, 227(20):8792–8806, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108003446.

Perot:2003:MUS

[PN03] Blair Perot and Ramesh Nallapati. A moving unstructured staggered mesh method for the simulation of incompressible free-surface flows. Journal of Computational Physics, 184(1):192–214, January 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910200027X.

Park:2009:PBP

[PNMK09] HyeongKae Park, Robert R. Nourgaliev, Richard C. Martineau, and Dana A. Knoll. On physics-based preconditioning of the Navier–Stokes equations. Journal of Computational Physics, 228(24):9131–9146, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005002.

Prosperetti:2001:PNM

[PO01] A. Prosperetti and H. N. Ogğuz. Physalis: a new o(N) method for the numerical simulation of disperse systems: Potential flow of spheres. Journal of Computational Physics, 167(1):196–216, February 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966671.

Ponta:2005:KLE

[Pon05] F. L. Ponta. The kinematic Laplacian equation method. Journal of Computational Physics, 207(2):405–426, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000355.

Pontaza:2006:LSF

[Pon06] J. P. Pontaza. A least-squares finite element formulation for unsteady incompressible flows with improved velocitypressure coupling. Journal of Computational Physics, 217(2):563–588, September 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000234.

Pontaza:2007:NCS

[Pon07a] J. P. Pontaza. A new consistent splitting scheme for incompressible Navier–Stokes flows: a least-squares spectral element implementation. Journal of Computational Physics, 225(2):1590–1602, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000708.

Pontaza:2007:SEL

[Pon07b] J. P. Pontaza. A spectral element least-squares formulation for incompressible Navier–Stokes flows using triangular nodal elements. Journal of Computational Physics, 221 (2):649–665, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910600310X.

Poncet:2009:AIB

[Pon09] Philippe Poncet. Analysis of an immersed boundary method for three-dimensional flows in vorticity formulation. Journal of Computational Physics, 228(19):7268–7288, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003465.

Popescu:2000:EAS

[Pop00] Lucretiu M. Popescu. An extension of alias sampling method for parametrized probability distributions. Journal of Computational Physics, 160(2):612–622, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964775.

Popinet:2003:GTB

[Pop03] Stéphane Popinet. Gerris: a tree-based adaptive solver for the incompressible Euler equations in complex geometries. Journal of Computational Physics, 190(2):572–600, September 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002985.

Popinet:2009:AAS

[Pop09] Stéphane Popinet. An accurate adaptive solver for surfacetension-driven interfacial flows. Journal of Computational Physics, 228(16):5838–5866, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900240X.

Potter:2000:LFF

[POS00] Michael E. Potter, Michal Okoniewski, and Maria A. Stuchly. Low frequency finite difference time domain (FDTD) for modeling of induced fields in humans close to line sources. Journal of Computational Physics, 162(1):82–103, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965239.

Pozrikidis:2001:IDS

[Poz01a] C. Pozrikidis. Interfacial dynamics for Stokes flow. Journal of Computational Physics, 169(2):250–301, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965823.

Pozrikidis:2001:NRD

[Poz01b] C. Pozrikidis. A note on the regularization of the discrete Poisson–Neumann problem. Journal of Computational Physics, 172(2):917–923, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968573.

Pilliod:2004:SOA

[PP04] James Edward Pilliod, Jr. and Elbridge Gerry Puckett. Second-order accurate volume-of-fluid algorithms for tracking material interfaces. Journal of Computational Physics, 199(2):465–502, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000920.

Parussini:2009:FDA

[PP09] Lucia Parussini and Valentino Pediroda. Fictitious Domain approach with hp-finite element approximation for incompressible fluid flow. Journal of Computational Physics, 228 (10):3891–3910, June 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001107.

Premnath:2009:SSC

[PPB09] Kannan N. Premnath, Martin J. Pattison, and Sanjoy Banerjee. Steady state convergence acceleration of the generalized lattice Boltzmann equation with forcing term through preconditioning. Journal of Computational Physics, 228(3):746–769, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005111.

Pascarelli:2000:MBL

[PPC00] Andrea Pascarelli, Ugo Piomelli, and Graham V. Candler. Multi-block large-eddy simulations of turbulent boundary layers. Journal of Computational Physics, 157(1):256–279, January 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199919996374X.

Peinetti:2006:PCM

[PPCW06] F. Peinetti, F. Peano, G. Coppa, and J. Wurtele. Particlein-cell method for parallel dynamics in magnetized electron plasmas: Study of high-amplitude BGK modes. Journal of Computational Physics, 218(1):102–122, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000623.

Pantano:2008:AHR

[PPDM08] C. Pantano, D. I. Pullin, P. E. Dimotakis, and G. Matheou. LES approach for high Reynolds number wall-bounded flows with application to turbulent channel flow. Journal of Computational Physics, 227(21):9271–9291, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002209.

Palmer:2000:LBA

[PR00] Bruce J. Palmer and David R. Rector. Lattice Boltzmann algorithm for simulating thermal flow in compressible fluids. Journal of Computational Physics, 161(1):1–20, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964258.

Park:2001:SMN

[PR01a] H. M. Park and D. H. Ryu. A solution method of nonlinear convective stability problems in finite domains. Journal of Computational Physics, 170(1):141–160, June 10, 2001.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967282.

Persson:2001:SWF

[PR01b] Per-Olof Persson and Olof Runborg. Simulation of a waveguide filter using wavelet-based numerical homogenization. Journal of Computational Physics, 166(2):361–382, January 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966622.

Pontaza:2003:SHL

[PR03] J. P. Pontaza and J. N. Reddy. Spectral/hp least-squares finite element formulation for the Navier–Stokes equations. Journal of Computational Physics, 190(2):523–549, September 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002961.

Pasquetti:2004:SEM

[PR04a] Richard Pasquetti and Francesca Rapetti. Spectral element methods on triangles and quadrilaterals: comparisons and applications. Journal of Computational Physics, 198(1):349–362, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104000324.

Pontaza:2004:STC

[PR04b] J. P. Pontaza and J. N. Reddy. Space–time coupled spectral/hp least-squares finite element formulation for the incompressible Navier–Stokes equations. Journal of Computational Physics, 197(2):418–459, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006466.

Prabhakar:2006:SHP

[PR06] V. Prabhakar and J. N. Reddy. Spectral/hp penalty least-squares finite element formulation for the steady incompressible Navier–Stokes equations. Journal of Computational Physics, 215(1):274–297, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004882.

Price:2008:MDK

[Pri08] Daniel J. Price. Modelling discontinuities and Kelvin– Helmholtz instabilities in SPH. Journal of Computational Physics, 227(24):10040–10057, December 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004270.

Pruett:2003:ABA

[PRL03] C. David Pruett, Joseph W. Rudmin, and Justin M. Lacy. An adaptive N-body algorithm of optimal order. Journal of Computational Physics, 187(1):298–317, May 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001013.

Probert:2003:IAG

[Pro03] M. I. J. Probert. Improved algorithm for geometry optimisation using damped molecular dynamics. Journal of Computational Physics, 191(1):130–146, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003085.

Prosser:2005:IBC

[Pro05] R. Prosser. Improved boundary conditions for the direct numerical simulation of turbulent subsonic flows. I. Inviscid flows. Journal of Computational Physics, 207(2):736–768, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000495.

Prosser:2007:TIB

[Pro07] R. Prosser. Towards improved boundary conditions for the DNS and LES of turbulent subsonic flows. Journal of Computational Physics, 222(2):469–474, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004438.

Protas:2008:ABOb

[Pro08] Bartosz Protas. Adjoint-based optimization of PDE systems with alternative gradients. Journal of Computational Physics, 227(13):6490–6510, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001575.

Pareschi:2000:FSM

[PRT00] L. Pareschi, G. Russo, and G. Toscani. Fast spectral methods for the Fokker–Planck–Landau collision operator. Journal of Computational Physics, 165(1):216–236, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966129.

Pillapakkam:2001:LSM

[PS01] S. B. Pillapakkam and P. Singh. A level-set method for computing solutions to viscoelastic two-phase flow. Journal of Computational Physics, 174(2):552–578, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910196927X.

Parent:2002:UDD

[PS02] Bernard Parent and Jean P. Sislian. The use of domain decomposition in accelerating the convergence of quasihyperbolic systems. Journal of Computational Physics, 179(1):140–169, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970488.

Park:2003:SIN

[PS03a] H. M. Park and H. J. Shin. Shape identification for natural convection problems using the adjoint variable method. Journal of Computational Physics, 186(1):198–211, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000469.

Prusa:2003:ASA

[PS03b] Joseph M. Prusa and Piotr K. Smolarkiewicz. An allscale anelastic model for geophysical flows: dynamic grid deformation. Journal of Computational Physics, 190(2):601–622, September 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002997.

Piller:2004:FVC

[PS04] M. Piller and E. Stalio. Finite-volume compact schemes on staggered grids. Journal of Computational Physics, 197(1): 299–340, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103006314.

Perigaud:2005:CFM

[PS05] Guillaume Perigaud and Richard Saurel. A compressible flow model with capillary effects. Journal of Computational Physics, 209(1):139–178, October 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001853.

Perot:2007:DCA

[PS07a] J. B. Perot and V. Subramanian. A discrete calculus analysis of the Keller box scheme and a generalization of the method to arbitrary meshes. Journal of Computational Physics, 226(1):494–508, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001696.

Perot:2007:DCM

[PS07b] J. B. Perot and V. Subramanian. Discrete calculus methods for diffusion. Journal of Computational Physics, 224(1):59–81, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106006127.

Pinnau:2007:MRT

[PS07c] René Pinnau and Alexander Schulze. Model reduction techniques for frequency averaging in radiative heat transfer. Journal of Computational Physics, 226(1):712–731, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002033.

Pyo:2007:GUM

[PS07d] Jae-Hong Pyo and Jie Shen. Gauge-Uzawa methods for incompressible flows with variable density. Journal of Computational Physics, 221(1):181–197, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002816.

Piller:2008:CFV

[PS08] M. Piller and E. Stalio. Compact finite volume schemes on boundary-fitted grids. Journal of Computational Physics, 227 (9):4736–4762, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910800048X.

Peng:2004:ITL

[PSC04] Y. Peng, C. Shu, and Y. T. Chew. A 3D incompressible thermal lattice Boltzmann model and its application to simulate natural convection in a cubic cavity. Journal of Computational Physics, 193(1):260–274, January 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004303.

Peng:2006:AMB

[PSC^{+}06] Y. Peng, C. Shu, Y. T. Chew, X. D. Niu, and X. Y. Lu. Application of multi-block approach in the immersed boundarylattice Boltzmann method for viscous fluid flows. Journal of Computational Physics, 218(2):460–478, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000933.

Pancheshnyi:2008:NSF

[PSCB08] S. Pancheshnyi, P. Ségur, J. Capeillère, and A. Bourdon. Numerical simulation of filamentary discharges with parallel adaptive mesh refinement. Journal of Computational Physics, 227(13):6574–6590, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001691.

Peng:2003:NIF

[PSCQ03] Y. Peng, C. Shu, Y. T. Chew, and J. Qiu. Numerical investigation of flows in Czochralski crystal growth by an axisymmetric lattice Boltzmann method. Journal of Computational Physics, 186(1):295–307, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000676.

Parand:2009:RLP

[PSD09] K. Parand, M. Shahini, and Mehdi Dehghan. Rational Legendre pseudospectral approach for solving nonlinear differential equations of Lane–Emden type. Journal of Computational Physics, 228(23):8830–8840, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004720.

Popescu:2005:FVT

[PSG05] Mihaela Popescu, Wei Shyy, and Marc Garbey. Finite volume treatment of dispersion-relation-preserving and optimized prefactored compact schemes for wave propagation. Journal of Computational Physics, 210(2):705–729, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105002585.

Petersen:2008:BTM

[PSH^{+}08] Dan Erik Petersen, Hans Henrik B. Sørensen, Per Christian Hansen, Stig Skelboe, and Kurt Stokbro. Block tridiagonal matrix inversion and fast transmission calculations. Journal of Computational Physics, 227(6):3174–3190, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005177.

Posta:2008:COG

[PSM08] F. Posta, S. Y. Shvartsman, and C. B. Muratov. Compensated optimal grids for elliptic boundary-value problems.

Journal of Computational Physics, 227(19):8622–8635, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108003331.

Piotrowski:2009:NRT

[PSMW09] Zbigniew P. Piotrowski, Piotr K. Smolarkiewicz, Szymon P. Malinowski, and Andrzej A. Wyszogrodzki. On numerical realizability of thermal convection. Journal of Computational Physics, 228(17):6268–6290, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002745.

Protas:2000:EAC

[PSN00] B. Protas, A. Styczek, and A. Nowakowski. An effective approach to computation of forces in viscous incompressible flows. Journal of Computational Physics, 159(2):231–245, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964337.

Pavliotis:2009:CED

[PSZ09] G. A. Pavliotis, A. M. Stuart, and K. C. Zygalakis. Calculating effective diffusivities in the limit of vanishing molecular diffusion. Journal of Computational Physics, 228(4):1030–1055, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005378.

Pudykiewicz:2006:NSR

[Pud06] Janusz A. Pudykiewicz. Numerical solution of the reactionadvection-diffusion equation on the sphere. Journal of Computational Physics, 213(1):358–390, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003840.

Pesch:2008:DGF

[PvdV08] L. Pesch and J. J. W. van der Vegt. A discontinuous Galerkin finite element discretization of the Euler equations for compressible and incompressible fluids. Journal of Computational Physics, 227(11):5426–5446, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000831.

Preis:2009:GAM

[PVPS09] Tobias Preis, Peter Virnau, Wolfgang Paul, and Johannes J. Schneider. GPU accelerated Monte Carlo simulation of the 2D and 3D Ising model. Journal of Computational Physics, 228(12):4468–4477, July 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001387.

Pollet:2007:ELO

[PVR07] Lode Pollet, Kris Van Houcke, and Stefan M. A. Rombouts. Engineering local optimality in quantum Monte Carlo algorithms. Journal of Computational Physics, 225(2):2249–2266, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107001295.

Pavarino:2000:OSM

[PW00a] Luca F. Pavarino and Timothy Warburton. Overlapping Schwarz methods for unstructured spectral elements. Journal of Computational Physics, 160(1):298–317, May 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964635.

Ploumhans:2000:VMH

[PW00b] P. Ploumhans and G. S. Winckelmans. Vortex methods for high-resolution simulations of viscous flow past bluff bodies of general geometry. Journal of Computational Physics, 165(2):354–406, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966142. See erratum [PW01].

Ploumhans:2001:EVM

[PW01] P. Ploumhans and G. S. Winckelmans. Erratum: “Vortex Methods for High-Resolution Simulations of Viscous Flow Past Bluff Bodies of General Geometry”: Volume 165, Number 2 (2000), pages 354–406. Journal of Computational Physics, 170(1):449–456, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968329. See [PW00b].

Paffrath:2007:APC

[PW07] M. Paffrath and U. Wever. Adapted polynomial chaos expansion for failure detection. Journal of Computational Physics, 226(1):263–281, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001568.

Polifke:2006:PRN

[PWM06] Wolfgang Polifke, Clifton Wall, and Parviz Moin. Partially reflecting and non-reflecting boundary conditions for simulation of compressible viscous flow. Journal of Computational Physics, 213(1):437–449, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003876.

Ploumhans:2002:VMD

[PWS^{+}02] P. Ploumhans, G. S. Winckelmans, J. K. Salmon, A. Leonard, and M. S. Warren. Vortex methods for direct numerical simulation of three-dimensional bluff body flows: Application to the sphere at Re = 300, 500, and 1000. Journal of Computational Physics, 178(2):427–463, May 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910297035X.

Plaskota:2000:NAW

[PWW00] Leszek Plaskota, Grzegorz W. Wasilkowski, and Henryk Woźniakowski. A new algorithm and worst case complexity for Feynman–Kac path integration. Journal of Computational Physics, 164(2):335–353, November 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965999.

Pasquetti:2002:CFB

[PX02] R. Pasquetti and C. J. Xu. Comments on “Filter-Based Stabilization of Spectral Element Methods”. Journal of Computational Physics, 182(2):646–650, November 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971780. See P. Fischer and J.

Mullen, C. R. Acad. Sci. Paris, 332 (2001), p. 265.

Park:2004:DEL

[PYC04] Noma Park, Jung Yul Yoo, and Haecheon Choi. Discretization errors in large eddy simulation: on the suitability of centered and upwind-biased compact difference schemes. Journal of Computational Physics, 198(2):580–616, August 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104000531.

Qamar:2009:HOK

[QA09] Shamsul Qamar and Munshoor Ahmed. A high order kinetic flux-vector splitting method for the reduced five-equation model of compressible two-fluid flows. Journal of Computational Physics, 228(24):9059–9078, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004951.

Quartapelle:2003:SRP

[QCGQ03] L. Quartapelle, L. Castelletti, A. Guardone, and G. Quaranta. Solution of the Riemann problem of classical gasdynamics. Journal of Computational Physics, 190(1):118–140, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103002675.

Qiang:2004:PPC

[QFR04] Ji Qiang, Miguel A. Furman, and Robert D. Ryne. A parallel particle-in-cell model for beam-beam interaction in high energy ring colliders. Journal of Computational Physics, 198(1):278–294, July 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000282.

Qiu:2006:NSP

[QKS06] Jianxian Qiu, Boo Cheong Khoo, and Chi-Wang Shu. A numerical study for the performance of the Runge–Kutta discontinuous Galerkin method based on different numerical fluxes. Journal of Computational Physics, 212(2):540–565, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003451.

Qin:2001:MMM

[QL01] Zhaohui S. Qin and Jun S. Liu. Multipoint Metropolis method with application to hybrid Monte Carlo. Journal of Computational Physics, 172(2):827–840, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968603.

Qian:2004:LSB

[QL04] Jianliang Qian and Shingyu Leung. A level set based Eulerian method for paraxial multivalued traveltimes. Journal of Computational Physics, 197(2):711–736, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000154.

Qiu:2007:RKD

[QLK07] Jianxian Qiu, Tiegang Liu, and Boo Cheong Khoo. Runge– Kutta discontinuous Galerkin methods for compressible twomedium flow simulations: One-dimensional case. Journal of Computational Physics, 222(1):353–373, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003561.

Quan:2009:MMB

[QLS09] Shaoping Quan, Jing Lou, and David P. Schmidt. Modeling merging and breakup in the moving mesh interface tracking method for multiphase flow simulations. Journal of Computational Physics, 228(7):2660–2675, April 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800675X.

Quispel:2003:EVP

[QM03] G. R. W. Quispel and D. I. McLaren. Explicit volumepreserving and symplectic integrators for trigonometric polynomial flows. Journal of Computational Physics, 186(1): 308–316, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103000688.

Quecedo:2003:FEM

[QP03] M. Quecedo and M. Pastor. Finite element modelling of free surface flows on inclined and curved beds. Journal of Computational Physics, 189(1):45–62, July 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002006.

Qiang:2000:OOP

[QRHD00] Ji Qiang, Robert D. Ryne, Salman Habib, and Viktor Decyk. An object-oriented parallel particle-in-cell code for beam dynamics simulation in linear accelerators. Journal of Computational Physics, 163(2):434–451, September 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965707.

Qian:2001:PES

[QS01] Jianliang Qian and William W. Symes. Paraxial eikonal solvers for anisotropic quasi-P travel times. Journal of Computational Physics, 173(1):256–278, October 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968755.

Qiu:2002:CCL

[QS02] Jianxian Qiu and Chi-Wang Shu. On the construction, comparison, and local characteristic decomposition for highorder central WENO schemes. Journal of Computational Physics, 183(1):187–209, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971913.

Qiu:2004:HWS

[QS04] Jianxian Qiu and Chi-Wang Shu. Hermite WENO schemes and their application as limiters for Runge–Kutta discontinuous Galerkin method: one-dimensional case. Journal of Computational Physics, 193(1):115–135, January 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004212.

Qiu:2005:HWS

[QS05] Jianxian Qiu and Chi-Wang Shu. Hermite WENO schemes for Hamilton–Jacobi equations. Journal of Computational Physics, 204(1):82–99, March 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004085.

Quan:2007:MMI

[QS07] Shaoping Quan and David P. Schmidt. A moving mesh interface tracking method for 3D incompressible two-phase flows. Journal of Computational Physics, 221(2):761–780, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003159.

Qasimov:2008:LBS

[QT08] H. Qasimov and S. Tsynkov. Lacunae based stabilization of PMLs. Journal of Computational Physics, 227(15):7322–7345, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108002301.

Qian:2001:VBM

[QV01] L. Qian and M. Vezza. A vorticity-based method for incompressible unsteady viscous flows. Journal of Computational Physics, 172(2):515–542, September 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968354.

Qamar:2005:HOK

[QW05] Shamsul Qamar and Gerald Warnecke. A high-order kinetic flux-splitting method for the relativistic magnetohydrodynamics. Journal of Computational Physics, 205(1):182–204, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004632.

Rendall:2009:EMM

[RA09] T. C. S. Rendall and C. B. Allen. Efficient mesh motion using radial basis functions with data reduction algorithms. Journal of Computational Physics, 228(17):6231–6249, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002721.

Ramiere:2007:GFD

[RAB07] Isabelle Ramière, Philippe Angot, and Michel Belliard. A general fictitious domain method with immersed jumps and multilevel nested structured meshes. Journal of Computational Physics, 225(2):1347–1387, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000460.

Ricchiuto:2007:ACR

[RAD07] M. Ricchiuto, R. Abgrall, and H. Deconinck. Application of conservative residual distribution schemes to the solution of the shallow water equations on unstructured meshes. Journal of Computational Physics, 222(1):287–331, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002853.

Rahmouni:2004:AMD

[Rah04] Adib N. Rahmouni. An algebraic method to develop wellposed PML models: Absorbing layers, perfectly matched layers, linearized Euler equations. Journal of Computational Physics, 197(1):99–115, June 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103006211.

Ramos:2003:NAU

[Ram03] Rubens Viana Ramos. Numerical algorithms for use in quantum information. Journal of Computational Physics, 192 (1):95–104, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003486.

Ramirez:2006:MCA

[Ram06] Jorge M. Ramirez. Multiplicative cascades applied to PDEs (two numerical examples). Journal of Computational Physics, 214(1):122–136, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004262.

Rylander:2002:SEI

[RB02] Thomas Rylander and Anders Bondeson. Stability of explicitimplicit hybrid time-stepping schemes for Maxwell’s equations. Journal of Computational Physics, 179(2):426–438, July 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970634.

Raad:2005:TDE

[RB05] Peter E. Raad and Razvan Bidoae. The three-dimensional Eulerian–Lagrangian marker and micro cell method for the simulation of free surface flows. Journal of Computational Physics, 203(2):668–699, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003985.

Rahul:2006:OSF

[RB06] Kumar Rahul and S. N. Bhattacharyya. One-sided finitedifference approximations suitable for use with Richardson extrapolation. Journal of Computational Physics, 219(1):13– 20, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910600266X.

Ricchiuto:2009:SRD

[RB09a] Mario Ricchiuto and Andreas Bollermann. Stabilized residual distribution for shallow water simulations. Journal of Computational Physics, 228(4):1071–1115, March 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005391.

Rieper:2009:ICG

[RB09b] Felix Rieper and Georg Bader. The influence of cell geometry on the accuracy of upwind schemes in the low Mach number regime. Journal of Computational Physics, 228(8):2918–2933, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109000096.

Remaki:2003:IAI

[RBH03] L. Remaki, H. Beaugendre, and W. G. Habashi. ISOD — an anisotropic isovalue-oriented diffusion artificial viscosity for the Euler and Navier–Stokes equations. Journal of Computational Physics, 186(1):279–294, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000664.

Ruf:2009:RSS

[RBK09] Matthias Ruf, Heiko Bauke, and Christoph H. Keitel. A real space split operator method for the Klein–Gordon equation. Journal of Computational Physics, 228(24):9092–9106, December 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004975.

Rossmanith:2004:WPA

[RBL04] James A. Rossmanith, Derek S. Bale, and Randall J. LeVeque. A wave propagation algorithm for hyperbolic systems on curved manifolds. Journal of Computational Physics, 199(2):631–662, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400110X.

Restelli:2006:SLD

[RBS06] M. Restelli, L. Bonaventura, and R. Sacco. A semi-Lagrangian discontinuous Galerkin method for scalar advection by incompressible flows. Journal of Computational Physics, 216(1):195–215, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005486.

Ramboer:2006:OTI

[RBSL06] Jan Ramboer, Tim Broeckhoven, Sergey Smirnov, and Chris Lacor. Optimization of time integration schemes coupled to spatial discretization for use in CAA applications. Journal of Computational Physics, 213(2):777–802, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500416X.

Rogers:2003:MBF

[RBT03] Benedict D. Rogers, Alistair G. L. Borthwick, and Paul H. Taylor. Mathematical balancing of flux gradient and source terms prior to using Roe’s approximate Riemann solver. Journal of Computational Physics, 192(2):422–451, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004029.

Rhebergen:2008:DGF

[RBvdV08] S. Rhebergen, O. Bokhove, and J. J. W. van der Vegt. Discontinuous Galerkin finite element methods for hyperbolic nonconservative partial differential equations. Journal of Computational Physics, 227(3):1887–1922, January 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107004391.

Rowley:2000:DNB

[RC00] Clarence W. Rowley and Tim Colonius. Discretely nonreflecting boundary conditions for linear hyperbolic systems. Journal of Computational Physics, 157(2):500–538, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963830.

Rochette:2006:TDC

[RC06] D. Rochette and S. Clain. Two-dimensional computation of gas flow in a porous bed characterized by a porosity jump. Journal of Computational Physics, 219(1):104–119, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106001422.

Raga:2009:FFM

[RC09a] A. C. Raga and J. Cantó. A family of functions for mass and energy flux splitting of the Euler equations. Journal of Computational Physics, 228(23):8908–8918, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004914.

Ryan:2009:LDP

[RC09b] Jennifer K. Ryan and Bernardo Cockburn. Local derivative post-processing for the discontinuous Galerkin method. Journal of Computational Physics, 228(23):8642–8664, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004616.

Rosatti:2005:SIS

[RCB05] G. Rosatti, D. Cesari, and L. Bonaventura. Semi-implicit, semi-Lagrangian modelling for environmental problems on staggered Cartesian grids with cut cells. Journal of Computational Physics, 204(1):353–377, March 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004206.

Ricchiuto:2005:RDG

[RCD05] Mario Ricchiuto, Árpád Csík, and Herman Deconinck. Residual distribution for general time-dependent conservation laws. Journal of Computational Physics, 209(1):249–289, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910500118X.

Rangan:2007:NMS

[RCT07] Aaditya V. Rangan, David Cai, and Louis Tao. Numerical methods for solving moment equations in kinetic theory of neuronal network dynamics. Journal of Computational Physics, 221(2):781–798, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003160.

Rossiello:2007:TOA

[RDPN07] G. Rossiello, P. De Palma, G. Pascazio, and M. Napolitano. Third-order-accurate fluctuation splitting schemes for unsteady hyperbolic problems. Journal of Computational Physics, 222(1):332–352, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003627.

Ren:2005:HMM

[RE05] Weiqing Ren and Weinan E. Heterogeneous multiscale method for the modeling of complex fluids and micro-fluidics. Journal of Computational Physics, 204(1):1–26, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004048.

Rathinam:2007:REM

[RE07] Muruhan Rathinam and Hana El Samad. Reversibleequivalent-monomolecular tau: a leaping method for “small number and stiff” stochastic chemical systems. Journal of Computational Physics, 224(2):897–923, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005547.

Reich:2000:MSR

[Rei00] Sebastian Reich. Multi-symplectic Runge–Kutta collocation methods for Hamiltonian wave equations. Journal of Computational Physics, 157(2):473–499, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963726.

Remis:2000:SFD

[Rem00] Rob F. Remis. On the stability of the finite-difference timedomain method. Journal of Computational Physics, 163 (1):249–261, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100965732.

Remis:2006:SFN

[Rem06] Rob F. Remis. Stability of FDTD on nonuniform grids for Maxwell’s equations in lossless media. Journal of Computational Physics, 218(2):594–606, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001227.

Ren:2007:ANS

[Ren07] Weiqing Ren. Analytical and numerical study of coupled atomistic-continuum methods for fluids. Journal of Computational Physics, 227(2):1353–1371, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700397X.

Rosatti:2006:WBA

[RF06] G. Rosatti and L. Fraccarollo. A well-balanced approach for flows over mobile-bed with high sediment-transport. Journal of Computational Physics, 220(1):312–338, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002439.

Rosenberg:2006:GAS

[RFFP06] Duane Rosenberg, Aimé Fournier, Paul Fischer, and Annick Pouquet. Geophysical-astrophysical spectral-element adaptive refinement (GASpAR): Object-oriented h-adaptive fluid dynamics simulation. Journal of Computational Physics, 215(1):59–80, June 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004791.

Rouizi:2009:NMR

[RFVP09] Y. Rouizi, Y. Favennec, J. Ventura, and D. Petit. Numerical model reduction of 2D steady incompressible laminar flows: Application on the flow over a backward-facing step. Journal of Computational Physics, 228(6):2239–2255, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800630X.

Rider:2007:AME

[RGK07] William J. Rider, Jeffrey A. Greenough, and James R. Kamm. Accurate monotonicity- and extrema-preserving methods through adaptive nonlinear hybridizations. Journal of Computational Physics, 225(2):1827–1848, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000897.

Reis:2004:NSI

[RGS04] Neyval C. Reis, Jr., Richard F. Griffiths, and Jane M. Santos. Numerical simulation of the impact of liquid droplets on porous surfaces. Journal of Computational Physics, 198(2):747–770, August 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000610.

Rasetarinera:2001:EID

[RH01a] Patrick Rasetarinera and M. Y. Hussaini. An efficient implicit discontinuous spectral Galerkin method. Journal of Computational Physics, 172(2):718–738, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968536.

Rivoalen:2001:PSE

[RH01b] E. Rivoalen and S. Huberson. The particle strength exchange method applied to axisymmetric viscous flows. Journal of Computational Physics, 168(2):519–526, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967129.

Ruuth:2005:HOL

[RH05] Steven J. Ruuth and Willem Hundsdorfer. High-order linear multistep methods with general monotonicity and boundedness properties. Journal of Computational Physics, 209 (1):226–248, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105000951.

Reynolds:2009:SCS

[RHPN09] Daniel R. Reynolds, John C. Hayes, Pascal Paschos, and Michael L. Norman. Self-consistent solution of cosmological radiation-hydrodynamics and chemical ionization. Journal of Computational Physics, 228(18):6833–6854, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003258.

Rider:2000:RWH

[Rid00] William J. Rider. Revisiting Wall heating. Journal of Computational Physics, 162(2):395–410, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965446.

Rylander:2004:PML

[RJ04] Thomas Rylander and Jian-Ming Jin. Perfectly matched layer for the time domain finite element method. Journal of Computational Physics, 200(1):238–250, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001652.

Rembold:2006:MJP

[RJ06] B. Rembold and P. Jenny. A multiblock joint PDF finitevolume hybrid algorithm for the computation of turbulent flows in complex geometries. Journal of Computational Physics, 220(1):59–87, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002221.

Rosam:2007:FIF

[RJM07] J. Rosam, P. K. Jimack, and A. Mullis. A fully implicit, fully adaptive time and space discretisation method for phase-field simulation of binary alloy solidification. Journal of Computational Physics, 225(2):1271–1287, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000393.

Rospsha:2007:CFF

[RK07] Nimrod Rospsha and Raphael Kastner. Closed form FDTDcompatible Green’s function based on combinatorics. Journal of Computational Physics, 226(1):798–817, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002070.

Riyanti:2007:PMB

[RKE^{+}07] C. D. Riyanti, A. Kononov, Y. A. Erlangga, C. Vuik, C. W. Oosterlee, R.-E. Plessix, and W. A. Mulder. A parallel multigrid-based preconditioner for the 3D heterogeneous high-frequency Helmholtz equation. Journal of Computational Physics, 224(1):431–448, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001477.

Ricci:2002:SIM

[RLB02] Paolo Ricci, Giovanni Lapenta, and J. U. Brackbill. A simplified implicit Maxwell solver. Journal of Computational Physics, 183(1):117–141, November 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971706.

Ren:2003:CWH

[RLZ03] Yu-Xin Ren, Miao’er Liu, and Hanxin Zhang. A characteristicwise hybrid compact-WENO scheme for solving hyperbolic conservation laws. Journal of Computational Physics, 192 (2):365–386, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910300398X.

Rider:2001:SMM

[RM01a] William J. Rider and Len G. Margolin. Simple modifications of monotonicity-preserving limiter. Journal of Computational Physics, 174(1):473–488, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969141.

Ruuth:2001:CTM

[RM01b] Steven J. Ruuth and Barry Merriman. Convolutionthresholding methods for interface motion. Journal of Computational Physics, 169(2):678–707, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910096580X.

Rupert:2007:APC

[RM07] C. P. Rupert and C. T. Miller. An analysis of polynomial chaos approximations for modeling single-fluid-phase flow in porous medium systems. Journal of Computational Physics, 226(2):2175–2205, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700304X.

Ruuth:2008:SEM

[RM08] Steven J. Ruuth and Barry Merriman. A simple embedding method for solving partial differential equations on surfaces. Journal of Computational Physics, 227(3):1943–1961, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700441X.

Raessi:2007:ANV

[RMB07] M. Raessi, J. Mostaghimi, and M. Bussmann. Advecting normal vectors: a new method for calculating interface normals and curvatures when modeling two-phase flows. Journal of Computational Physics, 226(1):774–797, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002069.

Rosatti:2008:GRS

[RMF08] G. Rosatti, J. Murillo, and L. Fraccarollo. Generalized Roe schemes for 1D two-phase, free-surface flows over a mobile bed. Journal of Computational Physics, 227 (24):10058–10077, December 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004282.

Riber:2009:ENS

[RMG^{+}09] E. Riber, V. Moureau, M. García, T. Poinsot, and O. Simonin. Evaluation of numerical strategies for large eddy simulation of particulate two-phase recirculating flows. Journal of Computational Physics, 228(2):539–564, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005238.

Rico-Martinez:2004:CPK

[RMGK04] R. Rico-Martínez, C. W. Gear, and Ioannis G. Kevrekidis. Coarse projective kMC integration: forward/reverse initial and boundary value problems. Journal of Computational Physics, 196(2):474–489, May 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006053.

Ruuth:2000:FGM

[RMO00] Steven J. Ruuth, Barry Merriman, and Stanley Osher. A fixed grid method for capturing the motion of selfintersecting wavefronts and related PDEs. Journal of Computational Physics, 163(1):1–21, September 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964805.

Ricot:2009:LBM

[RMSB09] Denis Ricot, Simon Marié, Pierre Sagaut, and Christophe Bailly. Lattice Boltzmann method with selective viscosity filter. Journal of Computational Physics, 228(12):4478–4490, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001399.

Rylander:2003:ERF

[RMV03] Thomas Rylander, Tomas McKelvey, and Mats Viberg. Estimation of resonant frequencies and quality factors from time domain computations. Journal of Computational Physics, 192(2):523–545, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004078.

Romano:2002:SSM

[Rom02] Vittorio Romano. 2D simulation of a silicon MESFET with a nonparabolic hydrodynamical model based on the maximum entropy principle. Journal of Computational Physics, 176(1):70–92, February 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969682.

Romano:2007:NSM

[Rom07] V. Romano. 2D numerical simulation of the MEP energytransport model with a finite difference scheme. Journal of Computational Physics, 221(2):439–468, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600297X.

Rossow:2000:FSS

[Ros00] Cord-Christian Rossow. A flux-splitting scheme for compressible and incompressible flows. Journal of Computational Physics, 164(1):104–122, October 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965860.

Rossow:2003:BPD

[Ros03] C.-C. Rossow. A blended pressure/density based method for the computation of incompressible and compressible flows. Journal of Computational Physics, 185(2):375–398, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000591.

Rossmanith:2006:WPM

[Ros06] James A. Rossmanith. A wave propagation method for hyperbolic systems on the sphere. Journal of Computational Physics, 213(2):629–658, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003967.

Rossow:2007:ECC

[Ros07] Cord-Christian Rossow. Efficient computation of compressible and incompressible flows. Journal of Computational Physics, 220(2):879–899, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002658.

Rossmanith:2008:CRD

[Ros08] James A. Rossmanith. A class of residual distribution schemes and their relation to relaxation systems. Journal of Computational Physics, 227(22):9527–9553, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003665.

Rossi:2009:DNS

[Ros09] Riccardo Rossi. Direct numerical simulation of scalar transport using unstructured finite-volume schemes. Journal of Computational Physics, 228(5):1639–1657, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005871.

Roy:2005:RCS

[Roy05] Christopher J. Roy. Review of code and solution verification procedures for computational simulation. Journal of Computational Physics, 205(1):131–156, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004619.

Ren:2008:SOS

[RP08a] Zhuyin Ren and Stephen B. Pope. Second-order splitting schemes for a class of reactive systems. Journal of Computational Physics, 227(17):8165–8176, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108003057.

Rosakis:2008:D

[RP08b] Ares Rosakis and Dale Pullin. Dedication. Journal of Computational Physics, 227(21):9006–9007, November 10, 2008.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004269.

Renardy:2002:PPR

[RR02] Yuriko Renardy and Michael Renardy. PROST: a parabolic reconstruction of surface tension for the volume-of-fluid method. Journal of Computational Physics, 183(2):400– 421, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971901.

Ramlau:2007:MSL

[RR07] Ronny Ramlau and Wolfgang Ring. A Mumford–Shah level-set approach for the inversion and segmentation of X-ray tomography data. Journal of Computational Physics, 221(2):539–557, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003020.

Recine:2005:NST

[RRC05] Greg Recine, Bernard Rosen, and Hong-Liang Cui. Numerical simulation of two-dimensional electron transport in cylindrical nanostructures using Wigner function methods. Journal of Computational Physics, 209(2):421–447, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001361.

Renardy:2001:NSM

[RRL01] Michael Renardy, Yuriko Renardy, and Jie Li. Numerical simulation of moving contact line problems using a volume-of-fluid method. Journal of Computational Physics, 171(1):243–263, July 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967853.

Reese:2001:MCA

[RRV01] J. S. Reese, S. Raimondeau, and D. G. Vlachos. Monte Carlo algorithms for complex surface reaction mechanisms: Efficiency and accuracy. Journal of Computational Physics, 173(1):302–321, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968779.

Ratz:2006:SEE

[RRV06] Andreas Rätz, Angel Ribalta, and Axel Voigt. Surface evolution of elastically stressed films under deposition by a diffuse interface model. Journal of Computational Physics, 214(1):187–208, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004304.

Rieben:2005:HOM

[RRW05] R. N. Rieben, G. H. Rodrigue, and D. A. White. A high order mixed vector finite element method for solving the time dependent Maxwell equations on unstructured grids. Journal of Computational Physics, 204(2):490–519, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004255.

Russo:2000:RCD

[RS00] Giovanni Russo and Peter Smereka. A remark on computing distance functions. Journal of Computational Physics, 163(1):51–67, September 1, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965537.

Ramos:2002:CQE

[RS02] Rubens Viana Ramos and Rui Fragassi Souza. Calculation of the quantum entanglement measure of bipartite states, based on relative entropy, using genetic algorithms. Journal of Computational Physics, 175(2):576–583, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969578.

Ropp:2005:SOS

[RS05] David L. Ropp and John N. Shadid. Stability of operator splitting methods for systems with indefinite operators: reaction-diffusion systems. Journal of Computational Physics, 203(2):449–466, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003870.

Ren:2006:MDU

[RS06a] Yu-Xin Ren and Yutao Sun. A multi-dimensional upwind scheme for solving Euler and Navier–Stokes equations. Journal of Computational Physics, 219(1):391–403, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001689.

Russo:2006:CSE

[RS06b] Giovanni Russo and Peter Smereka. Computation of strained epitaxial growth in three dimensions by kinetic Monte Carlo. Journal of Computational Physics, 214(2):809–828, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004742.

Ropp:2009:SOS

[RS09a] David L. Ropp and John N. Shadid. Stability of operator splitting methods for systems with indefinite operators: Advection-diffusion-reaction systems. Journal of Computational Physics, 228(9):3508–3516, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000679.

Roy:2009:GES

[RS09b] Christopher J. Roy and Andrew J. Sinclair. On the generation of exact solutions for evaluating numerical schemes and estimating discretization error. Journal of Computational Physics, 228(5):1790–1802, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005949.

Rasin:2005:MRL

[RSM05] Igor Rasin, Sauro Succi, and Wolfram Miller. A multirelaxation lattice kinetic method for passive scalar diffusion.

Journal of Computational Physics, 206(2):453–462, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104005224.

Ropp:2004:SAT

[RSO04] David L. Ropp, John N. Shadid, and Curtis C. Ober. Studies of the accuracy of time integration methods for reaction-diffusion equations. Journal of Computational Physics, 194(2):544–574, March 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004947.

Raza:2009:SGA

[RSS09] Nauman Raza, Sultan Sial, and Shahid S. Siddiqi. Sobolev gradient approach for the time evolution related to energy minimization of Ginzburg–Landau functionals. Journal of Computational Physics, 228(7):2566–2571, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006529.

Raza:2009:EMR

[RSSL09] Nauman Raza, Sultan Sial, Shahid S. Siddiqi, and Turab Lookman. Energy minimization related to the nonlinear Schrödinger equation. Journal of Computational Physics, 228(7):2572–2577, April 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006530.

Roussel:2003:CFA

[RSTB03] Olivier Roussel, Kai Schneider, Alexei Tsigulin, and Henning Bockhorn. A conservative fully adaptive multiresolution algorithm for parabolic PDEs. Journal of Computational Physics, 188(2):493–523, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300189X.

Reynolds:2006:FIN

[RSW06] Daniel R. Reynolds, Ravi Samtaney, and Carol S. Woodward. A fully implicit numerical method for single-fluid resistive magnetohydrodynamics. Journal of Computational Physics, 219(1):144–162, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001446.

Ryabenkii:2001:GDA

[RTT01] V. S. Ryaben’kii, S. V. Tsynkov, and V. I. Turchaninov. Global discrete artificial boundary conditions for time-dependent wave propagation. Journal of Computational Physics, 174(2):712–758, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969360.

Raimondeau:2000:LDA

[RV00] S. Raimondeau and D. G. Vlachos. Low-dimensional approximations of multiscale epitaxial growth models for microstructure control of materials. Journal of Computational Physics, 160(2):564–576, May 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964738.

Rus:2007:SSR

[RV07] Francisco Rus and Francisco R. Villatoro. Self-similar radiation from numerical Rosenau–Hyman compactons. Journal of Computational Physics, 227(1):440–454, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003415.

Rus:2009:APC

[RV09] Francisco Rus and Francisco R. Villatoro. Adiabatic perturbations for compactons under dissipation and numericallyinduced dissipation. Journal of Computational Physics, 228 (11):4291–4302, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001314.

Rauwoens:2009:CDC

[RVDM09] Pieter Rauwoens, Jan Vierendeels, Erik Dick, and Bart Merci. A conservative discrete compatibility-constraint low-Mach pressure-correction algorithm for time-accurate simulations of variable density flows. Journal of Computational Physics, 228(13):4714–4744, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001466.

Rauwoens:2007:SOE

[RVM07] Pieter Rauwoens, Jan Vierendeels, and Bart Merci. A solution for the odd-even decoupling problem in pressure-correction algorithms for variable density flows. Journal of Computational Physics, 227(1):79–99, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700321X.

Rubinacci:2009:FTA

[RVVL09] Guglielmo Rubinacci, Salvatore Ventre, Fabio Villone, and Yueqiang Liu. A fast technique applied to the analysis of resistive Wall modes with 3D conducting structures. Journal of Computational Physics, 228(5):1562–1572, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005822.

Ren:2000:IGR

[RW00] Weiqing Ren and Xiao-Ping Wang. An iterative grid redistribution method for singular problems in multiple dimensions. Journal of Computational Physics, 159(2):246–273, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964350.

Russell:2003:CGM

[RW03] David Russell and Z. Jane Wang. A Cartesian grid method for modeling multiple moving objects in 2D incompressible viscous flow. Journal of Computational Physics, 191(1):177–205, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003103.

Ripoll:2008:DMC

[RW08] J.-F. Ripoll and A. A. Wray. A 3-D multiband closure for radiation and neutron transfer moment models. Journal of Computational Physics, 227(4):2212–2237, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003853.

Reisner:2003:EPB

[RWMK03] Jon Reisner, Andrzej Wyszogrodzki, Vincent Mousseau, and Dana Knoll. An efficient physics-based preconditioner for the fully implicit solution of small-scale thermally driven atmospheric flows. Journal of Computational Physics, 189(1):30– 44, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103001980.

Rieben:2007:ALE

[RWWS07] R. N. Rieben, D. A. White, B. K. Wallin, and J. M. Solberg. An arbitrary Lagrangian–Eulerian discretization of MHD on 3D unstructured grids. Journal of Computational Physics, 226(1):534–570, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001714.

Rognlien:2002:API

[RXH02] T. D. Rognlien, X. Q. Xu, and A. C. Hindmarsh. Application of parallel implicit methods to edge-plasma numerical simulations. Journal of Computational Physics, 175(1):249–268, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910196944X.

Ryckelynck:2005:PHM

[Ryc05] D. Ryckelynck. A priori hyperreduction method: an adaptive approach. Journal of Computational Physics, 202(1):346–366, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910400289X.

Shim:2006:HAA

[SA06] Yunsic Shim and Jacques G. Amar. Hybrid asynchronous algorithm for parallel kinetic Monte Carlo simulations of thin film growth. Journal of Computational Physics, 212 (1):305–317, February 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105003220.

Sircombe:2009:VSC

[SA09] N. J. Sircombe and T. D. Arber. VALIS: a split-conservative scheme for the relativistic 2D Vlasov–Maxwell system. Journal of Computational Physics, 228(13):4773–4788, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001636.

Sacchetti:2007:SSM

[Sac07] Andrea Sacchetti. Spectral splitting method for nonlinear Schrödinger equations with singular potential. Journal of Computational Physics, 227(2):1483–1499, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004196.

Safouhi:2000:HHM

[Saf00] Hassan Safouhi. The HD and HD^{−} methods for accelerating the convergence of three-center nuclear attraction and four-center two-electron Coulomb integrals over B functions and their convergence properties. Journal of Computational Physics, 165(2):473–495, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096621X.

Safouhi:2002:ERN

[Saf02] Hassan Safouhi. Efficient and rapid numerical evaluation of the two-electron, four-center Coulomb integrals using nonlinear transformations and useful properties of sine and Bessel functions. Journal of Computational Physics, 176(1):1–19, February 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969256.

Saito:2002:NAD

[Sai02] T. Saito. Numerical analysis of dusty-gas flows. Journal of Computational Physics, 176(1):129–144, February 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969712.

Schlatter:2005:WMP

[SAK05] P. Schlatter, N. A. Adams, and L. Kleiser. A windowing method for periodic inflow/outflow boundary treatment of non-periodic flows. Journal of Computational Physics, 206(2):505–535, July 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000021.

Shin:2005:ARS

[SAKDJ05] Seungwon Shin, S. I. Abdel-Khalik, Virginie Daru, and Damir Juric. Accurate representation of surface tension using the level contour reconstruction method. Journal of Computational Physics, 203(2):493–516, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003894.

Shirani:2005:IPC

[SAM05] E. Shirani, N. Ashgriz, and J. Mostaghimi. Interface pressure calculation based on conservation of momentum for front capturing methods. Journal of Computational Physics, 203(1):154–175, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104003225.

Samtaney:2009:MSL

[Sam09] Ravi Samtaney. A method to simulate linear stability of impulsively accelerated density interfaces in ideal-MHD and gas dynamics. Journal of Computational Physics, 228(18):6773–6783, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003118.

Sandu:2001:PNI

[San01] Adrian Sandu. Positive numerical integration methods for chemical kinetic systems. Journal of Computational Physics, 170(2):589–602, July 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967506.

SanMiguel:2003:SAT

[San03] A. San Miguel. Stabilization and attitude of a triaxial rigid body by Lie group methods. Journal of Computational Physics, 191(1):75–96, October 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003061.

Sarra:2003:CSS

[Sar03] Scott A. Sarra. Chebyshev super spectral viscosity method for a fluidized bed model. Journal of Computational Physics, 186(2):630–651, April 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000895.

Saut:2004:CMU

[Sau04] Olivier Saut. Computational modeling of ultrashort powerful laser pulses in a nonlinear crystal. Journal of Computational Physics, 197(2):624–646, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006557.

Sun:2002:DSM

[SB02] Quanhua Sun and Iain D. Boyd. A direct simulation method for subsonic, microscale gas flows. Journal of Computational Physics, 179(2):400–425, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970610.

Sridar:2003:UFD

[SB03] D. Sridar and N. Balakrishnan. An upwind finite difference scheme for meshless solvers. Journal of Computational Physics, 189(1):1–29, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001979.

Safouhi:2006:FTM

[SB06a] Hassan Safouhi and Lilian Berlu. The Fourier transform method and the SD̄ approach for the analytical and numerical treatment of multicenter overlap-like quantum similarity integrals. Journal of Computational Physics, 216(1):19–36, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005401.

Schwartzentruber:2006:HPC

[SB06b] T. E. Schwartzentruber and I. D. Boyd. A hybrid particlecontinuum method applied to shock waves. Journal of Computational Physics, 215(2):402–416, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004936.

Sert:2006:SEF

[SB06c] Cuneyt Sert and Ali Beskok. Spectral element formulations on non-conforming grids: a comparative study of pointwise matching and integral projection methods. Journal of Computational Physics, 211(1):300–325, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002767.

Sun:2007:SIT

[SB07] Y. Sun and C. Beckermann. Sharp interface tracking using the phase-field equation. Journal of Computational Physics, 220(2):626–653, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002531.

Sharipov:2009:NSL

[SB09] Felix Sharipov and Guilherme Bertoldo. Numerical solution of the linearized Boltzmann equation for an arbitrary intermolecular potential. Journal of Computational Physics, 228(9):3345–3357, May 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000394.

Sahmim:2007:SMB

[SBA07] Slah Sahmim, Fayssal Benkhaldoun, and Francisco Alcrudo. A sign matrix based scheme for non-homogeneous PDE’s with an analysis of the convergence stagnation phenomenon. Journal of Computational Physics, 226(2):1753–1783, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002604.

Sun:2004:HCP

[SBC04] Quanhua Sun, Iain D. Boyd, and Graham V. Candler. A hybrid continuum/particle approach for modeling subsonic, rarefied gas flows. Journal of Computational Physics, 194(1):256–277, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004698.

Schwartz:2006:CGE

[SBCL06] Peter Schwartz, Michael Barad, Phillip Colella, and Terry Ligocki. A Cartesian grid embedded boundary method for the heat equation and Poisson’s equation in three dimensions. Journal of Computational Physics, 211(2):531–550, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910500286X.

Schneider:2000:EFV

[SBGK00] T. Schneider, N. Botta, K. J. Geratz, and R. Klein. Extension of finite volume compressible flow solvers to multidimensional, variable density zero Mach number flows: Volume 155, number 2 (1999), pages 248–286. Journal of Computational Physics, 158(2):262, March 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964568.

Sherwin:2001:LEB

[SC01] Spencer J. Sherwin and Mario Casarin. Low-energy basis preconditioning for elliptic substructured solvers based on unstructured spectral/hp element discretization. Journal of Computational Physics, 171(1):394–417, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968056.

Sanyasiraju:2008:LRB

[SC08a] Y. V. S. S. Sanyasiraju and G. Chandhini. Local radial basis function based gridfree scheme for unsteady incompressible viscous flows. Journal of Computational Physics, 227(20):8922–8948, October 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108003690.

Surana:2008:VDC

[SC08b] Amit Surana and Darren Crowdy. Vortex dynamics in complex domains on a spherical surface. Journal of Computational Physics, 227(12):6058–6070, June 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001204.

Sijoy:2009:FDT

[SC09a] C. D. Sijoy and Shashank Chaturvedi. Finite difference time domain algorithm for electromagnetic problems involving material movement. Journal of Computational Physics, 228(6):2282–2295, April 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108006347.

Subbareddy:2009:FDK

[SC09b] Pramod K. Subbareddy and Graham V. Candler. A fully discrete, kinetic energy consistent finite-volume scheme for compressible flows. Journal of Computational Physics, 228 (5):1347–1364, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005573.

Shestakov:2003:CSC

[SCC^{+}03a] A. I. Shestakov, R. H. Cohen, J. A. Crotinger, L. L. LoDestro, A. Tarditi, and X. Q. Xu. Corrigendum to: “Self-consistent modeling of turbulence and transport” [J. Comput. Phys. 185 (2003) 399–426]. Journal of Computational Physics, 186(1):360, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000470. See [SCC^{+}03b].

Shestakov:2003:SCM

[SCC^{+}03b] A. I. Shestakov, R. H. Cohen, J. A. Crotinger, L. L. LoDestro, A. Tarditi, and X. Q. Xu. Self-consistent modeling of turbulence and transport. Journal of Computational Physics, 185(2):399–426, March 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999102000633. See corrigendum [SCC^{+}03a].

Smithe:2009:DPA

[SCC09] David N. Smithe, John R. Cary, and Johan A. Carlsson. Divergence preservation in the ADI algorithms for electromagnetics. Journal of Computational Physics, 228(19):7289–7299, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003489.

Shu:2000:FVA

[SCD00] C. Shu, W. Chen, and H. Du. Free vibration analysis of curvilinear quadrilateral plates by the differential quadrature method. Journal of Computational Physics, 163(2):452–466, September 20, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965768.

Schulze:2008:EKM

[Sch08] Tim P. Schulze. Efficient kinetic Monte Carlo simulation. Journal of Computational Physics, 227(4):2455–2462, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004755.

Svard:2007:SHO

[SCN07] Magnus Svärd, Mark H. Carpenter, and Jan Nordström. A stable high-order finite difference scheme for the compressible Navier–Stokes equations, far-field boundary conditions. Journal of Computational Physics, 225(1):1020–1038, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000174.

Sui:2008:HMS

[SCRL08] Y. Sui, Y. T. Chew, P. Roy, and H. T. Low. A hybrid method to study flow-induced deformation of three-dimensional capsules. Journal of Computational Physics, 227(12):6351–6371, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001605.

Shao:2006:ACG

[SCT06] Sihong Shao, Wei Cai, and Huazhong Tang. Accurate calculation of Green’s function of the Schrödinger equation in a block layered potential. Journal of Computational Physics, 219(2):733–748, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002026.

Servan-Camas:2009:NNS

[SCT09] Borja Servan-Camas and Frank T.-C. Tsai. Non-negativity and stability analyses of lattice Boltzmann method for advection-diffusion equation. Journal of Computational Physics, 228(1):236–256, January 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004749.

Smith:2009:IQD

[SCW^{+}09] M. R. Smith, H. M. Cave, J.-S. Wu, M. C. Jermy, and Y.-S. Chen. An improved quiet direct simulation method for Eulerian fluids using a second-order scheme. Journal of Computational Physics, 228(6):2213–2224, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006281.

Sportisse:2000:RCK

[SD00] Bruno Sportisse and Rafik Djouad. Reduction of chemical kinetics in air pollution modeling. Journal of Computational Physics, 164(2):354–376, November 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966014.

Shapiro:2005:ACCa

[SD05a] Evgeniy Shapiro and Dimitris Drikakis. Artificial compressibility, characteristics-based schemes for variable density, incompressible, multi-species flows. Part I. Derivation of different formulations and constant density limit. Journal of Computational Physics, 210(2):584–607, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002536.

Shapiro:2005:ACCb

[SD05b] Evgeniy Shapiro and Dimitris Drikakis. Artificial compressibility, characteristics-based schemes for variable-density, incompressible, multispecies flows: Part II. Multigrid implementation and numerical tests. Journal of Computational Physics, 210(2):608–631, December 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002548.

Smith:2006:RRH

[SD06] Robert K. Smith and David G. Dritschel. Revisiting the Rossby–Haurwitz wave test case with contour advection. Journal of Computational Physics, 217(2):473–484, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000180.

Sandu:2005:ASA

[SDCC05] Adrian Sandu, Dacian N. Daescu, Gregory R. Carmichael, and Tianfeng Chai. Adjoint sensitivity analysis of regional air quality models. Journal of Computational Physics, 204(1):222–252, March 20, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004140.

Shay:2007:EFP

[SDD07] M. A. Shay, J. F. Drake, and B. Dorland. Equation free projective integration: a multiscale method applied to a plasma ion acoustic wave. Journal of Computational Physics, 226(1):571–585, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001726.

Samulyak:2007:NAM

[SDGX07] Roman Samulyak, Jian Du, James Glimm, and Zhiliang Xu. A numerical algorithm for MHD of free surface flows at low magnetic Reynolds numbers. Journal of Computational Physics, 226(2):1532–1549, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002483.

Schwartzkopff:2004:FHO

[SDM04] Thomas Schwartzkopff, Michael Dumbser, and Claus-Dieter Munz. Fast high order ADER schemes for linear hyperbolic equations. Journal of Computational Physics, 197(2):532–539, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006508.

Sengupta:2007:NCS

[SDR07] T. K. Sengupta, A. Dipankar, and A. Kameswara Rao. A new compact scheme for parallel computing using domain decomposition. Journal of Computational Physics, 220(2):654–677, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002543.

Sengupta:2007:EDB

[SDS07] T. K. Sengupta, A. Dipankar, and P. Sagaut. Error dynamics: Beyond von Neumann analysis. Journal of Computational Physics, 226(2):1211–1218, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002537.

Stock:2008:IVR

[SDT08] Mark J. Stock, Werner J. A. Dahm, and Grétar Tryggvason. Impact of a vortex ring on a density interface using a regularized inviscid vortex sheet method. Journal of Computational Physics, 227(21):9021–9043, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002969.

Shen:2004:FSH

[SE04] Colin Y. Shen and Thomas E. Evans. A free-surface hydrodynamic model for density-stratified flow in the weakly to strongly non-hydrostatic regime. Journal of Computational Physics, 200(2):695–717, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001858.

Sengil:2009:HEV

[SE09] Nevsan Sengil and Firat O. Edis. Highly efficient volume generation reservoirs in molecular simulations of gas flows. Journal of Computational Physics, 228(12):4303–4308, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001442.

Serna:2009:CBN

[Ser09] Susana Serna. A characteristic-based nonconvex entropy-fix upwind scheme for the ideal magnetohydrodynamic equations. Journal of Computational Physics, 228(11):4232–4247, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001120.

Sethian:2001:EIA

[Set01] J. A. Sethian. Evolution, implementation, and application of level set and fast marching methods for advancing fronts. Journal of Computational Physics, 169(2):503–555, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966579.

Sears:2003:NEM

[SF03] Mark P. Sears and Laura J. D. Frink. A new efficient method for density functional theory calculations of inhomogeneous fluids. Journal of Computational Physics, 190 (1):184–200, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103002705.

Saurel:2007:RPM

[SFDL07] Richard Saurel, Erwin Franquet, Eric Daniel, and Olivier Le Métayer. A relaxation-projection method for compressible flows. Part I: The numerical equation of state for the Euler equations. Journal of Computational Physics, 223(2):822–845, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004761.

Shahbazi:2007:HOD

[SFE07] Khosro Shahbazi, Paul F. Fischer, and C. Ross Ethier. A high-order discontinuous Galerkin method for the unsteady incompressible Navier–Stokes equations. Journal of Computational Physics, 222(1):391–407, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003585.

Suh:2006:CLE

[SFMP06] Jungsoo Suh, Steven H. Frankel, Luc Mongeau, and Michael W. Plesniak. Compressible large eddy simulations of wall-bounded turbulent flows using a semi-implicit numerical scheme for low Mach number aeroacoustics. Journal of Computational Physics, 215(2):526–551, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005036.

Shin:2006:NMI

[SFVK06] C. S. Shin, M. C. Fivel, M. Verdier, and S. C. Kwon. Numerical methods to improve the computing efficiency of discrete dislocation dynamics simulations. Journal of Computational Physics, 215(2):417–429, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004948.

Steinhoff:2000:NEM

[SFW00] John Steinhoff, Meng Fan, and Lesong Wang. A new Eulerian method for the computation of propagating short acoustic and electromagnetic pulses. Journal of Computational Physics, 157(2):683–706, January 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963891.

Shen:2003:SSR

[SFX03] Ching Shen, Jing Fan, and Chong Xie. Statistical simulation of rarefied gas flows in micro-channels. Journal of Computational Physics, 189(2):512–526, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002316.

Schmidt:2001:TBC

[SFY01] Frank Schmidt, Tilmann Friese, and David Yevick. Transparent boundary conditions for split-step Padé approximations of the one-way Helmholtz equation. Journal of Computational Physics, 170(2):696–719, July 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967579.

Sprague:2003:SEF

[SG03a] Michael A. Sprague and Thomas L. Geers. Spectral elements and field separation for an acoustic fluid subject to cavitation. Journal of Computational Physics, 184(1):149–162, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000244.

Sultanov:2003:IDE

[SG03b] Renat A. Sultanov and Dennis Guster. Integral-differential equations approach to atomic three-body systems. Journal of Computational Physics, 192(1):231–243, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103003735.

Schmitz:2006:DVS

[SG06]  H. Schmitz and R. Grauer.  Darwin–Vlasov simula-

tions of magnetised plasmas. Journal of Computational Physics, 214(2):738–756, May 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004717.

Sengupta:2003:ACU

[SGD03] T. K. Sengupta, G. Ganeriwal, and S. De. Analysis of central and upwind compact schemes. Journal of Computational Physics, 192(2):677–694, December 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004133.

Schofield:2009:SOA

[SGFL09] Samuel P. Schofield, Rao V. Garimella, Marianne M. François, and Raphaël Loubère. A second-order accurate material-order-independent interface reconstruction technique for multi-material flow simulations. Journal of Computational Physics, 228(3):731–745, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108005081.

Sovinec:2004:NMS

[SGG^{+}04] C. R. Sovinec, A. H. Glasser, T. A. Gianakon, D. C. Barnes, R. A. Nebel, S. E. Kruger, D. D. Schnack, S. J. Plimpton, A. Tarditi, M. S. Chu, and the NIMROD Team. Nonlinear magnetohydrodynamics simulation using high-order finite elements. Journal of Computational Physics, 195(1):355–386, March 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005369.

Schlamp:2007:HMV

[SH07a] Stefan Schlamp and Bryan C. Hathorn. Higher moments of the velocity distribution function in dense-gas shocks. Journal of Computational Physics, 223(1):305–315, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004384.

Shanker:2007:ACE

[SH07b] B. Shanker and H. Huang. Accelerated Cartesian expansions — a fast method for computing of potentials of the form R^{−ν} for all real ν. Journal of Computational Physics, 226(1):732–753, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002045.

Sharma:2007:PMA

[SH07c] Prateek Sharma and Gregory W. Hammett. Preserving monotonicity in anisotropic diffusion. Journal of Computational Physics, 227(1):123–142, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003233.

Shahbazi:2005:EEP

[Sha05] Khosro Shahbazi. An explicit expression for the penalty parameter of the interior penalty method. Journal of Computational Physics, 205(2):401–407, May 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004796.

Sescu:2008:MOF

[SHA08] Adrian Sescu, Ray Hixon, and Abdollah A. Afjeh. Multidimensional optimization of finite difference schemes for computational aeroacoustics. Journal of Computational Physics, 227(9):4563–4588, April 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000363.

Sherlock:2008:MCM

[She08] M. Sherlock. A Monte–Carlo method for Coulomb collisions in hybrid plasma models. Journal of Computational Physics, 227(4):2286–2292, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005359.

Shin:2007:CCF

[Shi07] Seungwon Shin. Computation of the curvature field in numerical simulation of multiphase flow. Journal of Computational Physics, 222(2):872–878, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004190.

Skeel:2007:CMB

[SHP07] Robert D. Skeel, David J. Hardy, and James C. Phillips. Correcting mesh-based force calculations to conserve both energy and momentum in molecular dynamics simulations. Journal of Computational Physics, 225(1):1–5, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001271.

Soria-Hoyo:2009:PBP

[SHPC09] C. Soria-Hoyo, F. Pontiga, and A. Castellanos. A PIC based procedure for the integration of multiple time scale problems in gas discharge physics. Journal of Computational Physics, 228(4):1017–1029, March 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005366.

Shi:2002:TTN

[SHS02] Jing Shi, Changqing Hu, and Chi-Wang Shu. A technique of treating negative weights in WENO schemes. Journal of Computational Physics, 175(1):108–127, January 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968925.

Schnack:2008:EIL

[SHS08] Jürgen Schnack, Peter Hage, and Heinz-Jürgen Schmidt. Efficient implementation of the Lanczos method for magnetic systems. Journal of Computational Physics, 227(9):4512–4517, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108000338.

Sivakumar:2009:PVR

[SHTB09] P. Sivakumar, D. G. Hyams, L. K. Taylor, and W. R. Briley. A primitive-variable Riemann method for solution of the shallow water equations with wetting and drying. Journal of Computational Physics, 228(19):7452–7472, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003799.

Sha:2007:ASF

[SHWC07] Wei Sha, Zhixiang Huang, Xianliang Wu, and Mingsheng Chen. Application of the symplectic finite-difference timedomain scheme to electromagnetic simulation. Journal of Computational Physics, 225(1):33–50, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005754.

Sanchez:2000:COC

[SHWW00] R. Sanchez, S. P. Hirshman, J. C. Whitson, and A. S. Ware. COBRA: An optimized code for fast analysis of ideal ballooning stability of three-dimensional magnetic equilibria. Journal of Computational Physics, 161(2):576–588, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965148.

Shyue:2001:FMT

[Shy01] Keh-Ming Shyue. A fluid-mixture type algorithm for compressible multicomponent flow with Mie–Grüneisen equation of state. Journal of Computational Physics, 171(2):678–707, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968019.

Shyue:2004:FMT

[Shy04] Keh-Ming Shyue. A fluid-mixture type algorithm for barotropic two-fluid flow problems. Journal of Computational Physics, 200(2):718–748, November 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400186X.

Shyue:2006:WPB

[Shy06] Keh-Ming Shyue. A wave-propagation based volume tracking method for compressible multicomponent flow in two space dimensions. Journal of Computational Physics, 215(1):219–244, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004869.

Shi:2007:HRK

[SHY07] Yu-Hsin Shi, J. C. Huang, and J. Y. Yang. High resolution kinetic beam schemes in generalized coordinates for ideal quantum gas dynamics. Journal of Computational Physics, 222(2):573–591, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003688.

Siewert:2000:PTC

[Sie00] C. E. Siewert. Poiseuille and thermal-creep flow in a cylindrical tube. Journal of Computational Physics, 160(2):470–480, May 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964647.

Shin:2002:MTD

[SJ02] Seungwon Shin and Damir Juric. Modeling three-dimensional multiphase flow using a level contour reconstruction method for front tracking without connectivity. Journal of Computational Physics, 180(2):427–470, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970865.

Schmitt:2004:NSR

[SJ04] D. Schmitt and D. Jault. Numerical study of a rotating fluid in a spheroidal container. Journal of Computational Physics, 197(2):671–685, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006582.

St-Jean:2007:SCM

[SJC07] Cédric St-Jean and Paul Charbonneau. Solar cycle modelling using spatiotemporal decomposition schemes. Journal of Computational Physics, 223(1):50–66, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004219.

Sengupta:2005:NFV

[SJD05] T. K. Sengupta, R. Jain, and A. Dipankar. A new fluxvector splitting compact finite volume scheme. Journal of Computational Physics, 207(1):261–281, July 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000264.

Simens:2009:HRC

[SJHM09] Mark P. Simens, Javier Jiménez, Sergio Hoyas, and Yoshinori Mizuno. A high-resolution code for turbulent boundary layers. Journal of Computational Physics, 228(11):4218–4231, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001119.

Sutherland:2003:IBC

[SK03] James C. Sutherland and Christopher A. Kennedy. Improved boundary conditions for viscous, reacting, compressible flows. Journal of Computational Physics, 191(2):502–524, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003280.

Sirisup:2004:SVM

[SK04a] S. Sirisup and G. E. Karniadakis. A spectral viscosity method for correcting the long-term behavior of POD models. Journal of Computational Physics, 194(1):92–116, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004625.

Suda:2004:APA

[SK04b] Reiji Suda and Shingo Kuriyama. Another preprocessing algorithm for generalized one-dimensional fast multipole method.

Journal of Computational Physics, 195(2):790–803, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005849.

Shi:2005:CPA

[SK05] Fengyan Shi and James T. Kirby. Curvilinear parabolic approximation for surface wave transformation with wavecurrent interaction. Journal of Computational Physics, 204 (2):562–586, April 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104004280.

Symeonidis:2006:FTS

[SK06] Vasileios Symeonidis and George Em Karniadakis. A family of time-staggered schemes for integrating hybrid DPD models for polymers: Algorithms and applications. Journal of Computational Physics, 218(1):82–101, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000611.

Spira:2007:GCF

[SK07a] Alon Spira and Ron Kimmel. Geometric curve flows on parametric manifolds. Journal of Computational Physics, 223(1):235–249, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106004347.

Sugiyama:2007:MSP

[SK07b] Tooru Sugiyama and Kanya Kusano. Multi-scale plasma simulation by the interlocking of magnetohydrodynamic model and particle-in-cell kinetic model. Journal of Computational Physics, 227(2):1340–1352, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003968.

Sentoku:2008:NMP

[SK08a] Y. Sentoku and A. J. Kemp. Numerical methods for particle simulations at extreme densities and temperatures: Weighted particles, relativistic collisions and reduced currents. Journal of Computational Physics, 227(14):6846–6861, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001988.

Skamarock:2008:TSN

[SK08b] William C. Skamarock and Joseph B. Klemp. A timesplit nonhydrostatic atmospheric model for weather research and forecasting applications. Journal of Computational Physics, 227(7):3465–3485, March 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000459.

Sheng:2001:SGN

[SKAS01] Q. Sheng, A. Q. M. Khaliq, and E. A. Al-Said. Solving the generalized nonlinear Schrödinger equation via quartic spline approximation. Journal of Computational Physics, 166(2):400–417, January 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966683.

Seiwald:2008:OEC

[SKK^{+}08] B. Seiwald, S. V. Kasilov, W. Kernbichler, V. N. Kalyuzhnyj, V. V. Nemov, V. Tribaldos, and J. A. Jiménez. Optimization of energy confinement in the 1/ν regime for stellarators. Journal of Computational Physics, 227(12):6165–6183, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001265.

Kim:2003:CSI

[sKKRH03] Sung soo Kim, Chongam Kim, Oh-Hyun Rho, and Seung Kyu Hong. Cures for the shock instability: Development of a shock-stable Roe scheme. Journal of Computational Physics, 185(2):342–374, March 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000372.

Samaey:2006:PDB

[SKR06] Giovanni Samaey, Ioannis G. Kevrekidis, and Dirk Roose. Patch dynamics with buffers for homogenization problems.

Journal of Computational Physics, 213(1):264–287, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003803.

Sansour:2008:NIF

[SKS08] Carlo Sansour, Igor Karsaj, and Jurica Soríc. On a numerical implementation of a formulation of anisotropic continuum elastoplasticity at finite strains. Journal of Computational Physics, 227(16):7643–7663, August 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002544.

Stanescu:2005:NSI

[SKW05] Dan Stanescu, Dongjin Kim, and Wojbor A. Woyczynski. Numerical study of interacting particles approximation for integro-differential equations. Journal of Computational Physics, 206(2):706–726, July 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000124.

Schmidt:2003:NWC

[SKWN03] Rodney C. Schmidt, Alan R. Kerstein, Scott Wunsch, and Vebjorn Nilsen. Near-wall LES closure based on onedimensional turbulence modeling. Journal of Computational Physics, 186(1):317–355, March 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000718.

Sirisup:2005:EFG

[SKXK05] Sirod Sirisup, George Em Karniadakis, Dongbin Xiu, and Ioannis G. Kevrekidis. Equation-free/Galerkin-free PODassisted computation of incompressible flows. Journal of Computational Physics, 207(2):568–587, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000410.

Shumlak:2003:ARS

[SL03] U. Shumlak and J. Loverich. Approximate Riemann solver for the two-fluid plasma model. Journal of Computational Physics, 187(2):620–638, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001517.

Sharma:2004:NSB

[SL04] Anupam Sharma and Lyle N. Long. Numerical simulation of the blast impact problem using the direct simulation Monte Carlo (DSMC) method. Journal of Computational Physics, 200(1):211–237, October 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104001512.

Su:2006:CKR

[SL06] C. H. Su and Didier Lucor. Covariance kernel representations of multidimensional second-order stochastic processes. Journal of Computational Physics, 217(1):82–99, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000842.

Shi:2007:LDF

[SL07a] Xing Shi and Siak Piang Lim. A LBM–DLM/FD method for 3D fluid-structure interactions. Journal of Computational Physics, 226(2):2028–2043, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002768.

Shi:2007:FVT

[SL07b] Yan Shi and Chang-Hong Liang. The finite-volume timedomain algorithm using least square method in solving Maxwell’s equations. Journal of Computational Physics, 226 (2):1444–1457, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107002422.

Simpson:2007:NCI

[SL07c] Matthew J. Simpson and Kerry A. Landman. Nonmonotone chemotactic invasion: High-resolution simulations, phase plane analysis and new benchmark problems. Journal of Computational Physics, 225(1):6–12, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001374.

Shu:2007:NIB

[SLC07] C. Shu, N. Liu, and Y. T. Chew. A novel immersed boundary velocity correction-lattice Boltzmann method and its application to simulate flow past a circular cylinder. Journal of Computational Physics, 226(2):1607–1622, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002513.

Shariff:2008:CDA

[SLF08] Karim Shariff, Anthony Leonard, and Joel H. Ferziger. A contour dynamics algorithm for axisymmetric flow. Journal of Computational Physics, 227(21):9044–9062, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004457.

Shellman:2003:MPL

[SLG^{+}03] Spencer D. Shellman, James P. Lewis, Kurt R. Glaesemann, Krzysztof Sikorski, and Gregory A. Voth. Massively parallel linear-scaling algorithm in an ab initio local-orbital totalenergy method. Journal of Computational Physics, 188(1):1– 15, June 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910300069X.

Sengupta:2009:NCS

[SLV09] T. K. Sengupta, V. Lakshmanan, and V. V. S. N. Vijay. A new combined stable and dispersion relation preserving compact scheme for non-periodic problems. Journal of Computational Physics, 228(8):3048–3071, May 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000151.

Li:2008:ASR

[sLwG08] Xue song Li and Chun wei Gu. An all-speed Roe-type scheme and its asymptotic analysis of low Mach number behaviour. Journal of Computational Physics, 227(10):5144–5159, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000697.

Sandham:2002:ESH

[SLY02] N. D. Sandham, Q. Li, and H. C. Yee. Entropy splitting for high-order numerical simulation of compressible turbulence. Journal of Computational Physics, 178(2):307–322, May 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102970221.

Serna:2004:PEM

[SM04] Susana Serna and Antonio Marquina. Power ENO methods: a fifth-order accurate weighted power ENO method. Journal of Computational Physics, 194(2):632–658, March 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004996.

Serna:2005:CSW

[SM05] Susana Serna and Antonio Marquina. Capturing shock waves in inelastic granular gases. Journal of Computational Physics, 209(2):787–795, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002019.

Seo:2006:LPC

[SM06a] Jung H. Seo and Young J. Moon. Linearized perturbed compressible equations for low Mach number aeroacoustics. Journal of Computational Physics, 218(2):702–719, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106001276.

Soares:2006:DAF

[SM06b] D. Soares, Jr. and W. J. Mansur. Dynamic analysis of fluid-soil-structure interaction problems by the boundary element method. Journal of Computational Physics, 219(2):498–512, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002002.

Sahin:2009:ALE

[SM09a]  Mehmet Sahin and Kamran Mohseni.  An arbitrary

Lagrangian–Eulerian formulation for the numerical simulation of flow patterns generated by the hydromedusa Aequorea victoria. Journal of Computational Physics, 228(12):4588–4605, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001570.

Shchepetkin:2009:CCO

[SM09b] Alexander F. Shchepetkin and James C. McWilliams. Correction and commentary for “Ocean forecasting in terrainfollowing coordinates: Formulation and skill assessment of the regional ocean modeling system” by Haidvogel et al., J. Comp. Phys. 227, pp. 3595–3624. Journal of Computational Physics, 228(24):8985–9000, December 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004872. See [Ano08-50].

Smith:2008:EDD

[SMAj08] M. R. Smith, M. N. Macrossan, and M. M. Abdel-jawad. Effects of direction decoupling in flux calculation in finite volume solvers. Journal of Computational Physics, 227(8):4142–4161, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005670.

Shahbazi:2009:MAH

[SMB09] Khosro Shahbazi, Dimitri J. Mavriplis, and Nicholas K. Burgess. Multigrid algorithms for high-order discontinuous Galerkin discretizations of the compressible Navier– Stokes equations. Journal of Computational Physics, 228(21):7917–7940, November 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003878.

Smereka:2006:NAD

[Sme06] Peter Smereka. The numerical approximation of a delta function with application to level set methods. Journal of Computational Physics, 211(1):77–90, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002627.

Srinivasan:2009:TMR

[SMGJ09] K. R. Srinivasan, K. Matous, P. H. Geubelle, and T. L. Jackson. Thermomechanical modeling of regressing heterogeneous solid propellants. Journal of Computational Physics, 228(21):7883–7901, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003775.

Sanchez:2002:CBT

[SML02] J. Sanchez, F. Marques, and J. M. Lopez. A continuation and bifurcation technique for Navier–Stokes flows. Journal of Computational Physics, 180(1):78–98, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970725.

Srivastava:2000:AGA

[SMO00] R. K. Srivastava, D. S. McRae, and M. T. Odman. An adaptive grid algorithm for air-quality modeling. Journal of Computational Physics, 165(2):437–472, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966208.

Shestakov:2001:CCP

[SMP01] A. I. Shestakov, J. L. Milovich, and M. K. Prasad. Combining cell- and point-centered methods in 3D, unstructuredgrid radiation-hydrodynamic codes. Journal of Computational Physics, 170(1):81–111, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967233.

Shirgaonkar:2009:NMF

[SMP09] Anup A. Shirgaonkar, Malcolm A. MacIver, and Neelesh A. Patankar. A new mathematical formulation and fast algorithm for fully resolved simulation of self-propulsion. Journal of Computational Physics, 228(7):2366–2390, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006311.

Shen:2004:CGF

[SMS04] Wen Zhong Shen, Jess A. Michelsen, and Jens Nørkær Sørensen. A collocated grid finite volume method for aeroacoustic computations of low-speed flows. Journal of Computational Physics, 196(1):348–366, May 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006004.

Seo:2008:PCF

[SMS08] Jung H. Seo, Young J. Moon, and Byeong Rog Shin. Prediction of cavitating flow noise by direct numerical simulation. Journal of Computational Physics, 227(13):6511–6531, June 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001630.

Sugawara:2007:TTD

[SMSS07] Hirotake Sugawara, Naoki Mori, Yosuke Sakai, and Yoshiyuki Suda. Timesaving techniques for decision of electron-molecule collisions in Monte Carlo simulation of electrical discharges. Journal of Computational Physics, 223(1):298–304, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106004372.

Satoh:2008:NIA

[SMT^{+}08] M. Satoh, T. Matsuno, H. Tomita, H. Miura, T. Nasuno, and S. Iga. Nonhydrostatic icosahedral atmospheric model (NICAM) for global cloud resolving simulations. Journal of Computational Physics, 227(7):3486–3514, March 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000654.

Svard:2006:OAD

[SN06] Magnus Svärd and Jan Nordström. On the order of accuracy for difference approximations of initial-boundary value problems. Journal of Computational Physics, 218(1):333–352, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000878.

Svard:2008:SHO

[SN08] Magnus Svärd and Jan Nordström. A stable high-order finite difference scheme for the compressible Navier–Stokes equations: No-slip wall boundary conditions. Journal of Computational Physics, 227(10):4805–4824, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005682.

Sanchez:2004:NKC

[SNGAS04] J. Sánchez, M. Net, B. García-Archilla, and C. Simó. Newton– Krylov continuation of periodic orbits for Navier–Stokes flows. Journal of Computational Physics, 201(1):13–33, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001895.

Snider:2001:ITD

[Sni01] D. M. Snider. An incompressible three-dimensional multiphase particle-in-cell model for dense particle flows. Journal of Computational Physics, 170(2):523–549, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967476.

Sial:2003:EMU

[SNLS03] S. Sial, J. Neuberger, T. Lookman, and A. Saxena. Energy minimization using Sobolev gradients: application to phase separation and ordering. Journal of Computational Physics, 189(1):88–97, July 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910300202X.

Shestakov:2008:MDS

[SO08] Aleksei I. Shestakov and Stella S. R. Offner. A multigroup diffusion solver using pseudo transient continuation for a radiation-hydrodynamic code with patch-based AMR. Journal of Computational Physics, 227(3):2154–2186, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004184.

Sockol:2003:MSN

[Soc03] Peter M. Sockol. Multigrid solution of the Navier–Stokes equations at low speeds with large temperature variations. Journal of Computational Physics, 192(2):570–592, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103004091.

Sofonea:2009:IDR

[Sof09] Victor Sofonea. Implementation of diffuse reflection boundary conditions in a thermal lattice Boltzmann model with flux limiters. Journal of Computational Physics, 228(17):6107–6118, September 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002617.

Sousa:2009:FDA

[Sou09] Ercília Sousa. Finite difference approximations for a fractional advection diffusion problem. Journal of Computational Physics, 228(11):4038–4054, June 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000904.

Sussman:2000:CLS

[SP00] Mark Sussman and Elbridge Gerry Puckett. A coupled level set and volume-of-fluid method for computing 3D and axisymmetric incompressible two-phase flows. Journal of Computational Physics, 162(2):301–337, August 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965379.

Sharma:2004:DNS

[SP04] Nitin Sharma and Neelesh A. Patankar. Direct numerical simulation of the Brownian motion of particles by using fluctuating hydrodynamic equations. Journal of Computational Physics, 201(2):466–486, December 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002384.

Sharma:2005:FCT

[SP05a] Nitin Sharma and Neelesh A. Patankar. A fast computation technique for the direct numerical simulation of rigid particulate flows. Journal of Computational Physics, 205(2):439–457, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004747.

Sjogreen:2005:PML

[SP05b] Björn Sjögreen and N. Anders Petersson. Perfectly matched layers for Maxwell’s equations in second order formulation. Journal of Computational Physics, 209(1):19–46, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001336.

Spence:2005:PBS

[SP05c] Alastair Spence and Chris Poulton. Photonic band structure calculations using nonlinear eigenvalue techniques. Journal of Computational Physics, 204(1):65–81, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004073.

Stuhne:2006:RUG

[SP06a] G. R. Stuhne and W. R. Peltier. A robust unstructured grid discretization for 3-dimensional hydrostatic flows in spherical geometry: a new numerical structure for ocean general circulation modeling. Journal of Computational Physics, 213(2):704–729, April 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004134.

Subramanian:2006:HOM

[SP06b] V. Subramanian and J. B. Perot. Higher-order mimetic methods for unstructured meshes. Journal of Computational Physics, 219(1):68–85, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001409.

Schilder:2007:CAT

[SP07] Frank Schilder and Bruce B. Peckham. Computing Arnol’d tongue scenarios. Journal of Computational Physics, 220 (2):932–951, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106002695.

Saurel:2009:SER

[SPB09] Richard Saurel, Fabien Petitpas, and Ray A. Berry. Simple and efficient relaxation methods for interfaces separating compressible fluids, cavitating flows and shocks in multiphase mixtures. Journal of Computational Physics, 228(5):1678–1712, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005895.

Soria:2001:PCS

[SPC01] C. Soria, F. Pontiga, and A. Castellanos. Particle-in-cell simulation of electrical gas discharges. Journal of Computational Physics, 171(1):47–78, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967634.

Spelt:2005:LSA

[Spe05] Peter D. M. Spelt. A level-set approach for simulations of flows with multiple moving contact lines with hysteresis. Journal of Computational Physics, 207(2):389–404, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105000343.

Sokolov:2006:TPC

[SPGR06] I. V. Sokolov, K. G. Powell, T. I. Gombosi, and I. I. Roussev. A TVD principle and conservative TVD schemes for adaptive Cartesian grids. Journal of Computational Physics, 220(1):1–5, December 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106003548.

Stevens:2009:UPC

[SPLM09] David Stevens, Henry Power, Michael Lees, and Herve Morvan. The use of PDE centres in the local RBF Hermitian method for 3D convective-diffusion problems. Journal of Computational Physics, 228(12):4606–4624, July 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001582.

Shahbazi:2003:SOA

[SPM03] Khosro Shahbazi, Marius Paraschivoiu, and Javad Mostaghimi. Second order accurate volume tracking based on remapping for triangular meshes. Journal of Computational Physics, 188(1):100–122, June 10, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001566.

Sportisse:2000:AOS

[Spo00] Bruno Sportisse. An analysis of operator splitting techniques in the stiff case. Journal of Computational Physics, 161(1):140–168, June 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999100964957.

Shi:2005:DLM

[SPT05] Xing Shi and Nhan Phan-Thien. Distributed Lagrange multiplier/fictitious domain method in the framework of lattice Boltzmann method for fluid-structure interactions. Journal of Computational Physics, 206(1):81–94, June 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004942.

Sytine:2000:CTP

[SPW^{+}00] Igor V. Sytine, David H. Porter, Paul R. Woodward, Stephen W. Hodson, and Karl-Heinz Winkler. Convergence tests for the piecewise parabolic method and Navier–Stokes solutions for homogeneous compressible turbulence. Journal of Computational Physics, 158(2):225–238, March 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964161.

Schmidt:2000:NDC

[SR00a] David P. Schmidt and C. J. Rutland. A new droplet collision algorithm. Journal of Computational Physics, 164(1):62–80, October 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965689.

Spivack:2000:SRC

[SR00b] M. Spivack and D. E. Reeve. Source reconstruction in a coastal evolution equation. Journal of Computational Physics, 161(1):169–181, June 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964969.

Sakai:2009:AOS

[SR09a] T. Sakai and L. G. Redekopp. An application of onesided Jacobi polynomials for spectral modeling of vector fields in polar coordinates. Journal of Computational Physics, 228(18):7069–7085, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003398.

Sun:2009:FVL

[SR09b] Yutao Sun and Yu-Xin Ren. The finite volume local evolution Galerkin method for solving the hyperbolic conservation laws. Journal of Computational Physics, 228(13):4945–4960, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001922.

Sanchez-Rocha:2009:CHR

[SRM09] Martín Sánchez-Rocha and Suresh Menon. The compressible hybrid RANS/LES formulation using an additive operator. Journal of Computational Physics, 228(6):2037–2062, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108006141.

Seynaeve:2007:FMA

[SRNV07] Bert Seynaeve, Eveline Rosseel, Bart Nicolaï, and Stefan Vandewalle. Fourier mode analysis of multigrid methods for partial differential equations with random coefficients. Journal of Computational Physics, 224(1):132–149, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006218.

Sanmiguel-Rojas:2005:CGF

[SROCdPFF05] E. Sanmiguel-Rojas, J. Ortega-Casanova, C. del Pino, and R. Fernandez-Feria. A Cartesian grid finite-difference method for 2D incompressible viscous flows in irregular geometries. Journal of Computational Physics, 204(1):302–318, March 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004188.

Sanmiguel-Rojas:2003:ENM

[SROCFF03] E. Sanmiguel-Rojas, J. Ortega-Casanova, and R. Fernandez-Feria. On the efficiency of a numerical method with periodic time strides for solving incompressible flows. Journal of Computational Physics, 186(1):212–229, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000494.

Sun:2007:NAL

[SRX07] Pengtao Sun, Robert D. Russell, and Jinchao Xu. A new adaptive local mesh refinement algorithm and its application on fourth order thin film flow problem. Journal of Computational Physics, 224(2):1021–1048, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600564X.

Swarztrauber:2000:GDS

[SS00] Paul N. Swarztrauber and William F. Spotz. Generalized discrete spherical harmonic transforms. Journal of Computational Physics, 159(2):213–230, April 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964313.

Spotz:2001:PCA

[SS01a] William F. Spotz and Paul N. Swarztrauber. A performance comparison of associated Legendre projections. Journal of Computational Physics, 168(2):339–355, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967002.

Suzuki:2001:LIT

[SS01b] Alfredo T. Suzuki and Alexandre G. M. Schmidt. Loop integrals in three outstanding gauges: Feynman, light-cone, and Coulomb. Journal of Computational Physics, 168(1):207–218, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966944.

Senocak:2002:PBM

[SS02] Inanc Senocak and Wei Shyy. A pressure-based method for turbulent cavitating flow computations. Journal of Computational Physics, 176(2):363–383, March 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102969925.

Sankaranarayanan:2003:SEG

[SS03a] S. Sankaranarayanan and Malcolm L. Spaulding. A study of the effects of grid non-orthogonality on the solution of shallow water equations in boundary-fitted coordinate systems. Journal of Computational Physics, 184(1):299–320, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000426.

Sofonea:2003:VFD

[SS03b] Victor Sofonea and Robert F. Sekerka. Viscosity of finite difference lattice Boltzmann models. Journal of Computational Physics, 184(2):422–434, January 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000268.

Surzhikov:2004:TCP

[SS04] Sergey T. Surzhikov and Joseph S. Shang. Two-component plasma model for two-dimensional glow discharge in magnetic field. Journal of Computational Physics, 199(2):437– 464, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104000919.

Sherer:2005:HOC

[SS05a] Scott E. Sherer and James N. Scott. High-order compact finite-difference methods on general overset grids. Journal of Computational Physics, 210(2):459–496, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002366.

Smolarkiewicz:2005:MEB

[SS05b] Piotr K. Smolarkiewicz and Joanna Szmelter. MPDATA: An edge-based unstructured-grid formulation. Journal of Computational Physics, 206(2):624–649, July 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000082.

Sofonea:2005:BCU

[SS05c] Victor Sofonea and Robert F. Sekerka. Boundary conditions for the upwind finite difference lattice Boltzmann model: Evidence of slip velocity in micro-channel flow. Journal of Computational Physics, 207(2):639–659, August 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000458.

Shepel:2006:NFE

[SS06a] Sergey V. Shepel and Brian L. Smith. New finiteelement/finite-volume level set formulation for modelling two-phase incompressible flows. Journal of Computational Physics, 218(2):479–494, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000945.

Shishkova:2006:NAK

[SS06b] I. N. Shishkova and S. S. Sazhin. A numerical algorithm for kinetic modelling of evaporation processes. Journal of Computational Physics, 218(2):635–653, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001240.

Severac:2007:SVV

[SS07a] E. Severac and E. Serre. A spectral vanishing viscosity for the LES of turbulent flows within rotating cavities. Journal of Computational Physics, 226(2):1234–1255, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002318.

Singh:2007:TDA

[SS07b] Rajkeshar Singh and Wei Shyy. Three-dimensional adaptive Cartesian grid method with conservative interface restructuring and reconstruction. Journal of Computational Physics, 224(1):150–167, May 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600622X.

Soffer:2007:OBN

[SS07c] A. Soffer and C. Stucchio. Open boundaries for the nonlinear Schrödinger equation. Journal of Computational Physics, 225(2):1218–1232, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000368.

Sethian:2008:SPD

[SS08] J. A. Sethian and Ying Shan. Solving partial differential equations on irregular domains with moving interfaces, with applications to superconformal electrodeposition in semiconductor manufacturing. Journal of Computational Physics, 227(13):6411–6447, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001381.

Sammis:2009:GNF

[SS09a] Ian Sammis and John Strain. A geometric nonuniform fast Fourier transform. Journal of Computational Physics, 228(18):7086–7108, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003404.

Sani:2009:SPL

[SS09b] M. Sani and M. S. Saidi. A set of particle locating algorithms not requiring face belonging to cell connectivity data. Journal of Computational Physics, 228(19):7357–7367, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003611.

Smolarkiewicz:2009:IUS

[SS09c] Piotr K. Smolarkiewicz and Joanna Szmelter. Iterated upwind schemes for gas dynamics. Journal of Computational Physics, 228(1):33–54, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004294.

Schwartzentruber:2007:MPC

[SSB07] T. E. Schwartzentruber, L. C. Scalabrin, and I. D. Boyd. A modular particle-continuum numerical method for hypersonic non-equilibrium gas flows. Journal of Computational Physics, 225(1):1159–1174, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000356.

Slepchenko:2000:NAF

[SSC00] Boris M. Slepchenko, James C. Schaff, and Y. S. Choi. Numerical approach to fast reactions in reaction-diffusion systems: Application to buffered calcium waves in bistable models. Journal of Computational Physics, 162(1):186–218, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910096532X.

Slimon:2000:DCA

[SSD00] Scot A. Slimon, Marios C. Soteriou, and Donald W. Davis. Development of computational aeroacoustics equations for subsonic flows using a Mach number expansion approach. Journal of Computational Physics, 159(2):377–406, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100964490.

Schulze:2003:CKM

[SSE03] Tim P. Schulze, Peter Smereka, and Weinan E. Coupling kinetic Monte–Carlo and continuum models with application to epitaxial growth. Journal of Computational Physics, 189(1):197–211, July 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103002080.

Sussman:2007:SIM

[SSH^{+}07] M. Sussman, K. M. Smith, M. Y. Hussaini, M. Ohta, and R. Zhi-Wei. A sharp interface method for incompressible two-phase flows. Journal of Computational Physics, 221(2):469–505, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002981.

Shen:2000:SSR

[SSL00] Hung Tao Shen, Junshan Su, and Lianwu Liu. SPH simulation of river ice dynamics. Journal of Computational Physics, 165(2):752–770, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966397.

Sabetghadam:2009:FSE

[SSN09] Feriedoun Sabetghadam, Shervin Sharafatmandjoor, and Farhang Norouzi. Fourier spectral embedded boundary solution of the Poisson’s and Laplace equations with Dirichlet boundary conditions. Journal of Computational Physics, 228(1):55–74, January 10, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108004336.

Silva:2003:NST

[SSND03] A. L. F. Lima E. Silva, A. Silveira-Neto, and J. J. R. Damasceno. Numerical simulation of two-dimensional flows over a circular cylinder using the immersed boundary method. Journal of Computational Physics, 189(2):351–370, August 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103002146.

Schrimpf:2000:RBO

[SSSWD00] Gerhard Schrimpf, Johannes Schneider, Hermann Stamm-Wilbrandt, and Gunter Dueck. Record breaking optimization results using the ruin and recreate principle. Journal of Computational Physics, 159(2):139–171, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964136.

Sigurgeirsson:2001:APF

[SSW01] Hersir Sigurgeirsson, Andrew Stuart, and Wing-Lok Wan. Algorithms for particle-field simulations with collisions. Journal of Computational Physics, 172(2):766–807, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968585.

Smolarkiewicz:2007:BRL

[SSW^{+}07] Piotr K. Smolarkiewicz, Robert Sharman, Jeffrey Weil, Steven G. Perry, David Heist, and George Bowker. Building resolving large-eddy simulations and comparison with wind tunnel experiments. Journal of Computational Physics, 227(1):633–653, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003506.

Seidl:2001:APS

[ST01] Albert Seidl and Mikio Takai. Accuracy problems in simulation of field emitter devices using finite elements. Journal of Computational Physics, 166(1):159–164, January 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966439.

Sun:2003:AUF

[ST03a] M. Sun and K. Takayama. An artificially upstream flux vector splitting scheme for the Euler equations. Journal of Computational Physics, 189(1):305–329, July 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002122.

Sun:2003:ELS

[ST03b] M. Sun and K. Takayama. Error localization in solutionadaptive grid methods. Journal of Computational Physics, 190(1):346–350, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910300278X.

Singer:2004:PML

[ST04] I. Singer and E. Turkel. A perfectly matched layer for the Helmholtz equation in a semi-infinite strip. Journal of Computational Physics, 201(2):439–465, December 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002372.

Schwab:2006:KLA

[ST06] Christoph Schwab and Radu Alexandru Todor. Karhunen– Loève approximation of random fields by generalized fast multipole methods. Journal of Computational Physics, 217 (1):100–122, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106000349.

Shadid:2005:PFC

[STD^{+}05] J. N. Shadid, R. S. Tuminaro, K. D. Devine, G. L. Hennigan, and P. T. Lin. Performance of fully coupled domain decomposition preconditioners for finite element transport/reaction simulations. Journal of Computational Physics, 205(1):24–47, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004474.

Stiriba:2002:NFS

[Sti02] Youssef Stiriba. A nonlinear flux split method for hyperbolic conservation laws. Journal of Computational Physics, 176(1):20–39, February 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969591.

Stinis:2005:MLA

[Sti05] Panagiotis Stinis. A maximum likelihood algorithm for the estimation and renormalization of exponential densities. Journal of Computational Physics, 208(2):691–703, September 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001269.

Sokolov:2002:AWN

[STiST02] Igor’ V. Sokolov, Eugene V. Timofeev, Jun ichi Sakai, and Kazuyoshi Takayama. Artificial wind — a new framework to construct simple and efficient upwind shock-capturing schemes. Journal of Computational Physics, 181(1):354– 393, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971305.

Stoltz:2007:PSS

[Sto07] Gabriel Stoltz. Path sampling with stochastic dynamics: Some new algorithms. Journal of Computational Physics, 225(1):491–508, July 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106006024.

Strain:2000:FMS

[Str00] John Strain. A fast modular semi-Lagrangian method for moving interfaces. Journal of Computational Physics, 161(2):512– 536, July 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965082.

Strain:2001:FSL

[Str01a] John Strain. A fast semi-Lagrangian contouring method for moving interfaces. Journal of Computational Physics, 170(1):373–394, June 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101967403.

Straughan:2001:STD

[Str01b] B. Straughan. Surface-tension-driven convection in a fluid overlying a porous layer. Journal of Computational Physics, 170(1):320–337, June 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967397.

Strain:2007:LCS

[Str07a] John Strain. Locally-corrected spectral methods and overdetermined elliptic systems. Journal of Computational Physics, 224(2):1243–1254, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005742.

Swanson:2007:CAR

[STR07b] R. C. Swanson, E. Turkel, and C.-C. Rossow. Convergence acceleration of Runge–Kutta schemes for solving the Navier–Stokes equations. Journal of Computational Physics, 224(1):365–388, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000976.

Stupfel:2001:HFE

[Stu01] Bruno Stupfel. A hybrid finite element and integral equation domain decomposition method for the solution of the 3-D scattering problem. Journal of Computational Physics, 172(2):451–471, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968147.

Shukla:2007:VHO

[STZ07] Ratnesh K. Shukla, Mahidhar Tatineni, and Xiaolin Zhong. Very high-order compact finite difference schemes on nonuniform grids for incompressible Navier–Stokes equations. Journal of Computational Physics, 224(2):1064–1094, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005663.

Summers:2000:RBV

[Sum00] D. M. Summers. A representation of bounded viscous flow based on Hodge decomposition of Wall impulse. Journal of Computational Physics, 158(1):28–50, February 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964045.

Sun:2000:SCF

[Sun00] Chenghai Sun. Simulations of compressible flows with strong shocks by an adaptive lattice Boltzmann model. Journal of Computational Physics, 161(1):70–84, June 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964878.

Suresh:2005:ISD

[Sur05] Ambady Suresh. Interaction of a shock with a density disturbance via shock fitting. Journal of Computational Physics, 206(1):6–15, June 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104005194.

Sussman:2003:SOC

[Sus03] Mark Sussman. A second order coupled level set and volumeof-fluid method for computing growth and collapse of vapor bubbles. Journal of Computational Physics, 187(1):110–136, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103000871.

Suslov:2006:NAS

[Sus06] Sergey A. Suslov. Numerical aspects of searching convective/absolute instability transition. Journal of Computational Physics, 212(1):188–217, February 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003177.

Schmidt:2001:SCD

[SUW01] M. Schmidt, D. Uhrlandt, and R. Winkler. Self-consistent description of radial space-charge confinement in DC column plasmas. Journal of Computational Physics, 168(1):26–46, March 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966762.

Scobelev:2000:SSC

[SV00] Boris Yu. Scobelev and Evgenii V. Vorozhtsov. Sufficient stability criteria and uniform stability of difference schemes. Journal of Computational Physics, 165(2):717–751, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966300.

Sommeijer:2007:SIT

[SV07] B. P. Sommeijer and J. G. Verwer. On stabilized integration for time-dependent PDEs. Journal of Computational Physics, 224(1):3–16, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005705.

Sengupta:2009:FIA

[SVB09] T. K. Sengupta, V. V. S. N. Vijay, and S. Bhaumik. Further improvement and analysis of CCD scheme: Dissipation discretization and de-aliasing properties. Journal of Computational Physics, 228(17):6150–6168, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002666.

Sen:2006:NNP

[SVH^{+}06] S. Sen, K. Veroy, D. B. P. Huynh, S. Deparis, N. C. Nguyen, and A. T. Patera. “natural norm” a posteriori error estimators for reduced basis approximations. Journal of Computational Physics, 217(1):37–62, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000830.

Sethian:2000:SBD

[SW00] J. A. Sethian and Andreas Wiegmann. Structural boundary design via level set and immersed interface methods. Journal of Computational Physics, 163(2):489–528, September 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965811.

Sethian:2004:NMS

[SW04a] J. A. Sethian and Jon Wilkening. A numerical model of stress driven grain boundary diffusion. Journal of Computational Physics, 193(1):275–305, January 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004315.

Shashkov:2004:RPA

[SW04b] Mikhail Shashkov and Burton Wendroff. The repair paradigm and application to conservation laws. Journal of Computational Physics, 198(1):265–277, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000270.

Schlottke:2008:DNS

[SW08a] Jan Schlottke and Bernhard Weigand. Direct numerical simulation of evaporating droplets. Journal of Computational Physics, 227(10):5215–5237, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000740.

Schober:2008:DPM

[SW08b] C. M. Schober and T. H. Wlodarczyk. Dispersive properties of multisymplectic integrators. Journal of Computational Physics, 227(10):5090–5104, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000600.

Staniforth:2008:ADC

[SW08c] Andrew Staniforth and Nigel Wood. Aspects of the dynamical core of a nonhydrostatic, deep-atmosphere, unified weather and climate-prediction model. Journal of Computational Physics, 227(7):3445–3464, March 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005626.

Sbalzarini:2006:PHE

[SWB^{+}06] I. F. Sbalzarini, J. H. Walther, M. Bergdorf, S. E. Hieber, E. M. Kotsalis, and P. Koumoutsakos. PPM

— a highly efficient parallel particle-mesh library for the simulation of continuum systems. Journal of Computational Physics, 215(2):566–588, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500505X.

Schlegel:2008:FPM

[SWG08] Fabrice Schlegel, Daehyun Wee, and Ahmed F. Ghoniem. A fast 3D particle method for the simulation of buoyant flow. Journal of Computational Physics, 227(21):9063– 9090, November 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108001782.

Schwendeman:2006:RPH

[SWK06] D. W. Schwendeman, C. W. Wahle, and A. K. Kapila. The Riemann problem and a high-resolution Godunov method for a model of compressible two-phase flow. Journal of Computational Physics, 212(2):490–526, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500330X.

Sheu:2000:ISS

[SWL00] Tony W. H. Sheu, S. K. Wang, and R. K. Lin. An implicit scheme for solving the convection-diffusion-reaction equation in two dimensions. Journal of Computational Physics, 164(1):123–142, October 10, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965884.

Sun:2006:SFV

[SWL06] Yuzhi Sun, Z. J. Wang, and Yen Liu. Spectral (finite) volume method for conservation laws on unstructured grids VI: Extension to viscous flow. Journal of Computational Physics, 215(1):41–58, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500478X.

Shepard:2001:SPA

[SWTM01] Ron Shepard, Albert F. Wagner, Jeffrey L. Tilson, and Michael Minkoff. The subspace projected approximate matrix (SPAM) modification of the Davidson method. Journal of Computational Physics, 172(2):472–514, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968287.

Shao:2003:DTD

[SWZ03] Zhenhai Shao, G. W. Wei, and Shan Zhao. DSC timedomain solution of Maxwell’s equations. Journal of Computational Physics, 189(2):427–453, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002262.

Sun:2009:DDM

[SXyWX09] Pengtao Sun, Guangri Xue, Chao yang Wang, and Jinchao Xu. A domain decomposition method for two-phase transport model in the cathode of a polymer electrolyte fuel cell. Journal of Computational Physics, 228(16):6016–6036, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002599.

Sjogreen:2003:GCH

[SY03] B. Sjögreen and H. C. Yee. Grid convergence of high order methods for multiscale complex unsteady viscous compressible flows. Journal of Computational Physics, 185(1):1–26, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910200044X.

Shi:2008:GKB

[SY08] Yu-Hsin Shi and J. Y. Yang. A gas-kinetic BGK scheme for semiclassical Boltzmann hydrodynamic transport. Journal of Computational Physics, 227(22):9389–9407, November 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108003392.

Shen:2009:EMM

[SY09a] Jie Shen and Xiaofeng Yang. An efficient moving mesh spectral method for the phase-field model of two-phase flows. Journal of Computational Physics, 228(8):2978–2992, May 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000126.

Shigeta:2009:MFS

[SY09b] Takemi Shigeta and D. L. Young. Method of fundamental solutions with optimal regularization techniques for the Cauchy problem of the Laplace equation with singular points. Journal of Computational Physics, 228(6):1903–1915, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006074.

Sheu:2009:DDA

[SYC09] Tony W. H. Sheu, C. H. Yu, and P. H. Chiu. Development of a dispersively accurate conservative level set scheme for capturing interface in two-phase flows. Journal of Computational Physics, 228(3):661–686, February 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005007.

Shen:2006:HRF

[SYG06] Yiqing Shen, Guowei Yang, and Zhi Gao. High-resolution finite compact difference schemes for hyperbolic conservation laws. Journal of Computational Physics, 216(1):114–137, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005450.

Scardovelli:2000:ARC

[SZ00] Ruben Scardovelli and Stephane Zaleski. Analytical relations connecting linear interfaces and volume fractions in rectangular grids. Journal of Computational Physics, 164(1):228–237, October 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965677.

Sampath:2001:NSC

[SZ01] Rajiv Sampath and Nicholas Zabaras. Numerical study of convection in the directional solidification of a binary alloy driven by the combined action of buoyancy, surface tension, and electromagnetic forces. Journal of Computational Physics, 168(2):384–411, April 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967063.

Shukla:2005:DHO

[SZ05] Ratnesh K. Shukla and Xiaolin Zhong. Derivation of highorder compact finite difference schemes for non-uniform grid using polynomial interpolation. Journal of Computational Physics, 204(2):404–429, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400422X.

Sandu:2008:DSO

[SZ08] Adrian Sandu and Lin Zhang. Discrete second order adjoints in atmospheric chemical transport modeling. Journal of Computational Physics, 227(12):5949–5983, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000983.

Schadle:2007:DDM

[SZB^{+}07] Achim Schädle, Lin Zschiedrich, Sven Burger, Roland Klose, and Frank Schmidt. Domain decomposition method for Maxwell’s equations: scattering off periodic structures. Journal of Computational Physics, 226(1):477–493, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001684.

Shen:2009:HOC

[SZC09] Yiqing Shen, Gecheng Zha, and Xiangying Chen. High order conservative differencing for viscous terms and the application to vortex-induced vibration flows. Journal of Computational Physics, 228(22):8283–8300, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004409.

Su:2007:CBA

[SZH07] Xiaohui Su, Yong Zhao, and Xiaoyang Huang. On the characteristics-based ACM for incompressible flows. Journal of Computational Physics, 227(1):1–11, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910700352X.

Sun:2006:WFP

[SZLW06] Yuhui Sun, Y. C. Zhou, Shu-Guang Li, and G. W. Wei. A windowed Fourier pseudospectral method for hyperbolic conservation laws. Journal of Computational Physics, 214(2):466–490, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004547.

Sokolov:2001:SEG

[SZS01] I. V. Sokolov, H.-M. Zhang, and J. I. Sakai. Simple and efficient Godunov scheme for computational relativistic gas dynamics. Journal of Computational Physics, 172(1):209– 234, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968214.

Shi:2003:RHO

[SZS03] Jing Shi, Yong-Tao Zhang, and Chi-Wang Shu. Resolution of high order WENO schemes for complicated flow structures. Journal of Computational Physics, 186(2):690–696, April 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000949.

Touma:2006:CFV

[TA06]  R. Touma and P. Arminjon.  Central finite volume

schemes with constrained transport divergence treatment for three-dimensional ideal MHD. Journal of Computational Physics, 212(2):617–636, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003487.

Takakura:2006:DEF

[Tak06] Yoko Takakura. Direct-expansion forms of ADER schemes for conservation laws and their verification. Journal of Computational Physics, 219(2):855–878, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002191.

Traore:2009:REF

[TAL09] Philippe Traoré, Yves Marcel Ahipo, and Christophe Louste. A robust and efficient finite volume scheme for the discretization of diffusive flux on extremely skewed meshes in complex geometries. Journal of Computational Physics, 228(14):5148–5159, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109001958.

Tan:2005:ASR

[Tan05a] Xiaobo Tan. Almost symplectic Runge–Kutta schemes for Hamiltonian systems. Journal of Computational Physics, 203 (1):250–273, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003262.

Tang:2005:SPG

[Tan05b] Huazhong Tang. On the sonic point glitch. Journal of Computational Physics, 202(2):507–532, January 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002967.

Tang:2008:FDA

[Tan08] Shaoqiang Tang. A finite difference approach with velocity interfacial conditions for multiscale computations of crystalline solids. Journal of Computational Physics, 227(8):4038–4062, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005621.

Tausch:2007:FMS

[Tau07] Johannes Tausch. A fast method for solving the heat equation by layer potentials. Journal of Computational Physics, 224(2):956–969, June 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106005602.

Tausch:2000:FMP

[TB00a] Johannes Tausch and Jerome Butler. Floquet multipliers of periodic waveguides via Dirichlet-to-Neumann maps. Journal of Computational Physics, 159(1):90–102, March 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964349.

Torres:2000:PSM

[TB00b] D. J. Torres and J. U. Brackbill. The point-set method: Front-tracking without connectivity. Journal of Computational Physics, 165(2):620–644, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096635X.

Torrilhon:2004:HOW

[TB04] M. Torrilhon and D. S. Balsara. High order WENO schemes: investigations on non-uniform convergence for MHD Riemann problems. Journal of Computational Physics, 201(2):586–600, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104002566.

Tang:2006:IGT

[TB06] Lei Tang and James D. Baeder. Improving Godunov-type reconstructions for simulation of vortex-dominated flows. Journal of Computational Physics, 213(2):659–675, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004110.

Torres:2009:FRB

[TB09] Claudio E. Torres and L. A. Barba. Fast radial basis function interpolation with Gaussians by localization and iteration. Journal of Computational Physics, 228(14):4976–4999, August 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109001156.

Tryggvason:2001:FTM

[TBE^{+}01] G. Tryggvason, B. Bunner, A. Esmaeeli, D. Juric, N. Al-Rawahi, W. Tauber, J. Han, S. Nas, and Y.-J. Jan. A fronttracking method for the computations of multiphase flow. Journal of Computational Physics, 169(2):708–759, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967269.

Thommes:2009:LBM

[TBJ^{+}09] G. Thömmes, J. Becker, M. Junk, A. K. Vaikuntam, D. Kehrwald, A. Klar, K. Steiner, and A. Wiegmann. A lattice Boltzmann method for immiscible multiphase flow simulations using the level set method. Journal of Computational Physics, 228(4):1139–1156, March 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800541X.

Tegze:2009:AOS

[TBT^{+}09] György Tegze, Gurvinder Bansel, Gyula I. Tóth, Tamás Pusztai, Zhongyun Fan, and László Gránásy. Advanced operator splitting-based semi-implicit spectral method to solve the binary phase-field crystal equations with variable coefficients. Journal of Computational Physics, 228(5):1612–1623, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005858.

Tolstykh:2001:TSL

[TC01a] A. I. Tolstykh and E. N. Chigirev. On thin shear layers numerical simulation. Journal of Computational Physics, 166(1):152–158, January 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966427.

Trebotich:2001:PMI

[TC01b] David P. Trebotich and Phillip Colella. A projection method for incompressible viscous flow on moving quadrilateral grids. Journal of Computational Physics, 166(2):191–217, January 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966543.

Tavener:2002:TFM

[TC02] S. J. Tavener and K. A. Cliffe. Two-fluid Marangoni–Bénard convection with a deformable interface. Journal of Computational Physics, 182(1):277–300, October 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971676.

Taira:2007:IBM

[TC07a] Kunihiko Taira and Tim Colonius. The immersed boundary method: a projection approach. Journal of Computational Physics, 225(2):2118–2137, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001234.

Tong:2007:NME

[TC07b] Mei Song Tong and Weng Cho Chew. Nyström method for elastic wave scattering by three-dimensional obstacles. Journal of Computational Physics, 226(2):1845–1858, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002653.

Tlupova:2009:BIS

[TC09a] Svetlana Tlupova and Ricardo Cortez. Boundary integral solutions of coupled Stokes and Darcy flows. Journal of Computational Physics, 228(1):158–179, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004713.

Tong:2009:MFM

[TC09b] Mei Song Tong and Weng Cho Chew. Multilevel fast multipole algorithm for elastic wave scattering by large threedimensional objects. Journal of Computational Physics, 228 (3):921–932, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005330.

Tome:2000:NSA

[TCM^{+}00] M. F. Tome, A. Castelo, J. Murakami, J. A. Cuminato, R. Minghim, M. C. F. Oliveira, N. Mangiavacchi, and S. Mc-Kee. Numerical simulation of axisymmetric free surface flows. Journal of Computational Physics, 157(2):441–472, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963489.

Trebotich:2005:SCS

[TCM05] D. Trebotich, P. Colella, and G. H. Miller. A stable and convergent scheme for viscoelastic flow in contraction channels. Journal of Computational Physics, 205(1):315–342, May 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104004693.

Thalhammer:2009:HOT

[TCN09] Mechthild Thalhammer, Marco Caliari, and Christof Neuhauser. High-order time-splitting Hermite and Fourier spectral methods. Journal of Computational Physics, 228(3):822–832, February 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108005251.

Tsai:2004:VDP

[TCO^{+}04] Y.-H. R. Tsai, L.-T. Cheng, S. Osher, P. Burchard, and G. Sapiro. Visibility and its dynamics in a PDE based implicit framework. Journal of Computational Physics, 199(1):260–290, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000841.

Tian:2007:HOC

[TD07] Z. F. Tian and S. Q. Dai. High-order compact exponential finite difference methods for convection-diffusion type problems. Journal of Computational Physics, 220(2):952–974, January 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002701.

Tome:2008:NSV

[TdAAP08] M. F. Tomé, M. S. B. de Araujo, M. A. Alves, and F. T. Pinho. Numerical simulation of viscoelastic flows using integral constitutive equations: a finite difference approach. Journal of Computational Physics, 227(8):4207–4243, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005712.

Toth:2006:PEI

[TDGP06] Gábor Tóth, Darren L. De Zeeuw, Tamas I. Gombosi, and Kenneth G. Powell. A parallel explicit/implicit time stepping scheme on block-adaptive grids. Journal of Computational Physics, 217(2):722–758, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000337.

Tromeur-Dervout:2006:CIG

[TDV06] Damien Tromeur-Dervout and Yuri Vassilevski. Choice of initial guess in iterative solution of series of systems arising in fluid flow simulations. Journal of Computational Physics, 219(1):210–227, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001483.

Thornber:2008:EGD

[TDWY08] B. Thornber, D. Drikakis, R. J. R. Williams, and D. Youngs. On entropy generation and dissipation of kinetic energy in high-resolution shock-capturing schemes. Journal of Computational Physics, 227(10):4853–4872, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000405.

Tornberg:2004:NAS

[TE04] Anna-Karin Tornberg and Björn Engquist. Numerical approximations of singular source terms in differential equations. Journal of Computational Physics, 200(2):462–488, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104001767.

Tornberg:2008:CBC

[TE08] Anna-Karin Tornberg and Björn Engquist. Consistent boundary conditions for the Yee scheme. Journal of Computational Physics, 227(14):6922–6943, July 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002039.

Temam:2006:SIC

[Tem06] Roger Temam. Suitable initial conditions. Journal of Computational Physics, 218(2):443–450, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001884.

Teng:2003:EBC

[Ten03] Zhen-Huan Teng. Exact boundary condition for timedependent wave equation based on boundary integral. Journal of Computational Physics, 190(2):398–418, September 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910300281X.

Tanushev:2009:GBD

[TET09] Nicolay M. Tanushev, Björn Engquist, and Richard Tsai. Gaussian beam decomposition of high frequency wave fields. Journal of Computational Physics, 228(23):8856–8871, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004768.

Tseng:2003:GCI

[TF03] Yu-Heng Tseng and Joel H. Ferziger. A ghost-cell immersed boundary method for flow in complex geometry. Journal of Computational Physics, 192(2):593–623, December 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004108.

Teyssier:2006:KDU

[TFD06] Romain Teyssier, Sébastien Fromang, and Emmanuel Dormy. Kinematic dynamos using constrained transport with high order Godunov schemes and adaptive mesh refinement. Journal of Computational Physics, 218(1):44–67, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000593.

Turpault:2004:MHS

[TFDK04] R. Turpault, M. Frank, B. Dubroca, and A. Klar. Multigroup half space moment approximations to the radiative heat transfer equations. Journal of Computational Physics, 198(1):363–371, July 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104000336.

Tysanner:2004:MBF

[TG04] Martin W. Tysanner and Alejandro L. Garcia. Measurement bias of fluid velocity in molecular simulations. Journal of Computational Physics, 196(1):173–183, May 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005928.

Tornberg:2006:NMS

[TG06] Anna-Karin Tornberg and Katarina Gustavsson. A numerical method for simulations of rigid fiber suspensions. Journal of Computational Physics, 215(1):172–196, June 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004845.

Tornberg:2008:FMM

[TG08] Anna-Karin Tornberg and Leslie Greengard. A fast multipole method for the three-dimensional Stokes equations. Journal of Computational Physics, 227(3):1613–1619, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002744.

Tomar:2007:TPE

[TGB^{+}07] G. Tomar, D. Gerlach, G. Biswas, N. Alleborn, A. Sharma, F. Durst, S. W. J. Welch, and A. Delgado. Two-phase electrohydrodynamic simulations using a volume-of-fluid approach. Journal of Computational Physics, 227(2):1267–1285, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003932.

Thuburn:2001:ANA

[TH01] John Thuburn and Thomas W. N. Haine. Adjoints of nonoscillatory advection schemes. Journal of Computational Physics, 171(2):616–631, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967993.

Toro:2009:FSU

[THD09] Eleuterio F. Toro, Arturo Hidalgo, and Michael Dumbser. FORCE schemes on unstructured meshes I: Conservative hyperbolic systems. Journal of Computational Physics, 228(9):3368–3389, May 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000412.

Tang:2006:PSM

[THL06] Shaoqiang Tang, Thomas Y. Hou, and Wing Kam Liu. A pseudo-spectral multiscale method: Interfacial conditions and coarse grid equations. Journal of Computational Physics, 213(1):57–85, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003591.

Tian:2007:DRT

[THN^{+}07] Yue Tian, S.-H. Hung, Guust Nolet, Raffaella Montelli, and F. A. Dahlen. Dynamic ray tracing and traveltime corrections for global seismic tomography. Journal of Computational Physics, 226(1):672–687, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700201X.

Thommes:2004:LIS

[Thö04] Guido Thömmes. A linear iterative scheme for the fast solution of the radiative heat transfer equations for glass. Journal of Computational Physics, 193(2):544–562, January 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004182.

Touil:2007:TDH

[THS07] H. Touil, M. Y. Hussaini, and M. Sussman. Tracking discontinuities in hyperbolic conservation laws with spectral accuracy. Journal of Computational Physics, 225(2):1810–1826, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000885.

Thuburn:2008:NWP

[Thu08a] J. Thuburn. Numerical wave propagation on the hexagonal Cgrid. Journal of Computational Physics, 227(11):5836–5858, May 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108001162.

Thuburn:2008:SCI

[Thu08b] J. Thuburn. Some conservation issues for the dynamical cores of NWP and climate models. Journal of Computational Physics, 227(7):3715–3730, March 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003755.

Tam:2009:FDC

[TJ09] Christopher K. W. Tam and Hongbin Ju. Finite difference computation of acoustic scattering by small surface inhomogeneities and discontinuities. Journal of Computational Physics, 228(16):5917–5932, September 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002447.

Tyagi:2008:LSM

[TJLT08] Manav Tyagi, Patrick Jenny, Ivan Lunati, and Hamdi A. Tchelepi. A Lagrangian, stochastic modeling framework for multi-phase flow in porous media. Journal of Computational Physics, 227(13):6696–6714, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001745.

Tang:2003:OGM

[TJS03] H. S. Tang, S. Casey Jones, and Fotis Sotiropoulos. An overset-grid method for 3D unsteady incompressible flows. Journal of Computational Physics, 191(2):567–600, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003310.

Tam:2000:WBE

[TK00] Christopher K. W. Tam and Konstantin A. Kurbatskii. A wavenumber based extrapolation and interpolation method for use in conjunction with high-order finite difference schemes. Journal of Computational Physics, 157(2):588–617, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199963933.

Timonov:2002:EAS

[TK02] Alexandre Timonov and Michael V. Klibanov. An efficient algorithm for solving the inverse problem of locating the interfaces using the frequency sounding data. Journal of Computational Physics, 183(2):422–437, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972001.

Trangenstein:2004:OSA

[TK04] John A. Trangenstein and Chisup Kim. Operator splitting and adaptive mesh refinement for the Luo–Rudy I model. Journal of Computational Physics, 196(2):645–679, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103006119.

Tiwari:2009:PPH

[TKH09] Sudarshan Tiwari, Axel Klar, and Steffen Hardt. A particleparticle hybrid method for kinetic and continuum equations. Journal of Computational Physics, 228(18):7109–7124, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003428.

Tang:2006:NCS

[TL06] Huazhong Tang and Tiegang Liu. A note on the conservative schemes for the Euler equations. Journal of Computational Physics, 218(2):451–459, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001896.

Tang:2004:ROD

[TLAD04] Deman Tang, Aiqin Li, Peter Attar, and Earl H. Dowell. Reduced order dynamic model for polysaccharides molecule attached to an atomic force microscope. Journal of Computational Physics, 201(2):723–752, December 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002621.

Tan:2007:AMR

[TLK07] Zhijun Tan, K. M. Lim, and B. C. Khoo. An adaptive mesh redistribution method for the incompressible mixture flows using phase-field model. Journal of Computational Physics, 225(1):1137–1158, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000344.

Tan:2009:IIM

[TLK09] Zhijun Tan, K. M. Lim, and B. C. Khoo. An immersed interface method for Stokes flows with fixed/moving interfaces and rigid boundaries. Journal of Computational Physics, 228(18):6855–6881, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003271.

Tan:2008:IIM

[TLL^{+}08] Zhijun Tan, D. V. Le, Zhilin Li, K. M. Lim, and B. C. Khoo. An immersed interface method for solving incompressible viscous flows with piecewise constant viscosity across a moving elastic membrane. Journal of Computational Physics, 227(23):9955–9983, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004464.

Tartakovsky:2005:SPH

[TM05] Alexandre M. Tartakovsky and Paul Meakin. A smoothed particle hydrodynamics model for miscible flow in threedimensional fractures and the two-dimensional Rayleigh– Taylor instability. Journal of Computational Physics, 207 (2):610–624, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105000434.

Tadjeran:2007:SOA

[TM07] Charles Tadjeran and Mark M. Meerschaert. A secondorder accurate numerical method for the two-dimensional fractional diffusion equation. Journal of Computational Physics, 220(2):813–823, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002610.

Tanguy:2007:LSM

[TMB07] Sébastien Tanguy, Thibaut Ménard, and Alain Berlemont. A level set method for vaporizing two-phase flows. Journal of Computational Physics, 221(2):837–853, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003214.

Thornber:2007:ILE

[TMD07] Ben Thornber, Andrew Mosedale, and Dimitris Drikakis. On the implicit large eddy simulations of homogeneous decaying turbulence. Journal of Computational Physics, 226 (2):1902–1929, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107002690.

Thornber:2008:IRM

[TMD^{+}08] B. Thornber, A. Mosedale, D. Drikakis, D. Youngs, and R. J. R. Williams. An improved reconstruction method for compressible flows with low Mach number features. Journal of Computational Physics, 227(10):4873–4894, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108000429.

Toth:2008:HMB

[TMG08] Gábor Tóth, Yingjuan Ma, and Tamas I. Gombosi. Hall magnetohydrodynamics on block-adaptive grids. Journal of Computational Physics, 227(14):6967–6984, July 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002076.

Tian:2007:CTA

[TMND07] Yue Tian, Raffaella Montelli, Guust Nolet, and F. A. Dahlen. Computing traveltime and amplitude sensitivity kernels in finite-frequency tomography. Journal of Computational Physics, 226(2):2271–2288, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003129.

Tadjeran:2006:SOA

[TMS06] Charles Tadjeran, Mark M. Meerschaert, and Hans-Peter Scheffler. A second-order accurate numerical approximation for the fractional diffusion equation. Journal of Computational Physics, 213(1):205–213, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003773.

Tartakovsky:2007:SRT

[TMSW07] Alexandre M. Tartakovsky, Paul Meakin, Timothy D. Scheibe, and Rogene M. Eichler West. Simulations of reactive transport and precipitation with smoothed particle hydrodynamics. Journal of Computational Physics, 222(2):654–672, March 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600372X.

Tai:2002:SCF

[TNGH02] Y. C. Tai, S. Noelle, J. M. N. T. Gray, and K. Hutter. Shockcapturing and front-tracking methods for granular avalanches. Journal of Computational Physics, 175(1):269–301, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969463.

Tymczak:2002:SNM

[TNR02] C. J. Tymczak, Anders M. N. Niklasson, and Heinrich Röder. Separable and nonseparable multiwavelets in multiple dimensions. Journal of Computational Physics, 175(2):363–397, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967439.

Tokar:2006:NSC

[Tok06a] Mikhail Z. Tokar. Numerical solution of continuity equation with a flux non-linearly depending on the density gradient. Journal of Computational Physics, 220(1):175–183, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002270.

Tokman:2006:EIL

[Tok06b] M. Tokman. Efficient integration of large stiff systems of ODEs with exponential propagation iterative (EPI) methods. Journal of Computational Physics, 213(2):748–776, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004158.

Tolstykh:2002:EVD

[Tol02a] Mikhail A. Tolstykh. Erratum: “Vorticity-Divergence Semi-Lagrangian Shallow-Water Model of the Sphere Based on Compact Finite Differences”: Volume 179, Number 1 (June 10, 2002), pages 180–200. Journal of Computational Physics, 181(1):394, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971354. See [Tol02b].

Tolstykh:2002:VDS

[Tol02b] Mikhail A. Tolstykh. Vorticity-divergence semi-Lagrangian shallow-water model of the sphere based on compact finite differences. Journal of Computational Physics, 179(1):180–200, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970506. See erra-

tum [Tol02a].

Tolstykh:2007:DAO

[Tol07] Andrei I. Tolstykh. Development of arbitrary-order multioperatorsbased schemes for parallel calculations. 1: Higher-than-fifth-order approximations to convection terms. Journal of Computational Physics, 225(2):2333–2353, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001350.

Tolstykh:2008:DAO

[Tol08] Andrei I. Tolstykh. Development of arbitrary-order multioperatorsbased schemes for parallel calculations.: Part 2: Families of compact approximations with two-diagonal inversions and related multioperators. Journal of Computational Physics, 227(5):2922–2940, February 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107005207.

Torrilhon:2003:NUC

[Tor03] M. Torrilhon. Non-uniform convergence of finite volume schemes for Riemann problems of ideal magnetohydrodynamics. Journal of Computational Physics, 192(1):73–94, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003474.

Toth:2000:CSC

[Tót00] Gábor Tóth. The ∇ · B = 0 constraint in shockcapturing magnetohydrodynamics codes. Journal of Computational Physics, 161(2):605–652, July 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965197.

Toth:2002:COD

[Tót02] Gábor Tóth. Conservative and orthogonal discretization for the Lorentz force. Journal of Computational Physics, 182(1):346–354, October 10, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971779.

Towers:2007:TMD

[Tow07] John D. Towers. Two methods for discretizing a delta function supported on a level set. Journal of Computational Physics, 220(2):915–931, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002683.

Towers:2008:CRT

[Tow08] John D. Towers. A convergence rate theorem for finite difference approximations to delta functions. Journal of Computational Physics, 227(13):6591–6597, June 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001708.

Towers:2009:DDF

[Tow09a] John D. Towers. Discretizing delta functions via finite differences and gradient normalization. Journal of Computational Physics, 228(10):3816–3836, June 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109000825.

Towers:2009:FDM

[Tow09b] John D. Towers. Finite difference methods for approximating Heaviside functions. Journal of Computational Physics, 228(9):3478–3489, May 20, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000576.

Toda:2009:MDC

[TOY09] Kunihiko Toda, Youichi Ogata, and Takashi Yabe. Multidimensional conservative semi-Lagrangian method of characteristics CIP for the shallow water equations. Journal of Computational Physics, 228(13):4917–4944, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001910.

Takagi:2003:PNM

[TOZP03] S. Takagi, H. N. Ogğuz, Z. Zhang, and A. Prosperetti. PHYSALIS: a new method for particle simulation: Part II: two-dimensional Navier–Stokes flow around cylinders. Journal of Computational Physics, 187(2):371–390, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000779.

Tsidulko:2005:MPC

[TPR05] Yu. Tsidulko, R. Pozzoli, and M. Romé. MEP: a 3D PIC code for the simulation of the dynamics of a non-neutral plasma. Journal of Computational Physics, 209(2):406–420, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001312.

Thanh:2007:VFM

[TPV07] K.-C. Le Thanh, C. Parzani, and M.-H. Vignal. A volume of fluid method for a two-dimensional plasma expansion problem. Journal of Computational Physics, 225(2):1937–1960, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000940.

Turiel:2006:NME

[TPVG06] Antonio Turiel, Conrad J. Pérez-Vicente, and Jacopo Grazzini. Numerical methods for the estimation of multifractal singularity spectra on sampled data: a comparative study. Journal of Computational Physics, 216(1):362–390, July 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005565.

Toledo:2002:VLE

[TR02a] Sivan Toledo and Eran Rabani. Very large electronic structure calculations using an out-of-core filter-diagonalization method. Journal of Computational Physics, 180(1):256–269, July 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970907.

Toth:2002:DCP

[TR02b] G. Tóth and P. L. Roe. Divergence- and curl-preserving prolongation and restriction formulas. Journal of Computational Physics, 180(2):736–750, August 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971202.

Toy:2007:CAV

[TR07] Michael D. Toy and David A. Randall. Comment on the article “Vertical discretizations for compressible Euler equation atmospheric models giving optimal representation of normal modes” by thuburn and woollings. Journal of Computational Physics, 223(1):82–88, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004256. See [TW05].

Tzivion:2001:NFS

[TRL01] Shalva Tzivion, Tamir G. Reisin, and Zev Levin. A new formulation of the spectral multi-moment method for calculating the kinetic collection equation: More accuracy with fewer bins. Journal of Computational Physics, 171(1):418–422, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101967762.

Thuburn:2009:NRG

[TRSK09] J. Thuburn, T. D. Ringler, W. C. Skamarock, and J. B. Klemp. Numerical representation of geostrophic modes on arbitrarily structured C-grids. Journal of Computational Physics, 228(22):8321–8335, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004434.

Thomas:2001:SOR

[TS01] Ingo Thomas and Thomas Sonar. On a second order residual estimator for numerical schemes for nonlinear hyperbolic conservation laws. Journal of Computational Physics, 171(1):227–242, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967841.

Tran:2002:HOM

[TS02] Quang Huy Tran and Bruno Scheurer. High-order monotonicitypreserving compact schemes for linear scalar advection on 2-D irregular meshes. Journal of Computational Physics, 175(2):454–486, January 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101969529.

Tornberg:2004:SDI

[TS04] Anna-Karin Tornberg and Michael J. Shelley. Simulating the dynamics and interactions of flexible fibers in Stokes flows. Journal of Computational Physics, 196(1):8–40, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005837.

Tskhakaya:2007:OPC

[TS07] D. Tskhakaya and R. Schneider. Optimization of PIC codes by improved memory management. Journal of Computational Physics, 225(1):829–839, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700006X.

Torrilhon:2008:BCR

[TS08] Manuel Torrilhon and Henning Struchtrup. Boundary conditions for regularized 13-moment-equations for micro-channel-flows. Journal of Computational Physics, 227(3):1982–2011, January 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004433.

Terracol:2001:MAL

[TSB01] Marc Terracol, Pierre Sagaut, and Claude Basdevant. A multilevel algorithm for large-eddy simulation of turbulent compressible flows. Journal of Computational Physics, 167(2):439–474, March 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966877.

Terracol:2003:TSA

[TSB03] M. Terracol, P. Sagaut, and C. Basdevant. A time selfadaptive multilevel algorithm for large-eddy simulation. Journal of Computational Physics, 184(2):339–365, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000177.

Tomita:2002:OIG

[TSG02] Hirofumi Tomita, Masaki Satoh, and Koji Goto. An optimization of the icosahedral grid modified by spring dynamics. Journal of Computational Physics, 183(1):307–331, November 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971937.

Teleaga:2006:RMT

[TSG^{+}06] Ioan Teleaga, Mohammed Seaïd, Ingenuin Gasser, Axel Klar, and Jens Struckmeier. Radiation models for thermal flows at low Mach number. Journal of Computational Physics, 215(2):506–525, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005012.

Tsukerman:2006:CDS

[Tsu06] Igor Tsukerman. A class of difference schemes with flexible local approximation. Journal of Computational Physics, 211(2):659–699, January 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002925.

Tsynkov:2003:ABC

[Tsy03] S. V. Tsynkov. Artificial boundary conditions for the numerical simulation of unsteady acoustic waves. Journal of Computational Physics, 189(2):626–650, August 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002493.

Tsynkov:2004:ALB

[Tsy04] S. V. Tsynkov. On the application of lacunae-based methods to Maxwell’s equations. Journal of Computational Physics, 199(1):126–149, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000798.

Titarev:2004:FVW

[TT04] V. A. Titarev and E. F. Toro. Finite-volume WENO schemes for three-dimensional conservation laws. Journal of Computational Physics, 201(1):238–260, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002281.

Titarev:2005:AST

[TT05a] V. A. Titarev and E. F. Toro. ADER schemes for threedimensional non-linear hyperbolic systems. Journal of Computational Physics, 204(2):715–736, April 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004358.

Toro:2005:ASS

[TT05b] E. F. Toro and V. A. Titarev. ADER schemes for scalar non-linear hyperbolic conservation laws with source terms in three-space dimensions. Journal of Computational Physics, 202(1):196–215, January 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104002736.

Toro:2006:DRS

[TT06a] E. F. Toro and V. A. Titarev. Derivative Riemann solvers for systems of conservation laws and ADER methods. Journal of Computational Physics, 212(1):150–165, February 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003141.

Toro:2006:MFS

[TT06b] E. F. Toro and V. A. Titarev. MUSTA fluxes for systems of conservation laws. Journal of Computational Physics, 216(2):403–429, August 10, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105005589.

Torres:2006:KUA

[TT06c] David J. Torres and Mario F. Trujillo. KIVA-4: An unstructured ALE code for compressible gas flow with sprays. Journal of Computational Physics, 219(2):943–975, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910600338X.

Terashima:2009:FTG

[TT09] Hiroshi Terashima and Grétar Tryggvason. A front-tracking/ ghost-fluid method for fluid interfaces in compressible flows. Journal of Computational Physics, 228(11):4012–4037, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000898.

Tomita:2001:SWM

[TTSG01] Hirofumi Tomita, Motohiko Tsugawa, Masaki Satoh, and Koji Goto. Shallow water model on a modified icosahedral geodesic grid by using spring dynamics. Journal of Computational Physics, 174(2):579–613, December 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968974.

Tang:2003:AMR

[TTZ03] H.-Z. Tang, Tao Tang, and Pingwen Zhang. An adaptive mesh redistribution method for nonlinear Hamilton– Jacobi equations in two- and three-dimensions. Journal of Computational Physics, 188(2):543–572, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300192X.

Tran:2004:PLS

[TU04] L. B. Tran and H. S. Udaykumar. A particle-level setbased sharp interface Cartesian grid method for impact, penetration, and void collapse. Journal of Computational Physics, 193(2):469–510, January 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003760.

Tucker:2003:DEB

[Tuc03] P. G. Tucker. Differential equation-based wall distance computation for DES and RANS. Journal of Computational Physics, 190(1):229–248, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002729.

Tosatto:2008:NSU

[TV08] Luca Tosatto and Luigi Vigevano. Numerical solution of under-resolved detonations. Journal of Computational Physics, 227(4):2317–2343, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004639.

Talon:2003:FVC

[TVMR03] Suzanne Talon, Alain Vincent, Georges Michaud, and Jacques Richer. A finite volume code for meridional circulation in stars. Journal of Computational Physics, 184(1):244–265, January 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102000293.

Trahan:2003:RBF

[TW03] Corey J. Trahan and Robert E. Wyatt. Radial basis function interpolation in the quantum trajectory method: optimization of the multi-quadric shape parameter. Journal of Computational Physics, 185(1):27–49, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000463.

Thuburn:2005:VDC

[TW05] J. Thuburn and T. J. Woollings. Vertical discretizations for compressible Euler equation atmospheric models giving optimal representation of normal modes. Journal of Computational Physics, 203(2):386–404, March 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003316. See comment [TR07].

Tong:2007:NMC

[TW07] Albert Y. Tong and Zhaoyuan Wang. A numerical method for capillarity-dominant free surface flows. Journal of Computational Physics, 221(2):506–523, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003007.

Taylor:2007:ONE

[TWM07] Ellen M. Taylor, Minwei Wu, and M. Pino Martín. Optimization of nonlinear error for weighted essentially non-oscillatory methods in direct numerical simulations of compressible turbulence. Journal of Computational Physics, 223(1):384–397, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106004426.

Tuminaro:2002:BFN

[TWS02] Raymond S. Tuminaro, Homer F. Walker, and John N. Shadid. On backtracking failure in Newton–GMRES methods with a demonstration for the Navier–Stokes equations. Journal of Computational Physics, 180(2):549–558, August 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971020.

Tang:2006:HOP

[TWYC06] Chen Tang, Wenping Wang, Haiqing Yan, and Zhanqing Chen. High-order predictor-corrector of exponential fitting for the N-body problems. Journal of Computational Physics, 214(2):505–520, May 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004560.

Tang:2000:HOG

[TX00] Hua-Zhong Tang and Kun Xu. A high-order gas-kinetic method for multidimensional ideal magnetohydrodynamics. Journal of Computational Physics, 165(1):69–88, November 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965975.

Tartakovsky:2006:SAT

[TX06] Daniel M. Tartakovsky and Dongbin Xiu. Stochastic analysis of transport in tubes with rough walls. Journal of Computational Physics, 217(1):248–259, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001215.

Tian:2007:TDM

[TXCD07] C. T. Tian, K. Xu, K. L. Chan, and L. C. Deng. A three-dimensional multidimensional gas-kinetic scheme for the Navier–Stokes equations under gravitational fields. Journal of Computational Physics, 226(2):2003–2027, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002756.

Tu:2007:CBS

[TY07] Guo-Hua Tu and Xiang-Jiang Yuan. A characteristic-based shock-capturing scheme for hyperbolic problems. Journal of Computational Physics, 225(2):2083–2097, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001210.

Tygert:2008:FAS

[Tyg08] Mark Tygert. Fast algorithms for spherical harmonic expansions, II. Journal of Computational Physics, 227(8): 4260–4279, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107005736.

Tai:2003:PUI

[TZ03] Chin Hoe Tai and Yong Zhao. Parallel unsteady incompressible viscous flow computations using an unstructured multigrid method. Journal of Computational Physics, 192(1):277–311, November 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003796.

Tan:2006:LSS

[TZ06] Lijian Tan and Nicholas Zabaras. A level set simulation of dendritic solidification with combined features of front-tracking and fixed-domain methods. Journal of Computational Physics, 211(1):36–63, January 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002603.

Tan:2007:LSS

[TZ07a] Lijian Tan and Nicholas Zabaras. A level set simulation of dendritic solidification of multi-component alloys. Journal of Computational Physics, 221(1):9–40, January 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002737.

Tan:2007:MGI

[TZ07b] Lijian Tan and Nicholas Zabaras. Modeling the growth and interaction of multiple dendrites in solidification using a level set method. Journal of Computational Physics, 226(1):131–155, September 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107001465.

Tan:2007:MMA

[TZ07c] Lijian Tan and Nicholas Zabaras. Multiscale modeling of alloy solidification using a database approach. Journal of Computational Physics, 227(1):728–754, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003555.

Tan:2004:MMM

[TZHT04] Zhijun Tan, Zhengru Zhang, Yunqing Huang, and Tao Tang. Moving mesh methods with locally varying time steps. Journal of Computational Physics, 200(1):347–367, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001706.

Tai:2005:PCU

[TZL05] C. H. Tai, Y. Zhao, and K. M. Liew. Parallel computation of unsteady incompressible viscous flows around moving rigid bodies using an immersed object method with overlapping grids. Journal of Computational Physics, 207(1):151–172, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000215.

Unfer:2007:ASL

[UBRT07] Thomas Unfer, Jean-Pierre Boeuf, François Rogier, and Frédéric Thivet. An asynchronous scheme with local time stepping for multi-scale transport problems: Application to gas discharges. Journal of Computational Physics, 227(2):898–918, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003312.

Ueki:2001:EMC

[UH01] Taro Ueki and J. Eduard Hoogenboom. Exact Monte Carlo perturbation analysis by forward-adjoint coupling in radiation transport calculations. Journal of Computational Physics, 171(2):509–533, August 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967889.

Uhlmann:2005:IBM

[Uhl05] Markus Uhlmann. An immersed boundary method with direct forcing for the simulation of particulate flows. Journal of Computational Physics, 209(2):448–476, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001385.

Ujevic:2006:SPD

[UL06] Maximiliano Ujevic and Patricio S. Letelier. Solving procedure for a 25-diagonal coefficient matrix: Direct numerical solutions of the three-dimensional linear Fokker–Planck equation. Journal of Computational Physics, 215(2):485–505, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005000.

Udaykumar:2001:SIC

[UMRK01] H. S. Udaykumar, R. Mittal, P. Rampunggoon, and A. Khanna. A sharp interface Cartesian grid method for simulating flows with complex moving boundaries. Journal of Computational Physics, 174(1):345–380, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969165.

Ustyugov:2009:PPM

[UPKN09] Sergey D. Ustyugov, Mikhail V. Popov, Alexei G. Kritsuk, and Michael L. Norman. Piecewise parabolic method on a local stencil for magnetized supersonic turbulence simulation. Journal of Computational Physics, 228(20):7614–7633, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003830.

Udaykumar:2003:EMC

[UTBV03] H. S. Udaykumar, L. Tran, D. M. Belk, and K. J. Vanden. An Eulerian method for computation of multimaterial impact with ENO shock-capturing and sharp interfaces. Journal of Computational Physics, 186(1):136–177, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000275.

Utnes:2008:SIP

[Utn08]  T. Utnes.  A segregated implicit pressure projection

method for incompressible flows. Journal of Computational Physics, 227(4):2198–2211, February 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700486X.

Utsumi:2004:NBS

[UYK^{+}04] Takayuki Utsumi, Takashi Yabe, James Koga, Takayuki Aoki, Masatoshi Sekine, Youichi Ogata, and Eiichi Matsunaga. A note on the basis set approach in the constrained interpolation profile method. Journal of Computational Physics, 196(1):1– 7, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005874.

Vasilyev:2000:HOF

[Vas00] Oleg V. Vasilyev. High order finite difference schemes on non-uniform meshes with good conservation properties. Journal of Computational Physics, 157(2):746–761, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963982.

Vay:2000:NAL

[Vay00] Jean-Luc Vay. A new absorbing layer boundary condition for the wave equation. Journal of Computational Physics, 165(2):511–521, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966233.

Vay:2001:EFS

[Vay01] Jean-Luc Vay. An extended FDTD scheme for the wave equation: Application to multiscale electromagnetic simulation. Journal of Computational Physics, 167(1):72–98, February 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100966592.

Vay:2002:APM

[Vay02] Jean-Luc Vay. Asymmetric perfectly matched layer for the absorption of waves. Journal of Computational Physics, 183 (2):367–399, December 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971755.

Vasilyev:2000:SGW

[VB00] Oleg V. Vasilyev and Christopher Bowman. Secondgeneration wavelet collocation method for the solution of partial differential equations. Journal of Computational Physics, 165(2):660–693, December 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100966385.

Veerapaneni:2008:CFG

[VB08] Shravan K. Veerapaneni and George Biros. The Chebyshev fast Gauss and nonuniform fast Fourier transforms and their application to the evaluation of distributed heat potentials. Journal of Computational Physics, 227(16):7768–7790, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108002635.

Vanella:2009:MLS

[VB09] Marcos Vanella and Elias Balaras. A moving-least-squares reconstruction for embedded-boundary formulations. Journal of Computational Physics, 228(18):6617–6628, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003246.

Vasil:2008:NMFa

[VBJ08a] Geoffrey M. Vasil, Nicholas H. Brummell, and Keith Julien. A new method for fast transforms in parity-mixed PDEs: Part I. Numerical techniques and analysis. Journal of Computational Physics, 227(17):7999–8016, September 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002441.

Vasil:2008:NMFb

[VBJ08b] Geoffrey M. Vasil, Nicholas H. Brummell, and Keith Julien. A new method for fast transforms in parity-mixed PDEs: Part II. Application to confined rotating convection. Journal of Computational Physics, 227(17):8017–8034, September 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108002428.

vanBrummelen:2003:PIC

[vBK03] E. H. van Brummelen and B. Koren. A pressure-invariant conservative Godunov-type method for barotropic two-fluid flows. Journal of Computational Physics, 185(1):289–308, February 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910200058X.

Vola:2003:LUF

[VBL03] D. Vola, L. Boscardin, and J. C. Latché. Laminar unsteady flows of Bingham fluids: a numerical strategy and some benchmark results. Journal of Computational Physics, 187(2):441–456, May 20, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103001189.

Vola:2004:NSC

[VBL04] D. Vola, F. Babik, and J.-C. Latché. On a numerical strategy to compute gravity currents of non-Newtonian fluids. Journal of Computational Physics, 201(2):397–420, December 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002359.

VandenAbeele:2007:DDP

[VBL07] Kris Van den Abeele, Tim Broeckhoven, and Chris Lacor. Dispersion and dissipation properties of the 1D spectral volume method and application to a p-multigrid algorithm. Journal of Computational Physics, 224(2):616–636, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005419.

vanBrummelen:2001:ENS

[vBRK01] E. H. van Brummelen, H. C. Raven, and B. Koren. Efficient numerical solution of steady free-surface Navier–Stokes flow. Journal of Computational Physics, 174(1):120–137, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101968809.

Vincent:2000:OCL

[VC00] Stéphane Vincent and Jean-Paul Caltagirone. A onecell local multigrid method for solving unsteady incompressible multiphase flows. Journal of Computational Physics, 163(1):172–215, September 1, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965665.

Vaithianathan:2003:NAS

[VC03] T. Vaithianathan and Lance R. Collins. Numerical approach to simulating turbulent flow of a viscoelastic polymer solution. Journal of Computational Physics, 187(1):1–21, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000287.

Vasconcelos:2003:DTA

[VCG03] Igor F. Vasconcelos, Isabelle Cantat, and James A. Glazier. Dynamics and topological aspects of a reconstructed twodimensional foam time series using Potts model on a pinned lattice. Journal of Computational Physics, 192(1):1–20, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003413.

Vosbeek:2000:ACD

[VCM00] P. W. C. Vosbeek, H. J. H. Clercx, and R. M. M. Mattheij. Acceleration of contour dynamics simulations with a hierarchical-element method. Journal of Computational Physics, 161(1):287–311, June 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965057.

Vanhille:2000:HOF

[VCP00] Christian Vanhille and Cleofé Campos-Pozuelo. A high-order finite-difference algorithm for the analysis of standing acoustic waves of finite but moderate amplitude. Journal of Computational Physics, 165(2):334–353, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966117.

Valerio:2007:FEI

[VCT07] J. V. Valério, M. S. Carvalho, and C. Tomei. Filtering the eigenvalues at infinite from the linear stability analysis of incompressible flows. Journal of Computational Physics, 227(1):229–243, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003282.

Valerio:2009:ECS

[VCT09] J. V. Valério, M. S. Carvalho, and C. Tomei. Efficient computation of the spectrum of viscoelastic flows. Journal of Computational Physics, 228(4):1172–1187, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108005433.

Vignolo:2002:EFD

[VCTS02] P. Vignolo, M. L. Chiofalo, M. P. Tosi, and Sauro Succi. Explicit finite-difference and particle method for the dynamics of mixed Bose-condensate and cold-atom clouds. Journal of Computational Physics, 182(2):368–391, November 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102971718.

Vukovic:2004:WSB

[VCZS04] Senka Vukovic, Nelida Crnjaric-Zic, and Luka Sopta. WENO schemes for balance laws with spatially varying flux. Journal of Computational Physics, 199(1):87–109, September 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104000774.

Venditti:2000:AEE

[VD00] David A. Venditti and David L. Darmofal. Adjoint error estimation and grid adaptation for functional outputs: Application to quasi-one-dimensional flow. Journal of Computational Physics, 164(1):204–227, October 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966002.

Venditti:2002:GAF

[VD02] David A. Venditti and David L. Darmofal. Grid adaptation for functional outputs: Application to two-dimensional inviscid flows. Journal of Computational Physics, 176(1):40–69, February 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969670.

Venditti:2003:AGA

[VD03] David A. Venditti and David L. Darmofal. Anisotropic grid adaptation for functional outputs: application to two-dimensional viscous flows. Journal of Computational Physics, 187(1):22–46, May 1, 2003. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103000743.

vanderBos:2009:NSP

[vdBG09]  Fedderik van der Bos and Volker Gravemeier.  Nu-

merical simulation of premixed combustion using an enriched finite element method. Journal of Computational Physics, 228(10):3605–3624, June 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109000606.

vandenDoel:2006:LSR

[vdDA06] K. van den Doel and U. M. Ascher. On level set regularization for highly ill-posed distributed parameter estimation problems. Journal of Computational Physics, 216(2):707–723, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000076.

vanderHolst:2007:HBA

[vdHK07] B. van der Holst and R. Keppens. Hybrid block-AMR in Cartesian and curvilinear coordinates: MHD applications. Journal of Computational Physics, 226(1):925–946, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910700215X.

Velichko:2002:SCD

[VDM^{+}02] O. I. Velichko, V. A. Dobrushkin, A. N. Muchynski, V. A. Tsurko, and V. A. Zhuk. Simulation of coupled diffusion of impurity atoms and point defects under nonequilibrium conditions in local domain. Journal of Computational Physics, 178(1):196–209, May 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970294.

vanderSman:2000:CDL

[vdSE00] R. G. M. van der Sman and M. H. Ernst. Convection-diffusion lattice Boltzmann scheme for irregular lattices. Journal of Computational Physics, 160(2):766–782, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096491X.

vanderVen:2008:AMM

[vdV08] H. van der Ven. An adaptive multitime multigrid algorithm for time-periodic flow simulations. Journal of Computational Physics, 227(10):5286–5303, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000788.

vanderVegt:2002:STD

[vdVvdV02] J. J. W. van der Vegt and H. van der Ven. Space–time discontinuous Galerkin finite element method with dynamic grid motion for inviscid compressible flows: I. General formulation. Journal of Computational Physics, 182(2):546–585, November 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102971858.

vanderVegt:2007:STD

[vdVX07] J. J. W. van der Vegt and Yan Xu. Space–time discontinuous Galerkin method for nonlinear water waves. Journal of Computational Physics, 224(1):17–39, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005894.

vanDam:2006:RMM

[vDZ06] A. van Dam and P. A. Zegeling. A robust moving mesh finite volume method applied to 1D hyperbolic conservation laws from magnetohydrodynamics. Journal of Computational Physics, 216(2):526–546, August 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005632.

vanErp:2005:ETI

[vEB05] Titus S. van Erp and Peter G. Bolhuis. Elaborating transition interface sampling methods. Journal of Computational Physics, 205(1):157–181, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004620.

Verboncoeur:2001:SSW

[Ver01] J. P. Verboncoeur. Symmetric spline weighting for charge and current density in particle simulation. Journal of Computational Physics, 174(1):421–427, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969232.

Valorani:2001:ETS

[VG01] Mauro Valorani and Dimitrios A. Goussis. Explicit timescale splitting algorithm for stiff problems: Auto-ignition of gaseous mixtures behind a steady shock. Journal of Computational Physics, 169(1):44–79, May 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967099.

Visbal:2002:UHO

[VG02] Miguel R. Visbal and Datta V. Gaitonde. On the use of higher-order finite-difference schemes on curvilinear and deforming meshes. Journal of Computational Physics, 181(1):155–185, September 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971172.

Veerapaneni:2009:NMS

[VGBZ09] Shravan K. Veerapaneni, Denis Gueyffier, George Biros, and Denis Zorin. A numerical method for simulating the dynamics of 3D axisymmetric vesicles suspended in viscous flows.

Journal of Computational Physics, 228(19):7233–7249, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003441.

Valorani:2005:HOC

[VGCN05] Mauro Valorani, Dimitris A. Goussis, Francesco Creta, and Habib N. Najm. Higher order corrections in the approximation of low-dimensional manifolds and the construction of simplified problems with the CSP method. Journal of Computational Physics, 209(2):754–786, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001981.

Veldman:2007:NSL

[VGL^{+}07] A. E. P. Veldman, J. Gerrits, R. Luppes, J. A. Helder, and J. P. B. Vreeburg. The numerical simulation of liquid sloshing on board spacecraft. Journal of Computational Physics, 224(1):82–99, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006139.

VandenAbeele:2009:SAS

[VGPL09] K. Van den Abeele, G. Ghorbaniasl, M. Parsani, and C. Lacor. A stability analysis for the spectral volume method on tetrahedral grids. Journal of Computational Physics, 228(2):257–265, February 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005263.

Vachal:2004:UMA

[VGS04] Pavel Vachal, Rao V. Garimella, and Mikhail J. Shashkov. Untangling of 2D meshes in ALE simulations. Journal of Computational Physics, 196(2):627–644, May 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006107.

Veerapaneni:2009:BIM

[VGZB09] Shravan K. Veerapaneni, Denis Gueyffier, Denis Zorin, and George Biros. A boundary integral method for simulating the dynamics of inextensible vesicles suspended in a viscous fluid in 2D. Journal of Computational Physics, 228(7):2334–2353, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108006244.

vanHeukelum:2002:DES

[vHBB02] A. van Heukelum, G. T. Barkema, and R. H. Bisseling. DNA electrophoresis studied with the cage model. Journal of Computational Physics, 180(1):313–326, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970956.

Visser:2005:CBM

[VHI05] D. C. Visser, H. C. J. Hoefsloot, and P. D. Iedema. Comprehensive boundary method for solid walls in dissipative particle dynamics. Journal of Computational Physics, 205(2):626–639, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104004838.

Visser:2006:MMV

[VHI06] D. C. Visser, H. C. J. Hoefsloot, and P. D. Iedema. Modelling multi-viscosity systems with dissipative particle dynamics. Journal of Computational Physics, 214(2):491–504, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004559.

Vikhansky:2003:NMI

[Vik03] A. Vikhansky. A new modification of the immersed boundaries method for fluid-solid flows: moderate Reynolds numbers. Journal of Computational Physics, 191(1):328–339, October 10, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103003139.

Vilmart:2008:RRE

[Vil08] Gilles Vilmart. Reducing round-off errors in rigid body dynamics. Journal of Computational Physics, 227(15):7083– 7088, July 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108002313.

Vikhansky:2004:MCM

[VK04] Alexander Vikhansky and Markus Kraft. A Monte Carlo methods for identification and sensitivity analysis of coagulation processes. Journal of Computational Physics, 200 (1):50–59, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104001317.

Vasilyev:2005:AMW

[VK05a] Oleg V. Vasilyev and Nicholas K.-R. Kevlahan. An adaptive multilevel wavelet collocation method for elliptic problems. Journal of Computational Physics, 206(2):412–431, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104005200.

Vikhansky:2005:CMR

[VK05b] Alexander Vikhansky and Markus Kraft. Conservative method for the reduction of the number of particles in the Monte Carlo simulation method for kinetic equations. Journal of Computational Physics, 203(2):371–378, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400395X.

Vire:2009:DES

[VK09] A. Viré and B. Knaepen. On discretization errors and subgrid scale model implementations in large eddy simulations. Journal of Computational Physics, 228(22):8203–8213, December 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006554.

VandenAbeele:2007:ASS

[VL07] Kris Van den Abeele and Chris Lacor. An accuracy and stability study of the 2D spectral volume method. Journal of Computational Physics, 226(1):1007–1026, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002197.

vanLoon:2006:FSI

[vLAvdV06] R. van Loon, P. D. Anderson, and F. N. van de Vosse. A fluid-structure interaction method with solid-rigid contact for heart valve dynamics. Journal of Computational Physics, 217(2):806–823, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000416.

Verhaeghe:2009:LBM

[VLB09] Frederik Verhaeghe, Li-Shi Luo, and Bart Blanpain. Lattice Boltzmann modeling of microchannel flow in slip flow regime. Journal of Computational Physics, 228(1):147–157, January 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004592.

Venkateswaran:2002:CMM

[VLKM02] Sankaran Venkateswaran, Jules W. Lindau, Robert F. Kunz, and Charles L. Merkle. Computation of multiphase mixture flows with compressibility effects. Journal of Computational Physics, 180(1):54–77, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970622.

VandenAbeele:2007:CBS

[VLW07] Kris Van den Abeele, Chris Lacor, and Z. J. Wang. On the connection between the spectral volume and the spectral difference method. Journal of Computational Physics, 227(2):877–885, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003889.

Veeramani:2007:FDF

[VMN07] C. Veeramani, P. D. Minev, and K. Nandakumar. A fictitious domain formulation for flows with rigid particles: a non-Lagrange multiplier version. Journal of Computational Physics, 224(2):867–879, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005523.

VandeWiele:2008:AFM

[VOD08] B. Van de Wiele, F. Olyslager, and L. Dupré. Application of the fast multipole method for the evaluation of magneto-static fields in micromagnetic computations. Journal of Computational Physics, 227(23):9913–9932, December 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004233.

Voller:2004:MCS

[Vol04a] Vaughan R. Voller. A Monte Carlo scheme for tracking filling fronts. Journal of Computational Physics, 200(2):399– 411, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104001731.

Vollmoller:2004:SCW

[Vol04b] P. Vollmöller. A shock-capturing wave-propagation method for dry and saturated granular flows. Journal of Computational Physics, 199(1):150–174, September 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000804.

vanOs:2004:SEM

[vOP04] R. G. M. van Os and T. N. Phillips. Spectral element methods for transient viscoelastic flow problems. Journal of Computational Physics, 201(1):286–314, November 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400230X.

Voss:2006:IPM

[Vos06] Heinrich Voss. Iterative projection methods for computing relevant energy states of a quantum dot. Journal of Computational Physics, 217(2):824–833, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000428.

Valle:2000:CAS

[VP00] R. G. Della Valle and P. Procacci. Computer-aided series expansion for phonon self-energy. Journal of Computational Physics, 165(2):428–436, December 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910096618X.

Valorani:2009:GSF

[VP09a] M. Valorani and S. Paolucci. The G-Scheme: a framework for multi-scale adaptive model reduction. Journal of Computational Physics, 228(13):4665–4701, July 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109001302.

Volkov:2009:IMF

[VP09b] Oleg Volkov and Bartosz Protas. An inverse model for a freeboundary problem with a contact line: Steady case. Journal of Computational Physics, 228(13):4893–4910, July 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001739.

Voller:2002:MLN

[VPA02] V. R. Voller and F. Porté-Agel. Moore’s law and numerical modeling. Journal of Computational Physics, 179(2):698–703, July 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910297083X.

Vaccari:2004:RES

[VPMC04] A. Vaccari, R. Pontalti, C. Malacarne, and L. Cristoforetti. A robust and efficient subgridding algorithm for finitedifference time-domain simulations of Maxwell’s equations. Journal of Computational Physics, 194(1):117–139, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004649.

Vu-Quoc:2004:ATF

[VQLZ04] L. Vu-Quoc, L. Lesburg, and X. Zhang. An accurate tangential force-displacement model for granular-flow simulations: Contacting spheres with plastic deformation, force-driven formulation. Journal of Computational Physics, 196(1):298–326, May 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103005989.

Vu-Quoc:2002:GUS

[VQSZ02] L. Vu-Quoc, V. Srinivas, and Y. Zhai. A generalized unit system for concise electromagnetic formulation and accurate numerical solutions. Journal of Computational Physics, 181(2):407–429, September 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971160.

Villamizar:2002:TDN

[VR02] Vianey Villamizar and Otilio Rojas. Time-dependent numerical method with boundary-conforming curvilinear coordinates applied to wave interactions with prototypical antennas. Journal of Computational Physics, 177(1):1–36, March 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969876.

Villamizar:2007:GCC

[VRM07] Vianey Villamizar, Otilio Rojas, and Joseph Mabey. Generation of curvilinear coordinates on multiply connected regions with boundary singularities. Journal of Computational Physics, 223(2):571–588, May 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004530.

Vukovic:2002:EWS

[VS02]  Senka Vukovic and Luka Sopta.  ENO and WENO

schemes with the exact conservation property for onedimensional shallow water equations. Journal of Computational Physics, 179(2):593–621, July 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970762.

Vikram:2007:FET

[VS07] M. Vikram and B. Shanker. Fast evaluation of time domain fields in sub-wavelength source/observer distributions using accelerated Cartesian expansions (ACE). Journal of Computational Physics, 227(2):1007–1023, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003750.

Valentini:2009:CED

[VS09] Paolo Valentini and Thomas E. Schwartzentruber. A combined Event-Driven/Time-Driven molecular dynamics algorithm for the simulation of shock waves in rarefied gases. Journal of Computational Physics, 228(23):8766–8778, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004690.

Vijalapura:2005:FSM

[VSG05] Prashanth K. Vijalapura, John Strain, and Sanjay Govindjee. Fractional step methods for index-1 differential-algebraic equations. Journal of Computational Physics, 203(1):305– 320, February 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104003298.

Verwer:2004:RTS

[VSH04] J. G. Verwer, B. P. Sommeijer, and W. Hundsdorfer. RKC time-stepping for advection-diffusion-reaction problems. Journal of Computational Physics, 201(1):61–79, November 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001925.

Vuik:2001:CPV

[VSMW01] C. Vuik, A. Segal, J. A. Meijerink, and G. T. Wijma. The construction of projection vectors for a deflated ICCG method applied to problems with extreme contrasts in the coefficients. Journal of Computational Physics, 172(2):426–450, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967956.

Vamos:2003:GRW

[VSV03] Calin Vamos, Nicolae Suciu, and Harry Vereecken. Generalized random walk algorithm for the numerical modeling of complex diffusion processes. Journal of Computational Physics, 186(2):527–544, April 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103000731.

Vidovic:2004:SCF

[VSW04] D. Vidovíc, A. Segal, and P. Wesseling. A superlinearly convergent finite volume method for the incompressible Navier– Stokes equations on staggered unstructured grids. Journal of Computational Physics, 198(1):159–177, July 20, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000233.

Vidovic:2006:SCM

[VSW06] D. Vidovíc, A. Segal, and P. Wesseling. A superlinearly convergent mach-uniform finite volume method for the Euler equations on staggered unstructured grids. Journal of Computational Physics, 217(2):277–294, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000301.

Valentini:2007:HVM

[VTC^{+}07] F. Valentini, P. Trávnícek, F. Califano, P. Hellinger, and A. Mangeney. A hybrid-Vlasov model based on the current advance method for the simulation of collisionless magnetized plasma. Journal of Computational Physics, 225(1):753–770, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000022.

Vomel:2008:SAE

[VTM^{+}08] Christof Vömel, Stanimire Z. Tomov, Osni A. Marques, A. Canning, Lin-Wang Wang, and Jack J. Dongarra. Stateof-the-art eigensolvers for electronic structure calculations of large scale nano-systems. Journal of Computational Physics, 227(15):7113–7124, July 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000326.

Vignoli:2008:ASS

[VTT08] G. Vignoli, V. A. Titarev, and E. F. Toro. ADER schemes for the shallow water equations in channel with irregular bottom elevation. Journal of Computational Physics, 227(4):2463–2480, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107004779.

Vomel:2007:UBS

[VTW^{+}07] Christof Vömel, Stanimire Z. Tomov, Lin-Wang Wang, Osni A. Marques, and Jack J. Dongarra. The use of bulk states to accelerate the band edge state calculation of a semiconductor quantum dot. Journal of Computational Physics, 223(2):774–782, May 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004736.

VanOs:2004:TUV

[VU04] J. J. A. M. Van Os and R. E. Uittenbogaard. Towards the ultimate variance-conserving convection scheme. Journal of Computational Physics, 197(1):197–214, June 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006260.

Verstappen:2003:SPD

[VV03] R. W. C. P. Verstappen and A. E. P. Veldman. Symmetrypreserving discretization of turbulent flow. Journal of Computational Physics, 187(1):343–368, May 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001268.

Valentini:2005:NSI

[VVM05] Francesco Valentini, Pierluigi Veltri, and André Mangeney. A numerical scheme for the integration of the Vlasov–Poisson system of equations, in the magnetized case. Journal of Computational Physics, 210(2):730–751, December 10, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002597.

Varoutis:2008:AIM

[VVS08] Stelios Varoutis, Dimitris Valougeorgis, and Felix Sharipov. Application of the integro-moment method to steadystate two-dimensional rarefied gas flows subject to boundary induced discontinuities. Journal of Computational Physics, 227(12):6272–6287, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001551.

Van:2002:TDF

[VW02] Tri Van and Aihua Wood. A time-domain finite element method for Helmholtz equations. Journal of Computational Physics, 183(2):486–507, December 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102972049.

Visher:2004:SHO

[VWW04] John Visher, Stephen Wandzura, and Amanda White. Stable, high-order discretization for evolution of the wave equation in 1 + 1 dimensions. Journal of Computational Physics, 194(2):395–408, March 1, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004881.

vanZuijlen:2007:HOT

[vZdBB07] A. H. van Zuijlen, A. de Boer, and H. Bijl. Higher-order time integration through smooth mesh deformation for 3D fluid-structure interaction simulations. Journal of Computational Physics, 224(1):414–430, May 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700143X.

vanZon:2007:NIE

[vZS07] Ramses van Zon and Jeremy Schofield. Numerical implementation of the exact dynamics of free rigid bodies. Journal of Computational Physics, 225(1):145–164, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600581X.

Vouvakis:2007:DDA

[VZSL07] Marinos Vouvakis, Kezhong Zhao, Seung-Mo Seo, and Jin-Fa Lee. A domain decomposition approach for non-conformal couplings between finite and boundary elements for unbounded electromagnetic problems in R^{3}. Journal of Computational Physics, 225(1):975–994, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000137.

Winiecki:2002:FSI

[WA02] T. Winiecki and C. S. Adams. A fast semi-implicit finitedifference method for the TDGL equations. Journal of Computational Physics, 179(1):127–139, June 10, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970476.

White:2008:HOF

[WA08] Laurent White and Alistair Adcroft. A high-order finite volume remapping scheme for nonuniform grids: The piecewise quartic method (PQM). Journal of Computational Physics, 227(15):7394–7422, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002593.

Waagan:2009:PMH

[Waa09] K. Waagan. A positive MUSCL–Hancock scheme for ideal magnetohydrodynamics. Journal of Computational Physics, 228(23):8609–8626, December 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004598.

Wagner:2005:TBN

[Wag05] Christopher L. Wagner. Theoretical basis for numerically exact three-dimensional time-domain algorithms. Journal of Computational Physics, 205(1):343–356, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910400470X.

White:2009:HOR

[WAH09] Laurent White, Alistair Adcroft, and Robert Hallberg. Highorder regridding-remapping schemes for continuous isopycnal and generalized coordinates in ocean models. Journal of Computational Physics, 228(23):8665–8692, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109004628.

Walther:2003:IMP

[Wal03] J. H. Walther. An influence matrix particle-particle particlemesh algorithm with exact particle-particle correction. Journal of Computational Physics, 184(2):670–678, January 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102000359.

Wang:2002:SFVa

[Wan02] Z. J. Wang. Spectral (finite) volume method for conservation laws on unstructured grids. Basic formulation: Basic formulation. Journal of Computational Physics, 178(1):210–251, May 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970415.

Wandzura:2004:SHO

[Wan04a] Stephen Wandzura. Stable, high-order discretization for evolution of the wave equation in 2 + 1 dimensions. Journal of Computational Physics, 199(2):763–775, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001305.

Wang:2004:VOR

[Wan04b] Qian Xi Wang. Variable order revised binary treecode. Journal of Computational Physics, 200(1):192–210, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104001500.

Wang:2005:UMM

[Wan05] Q. X. Wang. Unstructured MEL modelling of nonlinear unsteady ship waves. Journal of Computational Physics, 210(1):368–385, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500224X.

Wang:2004:TCF

[WAO^{+}04] Shao-Ping Wang, Mark H. Anderson, Jason G. Oakley, Michael L. Corradini, and Riccardo Bonazza. A thermodynamically consistent and fully conservative treatment of contact discontinuities for compressible multicomponent flows. Journal of Computational Physics, 195(2):528–559, April 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103005679.

Wu:2001:NSS

[WB01] Lin Wu and D. B. Bogy. Numerical simulation of the slider air bearing problem of hard disk drives by two multidimensional upwind residual distribution schemes over unstructured triangular meshes. Journal of Computational Physics, 172(2):640–657, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101968469.

Wang:2009:NAV

[WB09a] Jin Wang and Greg Baker. A numerical algorithm for viscous incompressible interfacial flows. Journal of Computational Physics, 228(15):5470–5489, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002253.

Witteveen:2009:ERM

[WB09b] Jeroen A. S. Witteveen and Hester Bijl. Effect of randomness on multi-frequency aeroelastic responses resolved by unsteady adaptive stochastic finite elements. Journal of Computational Physics, 228(18):7025–7045, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109003374.

Wagner:2009:QQM

[WBM09] Lucas K. Wagner, Michal Bajdich, and Lubos Mitas. QWalk: a quantum Monte Carlo program for electronic structure. Journal of Computational Physics, 228(9):3390–3404, May 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000424.

Wang:2001:OWE

[WC01] Z. J. Wang and R. F. Chen. Optimized weighted essentially nonoscillatory schemes for linear waves with discontinuity. Journal of Computational Physics, 174(1):381–404, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969189.

Werner:2007:SFA

[WC07] Gregory R. Werner and John R. Cary. A stable FDTD algorithm for non-diagonal, anisotropic dielectrics. Journal of Computational Physics, 226(1):1085–1101, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002227.

Werner:2008:EDM

[WC08] Gregory R. Werner and John R. Cary. Extracting degenerate modes and frequencies from time-domain simulations with filter-diagonalization. Journal of Computational Physics, 227(10):5200–5214, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000739.

Waterson:2007:DPB

[WD07] N. P. Waterson and H. Deconinck. Design principles for bounded higher-order convection schemes — a unified approach. Journal of Computational Physics, 224(1):182–207, May 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910700040X.

Wangard:2001:NSM

[WDM01] William Wangard III, David S. Dandy, and Brian J. Miller. A numerically stable method for integration of the multicomponent species diffusion equations. Journal of Computational Physics, 174(1):460–472, November 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196930X.

Wubs:2006:PIO

[WdND06] Fred W. Wubs, Arie C. de Niet, and Henk A. Dijkstra. The performance of implicit ocean models on B- and C-grids. Journal of Computational Physics, 211(1):210–228, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910500269X.

Weijer:2003:FIM

[WDÖ^{+}03] Wilbert Weijer, Henk A. Dijkstra, Hakan Öksüzoğlu, Fred W. Wubs, and Arie C. de Niet. A fully-implicit model of the global ocean circulation. Journal of Computational Physics, 192(2):452–470, December 10, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004042.

Weinbach:2005:RPS

[WE05] Yael Weinbach and Ron Elber. Revisiting and parallelizing SHAKE. Journal of Computational Physics, 209(1):193–206, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001828.

Weare:2009:PFP

[Wea09] Jonathan Weare. Particle filtering with path sampling and an application to a bimodal ocean current model. Journal of Computational Physics, 228(12):4312–4331, July 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109000801.

Weekes:2002:SSN

[Wee02] Suzanne L. Weekes. A stable scheme for the numerical computation of long wave propagation in temporal laminates. Journal of Computational Physics, 176(2):345–362, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999102969913.

Welfert:2007:AIA

[Wel07] B. D. Welfert. Analysis of iterated ADI–FDTD schemes for Maxwell curl equations. Journal of Computational Physics, 222(1):9–27, March 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002336.

Wen:2006:SSC

[Wen06] Xin Wen. A steady state capturing and preserving method for computing hyperbolic systems with geometrical source terms having concentrations. Journal of Computational Physics, 219(1):322–390, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001677.

Wen:2007:HON

[Wen07] Xin Wen. High order numerical methods to a type of delta function integrals. Journal of Computational Physics, 226(2):1952–1967, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002719.

Wen:2009:HON

[Wen09] Xin Wen. High order numerical methods to two dimensional delta function integrals in level set methods. Journal of Computational Physics, 228(11):4273–4290, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109001296.

Wright:2006:SNC

[WF06] Grady B. Wright and Bengt Fornberg. Scattered node compact finite difference-type formulas generated from radial basis functions. Journal of Computational Physics, 212(1):99– 123, February 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105003116.

Wang:2009:IBM

[WFC09] Zeli Wang, Jianren Fan, and Kefa Cen. Immersed boundary method for the simulation of 2D viscous flow based on vorticity-velocity formulations. Journal of Computational Physics, 228(5):1504–1520, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005706.

Waisman:2005:AGG

[WFTS05] Haim Waisman, Jacob Fish, Raymond S. Tuminaro, and John N. Shadid. Acceleration of the generalized global basis (GGB) method for nonlinear problems. Journal of Computational Physics, 210(1):274–291, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002202.

Wee:2006:MIK

[WG06] Daehyun Wee and Ahmed F. Ghoniem. Modified interpolation kernels for treating diffusion and remeshing in vortex methods. Journal of Computational Physics, 213(1):239–263, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105003797.

Wallstedt:2008:EET

[WG08] P. C. Wallstedt and J. E. Guilkey. An evaluation of explicit time integration schemes for use with the generalized interpolation material point method. Journal of Computational Physics, 227(22):9628–9642, November 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910800394X.

Wang:2009:ULC

[WG09] Z. J. Wang and Haiyang Gao. A unifying lifting collocation penalty formulation including the discontinuous Galerkin, spectral volume/difference methods for conservation laws on mixed grids. Journal of Computational Physics, 228(21):8161–8186, November 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004239.

Wang:2001:GSP

[WGCE01] Xiao-Ping Wang, Carlos J. García-Cervera, and Weinan E. A Gauss–Seidel projection method for micromagnetics simulations. Journal of Computational Physics, 171(1):357–372, July 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967932.

Welch:2007:APM

[WGCR07] D. R. Welch, T. C. Genoni, R. E. Clark, and D. V. Rose. Adaptive particle management in a particle-in-cell code. Journal of Computational Physics, 227(1):143–155, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003245.

Winter:2006:MSA

[WGNT06] C. L. Winter, A. Guadagnini, D. Nychka, and D. M. Tartakovsky. Multivariate sensitivity analysis of saturated flow through simulated highly heterogeneous groundwater aquifers. Journal of Computational Physics, 217(1):166–175, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000374.

Wasberg:2009:VMT

[WGRA09] Carl Erik Wasberg, Thor Gjesdal, Bjørn Anders Pettersson Reif, and Øyvind Andreassen. Variational multiscale turbulence modelling in a high order spectral element method. Journal of Computational Physics, 228(19):7333–7356, October 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003593.

Wang:2008:MCM

[WGS^{+}08] Qiqi Wang, David Gleich, Amin Saberi, Nasrollah Etemadi, and Parviz Moin. A Monte Carlo method for solving unsteady adjoint equations. Journal of Computational Physics, 227(12):6184–6205, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001393.

Wang:2006:PAD

[WGSL06] Zhiqiang Wang, Nasr Ghoniem, Sriram Swaminarayan, and Richard LeSar. A parallel algorithm for 3D dislocation dynamics. Journal of Computational Physics, 219(2):608–621, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910600194X.

Wang:2002:ENH

[WH02] Zhiping Wang and George P. Huang. An essentially nonoscillatory high-order Padé-type (ENO-Padé) scheme. Journal of Computational Physics, 177(1):37–58, March 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102969986.

Wichaidit:2005:PMP

[WH05] C. Wichaidit and W. N. G. Hitchon. Propagator methods for plasma simulations: application to breakdown. Journal of Computational Physics, 203(2):650–667, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003973.

White:2000:NMO

[Whi00] Daniel A. White. Numerical modeling of optical gradient traps using the vector finite element method. Journal of Computational Physics, 159(1):13–37, March 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964222.

Wang:2003:NMS

[WHLL03] Weichung Wang, Tsung-Min Hwang, Wen-Wei Lin, and Jinn-Liang Liu. Numerical methods for semiconductor heterostructures with band nonparabolicity. Journal of Computational Physics, 190(1):141–158, September 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002687.

Widmer:2008:SAF

[WHS08] G. Widmer, R. Hiptmair, and Ch. Schwab. Sparse adaptive finite elements for radiative transfer. Journal of Computational Physics, 227(12):6071–6105, June 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001216.

Willemsen:2000:MPC

[WHV^{+}00] S. M. Willemsen, H. C. J. Hoefsloot, D. C. Visser, P. J. Hamersma, and P. D. Iedema. Modelling phase change with dissipative particle dynamics using a consistent boundary condition. Journal of Computational Physics, 162(2):385–394, August 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100965434.

Wiaux:2007:FSS

[WJV07] Y. Wiaux, L. Jacques, and P. Vandergheynst. Fast spin ±2 spherical harmonics transforms and application in cosmology. Journal of Computational Physics, 226(2):2359–2371, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003191.

Walther:2001:TDV

[WK01a] J. H. Walther and P. Koumoutsakos. Three-dimensional vortex methods for particle-laden flows with two-way coupling. Journal of Computational Physics, 167(1):39–71, February 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966567.

Wilhelm:2001:SAD

[WK01b] D. Wilhelm and L. Kleiser. Stability analysis for different formulations of the nonlinear term in P_{N}–P_{N−2} spectral element discretizations of the Navier–Stokes equations. Journal of Computational Physics, 174(1):306–326, November 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969128.

Wang:2004:IBE

[WK04] C. Wang and B. C. Khoo. An indirect boundary element method for three-dimensional explosion bubbles. Journal of Computational Physics, 194(2):451–480, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103004911.

Wan:2005:AME

[WK05] Xiaoliang Wan and George Em Karniadakis. An adaptive multi-element generalized polynomial chaos method for stochastic differential equations. Journal of Computational Physics, 209(2):617–642, November 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001919.

Wan:2006:SEE

[WK06] Xiaoliang Wan and George Em Karniadakis. A sharp error estimate for the fast Gauss transform. Journal of Computational Physics, 219(1):7–12, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106002166.

Wackers:2007:MSM

[WK07] Jeroen Wackers and Barry Koren. Multigrid solution method for the steady RANS equations. Journal of Computational Physics, 226(2):1784–1807, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107002616.

Witteveen:2007:IFT

[WKB07] J. A. S. Witteveen, B. Koren, and P. G. Bakker. An improved front tracking method for the Euler equations. Journal of Computational Physics, 224(2):712–728, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005468.

Wells:2006:DGM

[WKG06] Garth N. Wells, Ellen Kuhl, and Krishna Garikipati. A discontinuous Galerkin method for the Cahn–Hilliard equation. Journal of Computational Physics, 218(2):860–877, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106001367.

Wise:2007:SRS

[WKL07] Steven Wise, Junseok Kim, and John Lowengrub. Solving the regularized, strongly anisotropic Cahn–Hilliard equation by an adaptive nonlinear multigrid method. Journal of Computational Physics, 226(1):414–446, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001647.

Wang:2002:SFVb

[WL02] Z. J. Wang and Yen Liu. Spectral (finite) volume method for conservation laws on unstructured grids: II. Extension to two-dimensional scalar equation. Journal of Computational Physics, 179(2):665–697, July 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970828.

Wagner:2003:CAC

[WL03] Gregory J. Wagner and Wing Kam Liu. Coupling of atomistic and continuum simulations using a bridging scale decomposition. Journal of Computational Physics, 190(1):249– 274, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103002730.

Wang:2006:ESV

[WL06] Z. J. Wang and Yen Liu. Extension of the spectral volume method to high-order boundary representation. Journal of Computational Physics, 211(1):154–178, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002664.

Wu:2006:DVC

[WLC^{+}06] J.-S. Wu, Y.-Y. Lian, G. Cheng, R. P. Koomullil, and K.-C. Tseng. Development and verification of a coupled DSMC– NS scheme using unstructured mesh. Journal of Computational Physics, 219(2):579–607, December 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001938.

Wang:2008:PSC

[WLC^{+}08] Chiaming Wang, Tungyou Lin, Russel Caflisch, Bruce I. Cohen, and Andris M. Dimits. Particle simulation of Coulomb collisions: Comparing the methods of Takizuka & Abe and Nanbu. Journal of Computational Physics, 227(9):4308–4329, April 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107005633.

Wang:2000:EMS

[WLE^{+}00] Hong Wang, Dong Liang, Richard E. Ewing, Stephen L. Lyons, and Guan Qin. An ELLAM–MFEM solution technique for compressible fluid flows in porous media with point sources and sinks. Journal of Computational Physics, 159(2):344–376, April 10, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999100964507.

Wang:2007:ELS

[WLKW07] S. Y. Wang, K. M. Lim, B. C. Khoo, and M. Y. Wang. An extended level set method for shape and topology optimization. Journal of Computational Physics, 221(1):395–421, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106002968.

Wang:2008:ECD

[WLT08] Heyu Wang, Ruo Li, and Tao Tang. Efficient computation of dendritic growth with r-adaptive finite element methods. Journal of Computational Physics, 227(12):5984–6000, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001034.

Wang:2007:ISU

[WM07] Li Wang and Dimitri J. Mavriplis. Implicit solution of the unsteady Euler equations for high-order accurate discontinuous Galerkin discretizations. Journal of Computational Physics, 225(2):1994–2015, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700099X.

Wang:2009:ABA

[WM09] Li Wang and Dimitri J. Mavriplis. Adjoint-based h– p adaptive discontinuous Galerkin methods for the 2D compressible Euler equations. Journal of Computational Physics, 228(20):7643–7661, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003854.

Wu:2007:HOP

[WMH07] Y. Wu, M. F. Modest, and D. C. Haworth. A highorder photon Monte Carlo method for radiative transfer in direct numerical simulation. Journal of Computational Physics, 223(2):898–922, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106004797.

Wollman:2005:NAV

[WO05] Stephen Wollman and Ercument Ozizmir. Numerical approximation of the Vlasov–Poisson–Fokker–Planck system in one dimension. Journal of Computational Physics, 202 (2):602–644, January 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104002992.

Wollman:2009:NAV

[WO09] Stephen Wollman and Ercument Ozizmir. Numerical approximation of the Vlasov–Poisson–Fokker–Planck system in two dimensions. Journal of Computational Physics, 228(18):6629–6669, October 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002605.

Wood:2006:AES

[Woo06] Aihua Wood. Analysis of electromagnetic scattering from an overfilled cavity in the ground plane. Journal of Computational Physics, 215(2):630–641, July 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005085.

Wang:2009:EPM

[WP09] Moran Wang and Ning Pan. Elastic property of multiphase composites with random microstructures. Journal of Computational Physics, 228(16):5978–5988, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002575.

Warburton:2000:PSS

[WPH00] T. Warburton, L. F. Pavarino, and J. S. Hesthaven. A pseudo-spectral scheme for the incompressible Navier–Stokes equations using unstructured nodal elements. Journal of Computational Physics, 164(1):1–21, October 10, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100965872.

Wall:2002:SIM

[WPM02a] Clifton Wall, Charles D. Pierce, and Parviz Moin. A semiimplicit method for resolution of acoustic waves in low Mach number flows. Journal of Computational Physics, 181(2):545–563, September 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910297141X.

Waltz:2002:PCT

[WPM^{+}02b] J. Waltz, G. L. Page, S. D. Milder, J. Wallin, and A. Antunes. A performance comparison of tree data structures for N-body simulation. Journal of Computational Physics, 178(1):1–14, May 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969438.

Wan:2002:DSC

[WPW02] D. C. Wan, B. S. V. Patnaik, and G. W. Wei. Discrete singular convolution-finite subdomain method for the solution of incompressible viscous flows. Journal of Computational Physics, 180(1):229–255, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970890.

Wimmer:2009:OBT

[WR09] Michael Wimmer and Klaus Richter. Optimal blocktridiagonalization of matrices for coherent charge transport. Journal of Computational Physics, 228(23):8548–8565, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004367.

Wang:2003:MNS

[WRu03] Fei Wang and Rizwan-uddin. A modified nodal scheme for the time-dependent, incompressible Navier–Stokes equations. Journal of Computational Physics, 187(1):168–196, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000937.

Wright:2001:EBM

[WS01] Jeffrey A. Wright and Richard W. Smith. An edgebased method for the incompressible Navier–Stokes equations on polygonal meshes. Journal of Computational Physics, 169(1):24–43, May 1, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967051.

Wedi:2004:EGC

[WS04] Nils P. Wedi and Piotr K. Smolarkiewicz. Extending Gal– Chen and somerville terrain-following coordinate transformation on time-dependent curvilinear boundaries. Journal of Computational Physics, 193(1):1–20, January 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003942.

Wu:2009:IVC

[WS09] J. Wu and C. Shu. Implicit velocity correction-based immersed boundary-lattice Boltzmann method and its applications. Journal of Computational Physics, 228(6):1963–1979, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108006116.

Watanabe:2008:FSF

[WSI08] Yasunori Watanabe, Ayumi Saruwatari, and David M. Ingram. Free-surface flows under impacting droplets. Journal of Computational Physics, 227(4):2344–2365, February 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107004640.

Wang:2009:AVC

[WSTW09] Yu Wang, Songhe Song, Zhijun Tan, and Desheng Wang. Adaptive variational curve smoothing based on level set method. Journal of Computational Physics, 228(17):6333– 6348, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109002770.

Wang:2009:HOW

[WSYS09] Wei Wang, Chi-Wang Shu, H. C. Yee, and Björn Sjögreen. High-order well-balanced schemes and applications to nonequilibrium flow. Journal of Computational Physics, 228(18):6682–6702, October 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002836.

Wan:2007:FBM

[WT07a] Decheng Wan and Stefan Turek. Fictitious boundary and moving mesh methods for the numerical simulation of rigid particulate flows. Journal of Computational Physics, 222(1):28–56, March 1, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002725.

Wang:2007:EAM

[WT07b] Han Wang and Huazhong Tang. An efficient adaptive mesh redistribution method for a non-linear Dirac equation. Journal of Computational Physics, 222(1):176–193, March 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003500.

Wang:2008:AGF

[WTL08] Chunwu Wang, Huazhong Tang, and Tiegang Liu. An adaptive ghost fluid finite volume method for compressible gas-water simulations. Journal of Computational Physics, 227(12):6385–6409, June 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108001642.

Wu:2001:TSM

[Wu01] Zi-Niu Wu. Transmission of a slowly moving shock across a nonconservative interface. Journal of Computational Physics, 171(2):579–615, August 10, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910196797X.

Wu:2002:NUC

[Wu02] Zi-Niu Wu. A note on the unified coordinate system for computing shock waves. Journal of Computational Physics, 180(1):110–119, July 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102970786.

Welch:2000:VFB

[WW00] Samuel W. J. Welch and John Wilson. A volume of fluid based method for fluid flows with phase change. Journal of Computational Physics, 160(2):662–682, May 20, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100964817.

Wang:2004:TDA

[WW04] Desheng Wang and Xiao-Ping Wang. A three-dimensional adaptive method based on the iterative grid redistribution. Journal of Computational Physics, 199(2):423–436, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104000890.

Wang:2007:RCE

[WWC07] Moran Wang, Jinku Wang, and Shiyi Chen. Roughness and cavitations effects on electro-osmotic flows in rough microchannels using the lattice Poisson–Boltzmann methods. Journal of Computational Physics, 226(1):836–851, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107002112.

Werder:2005:HAC

[WWK05] Thomas Werder, Jens H. Walther, and Petros Koumoutsakos. Hybrid atomistic-continuum method for the simulation of dense fluid flows. Journal of Computational Physics, 205(1):373–390, May 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104004723.

Wolny:2000:FBD

[WWVG00] Janusz Wolny, Anna Wnek, and Jean-Louis Verger-Gaugry. Fractal behaviour of diffraction pattern of Thue–Morse sequence. Journal of Computational Physics, 163(2):313– 327, September 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910096563X.

Wang:2007:IMS

[WXG07] Lian-Ping Wang, Yan Xue, and Wojciech W. Grabowski. A bin integral method for solving the kinetic collection equation. Journal of Computational Physics, 226(1):59–88, September 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107001428.

Wang:2009:IPC

[WYS09] Zhaoyuan Wang, Jianming Yang, and Frederick Stern. An improved particle correction procedure for the particle level set method. Journal of Computational Physics, 228(16):5819–5837, September 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109002393.

Wu:2000:GOI

[WZ00] Zi-Niu Wu and Hui Zou. Grid overlapping for implicit parallel computation of compressible flows. Journal of Computational Physics, 157(1):2–43, January 1, 2000. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963039.

Weiland:2002:MES

[WZ02] Thomas Weiland and Igor Zagorodnov. Maxwell’s equations for structures with symmetries. Journal of Computational Physics, 180(1):297–312, July 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970944.

Wu:2003:EPM

[WZ03] Zi-Niu Wu and Hui Zou. Efficient parallel method for implicit upwind schemes approximating the Euler equations in gas dynamics. Journal of Computational Physics, 187(2):683–715, May 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103001554.

Wei:2007:VPD

[WZ07] G. W. Wei and Shan Zhao. On the validity of “A proof that the discrete singular convolution (DSC)/Lagrangedistributed approximation function (LDAF) method is inferior to high order finite differences”. Journal of Computational Physics, 226(2):2389–2392, October 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910700246X. See [Boy06].

Wang:2009:SOC

[WZ09] Yin Wang and Jun Zhang. Sixth order compact scheme combined with multigrid method and extrapolation technique for 2D Poisson equation. Journal of Computational Physics, 228(1):137–146, January 10, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004580.

Wang:2004:SFV

[WZL04] Z. J. Wang, Laiping Zhang, and Yen Liu. Spectral (finite) volume method for conservation laws on unstructured grids IV: extension to two-dimensional systems. Journal of Computational Physics, 194(2):716–741, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005035.

Wang:2009:CII

[WZL09a] X. Sheldon Wang, L. T. Zhang, and Wing Kam Liu. On computational issues of immersed finite element methods. Journal of Computational Physics, 228(7):2535–2551, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006487. See erratum [WZL09b].

Wang:2009:ECI

[WZL09b] X. Sheldon Wang, Lucy T. Zhang, and Wing Kam Liu. Erratum to “On computational issues of immersed finite element methods” [J. Comput. Phys. 228 (2009) 2535–2551]. Journal of Computational Physics, 228(18):7125, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003106. See [WZL09a].

Xiao:2006:UFC

[XAI06] Feng Xiao, Ryosuke Akoh, and Satoshi Ii. Unified formulation for compressible and incompressible flows by using multi-integrated moments II: Multi-dimensional version for compressible and incompressible flows. Journal of Computational Physics, 213(1):31–56, March 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500358X.

Xiong:2008:HOF

[XCRX08] Z. Xiong, R. H. Cohen, T. D. Rognlien, and X. Q. Xu. A high-order finite-volume algorithm for Fokker–Planck collisions in magnetized plasmas. Journal of Computational Physics, 227(15):7192–7205, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002052.

Xiang:2006:ACM

[XCY06] Yang Xiang, Albert C. S. Chung, and Jian Ye. An active contour model for image segmentation based on elastic interaction. Journal of Computational Physics, 219(1):455–476, November 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001720.

Xie:2002:EFO

[XCZ02] Zhongqiang Xie, Chi-Hou Chan, and Bo Zhang. An explicit fourth-order orthogonal curvilinear staggered-grid FDTD method for Maxwell’s equations. Journal of Computational Physics, 175(2):739–763, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969657.

Xue:2007:UBC

[XD07] Changfeng Xue and Shaozhong Deng. An upwinding boundary condition capturing method for Maxwell’s equations in media with material interfaces. Journal of Computational Physics, 225(1):342–362, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005948.

Xie:2009:IAP

[XDB09] Dexuan Xie, Ranjan K. Dash, and Daniel A. Beard. An improved algorithm and its parallel implementation for solving a general blood-tissue transport and metabolism model. Journal of Computational Physics, 228(20):7850–7861, November 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900415X.

Xu:2009:ICA

[XDC09] Zhenli Xu, Shaozhong Deng, and Wei Cai. Image charge approximations of reaction fields in solvents with arbitrary ionic strength. Journal of Computational Physics, 228(6): 2092–2099, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108006189.

Xu:2009:NST

[XG09] Liwei Xu and Philippe Guyenne. Numerical simulation of three-dimensional nonlinear water waves. Journal of Computational Physics, 228(22):8446–8466, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004586.

Xu:2003:LBM

[XH03] Kun Xu and Xiaoyi He. Lattice Boltzmann method and gas-kinetic BGK scheme in the low-Mach number viscous flow simulations. Journal of Computational Physics, 190 (1):100–117, September 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103002559.

Xu:2008:MTK

[XHC08] Kun Xu, Xin He, and Chunpei Cai. Multiple temperature kinetic model and gas-kinetic method for hypersonic non-equilibrium flow computations. Journal of Computational Physics, 227(14):6779–6794, July 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001824.

Xu:2007:AAB

[XHW07] Zhenli Xu, Houde Han, and Xiaonan Wu. Adaptive absorbing boundary conditions for Schrödinger-type equations: Application to nonlinear and multi-dimensional problems. Journal of Computational Physics, 225(2):1577–1589, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000691.

Xiao:2004:UFC

[Xia04] Feng Xiao. Unified formulation for compressible and incompressible flows by using multi-integrated moments I: onedimensional inviscid compressible flow. Journal of Computational Physics, 195(2):629–654, April 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910300576X.

Xin:2007:BEF

[XJ07] Weidong Xin and André H. Juffer. A boundary element formulation of protein electrostatics with explicit ions. Journal of Computational Physics, 223(1):416–435, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004463.

Xiu:2001:SLH

[XK01] Dongbin Xiu and George Em Karniadakis. A semi-Lagrangian high-order method for Navier–Stokes equations. Journal of Computational Physics, 172(2):658–684, September 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968470.

Xiu:2003:MUF

[XK03] Dongbin Xiu and George Em Karniadakis. Modeling uncertainty in flow simulations via generalized polynomial chaos. Journal of Computational Physics, 187(1):137–167, May 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103000925.

Xu:2006:LSM

[XLLZ06] Jian-Jun Xu, Zhilin Li, John Lowengrub, and Hongkai Zhao. A level-set method for interfacial flows with surfactant. Journal of Computational Physics, 212(2):590–616, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003475.

Xia:2007:CPR

[XLM07] Guoping Xia, Ding Li, and Charles L. Merkle. Consistent properties reconstruction on adaptive Cartesian meshes for complex fluids computations. Journal of Computational Physics, 225(1):1175–1197, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107000605.

Xu:2005:CFV

[XLP05] Xiaofeng Xu, Joon Sang Lee, and Richard H. Pletcher. A compressible finite volume formulation for large eddy simulation of turbulent pipe flows at low Mach number in Cartesian coordinates. Journal of Computational Physics, 203(1):22–48, February 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910400316X.

Xu:2009:HRD

[XLS09a] Zhiliang Xu, Yingjie Liu, and Chi-Wang Shu. Hierarchical reconstruction for discontinuous Galerkin methods on unstructured grids with a WENO-type linear reconstruction and partial neighboring cells. Journal of Computational Physics, 228(6):2194–2212, April 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006256.

Xu:2009:HRS

[XLS09b] Zhiliang Xu, Yingjie Liu, and Chi-Wang Shu. Hierarchical reconstruction for spectral volume method on unstructured grids. Journal of Computational Physics, 228(16):5787–5802, September 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900237X.

Xie:2007:IDD

[XMP07] Xueying Xie, Lawrence C. Musson, and Matteo Pasquali. An isochoric domain deformation method for computing steady free surface flows with conserved volumes. Journal of Computational Physics, 226(1):398–413, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001635.

Xu:2005:MGK

[XMT05] Kun Xu, Meiliang Mao, and Lei Tang. A multidimensional gas-kinetic BGK scheme for hypersonic viscous flow. Journal of Computational Physics, 203(2):405–421, March 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104003845.

Xu:2006:NPS

[XMT06] Ying Xu, J. M. McDonough, and K. A. Tagavi. A numerical procedure for solving 2D phase-field model problems. Journal of Computational Physics, 218(2):770–793, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600132X.

Xiao:2004:CPS

[XP04a] Feng Xiao and Xindong Peng. A convexity preserving scheme for conservative advection transport. Journal of Computational Physics, 198(2):389–402, August 10, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000385.

Xu:2004:SSE

[XP04b] Chuanju Xu and Richard Pasquetti. Stabilized spectral element computations of high Reynolds number incompressible flows. Journal of Computational Physics, 196(2):680–704, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999103006120.

Xing:2005:HOF

[XS05a] Yulong Xing and Chi-Wang Shu. High order finite difference WENO schemes with the exact conservation property for the shallow water equations. Journal of Computational Physics, 208(1):206–227, September 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500094X.

Xu:2005:LDG

[XS05b] Yan Xu and Chi-Wang Shu. Local discontinuous Galerkin methods for nonlinear schrödinger equations. Journal of Computational Physics, 205(1):72–97, May 1, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004589.

Xu:2005:ADF

[XS05c] Zhengfu Xu and Chi-Wang Shu. Anti-diffusive flux corrections for high order finite difference WENO schemes. Journal of Computational Physics, 205(2):458–485, May 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104004760.

Xing:2006:HOW

[XS06] Yulong Xing and Chi-Wang Shu. High order well-balanced finite volume WENO schemes and discontinuous Galerkin methods for a class of hyperbolic systems with source terms. Journal of Computational Physics, 214(2):567–598, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105004626.

Xiu:2007:PUA

[XS07] Dongbin Xiu and Spencer J. Sherwin. Parametric uncertainty analysis of pulse wave propagation in a model of a human arterial network. Journal of Computational Physics, 226(2):1385–1407, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107002392.

Xiu:2009:ESG

[XS09] Dongbin Xiu and Jie Shen. Efficient stochastic Galerkin methods for random diffusion equations. Journal of Computational Physics, 228(2):266–281, February 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004774.

Xu:2004:IEM

[XSG04] M. Xu, F. Stefani, and G. Gerbeth. The integral equation method for a steady kinematic dynamo problem. Journal of Computational Physics, 196(1):102–125, May 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005898.

Xu:2008:IEA

[XSG08] M. Xu, F. Stefani, and G. Gerbeth. The integral equation approach to kinematic dynamo theory and its application to dynamo experiments in cylindrical geometry. Journal of Computational Physics, 227(17):8130–8144, September 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910800288X.

Xu:2009:ASI

[XSL09] Rui Xu, Peter Stansby, and Dominique Laurence. Accuracy and stability in incompressible SPH (ISPH) based on the projection method and a new approach. Journal of Computational Physics, 228(18):6703–6725, October 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002885.

Xu:2001:KMH

[Xu01a] K. Xu. A kinetic method for hyperbolic-elliptic equations and its application in two-phase flow. Journal of Computational Physics, 166(2):383–399, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966634.

Xu:2001:CDI

[Xu01b] Kun Xu. Comment on “Development of an Improved Gas-Kinetic BGK Scheme for Inviscid and Viscous Flows”. Journal of Computational Physics, 171(2):843–847, August 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967117. See [CKR00, CKR01].

Xu:2001:GKB

[Xu01c] Kun Xu. A gas-kinetic BGK scheme for the Navier– Stokes equations and its connection with artificial dissipation and Godunov method. Journal of Computational Physics, 171(1):289–335, July 20, 2001. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101967907.

Xu:2002:SUS

[Xu02a] Kun Xu. A slope-update scheme for compressible flow simulation. Journal of Computational Physics, 178(1):252–259, May 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999102970270.

Xu:2002:WBG

[Xu02b] Kun Xu. A well-balanced gas-kinetic scheme for the shallowwater equations with source terms. Journal of Computational Physics, 178(2):533–562, May 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970403.

Xu:2008:IIM

[Xu08] Sheng Xu. The immersed interface method for simulating prescribed motion of rigid objects in an incompressible viscous flow. Journal of Computational Physics, 227(10):5045–5071, May 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108000582.

Xu:2006:IIM

[XW06] Sheng Xu and Z. Jane Wang. An immersed interface method for simulating the interaction of a fluid with moving boundaries. Journal of Computational Physics, 216(2):454–493, August 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105005607.

Xia:2007:LDG

[XXS07] Yinhua Xia, Yan Xu, and Chi-Wang Shu. Local discontinuous Galerkin methods for the Cahn–Hilliard type equations. Journal of Computational Physics, 227(1):472–491, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107003439.

Xiao:2001:CCO

[XY01] Feng Xiao and Takashi Yabe. Completely conservative and oscillationless semi-Lagrangian schemes for advection transportation. Journal of Computational Physics, 170(2):498–522, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967464.

Xu:2005:DHP

[XYK05] Hongyi Xu, Weixing Yuan, and Mahmood Khalid. Design of a high-performance unsteady Navier–Stokes solver using a flexible-cycle additive-correction multigrid technique. Journal of Computational Physics, 209(2):504–540, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001816.

Younes:2005:SAD

[YA05] A. Younes and P. Ackerer. Solving the advection-diffusion equation with the Eulerian–Lagrangian localized adjoint method on unstructured meshes and non uniform time stepping. Journal of Computational Physics, 208(1):384– 402, September 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105001178.

Yamaleev:2001:MTE

[Yam01] Nail K. Yamaleev. Minimization of the truncation error by grid adaptation. Journal of Computational Physics, 170(2): 459–497, July 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101967452.

Yamamoto:2005:PMC

[Yam05] Satoru Yamamoto. Preconditioning method for condensate fluid and solid coupling problems in general curvilinear coordinates. Journal of Computational Physics, 207(1):240–260, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000252.

Yang:2008:IMS

[Yan08] Jianke Yang. Iteration methods for stability spectra of solitary waves. Journal of Computational Physics, 227(14):6862–6876, July 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910800199X.

Yang:2009:NCG

[Yan09]  Jianke Yang.  Newton-conjugate-gradient methods for

solitary wave computations. Journal of Computational Physics, 228(18):7007–7024, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003362.

Yang:2008:CFA

[YAvdB^{+}08] Jiaqi Yang, Aria Abubakar, Peter M. van den Berg, Tarek M. Habashy, and Fernando Reitich. A CG–FFT approach to the solution of a stress-velocity formulation of threedimensional elastic scattering problems. Journal of Computational Physics, 227(24):10018–10039, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108004117.

Yang:2006:EBF

[YB06] Jianming Yang and Elias Balaras. An embedded-boundary formulation for large-eddy simulation of turbulent flows interacting with moving boundaries. Journal of Computational Physics, 215(1):12–40, June 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004778.

Yatziv:2006:IFM

[YBS06] Liron Yatziv, Alberto Bartesaghi, and Guillermo Sapiro. O(N) implementation of the fast marching algorithm. Journal of Computational Physics, 212(2):393–399, March 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003736.

Ying:2004:KIA

[YBZ04] Lexing Ying, George Biros, and Denis Zorin. A kernelindependent adaptive fast multipole algorithm in two and three dimensions. Journal of Computational Physics, 196(2): 591–626, May 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103006090.

Ying:2006:HOB

[YBZ06] Lexing Ying, George Biros, and Denis Zorin. A highorder 3D boundary integral equation solver for elliptic PDEs in smooth domains. Journal of Computational Physics, 219(1):247–275, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001641.

Yamaleev:2002:AAG

[YC02] Nail K. Yamaleev and Mark H. Carpenter. On accuracy of adaptive grid methods for captured shocks. Journal of Computational Physics, 181(1):280–316, September 1, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102971251.

Ying:2006:FGC

[YC06a] Lexing Ying and Emmanuel J. Candès. Fast geodesics computation with the phase flow method. Journal of Computational Physics, 220(1):6–18, December 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003664.

Ying:2006:PFM

[YC06b] Lexing Ying and Emmanuel J. Candès. The phase flow method. Journal of Computational Physics, 220(1):184– 215, December 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106002282.

Yamaleev:2009:SMC

[YC09a] Nail K. Yamaleev and Mark H. Carpenter. A systematic methodology for constructing high-order energy stable WENO schemes. Journal of Computational Physics, 228 (11):4248–4272, June 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109001132.

Yamaleev:2009:TOE

[YC09b] Nail K. Yamaleev and Mark H. Carpenter. Third-order energy stable WENO scheme. Journal of Computational Physics, 228(8):3025–3047, May 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900014X.

Young:2005:NMM

[YCL05] D. L. Young, K. H. Chen, and C. W. Lee. Novel meshless method for solving the potential problems with arbitrary domain. Journal of Computational Physics, 209(1):290–321, October 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105001294.

Yue:2005:NMM

[YE05] Xingye Yue and Weinan E. Numerical methods for multiscale transport equations and application to twophase porous media flow. Journal of Computational Physics, 210(2):656–675, December 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002561.

Yue:2007:LMP

[YE07] Xingye Yue and Weinan E. The local microscale problem in the multiscale modeling of strongly heterogeneous media: Effects of boundary conditions and cell size. Journal of Computational Physics, 222(2):556–572, March 20, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003676.

Yeh:2007:SSI

[Yeh07] Kao-San Yeh. The streamline subgrid integration method: I. Quasi-monotonic second-order transport schemes. Journal of Computational Physics, 225(2):1632–1652, August 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107000745.

Yu:2009:IPB

[YF09] Zhao Yu and Liang-Shih Fan. An interaction potential based lattice Boltzmann method with adaptive mesh refinement (AMR) for two-phase flow simulation. Journal of Computational Physics, 228(17):6456–6478, September 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002903.

Yue:2007:ALE

[YFBH07] Pengtao Yue, James J. Feng, Christopher A. Bertelo, and Howard H. Hu. An arbitrary Lagrangian–Eulerian method for simulating bubble growth in polymer foaming. Journal of Computational Physics, 226(2):2229–2249, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003099.

Yang:2006:NSJ

[YFLS06] Xiaofeng Yang, James J. Feng, Chun Liu, and Jie Shen. Numerical simulations of jet pinching-off and drop formation using an energetic variational phase-field method. Journal of Computational Physics, 218(1):417–428, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600091X.

Yevick:2001:CTB

[YFS01] David Yevick, Tilmann Friese, and Frank Schmidt. A comparison of transparent boundary conditions for the Fresnel equation. Journal of Computational Physics, 168(2):433–444, April 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101967087.

Yu:2005:DDI

[YGL05] Huidan Yu, Sharath S. Girimaji, and Li-Shi Luo. DNS and LES of decaying isotropic turbulence with and without frame rotation using lattice Boltzmann method. Journal of Computational Physics, 209(2):599–616, November 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105001907.

Ying:2007:KFB

[YH07a] Wenjun Ying and Craig S. Henriquez. A kernel-free boundary integral method for elliptic boundary value problems. Journal of Computational Physics, 227(2):1046–1074, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107003774.

Yvonnet:2007:RMM

[YH07b] J. Yvonnet and Q.-C. He. The reduced model multiscale method (R3M) for the non-linear homogenization of hyperelastic media at finite strains. Journal of Computational Physics, 223(1):341–368, April 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004402.

Yu:2005:IPS

[YHCD05] P. Yu, S. Y. Hu, L. Q. Chen, and Q. Du. An iterativeperturbation scheme for treating inhomogeneous elasticity in phase-field models. Journal of Computational Physics, 208(1):34–50, September 1, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105000483.

Yang:2007:HOK

[YHSX07] Jaw-Yen Yang, Tse-Yang Hsieh, Yu-Hsin Shi, and Kun Xu. High-order kinetic flux vector splitting schemes in general coordinates for ideal quantum gas dynamics. Journal of Computational Physics, 227(2):967–982, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003579.

Ying:2006:KIF

[Yin06] Lexing Ying. A kernel independent fast multipole algorithm for radial basis functions. Journal of Computational Physics, 213(2):451–457, April 10, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105004298.

Yang:2006:ARR

[YJ06] Xiaofeng Yang and Ashley J. James. Analytic relations for reconstructing piecewise linear interfaces in triangular and tetrahedral grids. Journal of Computational Physics, 214(1):41–54, May 1, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004225.

Young:2006:MFS

[YJF^{+}06] D. L. Young, S. J. Jane, C. M. Fan, K. Murugesan, and C. C. Tsai. The method of fundamental solutions for 2D and 3D Stokes problems. Journal of Computational Physics, 211(1):1–8, January 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910500272X.

Yang:2006:ACL

[YJL^{+}06] Xiaofeng Yang, Ashley J. James, John Lowengrub, Xiaoming Zheng, and Vittorio Cristini. An adaptive coupled level-set/volume-of-fluid interface capturing method for unstructured triangular grids. Journal of Computational Physics, 217(2):364–394, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106000143.

Yosibash:2004:CMS

[YKG04] Z. Yosibash, R. M. Kirby, and D. Gottlieb. Collocation methods for the solution of von-Kármán dynamic non-linear plate systems. Journal of Computational Physics, 200(2): 432–461, November 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999104001755.

Yoon:2008:MDL

[YKK08] Sung-Hwan Yoon, Chongam Kim, and Kyu-Hong Kim. Multi-dimensional limiting process for three-dimensional flow physics analyses. Journal of Computational Physics, 227 (12):6001–6043, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108001174.

Yu:2001:SOA

[YL01] Heng Yu and Yu-Ping Liu. A second-order accurate, component-wise TVD scheme for nonlinear, hyperbolic conservation laws. Journal of Computational Physics, 173(1): 1–16, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968342.

Yuan:2008:MPC

[YLA08] Jianhua Yuan, Ya Yan Lu, and Xavier Antoine. Modeling photonic crystals by boundary integral equations and Dirichlet-to-Neumann maps. Journal of Computational Physics, 227(9):4617–4629, April 20, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000430.

Yeckel:2009:ABN

[YLD09] Andrew Yeckel, Lisa Lun, and Jeffrey J. Derby. An approximate block Newton method for coupled iterations of nonlinear solvers: Theory and conjugate heat transfer applications. Journal of Computational Physics, 228(23):8566–8588, December 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004379.

Yan:2007:NSF

[YM07] S. Yan and Q. W. Ma. Numerical simulation of fully nonlinear interaction between steep waves and 2D floating bodies using the QALE–FEM method. Journal of Computational Physics, 221(2):666–692, February 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003111.

Youn:2001:ASO

[YMF01] S. J. Youn, Wolfgang Mannstadt, and A. J. Freeman. Analytic spin-orbit coupling matrix element formulae in FLAPW calculations. Journal of Computational Physics, 172(1):387– 391, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999101968305.

Yabe:2004:HOS

[YMT^{+}04] Takashi Yabe, Hiroki Mizoe, Kenji Takizawa, Hiroshi Moriki, Hyo-Nam Im, and Youichi Ogata. Higher-order schemes with CIP method and adaptive soroban grid towards mesh-free scheme. Journal of Computational Physics, 194(1):57–77, February 10, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103004601.

Yang:2006:COA

[YMW06] Chao Yang, Juan C. Meza, and Lin-Wang Wang. A constrained optimization algorithm for total energy minimization in electronic structure calculations. Journal of Computational Physics, 217(2):709–721, September 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106000325.

You:2006:ASA

[YMWM06] Donghyun You, Rajat Mittal, Meng Wang, and Parviz Moin. Analysis of stability and accuracy of finite-difference schemes on a skewed mesh. Journal of Computational Physics, 213(1):184–204, March 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105003761.

Yokoi:2007:EIT

[Yok07] Kensuke Yokoi. Efficient implementation of THINC scheme: a simple and practical smoothed VOF algorithm. Journal of Computational Physics, 226(2):1985–2002, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002732.

Yoneya:2001:GNI

[Yon01] Makoto Yoneya. A generalized non-iterative matrix method for constraint molecular dynamics simulations. Journal of Computational Physics, 172(1):188–197, September 1, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101968196.

You:2006:HOP

[You06] Donghyun You. A high-order Padé ADI method for unsteady convection-diffusion equations. Journal of Computational Physics, 214(1):1–11, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105004638.

Yefet:2001:SFO

[YP01] Amir Yefet and Peter G. Petropoulos. A staggered fourthorder accurate explicit finite difference scheme for the time-domain Maxwell’s equations. Journal of Computational Physics, 168(2):286–315, April 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999101966914.

Yang:2006:SOB

[YP06] B. Yang and A. Prosperetti. A second-order boundaryfitted projection method for free-surface flow computations. Journal of Computational Physics, 213(2):574–590, April 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003931.

Yuan:2006:DGM

[YS06] Ling Yuan and Chi-Wang Shu. Discontinuous Galerkin method based on non-polynomial approximation spaces. Journal of Computational Physics, 218(1):295–323, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000854.

Yee:2007:DLD

[YS07a] H. C. Yee and B. Sjögreen. Development of low dissipative high order filter schemes for multiscale Navier–Stokes/MHD systems. Journal of Computational Physics, 225(1):910–934, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107000101.

Yu:2007:DFF

[YS07b] Zhaosheng Yu and Xueming Shao. A direct-forcing fictitious domain method for particulate flows. Journal of Computational Physics, 227(1):292–314, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003348.

Yuan:2007:AAF

[YS07c] Guangwei Yuan and Zhiqiang Sheng. Analysis of accuracy of a finite volume scheme for diffusion equations on distorted meshes. Journal of Computational Physics, 224(2):1170–1189, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005717.

Yuan:2008:MFV

[YS08] Guangwei Yuan and Zhiqiang Sheng. Monotone finite volume schemes for diffusion equations on polygonal meshes. Journal of Computational Physics, 227(12):6288–6312, June 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108001587.

Yang:2009:SII

[YS09] Jianming Yang and Frederick Stern. Sharp interface immersed-boundary/level-set method for wave-body interactions. Journal of Computational Physics, 228(17):6590– 6616, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109003209.

Ye:2001:FGS

[YSC01] Tao Ye, Wei Shyy, and Jacob N. Chung. A fixed-grid, sharpinterface method for bubble dynamics and phase change. Journal of Computational Physics, 174(2):781–815, December 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999101969384.

Yokota:2007:CIT

[YSO07] R. Yokota, T. K. Sheel, and S. Obi. Calculation of isotropic turbulence using a pure Lagrangian vortex method. Journal of Computational Physics, 226(2):1589–1606, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107002501.

Yu:2005:CQG

[YSS05] Jiun-Der Yu, Shinri Sakai, and James Sethian. A coupled quadrilateral grid level set projection method applied to ink jet simulation. Journal of Computational Physics, 206(1):227–251, June 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104005108.

Yu:2007:TPV

[YSS07] Jiun-Der Yu, Shinri Sakai, and J. A. Sethian. Twophase viscoelastic jetting. Journal of Computational Physics, 220(2):568–585, January 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106002506.

Yu:2006:FDM

[YSW06] Zhaosheng Yu, Xueming Shao, and Anthony Wachs. A fictitious domain method for particulate flows with heat transfer. Journal of Computational Physics, 217(2):424–452, September 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106000167.

Yuan:2007:RSI

[YT07] Li Yuan and Tao Tang. Resolving the shock-induced combustion by an adaptive mesh redistribution method. Journal of Computational Physics, 224(2):587–600, June 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106004840.

Yang:2005:SIC

[YU05a] Yi Yang and H. S. Udaykumar. Sharp interface Cartesian grid method III: Solidification of pure materials and binary solutions. Journal of Computational Physics, 210 (1):55–74, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105002524.

Yu:2005:DFM

[Yu05b] Zhaosheng Yu. A DLM/FD method for fluid/flexible-body interactions. Journal of Computational Physics, 207(1):1–27, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000148.

Yuan:2002:CIM

[Yua02] Li Yuan. Comparison of implicit multigrid schemes for three-dimensional incompressible flows. Journal of Computational Physics, 177(1):134–155, March 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102970075.

Yuste:2006:WAF

[Yus06] S. B. Yuste. Weighted average finite difference methods for fractional diffusion equations. Journal of Computational Physics, 216(1):264–274, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105005516.

Yee:2000:ESN

[YVD00] H. C. Yee, M. Vinokur, and M. J. Djomehri. Entropy splitting and numerical dissipation. Journal of Computational Physics, 162(1):33–81, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100965173.

Yu:2007:TDM

[YW07] Sining Yu and G. W. Wei. Three-dimensional matched interface and boundary (MIB) method for treating geometric singularities. Journal of Computational Physics, 227(1):602–632, November 10, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910700349X.

Yang:2007:GAM

[YWC07] Juekuan Yang, Yujuan Wang, and Yunfei Chen. GPU accelerated molecular dynamics simulation of thermal conductivities. Journal of Computational Physics, 221(2):799–804, February 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003172.

Yokoi:2005:TDN

[YXLF05] Kensuke Yokoi, Feng Xiao, Hao Liu, and Kazuaki Fukasaku. Three-dimensional numerical simulation of flows with complex geometries in a regular Cartesian grid and its application to blood flow in cerebral artery with multiple aneurysms. Journal of Computational Physics, 202(1):1–19, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002657.

Yabe:2001:CIP

[YXU01] Takashi Yabe, Feng Xiao, and Takayuki Utsumi. The constrained interpolation profile method for multiphase analysis. Journal of Computational Physics, 169(2):556–593, May 20, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100966257.

Yan:2009:MMS

[YYF09] Liang Yan, Feng-Lian Yang, and Chu-Li Fu. A meshless method for solving an inverse spacewise-dependent heat source problem. Journal of Computational Physics, 228 (1):123–136, January 10, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999108004579.

Yin:2005:NPS

[YYT05] Z. Yin, Li Yuan, and Tao Tang. A new parallel strategy for two-dimensional incompressible flow simulations using pseudo-spectral methods. Journal of Computational Physics, 210(1):325–341, November 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002226.

Yan:2007:LBM

[YZ07] Y. Y. Yan and Y. Q. Zu. A lattice Boltzmann method for incompressible two-phase flows on partial wetting surface with large density ratio. Journal of Computational Physics, 227(1):763–775, November 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003580.

Yue:2006:PFS

[YZF^{+}06] Pengtao Yue, Chunfeng Zhou, James J. Feng, Carl F. Ollivier-Gooch, and Howard H. Hu. Phase-field simulations of interfacial dynamics in viscoelastic fluids using finite elements with adaptive meshing. Journal of Computational Physics, 219(1):47–67, November 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001392.

Yue:2007:SSD

[YZF07] Pengtao Yue, Chunfeng Zhou, and James J. Feng. Spontaneous shrinkage of drops and mass conservation in phase-field simulations. Journal of Computational Physics, 223(1):1–9, April 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106005857.

Yu:2006:UQC

[YZL^{+}06] Y. Yu, M. Zhao, T. Lee, N. Pestieau, W. Bo, J. Glimm, and J. W. Grove. Uncertainty quantification for chaotic computational fluid dynamics. Journal of Computational Physics, 217(1):200–216, September 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001732.

Yang:2009:STD

[YZLH09] Xiaolei Yang, Xing Zhang, Zhilin Li, and Guo-Wei He. A smoothing technique for discrete delta functions with application to immersed boundary method in moving boundary simulations. Journal of Computational Physics, 228(20):7821–7836, November 1, 2009. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109004136.

Yu:2005:LST

[YZW05] Sining Yu, Shan Zhao, and G. W. Wei. Local spectral time splitting method for first- and second-order partial differential equations. Journal of Computational Physics, 206(2):727–780, July 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105000136.

Yu:2007:MIB

[YZW07] Sining Yu, Yongcheng Zhou, and G. W. Wei. Matched interface and boundary (MIB) method for elliptic problems with sharp-edged interfaces. Journal of Computational Physics, 224(2):729–756, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910600547X.

Zadeh:2008:PEF

[Zad08] Kouroush Sadegh Zadeh. Parameter estimation in flow through partially saturated porous materials. Journal of Computational Physics, 227(24):10243–10262, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108004762.

Zhang:2007:WPF

[ZB07] G. M. Zhang and R. C. Batra. Wave propagation in functionally graded materials by modified smoothed particle hydrodynamics (MSPH) method. Journal of Computational Physics, 222(1):374–390, March 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106003573.

Zhang:2009:RSF

[ZC09] Dali Zhang and Elena Cherkaev. Reconstruction of spectral function from effective permittivity of a composite material using rational function approximations. Journal of Computational Physics, 228(15):5390–5409, August 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109002101.

Zhou:2001:SGM

[ZCMI01] J. G. Zhou, D. M. Causon, C. G. Mingham, and D. M. Ingram. The surface gradient method for the treatment of source terms in the shallow-water equations. Journal of Computational Physics, 168(1):1–25, March 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966701.

Zinchenko:2000:EAH

[ZD00] Alexander Z. Zinchenko and Robert H. Davis. An efficient algorithm for hydrodynamical interaction of many deformable drops. Journal of Computational Physics, 157(2):539–587, January 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999199963842.

Zinchenko:2005:MAA

[ZD05] Alexander Z. Zinchenko and Robert H. Davis. A multipoleaccelerated algorithm for close interaction of slightly deformable drops. Journal of Computational Physics, 207 (2):695–735, August 10, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105000471.

Zinchenko:2008:ADN

[ZD08] Alexander Z. Zinchenko and Robert H. Davis. Algorithm for direct numerical simulation of emulsion flow through a granular material. Journal of Computational Physics, 227(16):7841–7888, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002660.

Zhang:2009:PFM

[ZDD09] Jian Zhang, Sovan Das, and Qiang Du. A phase field model for vesicle-substrate adhesion. Journal of Computational Physics, 228(20):7837–7849, November 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004148.

Zingg:2000:CSS

[ZDNP00] D. W. Zingg, S. De Rango, M. Nemec, and T. H. Pulliam. Comparison of several spatial discretizations for the Navier–Stokes equations. Journal of Computational Physics, 160(2):683–704, May 20, 2000. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999100964829.

Zhang:2006:FTM

[ZEA06] J. Zhang, D. M. Eckmann, and P. S. Ayyaswamy. A front tracking method for a deformable intravascular bubble in a tube with soluble surfactant transport. Journal of Computational Physics, 214(1):366–396, May 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910500450X.

Zhu:2002:SPS

[ZF02] Yi Zhu and Patrick J. Fox. Simulation of pore-scale dispersion in periodic porous media using smoothed particle hydrodynamics. Journal of Computational Physics, 182(2): 622–645, November 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999102971895.

Zhao:2008:FMM

[ZFM08] Hong Zhao, Jonathan B. Freund, and Robert D. Moser. A fixed-mesh method for incompressible flow-structure systems with finite solid deformations. Journal of Computational Physics, 227(6):3114–3140, March 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107005116.

Zabaras:2008:SFS

[ZG08] N. Zabaras and B. Ganapathysubramanian. A scalable framework for the solution of stochastic inverse problems using a sparse grid collocation approach. Journal of Computational Physics, 227(9):4697–4735, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000478.

Zabusky:2003:LSC

[ZGG03] N. J. Zabusky, S. Gupta, and Y. Gulak. Localization and spreading of contact discontinuity layers in simulations of compressible dissipationless flows. Journal of Computational Physics, 188(2):348–364, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999102000451.

Zahedi:2009:CLS

[ZGK09] Sara Zahedi, Katarina Gustavsson, and Gunilla Kreiss. A conservative level set method for contact line dynamics. Journal of Computational Physics, 228(17):6361–6375, September 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109002794.

Zou:2006:TEA

[ZGSD06] Jing Zou, Anna Gilbert, Martin Strauss, and Ingrid Daubechies. Theoretical and experimental analysis of a randomized algorithm for Sparse Fourier Transform analysis. Journal of Computational Physics, 211(2):572–595, January 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105002883.

Zabaras:2006:MDS

[ZGT06] Nicholas Zabaras, Baskar Ganapathysubramanian, and Lijian Tan. Modelling dendritic solidification with melt convection using the extended finite element method. Journal of Computational Physics, 218(1):200–227, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106000787.

Zhao:2001:FTF

[ZH01] P. Zhao and J. C. Heinrich. Front-tracking finite element method for dendritic solidification. Journal of Computational Physics, 173(2):765–796, November 1, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969116.

Zhang:2004:GMC

[ZH04] Y. Z. Zhang and D. C. Haworth. A general mass consistency algorithm for hybrid particle/finite-volume PDF methods.

Journal of Computational Physics, 194(1):156–193, February 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999103004650.

Zhao:2009:FWP

[ZH09] Yan Zhao and Yang Hao. Full-wave parallel dispersive finite-difference time-domain modeling of three-dimensional electromagnetic cloaking structures. Journal of Computational Physics, 228(19):7300–7312, October 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900357X.

Zhang:2002:MMF

[Zha02] Jun Zhang. Multigrid method and fourth-order compact scheme for 2D Poisson equation with unequal mesh-size discretization. Journal of Computational Physics, 179(1): 170–179, June 10, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199910297049X.

Zheng:2006:ENB

[Zhe06] Chunxiong Zheng. Exact nonreflecting boundary conditions for one-dimensional cubic nonlinear Schrödinger equations. Journal of Computational Physics, 215(2):552–565, July 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105005048.

Zheng:2007:PML

[Zhe07] Chunxiong Zheng. A perfectly matched layer approach to the nonlinear Schrödinger wave equations. Journal of Computational Physics, 227(1):537–556, November 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003464.

Zhong:2007:NHO

[Zho07] Xiaolin Zhong. A new high-order immersed interface method for solving elliptic equations with imbedded interface of discontinuity. Journal of Computational Physics, 225(1):

1066–1099, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999107000307.

Zheng:2009:FEV

[ZHSS09] Haibiao Zheng, Yanren Hou, Feng Shi, and Lina Song. A finite element variational multiscale method for incompressible flows based on two local Gauss integrations. Journal of Computational Physics, 228(16):5961–5977, September 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002460.

Ziegler:2004:CCT

[Zie04] U. Ziegler. A central-constrained transport scheme for ideal magnetohydrodynamics. Journal of Computational Physics, 196(2):393–416, May 20, 2004. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103006028.

Zhang:2006:CEA

[ZIP06] Q. Zhang, K. Ichiki, and A. Prosperetti. On the computation of ensemble averages for spatially non-uniform particle systems. Journal of Computational Physics, 212(1): 247–267, February 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999105003190.

Zhang:2009:HOI

[ZJ09] Ju Zhang and Thomas L. Jackson. A high-order incompressible flow solver with WENO. Journal of Computational Physics, 228(7):2426–2442, April 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108006360.

Zhang:2008:DNW

[ZJS08] Shuhai Zhang, Shufen Jiang, and Chi-Wang Shu. Development of nonlinear weighted compact schemes with increasingly higher order accuracy. Journal of Computational Physics, 227(15):7294–7321, July 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002295.

Zhang:2006:NOM

[ZJW06] Yaoxin Zhang, Yafei Jia, and Sam S. Y. Wang. 2D nearly orthogonal mesh generation with controls on distortion function. Journal of Computational Physics, 218(2):549–571, November 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999106001185.

Zhang:2008:BTE

[ZJWC08] Yaoxin Zhang, Yafei Jia, Sam S. Y. Wang, and H. C. Chan. Boundary treatment for 2D elliptic mesh generation in complex geometries. Journal of Computational Physics, 227(16):7977–7997, August 10, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108002878.

Zhang:2004:ASN

[ZK04] Z. Zhang and C. Kleinstreuer. Airflow structures and nanoparticle deposition in a human upper airway model. Journal of Computational Physics, 198(1):178–210, July 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999104000245.

Zhang:2005:LBS

[ZK05] Junfeng Zhang and Daniel Y. Kwok. A 2D lattice Boltzmann study on electrohydrodynamic drop deformation with the leaky dielectric theory. Journal of Computational Physics, 206(1):150–161, June 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104005054.

Zheng:2006:AAI

[ZK06] L.-J. Zheng and M. Kotschenreuther. AEGIS: An adaptive ideal-magnetohydrodynamics shooting code for axisymmetric plasma stability. Journal of Computational Physics, 211(2):748–766, January 20, 2006. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999105002950.

Zacharioudaki:2007:DCB

[ZKDT07] Maria Zacharioudaki, Charalampos Kouris, Yannis Dimakopoulos, and John Tsamopoulos. A direct comparison between volume and surface tracking methods with a boundaryfitted coordinate transformation and third-order upwinding. Journal of Computational Physics, 227(2):1428–1469, December 10, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999107004019.

Zimmermann:2001:SPT

[ZKK01] S. Zimmermann, P. Koumoutsakos, and W. Kinzelbach. Simulation of pollutant transport using a particle method. Journal of Computational Physics, 173(1):322–347, October 10, 2001. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968792.

Zhao:2007:VFM

[ZKL^{+}07] Zhiqin Zhao, Narayan Kovvali, Wenbin Lin, Chang-Hoi Ahn, Luise Couchman, and Lawrence Carin. Volumetric fast multipole method for modeling Schrödinger’s equation. Journal of Computational Physics, 224(2):941–955, June 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106005596.

Zhang:2009:SSF

[ZKS^{+}09] Shuai Zhang, Shigeyuki Kuwabara, Takahito Suzuki, Yoshio Kawano, Koji Morita, and Kenji Fukuda. Simulation of solidfluid mixture flow using moving particle methods. Journal of Computational Physics, 228(7):2552–2565, April 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999108006499.

Zhang:2005:LSM

[ZKY05] Xiang Kun Zhang, Kie-Chan Kwon, and Sung-Kie Youn. The least-squares meshfree method for the steady incompressible viscous flow. Journal of Computational Physics, 206(1):182–207, June 10, 2005. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999104005078.

Zhang:2004:EHO

[ZL04] Dongxiao Zhang and Zhiming Lu. An efficient, high-order perturbation approach for flow in random porous media via Karhunen–Loève and polynomial expansions. Journal of Computational Physics, 194(2):773–794, March 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103005072.

Zhang:2008:NIT

[ZL08a] Qinghai Zhang and Philip L.-F. Liu. A new interface tracking method: The polygonal area mapping method. Journal of Computational Physics, 227(8):4063–4088, April 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005645.

Zhou:2008:VLS

[ZL08b] Shiwei Zhou and Qing Li. A variational level set method for the topology optimization of steady-state Navier–Stokes flow. Journal of Computational Physics, 227(24):10178–10195, December 20, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910800449X.

Zhang:2009:HHC

[ZL09] Qinghai Zhang and Philip L.-F. Liu. HyPAM: a hybrid continuum-particle model for incompressible free-surface flows. Journal of Computational Physics, 228(4):1312–1342, March 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108005561.

Zheng:2005:AUV

[ZLAC05] Xiaoming Zheng, John Lowengrub, Anthony Anderson, and Vittorio Cristini. Adaptive unstructured volume remeshing — II: Application to two- and three-dimensional levelset simulations of multiphase flow. Journal of Computational Physics, 208(2):626–650, September 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105001233.

Zhu:2002:SFF

[ZP02] Luoding Zhu and Charles S. Peskin. Simulation of a flapping flexible filament in a flowing soap film by the immersed boundary method. Journal of Computational Physics, 179(2):452– 468, July 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910297066X.

Zhang:2005:SOM

[ZP05] Z. Zhang and A. Prosperetti. A second-order method for three-dimensional particle simulation. Journal of Computational Physics, 210(1):292–324, November 20, 2005. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105002214.

Zheng:2006:RKC

[ZP06] Zheming Zheng and Linda Petzold. Runge–Kutta–Chebyshev projection method. Journal of Computational Physics, 219 (2):976–991, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999106003391.

Zhu:2009:ARK

[ZQ09] Hongqiang Zhu and Jianxian Qiu. Adaptive Runge– Kutta discontinuous Galerkin methods using different indicators: One-dimensional case. Journal of Computational Physics, 228(18):6957–6976, October 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109003325.

Zhu:2008:RKD

[ZQSD08] Jun Zhu, Jianxian Qiu, Chi-Wang Shu, and Michael Dumbser. Runge–Kutta discontinuous Galerkin method using WENO limiters II: Unstructured meshes. Journal of Computational Physics, 227(9):4330–4353, April 20, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000120.

Zamzamian:2008:MUI

[ZR08] Kamiar Zamzamian and Seyed Esmail Razavi. Multidimensional upwinding for incompressible flows based on characteristics. Journal of Computational Physics, 227(19):8699–8713, October 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999108003409.

Zwart:2000:IST

[ZRR00] P. J. Zwart, G. D. Raithby, and M. J. Raw. The integrated space–time finite volume method and its application to moving boundary problems: Volume 154, number 2 (1999), pages 497–519. Journal of Computational Physics, 162(1):298, July 20, 2000. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999100965549.

Zheng:2006:CMC

[ZRS06] Yingsong Zheng, Jason M. Reese, and Henning Struchtrup. Comparing macroscopic continuum models for rarefied gas dynamics: a new test method. Journal of Computational Physics, 218(2):748–769, November 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106001318.

Zhuang:2001:HOF

[ZS01] Yu Zhuang and Xian-He Sun. A high-order fast direct solver for singular Poisson equations. Journal of Computational Physics, 171(1):79–94, July 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101967713.

Sun:2006:SCE

[zS06] Zhi zhong Sun. The stability and convergence of an explicit difference scheme for the Schrödinger equation on an infinite domain by using artificial boundary conditions. Journal of Computational Physics, 219(2):879–898, December 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106003342.

Zheng:2008:HMK

[ZSB^{+}08] Zheming Zheng, Ryan M. Stephens, Richard D. Braatz, Richard C. Alkire, and Linda R. Petzold. A hybrid multiscale kinetic Monte Carlo method for simulation of copper electrodeposition. Journal of Computational Physics, 227(10):5184–5199, May 1, 2008. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999108000727.

Zheng:2006:LBM

[ZSC06] H. W. Zheng, C. Shu, and Y. T. Chew. A lattice Boltzmann model for multiphase flows with large density ratio. Journal of Computational Physics, 218(1):353–371, October 10, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910600088X.

Zhang:2007:RIL

[ZSC07] Raoyang Zhang, Ilya Staroselsky, and Hudong Chen. Realization of isotropy of the lattice Boltzmann method via rotation of lattice velocity bases. Journal of Computational Physics, 225(2):1262–1270, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000381.

Zheng:2008:OOQ

[ZSC08] H. W. Zheng, C. Shu, and Y. T. Chew. An objectoriented and quadrilateral-mesh based solution adaptive algorithm for compressible multi-fluid flows. Journal of Computational Physics, 227(14):6895–6921, July 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108002027.

Zhang:2002:ACP

[ZSP02] Xing Zhang, David Schmidt, and Blair Perot. Accuracy and conservation properties of a three-dimensional unstructured staggered mesh scheme for fluid dynamics. Journal of Computational Physics, 175(2):764–791, January 20, 2002. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999101969736.

Zheng:2008:SEL

[ZSP08] Zheming Zheng, Bernd Simeon, and Linda Petzold. A stabilized explicit Lagrange multiplier based domain decomposition method for parabolic problems. Journal of Computational Physics, 227(10):5272–5285, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000776.

Zhou:2006:SCF

[ZSTC06] Yunkai Zhou, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Self-consistent-field calculation using Chebyshevfiltered subspace iteration. Journal of Computational Physics, 219(1):172–184, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910600146X.

Zagorodnov:2003:LTN

[ZSW03] Igor Zagorodnov, Rolf Schuhmann, and Thomas Weiland. Long-time numerical computation of electromagnetic fields in the vicinity of a relativistic source. Journal of Computational Physics, 191(2):525–541, November 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103003292.

Sun:2006:SCD

[zSW06] Zhi zhong Sun and Xiaonan Wu. The stability and convergence of a difference scheme for the Schrödinger equation on an infinite domain by using artificial boundary conditions. Journal of Computational Physics, 214(1):209–223, May 1, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999105004316.

Zagorodnov:2007:CFM

[ZSW07] Igor Zagorodnov, Rolf Schuhmann, and Thomas Weiland. Conformal FDTD-methods to avoid time step reduction with and without cell enlargement. Journal of Computational Physics, 225(2):1493–1507, August 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000642.

Zhang:2003:WEN

[ZSWW03] Mengping Zhang, Chi-Wang Shu, George C. K. Wong, and S. C. Wong. A weighted essentially non-oscillatory numerical scheme for a multi-class Lighthill–Whitham– Richards traffic flow model. Journal of Computational Physics, 191(2):639–659, November 1, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999103003449.

Zhong:2003:HON

[ZT03] Xiaolin Zhong and Mahidhar Tatineni. High-order nonuniform grid schemes for numerical simulation of hypersonic boundary-layer stability and transition. Journal of Computational Physics, 190(2):419–458, September 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002821.

Zhang:2007:ASD

[ZT07a] J. M. Zhang and Masa. Tanaka. Adaptive spatial decomposition in fast multipole method. Journal of Computational Physics, 226(1):17–28, September 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107001404.

Zygiridis:2007:OTD

[ZT07b] Theodoros T. Zygiridis and Theodoros D. Tsiboukis. Optimized three-dimensional FDTD discretizations of Maxwell’s equations on Cartesian grids. Journal of Computational Physics, 226(2):2372–2388, October 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999107003208.

Zhu:2005:SFS

[ZTPM05] Luoding Zhu, Derek Tretheway, Linda Petzold, and Carl Meinhart. Simulation of fluid slip at 3D hydrophobic microchannel walls by the lattice Boltzmann method. Journal of Computational Physics, 202(1):181–195, January 1, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999104002724.

Zhao:2002:HRC

[ZTZ02] Yong Zhao, Hsiang Hui Tan, and Baili Zhang. A high-resolution characteristics-based implicit dual timestepping VOF method for free surface flow simulation on unstructured grids. Journal of Computational Physics, 183(1):233–273, November 20, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102971962.

Zhao:2003:MDG

[ZVHP03] P. Zhao, M. Vénere, J. C. Heinrich, and D. R. Poirier. Modeling dendritic growth of a binary alloy. Journal of Computational Physics, 188(2):434–461, July 1, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103001852.

Zhang:2007:AEP

[ZVQ07] Xiang Zhang and Loc Vu-Quoc. An accurate elasto-plastic frictional tangential force-displacement model for granularflow simulations: Displacement-driven formulation. Journal of Computational Physics, 225(1):730–752, July 1, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999106006206.

Zhou:2003:HRC

[ZW03] Y. C. Zhou and G. W. Wei. High resolution conjugate filters for the simulation of flows. Journal of Computational Physics, 189(1):159–179, July 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103002067.

Zhao:2004:HOF

[ZW04] Shan Zhao and G. W. Wei. High-order FDTD methods via derivative matching for Maxwell’s equations with material interfaces. Journal of Computational Physics, 200(1):60–103, October 10, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104001330.

Zagorodnov:2005:TTS

[ZW05] Igor Zagorodnov and Thomas Weiland. TE/TM scheme for computation of electromagnetic fields in accelerators. Journal of Computational Physics, 207(1):69–91, July 20, 2005. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105000173.

Zhou:2006:FDI

[ZW06] Y. C. Zhou and G. W. Wei. On the fictitious-domain and interpolation formulations of the matched interface and boundary (MIB) method. Journal of Computational Physics, 219(1):228–246, November 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999106001495.

Zhang:2002:PMM

[ZWL02] Lucy T. Zhang, Gregory J. Wagner, and Wing K. Liu. A parallelized meshfree method with boundary enrichment for large-scale CFD. Journal of Computational Physics, 176(2):483–506, March 1, 2002. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999102969998.

Zhang:2006:WEN

[ZWS06] Peng Zhang, S. C. Wong, and Chi-Wang Shu. A weighted essentially non-oscillatory numerical scheme for a multi-class traffic flow model on an inhomogeneous highway. Journal of Computational Physics, 212(2):739–756, March 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105003542.

Zerroukat:2007:APS

[ZWS07] M. Zerroukat, N. Wood, and A. Staniforth. Application of the parabolic spline method (PSM) to a multi-dimensional conservative semi-Lagrangian transport scheme (SLICE). Journal of Computational Physics, 225(1):935–948, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107000113.

Zhang:2008:DVD

[ZXQX08] Zhenyu Zhang, Jianzhong Xu, Zhiguo Qi, and Guang Xi. A discrete velocity direction model for the Boltzmann equation and applications to micro gas flows. Journal of Computational Physics, 227(10):5256–5271, May 1, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108000764.

Zhang:2002:STC

[ZYC02] Zeng-Chan Zhang, S. T. John Yu, and Sin-Chung Chang. A space–time conservation element and solution element method for solving the two- and three-dimensional unsteady Euler equations using quadrilateral and hexahedral meshes. Journal of Computational Physics, 175(1):168–199, January 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101969347.

Zhou:2007:NFH

[ZYHS07] Qiang Zhou, Zhaohui Yao, Feng He, and M. Y. Shen. A new family of high-order compact upwind difference schemes with good spectral resolution. Journal of Computational Physics, 227(2):1306–1339, December 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999107003956.

Zhang:2001:JIT

[ZYKW01] Y. L. Zhang, K. S. Yeo, B. C. Khoo, and C. Wang. 3D jet impact and toroidal bubbles. Journal of Computational Physics, 166(2):336–360, January 20, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999100966580.

Zhang:2006:SME

[ZYL^{+}06] Moujin Zhang, S.-T. John Yu, S.-C. Henry Lin, Sin-Chung Chang, and Isaiah Blankson. Solving the MHD equations by the space–time conservation element and solution element method. Journal of Computational Physics, 214(2):599–617, May 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910500464X.

Zhang:2001:USF

[ZZ01] Jun Zhang and Jennifer J. Zhao. Unconditionally stable finite difference scheme and iterative solution of 2D microscale heat transport equation. Journal of Computational Physics, 170(1):261–275, June 10, 2001. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910196735X.

Zhang:2007:IDF

[ZZ07] N. Zhang and Z. C. Zheng. An improved direct-forcing immersed-boundary method for finite difference applications. Journal of Computational Physics, 221(1):250–268, January 20, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106002877.

Zhang:2008:ENM

[ZZ08] Pingwen Zhang and Xinwei Zhang. An efficient numerical method of Landau–Brazovskii model. Journal of Computational Physics, 227(11):5859–5870, May 10, 2008. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108001198.

Zhao:2009:NED

[ZZ09] Haibo Zhao and Chuguang Zheng. A new event-driven constant-volume method for solution of the time evolution of particle size distribution. Journal of Computational Physics, 228(5):1412–1428, March 20, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108005640.

Zhou:2006:HOM

[ZZFW06] Y. C. Zhou, Shan Zhao, Michael Feig, and G. W. Wei. High order matched interface and boundary method for elliptic equations with discontinuous coefficients and singular sources. Journal of Computational Physics, 213(1):1–30, March 20, 2006. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105003578.

Zhang:2008:MPM

[ZZVM08] Duan Z. Zhang, Qisu Zou, W. Brian VanderHeyden, and Xia Ma. Material point method applied to multiphase flows. Journal of Computational Physics, 227(6):3159–3173, March 1, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999107005165.